A Complete Bibliography of Statistics in Medicine (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

18 January 2025 Version 1.06

Title word cross-reference

$150.00 [445]. 2 [2856, 3022, 3076]. 2/3 [3211]. $21.99 [289]. 2 × 2 [3648, 2202, 2449, 2935, 407, 3555, 3650, 406]. 3 [3082, 2489]. $49.70 [398]. 4 × 4 [2433]. $69.95 [567]. $69.95/£ [432]. $79.95 [460]. $83.95 [447, 412]. $85.46 [2012]. $89.95 [106, 446, 380]. $99.95 [118, 506, 433]. _{2.5} [2321]. C [2856, 1838, 307, 840]. D [952, 3448]. Δ [2801]. ℓ_{0} [2969]. F [1820, 576, 1033]. G [1601, 213, 794, 1480]. h [319]. Var(β̂_{1}) [3569]. I × J [2468]. K [423, 1125, 1159, 817, 642, 1503, 1189]. L_{1} [754]. M [1743]. n [2183, 3187, 1503]. P [2768, 1825, 3, 1, 83, 584, 235, 2415, 2183, 2413, 2753, 1241, 2633, 2022, 2719, 1419, 2414]. Q [2128, 2127, 2130, 2129]. R^{2} [3014, 700]. S [483]. t [3185, 325, 2750, 1738, 3152, 1617]. τ [3313]. U [2801, 581, 3163, 2322, 2620, 68].

-component [1189]. -curve [3014]. -estimation [213]. -estimators [483]. -formula [794, 1480]. -index [307]. -inflated [423]. -level [3082]. -likelihood [319]. -nearest [1159]. -norm [2969]. -optimal [952, 3448].

-out-of- [1503]. -quantiles [1743]. -stage [3022, 3076]. -statistic [840, 581]. -statistics [325, 2801, 3163, 2620]. -statistics-based [68]. -test [576, 1033]. -tests [1738, 1820]. -value [83, 584, 3, 2768, 1825]. -values [235, 2415, 2413, 2753, 1241, 2633, 2022, 2719, 1419, 2414, 1].

. [2591].

/SIAM [398].

0 [2725]. 0-19-516865-8 [445]. 0752 [789].

1 [460, 206, 118, 3528, 2504]. 1-4200-6060-0 [433]. 1-58488-887-3 [118]. 1-58488-921-7 [460]. 1-58488-962-4 [106]. 10.1002 [3185, 72, 459, 586, 287]. 10.1002/sim.3607 [287]. 10.1002/sim.3722 [72]. 10.1002/sim.4088 [586]. 10.1002/sim.4105 [459]. 10.1002/sim.7315 [3185]. 13th [2811]. 1553 [584, 3]. 1659 [408]. 1679 [184]. 1702 [175].

2 [2505, 289, 3059]. 2008 [520, 412, 506, 445]. 2009 [398]. 2010 [787, 460, 446]. 2011 [1166]. 2013 [1440]. 2016 [2539]. 21st [3499]. 2268 [557]. 231 [208]. 2331 [3649]. 25 [208]. 2530 [150]. 26 [347]. 27 [175]. 273 [347]. 28 [584, 150, 3, 194, 184]. 29 [1497, 557, 408]. 2B [108]. 2nd [118, 380].

33 [2669]. 3523 [495]. 36 [3185, 2552]. 3682 [194]. 38 [3648, 3656, 3649].

4 [520, 412, 446, 1522]. 4249 [2669]. 4269 [3648]. 4th [289].

5 [106, 567]. 59.99GBP [2012]. 5D [164].

65 [1497].

7 [398, 460, 432, 1457, 411, 506, 2691, 433].

8 [2012, 380].

9 [447]. 978 [398, 2012, 447, 432, 118, 520, 106, 1457, 411, 412, 446, 2725, 289, 506, 2691, 3059, 1522, 567, 433, 380]. 978-0-19-516865-5 [445]. 978-0-387-68639-4 [520]. 978-0-470-02519-2 [289]. 978-0-898716-75-7 [398]. 978-1-4200-6060-7 [433]. 978-1-4200-6117-8 [380]. 978-1-4200-7287-7 [411]. 978-1-4200-7402-4 [446]. 978-1-4398-0354-7 [432]. 978-1-4665-6497-8 [2012]. 978-1-4665-8817-2 [3059]. 978-1-4822-4936-1 [2846]. 978-1-4987-0419-9 [2591]. 978-1-4987-2096-0 [2725]. 978-1-4987-7571-7 [2691]. 978-1-58488-087-5 [567]. 978-1-58488-609-9 [447]. 978-1-58488-658-7 [506]. 978-1-58488-828-4 [412]. 978-1-58488-887-1 [118]. 978-1-58488-921-2 [460].

978-1-58488-962-5 [106]. 978-1-84668-620-7 [1457]. 978-82-05-39959-4 [1522].

Aalen [2102]. AB/BA [1400, 770, 1401, 1399]. Abdul [2106]. aberration [370]. Abhaya [118]. abilities [3502]. ability [1781, 841, 842, 1311, 2789, 2865]. abnormal [2276]. absence [264, 212, 695, 2933, 1196, 481, 2203]. Absolute [953, 216, 442, 678, 1856]. absorbing [803]. absorptiometry [344]. abuse [529, 964, 1767, 3418, 2622]. Accelerated [2200, 1895, 2964, 2937, 3114, 1478, 1990, 818, 2852, 1446, 3324, 3167, 747]. accelerometer [2512]. accelerometers [2855]. acceptably [1476]. access [1203]. accidents [3443]. accommodates [1565]. Accommodating [2933]. according [1385, 1207]. account [935, 1001, 1091, 3163, 393, 1937]. Accounting [2648, 912, 1404, 3573, 1841, 3395, 1926, 2554, 351, 1899, 2142, 1995, 1925, 3015, 1788, 261, 2141, 3396, 3594, 2967]. accounts [494]. accrual [1869, 1834, 515, 67]. accumulated [2037]. accumulating [2547]. accuracies [1437]. Accuracy [624, 2789, 285, 491, 1053, 236, 2571, 2516, 842, 1698, 3014, 1486, 2306, 1798, 2162, 2090, 3591, 3655, 1051, 675, 2196, 2823, 1405, 3398, 3431, 502, 612, 844, 1573, 3127, 2064, 1938, 299, 3120, 2212, 1607, 729, 695, 3148, 1257, 1196, 3053, 2063, 1594]. Accurate [2753, 1252, 3438, 2724, 2279]. accused [3497]. acid [2518]. acquired [715, 3629]. across [1199, 2844, 376, 3079, 2618, 940, 2608, 913, 379, 3493, 3481, 1938, 1273, 1538]. across-trial [2618]. action [873]. active [1808, 1475, 1226, 531, 631, 601, 794, 1855, 1635]. activity [376, 1246, 2855, 480]. acuity [3492]. acupuncture [690, 689]. acute [1618, 1331, 2665, 677, 1690, 1096]. adaptation [1099, 1794]. adaptations [975]. adapted [476, 1033]. Adaptive [1858, 2465, 1882, 2383, 241, 325, 2061, 1149, 3306, 1099, 1101, 2115, 1998, 613, 1100, 3533, 1042, 796, 797, 3280, 1170, 786, 160, 1370, 2764, 2111, 3237, 83, 1132, 719, 239, 1674, 3331, 2566, 1891, 1130, 3447, 1500, 435, 1897, 1295, 988, 1191, 1869, 1299, 1476, 2167, 2002, 467, 2753, 680, 2021, 2112, 2404, 297, 1153, 1834, 3378, 1689, 3186, 799, 1976, 3151, 705, 2849, 2633, 313, 1013, 1461, 2593, 1448, 1279, 1284, 2617, 3048, 615, 1723, 2522, 2113, 177, 2161, 26, 3304, 2578, 3472, 1765, 1367, 1738, 3208, 2122, 414, 148, 2432, 2195, 2861, 2493, 3103, 3165, 580, 370, 3343]. adaptive [2583, 468, 399, 2114, 1548, 877, 3514, 1563, 3190, 1596, 3350, 436, 3456, 2934, 1690, 2117, 614, 649, 106]. adaptively [3305]. add [1867, 3484]. added [1375]. Addendum [603, 297]. Adding [84]. Additional [1321, 1682, 1851, 2898, 981, 3438, 992, 1305, 1413, 2448, 1389]. Additive [34, 2318, 2355, 1447, 2340, 2103, 3253, 3304, 2733, 1407, 3359, 3608, 3595, 3514, 3582, 3116, 3007]. additivity [1771]. address [1677, 2485, 1616, 2859, 1577]. Addressing [3062, 1697, 1692]. adequacy [428]. Adherence [11, 1514, 114, 3454, 3005, 3621, 1655]. adjust [188, 1959, 3214, 611, 2230]. Adjusted [3523, 160, 3459, 3322, 1370, 2355, 574, 240, 438, 3526, 3413, 88, 1464, 1220, 1662, 1843, 700, 677, 2908, 3163, 3641, 2709, 276, 50, 3532, 2749, 304, 3427, 3105, 1755, 3124, 944, 625, 1538, 2029, 2745, 1054, 3652]. Adjusting [417, 1104, 3446, 3512, 2356, 1624, 1879, 426, 250, 2086, 352, 2772]. Adjustment [1600, 234, 1348, 3382, 1454, 2964, 101, 1379, 2271, 2732, 154, 2730, 207, 1591, 1411, 813, 2511, 3633, 2442, 3302, 641, 1183, 1919, 113, 177, 1444, 1868, 2631, 1691, 1859, 1727, 2386]. adjustments [2296, 155, 328, 2051, 2089, 1807, 3018]. Administration [1178, 2579]. administrative [3512, 2090]. Admissible [3342]. admission [3527, 1203]. admissions [598]. admit [2437]. ado [3599]. adolescent [2803]. adolescents [345]. adopted [999]. adoption [95]. ADR [505]. adults [2595]. Advanced [1454, 2795, 172, 3134]. advances [461]. Adverse [966, 720, 1490, 505, 1476, 2547, 2570, 3118, 3197, 3042, 3293, 788, 3563, 1577]. advice [475]. AERS [1987]. affected [236, 1810]. affecting [281, 3375]. Africa [390, 2040]. AFT [1317]. after [147, 3495, 2732, 1371, 932, 1976, 674, 1151, 2036, 3574, 60, 1772]. aftermath [2330]. again [152, 407, 406]. against [3004, 2370, 1129]. Age [2290, 1447, 1928, 3074, 178, 1219, 1160, 2965, 3458, 1914, 1450, 330, 1468, 1730, 3057, 2748, 3499, 3251, 126, 61, 171, 1432, 939, 2280, 1155, 316, 1538]. age-adjusted [1538]. age-at-onset [1155, 316]. age-incidence [939]. age-period-cohort [2965, 1468, 1730, 2748]. age-related [3499, 2280]. Age-space-time [2290]. age-specific [1160, 61, 1432]. age-stratified [330]. ageing [321, 3051]. agent [1942, 1878, 110, 237, 2579]. agents [753, 1848, 1334, 1706, 2432, 486, 1558, 1052, 2416, 3456]. ages [1162]. aggregate [947, 1073, 2615, 2998, 3433, 3334, 221, 1512, 924]. aggregated [3597, 2634, 2792]. Aggregating [846]. aggregation [1571, 1391, 3222, 3108]. Aging [802, 557, 228]. agonistic [3430]. agreement [2354, 1412, 636, 3218, 675, 1780, 1378, 64, 2015, 2790, 192, 402, 812, 1426, 890, 2825, 934, 3496, 1847]. AIDS [1550, 2519, 238, 2372, 794, 2946]. aiming [2620]. air [519, 3015, 913, 268]. airway [2678]. Akihiro [1558]. al [543, 2033, 1946, 1559, 3549]. al. [1059, 3647]. Alan [411]. alarms [3403, 478]. Alcohol [2368, 1025, 128, 709, 263, 606]. alerts [366]. algorithm [3008, 552, 818, 1600, 1994, 1330, 273, 3375, 1842, 1095, 76, 3384, 293, 2805, 504, 2645]. Algorithms [2881, 1484, 3196, 408, 3474, 166, 409, 577]. alive [2845]. all-cause [3032]. all-or-nothing [2911]. all-subset [484]. allele [2249]. allocation [2925, 1370, 1923, 476, 400, 698, 969, 1516, 2740, 925, 2449, 2935, 1013, 1284, 1586, 2041, 968, 1812, 3145, 3103, 3165, 1050, 2736, 743]. Allowing [1840, 3314, 606, 2868, 2349, 2919, 702, 1394, 2338]. allows [674]. Alonso [2846]. alopecia [3149]. Alpha [3408, 160, 1722, 1013, 1136, 3154, 3165, 68]. alpha-adaptive [160]. alternating [3006, 65, 2609]. Alternative [3021, 497, 720, 3648, 1816, 909, 3244, 3020, 784, 3555, 3650, 652, 2500, 2871, 2632, 1776, 2234]. alternatives [1829, 933, 1129]. Altman [2287, 2825]. always [2301]. Alzheimer [3373, 2758, 1470, 3001, 2818, 3592, 3585, 3144]. Amazon [3529]. ambiguous [3033]. AMD [3159]. among [996, 391, 2819, 1229, 3590, 1665, 3218, 1833, 261, 2038, 274, 3113, 3012, 939, 205, 2928, 3453, 104, 246]. analyse [3244]. analysed [722]. Analyses [607, 587, 442, 2882, 2978, 3199, 463, 759, 2209, 170, 105, 552, 640, 134, 945, 2160, 3010, 756, 150, 1991, 2334, 1611, 1867, 2022, 3420, 3181, 3308, 8, 1335, 396, 3358, 1984, 595, 57, 2295, 2835, 1511, 734, 2631, 2668, 3284, 2322, 94, 1076, 1594]. analysi [586, 194, 287]. Analysing [1224, 292]. Analysis [2997, 815, 761, 488, 122, 239, 128, 2717, 505, 2621, 3397, 1990, 861, 1091, 1702, 2007, 3406, 349, 2047, 1530, 3581, 3234, 1931, 978, 2532, 2275, 1787, 604, 3629, 3492, 3531, 422, 3621, 343, 1732, 618, 2793, 787, 3285, 1062, 560, 165, 3560, 3273, 2657, 555, 1313, 2755, 2636, 3657, 2551, 2795, 2630, 3316, 1541, 2684, 1682, 1851, 987, 447, 2247, 457, 3027, 243, 1970, 2868, 401, 543, 2590, 1219, 2182, 97, 389, 949, 2210, 3325, 3279, 2073, 1404, 2509, 1294, 3485, 2423, 135, 763, 1333, 1349, 2624, 1621, 2469, 1731, 318, 186, 211, 3534, 1692, 323, 904, 780, 929, 1119, 1521, 1462]. analysis [1816, 2083, 2594, 2744, 2804, 2571, 55, 432, 2965, 3374, 434, 2343, 1304, 25, 6, 2055, 2006, 1580, 326, 542, 2397, 1779, 1674, 2300, 962, 1234, 1571, 454, 2979, 1315, 96, 1841, 2384, 2894, 1497, 947, 1073, 1208, 2841, 2596, 2649, 3465, 141, 528, 1470, 2277, 2844, 836, 3424, 2488, 3326, 300, 3639, 3654, 785, 685, 804, 2264, 1650, 544, 946, 3300, 571, 1957, 279, 330, 1001, 1995, 921, 2398, 3217, 90, 2796, 2439, 1269, 713, 2025, 526, 2780, 666, 1002, 1553, 545, 174, 2480, 2960, 541, 659, 1677, 290, 413, 2127, 1797, 2869, 800, 2615, 2998]. analysis [1992, 139, 1554, 1925, 2989, 2618, 645, 1803, 2601, 2842, 3156, 3135, 119, 3062, 540, 546, 1074, 1661, 1443, 2848, 1954, 3561, 1815, 2044, 1920, 3364, 3246, 3591, 3655, 303, 1874, 1745, 667, 3462, 3463, 5, 348, 286, 1298, 2205, 2781, 807, 317, 1226, 280, 1395, 1149, 2457, 1507, 3653, 3085, 2815, 1999, 1405, 2119, 14, 30, 1695, 1026, 1166, 3433, 1390, 1411, 3006, 3603, 87, 3231, 1249, 58, 187, 347, 440, 773, 746, 1837, 2224, 2464, 2348, 2837, 2852, 3134, 3086, 3202, 3478, 3271, 2866, 3128, 3212, 3572, 3058, 2573, 1960, 694, 1260, 2278, 2731, 3188, 1699, 42, 600, 48, 1830, 53, 1183, 1602, 2984, 1376]. analysis [1446, 3464, 407, 231, 830, 620, 581, 1506, 1210, 311, 1144, 3189, 472, 3416, 3066, 1003, 1788, 1840, 3314, 671, 2769, 3515, 3334, 1330, 2907, 784, 1394, 312, 769, 3045, 1297, 750, 1663, 2586, 3037, 834, 2132, 3372, 2554, 2357, 915, 1919, 2175, 2064, 2086, 616, 2629, 3018, 2697, 416, 1043, 3025, 1200, 3098, 3063, 2585, 3013, 137, 340, 2332, 2510, 2734, 1978, 1832, 679, 274, 2243, 331, 159, 1084, 1442, 2244, 3360, 2338, 1421, 2705, 2990, 1512, 305, 2060, 1359, 2537, 2938, 406, 357, 1207, 1968, 3579, 2359, 1005, 2227, 506, 601, 1273, 1000, 2136, 126]. analysis [3059, 630, 2353, 583, 1285, 1735, 2078, 2293, 299, 1713, 324, 1537, 2661, 619, 594, 1152, 2152, 1346, 1357, 24, 171, 714, 3269, 3482, 169, 3632, 717, 1718, 2533, 1583, 2709, 689, 129, 1498, 295, 354, 2035, 3521, 3129, 2242, 775, 350, 1227, 276, 277, 3356, 2569, 1482, 1382, 2318, 2528, 1940, 3557, 296, 52, 2065, 3022, 3282, 1164, 1823, 2104, 257, 3469, 2666, 3040, 1338, 806, 3123, 455, 2828, 3107, 3330, 2825, 2723, 3125, 749, 1937, 1857, 3357, 1749, 835, 3594, 301, 567, 802, 293, 1742, 575, 1856, 2639, 658, 951, 704, 2232, 2521, 2400, 2250, 2671]. analysis [2766, 3320, 3409, 3349, 1577, 3033, 1217, 1094, 1997, 3187, 3137, 1242, 2632, 283, 3618, 3287, 2301, 2798, 880, 1332, 433, 1757, 3112, 3580, 1639, 3078, 1368, 2440, 3158, 1831, 2803, 3081, 3635, 3350, 2199, 316, 3543, 2680, 3292, 3535, 3064, 2063, 2714, 3083, 2020, 2455, 380, 829, 1306, 1363, 1985, 3133, 2991, 2385, 2695, 1641, 456, 589, 1150, 3147, 513, 2386, 2520, 2128, 2130, 130, 133, 2129, 132, 131]. Analytic [1561, 1059, 1430, 1489, 3298, 2060, 700, 1145, 2197]. Analytical [1789, 2180, 2878]. analyze [2218, 3495, 785, 1495, 2855, 3017, 3644, 205, 1264]. analyzed [155]. Analyzing [3638, 2036, 215, 583, 3207, 1607, 2184, 2335, 2616, 827, 1342, 2994, 2915, 1427, 982, 2747, 1066, 252, 682, 2737, 1356, 1190, 158, 1364, 191, 2926, 2489]. anatomic [3153]. ancestry [284]. Ancillarity [3262]. and/or [3584, 1968]. Andersen [298, 39, 3593]. Andrés [3096]. Andrew [900]. androgenetic [3149]. Anemona [767]. angle [547]. angles [1138]. animal [3509]. animals [239]. ankylosing [1462]. Annual [107, 6, 2781, 194]. anomalous [3179]. ANOVA [1427]. Antedependence [1243]. anthropometry [3251]. anti [262, 3264]. anti-cancer [262]. anti-infective [3264]. antibacterial [1711]. Antibiotic [1112]. antibiotics [1662, 3418]. antibody [3571, 1995]. anticipated [3396]. antidepressant [1215]. antidiabetic [1521]. antigen [591, 2948]. antigenic [2949]. antihypertensive [1070]. antimicrobial [2530]. antipsychotic [2277]. antipsychotics [500]. antiretroviral [2450, 11, 794]. Antonio [3096]. any [2848, 1867]. Apgar [2263]. Application [266, 3443, 189, 3415, 3373, 2916, 1055, 1283, 41, 267, 1023, 59, 1037, 876, 523, 732, 511, 3375, 2368, 792, 3585, 3113, 2275, 3281, 3149, 1281, 628, 1550, 2263, 260, 2204, 321, 3391, 1323, 457, 1618, 2389, 2758, 761, 1886, 2423, 2976, 318, 3604, 3428, 3495, 3142, 1318, 1025, 1272, 1462, 2164, 942, 1674, 2919, 2519, 1686, 962, 3121, 479, 2315, 952, 715, 2813, 2225, 3159, 1988, 753, 2277, 2437, 1017, 2419, 300, 2388, 2048, 2236, 466, 206, 2548, 302, 808, 1473, 3571, 2944, 1001, 2760, 2067, 2017, 2439, 2854, 3559, 2089, 2371, 3528, 1209, 1677, 1680, 85, 1992]. application [3171, 238, 2468, 297, 563, 1401, 1399, 538, 2880, 3093, 1835, 2901, 183, 3381, 3656, 3293, 1771, 1909, 3036, 2154, 690, 3232, 3239, 3638, 3442, 503, 3031, 896, 2818, 3017, 3082, 1662, 867, 2442, 2512, 2274, 916, 1068, 1376, 1707, 3254, 893, 3051, 3558, 2036, 3066, 709, 1469, 1354, 1003, 1344, 3336, 3441, 3172, 2262, 834, 2452, 1383, 1022, 1703, 2086, 192, 307, 3251, 1200, 1519, 1215, 274, 2233, 2652, 3068, 3574, 2136, 677, 263, 1261, 997, 607, 3627, 1744, 3629, 894, 1632, 1616, 532, 1345, 793, 2216, 3282, 3467, 121, 590, 2775, 1879, 3518, 1966, 2723, 2393, 1302, 1857, 782, 802]. application [2722, 258, 493, 658, 2794, 1997, 606, 789, 2068, 3622, 576, 1400, 3387, 1189, 2279, 3292, 436, 1601, 2946, 374, 3117, 2954, 691, 456, 1150, 681, 3022]. Applications [566, 3602, 3012, 3496, 1818, 529, 964, 2721, 929, 1631, 2820, 2970, 462, 1767, 117, 1754, 2951, 2595, 240, 3577, 465, 2082, 1491, 2057, 1706, 2360, 1298, 2694, 1107, 3001, 3202, 1872, 1308, 1177, 3464, 662, 3368, 954, 1485, 776, 2293, 2120, 1382, 296, 931, 3384, 2678, 1772, 3640, 3116, 1917, 1831, 2622, 1737, 2695]. Applied [460, 3277, 262, 935, 270, 461, 1874, 1079, 583, 2846]. Applying [1142, 640]. approach [2176, 1205, 1517, 3074, 1901, 2641, 3238, 178, 2484, 728, 912, 1595, 3562, 51, 2218, 1430, 2685, 1294, 368, 777, 2365, 1692, 1862, 489, 1463, 1814, 2083, 3361, 2640, 1895, 2994, 716, 3527, 1580, 170, 2534, 3159, 1300, 1568, 2339, 2862, 621, 1090, 988, 1191, 2336, 2548, 102, 487, 330, 2149, 2742, 2407, 270, 3197, 1759, 3275, 1553, 1766, 1408, 174, 2310, 430, 1097, 1665, 932, 1427, 680, 1334, 2747, 2989, 224, 2273, 3268, 1706, 1447, 1920, 2340, 1838, 3016, 3309, 3656, 807, 1626, 280, 2457, 1151, 2665, 1241, 70, 1780, 252, 3442, 1837, 1870, 3222, 3468, 3271, 2511, 1918, 250, 1446, 2737]. approach [3362, 581, 2047, 1124, 471, 3045, 1230, 2940, 193, 1356, 2806, 3235, 123, 113, 656, 17, 3030, 889, 2028, 3545, 3472, 2118, 3360, 631, 3355, 1512, 329, 3396, 2441, 1725, 169, 655, 3335, 2448, 1590, 175, 1712, 933, 2569, 2528, 611, 928, 1407, 1801, 3595, 3518, 2825, 634, 278, 1638, 734, 518, 1719, 301, 1558, 3076, 951, 2502, 2620, 1564, 2632, 2741, 927, 2058, 2001, 3000, 2074, 2967, 54, 1257, 68, 3453, 1533, 3387, 2600, 2231, 625, 1328, 1508, 3319, 2085, 513, 7, 176, 2788, 2712, 2599]. Approaches [3451, 2740, 112, 401, 2624, 3140, 2434, 846, 2644, 3106, 608, 3231, 3431, 2134, 694, 2426, 2984, 1581, 2255, 1928, 1713, 2179, 986, 890, 2217, 1615, 2385, 1641, 1578]. appropriate [2586, 2905]. Approximate [221, 162, 1158, 556, 1638]. approximation [3534, 552, 2443, 3597, 1245, 3317]. approximations [1825, 137, 2337]. arbitrary [774, 1006]. Archimedean [2979]. arcsinh [3492]. area [2354, 3380, 716, 1768, 454, 2406, 967, 1681, 156, 249, 791, 1255, 484, 2830, 927, 2982, 1479, 2279]. areal [2854, 1022]. areas [1393, 716, 3291, 1372]. ARIC [3391, 2904]. Ariel [2846]. arising [3603]. arithmetic [785]. arm [147, 1323, 2787, 434, 239, 1580, 798, 1165, 487, 2149, 2987, 341, 1092, 2133, 2604, 5, 348, 1999, 2740, 863, 2194, 3433, 3484, 347, 111, 1699, 1636, 273, 2132, 2826, 2627, 1075, 1714, 3355, 1713, 594, 1152, 1358, 1939, 928, 983, 2545, 985, 1253, 1563, 2620, 2786, 3618]. arm-based [3618]. armed [1514, 2259, 580]. Armitage [2811, 787, 2319, 593, 1166]. Armitage-type [2319]. arms [1982, 1324, 1634]. array [163]. arterial [893]. artery [2198]. arthritis [418, 1998, 1049, 354, 1966, 2555]. ascending [3289]. ascertained [645, 2723, 1425]. Ascertainment [2438, 1413]. ask [2061]. aspart [2017]. aspect [2676]. assay [2974, 1780, 1872, 1037, 2757, 1306]. assays [2619, 978]. assembled [2820]. assess [3459, 3322, 2342, 323, 1463, 227, 1665, 503, 3247, 3589, 2150, 1050, 1185, 2514, 2610, 2501]. assessed [830]. Assessing [560, 2641, 3067, 3371, 2354, 2988, 636, 1839, 1027, 3385, 947, 3616, 3236, 487, 3571, 1629, 3422, 2527, 1585, 146, 3309, 1810, 1360, 1780, 168, 3433, 2285, 3541, 3619, 1428, 2790, 192, 245, 480, 2754, 3120, 1179, 1155, 1584, 1917, 1847, 1737, 1534, 1029, 1036, 2380, 1300, 2149, 2403, 1032, 2306, 3218, 3511, 1509, 402, 812, 2552, 679, 1773, 1426, 248, 3184, 695, 1414, 1368, 1034, 1887, 2776, 2011]. Assessment [3285, 418, 1134, 3170, 3557, 484, 828, 2517, 1439, 1337, 2188, 3310, 2841, 941, 1085, 3217, 1759, 673, 910, 1491, 1030, 1835, 738, 3602, 896, 3468, 1037, 3082, 1202, 3139, 1079, 1592, 1574, 429, 848, 1133, 1562, 994, 1086, 333, 386, 590, 3518, 2393, 729, 2902, 2646, 1275, 2020, 3072, 2922]. assessments [795, 2731, 1328]. Assignment [257, 786, 1515, 2593, 1701, 2832, 3187]. assist [706]. Assisted [3424, 3230, 482, 3089]. associated [3658, 3136, 1409, 3247, 539, 2275, 1856, 2639, 1755]. association [2484, 431, 2347, 3391, 3220, 346, 744, 3423, 1535, 2873, 12, 284, 819, 1536, 22, 3463, 3253, 388, 2166, 733, 1067, 3511, 605, 1506, 2831, 1467, 2782, 3360, 2060, 599, 1261, 717, 1012, 1800, 3138, 521, 890, 167, 3313, 2322, 2843, 1332, 2440, 2772, 1425, 3210]. Associations [849, 935, 670, 544, 946, 1471, 3118, 3635]. assuming [2484, 1197, 1848, 252]. assumption [454, 947, 1734, 3127, 3172, 535, 1761, 2613, 181]. assumptions [2589, 3224, 2708, 2983]. Assuncão [1865]. Assurance [1406, 1597, 957]. assure [1476]. asthma [1677, 2677]. asthmatics [2678]. Asymmetric [2595, 2843, 2303]. asymmetrical [1031]. asymmetry [3066, 2489]. Asymptotic [2801, 1351, 1463, 1986, 1464]. Atherosclerosis [1600]. Atherton [1889]. Atkinson [76]. atlas [1657]. attacks [329]. attempts [426]. attention [782]. attention-deficit [782]. attention-deficit/hyperactivity [782]. attitude [205]. attributable [1762, 3168, 89, 908, 2352, 1070, 2579, 3024, 3525]. attributes [1038]. attrition [2051, 3100]. AUC [1682, 1851, 36, 2801, 3462, 2062]. AUCs [838]. audit [565]. augment [2323]. augmentation [3384]. Augmented [1668, 3191, 712, 2969, 256, 1913]. Austin [1592]. Australian [1277]. Australians [2038]. Author [639, 2116, 195, 1804, 1296, 3654, 140, 2123, 1401, 3646, 1101, 151, 615, 650, 185, 73, 648, 2903, 2885, 209, 558, 288, 337, 1981, 973, 1975, 585, 1606, 1605, 629, 407]. autism [1176]. autocorrelation [1933, 104]. automated [366, 104]. Automatic [3291]. autopsy [3507]. autopsy-confirmed [3507]. Autoregressive [2784, 685, 3410, 3190]. autosomal [3144]. autosomal-dominant [3144]. Auxiliary [3243, 1528, 3347, 1473, 1990, 826, 976, 3119, 69, 2682, 2231]. auxiliary-covariate-based [1473]. availability [1207]. available [2824]. average [2401, 1198, 3429, 2640, 6, 697, 2781, 1143, 1652, 3420, 2169, 2281, 3048, 2923, 194, 908, 50, 3286, 1659, 2881, 2351, 3453, 2303]. averaged [2876]. averaging [170, 469, 2841, 2464, 2582, 3311, 2425]. avoid [2618]. Avoiding [1304, 3613, 562]. AWNCut [3230]. axillary [1122].

B [2106, 647, 72, 557, 1522, 3430, 327]. B-splines [327]. BA [1400, 770, 1401, 1399]. babies [1162]. back [3235, 1928, 1719, 1775]. back-calculation [3235, 1928, 1719]. background [2504]. Bacon [1889]. bacterial [1631, 1711]. Bagged [3228]. Baker [1320]. balance [3277, 1526, 156, 772, 2077, 2497]. balanced [1971, 207, 2751, 2337, 1564, 436, 2963]. Balancing [791]. Bandyopadhyay [628]. barrier [1996]. barriers [2159]. BART [2318, 3608]. Bartlett [616, 3018]. Bartlett-type [616, 3018]. basal [2802, 3410]. base [142]. based [3238, 2636, 728, 3459, 3322, 1593, 2427, 3034, 1492, 2973, 3035, 2138, 112, 2953, 1970, 1539, 1781, 401, 2589, 1434, 3415, 2210, 2469, 1688, 1959, 2365, 1483, 1862, 1463, 3140, 3423, 2640, 1895, 71, 898, 284, 2307, 3121, 2801, 2838, 552, 3255, 1423, 1211, 2696, 3565, 2287, 2802, 2779, 808, 3221, 1473, 2944, 2732, 1759, 2945, 2467, 3157, 1808, 1334, 3529, 982, 706, 3097, 2747, 2584, 2989, 1032, 1915, 381, 645, 1575, 2553, 1798, 2162, 2245, 2858, 1722, 2316, 3466, 3576, 1030, 992, 3410, 1689, 1922, 2817, 2972, 2105, 2694, 1066, 2166, 93, 2154, 783, 678, 2022, 3248, 3603, 901]. based [1461, 1870, 1994, 2593, 3230, 2969, 3631, 3583, 1037, 1662, 553, 449, 867, 2043, 1335, 1489, 437, 2325, 830, 2050, 3011, 581, 110, 1569, 1904, 758, 1230, 1687, 2447, 1574, 64, 2175, 1174, 1703, 616, 3018, 508, 402, 812, 2552, 492, 1524, 1598, 86, 1614, 1084, 1442, 3360, 1485, 1487, 1512, 2703, 1000, 1883, 125, 299, 617, 958, 997, 57, 443, 826, 1413, 2179, 727, 49, 726, 2677, 524, 843, 521, 793, 2690, 1164, 3119, 554, 16, 3152, 3211, 455, 486, 2948, 2335, 2883, 3049, 2825, 304, 2054, 831, 782, 3280, 3552, 1558, 3226, 2502, 3614, 3439, 3489, 167, 3033, 2843, 2620, 606]. based [1264, 3618, 181, 927, 1615, 2068, 2646, 1332, 536, 1679, 94, 1243, 1275, 68, 2062, 2198, 3350, 1510, 316, 626, 3083, 3072, 2535, 395, 1425, 2203, 2745, 2776, 693, 1011, 1671, 3624, 2673, 2922, 3407, 334, 3268, 2906]. baseline [1062, 1404, 780, 1977, 1002, 1553, 150, 791, 8, 1830, 1183, 312, 1207, 3158, 1076, 3383]. baselines [2550]. basic [3236, 688, 1763]. basis [1224, 730, 3262, 497]. basket [2674, 3176, 3634]. batch [1591, 42, 951]. batch-specific [951]. batch/experiment [42]. batch/experiment-specific [42]. Bayes [1637, 1430, 1731, 2407, 669, 1758, 261, 656, 1148, 2697, 3113, 3521, 452, 2001]. Bayesian [3185, 447, 2796, 3082, 1945, 2213, 2484, 1824, 2517, 1541, 2031, 2247, 51, 2590, 3415, 294, 2685, 2916, 1294, 2180, 2976, 135, 186, 1274, 1688, 2721, 3380, 2770, 3429, 2853, 1862, 904, 936, 1310, 1521, 1993, 3055, 3361, 653, 719, 2994, 101, 25, 716, 2308, 170, 326, 1923, 2566, 2949, 1291, 2519, 469, 1292, 1041, 2970, 264, 745, 1199, 2924, 798, 2339, 938, 1505, 1897, 1914, 1018, 3363, 3473, 2264, 1877, 1195, 1957, 487, 330, 2149, 2290, 1476, 1944, 1948, 858, 1764, 2854, 1848, 1910, 2780, 3275, 405, 1700, 1766, 1478, 1468, 41, 2529, 659, 3176, 2869, 238, 582, 2553, 2306, 3268]. Bayesian [1491, 1809, 3056, 563, 737, 212, 2162, 1153, 1834, 592, 756, 3576, 1730, 1745, 3378, 1835, 711, 164, 21, 766, 1298, 3016, 3186, 799, 1626, 3253, 2823, 2119, 313, 510, 873, 2095, 1600, 3181, 1649, 1837, 2464, 2852, 3222, 3468, 3271, 2442, 3170, 3247, 250, 1602, 2750, 1177, 1287, 1446, 1707, 2047, 127, 3589, 3180, 3077, 3195, 2792, 2372, 2617, 1003, 671, 2769, 2743, 1394, 1663, 1687, 373, 1020, 564, 367, 2697, 2294, 2368, 1200, 3223, 2088, 2582, 792, 137, 2177, 2352, 1367, 3193, 3208, 1362, 3366, 3607, 329, 700, 451, 1176, 736, 825, 2293, 1989, 268, 148, 171, 169, 1163, 3335, 1939]. Bayesian [1590, 350, 2733, 2318, 2528, 3524, 611, 3492, 1933, 554, 3608, 1846, 3343, 3123, 983, 2008, 2583, 1937, 1857, 1719, 301, 2229, 1222, 658, 1175, 3514, 3346, 3437, 3320, 2321, 1563, 1217, 3187, 1520, 3137, 2483, 181, 499, 1936, 1884, 3640, 821, 3002, 54, 1275, 865, 1533, 327, 3149, 2568, 512, 436, 2416, 2231, 3456, 625, 1170, 1306, 1328, 1162, 1508, 2453, 2494, 3536, 781, 1451, 1911, 2385, 2926, 2085, 2799, 3624, 3488, 3219, 2788, 777, 432]. Bayesian-frequentist [798]. be [3382, 639, 638, 2291, 1073, 150, 8]. because [2260, 672]. becoming [858]. BED [1532]. been [1696]. before [147, 975, 66]. before-after [147]. beginning [569, 570]. behavior [1044, 1355, 1046, 1302, 1045, 1011]. behavioral [2431, 3422, 2047]. behaviour [3003]. Behrens [3601]. Being [970, 3329]. beliefs [532]. believing [2009]. belong [1561]. below [986, 1813]. Belson [1300, 1833]. belt [1574]. beneficial [1073]. beneficiaries [908]. benefit [1762, 2431, 579, 2133, 2042, 1810, 3468, 455, 3434, 683, 1732]. benefit-risk [3468, 455]. benefits [435, 652]. bent [2519]. bent-cable [2519]. Berger [969, 639]. Berkson [1453]. Bernoulli [1821, 1245, 2029]. Best [282, 3450, 2053, 2767, 3229, 297, 2912]. Beta [2684, 111, 1302, 246, 2083, 1668, 2971, 1405, 3192, 1878, 1885, 2637]. Beta-binomial [2684, 111, 2083, 2971, 1405, 3192, 2637]. Beta-mapping [246]. beta-regression [246]. Better [2117, 3394, 631, 1776, 2234]. Between [1227, 3543, 1062, 431, 2347, 3391, 1173, 1304, 275, 1768, 2873, 2397, 81, 849, 2339, 3616, 661, 279, 858, 910, 3080, 945, 1030, 388, 791, 1506, 2015, 1919, 402, 812, 26, 954, 1614, 1738, 595, 3396, 599, 392, 2179, 3207, 2859, 611, 796, 16, 868, 1857, 3357, 167, 3313, 814, 1181, 2772, 3635, 3554, 2538, 1847, 2776, 2788, 2197, 3115]. between-studies [2397]. Between-study [3543, 1304, 945, 595, 1857, 3115]. between-subject [910, 2859]. between-trial [3357]. Between-within [1227]. beyond [875, 3330]. bi [2261]. bi-chi-squared [2261]. bia [1441, 1440]. Bias [491, 922, 444, 1119, 2516, 1409, 80, 1336, 3435, 3025, 3507, 2212, 353, 2828, 1951, 283, 2020, 2059, 920, 10, 2266, 97, 763, 1621, 488, 1692, 403, 434, 494, 1188, 3225, 3109, 1195, 138, 2016, 1764, 562, 1098, 2960, 377, 290, 2618, 3156, 1815, 2163, 551, 1991, 807, 421, 1249, 1656, 1811, 3398, 250, 1003, 1788, 671, 2800, 2485, 371, 1161, 2086, 679, 1263, 1888, 2754, 1357, 2709, 333, 1345, 2675, 2418, 2736, 351, 3164, 2881, 2482, 2668, 3076, 69, 959, 1213, 2682, 1721, 3488, 459]. bias-adjusted [2709]. bias-corrected [3225, 1811, 3164, 959]. bias-correction [920]. Bias-reduced [80, 3435]. biased [922, 698, 2740, 1174, 1588, 3550]. bibliographical [2146]. bidimensional [3141]. Big [2558]. Bigirumurame [2846]. bilateral [1714]. Binary [1616, 1902, 3285, 654, 491, 555, 3657, 728, 3067, 1059, 3443, 3345, 335, 2764, 147, 112, 949, 2961, 761, 763, 3150, 155, 1337, 744, 897, 2083, 3070, 2235, 1021, 2979, 2264, 3300, 574, 2398, 2966, 713, 1141, 2034, 1566, 2753, 680, 2783, 1491, 1074, 869, 1670, 2992, 5, 348, 766, 3293, 280, 3241, 1065, 2633, 1956, 3628, 1600, 1656, 347, 1067, 1811, 2169, 2281, 2381, 2837, 3086, 1448, 1037, 2977, 2660, 2808, 3516, 1699, 2433, 2574, 1177, 3110, 830, 281, 273, 3435, 2586, 1383, 3481, 1919, 217, 402, 3318, 595, 1853, 2625, 3647, 3353, 2227, 3396, 958, 1093, 826, 714]. binary [3482, 655, 2807, 3039, 3533, 220, 251, 3040, 157, 404, 1937, 1145, 3425, 339, 56, 3000, 3002, 2943, 1437, 3490, 2714, 550, 2605, 2954, 1594, 2085]. binary-continuous [2264, 3293]. binary/categorical [3318]. binding [54]. binning [191]. Binocular [2686]. binomial [3095, 3370, 2684, 2423, 1743, 3142, 2083, 2924, 2355, 1603, 3009, 3096, 2971, 3576, 3426, 1626, 1943, 1065, 1405, 1016, 3631, 111, 3192, 1569, 3441, 3494, 1614, 2233, 1178, 1989, 2942, 2637, 1606, 1545, 1429]. binomially [2132]. binormal [2261]. Bio [271]. Bio-creep [271]. bioassay [2697]. bioassays [2974]. BioCycle [931]. bioequivalence [1463, 2640, 552, 1807, 1401, 1399, 1962, 3619, 3048, 2923, 2255, 2910, 2881, 1638, 1610, 1400, 2776, 2303, 2296]. biographical [2146]. biography [2544]. biologic [1035]. biological [1090, 1033, 2565, 3624]. biologically [1307]. biologics [1032, 1030, 1777, 2007, 1034]. Biomarker [1697, 1723, 2522, 1687, 811, 827, 647, 1623, 3344, 2732, 2960, 3402, 2042, 674, 2196, 3244, 783, 187, 3086, 3398, 1624, 2043, 2731, 600, 1998, 1972, 2554, 1765, 848, 2782, 1812, 832, 3356, 2493, 2392, 3243, 2374, 395, 829, 1011]. Biomarker-based [1687]. Biomarker-driven [2522]. biomarker-strategy [3244]. biomarker-stratified [3243]. Biomarkers [874, 2840, 2138, 2066, 2767, 1311, 823, 2406, 1817, 2809, 2777, 2467, 3438, 3157, 3528, 1645, 2097, 3577, 2162, 1051, 3038, 1289, 2730, 1082, 3001, 234, 1271, 3633, 502, 612, 42, 3362, 464, 322, 3394, 2479, 1782, 62, 2986, 951, 3270, 1348, 1584, 814, 2062, 2982, 2198, 1510, 2865, 1911]. biomedical [1281, 2812, 1017, 1283, 566, 1282, 976, 2216]. biometrical [1259]. biomonitoring [372, 1758]. Biopharmaceutical [1127]. biopharmacy [2004, 461]. biopsy [2120]. biosimilar [1027, 1028, 1029, 1031, 1037, 3137]. biosimilarity [1027, 1032, 1030, 1038, 1034]. biosimilars [1036]. biospecimen [823]. biospecimens [899, 2215]. Biostatics [2846]. Biostatistical [569, 2040, 2091, 3327, 999, 570, 568, 686]. biostatistician [3312]. Biostatistics [2012, 3299, 390, 869, 2573, 2040, 2159, 1648, 118]. biosurveillance [3620, 365, 359, 361, 363, 362, 360, 364]. bioterror [329]. biotherapeutics [2880]. bipolar [3577]. birth [3074, 850, 1219, 519, 3293, 1220, 564, 1190]. births [143]. birthweight [171]. bisulfite [2423]. Bivariate [3485, 1823, 1083, 260, 2083, 275, 71, 1292, 2382, 2264, 2780, 2784, 680, 2702, 3426, 1405, 1316, 3181, 773, 2056, 2357, 2064, 3337, 137, 1039, 357, 1493, 1352, 2375, 2659, 301, 32, 1594, 334]. BK [194]. bladder [2854, 664, 1655]. Bland [2287, 2825]. Blastland [1457]. BLCM [2823]. bleeding [2813]. blind [3264, 975, 2257]. Blinded [117, 2826, 1397, 1296, 2547, 2570, 1270]. blinding [3602]. block [2655, 1203, 1776, 2234]. blocked [2148, 2265, 2304]. blocks [667, 772, 1129]. Blood [3299, 2867, 2518, 3391, 932, 3054, 396, 2137, 3030, 1996, 1070, 607, 3129, 2904]. blood-brain [1996]. Boca [2012, 447, 432, 118, 106, 411, 2591, 412, 446, 2725, 506, 2691, 3059, 567, 433, 380]. Bock [3651]. body [3459, 3322, 3371, 2802, 3410, 59, 1124, 3337, 2803]. BOIN [3624]. bone [2388, 3093]. Bonferroni [2426]. Book [398, 2012, 447, 460, 432, 118, 520, 106, 1457, 3082, 411, 2591, 412, 2846, 446, 2725, 289, 3369, 506, 2691, 445, 3059, 3022, 1522, 749, 567, 2489, 3076, 433, 380].

Books [1457]. boosted [1251]. Boosting [1019, 2631, 3615, 3580, 2993]. Bootstrap [1593, 403, 898, 3092, 94, 3279, 2717, 494, 2388, 3170, 2751, 3018, 3418]. Bootstrap-based [898, 94]. Bootstrapping [3517, 1709]. BOP2 [2799]. boring [152]. borrowing [2628, 2147]. both [1781, 2442, 2918, 392, 1655, 2650]. boundaries [1643, 878]. boundary [3095, 3009, 3096, 3582, 3543]. boundary-optimized [3095, 3009, 3096]. bounded [35, 415, 355]. Bounding [421]. bounds [2045, 2851, 314, 1322, 3248, 247, 2741]. Box [1298, 1904, 2735, 277]. Bradford [2240]. brain [2925, 2499, 3130, 935, 3254, 3066, 1996, 2136, 1415, 2453]. branching [1770, 563]. Branscum [432]. Branson [72]. Brazil [1069]. breadth [1381]. breaking [975]. breast [2517, 1122, 427, 1979, 2498, 867, 2442, 1177, 1885, 1256, 792, 2703, 997, 1345, 351, 205, 3292, 232, 1258]. BreastScreen [3044]. breath [3077]. Brian [1442]. Brick [400, 2075, 1516]. Bridging [1127, 1907, 1745, 518]. bring [688]. broadening [2561, 2563, 2562, 2564]. Broemeling [2012]. bronchiolitis [3399]. Brookhart [1265]. brothers [10]. Brownian [2216, 537]. budget [1139]. Building [55, 1869, 693, 1173, 2027, 3394, 1765, 1387, 1583, 1639, 1775]. bump [735]. burden [3423, 369, 501]. Burzykowski [2846]. Buyse [2846].

C [584, 649, 900, 3, 1592, 428, 424, 3430, 760, 121]. C-statistics [428]. cable [2519]. caesarean [2263]. CAG [1506]. CAHPS [3380]. calculated [1696]. calculating [3505, 3503, 1464, 3504]. calculation [1466, 18, 2100, 2478, 1685, 2914, 3198, 3631, 1581, 3235, 1928, 3355, 1860, 928, 2690, 3593, 1719, 956, 1975, 1974, 2605, 1606, 1545, 1429]. calculations [23, 653, 275, 647, 1231, 2403, 740, 572, 3486, 2205, 161, 338, 2256, 641, 1843, 1819, 322, 1406, 2650]. calibrate [1528, 333]. calibrated [1646, 3318]. Calibrating [2420, 1532, 1305, 3149]. Calibration [3538, 1908, 2271, 1647, 1630, 510, 2368, 1593, 1439, 1886, 2768, 1906, 1914, 2415, 2292, 2413, 3446, 1585, 847, 1592, 1574, 2897, 1419, 2414, 1435, 1453, 1717, 2393, 1170, 1258]. California [376]. call [235]. Camblor [3021]. Campbell [289, 2202]. Can [639, 638, 3289, 3491, 2291, 3298, 247]. Canadian [1509]. canary [2570]. Cancer [1886, 3476, 996, 1386, 2636, 2517, 1595, 3415, 2210, 1818, 771, 3495, 262, 1449, 716, 1122, 942, 1531, 2566, 172, 1247, 2419, 1547, 427, 2271, 3109, 330, 808, 1980, 2854, 2371, 1408, 1015, 1665, 1979, 240, 2324, 2601, 2498, 1194, 2090, 3153, 3364, 303, 1730, 790, 3057, 91, 622, 678, 3638, 111, 867, 1260, 2442, 1177, 620, 3558, 1354, 1392, 75, 3235, 1703, 1928, 163, 1256, 792, 340, 515, 664, 591, 1485, 2703, 126, 997, 1785, 272, 2248, 3317, 2120, 1590, 3461, 60, 3348, 3467, 590, 2775, 1422, 486, 1672, 2583, 351, 1719, 205, 3346, 1343, 2741, 789, 2646, 3281, 2606]. cancer [3242, 1694, 1293, 3292, 3072, 232, 1737, 2695, 1150, 1258]. cancer-trend [330]. cancers [3109]. CanCORS [1308]. candidate [1640, 734, 1533]. capacities [2467]. capacity [2040]. capture [2140]. capture-recapture [2140]. captured [1640]. Caputo [3569]. CAR [2290, 2408]. carcinogenicity [3509]. carcinoma [3430, 1885, 263]. cardiac [3217, 274, 1772]. cardiopulmonary [2018]. cardiovascular [3391, 1521, 1070, 1949, 1715, 2904, 1368, 1011]. Care [1264, 3142, 2915, 3385, 1957, 2944, 376, 3079, 1494, 3566, 2957, 1111, 2917, 1364, 1856, 2639, 3160]. career [1560]. caries [345, 2233]. Carlo [2796, 737, 1149, 2769, 1622, 1989, 2211, 2104]. carotid [1600]. Carrasco [727]. carriage [753]. carried [1841]. carry [3630]. carry-over [3630]. carryover [245]. cartilage [2667]. Casale [967]. Case [208, 815, 3417, 3513, 1781, 389, 1818, 1882, 2676, 1119, 1232, 2323, 101, 2215, 3333, 2200, 2844, 2271, 768, 3551, 102, 1195, 1450, 2779, 808, 3275, 1979, 2869, 3512, 139, 1685, 146, 1331, 3561, 2340, 1301, 490, 1702, 2103, 3656, 388, 1796, 1611, 3060, 30, 1026, 252, 977, 1411, 3442, 87, 746, 930, 99, 1896, 472, 2617, 907, 1063, 642, 2505, 3499, 2635, 2510, 1190, 2734, 95, 274, 1996, 2276, 2359, 1178, 741, 1787, 1537, 2597, 1261, 443, 3549, 3548, 475, 1609, 1389, 767, 1164, 554, 3436, 352, 3595, 2229, 2039, 3270, 3349, 2983, 828, 283, 3000, 2440, 394, 1244]. case [3294, 316, 497, 609, 710, 1328, 1594, 1258, 2599]. case-association [1261]. case-cohort [3417, 2200, 1685, 2340, 1301, 2103, 746, 472, 95, 2599]. case-control [389, 1818, 1119, 1232, 2323, 101, 2215, 2844, 2271, 3551, 102, 1195, 808, 2869, 3512, 146, 3561, 1301, 1702, 3060, 1026, 252, 87, 930, 1896, 907, 642, 1190, 2734, 1996, 741, 1787, 1537, 2597, 3549, 3548, 1389, 1164, 352, 3595, 3270, 3349, 828, 3000, 2440, 316, 710, 1328]. case-control-family [2676]. case-crossover [815, 3333, 3275, 490, 2635]. case-only [252]. case-parents [609]. cases [2466, 2710]. catch [1247]. catch-up [1247]. categorical [1434, 3237, 124, 2006, 2016, 791, 1236, 1589, 1660, 2050, 709, 2447, 3318, 1823, 3567, 627, 1511, 1243, 626, 1508]. categories [138, 1051, 1756]. categorization [1854, 2878]. categorizing [1854]. category [2097]. Causal [2755, 3658, 1319, 1323, 3136, 1316, 531, 748, 3063, 1653, 1679, 1499, 1546, 3382, 3560, 2641, 3238, 2053, 1747, 1570, 1321, 1619, 3371, 1934, 3027, 2868, 3213, 135, 444, 488, 2235, 2873, 1488, 1526, 1650, 1957, 3221, 270, 1141, 2034, 888, 1098, 1494, 3131, 3016, 1322, 3248, 3181, 2169, 2281, 2850, 2134, 2424, 3619, 3065, 2726, 3401, 1142, 1190, 2734, 2153, 3365, 2500, 3393, 3474, 2996, 3315, 1288, 3390, 1648, 1338, 1659, 2482, 762, 1951, 1312, 2351, 3437, 993, 3174, 2085, 2197, 1320]. causal-inference [2197]. causality [3069, 3188, 3172]. causally [3374, 3610]. cause [2176, 308, 1627, 1844, 3587, 3526, 2663, 1729, 3509, 2940, 2879, 1883, 169, 3032, 2616]. cause-effect [3587]. cause-of-death [3509]. cause-specific [3526, 2663, 2940, 2879, 1883, 3032]. caused [598]. causes [1385, 91, 678, 1509, 1237, 611, 1720]. causing [1601]. caution [1331]. cautionary [1889, 1981, 1845]. CD4 [800, 192, 2216, 3280]. CDER [2333]. cell [932, 263]. cells [1828, 1519]. Censored [1214, 1929, 1828, 728, 2829, 1393, 633, 122, 1317, 1726, 2820, 20, 2689, 715, 3477, 3159, 1300, 941, 2317, 2696, 1547, 2927, 3023, 3197, 2371, 2856, 1478, 3097, 355, 2858, 2316, 730, 1838, 286, 2205, 1226, 1066, 3340, 70, 1780, 783, 3239, 3420, 2069, 1837, 2258, 3086, 99, 3128, 449, 1826, 1187, 1210, 3051, 1819, 1223, 820, 3223, 1519, 2687, 1892, 604, 1952, 171, 2189, 1527, 2569, 3103, 1501, 3119, 3467, 428, 2351, 2496, 3349, 167, 2001, 3148, 420, 1813, 327, 3053, 498, 2946, 2991, 533, 1852, 2645, 1578, 2673, 3147]. censoring [321, 2470, 923, 1462, 2804, 2964, 963, 71, 1291, 702, 2471, 2696, 621, 1167, 3623, 1995, 224, 1990, 1734, 1056, 1353, 440, 876, 2274, 3581, 3515, 2800, 2441, 1352, 354, 1625, 43, 2217, 2623, 814, 3091, 2930, 1733]. cent [6, 194]. centennial [850]. Center [2040, 1552, 3397, 473, 340, 2099, 3614]. center-patient [2099]. center-specific [1552, 3397]. centered [883]. centers [1887]. Centile [2165]. centiles [3503, 3504, 3505]. central [883, 1778, 1847]. centre [2655, 1398, 1618, 516, 1091]. centre-stratified [2655]. CER [882]. cerebral [3068]. cerebrovascular [1874]. certificates [2875]. cervical [2452]. cessation [65, 1616]. chain [1122, 737, 1149, 2047, 2908]. chained [1016, 356]. chains [3325, 2488, 3326]. challenge [2560, 1912, 1931, 2859, 2678]. challenges [2765, 3274, 141, 688, 694, 110, 1487, 1784, 2893]. Chambers [2691]. chance [3625]. chances [1733]. Chandrika [398]. Chang [106]. Change [1894, 2608, 1366, 426, 23, 1404, 1556, 2509, 239, 6, 1097, 1194, 150, 2781, 1481, 3057, 8, 2050, 2262, 194, 2432, 2757, 205, 1175, 906, 3003, 1083, 246, 383]. change-in-status [2050]. Change-point [2608, 1366, 2432]. change-points [3057]. changepoint [3520]. changepoints [384, 342]. Changes [2891, 1768, 2759, 1024, 3636, 2047, 213, 416, 393, 246]. Chapman [2012, 447, 460, 118, 106, 411, 2591, 412, 2846, 446, 2725, 2691, 3059, 567, 433, 380]. characteristic [1393, 212, 249, 1477, 843, 484, 2250, 3294, 1372]. characteristic-based [843]. Characteristics [999, 2480, 136, 2404, 1956, 1279, 2642, 600, 2452, 2790, 3536]. characterize [563, 1607]. characterized [3589]. Characterizing [577, 1395]. Chart [3379, 50, 3105, 2906]. charts [574, 2547, 3413, 2190, 3500, 3499, 2029, 2745]. Checking [96]. checks [1696, 2501]. chemical [372, 3375]. chemotherapies [1595, 486]. Chen [496]. chi [2261, 2538]. chi-square [2538]. Chichester [289]. Chikuma [1558]. child [3502, 3501, 1259, 2503, 3031, 916, 2532, 3498, 1879]. childhood [1452, 3392, 1677, 1246, 3336, 3518]. Children [142, 141, 345, 2595, 2262, 677, 1772, 143]. China [3418]. Chinese [692, 688, 30, 690, 696, 694, 695, 693, 691]. Chlamydia [1123]. choice [469, 2304, 756, 1131, 2078, 2557, 2833, 2827, 1753, 1034]. cholesterol [929, 1298]. cholesterol-lowering [929, 1298]. choose [2923, 3125]. Choosing [1550, 2586, 1, 1694, 1923, 2044, 584, 3]. Christakis [1045]. Christensen [432]. Christine [445]. chronic [3036, 522, 3394, 2867, 3552, 1264, 2706, 681]. Chronicles [1457]. circuits [2121]. Circular [2939, 1109]. Circular-circular [2939]. claims [1798]. claims-based [1798]. Clarification [778]. class [1710, 2091, 3327, 1370, 1642, 124, 1698, 1705, 745, 300, 3233, 3268, 2823, 1649, 397, 3086, 1589, 3166, 646, 1373, 285, 1079, 2137, 2086, 2118, 2782, 2003, 2711, 1185, 1264, 1615, 232]. class-level [2782]. classes [3563, 2250]. classical [1453]. Classification [942, 1798, 844, 2026, 3400, 3412, 1146, 3206, 1505, 2406, 2667, 2842, 3303, 2730, 3215, 3431, 2977, 3471, 285, 1079, 631, 1850, 3270, 1437, 395, 1306, 2203, 2954]. classifications [1740, 2015, 1577]. classifying [1486, 610]. Clegg [194]. client [2089]. climate [1069]. climate-sensitive [1069]. Clinical [3258, 2900, 3260, 107, 879, 3543, 3285, 2561, 3546, 1454, 2657, 391, 2655, 1863, 2106, 2911, 1608, 3370, 2091, 3327, 1398, 1370, 2764, 2563, 1806, 902, 1781, 3415, 1982, 77, 2004, 477, 777, 1688, 2365, 565, 1862, 771, 1310, 1612, 298, 39, 719, 3227, 2383, 3178, 2006, 1923, 887, 1021, 1490, 2560, 872, 1234, 172, 1778, 462, 1088, 1379, 2890, 2884, 798, 1500, 1396, 882, 860, 271, 2419, 1547, 874, 3264, 578, 466, 988, 1191, 661, 1817, 2809, 427, 3473, 2628, 579, 2777, 1829, 571, 1957, 1473, 2403, 461, 1269, 1910, 2147, 516, 1700, 1795, 174, 1986, 884, 2998]. clinical [3265, 3097, 1324, 2341, 381, 2662, 465, 2082, 1809, 2302, 1092, 2133, 2604, 2061, 1834, 3246, 303, 2241, 869, 948, 3038, 992, 551, 150, 2901, 1134, 766, 799, 1216, 3151, 2958, 2334, 725, 1040, 1360, 1151, 1549, 3332, 3628, 1777, 2007, 690, 881, 873, 1099, 1101, 187, 1031, 1950, 3631, 1448, 791, 2863, 3633, 2573, 1131, 918, 1960, 8, 111, 694, 876, 867, 1624, 1907, 2285, 3073, 3542, 1699, 1252, 3619, 3261, 53, 641, 1116, 3464, 1636, 3203, 3029, 1904, 2617, 273, 788, 3175, 1100, 509, 464, 769, 662, 1687, 573, 177, 2582, 1714, 2331, 2885, 2562, 2578, 2864, 1634, 3419, 3585, 1133, 1562, 1268]. clinical [2990, 1039, 1406, 700, 1938, 1790, 1793, 2564, 445, 601, 3257, 630, 535, 414, 148, 1736, 2107, 2892, 15, 607, 2152, 1785, 1357, 1115, 2195, 1939, 2986, 2709, 689, 2295, 2242, 530, 2893, 3461, 220, 3214, 98, 721, 2014, 2065, 2659, 3165, 2526, 2418, 3460, 708, 2775, 1050, 1522, 983, 2335, 3330, 1672, 2583, 3125, 3552, 1792, 704, 2039, 2671, 2794, 3144, 3437, 665, 1563, 181, 56, 499, 1673, 1936, 2798, 859, 3259, 1193, 1596, 3078, 1917, 3635, 3305, 3301, 1791, 512, 436, 944, 2567, 2565, 67, 380, 1737, 1893, 2605, 3133, 2934, 3383, 781, 1776, 871, 2695, 2799, 3089, 3088]. clinical [1903]. clinically [2919]. clonal [163]. clone [1224]. cloning [2068]. Closed [2713, 2905, 258]. Closed-form [2713]. Cloud [3529]. Cloud-based [3529]. clump [45]. clumping [1813]. Cluster [291, 2643, 1514, 1959, 862, 155, 2259, 328, 489, 3140, 25, 2382, 3255, 2862, 233, 2272, 3245, 3395, 3606, 3342, 3209, 1805, 2999, 44, 2478, 2783, 1597, 3389, 1460, 1061, 2044, 961, 2546, 3486, 286, 2449, 1249, 1656, 657, 1811, 2852, 3308, 3516, 3483, 349, 1428, 3195, 1581, 3609, 2586, 2751, 915, 978, 3579, 1261, 1786, 712, 806, 3107, 2766, 809, 2366, 3490, 1852, 672, 3143, 1635]. cluster-correlated [3579]. cluster-level [1514, 1959, 712, 3107]. cluster-randomized [1514, 1061, 1811, 2766, 1852]. Clustered [2353, 1309, 1214, 2263, 3491, 1105, 1839, 20, 2994, 2101, 81, 3255, 3209, 3596, 2011, 1460, 592, 1276, 286, 280, 1066, 1956, 70, 605, 3483, 1354, 1456, 2, 1180, 2716, 1786, 1407, 3103, 3531, 931, 3338, 2496, 3000, 420, 94, 184, 1587, 3007, 1576, 1455]. Clustering [1693, 2907, 3132, 1109, 2978, 2259, 1525, 548, 2949, 2999, 240, 3510, 3230, 3254, 2038, 2814, 1856, 2639, 2309, 104]. clusters [1754, 233, 1759, 3486, 425, 958, 1114, 980, 3554, 1635]. CMaxSPRT [3408]. Co [1509, 1608, 2891, 2025, 711, 207, 2314, 485, 3533, 220, 2117]. co-data [2117]. co-development [711]. Co-infection [1509, 485]. co-occurring [2025]. co-primary [1608, 2891, 207, 2314, 3533, 220]. coal [2570]. coarseness [2018]. Cochran [2128, 2130, 2129, 2127, 2319, 593]. Cochrane [1991]. code [3513]. coefficient [1820, 454, 1438, 2481, 3202, 3247, 3154, 682, 3122, 2635, 1527, 1156, 1615, 1757, 68]. coefficients [1274, 2437, 2643, 217, 525, 2656, 343, 2878, 1081, 2993, 2453, 957]. coexist [3298]. cognition [1079]. cognitive [321, 267, 2608, 2118, 802, 3144]. coherent [3097]. Cohort [270, 1928, 583, 3459, 3322, 3417, 2965, 2820, 2200, 141, 1526, 1468, 1685, 1447, 2340, 1730, 1301, 2103, 89, 977, 2957, 3406, 746, 3398, 472, 2748, 95, 274, 2359, 126, 1347, 2104, 352, 351, 11, 2599, 1509]. cohorts [453, 603, 2577, 3100, 230]. coin [2059, 922, 698, 2740]. Cole [1265, 3505, 3504]. Collaboration [569, 1262]. collaborations [390, 570, 568, 686]. Collaborative [2975]. collapsed [138, 1107, 1955]. collected [2956]. Collection [1128]. collective [3254]. colon [3415, 1531, 735, 1980, 3364]. colony [2976]. colorectal [3495, 808, 2371, 678]. combination [3479, 1350, 83, 1147, 2767, 1848, 1795, 174, 2021, 2097, 3196, 502, 612, 1703, 282, 1790, 1793, 2122, 779, 1326, 52, 3343, 486, 157, 1548, 1792, 3281, 1791, 2494]. combinations [2925, 3344, 1334, 224, 1706, 1023, 2550, 1051, 110, 1377, 2845, 1672, 2583, 1558, 499, 2374, 1510]. combinatory [2985]. combine [584, 284, 3, 2984, 1]. Combined [3546, 949, 3406, 2357, 1487, 2990, 1512, 56, 1917]. Combining [3279, 1333, 2249, 1073, 1502, 2649, 2406, 905, 1665, 2730, 1248, 3398, 3203, 2027, 173, 2062, 2982, 481, 1146, 1361, 1423, 1817, 2809, 102, 1967, 3613, 1371, 940, 1151, 2495, 3356, 2229]. come [2845]. comes [1682, 1851, 2557, 2833]. commensurate [1466, 2784, 2035]. Comment [628, 3325, 853, 1085, 854, 541, 2128, 459, 2821, 615, 194, 1320, 1071, 1086, 2142, 542, 3551, 2625, 1046, 3033]. commentaries [1048, 546, 886]. Commentary [3503, 882, 545, 2334, 2130, 881, 1947, 3259, 3505, 1866, 2796, 569, 2113, 3504, 570]. Comments [1281, 3095, 3648, 1057, 2296, 2106, 852, 3444, 2563, 543, 1982, 361, 363, 3354, 584, 1028, 647, 2884, 614, 2219, 3654, 72, 768, 649, 362, 884, 2112, 586, 2011, 2971, 2959, 3656, 3, 3653, 1865, 2539, 130, 347, 3134, 2325, 2574, 2115, 200, 1592, 2562, 3355, 2122, 1945, 2564, 360, 287, 408, 1440, 2807, 2333, 767, 1265, 1889, 132, 3297, 2114, 1558, 1792, 131, 1045, 364, 1974, 1400, 184, 1791, 2220, 3133, 1545, 2520, 3459, 2116, 1283, 1441, 3071, 2903, 2885, 3650, 1793, 1981, 202, 1975, 409, 3021, 1606, 1401]. common [3300, 1072, 2309]. commonality [3348]. commonly [419]. communities [3207]. community [3004, 1678, 725, 492, 3348, 1679]. community-based [492, 1679]. comorbidity [915]. companion [711, 1950]. Comparability [1038, 3139]. Comparative [885, 196, 202, 3089, 1806, 2466, 3178, 3477, 882, 1869, 197, 574, 884, 2021, 1074, 1999, 2154, 198, 881, 622, 1431, 2726, 2485, 2447, 1579, 1842, 2122, 2387, 201, 3145, 3469, 1741, 3330, 1653, 200]. compare [1899, 2142, 2006, 838, 1090, 1754, 3481, 2141, 1372]. compared [283, 2301]. Comparing [3502, 442, 2876, 2148, 46, 1229, 264, 2700, 1817, 2809, 1186, 940, 1838, 916, 2226, 3104, 2248, 1118, 2450, 2526, 2749, 2217, 3160, 1517, 23, 1393, 2365, 1060, 2856, 1552, 2880, 1151, 1248, 2968, 631, 15, 1868, 2659, 3593, 762, 1936, 1437, 2865, 374, 1606, 1545, 1429, 1713]. Comparison [720, 2451, 3206, 457, 1173, 401, 1731, 1536, 2644, 3462, 1956, 93, 977, 2593, 3542, 2096, 1467, 217, 2709, 370, 2699, 1422, 2535, 624, 1484, 1059, 2249, 1342, 963, 1669, 547, 846, 96, 2384, 947, 1073, 1208, 2339, 2696, 3225, 935, 785, 1238, 2760, 2796, 888, 740, 1965, 2989, 572, 3010, 1224, 3106, 164, 1077, 2815, 3506, 3231, 2134, 213, 2769, 2262, 1378, 1876, 2190, 3101, 2874, 256, 2705, 1953, 839, 1218, 393, 2487, 3039, 310, 2904, 13, 1511, 3644, 1145, 1563, 3618, 3287, 864, 1275, 2857, 246]. Comparisons [2182, 2597, 76, 315, 3620, 1199, 314, 1395, 1611, 269, 1236, 1660, 2041, 3046, 924, 868, 2776]. compartmental [723]. ComPAS [3343]. compatibility [3347]. compatible [3204, 2336]. Competing [1680, 725, 1790, 3047, 724, 1628, 2427, 2648, 2882, 3034, 3316, 3513, 308, 2829, 321, 243, 1393, 2937, 2101, 1229, 953, 1630, 1417, 1728, 3171, 473, 1810, 1228, 387, 3638, 58, 773, 3583, 1072, 2274, 916, 1230, 123, 1157, 2879, 274, 2118, 1268, 3084, 2711, 3283, 1493, 169, 43, 3125, 3148, 3032, 3387, 498, 1690, 2651, 2616, 1733, 3525, 1793, 1792, 1791]. Competing-risks [1680]. complete [2655, 139, 1401, 1399, 283, 1400]. complete-case [283]. completed [640, 514]. completely [2728, 2512, 1445]. completion [565]. complex [3444, 203, 3517, 188, 417, 1104, 1105, 1959, 2588, 1307, 2915, 1041, 836, 2017, 3087, 2119, 1921, 2095, 2981, 3017, 2512, 3247, 1007, 3278, 508, 446, 2515, 443, 1785, 3598, 1120, 1880, 3540, 2234, 1054, 2799]. complexity [3568, 86]. compliance [2911, 147, 1934, 1650, 1418, 531, 3609, 292, 1855]. complications [1188]. compliers [2153]. Component [3258, 3260, 183, 204, 2278, 3261, 127, 3054, 1273, 3257, 3259, 1189]. components [84, 483, 3266, 610, 2990, 3396, 2174, 2656]. Composite [3440, 2892, 2211, 2803, 2106, 3199, 2596, 1060, 30, 1202, 215, 2289, 1235, 1268, 2990, 3396, 1736, 2107, 2212, 422, 3644, 2502, 2058, 926]. compound [3354, 2308, 1374, 32]. compromising [788]. computation [3556, 3123, 452]. computational [412]. Computationally [1970, 3240, 3614]. computations [1012]. compute [3612]. computer [2055]. computer-intensive [2055]. computing [66]. concealment [969, 968]. concentration [2841, 1424, 3267, 3540]. concentration-response [2841]. concentrations [1214, 2595, 607]. concept [3155, 107, 474, 2864, 3289, 1128]. conception [141, 1733]. Concepts [3513, 108]. conceptual [1369]. concerning [1889, 1981, 1845]. concerns [2848]. concluding [536]. Concordance [1354, 2656, 1818, 3455, 1686, 1438, 1167, 645, 3511, 1456, 915, 1493, 1455, 2435]. concordant [2613]. concurrency [2458]. concurrent [2720, 3031, 2579]. condition [2387]. Conditional [1105, 596, 674, 1622, 1773, 3119, 815, 147, 963, 1705, 1315, 2246, 988, 1191, 102, 2407, 2612, 149, 80, 459, 2468, 1722, 3060, 33, 3511, 3127, 2050, 2084, 3472, 2995, 2932, 2181, 1000, 2448, 2709, 353, 1823, 1672, 3099, 951, 2928, 2598, 3033, 1666, 2135, 3582, 3588, 2922, 2788]. conditional-mean [1315]. Conditionally [1216, 923]. conditions [1219, 2025, 3603, 915, 2038, 159, 536]. condom [503]. conduct [918]. conducted [675, 2634]. Conference [107]. Confidence [2045, 1637, 1211, 3204, 3642, 1371, 290, 1569, 616, 16, 906, 1538, 3658, 1593, 1642, 403, 2235, 1113, 635, 1603, 2317, 3136, 3612, 1135, 3303, 1706, 2858, 2781, 314, 1613, 1464, 2169, 2281, 1252, 1174, 1675, 2910, 3494, 2585, 2352, 1614, 617, 958, 3627, 175, 79, 590, 13, 2875, 3295, 1106, 1479, 7, 176, 1372]. confidential [1543]. confidentiality [3087, 2013]. configuration [3272]. configurations [1314]. Confirmatory [2115, 2657, 2111, 3055, 448, 467, 2112, 3306, 2693, 2113, 15, 98, 3165, 2114, 2039, 3133]. confirmed [3507]. confounded [2966, 997]. Confounder [304, 3166, 993]. Confounder-adjusted [304]. confounders [2469, 1959, 3477, 2866, 1751, 1978, 1359]. Confounding [1600, 2176, 3451, 1854, 2182, 188, 417, 1104, 1126, 2380, 3023, 270, 1654, 979, 3446, 2503, 421, 1870, 2356, 2511, 3058, 671, 2769, 3610, 748, 3163, 714, 1786, 1801, 352, 2827, 835, 3594, 2454, 2796]. confounding-by-indication [1870]. congenital [2321]. Connecticut [2854]. connection [3582]. Connections [1738]. connectomes [3635]. consensus [3576]. consensus-based [3576]. consequences [1197, 1483, 2557, 2833, 2736]. conservative [2301]. Considerations [742, 1092, 3636, 3025, 3550, 2657, 77, 463, 1029, 2098, 3385, 2894, 823, 1060, 1611, 2539, 1031, 2226, 1063, 1049, 3500, 1290, 1014, 1303, 3356, 2286, 2540, 3490, 3133, 3376]. considering [471, 2322]. consistency [160, 96, 947, 768, 2285, 1133, 1562, 767, 518, 3621]. consistency-adjusted [160]. consistent [3169, 3197]. Consonant [258]. CONSORT [875]. constant [1057, 1056]. constitutes [742]. Constrained [483, 3295, 2954, 2224, 2837, 261]. constraint [1192]. constraints [2596, 3036, 1139, 3252]. construct [2304]. Constructing [2758, 3505, 3503, 3504, 3499]. Construction [7, 176, 1121, 2190, 3500, 2352, 1415, 13, 175]. consumer [2881, 2255]. consumption [1112, 709]. contact [426, 2952]. contagion [1048, 1044, 1046, 1045]. contamination [223]. contemplate [3172]. content [3584]. context [3513, 2707, 1431, 200, 196, 1253, 3367]. contextual [2630]. contingency [3648, 1828, 1899, 2142, 407, 1519, 3555, 3650, 2141, 406, 3059, 2843]. Continual [507, 469, 2584, 989, 510, 1907, 511, 2582, 515, 1009, 1170, 950]. continuation [754]. Continuous [3560, 1643, 3352, 555, 587, 3067, 2138, 3339, 84, 1173, 2916, 761, 763, 862, 1412, 1147, 2434, 547, 3019, 2000, 3205, 1208, 1206, 660, 2862, 785, 2264, 2398, 3395, 2439, 2960, 1566, 680, 1992, 415, 1491, 1091, 766, 2958, 3293, 2196, 556, 1339, 3231, 2981, 791, 2912, 2877, 3073, 523, 531, 2047, 450, 1840, 1344, 3617, 3102, 17, 1075, 889, 3472, 878, 2698, 1207, 2181, 125, 1285, 1735, 2908, 2441, 619, 1114, 223, 2493, 1345, 1749, 1653, 2230, 665, 1343, 1727, 2878, 481, 334]. continuous-monitoring [878]. Continuous-time [3560, 3339, 1339, 2908, 2441]. contour [2580, 2494]. contra [3262]. contrast [2584, 1482, 3618]. contrast-based [3618]. contrast-enhanced [1482]. Contrasting [984]. contrasts [774, 2482]. contribution [2287]. contributions [1452, 1648, 859]. Control [2547, 189, 389, 1818, 2469, 2676, 1119, 1232, 2323, 101, 448, 2215, 3616, 2844, 3452, 2367, 2271, 2628, 3551, 102, 1195, 808, 3023, 3379, 1269, 2945, 3403, 1808, 2869, 3512, 146, 3561, 1301, 1702, 3186, 314, 1226, 269, 3060, 1026, 252, 87, 930, 1013, 1660, 1308, 1896, 471, 907, 642, 1190, 2734, 1996, 601, 741, 1787, 1537, 2597, 3549, 3548, 1326, 50, 1389, 1164, 352, 3595, 1672, 2827, 3270, 3349, 2741, 828, 1936, 3000, 1855, 2440, 3291, 3294, 2513, 316, 710, 2906, 1328, 2029, 2745]. Controlled [3657, 3040, 2648, 2364, 291, 2765, 1829, 1450, 2779, 888, 1805, 1074, 1331, 1441, 3381, 388, 690, 1411, 1263, 1440, 1632, 689, 3533, 3469, 3384, 115, 2983, 394, 1244, 2922]. Controlling [138, 3048, 1751, 3563, 1136, 2071]. controls [2044, 2710, 3441, 642, 2697, 3208]. controversies [1215]. controversy [3599]. convenience [377, 3131]. conventional [3330]. convolutive [2572]. Copas [1003]. COPD [1128]. Coping [1238]. Copula [3296, 334, 71, 2979, 3616, 1018, 3293, 3200, 2064, 3492, 3531]. Copula-based [334, 71]. copulas [1055, 1925, 2989, 1405, 3539, 3091, 1733]. copy [431, 318, 163]. coregionalization [2408]. coregionalized [2408]. cores [2120]. Cornfield [855, 850, 852, 853, 860, 854, 858, 856, 859]. coronary [1632, 2198, 1096]. Correa [1865]. correct [940, 2918, 3475, 1419, 410]. Corrected [2714, 996, 12, 3225, 1811, 3164, 959]. Correcting [2423, 302, 2762, 1888, 1345, 1855, 2399, 1733, 996, 3562, 3271, 1413]. Correction [3658, 3657, 1087, 1747, 3530, 1683, 3327, 3432, 1458, 2887, 3652, 1188, 2236, 1191, 1472, 1620, 3655, 1890, 1465, 1402, 218, 3483, 1267, 1746, 1983, 2886, 37, 559, 1442, 3647, 2602, 686, 1658, 63, 219, 3649, 69, 2326, 3651, 1239, 1546, 943, 2438, 920, 403, 1119, 1553, 1557, 778, 1161, 3025, 82, 277, 1213, 668, 2669]. Corrections [2958, 616, 3018]. Correlated [2965, 2439, 2677, 1214, 751, 2451, 3206, 3604, 2588, 2164, 2209, 3361, 898, 1568, 660, 2287, 2398, 2856, 1680, 680, 2436, 3577, 2959, 1838, 1716, 3602, 1225, 3631, 29, 2920, 1177, 3071, 373, 1383, 1714, 357, 3579, 617, 251, 52, 1169, 1875, 1242, 2068, 668, 1372]. correlated-errors-in-variables [2287]. correlates [2403, 3571]. Correlation [2702, 2289, 3427, 3501, 1839, 1899, 2066, 2142, 1311, 1438, 3392, 3395, 2468, 2011, 2245, 378, 3005, 2546, 3122, 3556, 623, 2141, 3396, 2716, 247, 958, 1352, 2179, 1644, 182, 796, 2656, 2718, 1213, 1564, 2366, 957, 2788]. Correlation-adjusted [3427]. correlation-based [2179]. correlations [275, 814]. correspondence [318, 1677]. correspondences [858]. cortex [3068]. cosinor [3054]. Cost [1637, 624, 1205, 1541, 2422, 3279, 1318, 170, 1495, 1138, 2858, 2258, 3449, 1187, 1581, 500, 62, 1364, 727, 726, 532, 144, 2351, 1120, 2513, 2076, 533, 2954, 456]. cost-effective [2513]. cost-effectiveness [1541, 2422, 3279, 170, 1495, 1138, 1581, 144, 2076]. cost-efficient [62, 2954]. cost-offsets [500]. Costa [1865]. costs [1541, 2164, 2449, 2935, 500, 3460, 3117, 3143]. could [468]. count [676, 1309, 2978, 2247, 1105, 2976, 529, 964, 2588, 2320, 2994, 25, 1296, 2307, 2813, 3565, 117, 1238, 2547, 1269, 3233, 2959, 3597, 1716, 3566, 1225, 3631, 2792, 3416, 2549, 3071, 2504, 2505, 1367, 324, 1270, 1744, 3138, 2024, 1849, 2985, 1102, 3640, 2622, 1182, 1189, 2680, 1900]. counties [1538]. counting [2555]. countries [2147]. counts [628, 3370, 479, 800, 1943, 1955, 1565, 192, 429, 2216, 3280, 1293]. county [1293]. county-level [1293]. Coupled [1785]. course [2280]. Covariance [1591, 2232, 2521, 1727, 751, 597, 1351, 483, 1055, 3225, 1802, 3079, 1209, 3216, 1909, 1540, 53, 1183, 2078, 1933, 806, 493, 959, 1213, 3575, 2695, 2520]. covariances [1094]. Covariate [2201, 2964, 240, 438, 3641, 560, 1370, 2204, 328, 3050, 453, 603, 1291, 3159, 947, 3255, 1526, 1167, 3221, 1473, 138, 1553, 1409, 2783, 2618, 582, 1504, 2316, 1874, 150, 1958, 2730, 1327, 30, 813, 8, 1284, 3581, 1344, 385, 534, 1410, 1184, 1919, 113, 177, 17, 2762, 3318, 1071, 2497, 2181, 2908, 1952, 1444, 3532, 2631, 1859, 293, 1935, 3450, 1155, 283, 1102, 1639, 3295, 618, 1755, 436, 944, 3383, 561, 2386]. covariate-adaptive [1284]. Covariate-adjusted [438, 3641, 1370, 2908, 3532, 1755, 944]. covariate-adjustment [113]. covariate-balanced [436]. covariate-dependent [2783, 1504, 813, 2762, 1102]. covariate-missing [3295]. covariates [555, 3185, 955, 2477, 3028, 84, 1173, 294, 186, 2349, 1726, 2466, 2744, 1229, 3347, 2471, 1315, 1488, 947, 2200, 2339, 713, 526, 1433, 643, 3216, 139, 224, 1092, 3182, 2992, 1922, 1740, 92, 1411, 2816, 3239, 1600, 2348, 791, 3449, 1187, 2274, 1830, 2750, 127, 549, 3515, 1282, 1416, 3126, 256, 3167, 2338, 2479, 1739, 2353, 3163, 619, 1782, 230, 1080, 1364, 3390, 1156, 712, 342, 1453, 809, 2943, 576, 1076, 2046, 1831, 2682, 3635, 3149, 2878, 2946, 1641, 1087]. Coverage [2585, 2354, 763, 752, 3422, 2948, 2875, 1694]. Cox [1087, 955, 2477, 2072, 3374, 3169, 270, 909, 3157, 2529, 1433, 643, 1734, 3597, 2707, 3182, 2992, 1298, 3553, 1066, 1327, 3060, 88, 1464, 87, 1341, 2511, 1007, 1904, 2735, 1344, 784, 2806, 1416, 256, 1071, 3393, 3269, 2305, 2827, 3409, 801, 926, 94, 1181, 618, 2682, 1480, 1451]. Coxian [2407, 757]. CP [1722]. craniofacial [2489]. CRC [2012, 447, 460, 432, 118, 106, 411, 2591, 412, 2846, 446, 2725, 506, 2691, 3059, 567, 433, 380]. credible [2842]. credit [795]. creep [271]. CRIC [2867]. criteria [3546, 2973, 3035, 3576, 2806, 2289]. criterion [2133, 2245, 592, 474, 1110, 86, 3104, 1265, 2620]. critical [1864, 1866, 1865, 1038, 3367]. Cronbach [68]. Cross [3508, 3068, 2953, 1188, 3245, 3606, 2784, 588, 1962, 2719, 379, 1124, 311, 2036, 312, 2586, 3499, 2686, 2352, 61, 1432, 1368, 681]. cross-lagged [2784]. cross-over [1962, 311, 312, 2586]. Cross-sectional [3508, 3068, 1188, 3245, 379, 1124, 2036, 3499, 2686, 2352, 61, 1432, 1368, 681]. cross-validation-based [2953]. cross-validatory [2719]. crossed [2356, 2068, 1257]. crossing [2880, 193, 1724, 2991]. crossover [815, 2729, 633, 780, 3333, 552, 3606, 3275, 770, 297, 1401, 1399, 2550, 3486, 490, 2958, 2381, 2464, 2672, 3308, 437, 2325, 2433, 2574, 1469, 722, 2635, 245, 24, 2845, 3630, 3158, 1400, 1076]. Crouch [445]. crude [91]. CTT [401]. cubic [105]. Cumulative [1551, 2829, 2955, 1964, 2307, 2471, 2771, 3637, 336, 3526, 2663, 608, 1729, 791, 2831, 1237, 2940, 2879, 1625, 2805, 2185].

Cure [272, 2477, 2470, 2288, 3050, 2937, 2371, 2746, 3232, 65, 867, 1180, 1633, 340, 2189, 3531, 3467, 956, 1975, 1974, 420, 1451]. cured [1531, 3193, 1527, 1567, 2326]. Current [872, 860, 1462, 1113, 3414, 1944, 1948, 2880, 93, 2931, 1684, 1330, 1947, 2159, 1945, 1407, 3514, 1155]. curtailment [501]. curve [2354, 189, 454, 2406, 3014, 2761, 2888, 1764, 156, 2261, 514, 212, 438, 3462, 1649, 249, 1255, 226, 382, 299, 3120, 894, 2583, 927, 3294, 2982, 1880, 1479, 2279, 458, 691]. curves [1593, 2451, 2840, 457, 1393, 2758, 3361, 1580, 2291, 1770, 3217, 149, 1504, 59, 2654, 1573, 1773, 480, 2181, 1118, 2054, 939, 3323, 484, 2135, 374, 1372]. customized [3456]. CUSUM [3403, 3413, 393, 3105, 2029, 2745]. cut [844, 1826, 285, 1968, 1755]. cut-off [285]. cut-point [1826, 1755]. cut-points [1968]. Cutoff [492, 1765]. cutpoint [299]. cycle [3410]. cycles [3410, 825, 1590, 2568]. cystic [2262, 954]. cystitis [1655]. cystitis/painful [1655]. cytokine [1214]. cytostatic [1052]. cytotoxic [1595].

D [628, 777, 72, 150, 727, 1265, 1889, 2489, 184, 2271, 2012]. D’Agostino [2106, 647]. daily [761, 3454, 3470, 2335, 2321, 1831]. Dakar [11]. damage [2698, 2555]. damage-resistance [2698]. danger [1457]. Daniels [447]. Data [1433, 2334, 694, 3013, 3304, 2558, 749, 787, 1109, 398, 1517, 165, 426, 654, 827, 3276, 676, 1929, 1309, 3479, 1313, 3501, 2962, 2941, 1628, 3185, 2458, 2263, 266, 2882, 2978, 1492, 3417, 2973, 3035, 308, 2829, 2438, 3491, 3443, 912, 321, 987, 2201, 2664, 447, 3339, 3562, 920, 2247, 203, 457, 112, 1539, 3329, 1781, 2868, 3517, 389, 2688, 3325, 3279, 1971, 2961, 633, 3237, 188, 417, 1104, 1105, 2758, 2916, 761, 1886, 1333, 2249, 2624, 368, 529, 964, 3307, 1959, 2721, 3604, 1172, 1342, 403, 2952, 1412, 3354, 2676, 3495, 2349, 1820, 2588, 122, 489]. data [527, 897, 1310, 1318, 1317, 1351, 1272, 1726, 1462, 1993, 2164, 2102, 2594, 2744, 2804, 2839, 3055, 3140, 55, 432, 937, 1631, 20, 1113, 2994, 2343, 2101, 25, 716, 2006, 275, 1296, 2689, 1942, 1375, 1254, 998, 1307, 1674, 2188, 2915, 448, 3310, 2519, 2300, 2870, 3347, 1008, 1292, 1041, 2307, 3121, 846, 1234, 2000, 911, 715, 2979, 3477, 745, 2813, 2838, 3447, 18, 1073, 1396, 3414, 3255, 3637, 1470, 2437, 2246, 3424, 1946, 941, 2317, 2696, 2488, 3326, 3565, 2862, 423, 1547, 235, 578, 670, 117, 785, 2927, 1754, 685, 3363, 1085, 1668, 2628, 2777, 1957, 1802, 384, 3023]. data [2944, 752, 921, 2398, 1980, 1064, 2017, 1944, 1948, 2570, 1417, 2951, 270, 3197, 376, 3209, 2183, 2413, 1269, 713, 2854, 3275, 405, 516, 1681, 1728, 2371, 2956, 1795, 1766, 2856, 3626, 557, 1478, 2310, 430, 640, 1665, 1566, 3397, 910, 80, 3233, 1677, 2667, 1427, 3080, 3446, 982, 1797, 3512, 2615, 2998, 2436, 706, 1992, 3216, 3097, 2747, 1554, 3171, 224, 1990, 816, 2618, 582, 940, 645, 1024, 1575, 1803, 2037, 2553, 2601, 3596, 415, 3087, 1809, 2702, 2971, 2959, 1460, 3411, 119, 1074, 267, 1401, 1399, 1954, 2319, 2858, 1834, 2090, 2316, 3597, 1920, 730, 2880, 592, 869, 190, 961, 1745]. data [948, 1441, 2914, 3038, 1835, 992, 1716, 2163, 1301, 286, 183, 1298, 1693, 3553, 3656, 1626, 1226, 280, 1066, 2457, 1065, 1796, 1107, 2815, 3566, 1956, 2330, 2194, 1016, 3340, 556, 14, 1740, 70, 93, 1921, 3602, 2154, 1339, 1780, 204, 3232, 92, 977, 783, 678, 1316, 1390, 1591, 2095, 2607, 3239, 3006, 2957, 3001, 3406, 3638, 3442, 3231, 1418, 59, 249, 58, 65, 187, 746, 930, 896, 1067, 1250, 1225, 1551, 1833, 2348, 2654, 2818, 2850, 2855, 3198, 2981, 3017, 3592, 3480, 919, 1872, 29, 99, 1125, 1236, 1926, 3128, 3572, 2573, 918, 1960, 449, 1826, 2278, 2512, 2920, 3188, 3154, 1187, 379, 1308, 605, 3483, 53, 682, 641, 1602]. data [2931, 2634, 3065, 2433, 2821, 349, 1177, 1287, 1428, 1446, 1707, 3192, 3464, 2050, 3011, 3173, 2408, 3051, 179, 3077, 2587, 1530, 1124, 2792, 3558, 3509, 3203, 3416, 2372, 709, 1354, 1456, 1840, 732, 1344, 1955, 3334, 1684, 1330, 2800, 2907, 3441, 2549, 907, 3045, 1972, 385, 319, 3071, 3172, 1667, 3435, 1509, 306, 1663, 3037, 1223, 834, 1383, 1356, 2, 1022, 1876, 2504, 2505, 3234, 2554, 3101, 2357, 684, 215, 1174, 1180, 820, 1043, 210, 173, 292, 3499, 446, 3251, 163, 3223, 1543, 3098, 2005, 3030, 2687, 792, 1724, 1004, 1676, 2186, 889, 1714, 221, 2028, 2332, 2734, 954, 1010]. data [1614, 1263, 1947, 2995, 3193, 3208, 3399, 1084, 1442, 2244, 2531, 3365, 228, 1892, 2705, 1512, 329, 3084, 1359, 2537, 3240, 2938, 1207, 2337, 3579, 1945, 1176, 2532, 506, 2136, 776, 2716, 1374, 61, 2353, 583, 994, 1086, 1784, 1233, 324, 958, 741, 1537, 2597, 924, 604, 997, 15, 3641, 1493, 57, 443, 344, 607, 826, 1440, 1218, 1397, 1270, 392, 2414, 2189, 1786, 3482, 1221, 1744, 1527, 158, 1718, 2996, 1080, 2448, 1364, 2333, 2709, 49, 1616, 2035, 1303, 1012, 2942, 3408, 3129, 3630, 1607, 1800, 3356, 1750, 3461, 3470, 2569, 1482, 1179, 1382, 1609]. data [3141, 3557, 296, 2216, 52, 3132, 1435, 1407, 1933, 2690, 3103, 3531, 3359, 3022, 3282, 525, 1164, 1823, 2814, 3119, 3467, 890, 708, 257, 1741, 16, 627, 1169, 1846, 2666, 3384, 342, 1453, 3338, 3123, 811, 1966, 2392, 3330, 3164, 157, 404, 422, 1937, 428, 1185, 301, 831, 782, 567, 2722, 2489, 293, 2623, 3284, 69, 205, 493, 1617, 1849, 2049, 2502, 2351, 2400, 3514, 167, 1343, 2830, 3621, 1997, 1666, 1155, 3160, 1242, 1564, 606, 1264, 3287, 993, 2068, 2001, 343, 1102, 3000, 536, 1414, 2637, 3148, 3640, 420, 54, 1243, 3295, 191, 576, 618, 960, 1567, 1610, 1813, 2326, 2501, 2756, 3291, 68, 1400, 184, 1587]. data [1917, 1831, 2803, 3081, 3635, 3350, 1880, 327, 2309, 2391, 2650, 2793, 2199, 2568, 3387, 316, 1189, 2680, 3053, 2993, 3083, 395, 498, 644, 829, 1363, 1737, 1900, 2453, 3210, 3090, 2991, 2934, 1887, 2926, 3007, 1054, 2695, 1011, 1576, 1594, 1641, 2616, 2645, 3252, 456, 1150, 1578, 2459, 2673, 3147, 1903, 3219, 2930, 1258, 1455, 681, 1192, 2117, 2415, 2750, 3082]. Data-adaptive [3304, 3447, 3350]. data-based [1011]. data-driven [2573, 2934]. Data-generating [3013]. data-generation [3482]. Database [1991, 2700, 3551, 1987, 3549, 3548, 3605]. databases [3550, 3160]. dataset [565]. datasets [754, 1799, 3478, 1260, 1114]. Datta [3483]. David [1457, 289, 445]. Davidian [506]. de-identified [3307]. deal [943, 739]. Dealing [2339, 3125, 1126, 1815, 986]. death [1385, 2701, 1627, 1531, 172, 2246, 206, 967, 3023, 557, 3135, 3085, 388, 91, 1316, 3420, 3509, 1819, 1392, 1509, 2118, 228, 995, 2387, 1952, 169, 43, 1879, 2875, 2217, 1927, 794, 66]. death-censored [1819]. Deaths [908, 1874, 3574]. debate [882, 884, 881, 2251]. decade [901]. decay [1995, 2992]. decaying [3395, 1644]. Decision [2968, 3567, 1059, 1682, 1851, 938, 2415, 990, 706, 1153, 3462, 3484, 1279, 1068, 3298, 480, 455, 1145, 781, 2922]. decision-making [2415, 480]. decision-theoretic [938, 990, 1153, 3484, 1279]. Decisions [107, 3527, 551, 2754]. decline [904, 267, 2069, 1111, 3113]. decoding [2453]. decompose [3449]. Decomposition [1385, 3577, 3610]. deconvolution [482, 3183]. decreasing [1096]. deep [2925, 3196]. default [2016, 2582]. defence [1280]. deficit/hyperactivity [782]. defined [1955, 868]. Defining [2424, 486, 890]. definition [113]. definitions [1001]. degradation [1836]. degree [12]. degrees [2437, 3587, 1753, 1047]. delay [410]. Delayed [2266, 1688, 2188, 1689, 387, 274, 2242, 2606]. delays [3562]. Deletion [2214]. DeLong [838]. Delta [2388, 1843]. delta-adjusted [1843]. demand [3142]. demands [1690]. dementia [1680, 2154, 2118, 802, 66]. Demographics [802]. demonstrating [1711]. dengue [1069]. denoising [1161]. dense [3480]. densities [3492]. Density [3183, 976, 3093, 3254, 3564, 2932, 1345, 782, 2489, 3407]. density-based [782, 3407]. dental [345, 302, 730, 2233]. departed [986]. department [3527, 1203, 1973]. departments [1560]. departure [1771]. depend [941]. Dependence [3539, 2961, 1705, 128, 1969, 2928, 2598, 1567, 2326]. dependency [2748]. dependent [1517, 2204, 1934, 112, 2178, 2470, 1393, 3517, 3604, 1462, 2102, 2744, 2804, 2516, 2964, 71, 1229, 1126, 2471, 3565, 300, 1167, 270, 713, 2480, 336, 979, 1433, 2783, 224, 1990, 1504, 2503, 150, 3381, 1740, 387, 87, 813, 440, 1477, 2654, 8, 2442, 2274, 1573, 2634, 1582, 732, 3234, 3564, 1200, 3098, 2762, 2652, 2052, 1176, 2181, 3550, 3601, 826, 1346, 3335, 230, 3075, 1625, 2773, 1741, 1879, 634, 1742, 729, 2741, 1102, 1639, 2135, 3091, 1917, 2076, 2680, 3083, 1641, 537, 2673, 3143, 1733, 3525]. dependently [3147]. depending [2642]. depends [3591, 3655, 2274]. depression [3381]. Derivative [3129]. derivatives [480]. derived [2876, 761, 1797, 750, 1118, 3105]. Deriving [1160]. dermatomyositis [3056]. dermoscopic [548]. DerSimonian [134, 2160]. DeSantis [628]. description [3443, 3301]. descriptors [548]. Design [3227, 1891, 424, 3561, 1226, 3278, 1063, 2132, 722, 2161, 3500, 1005, 2227, 2242, 3356, 3072, 2925, 815, 3417, 3444, 1678, 949, 477, 1688, 3150, 1862, 3201, 1119, 1521, 1612, 2051, 2728, 1942, 3385, 3331, 2566, 2674, 798, 1500, 2200, 1208, 1396, 2009, 141, 1623, 938, 494, 3452, 3551, 108, 1473, 2272, 1269, 2421, 1910, 1700, 770, 3176, 2998, 703, 1685, 1915, 465, 1711, 3135, 297, 572, 2302, 1661, 1153, 3246, 588, 711, 106, 2103, 3186, 3244, 89, 790, 1777, 690, 1166, 2224, 2464, 2837, 2863, 3633, 1279, 918, 111, 3073, 3542, 1830, 2634, 3065, 2433, 1204, 1878, 2617, 1469, 2836, 273, 2113, 2030, 1760]. design [143, 1687, 2826, 2627, 2619, 3098, 1190, 1978, 2864, 274, 1765, 3193, 3208, 1562, 3355, 1039, 1953, 148, 62, 3627, 3549, 3548, 470, 1939, 689, 3039, 3533, 2493, 98, 928, 309, 632, 1050, 3343, 983, 2948, 2583, 468, 399, 1445, 1548, 743, 3421, 3649, 3243, 3346, 985, 789, 1244, 1437, 3124, 3490, 2568, 609, 512, 436, 2416, 2565, 3456, 47, 2373, 2494, 2963, 781, 1776, 2234, 1671, 2799, 3624, 3143, 433, 380]. design-by-treatment [1661]. designed [1504, 144]. Designing [786, 114, 466, 990, 1194, 2133, 2042, 2606, 3242, 1595, 1518]. Designs [2581, 1324, 2115, 1550, 2111, 1806, 3220, 3277, 77, 3237, 83, 1132, 1882, 180, 833, 1232, 1490, 952, 1130, 435, 1941, 2219, 1206, 660, 1869, 1299, 1957, 3245, 2167, 3342, 1848, 2002, 1805, 2753, 100, 2021, 2584, 1597, 2112, 507, 866, 723, 1399, 2604, 1023, 2304, 2550, 2044, 1670, 3584, 1962, 1702, 1689, 2849, 2633, 1549, 3306, 2166, 313, 3448, 1099, 1101, 3634, 1448, 3516, 1072, 1116, 1636, 110, 2836, 3048, 1998, 1100, 75, 509, 2923, 2447, 508, 78, 245, 492, 3545, 679, 3472, 3068, 3579, 1790, 1793, 2122, 2405, 2010, 1787, 1358, 1218, 2195, 2986, 2845, 3630, 2773, 580]. designs [13, 3211, 2008, 2723, 2545, 2114, 1792, 877, 3450, 665, 1563, 1673, 1332, 864, 3582, 1076, 2682, 1791, 2567, 3083, 1963, 2455, 3177, 3651, 2934, 3536, 3089, 1401, 1400]. detect [431, 318, 3042, 44, 3595]. Detecting [51, 2000, 1935, 3635, 710, 1985, 2031, 1097, 3577, 490, 1943, 788, 662, 655, 980, 2046, 2538, 2906, 2385]. Detection [3153, 2817, 2972, 425, 2678, 1533, 2840, 3620, 328, 2767, 548, 2519, 1502, 2200, 1897, 233, 376, 1759, 1987, 1024, 212, 961, 2914, 2608, 2643, 892, 549, 1366, 367, 1168, 1095, 1019, 599, 1261, 1175, 814, 2943, 1813, 3291, 2772, 2865, 2231, 458]. detects [261]. determinant [2245]. determinant-based [2245]. determinants [802]. determination [1309, 1608, 1820, 527, 2640, 2315, 3114, 1476, 1965, 703, 1504, 1670, 269, 220, 2713, 1638, 3032]. determined [192]. determining [3323]. deterministic [2292]. develop [3501, 1082]. Developing [2427, 3034, 2617, 786, 731, 1234, 795, 1431, 3647, 3352, 3353, 3214]. Development [2138, 2004, 2180, 1521, 1814, 872, 3155, 1088, 2662, 563, 711, 1127, 397, 1069, 200, 226, 382, 3372, 1139, 3068, 475, 196, 1253]. developmental [528]. developments [1619]. deviance [3104]. deviation [3139, 2028]. device [3217, 1950]. devices [2761, 2888, 1999]. DFS [232]. DIA [1945, 1944, 1948]. diabetes [2916, 114, 1992, 3203, 1190, 2229]. diabetic [2456]. diagnoses [1599]. Diagnosing [640]. diagnosis [3109, 2097, 3362, 333, 481]. diagnostic [1059, 1710, 2138, 1637, 2087, 2342, 2961, 1621, 236, 1534, 2571, 1698, 264, 2384, 635, 2421, 2780, 156, 2527, 2960, 3080, 1186, 1925, 2989, 688, 2306, 2498, 212, 2162, 3591, 3655, 1051, 675, 3462, 2936, 674, 161, 338, 1077, 2823, 1405, 556, 1950, 3398, 2977, 502, 612, 2642, 250, 1116, 3127, 1582, 1255, 1394, 2030, 1373, 64, 2064, 217, 137, 2686, 282, 1888, 299, 617, 2212, 727, 726, 611, 931, 1169, 301, 1145, 729, 695, 2928, 2598, 1348, 2068, 2374, 1257, 1196, 1510, 2063, 1594, 2788]. Diagnostics [1108, 3028, 1434, 2214, 711, 3232, 1724, 2703]. dialysis [3249, 1744, 1715]. Diane [380]. diary [2335]. dichotomised [1725]. Dichotomising [889]. dichotomization [223, 829]. Dichotomizing [862]. dichotomous [3587, 1141, 3446, 703, 1276, 1574, 3013, 24, 1312, 1332, 1576]. dichotomously [1314]. dietary [804, 1246, 2463]. differ [2624]. difference [3491, 335, 495, 898, 1473, 1613, 88, 1464, 1569, 1904, 496, 1614, 601, 617, 2661, 3348, 16, 304, 1106]. Differences [1220, 2179, 216, 1404, 2919, 2355, 785, 2570, 3397, 100, 1513, 1065, 1919, 245, 3207, 2054, 1856, 2639, 2622, 2538]. Different [890, 2484, 216, 1053, 1212, 1697, 2182, 2434, 846, 2437, 706, 1101, 1907, 1124, 1100, 3101, 760, 1070, 2990, 1039, 741, 2878, 550]. differential [491, 1160, 914, 2018, 3512, 1999, 553, 1918, 2486, 1432, 1508]. differentially [1163, 1902, 504]. differing [1577]. diffusion [3112]. Diggle [411]. digital [3179]. dilation [2452]. dilution [2049, 2929, 47]. Dimension [2404, 2201, 824, 1713, 1664]. dimensional [3412, 1313, 754, 699, 1466, 745, 3487, 1064, 2019, 1566, 3171, 1896, 2907, 1842, 2186, 86, 1784, 386, 1164, 2305, 2580, 2985, 3409, 3427, 3116, 2199, 2993, 966]. dimensionally [1501]. dimensions [3061]. dioxin [1272]. Direct [2701, 1556, 2940, 251, 1552, 3058, 2424, 583, 1632, 1801]. direction [538]. directional [2489, 3291]. directions [1648]. directly [3267]. Dirichlet [1987, 556, 2814, 1887]. disability [1768, 1264]. disaster [3574]. Disclosure [1308, 178, 3184]. discontinuation [1960]. discontinuations [2884, 2331, 2885]. discontinuity [1957]. discordant [1186]. discoveries [733, 788]. Discovering [2456, 692, 1151, 3179, 1163]. Discovery [203, 1901, 12, 3452, 14, 656, 3563, 2305]. Discrete [517, 81, 660, 119, 2304, 3597, 1016, 3060, 3638, 876, 2285, 1530, 2041, 1663, 3372, 2010, 2152, 2620, 2786, 1831, 334]. discrete-state [119]. Discrete-time [517, 81, 119, 876, 2041, 2010]. discretized [864]. Discriminant [317, 318, 454, 3438, 2842]. Discrimination [840, 3066, 2897, 36, 2896, 1645, 1646, 651, 760, 248, 2899]. discriminatory [679, 839, 2865]. Discussants [2131, 1948]. Discussion [3258, 1946, 145, 3261, 3549, 2129, 887, 836, 2415, 133, 2039, 3525]. Disease [1743, 2667, 120, 491, 2819, 260, 3391, 3562, 1627, 1970, 1898, 3373, 2916, 368, 1525, 122, 1964, 2323, 2820, 124, 369, 596, 3161, 2307, 635, 1470, 2437, 302, 1995, 2290, 2019, 405, 803, 2811, 1925, 3402, 2162, 3597, 992, 2481, 21, 2710, 89, 930, 2719, 2818, 3592, 1573, 1069, 1707, 3362, 1506, 127, 522, 3471, 419, 425, 642, 1079, 210, 1543, 429, 3394, 2151, 3585, 3366, 2932, 1290, 1261, 485, 1850, 2867, 2184, 3250, 121, 980, 3012, 2514, 162, 3552, 484, 2250, 2155, 2933, 3144, 1520, 2483, 2068, 2706, 2198, 1363, 1601, 693, 691, 104, 681]. disease-causing [1601]. disease-progression [1290]. disease-related [1573]. disease-specific [1627, 2068]. diseases [2364, 914, 1700, 861, 1572, 2103, 598, 908, 1949, 3507, 2409, 3003]. disjoint [405]. disorder [782]. disorders [3577]. disparities [345, 1300, 168, 1833, 2038]. disparity [3292]. dispersed [2549, 2024]. dispersion [2994, 324]. display [3105]. Dissecting [2967]. dissemination [97, 2159]. dissolution [3139, 2535]. distal [3559]. distance [1030, 1909, 1220, 1426]. distinct [191]. Distinguishability [1412]. Distributed [225, 1471, 2132, 2678, 1610]. Distribution [774, 1182, 1057, 3185, 1854, 1539, 495, 2320, 1580, 453, 603, 2801, 2317, 2019, 3612, 3402, 1424, 3426, 317, 1708, 3631, 791, 111, 2750, 3173, 496, 1184, 2064, 2864, 2136, 2078, 485, 1501, 310, 2024, 2678, 278, 1414, 2772, 1900]. Distribution-free [774, 1182, 1057, 2320, 1580, 3173, 2024, 1414]. distributional [889, 1725]. distributions [1669, 2291, 2919, 2407, 45, 2273, 355, 2572, 1825, 1405, 556, 773, 1758, 1287, 3122, 1861, 1049, 2293, 3207, 3152, 1857, 2673, 2303]. diurnal [2023]. divergence [2291]. diverse [440, 3515]. diversity [3528]. Dmitrienko [777]. DNA [2423, 2710, 3605, 1164, 54]. do [1645]. doctors [695]. Dodd [1982]. does [970, 1146, 885]. DOI [72, 459, 586, 287, 3185]. domain [3471]. dominant [3144]. donated [2518]. done [217]. Donner [175]. donors [2137]. doorknobs [859]. doRedis [3529]. dosage [3036, 2459]. Dose [105, 1396, 1206, 2683, 471, 2005, 470, 1672, 165, 3546, 2045, 112, 77, 2988, 2383, 1942, 1490, 1158, 3155, 938, 660, 685, 3473, 1001, 1848, 1910, 673, 1808, 680, 1334, 2021, 507, 514, 2404, 2576, 2577, 297, 866, 1706, 2119, 1777, 2007, 1912, 1878, 509, 511, 1931, 508, 1524, 1367, 3289, 1790, 1793, 2122, 263, 2432, 1782, 2425, 894, 1590, 2859, 554, 2883, 3049, 2583, 1558, 1792, 2580, 258, 181, 1052, 1791, 2568, 512, 2565, 2494, 3624, 513]. dose-escalation [1942]. Dose-finding [1396, 1206, 471, 470, 165, 1490, 938, 660, 1848, 1910, 1808, 680, 1334, 2021, 507, 2404, 297, 866, 509, 511, 508, 1790, 1793, 2122, 2432, 1590, 3049, 1558, 1792, 181, 1791, 2568, 512, 513]. dose-modification [685]. Dose-response [105, 2683, 2045, 1158, 3473, 2119, 263, 894, 258]. dose-time-response [1777, 2007]. dose-toxicity [514]. doses [2988, 180, 1131]. Double [352, 2351, 1550, 3507, 984, 943, 739]. Double-robust [352, 984]. double-sampling [1550]. Doubly [2838, 1084, 1442, 2390, 3286, 993, 3588, 1054, 1762, 3447, 2258, 306, 762]. doubt [535]. DR [1127, 1559]. DR-IBS [1127]. Dr. [1058]. draft [1028, 1178]. dream [719]. drinking [128]. drive [1657]. driven [3272, 2573, 3636, 1723, 2522, 2417, 470, 704, 3149, 2934]. driver [3541, 1533]. Driving [3273, 2835, 3276, 3274, 3275]. drop [1430, 180, 2594, 573, 3214]. drop-in [3214]. drop-minimum [2594]. drop-the-loser [1430, 180]. drop-the-losers [573]. dropout [3657, 3339, 1310, 2813, 1873, 384, 1551, 2672, 3464, 179, 2275, 3141, 3040, 3640]. Dropouts [770, 165, 1910, 3448, 1961, 2531, 2376]. dropping [1982, 1634]. Drug [1178, 720, 1901, 2004, 2180, 1521, 262, 1088, 1035, 3114, 227, 505, 3197, 2595, 3042, 1575, 2662, 1224, 1023, 1930, 2530, 1127, 3468, 110, 3139, 1509, 1703, 1139, 1790, 1793, 475, 779, 1326, 2942, 3563, 3343, 455, 1672, 2583, 868, 1548, 1792, 2610, 2985, 1253, 2579, 499, 3281, 1791, 2535, 2494, 2459]. drug-combination [1326]. drug-count [2985]. drug-induced [227]. drug-related [3563]. drugs [929, 3264, 1711, 711, 1298, 2923, 2255, 1519]. dual [83, 1942, 1878, 110, 344, 2579, 3586]. dual-agent [1942, 1878, 2579]. dual-energy [344]. dual-frame [3586]. Dubois [1792, 1791]. due [1386, 2259, 1841, 637, 557, 3135, 91, 1551, 2356, 3464, 1840, 3314, 228, 2212, 351, 1937, 814, 2020]. Dunnett [643]. Dunnett-type [643]. duration [2389, 1680, 1684, 2824, 1123]. during [2853, 2018, 2609, 3251, 825, 3129]. dying [678]. Dynamic [2664, 548, 1041, 3155, 2842, 948, 3031, 3592, 1392, 1157, 3564, 3120, 2065, 2029, 2745, 3428, 1044, 3014, 2628, 3388, 41, 1572, 2037, 2273, 3085, 925, 1589, 3191, 2226, 523, 3643, 1416, 664, 3585, 1949, 677, 583, 3598, 2450, 1482, 1046, 1045, 1520, 3116, 2513]. dynamics [424, 582, 1248, 825, 811]. dysphagia [1712]. dysregulation [782]. dystrophy [962]. each [1136]. Early [107, 858, 2576, 1082, 3546, 2456, 1619, 3620, 2006, 635, 1560, 466, 3473, 1759, 1216, 674, 3628, 1069, 419, 2578, 848, 591, 1357, 2195, 370, 2668, 3434, 789, 859]. early-phase [2578]. EARS [370]. East [511]. Easy [2279, 252]. easy-to-implement [252]. eBook [2725]. Ecological [375, 3310, 138, 910, 2618, 1835, 896, 3082, 127, 452, 2933]. Ecological-type [375]. Economic [2765, 3363, 2443]. Edgewise [2499]. edition [2725]. Editor [2033, 208, 1497, 557, 1559, 150, 2669, 1058, 727, 175, 253]. Editorial [320, 2560, 2890]. Editors [1403, 1774]. edn [118, 289, 380]. eds [411]. educational [1165]. Edwards [194]. EEG [3641]. Effect [3610, 3368, 997, 729, 3285, 1062, 1112, 1651, 3451, 1053, 1496, 3067, 2973, 3035, 1858, 2401, 1197, 3371, 1618, 3213, 923, 1748, 1514, 3429, 1463, 2209, 937, 3178, 2919, 208, 1396, 1767, 2246, 697, 2454, 1450, 3221, 3587, 270, 2147, 1384, 2999, 44, 1924, 1074, 1331, 1605, 588, 3584, 254, 3194, 490, 2958, 1640, 1943, 3241, 1360, 3057, 1652, 421, 1591, 1600, 1418, 2169, 2281, 2258, 2693, 553, 2285, 48, 1203, 3589, 2735, 2800, 464, 3166, 1022, 2064, 1675, 416, 3337, 2228, 1190, 2153, 3061, 1070, 2724, 4, 1487, 2990, 1773, 2417, 1913, 535, 3228, 1114, 2179, 1604, 175]. effect [2242, 3075, 530, 251, 708, 632, 3286, 1422, 2828, 3280, 2668, 3434, 2351, 2496, 2400, 2985, 809, 794, 2301, 1102, 3024, 1813, 2606, 3242, 1917, 2622, 3301, 3149, 3543, 2399, 1054, 7, 176]. effect-measure [1054]. effect-modifying [421]. Effective [419, 2595, 1706, 1284, 2513]. effectively [3210]. Effectiveness [202, 3004, 2941, 1541, 2422, 3279, 2308, 170, 3477, 882, 2277, 1869, 197, 1495, 1138, 884, 885, 1999, 198, 881, 2511, 1581, 2485, 200, 1579, 1842, 201, 532, 144, 196, 1653, 2076]. Effects [1854, 2952, 1739, 787, 720, 654, 1057, 3185, 2053, 2630, 2701, 2590, 2976, 1361, 3150, 2853, 2349, 1816, 1993, 3050, 1535, 483, 298, 39, 2235, 2851, 1304, 1580, 2873, 1674, 1686, 2215, 1778, 2431, 1488, 1497, 3255, 637, 1770, 1526, 685, 2264, 2214, 1001, 2272, 2018, 1259, 562, 1141, 2034, 2956, 1478, 336, 910, 290, 1680, 413, 2436, 3216, 238, 2553, 3015, 3156, 586, 3062, 134, 1661, 2160, 2848, 3010, 2503, 3106, 2316, 3246, 1874, 1091, 1835, 473, 3182, 183, 766, 3288, 1626, 317, 2665, 1322, 1481, 130, 1143, 977, 2095, 3420, 187, 773, 896, 1540, 2850, 3249, 29, 3082, 2134]. effects [3058, 965, 2285, 3636, 3449, 1918, 2486, 3302, 2424, 1769, 1602, 2750, 3065, 2726, 3173, 3054, 450, 1144, 610, 396, 3358, 2831, 2743, 662, 750, 1108, 2137, 1184, 3101, 722, 2064, 616, 1075, 1305, 446, 2725, 245, 2515, 3324, 2585, 3013, 429, 340, 1416, 2510, 2734, 2084, 331, 1070, 159, 1133, 1562, 3607, 1853, 2625, 357, 263, 583, 287, 2010, 237, 3393, 604, 3482, 1221, 272, 3632, 2996, 3315, 894, 1288, 1632, 129, 133, 1498, 3630, 1529, 251, 296, 3111, 3615, 1801, 2713, 3338, 1659, 486, 132, 278, 734, 2229, 115, 3552, 131, 743, 891, 695, 1312, 1617, 2049, 2496]. effects [2598, 3614, 3489, 3160, 2369, 993, 2068, 1639, 2135, 3078, 420, 1257, 1567, 2326, 2756, 3453, 1831, 497, 1189, 550, 3625, 2085, 537, 104, 2185, 3407, 2712]. efficacy [391, 2911, 2729, 477, 2225, 3473, 108, 2403, 3571, 2527, 673, 1711, 297, 1031, 2863, 2274, 110, 1215, 1782, 470, 2008, 3330, 3125, 868, 499, 1596, 960, 1655, 1911, 950]. efficacy/tolerability [297]. efficiencies [3389].

Efficiency [3213, 22, 1116, 672, 2259, 3331, 1130, 2272, 3245, 3265, 2103, 534, 1154, 2986, 712, 283, 1076, 2682, 3383]. Efficient [833, 744, 1899, 2142, 800, 1026, 2449, 2935, 2969, 1161, 2141, 865, 3143, 1858, 2674, 1869, 3623, 1473, 1099, 1183, 784, 1760, 3100, 1356, 3240, 3228, 62, 655, 1365, 3614, 1332, 2954]. Efforts [188]. Efron [2059]. Egger [2918]. EHR [3268, 3470]. EHR-based [3268]. either [2356, 2620]. EJ [194]. Elastic [395, 3573]. elderly [3126]. electroencephalograms [1248]. electroencephalography [577]. electronic [2820, 2760, 3240, 1666]. elevated [3118]. elicit [532]. Eliminating [377]. elliptical [1022]. Eloise [1440]. Elst [2846]. Elston [584, 3]. EMA [1127, 2548, 2255]. Emanuel [445]. EMAX [3473, 2056]. embryonic [1884]. emergence [153]. Emergency [1550, 3527, 1203, 1973]. emerging [2004, 1711, 3003]. eminence [2270]. EMLasso [1233]. empiric [785]. Empirical [1430, 1085, 2858, 1758, 261, 1174, 1148, 1133, 994, 1086, 129, 1498, 2305, 332, 1528, 3225, 2415, 2413, 791, 3464, 539, 656, 3113, 2500, 1419, 2414, 3521, 1511, 782, 3295, 3407, 1497, 130, 287, 133, 132, 131]. empirical-likelihood [3295]. employing [2030]. emptying [183]. Emulating [3428]. encrypted [2089]. end [949, 44, 569, 570, 2892]. endogeneity [279]. endometriosis [3361]. endpoint [3285, 2973, 3035, 463, 2787, 3485, 963, 547, 435, 1206, 3229, 1060, 706, 1061, 1194, 2633, 1322, 3628, 347, 2912, 3619, 273, 464, 2846, 3046, 1367, 2990, 98, 2014, 2418, 422, 2455, 781]. endpoints [3238, 639, 2106, 1608, 3345, 335, 638, 1618, 2180, 2365, 2891, 1923, 2596, 874, 1817, 2809, 279, 27, 2987, 673, 2753, 1992, 1092, 5, 348, 1640, 207, 991, 1549, 1037, 3073, 1723, 2522, 758, 2041, 2132, 3617, 3472, 3585, 1268, 700, 2314, 3396, 2010, 3386, 1736, 2107, 655, 779, 3533, 2893, 3461, 220, 796, 797, 2775, 3594, 3644, 3552, 805, 56, 1673, 926, 3158, 1089, 1893, 2799, 2197]. energy [344, 3518]. engagement [2957]. England [2679]. English [689]. English-language [689]. enhanced [2672, 1482]. Enhancing [817, 114, 2040]. enough [55]. Enriched [679, 1612, 3542, 350, 2773, 3243]. enrichment [3237, 3151, 2849, 3306, 3634, 3633, 2984, 1204, 1723, 2522, 3472, 978, 2405, 2861, 3533, 3165, 1596, 2567, 2934]. enrollment [2302, 2088, 1963, 2219]. Ensemble [1799, 3494, 3382, 2953, 3400, 3453, 2309, 2954]. ensembles [764]. Ensor [3647]. entangled [1137]. entering [399]. enterocolitis [2679]. entire [2359]. entropy [3238, 758]. entropy-based [758]. entry [2188, 93, 387, 3001, 1223]. environment [1112, 1452, 833, 899, 2844, 836, 405, 2817, 2972, 252, 834, 1019, 2773, 1875, 3595, 3377, 1757, 2074, 2967, 3580, 3350, 1425]. environmental [368, 372, 909, 1259, 1752, 21, 2350, 3375, 1176, 3319]. enzyme [1306]. enzyme-linked [1306]. epidemic [1492, 1452, 1678, 2952, 2055, 1149, 2792, 3508, 2151]. epidemics [1069, 1201, 2513]. epidemiologic [823, 2339]. Epidemiological [569, 2100, 2119, 748, 1751, 2938, 570, 568, 686, 2152, 144, 708, 1522]. epidemiology [2048, 2236, 2811, 2240, 1289, 1557, 1275, 2538, 2599]. epidermal [2456, 2962, 575]. Epigenetic [1175]. epilepsy [1189]. episodes [298, 39]. episodic [3603, 159]. epistasis [819]. Epitope [1667, 2948]. EQ [164]. eQTL [1042]. equal [3145]. equality [2433, 707, 2054]. equally [3622]. equation [3008, 1219, 1160, 186, 1897, 2025, 1356, 2448, 1761, 1453, 658]. Equations [2521, 3329, 597, 1895, 1325, 3225, 2018, 2245, 3005, 2289, 247, 2376, 1800, 712, 52, 1717, 2656, 441, 2232, 2718, 1213, 1564, 2402, 356, 2520]. Equilibrium [2502, 833, 326, 235, 252]. Equivalence [454, 2261, 768, 808, 214, 2381, 437, 2325, 2574, 767]. equivalent [3105]. ERG [3113]. ERP [14]. Erratum [298, 381, 119, 382, 357, 1152]. Error [2089, 3074, 996, 1854, 920, 3213, 1404, 3445, 2423, 2315, 1158, 2200, 2048, 2236, 988, 1191, 476, 2421, 526, 1553, 3216, 816, 940, 2044, 154, 1289, 1557, 1958, 3413, 31, 3506, 234, 249, 1013, 2660, 2808, 3493, 3581, 3048, 259, 3166, 1063, 371, 3617, 1876, 2554, 3183, 3025, 3251, 2762, 3351, 3472, 86, 2531, 2293, 1989, 2807, 2942, 976, 1435, 1720, 3518, 1749, 951, 2230, 1348, 3582, 3291, 2046, 3149, 2257, 1328]. error-based [86]. error-prone [2421, 1328]. Errors [3510, 1291, 2970, 3205, 2287, 156, 2527, 582, 3303, 3062, 42, 127, 443, 1453, 1717, 831, 729, 3586, 94, 2943, 104]. Errors-in-variables [3510, 2970]. Escalation [1326, 1942, 1878, 471, 1672, 2741]. esophageal [263]. Essentials [2591]. establish [2620]. established [3573, 3043]. Establishing [81, 1038, 553]. Establishment [142]. establishments [3337]. estimand [3619, 3589]. Estimands [2561, 2560, 3189, 3658, 2884, 3136, 2334, 3603, 2331, 2885, 3525, 2563, 2562, 2564]. Estimate [1988, 391, 2053, 3371, 1637, 1627, 1514, 2952, 2349, 2388, 3010, 1504, 1441, 1065, 2665, 913, 1600, 1833, 876, 1366, 3235, 2028, 274, 2153, 1263, 535, 1440, 3330, 959, 794, 3543, 2386]. estimated [3329, 135, 1349, 2291, 1409, 2656, 2668, 3434]. estimates [444, 3380, 1304, 12, 905, 427, 981, 1187, 1984, 1588, 2724, 1413, 2179, 304]. Estimating [3074, 1386, 335, 3391, 1393, 2729, 1449, 6, 2308, 702, 715, 2471, 935, 2367, 3109, 1167, 752, 2067, 2956, 149, 2443, 415, 514, 730, 588, 2707, 675, 473, 3288, 91, 1921, 168, 88, 733, 2719, 2654, 3196, 2134, 2642, 1203, 2536, 385, 2452, 3101, 1075, 500, 416, 3251, 1190, 2734, 3399, 1725, 1493, 1352, 443, 3537, 3317, 3315, 1288, 3600, 3075, 3615, 310, 2482, 1719, 115, 2521, 2496, 2409, 1094, 2369, 3294, 1655, 681, 491, 2213, 216, 1212, 3067, 3329, 488, 597, 1895, 1580, 1325, 2919, 1247, 697, 3225, 670, 270, 3087, 2245, 3005, 1107, 1143, 1391, 3231, 1656, 2258, 3191, 2424]. estimating [907, 1356, 2289, 17, 3337, 3256, 1519, 329, 2441, 247, 2376, 2661, 3393, 2448, 2450, 1761, 1800, 712, 52, 3111, 1453, 1422, 3518, 811, 2583, 278, 2656, 1951, 441, 2232, 2718, 167, 1213, 1564, 2402, 2058, 3148, 821, 94, 1257, 3611, 2085, 957, 2520, 1635, 194]. Estimation [810, 2840, 923, 2787, 298, 39, 3168, 1698, 899, 2215, 2355, 3544, 206, 2034, 2612, 1384, 3402, 1401, 1399, 1424, 1976, 3526, 991, 89, 3057, 31, 2816, 3239, 1418, 1067, 2693, 48, 3192, 2549, 2030, 642, 1184, 2652, 2716, 677, 61, 229, 1632, 632, 2393, 831, 1927, 439, 891, 1312, 3425, 3586, 814, 960, 1567, 1813, 2326, 3453, 2772, 1880, 2673, 1096, 3004, 1651, 1550, 1214, 2263, 3370, 2551, 1492, 3568, 3443, 1858, 2401, 3008, 3027, 2955, 2590, 1430, 3373, 2072, 3307, 1337, 3429, 403, 1132, 744, 1351, 1025, 262, 3043, 937, 2516, 2640, 3227, 124, 2055, 369, 1296, 1055, 2420, 241]. estimation [3019, 3333, 2838, 914, 2771, 3637, 637, 1770, 1165, 1914, 1908, 1018, 3221, 3623, 1473, 1495, 2761, 2888, 2167, 3197, 2966, 1764, 1141, 1681, 2002, 1209, 3097, 2165, 2618, 2702, 3389, 119, 3062, 1331, 212, 723, 438, 2546, 2360, 2103, 3656, 1216, 3151, 674, 2849, 2633, 2119, 2154, 1780, 1652, 2358, 3442, 249, 1758, 1477, 2850, 2855, 1821, 2931, 2821, 2726, 3110, 3254, 3051, 2056, 2587, 3189, 3203, 213, 549, 259, 1251, 2800, 3175, 2113, 534, 3508, 2826, 573, 3183, 623, 3102, 1305, 2228, 2585, 1676, 3351, 2084, 3061, 1084, 1442, 2531, 2932, 3068, 2698, 1739, 2417, 2181, 324]. estimation [1415, 1989, 3228, 997, 1397, 1270, 3598, 727, 2487, 3129, 726, 1482, 712, 976, 257, 590, 797, 554, 627, 3286, 352, 1338, 1659, 3012, 1717, 1365, 1432, 2217, 2489, 762, 743, 493, 3450, 3160, 809, 1242, 1278, 1881, 765, 2068, 2805, 864, 2135, 1755, 1657, 2756, 3091, 3575, 327, 2391, 1479, 1189, 2279, 458, 1721, 2399, 1451, 1054, 1671, 3488, 1733, 1400]. estimator [2911, 147, 1762, 1098, 1704, 2669, 306, 750, 1133, 3521, 276, 3164, 441, 2351, 2279, 3115]. Estimators [2708, 79, 2178, 3213, 483, 963, 3447, 3169, 3225, 2683, 888, 377, 1811, 1872, 3558, 255, 3368, 245, 2874, 256, 839, 2248, 2709, 2713, 157, 2232, 2521, 1213, 993, 765, 618, 984, 533, 2520, 3115, 3076]. estrogen [2636]. ethical [571]. ethics [971, 1279, 445]. Ethnic [1600, 1907]. etiologic [1369, 2857]. etiologies [1850]. eUCI [2089]. EUR26.40 [289]. European [2923]. evaluate [2042, 387, 2824, 3405, 839, 2757]. Evaluating [236, 1705, 661, 156, 3397, 1277, 1922, 3628, 690, 3431, 2274, 3127, 2687, 1883, 237, 524, 410, 3270, 242, 1196, 3201, 1521, 1234, 3014, 3344, 3275, 766, 161, 338, 234, 795, 1573, 651, 1268, 3605, 1801, 2881, 428, 2099, 3076, 1348, 943, 739, 163]. Evaluation [2147, 2467, 1932, 569, 3038, 2515, 1154, 2497, 570, 568, 1093, 1666, 1911, 3238, 731, 1550, 1697, 2364, 1971, 3485, 761, 929, 2873, 2765, 3522, 935, 424, 673, 1645, 1032, 2498, 1298, 593, 3060, 14, 696, 1271, 2224, 2837, 1255, 2113, 75, 722, 684, 2846, 2186, 2052, 700, 357, 265, 843, 386, 2418, 1741, 518, 789, 2934, 2197, 686]. evaluations [3363, 2443]. even [90]. Event [1008, 2899, 165, 654, 266, 1496, 3417, 2347, 912, 2664, 3199, 1393, 3517, 2182, 2389, 2787, 2469, 3307, 3495, 1993, 2102, 2839, 3178, 2055, 71, 1923, 2870, 1490, 952, 2813, 2225, 1988, 1300, 435, 2437, 2388, 1817, 2809, 1476, 2547, 2570, 3118, 1417, 2421, 888, 1986, 2998, 3097, 582, 2553, 2273, 1061, 3561, 2546, 3553, 1226, 1549, 1016, 3006, 3001, 1837, 2348, 2818, 3592, 2511, 2920, 3541, 3636, 1582, 3589, 1530, 213, 732, 3175, 1723, 2522, 2452, 367, 1168, 3025, 1200, 2895, 515, 1268, 2990, 1406, 700, 3647, 3353, 451, 2754, 997, 15, 594, 1152, 839, 2152, 1397, 2375, 1080, 2448]. event [2295, 295, 3039, 3075, 530, 1006, 296, 3022, 3282, 2659, 525, 43, 2827, 422, 3594, 3297, 2722, 1009, 704, 1997, 2001, 3148, 94, 1181, 960, 1567, 2326, 2756, 3158, 3124, 3053, 561, 1852, 2616, 1150]. event-dependent [71]. event-driven [704]. event-related [952]. event-reporting [704]. event-time [3178]. Event-weighted [1008]. events [731, 2438, 2347, 211, 3495, 71, 1674, 2813, 3465, 2454, 2615, 2341, 3561, 2503, 1867, 3539, 2609, 3202, 3619, 3643, 2186, 2724, 2276, 305, 2275, 1715, 1632, 2690, 3359, 309, 3563, 3593, 1577, 3645, 3124, 1081, 2680, 2455, 561, 1664, 589]. events-add [1867]. every [698]. Evidence [2270, 2998, 2649, 138, 3559, 154, 1395, 3433, 3212, 2629, 199, 3311, 1128, 1179, 3436, 455, 3357, 2229, 3437]. evidence-based [455]. evolution [2949, 1090, 1930, 2530, 591]. evolving [3508]. Ewing [939]. EWMA [50]. Ewout [1592]. Exact [2059, 1299, 1986, 1708, 2169, 2281, 404, 1615, 2235, 1158, 715, 2924, 303, 1178, 655]. exam [2090]. Examination [2512, 344, 903, 57, 121]. examination-based [57]. examine [388, 564]. examined [298, 39]. Examining [3129, 1452, 1044, 2534, 2088, 2708, 1046, 1045, 2622]. example [345, 547, 905, 649, 3256, 1176]. examples [613, 614]. Excellence [2040]. excess [996, 2349, 1423, 981, 376, 2959, 1716, 3094, 3071, 2698, 3574, 1883, 3317]. excessive [1125]. Exchange [2077]. exchangeability [3176]. excited [3410]. exciting [3273, 1043]. exclusion [3457]. excursion [577]. executive [871]. Exemplary [18]. exercise [3126]. existing [2644, 2815]. exit [93, 1223]. expanding [3155, 2740]. expansion [2577]. expect [1756]. expectancies [443]. expectancy [1768]. Expectation [2645, 1386, 552, 818]. expectations [1662]. Expected [1453, 2187, 2443, 3170, 1187]. Expenditure [3117, 2164, 2310, 3518]. expenditures [1973]. experiences [467, 788, 877]. experiment [2925, 325, 1703, 3281]. experiment-specific [42]. experimental [147, 1323, 1830, 601, 3627, 1936]. experiments [3658, 1496, 2031, 262, 963, 952, 1488, 3136, 1085, 2304, 538, 1912, 2836, 2697, 1633, 994, 2631, 1859, 433]. expert [1901, 646, 532]. explain [1123]. Explained [841, 3041, 3094, 842, 2436, 2163, 2228]. explanatory [634]. exploiting [2844]. exploration [2017, 3000]. exploratory [3534, 2732, 1194, 1163, 983, 1936]. explore [3155, 1114, 3292]. Exploring [435, 3188, 2197, 936, 947, 1712, 2229]. exponent [182]. Exponential [2992, 2770, 1408, 746]. exponentially [50]. exponentials [1123]. exponentiated [1669]. Exposome [837]. Exposure [2240, 3564, 2773, 3112, 720, 996, 1350, 23, 920, 1272, 3168, 101, 3205, 2454, 1471, 2019, 19, 3015, 1921, 1758, 3271, 3247, 1769, 830, 2536, 3610, 1063, 3351, 2708, 1718, 49, 1868, 257, 1902, 352, 2827, 2496, 2280, 167, 2321, 828, 3287, 993, 1275, 2772, 1328]. exposure-attributable [3168]. exposure-induced [3610]. exposure-lag-response [1471]. exposure-outcome [352, 3287]. Exposure-wide [2240]. exposures [3067, 815, 3517, 3477, 909, 336, 87, 2866, 125, 1875, 2230, 2933, 3320, 3525]. expressed [1163, 504]. Expressing [587]. expression [431, 3424, 325, 85, 1591, 3230, 620, 1303, 2814, 54, 710]. expression-based [3230]. Extended [555, 508, 747, 1740, 1003, 3632, 3267]. Extending [1062, 840, 46, 1300, 1597, 134, 2160, 1704, 2669, 1600, 2918, 2294, 2725, 2691]. Extension [323, 863, 1013, 1833, 2831, 477, 2924, 908, 2008, 115, 3554]. Extensions [3150, 2743, 322, 796, 797, 1741, 2305, 647]. external [1439, 3043, 2098, 3637, 1592, 3357, 2435]. externally [1079]. extra [2588]. extraction [534, 1273]. extrapolation [1844, 2270, 892, 3311]. Extreme [3409, 3220, 128, 1759, 1825, 1787]. Ezekiel [445].

F [150, 1265, 1792, 1791]. Factor [3515, 555, 1313, 3534, 3055, 3422, 1968, 802, 439, 3024]. Factorial [2604, 2765, 1023, 2304, 1336, 2936]. Factors [281, 1053, 2182, 1731, 1988, 849, 3275, 888, 1695, 421, 2356, 3249, 450, 2697, 1155, 1368, 693, 681]. Fagerland [3059]. failing [1798]. Failure [3234, 728, 308, 2072, 2804, 1895, 2964, 2937, 2200, 1478, 1990, 818, 3578, 1729, 2672, 2852, 99, 1446, 1237, 3372, 3223, 3324, 3167, 1352, 3461, 3531, 3467, 3338, 747, 2991, 2673, 3147]. failure-time [3461]. Fairclough [380]. fall [373]. fallacy [2933]. Fallback [2314]. False [14, 3452, 3403, 788, 478, 3563, 2305]. familial [2979, 1391, 1258]. families [236, 1449, 1783]. family [1863, 2676, 2323, 284, 2315, 2817, 2710, 2972, 2746, 2166, 3506, 1013, 3173, 182, 521, 2723, 493, 2502, 1332, 316, 1425]. family-based [284, 2817, 2972, 2166, 521, 2502, 1332, 1425]. family-wise [2315, 1013]. Familywise [3506]. Faraway [2725]. Fast [367, 1168, 1012, 1107, 3337, 515]. fast-food [3337]. fatality [3656, 3442]. fatigue [637]. favorable [1219]. fcMRI [2136]. FDA [1028, 1987, 2255, 2333]. FDA/CDER [2333]. FDG [2646]. FDG-PET [2646]. feasibility [2741]. feasible [2424]. Feature [1273, 3577, 2969]. features [2553, 3364, 3463, 3592, 2088, 3129]. Feldman [1442]. femoroacetabular [2667]. fertilization [1530]. fetal [1219, 2190]. fetuses [226, 382]. Feuer [194]. few [495, 2466, 48, 496, 2186, 3537]. fiber [2456, 2962]. fibers [575]. fibre [1314]. fibrillation [2018]. fibrosis [2262, 954]. fiducial [2087]. field [529, 964]. fields [2685, 21]. file [2089]. Files [443]. Filipino [1219]. fillings [730]. filtering [3544]. find [2194]. Finding [2675, 2309, 165, 112, 77, 1490, 1396, 1206, 938, 660, 1869, 1848, 1910, 1808, 680, 1334, 2021, 507, 2404, 2576, 297, 866, 471, 509, 511, 508, 2005, 1524, 1790, 1793, 2122, 2432, 2425, 470, 1590, 2733, 2883, 3049, 1672, 1558, 1792, 2580, 181, 1052, 1791, 2568, 512, 513]. findings [587, 3161, 3436]. fine [3026, 2077, 2882]. fine-tuning [3026]. Finite [2317, 682, 3466, 3181, 1565, 3256, 171, 3141, 801]. finite-sample [801]. Finland [1468]. First [2901, 3199, 2389, 298, 39, 1560, 1268, 2990]. Firth [2806, 2724]. Fisher [2924, 3601]. fit [2589, 2976, 3070, 1906, 1171, 2473, 1629, 3241, 885, 795, 2150, 820, 1004, 933, 2699, 1414, 2501, 1737, 1286]. Fitting [38, 2340, 1339, 523, 328, 563, 2095, 1972, 3240, 550, 2645]. Fitzmaurice [506]. Five [2973, 3035, 2111, 2113, 2112, 2115, 2114]. fixed [375, 562, 3062, 3106, 3463, 3486, 2071, 601, 1358, 1940, 2301]. fixed-effect [2301]. FL [3059]. Flagging [788]. flatness [2489]. fleets [3529]. flexibility [864]. Flexible [243, 1806, 2073, 998, 2919, 1021, 336, 465, 3340, 2663, 3051, 1636, 450, 1039, 2014, 3165, 493, 1474, 2135, 1386, 2664, 2470, 1318, 2164, 2534, 2380, 2596, 3616, 494, 233, 1766, 174, 3097, 723, 3005, 3526, 3253, 1943, 2746, 3511, 1284, 2940, 193, 573, 215, 2032, 480, 2441, 2189, 980, 1859, 1193, 2373, 3117]. Flexibly [1002]. Florida [3422]. flow [1395]. fMRI [662, 3123, 2309, 2453]. focus [1064, 1557, 1069]. Focused [3026]. fold [1666]. follicular [732]. follow [141, 44, 1032, 2341, 1030, 1999, 1921, 3578, 1225, 443, 1413, 2786, 3645, 1033, 1655, 1034, 2673]. follow-on [1032, 1030, 1033, 1034]. follow-up [141, 44, 2341, 1999, 1921, 3578, 1225, 443, 1413, 2786, 3645, 1655, 2673]. following [1310, 303, 3151, 2585, 2678]. follows [1749]. Food [1178, 1452, 2360, 3337]. forcing [523]. Forecasting [3410, 2802, 2811, 2350, 1694]. forecasts [967, 1124]. forensic [1914]. forest [3303, 1230, 1850]. forests [2420, 3111, 2653]. Forgotten [3258, 3260, 3261, 3257, 3259]. Forlag [1522]. form [3055, 3050, 2713]. forming [2976]. forms [2497]. Formula [3569, 806, 1312, 794, 1480]. formulas [2315, 957]. formulations [736]. forward [1841, 1149, 3375]. Four [3498]. Fowler [1045]. Fractal [1979]. fraction [1531, 89, 65, 908, 3193, 272, 3531, 1567, 2326, 3525]. Fractional [1411, 2216, 1173, 1954, 1023, 2304, 2027, 125, 747]. fractions [3168]. fracture [1836]. Fractures [1836]. frailties [3232, 2139]. Frailty [566, 516, 3596, 1952, 939, 810, 2438, 1627, 2308, 3570, 3041, 1728, 643, 818, 3085, 65, 2609, 1354, 1392, 732, 307, 95, 2275, 3283, 2483, 32, 960]. frame [3586, 747]. Framework [2032, 1369, 1898, 2688, 1820, 2812, 1779, 1234, 2380, 670, 1905, 673, 1805, 1803, 1074, 280, 1611, 2957, 3484, 2731, 2408, 1955, 2038, 652, 1268, 3632, 296, 1050, 2825, 1937, 2049, 1332, 395]. Framingham [760, 49]. Franklin [445]. free [1057, 2320, 1580, 1768, 172, 1018, 2319, 1969, 3173, 3556, 774, 2418, 2024, 3313, 1414, 1755, 1182, 3053, 1847, 2020, 3072]. freedom [1753]. French [2004, 461, 996]. frequencies [2547, 3042, 1107]. frequency [3069, 1302, 1278]. Frequentist [1877, 313, 3536, 798]. frequently [761]. Fridge [3026]. Friede [347]. friendship [416]. front [705]. frontier [1043]. Fuentes [411]. full [2070, 2006, 1698, 635, 2339, 3363, 1301, 1788]. full-cohort [1301]. fully [1003, 256, 3099, 1666, 1533]. function [1929, 2829, 321, 2471, 2771, 3637, 988, 1191, 3599, 1722, 1515, 2663, 2816, 3077, 1237, 2262, 2940, 3183, 792, 3472, 3068, 727, 726, 1933, 3518, 493, 1242, 2805, 232, 561]. Functional [3121, 1554, 2818, 1744, 996, 898, 952, 1470, 1754, 2720, 2067, 430, 692, 1626, 2608, 3031, 2069, 2981, 3030, 3641, 894, 1382, 576, 3291, 1011]. functionals [724]. functions [38, 84, 1318, 105, 2696, 1869, 953, 703, 3547, 2880, 1220, 791, 523, 1569, 425, 215, 1724, 760, 878, 2699, 2054, 959, 809, 327, 1664]. Funnel [1485, 1662]. fused [3412]. fusion [368, 2472, 3348]. Fuster [727]. futility [1216, 674, 3628, 2668]. future [887, 882, 884, 881, 2175, 1000, 1648, 3033]. FWER [1136, 2071].

G [150, 1320, 1793, 1889, 3595]. G. [175]. gain [1977]. Gaining [2210]. gallstone [1662]. gamma [1463, 1669, 3558, 2824, 2864, 2677]. gamma-based [2677]. Ganju [767]. gap [1008, 349, 3011]. gaps [3359]. Garrett [506]. gastric [183]. gatekeeping [462, 1503]. Gauss [1674]. Gauss-Hermite [1674]. Gaussian [996, 2201, 2873, 3296, 1209, 2044, 3597, 1758, 1540, 3037, 894, 1501, 162, 2632, 626]. Gaussian-based [626]. GBP [520, 2591, 2846, 2691]. GEE [2971, 2232, 2521, 2520, 3308, 2821, 3435, 1676, 2637]. Geert [2846, 506]. Geertruida [3651]. GEEs [2376]. Gelfand [411]. gender [3565]. gender-based [3565]. Gene [3423, 12, 431, 833, 899, 2844, 836, 3424, 405, 85, 2817, 2972, 252, 1591, 3230, 2984, 620, 834, 1019, 2814, 2773, 1875, 3595, 734, 1753, 3377, 1757, 2074, 2967, 2980, 3580, 54, 3116, 504, 3350, 1425]. Gene-based [3423]. gene-environment [833, 899, 2844, 836, 405, 2817, 2972, 252, 834, 1019, 2773, 1875, 3377, 1757, 2074, 2967, 3580, 3350, 1425]. gene-gene [836, 834, 1753, 2980]. gene-set [734]. General [732, 3608, 2630, 1370, 3277, 1317, 1895, 1779, 670, 3344, 1065, 2731, 2408, 385, 3401, 999, 1524, 86, 3355, 175, 933, 928, 1050, 3595, 1365, 3076, 2049, 3146, 1332, 1503, 7, 176]. general/internal [999]. generalisation [3029]. generalised [2214]. Generalizability [3131, 2671, 1441, 1263, 1440, 2229]. Generalization [1079, 285]. Generalized [315, 697, 925, 2920, 414, 1800, 3644, 1849, 2521, 2943, 644, 3067, 3329, 1104, 1337, 597, 1103, 1463, 1839, 1669, 1686, 2315, 2431, 3225, 2310, 2011, 2245, 1447, 3005, 1920, 1730, 1077, 3239, 503, 795, 3173, 549, 1251, 2743, 750, 1329, 1356, 1184, 2289, 2725, 2864, 3126, 2071, 247, 2376, 3641, 1221, 265, 1761, 712, 296, 1453, 2656, 1749, 441, 2232, 2230, 2718, 1213, 1564, 2402, 2231, 550, 1129, 2695, 2645, 2520]. generated [2855, 1137]. Generating [955, 1071, 3013, 1087]. generation [1433, 3482]. generic [1805]. generics [2776]. genes [2031, 1231, 1536, 1163, 1533]. Genetic [3190, 316, 3220, 135, 236, 2323, 1535, 1041, 836, 566, 1259, 2601, 22, 1909, 2816, 3017, 3478, 396, 1467, 2782, 3138, 868, 2502, 2322, 880, 2440]. genome [284, 819, 494, 235, 1536, 733, 1356, 2060, 717, 1012, 1800, 521, 3210]. genome-wide [284, 819, 494, 235, 1536, 733, 1356, 2060, 717, 1012, 1800, 521, 3210]. genomic [1803, 2324, 3222, 1260]. genomics [3121, 1544]. genotype [1107, 3210]. genotypes [2816]. genotypic [2907]. Genshiro [460]. geo [204]. geo-referenced [204]. geocoding [104]. geographic [240]. geographical [485]. geometric [785, 1657, 537]. geometry [1979]. Germany [3142, 2154, 1520]. gestational [3074, 1219, 3499, 3251, 171, 1162]. gets [2454]. ghosts [986]. Gibbs [783, 2136]. Gill [298, 39, 3593]. GIS [19]. given [2819, 2451, 3206, 635, 2800, 2479, 1422]. glaucoma [547, 1554]. glaucomatous [2685]. GLM [19]. Global [699, 3487, 84, 2319, 738, 1316, 1461, 662, 577]. glucose [523, 607]. glycemic [3616]. GMM [2310, 3411, 1740]. GNE [3366]. Go [107, 706, 247]. Go/No [107, 706]. gold [1710, 2073, 1147, 264, 212, 161, 338, 1582, 2030, 2826, 2627, 192, 3355, 1358, 3627, 928, 695, 3425, 1257, 1196, 481, 1306, 2203]. gone [3524]. good [55, 2009, 2013, 2295]. Goodness [2589, 3070, 3241, 1286, 1906, 1171, 2473, 1629, 2150, 820, 1004, 933, 2699]. Goodness-of-fit [2589, 3070, 1286, 1906, 1171, 2473, 2699]. Gould [1946]. grade [1475]. graded [1396]. gradient [3615]. grading [509]. Grady [445]. Graf [3569]. Granger [3069]. Granger-causality [3069]. Grant [969]. Graph [2948, 2770, 3272, 2177]. Graph-based [2948]. graphic [2009].

Graphical [1439, 2342, 2426, 1329, 1321, 1319, 3443, 463, 3424, 2457, 3044, 1783, 2150, 1320, 2136, 2168, 1546, 1592]. Graphics [567, 749]. Graphing [3266]. graphs [1137, 3622]. gravity [2791]. Gray [2882]. Greenwood [1560, 2146, 2544]. Gregg [3483]. Group [3142, 2196, 1204, 2405, 1358, 2929, 864, 3047, 116, 1561, 1608, 1971, 3255, 1897, 1299, 27, 1944, 1948, 3209, 2595, 1965, 3599, 1331, 869, 3378, 3215, 2224, 2837, 3572, 1072, 3483, 1183, 273, 78, 2088, 1761, 3533, 1868, 3162, 796, 797, 2545, 2049, 339, 1264, 3582, 3645, 1275, 865, 2622, 1089, 2538, 1363, 1671, 2117, 1945, 2334]. group-based [1264, 1275]. group-randomized [1561, 2224, 2837]. group-regularized [2117]. group-sequential [1608, 27, 78, 3533]. grouped [1016, 3198, 1684, 257, 343]. Grouping [3087]. groups [1561, 2581, 2728, 3487, 635, 1060, 740, 692, 400, 1327, 93, 1143, 1833, 1725, 3323, 3453, 246]. Growth [1861, 3502, 3505, 3501, 3459, 3322, 1219, 2853, 937, 239, 963, 3503, 3458, 2519, 3236, 2611, 1649, 3031, 893, 564, 3504, 1773, 3498, 793, 1222, 1884]. growth-scaling [3236]. Guest [1774]. guidance [1028, 1178, 356]. guide [614, 649, 613, 2060]. guidelines [742, 799, 2823, 2376]. Gumbel [1245]. Guttorp [411]. GWAS [494]. Gyldendal [1522].

H [3651]. H1N1 [905]. Haenszel [3465, 1613, 593, 1016]. Haihong [2106]. Half [3264]. Hall [2012, 447, 460, 118, 106, 411, 2591, 412, 2846, 446, 2725, 2691, 3059, 567, 433, 380]. Hall/CRC [2012, 447, 460, 118, 106, 411, 2591, 412, 2846, 446, 2725, 2691, 3059, 567, 433, 380]. Hamada [1558]. Handbook [2558, 411]. handle [3363, 1551, 2672]. Handling [780, 2398, 892, 2434, 1941, 2862, 1359, 1786, 2830]. Hankey [194]. Hanson [432]. haplotype [1107, 1278]. happens [405]. happily [3298]. harasser [3497]. Hardback [460, 2846, 445]. hardcover [2591]. Hardy [833, 326, 235, 252, 2502]. Hardy-Weinberg [833, 235, 252]. harm [3434]. harmonizing [2035]. Harrell [307]. Hartung [2848, 2301]. having [2668]. Hazard [3175, 1212, 2630, 1971, 2349, 1829, 909, 1728, 336, 3157, 2880, 387, 3094, 215, 1724, 1387, 2699, 1951, 3514, 3112, 3091, 3032, 327, 1096]. hazards [955, 1087, 3491, 2477, 1844, 2589, 923, 294, 1483, 3374, 1008, 3623, 90, 1097, 520, 1433, 2436, 1734, 2340, 2914, 1056, 2103, 3526, 1327, 387, 1341, 2931, 3509, 193, 2879, 1071, 1176, 126, 535, 2168, 1156, 1407, 1365, 2623, 956, 1975, 801, 1974, 498, 2991, 1451, 1286]. HB [3059]. head [2489]. Health [2439, 1681, 377, 30, 2512, 2077, 289, 443, 344, 1917, 345, 368, 903, 3142, 2820, 1375, 2915, 3385, 911, 105, 849, 501, 423, 427, 2683, 2760, 3422, 3079, 2443, 1494, 3512, 3015, 355, 164, 738, 2665, 3566, 168, 204, 913, 2095, 1551, 1833, 1125, 965, 3636, 1308, 2047, 371, 2350, 1383, 3375, 1022, 2504, 2505, 915, 2038, 2013, 500, 416, 848, 3240, 3550, 268, 997, 57, 1364, 2393, 1222, 167, 1666, 3160, 242, 626]. health-related [1375, 738, 1222]. healthcare [2700, 1085, 3559, 2304, 994, 1086, 625]. healthy [825]. heaped [2791]. heaping [912]. hearing [776]. Heart [3299, 49, 2437, 1507, 1695, 1200, 82, 2321]. heat [3639, 3654, 3653]. Heckman [2336]. height [3459, 3322, 2853, 3458]. hematology [470]. hemodialysis [2760]. hemoglobin [2137]. Hepatitis [3430, 424, 121]. hepatocellular [3430]. herbal [692]. Hermite [1674]. Heterogeneity [3013, 3451, 2978, 3491, 1369, 1555, 2588, 2397, 735, 1165, 1967, 2999, 945, 3584, 1360, 1260, 1602, 3493, 2800, 750, 1565, 2874, 1123, 595, 2338, 2537, 3283, 3521, 2859, 3557, 757, 1857, 3357, 802, 2155, 2322, 3543, 2857, 1594]. heterogeneous [902, 2259, 1631, 1259, 3288, 2449, 2935, 682, 3358, 2151, 3061, 3460, 3160]. Heteroscedastic [2288, 526, 259, 1378, 1876, 3183, 2487, 1717]. Heuvel [3651]. Hidden [479, 3074, 3339, 2218, 3325, 3326, 1246, 3378, 1695, 1248, 2511, 709, 623, 1201, 3126, 2677, 1175, 821, 3149, 2793, 1508, 628, 2488]. Hierarchical [2962, 1452, 2101, 2979, 27, 519, 1711, 671, 736, 934, 1577, 2385, 2213, 633, 2685, 2916, 3380, 2349, 3455, 2308, 1505, 3473, 2017, 2761, 2888, 3217, 3596, 2665, 2023, 1707, 755, 1565, 373, 1148, 3104, 1114, 484, 1175, 3418, 2409, 2967, 1275, 3456, 2926, 3219]. Higgins [586]. High [3412, 3171, 1842, 178, 1313, 754, 902, 699, 1466, 745, 3487, 3452, 325, 1064, 1566, 1390, 1872, 1896, 2907, 2619, 86, 3061, 978, 1784, 386, 1164, 2305, 2985, 3409, 3427, 3116, 2199, 2993, 1900, 966]. high-age [178]. High-dimensional [3412, 3171, 1842, 1313, 754, 699, 1466, 745, 3487, 1064, 1566, 1896, 2907, 86, 1784, 386, 1164, 2305, 2985, 3409, 3427, 2199, 2993, 966]. high-throughput [3452, 1390, 2619, 978, 1900]. Higher [666]. Higher-order [666]. highly [266, 1898, 2030, 1760, 2923, 2255, 794]. Hill [153, 2240, 1703]. Hill-based [1703]. Hippocampal [1470]. Hirakawa [1558]. historical [3345, 448, 2628, 1269, 1809, 3186, 198, 3441, 2824, 2697, 3193, 3208, 3408, 1179]. histories [1680, 2710]. history [1863, 2458, 2471, 2951, 1166, 3578, 522, 1392, 2359, 776, 677]. history-adjusted [677]. Hitting [3621]. HIV [2213, 1678, 2389, 3428, 403, 2308, 206, 270, 2089, 3528, 582, 1024, 2957, 503, 2274, 916, 1532, 2617, 1344, 3508, 1020, 192, 2652, 583, 1599, 2859, 2216, 310, 1879, 3012, 2058, 3640]. HIV-1 [206, 3528]. HIV-infected [503, 192]. HIV-positive [3428, 2216]. HNTD [1711]. hoc [1331, 2022]. Hochberg [2082, 3622]. Hodgkin [274]. Hogan [447]. Holm [1089]. home [43]. Homogeneity [931, 3480, 1714, 1033]. homogeneous [709, 1201]. homophily [3272]. honor [850]. horizon [2364]. hormonal [825]. hormone [1342, 2747, 2023, 1917]. Hosmer [1004]. hospice [1665]. hospice-use [1665]. hospital [715, 81, 2944, 2407, 2572, 1187, 598, 1883, 3283, 607, 3629, 757, 2099, 2599]. hospital-acquired [715, 3629]. hospitalisations [2503]. hospitalization [2036]. hospitalizations [3249, 1744]. hospitals [1112, 1486]. hostile [10]. hot [577]. household [3004, 2743]. housekeeping [1231]. HPV [2468, 280]. human [368, 1044, 1149, 1758, 517, 226, 382, 3500, 1046, 1045]. hunting [735]. Huntington [1506]. Hurdle [319, 2389, 3161, 2959, 1716, 3071, 1715]. Hurricane [3574]. Hybrid [62, 1361, 1708, 662, 512, 3219]. hyperactivity [782]. hyperbolic [1211]. hyperinflation [2678]. hypertabastic [2699]. hypertension [547, 564]. hypnogram [708]. hypoglycemia [1988]. hypotheses [2768, 1783, 1615, 1893]. Hypothesis [1963, 699, 1466, 214, 1544, 2539, 1136, 2417, 3520, 1014, 1046, 2286, 2540, 2219]. hypothesis-driven [2417].

I-II [1595, 2566]. I/II [1910, 2863, 75, 470, 1548, 499, 2565, 950, 3624]. I/IIa [789]. IBS [1127]. ICC [3122]. ICI [3538]. ideas [432]. Identifiability [1488, 2850, 637]. identifiable [264]. Identification [2694, 2486, 3348, 1659, 2941, 12, 578, 2367, 1085, 2662, 538, 3480, 3430, 553, 2573, 3045, 2748, 2510, 994, 1086, 2859, 3532, 3567, 3563, 3049, 3439]. identifications [1704, 2669]. identified [3307]. identifier [2089]. identify [1988, 1420, 1687, 2038, 2494, 3624]. Identifying [764, 949, 1349, 3639, 3654, 737, 3364, 3653, 3566, 3036, 1481, 915, 2580, 3377, 2980, 394, 2513, 1162, 1231, 635, 325, 3157, 1494, 982, 1798, 297, 3410, 1918, 3375, 2678, 1244, 504, 2565, 1601]. IDI [2801]. if [1073, 90]. ignorability [179, 3180]. ignorable [186, 919, 604, 1413, 351, 157, 2778, 1243]. ignorably [294]. ignorance [883, 725]. Ignoring [755, 1347, 3507, 247, 2099]. II [3647, 797, 1595, 1806, 477, 1688, 1862, 771, 842, 1923, 2566, 2315, 798, 1500, 1941, 466, 2732, 1910, 1795, 990, 1324, 303, 799, 1216, 2105, 991, 2633, 1549, 790, 3017, 2912, 2863, 3073, 75, 1687, 474, 1039, 3353, 2417, 148, 470, 2195, 277, 98, 721, 2014, 468, 399, 1548, 1343, 499, 1673, 2567, 2565, 781, 950, 2799, 3624]. II-binary [3647]. II-III [790]. II/III [1500, 466, 1795, 799, 1216, 2105, 2417, 98, 2014]. IIA [111, 789]. III [3017, 1500, 466, 1795, 799, 1216, 2105, 790, 1131, 2417, 98, 2014, 399, 3137, 1690]. IL7 [3280]. illness [2701, 206, 3023, 995, 2387, 1879, 2217]. illness-death [2701, 206, 3023, 2387, 1879, 2217]. illusion [2628]. illustration [1759, 3557]. illustrations [2182]. image [3224, 1161]. images [548, 2136, 3179]. imaging [676, 1754, 1802, 2067, 2667, 1427, 1482, 395]. Imbalance [2655, 560, 291, 1516, 177, 1207, 809]. immortal [2485]. immune [3624]. Immunization [371, 374]. immunogenicity [2880]. immunosorbent [1306]. immunotherapy [3201, 75, 3242]. Impact [1311, 3570, 3493, 1639, 3003, 1034, 165, 2263, 862, 937, 2434, 637, 882, 2548, 981, 3224, 1209, 884, 945, 551, 1958, 807, 881, 3433, 1950, 3541, 1509, 684, 1919, 245, 2897, 3351, 630, 2754, 2757, 3434, 3586, 1076]. impairment [481]. imperatives [198]. imperfect [2138, 1394, 729, 2598, 2805]. impingement [2667]. implausible [1162]. implement [252]. Implementation [2578, 1916, 141, 3584, 106, 2376, 1644, 1106, 1690]. Implementing [2537, 3460, 1234, 2295]. implications [1146, 1678, 2952, 833, 909, 3609, 3033]. importance [1381, 724, 702, 1988, 3303, 2719, 251]. Important [372]. Improper [667]. Improve [2682, 1146, 3307, 1055, 1051, 1289, 3398, 502, 612, 534, 468, 1343, 1213, 3383]. Improved [2955, 771, 2744, 1408, 2550, 2781, 2307, 1613, 1909, 1095, 441, 959, 1242]. improvement [2648, 1542, 36, 323, 2896, 647, 454, 3522, 3438, 1645, 1646, 322, 651, 1756, 2897, 1117, 1185, 2899]. improvements [2306, 772, 3234]. Improving [875, 763, 3043, 2945, 3265, 2103, 40, 2912, 522, 64, 2294, 57, 144, 339, 1564, 2402, 3301, 3611, 712]. Imputation [1920, 1543, 178, 3657, 2477, 2342, 3279, 1333, 1688, 565, 2382, 1841, 3159, 2339, 2336, 921, 3613, 430, 640, 1665, 1566, 1371, 139, 1990, 2341, 1734, 2340, 1301, 3182, 1689, 1922, 2815, 70, 1780, 977, 783, 1729, 2607, 1551, 449, 2512, 600, 472, 1344, 1410, 2027, 3318, 256, 2338, 1359, 344, 3092, 2450, 1701, 2832, 333, 1616, 2035, 2757, 3040, 3164, 293, 1666, 283, 356, 1732, 2199]. imputation-based [1689]. imputations [2342, 3384]. impute [626]. imputed [1820, 1272, 3314]. Imputing [2636, 977, 3167]. INAR [3325, 3326, 2488]. INAR-hidden [3325, 3326, 2488]. incentives [3454]. Incidence [1432, 3074, 2829, 2955, 1160, 403, 716, 596, 715, 914, 2771, 3637, 1238, 280, 2663, 2154, 608, 1729, 1656, 1532, 2587, 1237, 3508, 2940, 3235, 1928, 1883, 61, 1787, 410, 2827, 939, 1719, 2805, 3081, 1694, 1293, 3611, 2185]. incident [2820, 3406, 794]. include [2831]. including [3370, 1805, 341, 1867, 2650]. inclusion [448, 3387]. income [3371]. Incomplete [1519, 3479, 1828, 761, 897, 1351, 2398, 405, 430, 1566, 1401, 1399, 1745, 471, 907, 1410, 1080, 2335, 1155, 606, 2501, 1400, 2199, 2605]. incompletely [2950, 2479]. incompleteness [1665]. Inconsistency [2594, 3639, 3654, 1661, 2160, 3653, 2385]. Inconsistent [1984]. inconsistently [2391]. incorporate [2180, 15]. incorporates [1942, 2048, 2236, 2049]. Incorporating [2051, 1490, 1259, 888, 3433, 397, 1218, 3357, 3158, 3043, 2188, 3236, 1957, 2407, 2644, 1204, 110, 522, 2697, 1139, 3208, 1039, 1590, 98, 2014, 2913, 3377, 2568]. Incorporation [1863, 2676, 1992, 3232, 2139]. increase [1682, 1851, 614, 649, 613, 615]. increased [3042]. increases [1943]. Increasing [1859, 1381]. increment [1084, 1442, 1761]. increment-based [1084, 1442]. incremental [3438, 1584, 2076, 2899]. Independence [3116, 833, 2844, 2468, 252, 3127, 1565, 3595]. independent [260, 923, 621, 2214, 1967, 1734, 3638, 2442, 1287, 1878, 163, 1761, 2514, 3146, 2650]. Index [3538, 3021, 3459, 3322, 3371, 2087, 345, 1818, 1899, 2142, 3455, 1121, 1295, 1167, 3157, 1645, 1032, 59, 1202, 179, 1124, 3020, 3122, 539, 285, 1079, 3337, 307, 2895, 2028, 3365, 2141, 299, 2871, 3297, 2610, 1849, 1755, 2803, 1479, 1510, 3540, 3053, 943, 739, 2899, 840]. indication [254, 1870, 1422]. indicator [2178]. indicators [1219, 2072, 377, 1720]. indices [1525, 3590, 2856, 3218, 1838, 2050, 2703, 3267]. indifference [1205]. indifferent [3421, 3649]. Indirect [119, 208, 1199, 3058, 1187, 2424, 2629, 4, 583, 1801, 2776]. Individual [3231, 2013, 793, 987, 3517, 1514, 2758, 1333, 2624, 1674, 2300, 846, 1234, 2431, 947, 1073, 3363, 2398, 1992, 2618, 1331, 1224, 1920, 3038, 1298, 1943, 2815, 3468, 3045, 3037, 221, 3113, 1512, 1359, 1207, 1273, 924, 2260, 2035, 712, 3557, 3615, 486, 2400, 2933, 1997, 3287, 3003]. individual-level [1514, 3363, 3468, 221, 712]. individual-patient [1992]. Individualized [3643, 1572, 3196, 2642, 3154, 3440, 3111, 3404]. Individualizing [2459]. individually [1561, 3265]. individuals [178, 3428, 939]. IndOR [819]. Indrayan [118]. Induced [3011, 1895, 227, 1018, 3610, 3507, 1801]. Inefficiency [772]. infant [1315, 3372, 2489]. infants [1219]. infarction [1323]. infected [503, 192]. infection [2204, 403, 1994, 3572, 2030, 1509, 1123, 485, 811, 917, 2085]. infections [715, 3629, 310, 1189, 1601]. Infectious [2932, 753, 2811, 210, 429, 1520, 3003]. infective [3264]. Inference [3276, 2209, 2951, 1752, 2090, 2316, 1515, 608, 1904, 539, 2861, 309, 1047, 1610, 2605, 2641, 3238, 2755, 3185, 1747, 1570, 1321, 1319, 1619, 2401, 375, 1323, 3329, 1310, 2851, 2101, 2767, 596, 898, 3347, 1891, 1158, 2838, 3447, 2225, 2596, 3255, 2246, 3236, 1299, 808, 2183, 666, 2945, 562, 405, 1209, 1986, 643, 1979, 703, 3131, 582, 2162, 3106, 1298, 705, 2936, 313, 1316, 1729, 3248, 3302, 2750, 1287, 3589, 3195, 3189, 2735, 3175, 748, 2940, 3401, 3018, 1142, 792, 1320, 2084, 3365, 2500, 1853, 2625, 776, 774, 3092, 3474, 3262, 2996, 3390, 1648, 1846, 2828, 1185, 162, 762, 452]. inference [1653, 3437, 3614, 1343, 1520, 339, 959, 809, 1213, 1242, 3146, 3174, 1679, 1499, 1639, 3002, 814, 1243, 2076, 3540, 944, 2231, 1328, 1546, 2197]. inferences [442, 2634, 3180, 927, 3267]. inferential [142]. inferior [2234]. inferiority [759, 271, 2149, 214, 1135, 341, 154, 5, 348, 1226, 863, 347, 2381, 437, 2325, 2574, 213, 2, 474, 78, 3355, 601, 1178, 617, 1358, 928, 1855, 1192, 184]. inferring [1999]. infertility [311, 3319]. inflated [628, 529, 964, 1555, 3142, 2588, 1103, 2320, 2994, 3161, 479, 423, 2959, 1716, 2981, 1769, 3192, 2549, 319, 3071, 1565, 1020, 2233, 1744, 2024, 891, 1772, 1849, 1102, 2622]. inflation [1854, 476, 3124, 2257]. inflection [3250]. Influence [2666, 3028, 1103, 3309, 2790, 1047]. influenced [621]. influential [737]. influenza [3004, 2949, 905, 376, 3559, 1201, 2276]. Info [2080, 2093, 2491, 2507, 2524, 2542, 2109, 2125, 2144, 2157, 2171, 2192, 2207, 2222, 2238, 2253, 2268, 2283, 2298, 2312, 2328, 2345, 2362, 2378, 2395, 2411, 2429, 2444, 2461, 2475]. inform [3114]. informant [1428]. informants [1452]. Information [2080, 2093, 2491, 2507, 2524, 2542, 2109, 2125, 2144, 2157, 2171, 2192, 2207, 2222, 2238, 2253, 2268, 2283, 2298, 2312, 2328, 2345, 2362, 2378, 2395, 2411, 2429, 2444, 2461, 2475, 2079, 2081, 2282, 2284, 2297, 2299, 2311, 2313, 2327, 2329, 2344, 2346, 2361, 2363, 2377, 2379, 2394, 2396, 2410, 2412, 2428, 2430, 2445, 2446, 2460, 2462, 2474, 2476, 2092, 2490, 2492, 2506, 2508, 2523, 2525, 2541, 2543, 2094, 2108, 2110, 2124, 2126, 2143, 2145, 2156, 2158, 2170, 2172, 2191, 2193, 2206, 2208]. Information [2221, 2223, 2237, 2239, 2252, 2254, 2267, 2269, 2559, 2575, 2592, 2872, 2889, 2909, 2921, 2603, 2614, 2626, 2638, 2647, 2658, 2670, 2681, 2692, 2704, 2715, 2727, 2739, 2752, 2763, 2774, 2785, 2797, 2810, 2822, 2834, 2847, 2860, 3328, 3341, 2764, 453, 603, 3584, 2806, 1671, 2249, 1528, 1534, 3043, 505, 2644, 1269, 2147, 405, 1665, 2443, 1371, 136, 2618, 592, 1867, 3044, 3509, 534, 2289, 1305, 2228, 2687, 1110, 3104, 2661, 15, 57, 2448, 98, 2014, 2184, 3119, 1265, 2736, 2827, 69, 877, 3377, 2231, 3383, 2712]. Information-adaptive [2764]. Information-based [1671]. Informative [1353, 321, 3339, 2590, 1642, 489, 346, 170, 702, 2813, 3414, 3255, 1873, 621, 384, 3209, 3389, 1460, 286, 2069, 1961, 3483, 3515, 2800, 2751, 3289, 2537, 1888, 2275, 1786, 121, 2616, 456]. informatively [1926]. informed [595, 1892]. infrastructure [196, 200]. inhaler [2677]. inherent [3507]. inhibitory [1424]. Initial [3381]. initiating [2389]. initiation [192]. Initiative [1917]. inject [1519]. injection [1509]. injury [3443]. inner [2595]. innovation [3273]. Innovative [3276]. insight [1320, 533, 1546]. insights [1321, 1319, 2364, 2836]. insomnia [1376]. inspection [122, 2706]. Institute [3299]. institutions [2968]. Instrumental [1747, 1570, 1619, 2533, 3518, 2651, 2091, 3327, 135, 763, 1349, 2249, 488, 1686, 3457, 2683, 1654, 3420, 1837, 1870, 1435, 1951, 3076, 3226, 3387]. instruments [444, 763, 1822, 1431, 3335, 2533]. insulin [2017, 523, 618]. insurance [3422]. intact [3635]. intake [1246, 2463, 278]. Integrated [3538, 36, 19, 368, 2896, 3438, 2443, 1645, 1646, 2719, 137, 651, 2897, 3317, 1533, 2899]. Integrating [3471, 2514, 787]. integration [940]. Integrative [1803, 2324, 1260, 2278, 620, 1362, 1757, 2601, 3478]. integrity [1996]. intensity [1203, 1084, 1442]. intensive [2055, 1589]. intent [1891, 213]. intent-to-treat [1891, 213]. Intention [1960, 2560]. Intention-to-treat [1960]. Inter [3321]. Inter-unit [3321]. Interaction [1285, 1735, 51, 833, 899, 2215, 3121, 1420, 2844, 1894, 661, 405, 1979, 1661, 2316, 1513, 3478, 3430, 2043, 817, 1832, 3440, 1093, 1529, 3532, 1435, 3111, 2773, 3595, 3580, 3350]. interaction-based [3121]. interactions [936, 947, 1420, 836, 1468, 2618, 1922, 2817, 2972, 1360, 396, 834, 3375, 26, 1019, 1875, 1753, 835, 2099, 2610, 1935, 3377, 1757, 2074, 2967, 2980, 1425]. intercept [369]. intercepts [1184]. interchangeability [1027, 1036, 1035, 1428, 2776]. intercurrent [3619]. interest [49]. INTERGROWTH [3499]. INTERGROWTH-21st [3499]. interhospital [2770, 2722]. Interim [1193, 3305, 777, 1882, 759, 614, 649, 476, 467, 3073, 613, 615, 594, 1152, 2709, 2668, 3088]. Intermediate [2795, 3134, 3643, 1892, 1801, 2482, 1653, 1196]. Intermittent [634, 3466, 3128, 2652, 3359, 2706, 3640, 2600, 2020, 3072]. intermittently [2348]. Internal [2963, 1439, 1592, 999, 1348]. International [461, 1127, 2004]. Internet [606]. Internet-based [606]. Interplay [989]. interpret [1682, 1851, 3495]. Interpretability [724]. Interpretable [2735, 2380, 2780]. Interpretation [1456, 1455, 1281, 1916, 494, 1017, 1283, 953, 1700, 997, 1128]. interpretations [1340, 2246]. Interpreting [1215, 1419, 3566, 1339]. interrogation [2060]. interrupted [2915, 3385]. interstitial [1655]. Interval [262, 2360, 249, 1210, 573, 324, 2829, 321, 2087, 122, 1317, 2804, 2820, 2640, 2235, 715, 941, 2696, 1547, 2927, 3023, 1495, 3612, 1478, 2584, 730, 1056, 286, 2205, 2781, 1613, 1066, 70, 2069, 99, 2863, 3051, 3581, 2352, 1614, 82, 617, 958, 2189, 1527, 3467, 13, 167, 420, 1813, 1106, 327, 2991, 2645, 3624, 2673, 3147]. Interval-censored [1210, 2829, 122, 1317, 2820, 715, 2696, 1547, 2927, 3023, 1478, 730, 286, 2205, 1066, 70, 2069, 99, 3051, 2189, 3467, 167, 420, 1813, 327, 2991, 2673, 3147]. interval-censoring [321]. intervals [3658, 1593, 1637, 1642, 403, 1113, 635, 1211, 1603, 2317, 3136, 3204, 2287, 1135, 1371, 290, 2842, 3303, 2858, 1464, 2169, 2281, 3058, 1569, 3336, 3441, 1174, 616, 1675, 3494, 2585, 3627, 79, 590, 16, 2875, 906, 1538, 1372]. intervention [298, 39, 2915, 1165, 1752, 44, 3265, 1246, 2463, 2047, 492, 2153, 1501, 1679, 1567, 1813, 2326, 2756]. interventions [114, 3385, 1568, 845, 1384, 1579, 1842, 2515, 3598, 3629, 3524, 606]. Interview [1681, 443]. intima [1600]. intima-media [1600]. Intra [2366, 3255, 3215, 1615]. intra-class [1615]. Intra-cluster [2366, 3255]. intra-group [3215]. intraclass [1340, 3122, 958, 957]. intracluster [2546]. intralesional [2388]. introduced [621]. Introduction [850, 1774, 210, 822, 432, 3265, 1382, 460]. intuitive [3612]. invalid [2533]. invariant [2066, 3200, 1586, 2938, 1739]. Inverse [1325, 3023, 1444, 1389, 1338, 1651, 2053, 2551, 2365, 3429, 1158, 1211, 1799, 1986, 2707, 1240, 1084, 1442, 2441, 3450, 1459, 1733]. inverse-intensity [1084, 1442]. inverse-weighted [1651]. investigate [3497, 2065]. Investigating [1781, 3283, 354, 2983, 1062, 1369, 901, 499, 2555]. Investigation [3282, 2688, 1357, 3022]. Investigations [1178]. investigator [621]. investigator-introduced [621]. investigators [2257]. involving [1806, 2482, 2598]. IPD [2338]. IPTW [2053, 2551]. Iris [602]. irregular [1041, 2244, 1933, 2673]. irregularly [1084, 1442]. irrelevance [725]. IRT [1824, 3534]. ISBN [398, 2012, 447, 460, 432, 118, 520, 106, 1457, 411, 2591, 412, 2846, 446, 2725, 289, 506, 2691, 445, 3059, 1522, 567, 433, 380]. ISBS [1127]. Issue [2080, 2093, 2491, 2507, 2524, 2542, 2109, 2125, 2144, 2157, 2171, 2192, 2207, 2222, 2238, 2253, 2268, 2283, 2298, 2312, 2328, 2345, 2362, 2378, 2395, 2411, 2429, 2444, 2461, 2475, 2079, 2081, 2282, 2284, 2297, 2299, 2311, 2313, 2327, 2329, 2344, 2346, 2361, 2363, 2377, 2379, 2394, 2396, 2410, 2412, 2428, 2430, 2445, 2446, 2460, 2462, 2474, 2476, 2092, 2490, 2492, 2506, 2508, 2523, 2525, 2541, 2543, 2094, 2108, 2110, 2124, 2126, 2143, 2145, 2156, 2158, 2170, 2172, 2191, 2193, 2206, 2208]. Issue [2221, 2223, 2237, 2239, 2252, 2254, 2267, 2269, 2559, 2575, 2592, 2872, 2889, 2909, 2921, 2603, 2614, 2626, 2638, 2647, 2658, 2670, 2681, 2692, 2704, 2715, 2727, 2739, 2752, 2763, 2774, 2785, 2797, 2810, 2822, 2834, 2847, 2860, 3328, 3341, 850, 1947, 3367]. Issues [107, 689, 356, 1398, 972, 3329, 361, 363, 372, 872, 1088, 2890, 141, 860, 365, 359, 108, 1944, 1948, 362, 2463, 143, 1945, 360, 880, 364]. Italy [967]. Item [1824, 1355, 740, 3537]. item-response [1355]. items [3537]. iterative [1106, 2399]. ix [432]. Izu [767].

J [3648, 2106, 447, 647, 1497, 72, 649, 858, 459, 586, 411, 2725, 289, 3647, 287, 727, 767, 1889, 567, 184, 289]. J. [853, 858]. jackknife [1821]. jackknifing [1438]. Jackson [543]. Jacquez [817]. James [1045]. Japanese [511]. Jerome [852, 854, 850, 860, 859]. Jiajia [1974]. Joachim [602]. John [520, 2691, 433]. Johnson [432]. joinpoint [1693]. Joint [1628, 3417, 2389, 2853, 1726, 2188, 2870, 172, 1946, 2419, 227, 1018, 1650, 384, 1944, 1948, 2341, 3005, 58, 773, 1111, 3086, 1376, 1582, 3556, 1633, 2177, 1947, 2118, 228, 1945, 171, 2996, 342, 2068, 1120, 3124, 1479, 2484, 2458, 2517, 2664, 3199, 1627, 2916, 3428, 489, 527, 1993, 2839, 418, 998, 962, 528, 1873, 1547, 3388, 673, 582, 2553, 1491, 3253, 3085, 3200, 1652, 2069, 2818, 2855, 2917, 2981, 3188, 1446, 2372, 2831, 1392, 732, 2907, 2086, 1724, 2995, 3585, 357, 451, 2275, 2908, 1715, 3022, 3282, 1846, 2666, 1222, 2483, 3640, 54, 865, 2555, 2600, 232, 1737, 561, 589, 557, 1127]. ‘Joint [1945]. Jointly [954, 15, 504, 3032]. Jonckheere [1129]. journals [999, 689]. Jover [727]. Joy [2846]. JSM [2039]. Judgment [3256]. Julian [2725]. jump [3510, 1181]. justify [1647]. juvenile [3056].

K. [557]. Kaitlin [969]. Kaizar [1440]. Kamath [398]. Kaplan [20, 2291, 224, 876, 255, 2054]. Kappa [1276, 1587, 3481, 217, 1615, 1576]. Keen [567]. Kendall [3313]. kernel [241, 3019, 3247, 2932, 3108, 2489, 3319]. Kevin [567]. Key [952, 1088]. KI [3647]. kidney [2182, 300, 3397, 3394, 2867, 3552, 481, 1887]. Kieser [347]. kinetics [3085]. Kingdom [164]. Kitagawa [460]. Knapp [2848, 2301]. Knowledge [2159, 1957, 397]. known [2470, 642, 1480]. Knox [1598]. Korean [3656, 3442]. Kotz [495]. kriging [3281]. Kroner [1522]. Kwon [184].

L [727, 1889, 380]. L. [858]. Laake [3059, 1522]. Label [2938]. Label-invariant [2938]. labels [2257]. labor [266]. laboratory [823, 376, 607]. lag [225, 1471, 704, 3242]. lagged [2784]. Laird [134, 2160]. Landmark [1256, 2485]. landmarking [1157]. Lang [446]. language [689]. Laplace [2443, 137, 3317]. Large [1297, 3534, 2183, 1799, 2956, 977, 1308, 416, 1004, 221, 57, 3179, 859, 3000, 2062, 1690]. Large-scale [1297, 3534, 2183, 1308, 57]. lasso [3412, 1897, 1915, 3378, 1233, 2472, 2305, 3514, 3190, 1295]. lasso-based [1915]. last [1841]. late [2210, 515]. late-onset [515]. latency [3351, 2280]. Latent [2091, 3327, 3310, 1695, 1373, 1529, 1875, 232, 391, 1313, 1710, 2201, 897, 124, 3527, 1698, 1705, 745, 528, 2621, 3233, 3268, 267, 738, 3016, 3288, 2823, 1339, 204, 1649, 397, 696, 3222, 3086, 1660, 646, 2350, 2137, 2086, 3223, 2118, 2003, 2479, 2711, 268, 2708, 333, 1712, 3138, 2794, 1264]. latent-variable [897]. later [1590]. laterality [935]. Latin [2433]. Latin-square [2433]. lattice [2408]. Lauer [882, 884]. Lawson [433]. layered [1679]. layout [324]. Leacky [2846]. lead [1247, 3109, 2532]. lead-time [3109]. learner [2707]. Learning [2761, 2888, 2953, 189, 2277, 3217, 1799, 2611, 1909, 40, 3431, 3196, 3191, 1579, 3030, 3440, 2975, 3108, 3470, 3409, 2001, 3400, 3383, 2954]. least [34, 963, 1055, 2241, 1390, 1644, 1940, 3162]. least-squares [34, 1390]. Lecture [787, 1166, 2811]. Lee [3656, 2539, 1696]. Lefebvre [1889]. left [1317, 1726, 2689, 1291, 2371, 1056, 1780, 783, 3086, 954, 604, 1952, 1080, 2569, 2217, 167, 814, 498]. left- [167]. left-censored [1726, 2371, 1780, 783, 3086, 604, 1952, 2569]. left-censoring [814]. left-truncated [1317, 2689, 1080, 498]. legacy [3299]. Legionnaires [1261]. Leiden [2723]. Lemeshow [1004]. length [1219, 2187, 2572, 1174, 1436, 757]. length-biased [1174]. Leon [900]. lesion [676, 1943]. lessons [857]. Letter [2033, 208, 1497, 557, 150, 2669, 1058, 727, 175, 253, 3326, 3096, 1559]. leukaemia [1618, 677]. leukemia [436]. Leurgans [557]. Level [1662, 3436, 2451, 1678, 1514, 1959, 2259, 897, 2499, 2343, 947, 3344, 3363, 1805, 2090, 3433, 2957, 657, 896, 3468, 3082, 3516, 3203, 221, 2782, 3113, 1285, 924, 712, 3107, 2772, 1293, 1508]. Level-adjusted [1662]. levels [1921, 1428, 60, 257, 814]. leverages [2231]. Leveraging [1977, 712, 3595, 3383]. Li [2106]. liability [2676]. liability-threshold [2676]. Lie [445]. Life [2819, 1929, 1385, 1386, 3008, 2323, 1768, 2689, 1375, 141, 3114, 981, 2621, 3596, 1515, 738, 1166, 3578, 1316, 3406, 1111, 2917, 919, 3464, 2552, 1256, 792, 443, 2280, 3105, 380, 2695]. life-adjusted [3105]. life-table [981]. life-threatening [2323]. lifetime [3373, 332]. like [2683]. Likelihood [645, 1575, 3248, 2381, 830, 3605, 167, 3083, 1578, 3004, 810, 815, 3568, 1025, 1528, 1827, 3361, 963, 2101, 3544, 102, 2167, 666, 2945, 1371, 2261, 3596, 2858, 1758, 2511, 867, 2043, 1007, 3110, 3464, 231, 549, 539, 319, 2940, 2806, 215, 2289, 1174, 3018, 1614, 3360, 2500, 2698, 958, 1537, 3598, 1715, 265, 2569, 980, 157, 831, 782, 258, 951, 2502, 2058, 1243, 3295, 3091, 747, 332, 3407, 2574]. Likelihood-based [645, 3248, 830, 167, 3083, 2945, 2858, 1174, 3018, 1614, 3360, 958, 1243]. likelihood-ratio [2261]. likelihoods [808]. Likert [246]. limit [1205, 2519, 1502, 212, 918, 892, 1366, 986, 1813, 2772, 2865, 458]. limitation [178]. Limitations [2413, 2265, 201, 2415]. limits [2840, 2767, 2200, 1024, 1516, 1252, 549, 599, 175, 2825, 2943, 2231, 2535, 1034, 2029, 2745, 7, 176]. Lin [412, 1793]. Lincoln [2669, 1704]. line [1037, 259]. line-segment [259]. Linear [260, 1778, 1051, 2408, 3084, 2845, 654, 1313, 1057, 1197, 1104, 1337, 328, 1147, 2349, 1820, 1528, 1839, 3043, 483, 369, 2767, 1686, 454, 3457, 2431, 1090, 2388, 3642, 685, 2214, 225, 544, 946, 1471, 2272, 1995, 2720, 224, 2097, 2324, 2011, 2550, 3106, 317, 3340, 1695, 1481, 977, 3239, 503, 3173, 3581, 549, 2743, 1108, 1329, 1184, 3368, 2748, 2725, 429, 2140, 3351, 2531, 3607, 2337, 2293, 3641, 1744, 1012, 182, 296, 60, 1823, 708, 1042, 1453, 2883, 1749, 1849, 2230, 2794, 3450, 3287, 927, 1757, 2374, 2135, 2943, 576, 1727, 1510, 2231, 550, 644, 1911, 3007, 2695]. linearization [1512, 3367]. linearly [2406, 2982]. link [235, 953, 1821]. linkage [1705, 2742, 982, 1797, 22, 3174]. Linked [443, 921, 3171, 3017, 3387, 1306]. linking [720, 2089]. Links [2873, 279]. lipidomics [1796]. LISA [425]. lists [3222]. literature [2812, 672]. Liu [150, 3]. liver [1609]. LoA [46]. load [2389, 800, 1024]. Local [1103, 735, 84, 328, 2067, 2147, 3510, 2043, 179, 3180, 3203, 1598, 2391, 1847]. localized [376]. locally [710]. location [3185, 1317, 910, 1835, 896, 3082, 2750, 385]. location-scale [1317]. locations [1543]. LOCF [3314]. loci [710]. loess [1593, 1439, 1592]. loess-based [1593]. Log [328, 1017, 1283, 1282, 751, 369, 3209, 3597, 1549, 1821, 3483, 2536, 173, 2140, 1989, 1860, 708, 1610, 3305, 1281]. log-binomial [1989]. log-Gaussian [3597]. Log-linear [328, 369, 2140, 708]. log-normal [1610]. log-odds [751]. log-rank [1549, 173, 1860, 3305]. log-transformed [2536]. logarithmic [1280]. Logistic [35, 1317, 3205, 829, 1053, 1439, 2795, 3538, 147, 1349, 155, 328, 81, 1894, 2016, 80, 459, 1371, 3411, 2481, 3060, 1740, 1833, 2356, 3134, 1926, 231, 1984, 1592, 2150, 17, 2084, 2724, 248, 2908, 1233, 1583, 79, 353, 404, 2778, 3588, 3376]. Logistic-AFT [1317]. logit [965]. loglinear [755]. lognormal [1900]. logrank [20, 941, 1327, 1461, 3519]. London [2012, 1457, 1768]. Long [1264, 2364, 3114, 1468, 3397, 3036, 3101, 1235, 115]. Long-Term [1264, 2364, 3114, 1468, 3397, 3101, 115]. longer [3625]. Longevity [30, 2723]. Longitudinal [1873, 2971, 30, 416, 2479, 775, 1222, 2637, 787, 3412, 426, 654, 827, 676, 178, 1313, 3502, 2941, 1628, 2484, 3185, 3657, 3370, 3417, 3008, 1350, 2664, 447, 3339, 23, 1934, 2218, 3329, 2868, 401, 244, 2961, 2389, 1556, 2509, 2685, 2758, 761, 964, 527, 744, 1351, 1726, 1993, 2320, 2744, 2839, 3140, 55, 937, 1631, 2006, 275, 998, 2188, 2519, 2870, 1292, 1041, 1199, 2200, 1396, 849, 2246, 836, 1946, 2419, 2862, 3225, 670, 685, 1650, 1802, 384, 3245, 1944, 1948, 713, 557, 2310, 430, 1209, 3233, 44, 800, 2478, 3216, 816, 703, 2341, 582, 1024, 1575, 2553, 3402, 2842, 3411]. longitudinal [267, 1954, 1932, 592, 1670, 1835, 254, 183, 3381, 1626, 1771, 2694, 280, 3200, 3340, 1695, 1740, 1316, 2957, 813, 58, 773, 1067, 1551, 2593, 2818, 2855, 3086, 3592, 3480, 29, 1589, 3128, 965, 3188, 2486, 53, 682, 641, 1602, 2750, 2821, 1287, 1446, 3464, 3173, 1210, 179, 2372, 3336, 3435, 2262, 2452, 373, 908, 3234, 3499, 3251, 3046, 1633, 2005, 1724, 1676, 954, 1947, 2995, 1084, 1442, 2244, 2531, 228, 2705, 3068, 1739, 451, 1945, 776, 2010, 3120, 2376, 2708, 604, 15, 607, 826, 3598, 1221, 2996, 2448, 1288, 1761, 2035, 1012, 3129, 1607, 1800, 2569, 3141, 2216, 1078, 1435]. longitudinal [1933, 60, 3022, 3282, 1501, 2814, 2773, 3469, 1741, 1846, 2666, 3040, 3384, 1875, 342, 157, 2656, 422, 2722, 69, 493, 658, 1653, 1617, 2232, 2521, 2794, 167, 3621, 1666, 1564, 2402, 1414, 3640, 1368, 1243, 191, 576, 68, 1083, 1831, 2198, 2803, 2391, 1182, 3083, 1737, 2946, 1911, 2926, 2695, 1671, 1641, 3252, 3088, 537, 456, 2459, 2520, 2712, 506]. longitudinal-survival [1873]. longitudinal/spatial [3480]. longitudinally [2066, 824]. look [1864, 1866, 476, 1865]. Lorenz [3483]. loser [1430, 180]. losers [573]. loss [1386, 922, 3604, 2259, 2838, 3578, 2737, 776, 1413, 3400, 1655, 672]. loss-based [2838]. lost [1385, 2819]. lot [768, 1597, 767]. lottery [3454]. Low [2295, 1112, 1476, 1656, 2660, 2808, 1912, 1896, 1931, 2186, 1190, 2807, 2859]. low-dimensional [2186]. low-dose [1912, 1931, 2859]. Low-event-rate [2295]. low-incidence [1656]. lower [599, 1813, 1601]. lowering [929, 1298]. Ltd [1457]. Lu [584, 150, 3185, 1974]. Lung [3299, 3476, 996, 3109, 240, 3077, 2262, 1703, 2736, 3281]. Luo [496]. LX [194]. Lydersen [3059, 1522]. Lyle [2012]. lymph [1122]. lymphoblastic [677]. Lymphoma [274, 732]. Lynch [1449].

M [628, 777, 647, 208, 72, 586, 150, 1865, 2539, 347, 1265, 1889, 1792, 1791, 447]. M-D [777]. M-K [1792, 1791]. Macaskill [3649]. Machin [289]. Machine [3383, 40, 3247, 3409, 2001, 3319]. machines [34]. made [1006]. magnetic [676, 1754, 2067, 2667, 1161, 1482]. magnetoencephalography [577]. magnetoencephalography/electroencephalography [577]. maintenance [677]. Maiti [459]. Major [1560, 2146, 2544]. make [970]. makes [3312]. Making [152, 442, 2415, 402, 812, 480, 455, 2922]. Malaria [1278, 1238, 1224]. Mallick [150]. malnutrition [3392, 3336]. management [2677, 749]. Managing [1690]. Mandel [3115]. Manhattan [618]. manipulated [639, 638]. Mann [3658, 18, 3136, 3204, 3290, 1819, 627, 3204, 1859, 1499]. MANOVA [2392]. Mantel [593, 3465, 1613, 1016]. many [2466, 1822, 3606, 3175, 908, 2015, 3640]. Mapping [1262, 1898, 1743, 3161, 1041, 2290, 2719, 127, 1022, 1543, 162, 2933, 246]. maps [2854]. Marc [2846]. margin [2149, 3221, 341, 113, 601, 1358, 1193]. margin-based [3221]. Marginal [1494, 605, 1407, 525, 1212, 375, 3199, 2365, 897, 2083, 2744, 3169, 1495, 1799, 2612, 979, 1427, 2503, 2707, 1067, 1662, 1509, 1180, 3046, 1190, 2734, 1110, 2153, 2995, 3396, 677, 3475, 3393, 2450, 49, 3214, 79, 3492, 2867, 1265, 1889, 1981, 1845, 801, 2803, 1852, 1641]. marginal-conditional [2995]. marginal-mean [1427]. Marginalized [2959, 1716, 3071, 2233, 1769, 3104, 1750]. marginally [1136]. margins [772, 1031]. Maria [3574]. Marie [506]. Marit [1522]. Mark [106]. marked [1314]. marker [1682, 1851, 2898, 1623, 3571, 1646, 1573, 2228, 2687, 1154, 392, 1050, 1169, 3456]. marker-stratified [3456]. markers [728, 1053, 1146, 2916, 1147, 208, 1502, 3616, 1467, 4, 3120, 1607, 524, 2859, 1078, 931, 2899]. market [2211, 2942]. marketing [1127]. Markov [628, 3325, 3326, 3074, 3339, 2218, 479, 2488, 2498, 663, 3056, 1246, 737, 3466, 3378, 1056, 1149, 1695, 1339, 1248, 2047, 709, 517, 1201, 331, 3126, 2387, 2908, 2677, 121, 1966, 1175, 3149, 66, 1508]. Markovianity [995]. Martín [3096]. Martínez [3021]. Martingale [2469]. Mary [3483]. masked [169]. mass [3459, 3322, 3371, 2075, 59, 1124, 3337, 2803, 395]. massive [911]. Mastering [2174]. Matched [3284, 442, 375, 1970, 1818, 1361, 1232, 2271, 102, 3626, 2869, 3512, 3135, 490, 1241, 3060, 930, 907, 642, 2, 2635, 309, 1164, 2104, 13, 3124, 3294, 184, 1587, 1900, 1192]. matched-pair [375, 3626, 2, 13, 184, 1587]. matched-pairs [3135]. matches [3605]. Matching [3474, 2263, 1484, 1709, 2070, 3316, 1858, 2465, 2401, 1375, 284, 1331, 3433, 1830, 3065, 684, 3564, 2077, 2497, 3228, 1347, 1389, 1868, 762]. materially [788]. maternal [2853, 1995, 564, 3251, 2321]. mathematical [1139]. mathematics [1763, 2530]. matrices [917]. matrix [3501, 1351, 3216, 1426, 2078, 1213]. Matsui [1558]. Matsuyama [1266]. matter [2665]. Maurice [1380]. max [502, 612]. maximally [895, 1708, 2653]. maximin [1581, 710]. maximisation [552]. Maximization [2645, 469]. maximize [2982]. maximized [33]. Maximizing [1147, 1139]. Maximum [3004, 3568, 476, 3528, 549, 2257, 3238, 963, 3544, 2167, 3110, 3139, 1537, 3598, 2569, 2583, 2580, 2494]. may [2624, 448, 1073]. McNemar [2294]. MCP [2883]. MCP-MOD [2883]. MD [1380]. mean [1651, 1929, 1057, 389, 1430, 744, 1318, 3523, 3652, 2689, 12, 2471, 1315, 785, 1871, 1473, 3612, 2621, 3397, 1427, 3097, 3596, 3389, 2858, 3194, 3426, 593, 3406, 1540, 449, 3449, 893, 1675, 2762, 535, 2661, 1632, 1856, 2639, 3614, 1997, 2369, 927, 683, 1664, 2695, 332]. mean-association [744]. mean-covariance [2695]. meaningfulness [530]. means [2738, 898, 785, 2468, 1926, 3619, 601]. measles [3544, 752]. Measure [3021, 2630, 2898, 345, 1748, 1147, 1103, 647, 2291, 2397, 2048, 2236, 3438, 1074, 2316, 1640, 3154, 3020, 322, 2990, 2227, 1953, 2871, 24, 2843, 1054, 2435]. measured [3213, 2916, 2469, 2066, 3205, 2439, 3001, 3636, 42, 1830, 2536, 3025, 824, 2918, 1718, 3129, 1435, 1749, 3594, 951, 191]. Measurement [3617, 3074, 996, 920, 1404, 1916, 3445, 2423, 46, 1291, 2200, 2048, 2236, 526, 1553, 3216, 816, 582, 940, 1289, 1557, 1958, 234, 1950, 1448, 3493, 3581, 3416, 259, 3166, 1063, 1876, 2554, 3183, 3251, 2762, 3351, 848, 2531, 1968, 2708, 62, 976, 1720, 3518, 1717, 1222, 831, 2230, 1348, 2943, 2046, 3149, 1903]. measurement-error-driven [3149]. measurements [1350, 920, 780, 2839, 547, 3631, 2912, 3464, 620, 312, 2262, 1108, 1378, 1876, 192, 276, 277, 2656, 1343, 2632, 3158, 246, 2788]. Measures [2978, 3522, 2015, 1053, 1496, 2422, 3069, 636, 699, 2516, 841, 842, 2919, 1208, 235, 2789, 3079, 1486, 800, 2436, 415, 538, 2163, 88, 2512, 605, 2226, 1707, 1354, 1456, 3336, 3556, 261, 834, 3235, 2790, 3368, 623, 3298, 2118, 1487, 774, 230, 2035, 175, 890, 2904, 2713, 1185, 1653, 906, 814, 1129, 2963, 7, 176, 1455]. Measuring [2458, 369, 2431, 823, 2974, 1378, 3496, 2798, 3554, 2548]. MEBoost [3445]. mechanism [2471, 3188, 1616, 351]. mechanisms [2336]. mechanistic [1122, 3544, 3280, 2946]. Med [3185]. media [1600]. median [2630, 1515, 1904, 3554, 332]. medians [1517, 3334, 707]. Mediation [3374, 3300, 2866, 1749, 3594, 3349, 3560, 3027, 2868, 3487, 1803, 3016, 3181, 2850, 3063, 3052, 2065, 1338, 2631, 891, 1312, 3320, 3292]. mediator [3027, 2424, 3610, 1749, 1312, 3349]. mediator-outcome [3610]. mediators [3374, 3487, 2866, 3610, 3063, 3594]. Medical [1381, 1085, 994, 624, 1318, 2164, 1254, 2431, 2009, 153, 1560, 2761, 2888, 2310, 1433, 2615, 2858, 1999, 1373, 3481, 1010, 1382, 1625, 2351, 1120, 3117, 456, 118, 289, 1522]. Medicare [3380, 3171, 908, 3387]. medication [1655]. medications [2277, 901, 500, 1070]. Medicine [3656, 1325, 41, 692, 688, 690, 696, 1950, 694, 999, 2174, 2986, 879, 3064, 693, 691, 3648, 584, 495, 208, 1497, 72, 557, 459, 586, 150, 2669, 3, 2539, 347, 194, 287, 408, 1440, 175, 3649, 184, 543, 3369]. medicines [2693]. meet [1798]. meeting [2004, 461]. meets [511]. Mehrotra [150]. Mehta [649]. Meier [20, 2291, 224, 876, 255, 2054]. melanoma [3232, 1773]. mellitus [1190]. Member [398]. members [2323]. Memory [1137, 802]. Mendelian [3371, 389, 2688, 444, 2249, 1686, 1822, 2000, 1098, 756, 2918, 3335, 2913]. menstrual [3410, 825]. menstruation [2802, 3410]. mental [2683, 168, 2038, 500]. merits [2070]. MERS [3656, 3442]. mesothelioma [967]. Meta [2624, 211, 929, 2083, 434, 2343, 1580, 1571, 2844, 136, 2615, 1925, 2989, 1999, 1405, 2984, 3037, 2243, 1207, 1357, 301, 1856, 2639, 1997, 3287, 2878, 1062, 560, 587, 2684, 987, 543, 2590, 97, 1430, 1294, 2180, 3485, 135, 1333, 1621, 1521, 1816, 2209, 2571, 1304, 542, 2397, 1674, 2300, 1234, 96, 1841, 2384, 1497, 947, 1073, 2649, 3465, 3639, 3654, 785, 2264, 544, 946, 1001, 2100, 2398, 2780, 666, 1002, 1553, 545, 2480, 2960, 541, 2128, 659, 290, 413, 2127, 1992, 2618, 3156, 586, 3062, 134, 540, 546, 945, 1074, 1661, 1443, 2160, 2848, 3010, 1954, 1815]. meta [1920, 3246, 3591, 3655, 1991, 1298, 807, 1395, 2457, 3653, 2130, 2815, 1867, 3433, 3231, 187, 3271, 3212, 48, 581, 1144, 3358, 1003, 1788, 1840, 3314, 3334, 1394, 3045, 750, 1663, 2175, 2064, 616, 2629, 3018, 2585, 3013, 137, 1215, 1832, 331, 595, 2338, 700, 1359, 2537, 2938, 357, 1968, 1000, 263, 2078, 287, 299, 1713, 237, 2661, 619, 3482, 3632, 129, 1498, 2295, 295, 2035, 3521, 2528, 1940, 3557, 296, 3022, 3282, 3469, 2825, 1511, 1937, 1857, 3357, 2400, 1577, 3033, 1217, 1094, 3618, 2301, 2798, 3078, 3543, 3535, 2063, 1985, 2385, 1594, 513, 2197, 2400, 130, 133, 2129]. meta [132, 131]. meta-analyses [587, 2209, 134, 945, 2160, 3010, 1991, 1867, 3358, 595, 2295, 1511, 1594]. meta-analysi [586, 287]. Meta-analysis [2624, 211, 929, 2083, 434, 2343, 1580, 1571, 2844, 2615, 1925, 2989, 1999, 1405, 2984, 3037, 2243, 1207, 1357, 301, 1856, 2639, 1997, 3287, 1062, 560, 2684, 987, 543, 2590, 97, 1294, 3485, 135, 1333, 1621, 1521, 1816, 2571, 1304, 542, 2397, 1674, 2300, 1234, 96, 1841, 2384, 1497, 947, 1073, 2649, 3465, 3639, 3654, 785, 2264, 544, 946, 1001, 2398, 2780, 666, 1002, 1553, 545, 2480, 2960, 541, 659, 290, 413, 2127, 1992, 2618, 3156, 3062, 540, 546, 1074, 1661, 1443, 2848, 1954, 1815, 1920, 3246, 3591, 3655, 1298, 807, 1395, 2457, 3653, 2815, 3433, 3231, 187, 3271, 3212]. meta-analysis [48, 581, 1144, 1003, 1788, 1840, 3314, 3334, 1394, 3045, 750, 1663, 2175, 2064, 616, 2629, 3018, 2585, 3013, 137, 1832, 331, 2338, 1359, 2537, 357, 1968, 1000, 2078, 299, 1713, 2661, 619, 3482, 3632, 129, 1498, 295, 2035, 3521, 2528, 1940, 3557, 296, 3022, 3282, 3469, 2825, 1937, 1857, 3357, 2400, 1577, 3033, 1217, 1094, 3618, 2301, 2798, 3078, 3543, 3535, 2063, 1985, 2385, 513, 130, 133, 2129, 132, 131]. meta-analytic [1430, 700, 2197]. Meta-analytical [2878, 2180]. meta-epidemiological [2100, 2938]. Meta-regression [136, 666, 1443, 1215, 2537, 263, 237, 2798]. Meta-STEPP [2400]. metastasis [163]. metastatic [1015]. Method [207, 391, 2070, 1710, 1854, 189, 2469, 1361, 3620, 584, 897, 1816, 2571, 369, 1254, 469, 2384, 1247, 2717, 3465, 494, 501, 2388, 3639, 3654, 661, 544, 1450, 2779, 2310, 3042, 1433, 2584, 381, 3596, 989, 3547, 1443, 2044, 303, 772, 154, 164, 1066, 3, 3653, 1241, 388, 1107, 3244, 925, 510, 783, 1551, 1833, 2855, 553, 1907, 3511, 1159, 1489, 349, 2485, 646, 907, 511, 1356, 1378, 1876, 684, 2294, 402, 812, 2582, 515, 848, 2705, 2918, 2500, 1426, 414, 1537, 2432, 15, 2152, 1715, 717, 1583, 3461, 2859, 370, 2883, 278, 2905, 3323, 1009, 3439, 2983, 2301]. method [1, 821, 191, 2865, 3367, 1170, 1601, 2399, 950]. methodological [1369, 361, 363, 365, 359, 362, 3500, 360, 364]. methodology [827, 2031, 2004, 465, 2576, 134, 2160, 690, 3499, 2661]. Methods [3307, 3503, 1126, 1941, 1518, 117, 1877, 1681, 1552, 982, 110, 3504, 2552, 2387, 532, 1501, 3496, 2230, 1062, 996, 1454, 731, 1561, 216, 442, 1212, 1496, 1747, 1570, 1619, 2876, 920, 2004, 1105, 2180, 135, 2249, 529, 488, 1337, 1342, 46, 323, 936, 929, 1534, 2744, 1036, 25, 2055, 1698, 1296, 1923, 2765, 2382, 1379, 2700, 935, 1085, 2944, 2760, 1944, 1948, 461, 2089, 888, 1700, 413, 146, 1798, 1523, 3087, 586, 563, 1224, 378, 1815, 772, 1922, 799, 1796, 2815, 1867, 622, 977, 1729, 2095, 1656, 1013, 1067, 2258, 2981, 3542, 42, 1431, 231, 830, 2586, 3405, 200, 3234, 2190]. methods [3101, 999, 1148, 3018, 292, 2846, 2186, 3394, 2578, 2352, 3061, 1947, 1738, 1084, 1442, 2159, 3113, 1485, 2497, 3240, 1968, 1945, 1285, 1735, 994, 1086, 2597, 15, 1270, 1786, 3390, 2487, 386, 3022, 3282, 310, 370, 13, 1422, 196, 1857, 3644, 831, 1951, 1935, 2155, 167, 3137, 3160, 3287, 3645, 94, 497, 609, 2857, 3383, 1852, 3088, 3505]. methylation [2423, 1784, 1164]. Metrics [1526, 569, 3538, 2901, 3358, 651, 570, 568, 686, 2280, 2902]. Metz [3080]. mHealth [2256]. MICE [1920]. Michael [882, 1457, 2591, 289]. micro [2256]. micro-randomized [2256]. Microarray [992, 325, 1148]. Microarray-based [992]. microbiome [3207]. microdata [3184]. MIDAS [2416]. midtrial [1132]. Miller [445]. Min [412, 502, 612]. min-max [502, 612]. mine [2570]. mineral [3093]. miners [996]. minimisation [667]. minimization [698, 2539, 2969, 76, 2286, 2540]. Minimum [3352, 3353, 1351, 2594, 2838, 1424, 1696, 814, 3647]. mining [398, 2812]. MIREC [3239]. mis [1984, 1282, 1721]. mis-specified [1984]. mis-substitution [1721]. miscategorized [2353]. misclassification [101, 3637, 637, 302, 2966, 3626, 2869, 3512, 3271, 3633, 1624, 2442, 231, 1383, 2554, 1718, 611, 2230, 3024, 3002, 3400, 1508, 1721]. misclassified [2218, 3197, 1177, 3390, 1902, 2714]. mismeasured [3185, 847, 2750, 3390, 2946]. misreported [1492]. misreporting [1665]. Missing [3477, 2783, 2334, 49, 2176, 426, 3568, 308, 920, 2477, 3028, 294, 2072, 186, 1726, 2343, 3310, 3347, 2838, 1488, 2339, 2862, 2264, 2777, 3296, 557, 2987, 1677, 139, 3135, 1920, 1922, 161, 338, 2457, 1796, 2330, 30, 1729, 2816, 1418, 1067, 1551, 2850, 919, 1926, 3128, 918, 1960, 2512, 1699, 600, 641, 3192, 3464, 1840, 3314, 2549, 385, 3172, 1237, 2332, 256, 3167, 3365, 2338, 228, 2500, 1359, 2337, 2532, 2353, 1233, 604, 344, 392, 230, 1718, 1364, 2333, 1616, 1156, 3524, 1078, 3132, 2184, 2875, 3164, 1937, 293, 69, 1617, 3450, 283, 606, 993, 1414, 3295, 618, 2616, 447]. missing-data [1718]. missing-not-at-random [2532]. missingness [346, 3140, 3363, 2783, 29, 144, 2713, 157, 2778, 3640, 1243]. Misspecification [926, 654, 937, 2188, 3570, 1209, 223, 762]. Misspecified [2183, 2961, 90, 3603, 1184, 2828]. Mistake [208]. Misunderstandings [2127, 2128, 2130, 2129]. Misuse [838, 1978]. mitigate [3156]. mixable [2225]. Mixed [1231, 547, 3293, 2748, 924, 894, 708, 3552, 654, 555, 1313, 1057, 3185, 1197, 260, 1104, 1294, 2976, 155, 1337, 1025, 1463, 1839, 1993, 55, 483, 71, 1674, 3310, 1778, 96, 2431, 1199, 947, 1073, 1396, 1090, 2388, 3642, 685, 2264, 1668, 2214, 2272, 1995, 2018, 562, 1478, 1566, 910, 2436, 238, 2553, 2011, 3106, 1835, 183, 1626, 317, 1943, 1395, 3200, 3340, 1695, 1481, 2095, 1649, 896, 1540, 2855, 2981, 3082, 1602, 2750, 179, 3416, 2735, 709, 2743, 2907, 1108, 2137, 1184, 722, 2064, 2725, 3607, 3240, 1739, 2337, 1913, 1218, 3520, 3632, 2996, 1012, 775, 296, 793]. mixed [3132, 60, 2713, 2828, 278, 734, 1617, 2794, 1881, 2943, 191, 2680, 2231, 550, 1508, 1150, 334, 2712, 446]. mixed-effect [1463, 1396, 1943, 1913, 2828]. Mixed-effects [894, 3552, 3185, 1993, 1778, 2431, 910, 2436, 238, 2553, 1835, 1626, 896, 1602, 2750, 2137, 3607, 3632, 2996, 278, 734, 1617, 550, 2712]. mixing [917]. Mixture [2820, 1355, 1532, 1049, 340, 1123, 2985, 420, 555, 912, 186, 1317, 1103, 1993, 1631, 2937, 1291, 2519, 2225, 528, 2862, 423, 804, 2777, 2407, 1797, 563, 3466, 2611, 1745, 317, 3566, 387, 3181, 1961, 682, 1489, 1602, 1843, 2426, 610, 385, 1667, 1565, 1861, 564, 123, 1180, 1148, 3256, 171, 2189, 158, 3207, 3141, 2814, 1453, 802, 960, 1189, 1162, 1887, 1911]. mixture-model [912]. Mixture-of-exponentials [1123]. Mixtures [462, 1209, 2572, 3247, 3375]. MK [789]. MK-0752 [789]. MNAR [2336]. mobile [935]. MOD [2883]. Model [112, 3028, 1959, 3140, 2841, 552, 1228, 1144, 110, 1724, 1524, 1387, 3311, 2425, 762, 3489, 3382, 654, 1112, 1651, 676, 555, 1313, 2213, 2484, 2458, 266, 2882, 2701, 2684, 1197, 912, 3008, 2876, 321, 2201, 3339, 1934, 2477, 1627, 2218, 2470, 1173, 2210, 294, 1349, 1274, 1483, 1172, 597, 2853, 2676, 328, 2349, 489, 1318, 1232, 1463, 1839, 2164, 2083, 2102, 2370, 3055, 3050, 55, 1535, 3043, 298, 39, 2964, 3374, 1449, 841, 842, 2308, 2434, 2098, 170, 71, 2188, 2915, 3385, 469, 1292, 1315, 1841, 81, 2200, 1208, 1396, 914, 3414, 3544, 1423, 528, 1171, 3424, 1505, 2473]. model [2862, 423, 1090, 2048, 2236, 206, 1018, 685, 3363, 804, 3473, 2336, 233, 2214, 3236, 3570, 3623, 3023, 1495, 2720, 1980, 2017, 3388, 2019, 562, 2371, 3224, 1478, 41, 1097, 1836, 2784, 2621, 3157, 803, 2529, 1433, 643, 3080, 1797, 19, 2436, 1925, 2601, 663, 2404, 3056, 1460, 1246, 119, 1661, 1734, 267, 2503, 2939, 1447, 3466, 818, 22, 1730, 948, 1056, 2707, 3410, 2901, 183, 1922, 2103, 3381, 3553, 3413, 1771, 317, 1943, 1066, 2457, 1360, 1065, 3085, 3200, 1405, 1016, 1327, 3060, 1695, 1481, 93, 88, 1464, 3232, 1391, 3406, 1646, 3181, 87, 669, 2069, 696, 896, 795, 1111, 1540]. model [1969, 2356, 2464, 2672, 2818, 2609, 2917, 919, 29, 99, 1341, 1589, 1660, 3082, 449, 1260, 1961, 2285, 3541, 1769, 1183, 1489, 1007, 1177, 1707, 1582, 2121, 3192, 2047, 3054, 2056, 3558, 610, 3029, 3416, 1904, 2735, 709, 1003, 1788, 732, 1344, 784, 509, 2549, 2824, 1509, 1565, 598, 2027, 2350, 1329, 1383, 2137, 1876, 564, 1973, 3372, 2357, 113, 2064, 1703, 1180, 508, 1142, 2879, 2725, 2635, 1885, 2582, 1139, 429, 2140, 340, 1117, 95, 2084, 1010, 664, 2151, 86, 591, 2995, 3585, 3365, 2003, 1362, 3366, 1892, 2556, 1512, 2698, 3647, 3352, 3353, 451, 1913, 995, 2387, 2275, 2711]. model [825, 248, 2908, 3283, 3052, 2708, 2432, 604, 3641, 1218, 3537, 2189, 272, 1527, 223, 1616, 1012, 1750, 1436, 1712, 532, 2677, 1156, 3141, 296, 3359, 1823, 2814, 3250, 3467, 1265, 1889, 2699, 554, 3608, 3593, 3338, 1879, 3518, 455, 3049, 410, 1365, 278, 2827, 734, 2217, 1749, 1222, 3280, 802, 439, 956, 1975, 1849, 2230, 2794, 2985, 3144, 3437, 3409, 3439, 2483, 3190, 32, 927, 1884, 1881, 1102, 3000, 1414, 1757, 1974, 2706, 3640, 1639, 420, 2943, 1181, 960, 1755, 2501, 3091, 1831, 2622, 3149, 66, 747, 1081, 232, 498, 458, 644, 1737, 2453, 2946, 3090, 561, 1451, 1887, 1911, 2926, 3117, 3007, 1054, 2695]. model [2651, 693, 1286, 2645, 3089, 1578, 1903, 2922, 3219, 2788, 2930, 2712, 1258, 1775, 2011, 2489]. model-adjusted [1054]. model-assisted [3089]. Model-based [112, 1959, 552, 110, 1524, 3489, 2210, 1463, 1922, 508, 1512, 554, 3049, 3439, 927, 2922]. model-building [1639]. model-free [1755]. model-informed [1892]. Modeling [2813, 753, 1471, 3392, 910, 2999, 3216, 3454, 1475, 1834, 556, 65, 3249, 1125, 2977, 2023, 2877, 3154, 371, 2504, 2505, 3337, 3574, 1221, 1303, 757, 2928, 1102, 3252, 537, 3074, 3276, 2791, 1901, 1628, 2517, 2701, 447, 243, 84, 460, 1898, 2685, 2916, 3380, 489, 527, 1310, 1726, 2949, 1231, 172, 2437, 2246, 1946, 1873, 227, 1355, 1018, 2802, 1650, 384, 2683, 1944, 1948, 2761, 2888, 90, 519, 1799, 2025, 1002, 1766, 1408, 557, 2310, 336, 673, 1209, 19, 1024, 1803, 415, 1491, 2971, 2162, 3005, 2611, 21, 1626, 280, 2665, 3200, 3332, 204, 2957, 3539, 58, 930, 1111, 2855, 3086, 2981]. modeling [2226, 1506, 450, 2372, 1955, 1684, 3556, 517, 1667, 1237, 373, 1020, 1861, 1022, 123, 215, 3256, 1724, 3304, 954, 1996, 1947, 2118, 228, 2705, 1945, 1176, 1785, 1744, 1715, 2120, 333, 3524, 793, 1933, 3250, 1741, 1846, 342, 1262, 939, 934, 301, 3418, 2321, 2637, 1368, 54, 1257, 1275, 1120, 3124, 68, 3456, 67, 1508, 681]. Modelling [426, 308, 97, 1768, 1314, 2018, 355, 2572, 2879, 125, 1114, 485, 1006, 2579, 1828, 3562, 2770, 1122, 998, 2188, 2919, 1008, 1568, 2419, 1238, 2407, 516, 932, 3015, 3597, 1149, 2663, 87, 2912, 3077, 2940, 1049, 2510, 3498, 268, 60, 3123, 634, 1520, 2600]. Models [2588, 769, 3172, 3496, 720, 3299, 628, 731, 724, 3502, 1386, 2962, 810, 3185, 754, 1824, 431, 955, 1087, 1053, 1439, 3538, 3417, 1452, 1541, 2438, 3443, 260, 2204, 2664, 2247, 3028, 1781, 2342, 1434, 1678, 389, 2073, 2961, 2389, 3373, 2072, 1104, 1514, 1556, 2976, 964, 1731, 186, 1555, 1337, 323, 3142, 122, 929, 1317, 1103, 1025, 1528, 1993, 2288, 2320, 2839, 3070, 1631, 2965, 483, 2820, 1895, 2937, 239, 124, 2055, 2006, 1531, 3161, 1229, 2689, 1674, 2919, 3310, 1705, 1291, 2519, 2870, 1686, 1021, 962, 1041, 3121, 2970, 479, 846, 1234, 1571, 952, 838, 264]. models [2979, 745, 1778, 2431, 1767, 3169, 1121, 3522, 2437, 3616, 2246, 1897, 2789, 1894, 1908, 1295, 2388, 2287, 3642, 845, 1668, 2777, 3570, 225, 1471, 1167, 1647, 1630, 1238, 330, 2292, 1001, 2272, 3041, 1995, 2290, 3296, 1629, 566, 2187, 909, 2644, 2183, 424, 1259, 3275, 1728, 740, 1965, 3079, 979, 1468, 640, 3321, 910, 3233, 1979, 1494, 1680, 85, 2261, 1585, 2618, 146, 238, 582, 2553, 2324, 1523, 1388, 2959, 3411, 563, 737, 3010, 3106, 1447, 1834, 730, 2340, 1745, 3378, 1835, 1716, 1289, 2992, 738, 490, 766, 1298, 1693, 1922, 3526, 3293, 3253, 1909, 3241, 2823, 2746, 3566, 3340]. models [1740, 91, 3602, 1339, 1777, 2007, 1248, 387, 31, 913, 2095, 3239, 3420, 1646, 3231, 1649, 503, 669, 65, 1225, 1994, 2719, 2852, 3583, 3044, 1872, 1821, 867, 2023, 2920, 3449, 682, 1602, 1843, 2931, 2750, 3493, 523, 531, 3173, 1210, 2408, 893, 3051, 1532, 2792, 3581, 3094, 3029, 2831, 1354, 1392, 549, 1366, 259, 1251, 2743, 3336, 755, 1972, 385, 1063, 319, 3071, 3100, 1373, 2262, 2452, 1108, 1329, 1020, 2806, 1592, 1574, 2150, 1184, 3368, 722, 1240, 1148, 1201, 820, 2748, 210, 307, 446, 3223, 1543, 3046, 1633, 2005, 1622, 3324, 1235, 3013, 221, 651, 1416, 2510, 1751, 3351, 2177, 3318]. models [1524, 1110, 86, 331, 1123, 3126, 3167, 2531, 3104, 2652, 3607, 652, 2052, 1290, 1071, 2032, 2479, 1773, 3240, 2938, 1739, 82, 2337, 736, 126, 677, 263, 2139, 3475, 1989, 2708, 3393, 1952, 1782, 839, 171, 3269, 3520, 2472, 1744, 1118, 158, 230, 2533, 2996, 1583, 2450, 894, 1288, 1632, 49, 2487, 775, 1227, 1529, 3214, 3492, 2867, 3557, 2216, 1435, 1407, 525, 121, 2904, 708, 43, 1981, 1845, 3469, 2024, 2713, 2666, 1875, 1453, 2736, 1720, 2305, 1966, 2828, 2393, 404, 2905, 3552, 2722, 917, 2610, 2623, 658, 891, 695, 1175, 1312, 1772, 1617, 2496, 2929, 1577, 2830]. models [1155, 801, 2369, 1264, 3618, 242, 2068, 926, 1414, 1474, 1855, 3078, 683, 3281, 1243, 576, 865, 3116, 1083, 1727, 2046, 2682, 2793, 2555, 1480, 1182, 3064, 2231, 550, 1162, 1737, 1508, 2946, 589, 3376, 334, 2435, 383, 2599, 1057, 2725]. moderated [3052]. moderation [1651, 2300]. moderators [1151]. modern [3612, 690, 3431]. modest [1982, 1634]. Modestly [3519]. modification [2383, 2454, 685, 2848, 2061, 490, 1054]. modifications [515]. Modified [3345, 441, 3403, 45, 388, 349, 2504, 2505, 17, 277, 2399, 1641]. modify [1384]. modifying [421, 1099, 1101, 1100, 464]. Mogg [150]. molecular [2006, 735, 823]. molecularly [1848, 2432, 2416, 3456]. Molenberghs [2846, 506]. moment [2468]. momentary [3310, 910, 1835, 896, 3082]. moments [2310]. momentum [2075]. Monferrato [967]. monitor [1625]. Monitoring [574, 2570, 3444, 189, 244, 477, 777, 3428, 180, 1778, 905, 2547, 3176, 2577, 523, 3278, 192, 878, 1939, 50, 1868, 2008, 3305, 625, 3611, 1011, 3088]. Monographs [460]. monotone [2589, 1528, 1295, 302, 919, 3464, 3365, 230, 393, 1078, 2713, 3384, 157]. Monotonic [2610, 2970]. monotonicity [181]. Monte [2796, 737, 1149, 2769, 1622, 1989, 2211, 2104]. Montserrat [411]. mood [1246, 782]. moral [198]. morbidity [904, 1772]. Morning [3054]. Morris [584, 3]. mortal [230]. mortality [3476, 996, 1627, 1315, 81, 914, 1423, 981, 1468, 240, 1921, 622, 3017, 1573, 1020, 443, 1413, 3317, 354, 611, 590, 2875, 1432]. Morten [3059]. most [710]. mother [916, 1879]. mother-to-child [916, 1879]. motion [2216, 537]. motor [3372]. mover [2793, 2555]. mover-stayer [2793, 2555]. Moving [2053, 3330, 1198, 50]. MPlus [2025]. MR [2918]. MR-Egger [2918]. MRC [321]. MRI [952, 1505, 3153, 1943]. MS [1380]. MSM [1480]. MTI [2148]. Much [3599, 1070]. Muchene [2846]. Multi [1531, 1995, 2987, 1246, 1600, 3634, 1562, 121, 2545, 1398, 1637, 2087, 2787, 122, 897, 434, 239, 548, 914, 2419, 946, 2147, 516, 2097, 861, 1803, 2037, 1388, 1092, 2133, 2604, 1932, 730, 3539, 1969, 1636, 1469, 1049, 3046, 340, 331, 1133, 2139, 1713, 2472, 521, 1501, 580, 518, 1617, 985, 1253, 1563, 1679, 1757, 1508, 2599]. Multi-arm [2987, 2545, 2787, 434, 239, 1092, 2133, 2604, 1636, 1713, 985, 1253, 1563]. multi-armed [580]. multi-category [2097]. multi-center [340]. multi-centre [1398, 516]. multi-dimensionally [1501]. Multi-disease [1995]. multi-endpoint [3046]. Multi-Ethnic [1600]. multi-layered [1679]. multi-level [897, 1508]. multi-outcome [1932, 1617]. multi-parameter [946]. multi-period [1469]. multi-platform [1803]. Multi-profile [1246]. Multi-regional [1562, 2147, 1133, 518]. multi-response [1757]. Multi-stage [3634, 914, 861, 2037, 1092, 2133, 2604, 521, 985, 1563]. Multi-state [1531, 121, 1637, 2087, 122, 2419, 1388, 730, 1969, 1049, 331, 2139, 2472, 2599]. multi-type [3539]. multi-valued [548]. multiarm [3342, 3484, 3448, 3145]. multicase [1427]. multicenter [1778, 2935, 1373, 1690]. multicentre [1618, 1091]. multicomponent [2892]. multicriteria [455]. Multidimensional [226, 382, 2794, 2950, 925, 2140]. multidistance [1896]. multidose [388]. Multilevel [1674, 738, 3052, 3338, 1368, 2795, 2630, 2978, 2349, 2959, 3389, 1716, 1250, 1540, 3134, 3249, 893, 319, 3071, 1383, 3317, 52, 343]. multimarker [2861]. multimorbidity [2038]. multinomial [3412, 1105, 1585, 645, 538, 190, 2481, 3602, 965, 1823, 2321, 3376]. multiparametric [3153]. multiplatform [3121]. Multiple [2477, 318, 565, 2382, 1544, 1138, 1566, 1388, 3182, 783, 1729, 2607, 1448, 449, 2512, 472, 464, 1410, 2244, 2338, 1359, 344, 779, 2035, 2757, 1720, 868, 356, 2199, 3292, 786, 676, 178, 3501, 1561, 815, 1452, 2973, 3035, 3345, 3371, 2138, 3199, 2342, 97, 3279, 463, 2180, 1333, 2249, 2768, 1688, 2365, 180, 2209, 3361, 1535, 298, 39, 2891, 3178, 2307, 1571, 1315, 462, 1568, 2339, 2844, 1206, 697, 1090, 117, 466, 845, 2336, 384, 27, 1495, 921, 2547, 3613, 2987, 430, 1097, 1665, 3157, 1371, 1805, 1992, 139, 1990, 2341, 381, 940, 2553, 3015, 1734, 2162, 2604, 3218, 3106]. multiple [3005, 2340, 3038, 1301, 1515, 2103, 2530, 314, 2608, 1943, 2694, 1040, 1704, 2669, 2815, 269, 863, 1327, 3340, 1481, 70, 1143, 1248, 1780, 3506, 930, 1833, 2593, 3592, 3478, 1660, 3572, 1260, 2278, 600, 2984, 1428, 1707, 3362, 620, 3054, 3077, 1530, 2036, 3029, 281, 1137, 3471, 549, 1344, 1251, 3610, 1509, 2027, 2350, 1079, 217, 656, 1095, 3046, 2515, 3063, 3318, 1996, 3585, 256, 282, 2003, 3289, 2052, 1968, 1938, 451, 1005, 2532, 617, 237, 408, 3092, 3474, 2450, 1701, 2832, 354, 333, 1616, 1303, 3108, 3356, 2893, 3461, 220, 3563, 3436, 1875, 3338, 410, 3164, 1638, 1047, 3323, 293]. multiple [805, 1312, 2928, 3320, 3425, 2409, 1217, 1094, 3187, 1666, 283, 1936, 166, 409, 2068, 3645, 1732, 668, 865, 2982, 3453, 504, 2198, 1089, 2309, 1510, 2568, 497, 481, 1328, 1363, 1893, 2203, 246]. multiple-breath [3077]. multiple-class [1079]. multiple-component [3054]. multiple-disease [1363]. multiple-domain [3471]. multiple-infection [3572]. multiple-informant [1428]. multiple-level [1805]. multiple-mediator [1312]. multiple-model [3029, 1616]. Multiple-objective [1448]. multiple-outcome [1509]. multiple-treatments [237]. multiple-variable [2532]. Multiplication [3479]. Multiplication-combination [3479]. Multiplicative [1832, 2533].

Multiplicity [2890, 2894, 1807, 2039, 1454, 1088, 1379, 641, 2296]. multiply [1820, 1272, 3060]. multiply-imputed [1272]. multireader [1427, 3080, 2936]. multiregional [2285]. multiresponse [2486, 2926]. multiresponses [2745]. Multisample [3163]. Multiscale [1421]. multisource [3336]. multistage [665, 1503]. Multistate [3373, 724, 2458, 2919, 816, 1523, 663, 93, 3051, 1223, 2652, 727, 726, 2867, 634, 3149, 2600]. Multistep [3115, 3573]. multithreshold [3439]. multitype [753]. multivalued [2134]. multivariable [84, 1173, 1140, 2098, 2918, 3647, 3352, 3353, 3269]. Multivariate [2509, 1294, 542, 544, 946, 545, 541, 3015, 3156, 540, 546, 3085, 3128, 581, 893, 1896, 3126, 1968, 2337, 1617, 1109, 1214, 1898, 1462, 1839, 1816, 2571, 2839, 3055, 1254, 2384, 3522, 2371, 1135, 1209, 2011, 134, 945, 1443, 2319, 1835, 21, 1298, 807, 317, 977, 1649, 1540, 2855, 2634, 1287, 1446, 2408, 179, 3416, 662, 3372, 367, 1168, 215, 3223, 2910, 429, 2995, 2118, 1949, 451, 776, 2078, 15, 2569, 2216, 1501, 1846, 2666, 3384, 2392, 1511, 1222, 2794, 1217, 1094, 2483, 1884, 1900, 589, 966, 543]. multivariate-t [2216]. mumps [752]. muscle [1314]. muscular [962]. must [2618]. mutation [1763, 2530]. myelodysplastic [2813]. myeloid [1618]. myeloma [2003]. myocardial [1323]. myopathy [3366, 2985]. myopia [658]. myths [1115].

N [3, 1793]. Nadarajah [495]. Naive [2001, 3226]. Nam [184]. NARFCS [3099]. Nathan [584]. National [3299, 141, 1681, 2512, 371, 443, 565, 903, 3397, 3203, 143, 2757, 374, 142, 344, 1264]. nationwide [1887]. natural [1497, 1763, 2530, 130, 3574, 776, 287, 129, 133, 1498, 1801, 132, 131, 1733]. naturalistic [3276, 3275]. NCI [278]. NCS [142]. nearby [716]. nearest [1159, 817]. nearly [2194]. necessarily [2234]. necessity [3587]. necrotising [2679]. needed [2891, 1417, 88, 1464, 1419]. needing [1627]. negative [3370, 1743, 3142, 2355, 45, 1943, 3631, 2132, 2233, 532, 1606, 1545, 1429]. negotiating [475]. neighbor [817]. neighborhood [188, 417, 1104, 3337]. neighbors [1159]. Neither [1940]. neonatal [2679, 3500]. nephrology [1469]. nerve [2456, 2962, 575]. nested [1651, 1514, 1827, 838, 3522, 2443, 1711, 3561, 1301, 1702, 137, 651, 2359, 2227, 2139, 741, 1537, 2597, 3317, 1632, 1389, 52, 404]. Net [1646, 1573, 2895, 3297, 2349, 3573, 647, 702, 3639, 3654, 1645, 3653, 322, 651, 1756, 652, 2248, 395, 2899, 681]. net-based [395]. Network [3465, 1001, 1954, 1164, 1062, 2997, 1678, 3485, 2952, 2300, 1568, 2649, 3639, 3654, 2951, 1992, 3156, 1661, 2160, 1395, 3653, 3433, 3212, 1003, 1788, 1840, 2175, 2629, 3018, 1043, 2177, 1832, 1713, 3482, 3629, 3524, 3469, 3357, 1262, 1047, 2722, 1577, 3618, 1985, 2385]. Network-based [1164]. network-structured [2177]. networks [2770, 2499, 1044, 2679, 416, 1415, 1046, 1045, 3112, 3116]. Neuenschwander [72]. neural [2121, 1421, 2556, 480]. neurodegenerative [3507, 3250]. neurodevelopmental [3247]. neuroimaging [2608]. neuronal [457]. neurons [2121]. neuropathy [2456]. neuroscience [3069]. nevertheless [1867]. newer [2902]. news [3611]. NHANES [3017, 2532]. NHLBI [871]. Nicholas [1045]. Nigeria [3392]. NINDS [58]. NIS [374]. NNT [88]. No [398, 2012, 447, 460, 432, 107, 1967, 118, 520, 106, 1457, 411, 2591, 412, 446, 2725, 289, 506, 2691, 445, 3059, 1522, 567, 433, 380, 2388, 706, 2030, 3351, 631]. node [1122]. Nolen [2846]. nominal [840]. nominated [3093]. Non [723, 2, 617, 1413, 2778, 1192, 996, 442, 147, 1934, 2178, 2590, 923, 294, 2988, 1886, 186, 2349, 759, 170, 264, 3477, 635, 2649, 271, 1946, 1754, 225, 544, 946, 1471, 2149, 1995, 3296, 2067, 1944, 1948, 2421, 45, 557, 1135, 2784, 341, 2503, 438, 154, 5, 348, 1226, 863, 2166, 977, 2607, 1418, 347, 2381, 2609, 919, 437, 2325, 2574, 179, 3180, 709, 213, 3609, 261, 1565, 373, 1861, 474, 1201, 292, 78, 2762, 429, 2028, 1947, 228, 3355, 305, 1945, 1005, 601, 1178, 1426, 1953, 604, 1358, 230, 2035, 928, 60, 351, 157, 3644, 2229, 831, 69]. non [2579, 3287, 1474, 1855, 2135, 1106, 1655, 184]. non-adherence [1655]. non-commensurate [2784, 2035]. Non-compartmental [723]. non-compliance [147, 1934, 1418, 3609, 1855]. non-concurrent [2579]. non-detects [261]. non-Gaussian [996, 3296]. non-homogeneous [709, 1201]. non-identifiable [264]. non-ignorability [179, 3180]. Non-ignorable [1413, 2778, 186, 919, 604, 351, 157]. non-ignorably [294]. non-independence [1565]. Non-inferiority [2, 617, 1192, 759, 271, 2149, 1135, 341, 154, 5, 348, 1226, 863, 347, 2381, 437, 2325, 2574, 213, 474, 78, 3355, 601, 1358, 928, 1855, 184]. non-inferiority-the [1178]. non-informative [2590, 170]. non-iterative [1106]. non-linear [2349, 225, 544, 946, 1471, 1995, 977, 429, 60, 3287, 2135]. non-missing [3477]. non-monotone [919, 230, 157]. non-negative [45]. non-normal [2166, 2607, 1861, 2028]. non-overlapping [2988]. non-paired [442]. non-parametric [2178, 635, 1754, 438, 2028, 305, 1426, 831]. non-participation [557, 228]. non-perfect [1886]. non-pharmacological [1005]. non-proportional [923, 2349]. non-random [179, 213, 292, 69]. non-randomized [2649, 373, 2229, 1474]. non-recurring [2421]. non-response [1953]. non-restraint [2609]. non-specific [2503]. non-stationary [2067]. non-survival [1946, 1944, 1948, 1947, 1945]. non-zero [2762]. nonadherence [3309]. noncompliance [1323, 3213, 3063, 1939, 3315, 775]. nondetects [2392]. nonignorable [3657, 122, 1488, 2548, 1910, 1067, 29, 876, 3040, 3295]. Noninferiority [3626, 100, 2924, 487, 3309, 1919, 113, 1179, 2620, 2786]. nonisotropic [657]. noniterative [3152]. Nonlinear [3253, 1344, 3250, 266, 2045, 3121, 952, 3321, 238, 2553, 183, 3085, 1872, 1602, 3375, 722, 2782, 3084, 263, 2487, 793, 1617, 1881, 1639, 3116, 2946]. nonlinearly [2406]. nonmixture [3467]. nonnegative [1915]. nonnested [3522]. nonnormal [3384]. nonordered [3610]. Nonparametric [244, 3237, 3637, 2696, 816, 1507, 2358, 600, 1506, 3254, 1223, 2181, 843, 2318, 1169, 3645, 3575, 2076, 2063, 3147, 728, 2087, 3050, 2516, 2851, 2949, 2771, 2717, 2854, 3042, 1990, 146, 2316, 1838, 164, 608, 1729, 59, 773, 758, 2725, 792, 707, 995, 3601, 1163, 1529, 2690, 3437, 2805, 1499, 3081, 2279, 561, 3488]. nonparametrics [653]. Nonproportional [3491, 1387]. nonresponse [2548, 2830, 3586]. nonstandard [3638]. nonstationary [1243, 66]. nonsusceptibility [1317]. nonzero [3449]. nor [1940]. Norm [1457, 2969]. normal [2365, 1018, 240, 238, 355, 317, 2166, 977, 2607, 29, 1287, 1861, 651, 2028, 158, 1823, 1846, 2666, 1610]. normal/independent [260, 1287]. normality [454, 3522]. normalization [1796]. Norsk [1522]. Northern [376, 618]. Norway [3044]. Norwegian [1522, 1886]. nosocomial [1178]. not-at-random [3099]. Note [1403, 1828, 2202, 2235, 3347, 1247, 1438, 72, 1645, 1605, 2163, 3486, 1464, 1783, 9, 2619, 1152, 1696, 1604, 223, 1644, 3521, 1889, 1981, 1845, 3226, 2501, 1576]. Notes [3020, 2871, 3021]. nothing [2911, 3599, 1867]. Novel [827, 651, 2213, 1053, 1206, 1645, 3542, 2617, 1955, 2751, 192, 848, 237, 3144, 789, 3000, 1584, 2538, 3089]. Nowcasting [1201]. NRC [2330]. NRI [1647]. NRIs [2801]. nucleic [2518]. nucleotide [1140]. Null [1014]. Number [1417, 1385, 431, 1492, 1350, 318, 3236, 3486, 88, 1464, 733, 1377, 163, 1519, 3298, 2120, 2062, 1192]. numbers [3463, 1457, 419, 1696]. Numeric [2050]. numerical [2342]. nurse [81]. nursing [43]. nutrients [278]. Nutrition [2512, 344, 903, 2047]. nutritional [2048, 2236, 1289, 1557].

obesity [1452, 240, 1246, 3518]. object [3635]. object-oriented [3635]. Objective [2529, 669, 1663, 639, 638, 1731, 326, 3599, 2464, 1448, 2113]. objectives [1040]. obligation [971]. observation [321, 1841, 3466, 2069, 1225, 2244, 2652, 2104, 1741, 634, 865, 456]. Observational [1085, 994, 1789, 216, 2053, 203, 3523, 3652, 2415, 2367, 1799, 2413, 3131, 1441, 3288, 2457, 901, 1461, 2511, 3065, 3101, 1978, 1263, 619, 1440, 1419, 2414, 1868, 2827, 3160, 1917]. observations [2755, 3055, 3178, 2760, 737, 1067, 1819, 471, 373, 1273, 1426, 986, 1006, 976, 2600, 2605]. observed [2218, 2388, 2950, 2348, 2912, 908, 1084, 1442, 2479, 536]. observers [636, 402, 812]. obtained [3605]. Obuchowski [3080]. occupancy [2652]. occupation [93, 1223]. occupational [1275]. occurrence [2019, 89, 3589]. occurring [2025, 997]. ocular [547]. odds [751, 2684, 1748, 1349, 488, 1172, 1232, 1396, 1568, 1211, 1171, 2612, 1098, 100, 3010, 1316, 2381, 437, 2325, 2574, 509, 907, 2879, 1235, 1190, 480, 79, 2557, 2833, 352, 1338, 1155, 3554, 2645]. odds-rate [2645]. ODE [3116]. off [3656, 3442, 285]. offers [90]. offsets [500]. offspring [236, 141, 95]. old [872]. older [369]. omics [1781, 3463, 3348]. omnibus [1829]. onco [470]. onco-hematology [470]. oncology [1882, 2006, 1396, 1941, 680, 2021, 991, 2693, 2912, 1723, 2522, 2122, 2418, 3211, 1548, 3313, 789, 1052, 512, 2399]. One [2618, 3334, 295, 3461, 1673, 2624, 1165, 1838, 675, 1704, 2669, 885, 1549, 3231, 2981, 17, 2879, 631, 305, 324, 3228, 1860, 3627, 2375, 3282, 2392]. One-sample [3334, 1549, 305, 1860]. one-shot [1838]. One-stage [2618, 295, 1673, 2624, 3231, 3282]. One-step [3461]. one-to-one [3228]. one-way [324, 2392]. ones [3363]. Online [2953, 3656, 3442, 596, 287]. only [2469, 719, 124, 1073, 1165, 675, 252, 3001, 3012, 181]. onset [1964, 1506, 1063, 3251, 515, 1155, 316]. onsets [3399]. open [547, 2088]. open-angle [547]. open-enrollment [2088]. operating [1393, 2480, 2404, 212, 249, 1477, 1279, 600, 843, 484, 2250, 3294, 3536, 1372]. operational [1956]. opinion [1901]. opioid [3036]. OPPERA [2072]. Opportunities [2115, 1487, 2111, 2112, 2114]. opportunity [3273]. Optimal [2070, 1350, 28, 660, 3452, 1473, 1015, 3290, 2097, 1777, 1826, 3516, 2041, 78, 1812, 2010, 3393, 2195, 721, 52, 580, 3211, 743, 985, 665, 1437, 1510, 2567, 3177, 3651, 2373, 1637, 77, 584, 952, 2367, 1871, 2272, 673, 3, 1549, 2194, 3036, 3448, 3196, 3191, 1581, 1469, 656, 1142, 3545, 3440, 299, 3600, 1050, 1, 3404, 394, 1244, 2565, 3536, 691, 2799, 3624]. Optimality [572]. Optimally [2495, 1084, 1442]. optimisation [3546, 722, 2032]. Optimising [3421, 3649]. Optimization [1572, 2037, 1869, 424, 2105, 2948]. optimized [3095, 3009, 3096]. Optimizing [3344, 1795, 366, 3044, 1893, 2566]. Optimum [2925, 726, 2059, 727]. optogenetic [2121]. O’Quigley [520]. oral [1383]. Order [776, 155, 666, 696, 3180, 825, 575, 3146, 1893, 3252]. Order-restricted [776, 825, 3146]. ordered [3374, 2581, 2728, 942, 2473, 1236, 1660, 2050, 627, 1445, 644, 1129]. ordering [1848]. orderings [2250]. orders [1009]. ordinal [3657, 754, 2517, 1970, 1105, 2916, 1172, 1631, 2851, 635, 1767, 2439, 3626, 267, 2319, 3005, 961, 1051, 977, 58, 773, 1961, 2877, 2096, 1955, 509, 2015, 2790, 1235, 2353, 3052, 1436, 890, 3040, 1222, 56, 1257, 3090, 246]. ordinal-rating [246]. ordinary [1160, 2241]. organ [3563]. organs [2686]. oriented [3635]. origin [163]. original [2179]. orphan [1850]. Oslo [1522]. osteoporosis [3256, 693, 691]. Osteoporotic [1836]. osteosarcoma [939]. other [946, 3245, 3118, 2478, 1016, 91, 678, 1137, 3500, 2699, 2280, 1499]. outbreak [3620, 3656, 3442, 419, 1095]. outbreaks [2276]. Outcome [1689, 826, 728, 1970, 761, 3604, 862, 2235, 2851, 25, 2000, 2838, 1841, 3637, 2862, 670, 1650, 3571, 888, 2869, 415, 1491, 3156, 1932, 961, 1838, 807, 1151, 3420, 1837, 2169, 2281, 3086, 3196, 1589, 2866, 3191, 3516, 1699, 2634, 3189, 1840, 1760, 3610, 1509, 2586, 3234, 1075, 3098, 1235, 1853, 2625, 1207, 2227, 3550, 2754, 3052, 958, 741, 1114, 1413, 24, 3482, 158, 296, 1501, 2773, 1741, 352, 1937, 1749, 1312, 1617, 2496, 3489, 167, 2620, 606, 3287, 993, 3083, 625, 1887, 334]. Outcome-adaptive [1689]. outcome-based [1970]. Outcome-dependent [826, 3604, 2634, 3234, 3098, 3550, 1741, 3083]. outcome-weighted [3191]. Outcomes [401, 1085, 994, 1628, 2484, 3657, 587, 1496, 2070, 3067, 3568, 2764, 147, 3199, 2955, 2218, 35, 1916, 1105, 1514, 1294, 763, 315, 3307, 840, 3150, 155, 1337, 1839, 2083, 3140, 3455, 699, 1466, 3527, 3477, 3447, 2382, 2225, 1208, 1988, 1300, 1568, 849, 1518, 2700, 466, 785, 2264, 3300, 1905, 3587, 574, 2398, 3296, 2966, 2439, 2421, 3559, 1141, 1965, 1552, 2784, 3397, 1494, 680, 3446, 2783, 703, 2341, 381, 3135, 2011, 3106, 538, 3005, 2044, 1091, 2546, 1835, 766, 1702, 2958, 3293, 1626, 2694, 3200, 3231, 1649, 1656, 773, 795, 1811, 2381, 2593, 2837, 2981, 3592, 1961, 3636, 2274, 916]. outcomes [2226, 1177, 2737, 1530, 281, 3314, 1467, 2447, 3375, 1574, 564, 3102, 1919, 3025, 2088, 2005, 3013, 1070, 595, 1487, 1406, 2698, 357, 3647, 3352, 3353, 451, 2353, 1725, 230, 2375, 3390, 295, 3039, 843, 3524, 2659, 1625, 3040, 3338, 2335, 1511, 2794, 1577, 665, 1217, 1094, 1997, 339, 2943, 865, 3490, 1083, 2622, 3305, 3301, 3149, 497, 3292, 2906, 2714, 1508, 2605, 2085, 3088, 966, 2712, 2574]. Outlier [1881]. outliers [2370, 2678]. outlying [1985]. output [2292]. outputation [2244]. outs [3628]. outstanding [972]. ovarian [232]. over-dispersed [2549, 2024]. over-dispersion [324]. Overall [3218, 1466, 206, 1018, 1969, 3212, 1769, 2050, 3556, 1179, 3330, 428, 2496, 3313]. overdispersed [1296, 1269, 2971, 2959, 1716, 3441, 3071, 1367, 1270, 2637]. Overdispersion [3602, 755, 2046, 1182, 701]. overdose [471, 1326, 1672, 2741]. overestimated [419]. Overlap [3021, 3221, 3020, 2871]. overlapping [2988, 138, 2025, 3638]. overrunning [1941]. overview [2504, 3557]. Oxford [445].

P [586, 3547, 3649]. PA [58]. Pablo [3021]. packages [550]. paediatric [1469, 3492]. PAF [89]. Page [2080, 2093, 2491, 2507, 2524, 2542, 2109, 2125, 2144, 2157, 2171, 2192, 2207, 2222, 2238, 2253, 2268, 2283, 2298, 2312, 2328, 2345, 2362, 2378, 2395, 2411, 2429, 2444, 2461, 2475, 2079, 2282, 2297, 2311, 2327, 2344, 2361, 2377, 2394, 2410, 2428, 2445, 2460, 2474, 2092, 2490, 2506, 2523, 2541, 2108, 2124, 2143, 2156, 2170, 2191, 2206, 2221, 2237, 2252, 2267, 2559, 3569]. Pages [2012, 398, 447, 460, 432, 118, 520, 106, 1457, 411, 2591, 412, 446, 2725, 289, 506, 2691, 445, 3059, 1522, 567, 433, 380]. pain [761, 3036]. painful [1655]. pair [1858, 2465, 2401, 3206, 375, 3626, 2869, 645, 2, 13, 184, 1587]. pair-matched [2869]. pair-matching [1858, 2465, 2401]. paired [3479, 442, 1323, 1603, 940, 907, 3481, 173, 2686, 1614, 2690, 890, 16, 782, 1813, 1106, 2650, 2605]. paired-sample [940]. pairs [2738, 3135, 610, 309, 2104]. Pairwise [2961, 1236, 1062, 315, 2511, 1840, 915, 1937, 3644]. palliative [1111, 2917]. Palys [969]. Pamela [2591]. pancreatic [789]. pandemic [905]. Panel [3117, 2051, 2164, 887, 1082, 1339, 1225, 684, 1966, 2039, 2793, 2680]. panoramic [659]. paper [853, 3020]. papers [3498, 2333]. papillomavirus [1149, 517]. Paradigms [877, 2197]. paradox [2896]. parallel [1208, 1805, 2999, 572, 1331, 869, 1037, 1183, 893, 1075, 1953, 3039, 632, 2766]. parallel-group [1331, 1183]. Parameter [2072, 3299, 3339, 3310, 3204, 3236, 946, 2404, 324, 1520, 1242]. parameterisation [756]. Parameterization [331]. parameters [2876, 1463, 3080, 723, 3248, 1067, 2821, 2056, 671, 1676, 3537, 1727]. Parametric [2919, 635, 387, 503, 1386, 2178, 1731, 653, 2534, 998, 1779, 1674, 1754, 1802, 3623, 267, 438, 280, 1151, 2663, 608, 1187, 1446, 1210, 1297, 2940, 215, 820, 2028, 2084, 1524, 305, 2032, 1426, 2441, 2189, 295, 2024, 410, 831, 794, 1480, 3535, 2399, 2673]. parenchymal [2678]. Parent [95, 1259]. parent-child [1259]. parental [346]. parents [609]. Pareto [128]. Parkinson [992, 1707, 1079]. Parsimonious [1315, 1288]. Part [3647, 3353, 2870, 1924, 1540, 2981, 3449, 1750, 458, 2504, 2505, 3352].

Partial [1820, 1007, 1310, 1055, 1848, 2443, 136, 1390, 795, 3222, 2043, 1255, 2806, 1933, 3162, 1009, 3024, 2982, 2279, 943, 739]. partial-likelihood [2043]. partially [2728, 982, 1241, 2022, 3638, 503, 1699, 1308, 2531, 2227, 1445, 1849, 1757, 644, 3007, 2695, 1192]. Participant [3524, 1333, 2624, 1674, 2300, 846, 1234, 2398, 2618, 1920, 3038, 2815, 3231, 1359, 1207, 924, 2035, 3557, 1997, 3287]. participants [1350, 987, 743, 3421, 3649]. participation [557, 228]. particle [3544]. particulate [2665]. partition [2204, 805]. Partitioned [3411]. Partitioning [3476, 553, 3045, 1078, 3049, 3635]. partly [1528, 2102]. Partnership [1085, 994]. partnerships [390]. past [857]. path [583, 2065]. pathogen [2970]. pathogens [1601]. pathology [3179]. paths [3155]. Pathway [1121, 3108, 766, 717]. pathways [203, 2538]. Patient [883, 401, 1916, 2662, 43, 2997, 1398, 3517, 2770, 1612, 1814, 947, 1073, 1121, 2407, 376, 1965, 136, 1680, 1992, 1711, 1331, 2042, 1298, 2957, 795, 553, 3636, 769, 3045, 1687, 3037, 2088, 515, 1512, 329, 1005, 2099, 2722, 2400, 2646]. Patient-centered [883]. patient-level [947, 2957]. Patient-reported [1916, 1965, 795]. patient-specific [1121]. patients [676, 2819, 2389, 114, 128, 2277, 3109, 740, 1665, 3364, 1276, 1943, 2694, 3057, 503, 3249, 511, 192, 3113, 1744, 2248, 3317, 2216, 486, 232, 1096]. Pattern [2777, 158, 3657, 2862, 804, 3364, 1874, 1745, 1961, 1489, 1843, 123, 3040, 1175, 2400]. Pattern-mixture [158, 2862, 1745, 1489, 1843]. patterns [2456, 1197, 233, 2371, 440, 1376, 915, 2038, 1751, 1785, 3179]. Paule [3115]. peaks [3399]. peanut [3201]. Peer [2368, 2077]. penalised [731, 2186]. Penalization [2016]. Penalized [1025, 1905, 397, 3478, 2931, 3324, 810, 754, 2247, 1507, 2969, 867, 3336, 2806, 2472, 493, 2074, 2967, 2980, 3091, 327, 2712]. penalties [1416]. penalty [2305]. Pencina [647]. Pennsylvania [107]. people [55, 1186, 1519]. per-clone [1224]. per-family [3506]. per-individual [1224]. per-protocol [213]. percent [2781, 906]. percentile [987, 937, 59]. perfect [1886, 2443, 2162]. perform [134, 2160]. Performance [728, 920, 2584, 813, 3122, 292, 2153, 166, 996, 216, 1053, 1212, 3067, 1146, 1637, 2182, 3620, 2434, 227, 2700, 1486, 139, 1798, 1277, 2901, 313, 1950, 1811, 3044, 1202, 3541, 3493, 1255, 3101, 2585, 1117, 679, 1883, 2024, 2749, 2099, 441, 3270, 809, 409, 1584, 3376, 408]. Period [1928, 2965, 1468, 1447, 1730, 1469, 2748, 126, 310, 115]. period-cohort [1447]. periodogram [2309]. periodontal [260, 2721, 1668]. periods [3584, 2609, 3107]. Permutation [1618, 3607, 14, 2627, 402, 812, 1738, 1426, 1742, 1935, 2766]. permutation-based [402, 812]. permuted [2655, 2265, 772, 1776, 2234]. permuted-block [2655]. persistence [280, 517]. person [2512, 2485, 3537]. person-time [2485]. personal [1629]. personalised [2174]. Personalized [427, 1782, 3630, 3026, 2194, 1950, 2693, 2642, 3625]. perspective [398, 2561, 2091, 3327, 2611, 2334, 2330, 3578, 1646, 1101, 3493, 2836, 1100, 2333, 2563, 2562, 2564]. Perspectives [521]. Perualila [2846]. pesticide [1176]. PET [2646]. Peter [411, 1592]. Peters [1300, 1833]. Petersen [2669, 1704]. Peto [1748]. Petter [3059, 1522]. pFDR [1415]. pFNR [1415]. PH [3060, 307]. Ph.D. [900]. pharmacoepidemiology [2176]. Pharmacogenetic [396]. pharmacogenomic [1356]. pharmacogenomics [412]. pharmacokinetic [2876, 1942, 3571, 723, 1356, 722, 1638, 2926]. pharmacokinetics [2017]. pharmacological [1005]. Pharmacovigilance [33]. Phase [866, 1549, 3017, 1548, 789, 3546, 1595, 1806, 477, 1688, 1862, 771, 239, 2383, 2581, 2728, 1923, 2566, 798, 1500, 1396, 1941, 1908, 466, 3473, 108, 1473, 2407, 2732, 1910, 1795, 990, 680, 1334, 2021, 706, 1324, 514, 2576, 303, 799, 1216, 2105, 991, 2633, 790, 2912, 2863, 1131, 2660, 2808, 111, 1907, 3073, 1878, 110, 2617, 1998, 75, 511, 1687, 474, 3183, 2582, 1139, 2578, 515, 1039, 3579, 1790, 1793, 2122, 2417, 109, 148, 2432, 470, 2195, 2807, 1326, 3533, 1609, 98, 721, 2014, 757, 3211, 1672, 2583, 468, 399, 1558, 1792, 3346, 3450, 3270, 1343, 3137, 2579]. phase [2741, 181, 499, 1673, 2706, 1437, 1052, 1791, 512, 2567, 2565, 781, 1690, 950, 2799, 3089, 3624, 537, 2577, 2302, 1445]. phase-dependent [537]. phase-type [2407, 757]. phenotype [3220, 397, 1800, 3138, 2723]. phenotypes [2907, 1332, 2440]. phenotypic [2907]. phenotyping [3268]. phone [935]. PHow [943]. phylogenetics [2949]. Physical [3126, 1246, 2855, 2038]. Physician [1202, 1262]. physicians [1276]. picking [3229]. Piecewise [746, 239, 1408, 1056, 1911]. piecewise-constant [1056]. Pigeot [602]. pigmentosa [3113]. pilot [786, 2194, 721, 2963]. pitfalls [2111, 3613, 2115, 2112, 2114]. Placebo [1953, 341, 1226, 1489, 2041, 601, 115, 1855]. placebo-based [1489]. placebo-controlled [115]. placebo-treatment [2041]. placed [1192]. placement [43]. Plan [1550]. planned [476, 2405]. Planning [5, 348, 311, 2175, 1268, 1000, 1253, 3033, 1064, 3290, 1406, 2620, 2786, 347]. platform [1803, 3343, 2416]. Platt [1265]. plausible [1307]. plea [3551, 3549, 3548, 1775]. pleiotropy [2688, 2000, 2918, 2913]. Pleural [967]. PLNseq [1900]. plot [2287, 3639, 3654, 3653, 2400]. plots [1630, 1662, 1485]. PM [2321]. pneumonia [1178, 1952]. pneumoniae [753]. Pocock [649]. Point [1874, 3151, 2849, 2633, 930, 2458, 949, 1894, 2019, 3364, 3309, 2608, 844, 1826, 2121, 1366, 2262, 285, 2432, 1755, 1083, 383]. point-process [2121]. points [2427, 3034, 1199, 1097, 1481, 3057, 1968, 2892, 3250, 191]. points-based [2427, 3034]. pointwise [2317]. Poisson [628, 3354, 996, 2978, 1103, 2965, 2308, 71, 3161, 479, 330, 2183, 3612, 3194, 1769, 2824, 1020, 1374, 3408, 611, 708, 32, 1189, 1900, 2930]. Poisson-gamma [2824]. policies [2513]. Policy [200, 196]. pollutants [3015, 3319]. pollution [519, 913, 268]. Polya [1346, 350]. polymorphism [1140]. polymorphisms [868]. polynomial [1411, 2027, 2391]. polynomials [1173, 911, 1954, 2587, 1240, 125, 747]. polytomous [840, 102, 1965, 3537, 840]. pool [2049]. Pooled [902, 2518, 372, 903, 2215, 2717, 1107, 1758, 2731, 1972, 62]. Pooled-testing [902]. pooled-unpooled [62]. Pooling [659, 899, 1241, 830, 1760, 3102, 2409, 828, 829, 3210]. pools [3205, 2536]. poor [1982, 1634]. popular [3590]. Population [3318, 2636, 3459, 3322, 1555, 1612, 369, 1423, 2067, 2956, 1468, 990, 2747, 2090, 2817, 2972, 89, 678, 3634, 1723, 2522, 1356, 64, 2352, 3126, 1485, 1715, 393, 3533, 3138, 3605, 3286, 468, 399, 2514, 3024, 821, 3525]. population-attributable [3525]. population-based [2636, 3459, 3322, 1423, 2747, 678, 64, 1485]. Population-calibrated [3318]. population-level [2090]. populations [457, 902, 942, 2225, 1711, 3389, 1907]. portfolios [1139]. PoS [3211]. Positing [1972]. positive [1913, 2120, 3428, 2216]. positives [124, 3012]. positivity [3172]. Possible [755]. possibly [2218, 2991]. post [141, 1188, 1195, 1331, 2022, 3256, 2060, 2211, 2942, 2646, 2085]. post-analytic [2060]. post-hoc [1331]. post-infection [2085]. post-market [2211, 2942]. post-stratification [1195, 3256]. post-surgical [1188]. post-therapy [2646]. Posterior [469, 808, 640, 2501]. postnatal [2853]. postoperative [1772]. postoutbreak [3403]. postrandomization [3589]. postselection [3302]. Potential [543, 541, 540, 546, 856, 3527, 2006, 3487, 1905, 2570, 3118, 545, 179, 396, 3234, 1357, 1609]. potentially [2533, 1917]. potentials [421]. Power [116, 23, 1466, 971, 2403, 1965, 2205, 593, 1016, 3308, 1843, 1819, 3269, 2690, 374, 3285, 3345, 2210, 463, 862, 527, 275, 3385, 448, 911, 72, 740, 3438, 706, 1685, 2057, 572, 2550, 538, 1722, 1730, 2539, 1327, 3170, 522, 281, 788, 9, 1885, 1004, 889, 2337, 1000, 1735, 839, 2675, 3119, 627, 2881, 1859, 3033, 2402, 2286, 2540, 2373, 2922]. power-integrated [3438]. Power/sample [2403]. Powerful [3220, 1642, 1535, 3350, 710]. pp [432]. PQL [155]. Practicable [1113]. Practical [2882, 3329, 77, 3274, 467, 799, 515, 398, 2182, 1942, 2291, 614, 649, 613, 143, 3208, 877, 2416, 2494]. practice [2053, 1146, 427, 1938, 3099, 2099, 356, 2798]. practices [2159, 2295]. pravastatin [1632]. pre [2465, 141, 2149, 476, 174, 1099, 1101, 1100, 2646]. pre-clinical [174]. pre-planned [476]. pre-specification [2465]. pre-specified [2149, 1099, 1101, 1100]. pre-therapy [2646]. Precision [2167, 2210, 1977, 2974, 2175, 2986, 2631, 3064, 957]. predefined [2623]. prediabetes [3203]. predict [2973, 3035, 2287, 1573, 3113]. predicted [1662]. Predicting [1964, 678, 187, 965, 1609, 704, 1293, 2712, 654, 1398, 716, 1885]. Prediction [3336, 3441, 2137, 2198, 3276, 754, 1682, 1851, 2898, 1781, 323, 3043, 1021, 846, 1234, 1571, 1158, 3544, 3642, 953, 1167, 1630, 2760, 1980, 1064, 3388, 2644, 1585, 3171, 146, 2273, 992, 2481, 1958, 3332, 31, 622, 3031, 3592, 1159, 3493, 3029, 1392, 3515, 1574, 1157, 3643, 1256, 2585, 2186, 1117, 86, 2932, 3647, 3352, 3353, 2597, 2260, 1583, 894, 3108, 3214, 3557, 386, 2904, 428, 1185, 2905, 3190, 2001, 3078, 3053, 67, 3376, 2435, 2117]. predictions [2664, 1542, 1229, 3014, 1468, 3026, 3462, 3085, 2724, 793, 3280]. Predictive [1053, 2767, 706, 429, 3386, 1857, 3088, 2641, 3502, 777, 1862, 1814, 841, 842, 596, 1311, 2789, 2467, 3438, 1334, 2196, 1077, 2719, 3633, 3073, 231, 148, 2986, 879, 3119, 983, 2008, 1558, 3148, 1584, 2501, 3376]. predictive-signature [1814]. predictor [1689, 3493, 1435, 2135]. predictors [2423, 3573, 2434, 1920, 2274, 261, 2177, 824, 2260, 251, 1584, 2231, 966]. predominantly [325]. Preface [2947, 2173, 74, 974, 718, 358, 870, 687]. preference [3150, 1870, 743]. preference-based [1870]. preferences [2159, 3421, 3649]. preferential [1684]. pregnancy [2853, 893, 3251, 3129, 2712]. premenopausal [232]. prenatal [793]. PrEP [2954]. Prescott [3648]. prescription [376]. presence [3451, 2427, 3034, 1541, 335, 1844, 3445, 3495, 2744, 3178, 3347, 2813, 3255, 1770, 1873, 2777, 1630, 384, 1967, 270, 2421, 2527, 2503, 378, 1056, 387, 1316, 2816, 2607, 1418, 3619, 1582, 1819, 213, 3610, 1383, 3183, 292, 1305, 3251, 3643, 3063, 274, 223, 2242, 3132, 121, 1801, 2482, 1009, 2058, 821, 1567, 2326, 3032]. present [3547, 3584, 1599]. presenting [3212]. preserve [3087]. Preserving [698, 2370, 2740, 2672]. President [1550]. prespecified [341, 601, 3165]. Press [432, 506, 445]. Pressler [1440]. pressure [3391, 893, 3054, 396, 3030, 1070, 3129, 532, 2867, 2904]. pressures [2530]. preterm [3074, 1162]. Prevalence [2557, 2833, 3012, 1721, 3476, 1160, 124, 1925, 2196, 2154, 168, 2587, 3203, 2030, 642, 3102, 3256, 229, 1432, 2409, 2058, 821]. prevalence-based [2154]. prevalences [3087]. prevalent [2820, 453, 603, 3406, 2654]. prevention [1678, 2308, 1829, 108, 3571, 1015, 373, 848, 1632, 3640]. preventive [845]. previous [2973, 3035, 3109]. Price [398, 2012, 447, 460, 432, 118, 520, 106, 1457, 2591, 412, 446, 2725, 289, 506, 2691, 445, 3059, 1522, 567, 433, 380]. PRIM [1814]. primary [1608, 2891, 547, 1957, 2987, 1061, 1194, 207, 848, 2314, 1632, 3533, 220, 796, 797, 3125, 1089]. prime [1302]. Principal [834, 1817, 2809, 1752, 2278, 916, 1335, 3589, 1142, 2153, 1273, 714, 2859, 1801, 2775]. principle [1891]. Principles [851, 571, 856, 2009, 1485, 852, 854]. Prior [2511, 2213, 3345, 2590, 72, 3402, 2057, 510, 9, 595, 631, 2078, 741, 3377, 1170]. Prioritized [3455, 315, 2737, 1487, 3644]. priors [170, 921, 2016, 2780, 989, 1834, 756, 671, 2537]. prisoners [2089]. privacy [2013]. Probabilistic [2422, 2811, 1657, 982, 3567]. probabilities [431, 2701, 2178, 2924, 2360, 1921, 965, 1878, 2652, 681]. Probability [460, 1340, 2120, 2053, 2551, 1321, 1319, 2354, 3027, 36, 1325, 2420, 635, 3023, 1799, 673, 1334, 2584, 1032, 1809, 2162, 2707, 3526, 2075, 91, 33, 3073, 3449, 1159, 2591, 3195, 143, 1320, 2441, 3386, 3393, 1444, 2942, 1389, 1823, 1558, 3450, 1459, 1694, 3367, 3611, 2029, 2745, 3088, 1546, 1733, 3211]. probability-sensitive [36]. probit [3414, 2439, 669, 99, 2321, 2453]. problem [3095, 1710, 315, 1822, 3009, 3096, 1677, 193, 3601, 1718, 2333, 3543]. Problems [3617, 2964, 1586, 755, 285, 1079, 3365, 2500, 3295]. procedural [3217]. procedure [2059, 2307, 819, 3097, 2082, 1516, 1660, 1284, 662, 656, 1367, 2032, 125, 3520, 251, 796, 797, 3152, 3099, 1185, 2905, 181, 1503, 3622, 1089, 2391]. procedures [3382, 2148, 902, 949, 463, 462, 1905, 2761, 2888, 2002, 1808, 1040, 269, 3506, 1137, 1783, 2071, 3145, 617, 393, 2749, 428, 258, 499].

Proceedings [141, 107]. process [1901, 2458, 2201, 1627, 189, 2180, 122, 1351, 2471, 302, 2019, 1979, 1987, 267, 556, 2121, 2824, 2350, 226, 382, 1010, 591, 2118, 2556, 894, 2814, 1831, 2555, 1887, 537]. processes [120, 898, 2419, 2498, 563, 3466, 3597, 3194, 2977, 1376, 893, 3234, 3585, 354, 2065, 162, 2706, 2600]. processing [1254, 694]. product [1878, 1644]. products [1027, 1028, 1029, 1035, 108, 1037, 1033]. Professor [3505, 1380, 3504]. Profile [1457, 958, 1550, 1371, 1246, 297, 1614, 2698]. profiles [236, 1121, 2023, 3139, 2535]. profiling [1667]. prognosis [1323, 620, 1354, 1787]. prognostic [731, 1140, 1977, 2434, 2098, 3522, 300, 2612, 888, 2480, 85, 224, 3364, 1652, 3302, 450, 3401, 1765, 652, 1968, 3120]. Program [684, 3299, 1521, 590, 1253]. programs [1597, 2105, 848, 3012]. progression [266, 2758, 122, 172, 3159, 1547, 1018, 3402, 1475, 2481, 1969, 3592, 1998, 3556, 2452, 664, 3366, 1290, 392, 2867, 2418, 3250, 121, 3144, 3313, 1847, 2020, 3072]. progression-free [172, 1018, 1969, 3556, 3313, 1847, 2020, 3072]. progressions [2419]. progressive [2701, 1449, 1836]. Project [3499]. projection [1149, 631]. projections [1730, 622]. prolongation [2841]. promise [543, 197, 545, 541, 540, 546]. promises [141]. promising [614, 649, 1092, 613, 615, 3436]. promote [3547]. Promoting [2601]. Promotion [3050]. prone [2421, 1328]. proof [3155, 107, 80, 3435, 474, 3289, 1128]. proof-of-concept [3155, 3289, 1128]. Propensity [2263, 3316, 1434, 1249, 1656, 1250, 2258, 3382, 216, 442, 1212, 1484, 1709, 1496, 2053, 3067, 2964, 1375, 697, 2700, 661, 3613, 1409, 1441, 1652, 40, 813, 901, 1830, 3065, 2726, 3189, 1251, 385, 3172, 1842, 1263, 3365, 2495, 1440, 1288, 79, 304, 1691, 3284, 765, 1459, 3174, 1727, 2386]. Propensity-score [3316, 216, 442, 1709]. propensity-score-matched [3284]. proper [1941, 591]. Properties [3199, 3297, 2655, 952, 1137, 2897, 2895, 801]. prophylactic [1662]. proportion [495, 2388, 1668, 149, 2165, 1640, 496, 2228, 1714, 1114, 1527, 3563, 3007]. Proportional [90, 3509, 509, 1235, 126, 3091, 955, 1087, 2477, 2589, 923, 294, 1483, 2349, 3374, 1008, 1396, 1171, 3623, 909, 1097, 2621, 3157, 1433, 2436, 3596, 1734, 2914, 1360, 1327, 3406, 1341, 2931, 3195, 2879, 2864, 1071, 535, 1156, 2623, 956, 1975, 1155, 801, 1974, 498, 1451, 2695, 1286, 1578, 520]. Proportional-odds [1235]. proportionality [2168]. proportions [728, 216, 1827, 1603, 3576, 2981, 1569, 3494, 1614, 1178, 617, 16, 1106]. proposal [1982, 2040, 1634, 2535]. proposed [3234]. Proschan [1982, 2539, 2591]. prospect [2628]. prospective [1864, 1866, 2307, 1623, 808, 1865, 3398, 1598, 2251]. prospectively [2405]. prostate [1247, 2419, 3153, 3558, 591, 2120, 60, 3467, 2775, 1422]. prostate-specific [591]. protecting [2013]. protection [142, 1912]. protective [960]. protein [54]. protein-DNA [54]. proteomics [2199]. protocol [213]. protocols [3280]. prototypes [646]. provided [3462]. provider [3315]. providers [1798, 2515, 1856, 2639]. provision [3434]. proximal [3559]. Proximity [2791, 3337]. proxy [158, 1718]. PSA [60]. Pseudo [1517, 2569, 2755, 1230, 1237, 2537, 157, 2185]. pseudo-likelihood [157]. pseudo-observations [2755]. Pseudo-value [1517, 2185]. pseudo-values [1230, 1237]. pseudolikelihood [1105, 1816]. pseudoscore [2950]. Pseudosibship [609]. PSICA [3567]. psoriatic [1049, 354, 1966, 2555]. psychiatric [1612, 1431]. psychotropic [901]. public [911, 105, 427, 2350, 2013]. Publication [1991, 1621, 2960, 290, 1815, 1003, 1788, 2675]. Published [287, 846, 1571, 3656]. publishing [999]. pulsatile [1342]. pulses [2747]. purposes [282]. pursuit [3480].

Q [2277, 3]. Q-learning [2277]. Qing [584]. QT [2841, 1183, 82]. QT/QTc [1183]. QTc [1183]. Qu [208]. quadratic [239, 703, 959, 809, 1242]. quadrature [1674]. qualified [742]. Qualitative [1420, 51, 936, 1513, 26]. quality [3008, 1375, 235, 3079, 1597, 738, 1316, 1111, 2917, 919, 1038, 1202, 3154, 3464, 2552, 3600, 380, 2695]. quality-of-life [919]. quantal [2697]. Quantifying [3178, 2434, 3587, 945, 3591, 3462, 3194, 3212, 371, 2629, 760, 4, 1599, 2668, 2280, 3313, 3148, 3090, 3625, 1901, 3538, 652, 1594, 208, 3655]. Quantile [2868, 3093, 3464, 3047, 35, 262, 1873, 2553, 1504, 1220, 59, 3583, 1973, 792, 1773, 1364, 2803, 3117]. quantiles [1743, 608, 539, 2230]. quantitation [986]. quantitative [3067, 621, 2667, 3547, 1327, 2166, 812, 1501, 1332, 710, 2203, 3210]. quantities [2178, 986, 532]. quasi [1206, 3596, 1644, 576, 747]. quasi-continuous [1206]. quasi-least [1644]. quasi-likelihood [3596, 747]. Quermass [1979]. Quermass-interaction [1979]. question [1474]. questionnaire [423]. questionnaires [501]. questions [1697].

R [432, 647, 649, 150, 3, 1865, 194, 2725, 1793, 1440, 1265, 749, 3651, 3513, 3529, 795, 2376, 106, 2846, 2691, 567]. racial [3292]. radiation [1782]. radiological [1427]. radon [996]. Raising [883]. Rajan [557]. Ralph [2106]. Random [1930, 2585, 263, 1850, 296, 3111, 695, 3078, 654, 3299, 1112, 116, 2204, 2590, 2770, 636, 1816, 1535, 937, 1304, 1580, 2420, 1497, 1770, 2214, 3623, 3296, 1478, 290, 1979, 1680, 413, 3216, 586, 3303, 134, 1661, 2160, 2848, 3010, 1706, 3246, 1605, 1763, 2530, 317, 3241, 161, 338, 130, 773, 29, 965, 1961, 2285, 600, 3173, 3054, 179, 1144, 610, 213, 1230, 1022, 1184, 2064, 616, 292, 1305, 3324, 3013, 429, 340, 2084, 3126, 3607, 2698, 357, 2532, 287, 604, 3520, 3482, 1604, 272, 129, 133, 1498, 1940, 1078, 3250, 3338, 132, 3099, 131, 69, 2496, 3489, 2068, 1102, 2653]. random [420, 1257, 3242, 1083, 3543, 1189, 1451, 2930, 383]. random-censoring [2930]. random-effect [937, 3543]. Random-effects [263, 3078, 1816, 1535, 1304, 413, 586, 2848, 3010, 29, 965, 3054, 610, 616, 3013, 357, 3482]. random-threshold [2698]. randomisation [2655, 2000, 756, 1115, 3335]. Randomised [3434, 875, 1934, 2765, 2382, 888, 1805, 2478, 2783, 1523, 1074, 667, 1336, 3486, 2800, 2586, 3617, 1207, 2631, 743, 3421, 3649, 1856, 2639, 665, 606, 1459, 2650, 1635]. Randomization [3258, 3260, 3261, 1830, 1853, 2625, 3257, 3262, 3259, 2059, 3371, 883, 2148, 389, 2688, 444, 2249, 2851, 2265, 1686, 1822, 2002, 1098, 467, 1153, 948, 772, 1689, 400, 698, 1516, 2075, 2740, 2022, 2224, 2837, 1919, 177, 1622, 3419, 2918, 630, 414, 3145, 3460, 2913, 1563, 3490, 2682, 436, 1776, 1690]. randomization-based [2022]. randomizations [705]. Randomized [3258, 3260, 1623, 786, 391, 1561, 3658, 1496, 2648, 1697, 1858, 2465, 2364, 972, 1514, 114, 862, 3150, 155, 2259, 291, 1862, 2235, 25, 1977, 1325, 2838, 1488, 2649, 3136, 2862, 578, 1355, 2367, 1829, 3395, 3606, 3342, 3559, 2999, 44, 3265, 3097, 1324, 1331, 1061, 2241, 1441, 2546, 551, 1515, 3381, 1810, 1151, 2194, 690, 2449, 1249, 1656, 1811, 2224, 2593, 2837, 3308, 2256, 3516, 1699, 3261, 53, 1116, 531, 1581, 1904, 1984, 464, 3609, 534, 1687, 373, 2447, 2161, 1714, 1263, 3257, 535, 1440, 1444, 1939, 1701, 2832, 3533, 3461, 712, 3111, 2526, 580, 3469, 806, 3330, 2229, 1859]. randomized [2668, 2766, 3437, 1997, 3187, 339, 2366, 859, 3259, 1474, 1855, 865, 1655, 944, 1129, 2399, 1852, 2922, 672, 3143]. randomly [141]. randomness [842]. range [3155]. Rank [2370, 1895, 3255, 1335, 633, 2365, 3361, 3209, 381, 1549, 593, 2672, 3483, 3011, 173, 2510, 1892, 2716, 1860, 2375, 1042, 2653, 1499, 3305, 710]. Rank-based [1895, 3255, 1335, 2365, 381, 3011]. Rank-preserving [2370, 2672]. ranked [3222, 1667]. ranking [494, 1536, 1148, 890, 486]. ranks [1538]. Rapid [2017, 2302, 3000]. Rare [880, 731, 2364, 2955, 211, 2055, 3465, 2700, 1700, 3087, 1107, 2166, 2724, 3075, 3350]. rare-event [2055]. rare-variant-based [3350]. Rasch [401, 1965, 2140]. Rasch-based [401]. rate [426, 1854, 2470, 3050, 2937, 12, 2315, 2355, 3452, 235, 1869, 2570, 476, 2945, 154, 3656, 2746, 1507, 1999, 14, 3232, 2358, 3603, 3442, 1013, 2511, 2587, 3048, 1928, 1724, 2895, 3113, 82, 1715, 2295, 3075, 309, 2699, 2305, 3297, 3081, 2257, 1664, 2645, 2899]. rate-based [3603]. rater [2977, 2790, 333, 1345, 3090]. raters [3218, 2015, 617]. rates [2213, 2978, 3206, 23, 1556, 2509, 2423, 715, 1986, 1695, 168, 3506, 2679, 1022, 126, 61, 3359, 3593, 811, 1927, 2623, 1293, 3267, 1538, 1429, 1606, 1545]. rates/count [2978]. rather [3459, 3322]. rating [3090, 246]. ratings [646, 2790, 3481, 1257]. Ratio [3266, 785, 754, 2630, 3027, 1971, 1748, 1349, 488, 1232, 1827, 1211, 1098, 2261, 1575, 3010, 698, 2740, 33, 2381, 2511, 3619, 437, 2325, 2574, 1007, 3175, 907, 1410, 1190, 2864, 2077, 480, 229, 265, 79, 309, 590, 2557, 2833, 980, 352, 1338, 2875, 782, 1951, 258, 906, 3554, 332, 1578, 3407]. ratio-of-mediator-probability [3027]. ratio-weighted [1338]. ratios [1212, 751, 2684, 987, 262, 1923, 1568, 2612, 100, 1513, 1999, 387, 1316, 1714, 2243, 3145, 1989, 3627, 3605, 327, 2076]. Rato [2591]. Raton [2012, 447, 432, 118, 106, 411, 412, 446, 2725, 506, 2691, 3059, 567, 433, 380]. ray [344]. RCTs [3315]. RD [88, 3647]. RDD [3586]. Re [3419, 1635, 3370, 298, 39, 1296, 2113, 2826, 1397, 1270, 797, 864, 1671]. Re-estimating [1635]. re-estimation [3370, 1296, 2113, 2826, 1397, 1270, 797, 864, 1671]. re-evaluation [2113]. re-examined [298, 39]. Re-randomization [3419]. reaction [505, 3197]. reactions [3042]. read [3628, 1900]. read-outs [3628]. reader [2068]. reader-based [2068]. readings [402, 812]. readmission [3495]. readmissions [2407]. Real [55, 2276, 694, 1579, 3101, 3022, 3282, 3437]. Real-time [2276]. real-world [694, 1579, 3101, 3437]. realistic [3234]. realities [197]. really [2454]. reappearance [431]. reappraisal [1574]. reasoning [3547]. reassessment [469, 2584, 507, 989, 510, 1926, 1907, 511, 2582, 515, 1009, 2257, 1170, 950]. rebounds [2652]. recalibration [1021]. recapture [2140]. received [1384]. Receiver [2250, 1393, 2480, 212, 249, 1477, 600, 843, 484, 3294, 1372]. recently [2089]. receptor [2636]. RECIST [2912]. Reclassification [1542, 647, 2900, 1645, 2306, 1646, 322, 651, 1756, 2895, 3297, 2899]. recognition [1175]. recombination [1930]. recommendation [2202]. Recommendations [2044, 2648, 2882, 2765]. Recommended [1603, 1331]. Reconciling [3559]. record [1705, 921, 2742, 1797, 1666, 3174]. record-linked [921]. recorded [761]. records [2820, 2760, 3240]. recovering [2736]. recovery [392]. recruitment [1398, 2764, 3395, 3332, 769, 2824, 1635]. recurrence [3074, 1627, 1531, 1010, 2184]. recurrences [1392, 2483]. Recurrent [2348, 165, 2438, 2347, 3495, 2102, 71, 1674, 1008, 1988, 2437, 2454, 2998, 861, 1572, 3561, 2503, 3553, 3006, 3539, 3603, 2609, 3202, 3541, 349, 3011, 732, 1200, 305, 2275, 1397, 2448, 3075, 2690, 3359, 309, 525, 3593, 3645, 960, 1567, 2326, 2756, 3124, 1081, 561, 1664, 589, 1150]. recurring [2421, 159]. Recursive [1078, 553, 3045, 3049, 3635]. rediscovery [2358]. redistributing [1964]. redress [1296]. Reduce [1713, 448, 501, 2675, 2305, 2682]. reduced [3043, 1408, 80, 3435, 2510]. Reducing [1205, 478, 1989, 807, 1870, 788, 2736, 3586]. reduction [2201, 2516, 2016, 459, 824, 353, 1459, 1664]. reductions [216, 442]. reestimation [117]. reference [1710, 2138, 1147, 1375, 2162, 1504, 250, 1394, 2190, 595, 2212, 2881, 729, 458, 943, 739]. referenced [204, 3641]. referral [2997, 2770]. refine [2003]. refined [1443]. refinement [2292, 2848]. refitting [1258]. reflect [2901]. Reflections [3260]. Regaining [1675]. regard [1577]. regarding [1605, 1696, 1604, 2893]. regime [1762]. regimen [2389]. regimens [3191]. regimes [2367, 1871, 1572, 2037, 3036, 2226, 255, 3401, 3600]. region [3095, 3009, 3096, 3066, 3641, 1479]. region-referenced [3641]. region-specific [3066]. regional [2147, 2665, 1133, 1562, 272, 3317, 518]. regions [431, 2307, 1706, 2910, 3295, 2309]. registration [716]. registries [3638]. registry [2636, 2183, 678, 1785]. regress [2287]. Regression [431, 1462, 2689, 1877, 2187, 526, 3426, 286, 847, 3572, 482, 1330, 1237, 1346, 1080, 2318, 1435, 1741, 2680, 1363, 2991, 456, 3047, 996, 1651, 1214, 1053, 1439, 2795, 2978, 3538, 2701, 2829, 1854, 3028, 84, 1140, 1434, 35, 294, 2073, 1556, 1886, 2423, 2976, 1349, 1731, 155, 46, 1743, 3142, 1820, 1317, 1103, 1528, 2288, 2370, 2804, 3043, 12, 128, 3159, 3205, 2355, 3616, 1171, 1873, 1894, 1668, 1957, 953, 1238, 330, 1905, 2183, 666, 2945, 1409, 149, 3224, 80, 459, 2529, 520, 3026, 136, 2950, 2553, 3596, 3056, 3411, 119, 3303, 1443, 2939, 3093, 2914, 3510, 2481]. regression [3182, 3194, 490, 766, 1693, 1922, 3526, 3293, 1943, 1065, 1228, 3566, 3060, 1740, 88, 1464, 3232, 92, 1390, 2643, 3006, 3181, 397, 3031, 59, 1833, 1994, 2356, 3134, 3202, 3017, 3583, 1872, 1926, 1821, 965, 2731, 2877, 3247, 1918, 3302, 1769, 2821, 3192, 3464, 231, 830, 3011, 1210, 127, 3558, 3509, 1984, 1760, 1972, 750, 1223, 2940, 2452, 2806, 1592, 2150, 1973, 113, 1240, 17, 2725, 1885, 2088, 1622, 1676, 2186, 2084, 2177, 1215, 3318, 2233, 2724, 3126, 256, 1773, 2537, 736, 1387, 263, 248, 2908, 2293, 237, 1989, 1785, 2975, 3269, 1715, 1118, 230, 1583, 1364, 894, 2733, 1436, 933, 611]. regression [79, 1868, 1407, 353, 525, 257, 3608, 757, 1902, 3518, 2305, 2883, 1365, 1302, 1691, 3226, 951, 891, 1772, 2049, 2496, 2778, 3514, 3450, 2929, 3427, 2798, 343, 926, 2980, 94, 3404, 618, 2803, 2682, 3081, 3149, 2878, 3588, 498, 829, 3319, 3117, 2695, 1641, 966, 2185, 246, 334, 2117]. Regression-assisted [482]. regression-based [2553, 3226]. regression-derived [750]. regressions [3457, 2016, 1371, 1288]. regressor [2950]. regularization [3142, 1164]. regularized [2278, 1482, 1772, 3320, 2117]. regulators [475]. Regulatory [3139, 475, 2884, 742, 465, 2334, 2330, 2255, 2332, 2331, 2885, 2333, 3116]. Reidar [445]. rejection [3095, 3009, 3096]. rejective [463]. Rejoinder [1048, 365, 546, 3263, 570]. related [3538, 2031, 2389, 372, 1375, 952, 2016, 507, 2463, 738, 1826, 1573, 3499, 480, 2871, 3563, 1222, 1856, 2639, 2280, 2309, 3021]. relatedness [1490]. Relating [3458, 210]. relation [996, 2019, 21, 943, 739]. Relational [2722]. relations [3590, 95]. Relationship [3080, 1062, 1158, 81, 3616, 954, 263, 392, 2197, 3115]. relationships [2045, 135, 3587, 1494, 2897, 2895, 3297, 3287]. Relative [3245, 663, 3389, 1240, 1588, 2986, 2354, 2534, 241, 3019, 1423, 1211, 2759, 139, 1030, 314, 1353, 91, 1656, 450, 2536, 642, 229, 1696, 13, 66, 3540]. relativism [855]. Relaxed [3221]. Relaxing [1734]. released [2089]. relevance [3415, 1134, 1646, 1738, 2990]. reliabilities [1340]. reliability [28, 3321, 3122, 64, 2552, 439, 1348]. reliable [3639, 3654, 3653]. Reliably [3229]. Relief [1550]. rely [1645]. remaining [2893]. remark [777]. Remarks [1059, 495, 496]. remedies [1721]. reminiscences [859]. REML [2585, 2713, 1189]. Remontet [2033]. removable [1093]. renewal [1010]. Repairing [2259]. repeatability [2619]. Repeated [2374, 266, 1350, 920, 112, 636, 699, 547, 1208, 2760, 415, 1460, 538, 1448, 2512, 379, 1912, 2226, 1707, 1124, 3416, 261, 834, 1108, 623, 1931, 192, 1235, 2118, 230, 276, 277, 2859, 2904, 2713, 2656, 1653, 3425, 2632, 2805, 814, 1129, 2963, 1903]. repeated-measurement [3416]. repeated-measures [2226]. repeatedly [2916, 2439, 3129, 3594]. repeats [1506]. replacement [1709]. replicate [1876, 976]. replicated [402, 812, 2788]. replicates [640, 3601]. replication [494, 1599]. replications [2514]. Reply [195, 1804, 2033, 140, 2123, 2131, 151, 1456, 185, 1266, 209, 558, 288, 133, 337, 973, 1975, 409, 2220, 176, 639, 629, 3654, 1401, 3646, 407, 650, 73, 648, 2885, 1981, 585, 1606]. Report [1944, 1945, 1948, 2330]. Reported [401, 912, 3325, 1916, 3604, 2488, 3326, 1965, 795, 3636, 57]. Reporting [2823, 875, 1496, 2882, 3562, 227, 138, 3156, 807, 1684, 370, 410, 704, 3611]. reports [2791, 505, 2421, 1599, 689, 2335]. representation [813, 2064]. representations [816]. representative [764]. reproducibility [1032]. reproduction [1492, 3236, 419]. require [1646]. required [2952, 1476]. requirement [44]. requirements [3308]. resampling [3140, 2098, 2717, 349, 386]. rescue [2677]. Research [851, 390, 856, 202, 3273, 852, 1697, 2364, 2976, 1025, 3274, 105, 2009, 1767, 882, 1017, 1283, 854, 302, 197, 571, 855, 1408, 2467, 884, 3547, 1127, 885, 198, 881, 2917, 837, 3051, 2485, 200, 1383, 2504, 2505, 1579, 1842, 1282, 2013, 445, 201, 2333, 1648, 196, 2598, 3033, 2622, 1281, 1522]. researchers [1433, 999]. resection [2388]. resetting [2556]. residential [522, 3337]. residential-workplace [3337]. Residual [2518, 1929, 2469, 2689, 2277, 2621, 3596, 2848, 1515, 3406, 1256, 792, 3387, 332]. residual-based [2469]. residuals [1651, 2589, 2961, 3070, 149]. resistance [1112, 1711, 1930, 2530, 2698, 11, 618]. resistant [563]. resolution [2292, 2600]. resonance [676, 1754, 2067, 2667, 1161, 1482]. resources [390, 1812]. respect [3564]. respiratory [1601]. Respondent [3272, 501]. Respondent-driven [3272]. respondents [3087, 158, 1718]. responder [1611]. responders [2694]. Response [3505, 3459, 2142, 1866, 1283, 3326, 3096, 602, 1559, 1441, 969, 1058, 1279, 3071, 496, 177, 3504, 3650, 2625, 1793, 1086, 3103, 202, 205, 3021, 754, 2045, 1370, 2116, 1806, 1173, 1434, 2988, 3070, 1158, 105, 1130, 2841, 637, 1355, 3473, 1471, 2683, 2002, 740, 2165, 366, 1153, 1605, 1670, 992, 2463, 3186, 1065, 2119, 1777, 2007, 313, 1101, 1448, 2912, 553, 2877, 2121, 3192, 615, 2549, 1703, 2903, 1039, 2500, 263, 1953, 826, 3537, 223, 894, 251, 580, 554, 1902, 11, 258, 2985, 880, 2646, 1474, 1757, 3002, 436, 1690]. Response-adaptive [1279, 177, 3103, 1370, 1130, 2002, 1153, 3186, 313, 1448, 580, 436, 1690]. Responses [1948, 886, 555, 260, 1370, 335, 186, 1688, 744, 2320, 1767, 660, 1276, 3241, 3340, 2855, 1448, 1125, 1660, 3128, 2096, 3464, 1913, 2024, 1638, 1222, 1182, 550, 537]. restraint [2609]. Restricted [683, 3523, 3652, 2689, 105, 804, 3097, 449, 2088, 776, 825, 535, 3145, 980, 3614, 1997, 2369, 3146]. restricting [2951]. restriction [3457]. Results [3459, 3322, 614, 3114, 1085, 649, 1681, 1186, 3547, 213, 613, 615, 2175, 2874, 1938, 994, 2450, 2229, 2798, 2805, 1196]. resuscitation [2018]. retaining [889]. retention [2957]. Rethinking [2974]. retinitis [3113]. Retirement [2439]. Retracted [2750]. Retraction [3185]. retrieved [1310]. retrospective [2458, 2006, 621, 3551, 808, 3131, 3549, 3548, 351]. retrospectively [912]. reusing [741]. reveals [735]. reverse [1623]. Review [398, 2012, 447, 432, 520, 1457, 3082, 411, 2591, 412, 2846, 446, 2725, 2186, 289, 3369, 506, 2691, 3059, 1786, 3022, 1522, 749, 3496, 567, 2489, 3076, 380, 2648, 1698, 621, 2944, 210, 1648, 3469, 3089, 2287]. Reviews [460, 118, 106, 1991, 445, 433, 659, 807]. revision [917]. revisited [579, 2082]. revisiting [2240]. revolution [832]. Reweighting [618]. rhythms [1090]. richer [2005]. ridge [3026, 1872, 1416, 2117]. right [20, 2689, 3477, 1300, 2317, 3197, 2371, 1838, 1056, 1223, 1892, 354, 3103, 3119, 2217, 498, 1578]. right-censored [20, 2689, 3477, 2317, 3197, 2371, 1838, 1223, 1892, 3103, 498, 1578]. Risk [3413, 50, 2902, 1062, 996, 1628, 2518, 2517, 216, 442, 1053, 2427, 2882, 3034, 1682, 1851, 2898, 1198, 1971, 2073, 1818, 3373, 323, 1521, 1964, 2900, 241, 3019, 1988, 849, 1121, 1211, 1085, 427, 845, 233, 953, 1630, 2403, 2760, 574, 1629, 1476, 3118, 3197, 855, 2644, 3275, 1002, 1553, 3422, 2759, 100, 1585, 146, 1491, 2133, 3462, 2481, 2901, 1958, 21, 1613, 1909, 1065, 88, 1464, 678, 2816, 901, 657, 3249, 3468, 2274, 916, 1069, 349, 1392, 2800, 2536, 642, 2255, 1240, 1256, 1139, 2186, 760, 1190, 274, 2352, 664, 1588, 2003, 1268, 994, 1086, 2597, 1989, 229, 485, 2867]. risk [2184, 386, 2904, 43, 13, 2749, 3184, 455, 2881, 304, 428, 1185, 2280, 1155, 3190, 242, 3105, 2001, 3024, 1368, 394, 618, 1244, 2198, 66, 625, 2029, 2745, 2651, 693, 1733, 2435, 2899, 681]. Risk-adjusted [3413, 50, 574, 2749, 3105, 625, 2029, 2745]. risk-benefit [2133]. risk-scoring [2427, 3034]. risk-set [349]. risks [3047, 724, 587, 2427, 2648, 3034, 3316, 3513, 308, 2829, 321, 243, 1393, 1449, 2937, 2101, 1229, 3590, 953, 1630, 1417, 1728, 1766, 1680, 3171, 473, 314, 725, 1810, 1228, 387, 913, 3638, 1656, 58, 773, 3583, 1072, 1230, 123, 1157, 2879, 274, 2118, 3084, 2711, 3283, 1493, 169, 1696, 3125, 3148, 3032, 3387, 498, 2616, 3525]. risky [1302]. Riviere [1792, 1791]. RNA [1900]. RNA-sequencing [1900]. road [3443]. Robert [584, 445]. Robin [3648]. Robust [2976, 962, 3169, 3623, 1141, 1872, 1187, 42, 1287, 534, 2228, 2531, 305, 2414, 60, 3107, 3580, 3404, 2440, 1189, 1762, 346, 2370, 2571, 2915, 2838, 3447, 2841, 224, 22, 1241, 2258, 3511, 1183, 581, 3166, 750, 1356, 1084, 1442, 3311, 3228, 15, 3138, 2390, 2690, 3286, 352, 762, 441, 2351, 1881, 993, 1332, 2074, 2967, 710, 2565, 3588, 984, 2373, 1054, 1852]. Robustness [489, 1296, 2241, 1270, 3371, 3213, 653, 290, 762, 1775]. ROC [2451, 2840, 3206, 2073, 1147, 3361, 653, 454, 2406, 300, 1764, 156, 2261, 440, 2654, 1573, 1255, 2181, 299, 3120, 1118, 1169, 2982, 1880, 1479, 2279, 2714, 691, 3488]. ROC-type [1147]. Rockette [3080]. Roe [3080]. Röhmel [602]. Rongling [412]. Rossi [1889]. rotavirus [1520]. routinely [2956]. routines [626]. rt [58]. rt-PA [58]. rubella [752]. rule [1795, 2545, 103, 2668, 3105, 2203]. rules [1099, 1101, 2642, 3471, 1100, 3440, 2417, 2405, 3404, 1437]. run [3656, 3442]. run-off [3656, 3442]. Russell [153]. Ryan [408].

S [628, 882, 649, 557, 459, 3, 2539, 1265, 1889, 1792, 1791]. sacrifice [3509]. Safeguarding [83]. safety [477, 3217, 673, 1575, 2360, 3036, 1643, 2274, 265, 2211, 2942, 1868, 2008, 499, 394, 1244]. Saharan [390, 2040]. Salmonella [1759]. salvage [254, 1422]. same [217]. Sample [1309, 1608, 3370, 155, 527, 653, 2098, 275, 3333, 1165, 2100, 1064, 44, 2478, 703, 1685, 1061, 1504, 1670, 161, 338, 269, 1327, 3060, 1477, 3198, 3631, 2256, 641, 1581, 1860, 220, 3593, 956, 1975, 2786, 3490, 3032, 2650, 1852, 1429, 957, 3376, 3095, 2401, 23, 3415, 2688, 2469, 315, 3604, 2640, 2343, 453, 603, 1296, 1923, 3331, 1231, 2315, 264, 18, 2717, 614, 2317, 3225, 117, 3009, 3096, 2403, 649, 1476, 476, 3612, 3290, 740, 1965, 381, 940, 572, 2061, 538, 2044, 3576, 869, 2914, 3486, 2205, 1549, 593, 1026, 1099, 1101, 1811, 3308, 2660, 2808, 2512, 2634, 2821, 1896, 1819]. sample [3334, 613, 615, 1100, 2113, 143, 2826, 17, 3046, 1676, 2243, 3355, 305, 1773, 3647, 3352, 3353, 2136, 1397, 1270, 2807, 721, 928, 2690, 797, 2713, 3152, 806, 983, 2008, 2392, 1638, 782, 1794, 441, 2620, 339, 2402, 801, 1348, 1936, 880, 864, 1120, 2257, 2373, 2605, 1671, 1635, 768, 767, 1974, 1606, 1545, 2807]. sampled [958, 2391]. samples [442, 372, 903, 942, 284, 804, 3093, 1241, 1758, 2036, 95, 2359, 2168, 3537, 276, 277, 2232, 2521, 2632, 1425, 2520]. sampling [1550, 2876, 3220, 3604, 3201, 2051, 3272, 1188, 1908, 1473, 3245, 3197, 377, 1986, 3131, 1597, 3389, 723, 783, 3638, 2719, 2654, 1926, 1007, 3195, 1998, 2030, 3100, 3564, 3098, 1996, 3550, 826, 2260, 3586, 1437, 3367, 3083]. sandwich [1811, 276]. Sankoh [2106]. sarcoma [939]. SAS [106, 2846, 3513, 433, 749]. SATS [504]. Savage [858]. SCAD [2817, 2972]. scalable [2548]. scalar [3224, 1978, 3518]. scalar-on-function [3518]. scalar-on-image [3224]. scale [3185, 3534, 1317, 2183, 2854, 910, 1835, 896, 3082, 1308, 2750, 3122, 1297, 57, 1345, 3090]. scaled [2923, 2881]. scales [909, 1388, 1235, 246]. scaling [3236, 2742, 2703]. Scan [1996, 431, 1539, 3354, 1864, 1866, 240, 190, 1825, 1865, 657, 367, 1168, 1588, 1245, 1374, 980, 2251, 1244, 701]. Scan-stratified [1996]. scans [2646]. scenarios [3444, 3278]. schedule [2566]. schedules [1015, 1039, 1445]. scheme [2751, 350, 2878]. schemes [3389, 3638, 222]. Schisterman [1265]. schizophrenia [2277, 500]. Schuemie [3549]. science [1763]. Sciences [289]. Scientific [1029, 397, 398]. scientists [432]. sclerosis [676, 117, 466, 1943, 2694, 1996, 410]. Score [2347, 3466, 3382, 2263, 216, 442, 1212, 1484, 1709, 1496, 2053, 3067, 3316, 3220, 1434, 2249, 1555, 328, 1375, 697, 2700, 661, 1409, 2612, 2468, 2319, 1077, 593, 3060, 40, 813, 901, 1656, 1250, 2381, 2258, 3154, 3065, 2574, 2726, 2050, 3189, 1251, 3172, 1842, 2782, 3365, 2495, 3537, 79, 304, 1691, 3284, 56, 765, 1459, 3146, 3174, 1106, 2386]. score-based [1434, 901, 2050, 304]. score-free [2319]. score-independent [3146]. score-type [1555]. scores [2964, 3613, 355, 1441, 1652, 1249, 1830, 385, 3401, 1978, 2153, 1263, 1440, 1288, 1345, 3427, 1601, 246]. scoring [2427, 3034]. Scotland [913, 1519]. screener [484]. Screening [1068, 491, 902, 262, 124, 1247, 3452, 3118, 2371, 1015, 2090, 1994, 1872, 2026, 1431, 1079, 2686, 978, 282, 1949, 3467, 590, 3012, 1302, 729, 1253, 3116]. SD [3649]. SDIE [538]. seamless [1882, 466, 1795, 799, 1976, 2863, 2417, 98, 3211]. search [553, 2835, 3605]. seasonal [2271, 598, 1751, 3399]. Second [3180, 575, 155, 696, 2725, 163]. Second-order [3180, 575, 155, 696]. Secondary [3098, 2723, 1119, 119, 1702, 991, 2734, 1537, 796, 797, 2440]. section [2263]. sectional [1188, 3245, 379, 1124, 2036, 3508, 3499, 2686, 2352, 3068, 61, 1432, 1368, 681]. sections [3364]. Seeing [2009]. SEER [3171, 3387]. SEER-Medicare [3171, 3387]. segment [259]. Segmental [1024]. segmentation [1657]. Segmented [2587]. Segregation [1525]. seizure [1189]. select [1607, 2905]. selected [1430, 141, 895, 1708, 2653]. Selecting [1802, 2710, 1092, 2042, 3463, 1972, 86, 736, 299]. Selection [10, 3100, 3375, 2718, 3047, 810, 1824, 2204, 2266, 3415, 3445, 2787, 529, 1731, 1274, 2721, 3150, 1483, 597, 1692, 1132, 1882, 1272, 2320, 2839, 3140, 3227, 434, 1315, 494, 1295, 466, 988, 1191, 1195, 1871, 1905, 2720, 3296, 2732, 1728, 1478, 673, 2529, 467, 3171, 1915, 3577, 2133, 2245, 1745, 3038, 551, 1976, 1763, 2530, 1909, 1796, 1228, 503, 669, 1249, 1656, 2464, 2969, 3634, 3583, 1341, 3572, 844, 867, 1431, 1204, 1003, 1788, 646, 1329, 285, 573, 2289, 1148, 1142, 1724, 3324, 2177, 1367, 1362, 3084, 1888, 2417, 2754, 408, 594, 1152, 1782, 1785, 2425, 2189, 1163, 3630, 524, 2014]. selection [3132, 222, 3467, 983, 2008, 1966, 1365, 2482, 1742, 743, 484, 3427, 1564, 2366, 166, 409, 2653, 1193, 944, 395, 691]. selections [1903]. selective [2764, 2729, 3186]. Self [1450, 2791, 3604, 3150, 2779, 2421, 3410, 3553, 388, 1411, 57, 2983, 394, 1244]. Self-controlled [1450, 2779, 388, 1411, 2983, 394, 1244]. self-excited [3410]. self-reported [3604, 57]. self-reports [2791, 2421]. self-selection [3150]. self-triggering [3553]. Semi [267, 2481, 2084, 321, 2771, 2498, 1056, 280, 517, 2579, 2616]. semi-attributable [2579]. semi-competing [321, 2616]. semi-Markov [2498, 1056, 517]. semi-nonparametric [2771]. Semi-parametric [267, 2084, 280]. Semi-varying [2481]. semicompeting [1766]. semicontinuous [2504, 2505, 1750, 3252]. Semiparametric [2829, 1232, 2804, 2937, 3414, 3275, 2950, 1523, 1491, 490, 3006, 1225, 2852, 3202, 1684, 3368, 2635, 2705, 2698, 1527, 230, 3359, 1846, 2756, 589, 1274, 2721, 744, 1025, 1993, 1291, 1292, 1041, 3159, 1897, 1908, 1473, 146, 2324, 3056, 818, 3378, 766, 3016, 1226, 3603, 1994, 2917, 3478, 99, 867, 3362, 1344, 3166, 306, 1180, 1148, 3223, 451, 2139, 3467, 2666, 1365, 2074, 3580, 2865, 1081, 944, 3064]. Senegal [1344]. Senn [3569]. sensitive [36, 3157, 1069, 1687, 1474]. sensitivities [2068]. Sensitivity [1886, 2960, 3420, 1271, 3058, 876, 1699, 231, 2332, 714, 354, 835, 2636, 3657, 2451, 447, 434, 635, 2380, 227, 2796, 3135, 1074, 1745, 2334, 1335, 179, 3180, 671, 2769, 769, 2030, 2686, 3040, 1879, 3425, 2374]. Separable [3079, 1802]. separating [2618]. separation [80, 3435, 1047]. separation-proof [80, 3435]. sequence [2071, 54]. sequences [2277]. sequencing [2423, 2710, 1900]. Sequential [3513, 413, 2577, 790, 3102, 265, 3039, 624, 1205, 786, 160, 1608, 2925, 244, 1521, 1612, 3423, 1208, 435, 1299, 27, 270, 2002, 2595, 1097, 572, 1153, 772, 1515, 705, 2196, 33, 1643, 1551, 2593, 791, 1072, 1636, 1204, 3048, 273, 78, 2405, 1953, 1358, 1939, 1761, 1701, 2832, 2211, 3533, 2528, 1868, 796, 797, 632, 2545, 3187, 864, 3582, 3645, 1089, 2567, 2203, 1671, 3407, 586]. sequentially [463, 114, 2161]. serial [3497, 47, 1893]. serially [3565]. Series [2012, 2846, 460, 1342, 3168, 2915, 3385, 2802, 1450, 2779, 2067, 990, 2747, 388, 1411, 1063, 2350, 1885, 429, 1751, 1273, 3123, 2983, 394, 1244, 47]. serious [2800]. serological [752, 811]. serve [3289]. service [168]. services [2504, 2505]. session [2088, 2120]. set [18, 1706, 2984, 349, 642, 976, 734]. sets [2925, 12, 80, 2468, 1543, 1004, 221, 504]. setting [2873, 1923, 1906, 1771, 2424, 1812, 727, 726, 1868, 68]. settings [1637, 2087, 435, 3225, 3493, 3234, 3579, 3598, 714, 386, 2435]. Seven [1115, 3010]. Seventh [2004]. Several [323, 2655, 1829, 1509, 1285, 1735, 1372, 2788]. severe [532, 782]. Severity [1098, 2916, 1677, 2667, 3381]. severity-dependent [3381]. sex [1220]. sexual [2458, 2951, 3497, 917]. sexually [1994, 917]. SF-36 [2552]. shape [1470]. Shared [3299, 3339, 3310, 3570, 3041, 127, 1354, 95, 1179]. shared-parameter [3339, 3310]. Sharp [2851, 1322]. Shaw [2591]. shelf [3114]. Shift [1586, 2393]. Shift-invariant [1586]. Shigeyuki [1558]. Shkedy [2846]. shock [1836]. Short [622, 2758, 1438, 1689, 98, 2014]. Short-term [622, 2758, 1689, 98, 2014]. shot [1838]. Should [3382, 8, 882, 3606, 884, 1091, 2901, 881, 150]. shoulders [860]. Shrinkage [1132, 1305, 2484, 1771, 3558, 1133, 2975, 3343]. shrunken [1334, 1558]. Shuster [1497, 287]. SIAM [398]. sibling [610]. sided [214]. sides [922]. Signal [1987]. signaling [3105]. signals [3130, 3254, 788]. signature [1814, 1765]. signatures [1140, 85]. significance [3648, 1134, 656, 3555, 3650, 1285, 2374, 577]. significant [3291]. sim.3607 [287]. sim.3722 [72]. sim.4088 [586]. sim.4105 [459]. sim.7315 [3185]. Simes [2426]. SIMEX [1553, 3025]. similar [1496, 2147, 2595]. similarity [2988, 2601]. Simmonds [586]. Simon [111]. SIMPLE [369, 623, 2376, 1059, 1197, 1970, 2571, 938, 1098, 588, 1241, 2863, 1284, 437, 2325, 907, 1377, 17, 24, 3537, 806, 1857, 1145, 3105, 2799]. simpler [2005]. simplicity [3568]. Simplified [3446, 2179]. Simpson [2896]. simulate [955, 1087, 1071]. Simulating [1307, 979, 2623]. Simulation [911, 2732, 538, 3482, 1868, 497, 1480, 1053, 1709, 2364, 1173, 1621, 3168, 841, 842, 2055, 785, 1759, 3080, 3529, 1652, 1249, 2485, 3609, 3405, 3013, 1751, 2874, 1285, 1735, 2078, 1114, 3537, 49, 1389, 3022, 3282, 2104, 2671, 801, 3536]. Simulation-based [2732, 1759]. simulations [1834, 997]. Simultaneous [214, 1135, 314, 2936, 2056, 2910, 3046, 3627, 3174, 1642, 20, 2596, 3106, 2530, 78, 3440, 2545]. simultaneously [250]. sine [1211]. Single [3298, 1140, 477, 1580, 798, 1295, 2547, 3157, 3433, 111, 2121, 110, 273, 3365, 485, 2014, 983, 1432, 2610, 1849]. single-arm [1580, 798, 3433, 111, 273, 983]. single-disease [485]. single-index [1295, 3157, 3365, 2610, 1849]. single-nucleotide [1140]. Single-number [3298]. singly [2702]. sinh [3492]. sinh-arcsinh [3492]. Sinha [459]. SIR [3236]. sisters [1258]. sisterships [1258]. sisVIVE [3371]. Sixth [461]. Size [627, 3285, 1309, 1608, 3370, 23, 155, 291, 489, 527, 653, 2640, 2098, 275, 1296, 3331, 1231, 2315, 264, 3333, 18, 3255, 2717, 614, 1165, 117, 768, 2403, 2100, 649, 1064, 1476, 476, 3209, 3290, 740, 1965, 44, 2478, 703, 1685, 381, 3389, 572, 1061, 1194, 1504, 2061, 538, 3576, 2914, 1670, 3486, 286, 2205, 2958, 161, 338, 885, 269, 2194, 593, 1327, 3060, 1099, 1101, 1477, 3198, 3308, 3631, 2256, 2660, 2808, 48, 3483, 641, 3195, 1581, 1896, 1819, 613, 615, 1100, 2113, 2826, 2751, 17, 3500, 3355, 1290, 3647, 3352, 3353, 1860, 1397, 1270, 1786, 393, 220, 251, 928]. size [2690, 767, 797, 2713, 3152, 3593, 806, 983, 2008, 1638, 1794, 956, 1975, 2620, 2786, 809, 880, 864, 1974, 3490, 3032, 2650, 2257, 2373, 2605, 1852, 1671, 1606, 1545, 1429, 957, 3376, 672, 1635, 2807]. size/power [1685]. sizes [116, 155, 2259, 2209, 1923, 3245, 1460, 869, 721, 1936]. Sizing [2659, 3396]. Skaltsa [727]. skeletons [989, 2582]. Skellam [2556]. Skew [238, 1022, 3185, 260, 355, 2750, 2293, 1846, 2666]. skew- [3185, 2750]. Skew-elliptical [1022]. Skew-normal [238, 260, 355, 1846, 2666]. skew-normal/independent [260]. skew-symmetric [2293]. skewed [1993, 1494, 2096, 3173, 828]. skewed-longitudinal [1993]. skewness [1291, 2519, 582, 1024, 2981, 1602]. skills [2977]. sleep [1248, 3541, 708]. sleepiness [227]. slippery [859]. Slope [259, 369, 2897, 3552]. slope-based [3552]. slope-intercept [369]. slopes [1184]. Small [1811, 2821, 1676, 1682, 1851, 3415, 1768, 2717, 3225, 804, 2183, 2854, 1681, 3612, 990, 1986, 80, 2061, 2360, 3435, 750, 1377, 3046, 1773, 1114, 3537, 276, 277, 1927, 441, 2232, 2521, 2830, 3187, 339, 2402, 2632, 47, 2520]. small-sample [3225, 3612, 3046, 441, 339, 2402]. small-scale [2854]. small-study [750]. smaller [672]. smart [2677, 2226]. smoking [2853, 1355, 1681, 1600, 65, 1376, 1616]. Smooth [2771, 327, 383, 2840, 1450]. smoother [1592]. smoothers [1439]. Smoothing [379, 1198, 1895, 1447, 818, 3011, 2830, 2391]. smoothly [2678]. smoothness [2437]. snapshot [210]. Snell [3647]. SNP [2771, 819, 235]. SNP-SNP [819]. SNPs [3371]. snSMARTs [3187]. Social [1044, 1048, 416, 1043, 1046, 1047, 1045, 3112, 1045]. Society [2004, 461, 3371]. sojourn [1247, 1551]. solution [1059, 1967, 1573, 1145]. solutions [755, 3617]. Some [1398, 1197, 180, 3161, 365, 359, 108, 1049, 3061, 2893, 533, 2218, 600, 1569, 3421, 3649, 859, 1721, 361, 363, 362, 360, 364]. Somnath [3483]. Songfeng [1974]. sore [2454]. sound [538]. source [2019, 1704, 2669, 930, 577]. sources [97, 1665, 940, 2665, 3638, 2278]. Southeast [1069]. Space [204, 1598, 2802, 233, 1238, 2290, 2951, 3015, 2404, 3410, 2023, 817, 1588, 1290, 2251, 577]. Space-time [204, 1598, 233, 3015, 817, 1588, 2251]. spaced [685]. Sparse [1390, 3162, 2993, 815, 1140, 2758, 1292, 735, 80, 3435, 1663, 296, 162, 3580]. sparsity [2601, 1577]. Spatial [1198, 961, 3123, 2830, 701, 104, 1539, 3354, 596, 2534, 241, 3019, 233, 3392, 519, 2019, 2759, 240, 378, 190, 1825, 21, 2643, 657, 3480, 411, 2372, 425, 1020, 1022, 2357, 2088, 1543, 2151, 1588, 1245, 1374, 980, 2678, 575, 3190, 2058]. Spatially [1482, 1274, 3597, 1874, 2792, 373, 1383, 3399, 2453]. spatially-temporally [1274]. spatio [2247, 803, 2514, 3575]. spatio-temporal [2247, 803, 2514, 3575]. Spatiotemporal [2292, 378, 3081, 3019, 1802, 3235, 1973, 577, 3219]. Special [850, 2510]. species [1303]. specific [1112, 1627, 1844, 1542, 1160, 2919, 1770, 1121, 1552, 3397, 2503, 3526, 2663, 2424, 42, 3066, 2030, 2940, 2879, 591, 1883, 61, 2661, 342, 1432, 951, 2068, 3032, 984, 537]. Specification [53, 3051, 2465, 2707, 1341, 3475, 3099, 1666]. specifications [2213]. specificities [2068]. specificity [2451, 635, 2686, 3425, 2374]. specified [2149, 1099, 1101, 1984, 1100, 933]. specify [2582]. specifying [3231]. specimens [902, 675]. Spectral [3130, 1507, 3254]. spectrometry [395]. spectrum [2006, 2067]. Spending [878, 38, 1722, 2942, 3408]. Spiegelhalter [72, 1457]. spike [2939, 2121, 1421, 2556, 125]. spikes [1273]. Spline [2779, 1808, 2247, 105, 913, 893, 493]. Spline-based [2779, 1808]. splines [1173, 1447, 327, 2712]. Split [3418, 2653]. spondylitis [1462]. spontaneous [505]. spot [1869, 3529]. spots [577]. spread [3392, 2151, 2514, 3003]. Springer [520]. squamous [263]. square [2433, 2538]. squared [2261]. squares [34, 963, 1055, 2241, 1507, 1390, 1644, 1940, 3162]. Sr [647, 2106]. stability [3114]. stabilize [3365]. stabilized [1889, 1981, 1845]. stable [1144]. stacked [2369]. staffing [81]. stage [2738, 987, 972, 949, 1430, 2210, 28, 2787, 2685, 1349, 2624, 488, 3150, 1132, 1862, 771, 1130, 635, 798, 1500, 914, 2219, 1518, 3446, 2618, 861, 2037, 2273, 989, 3389, 866, 1092, 2133, 2604, 303, 711, 3463, 1962, 3151, 674, 1082, 991, 2633, 1549, 3231, 2693, 3634, 3572, 3633, 111, 2026, 3048, 2030, 255, 2923, 3037, 573, 3545, 1765, 591, 148, 2195, 295, 251, 521, 222, 3022, 3282, 796, 797, 734, 743, 3421, 3649, 1951, 1794, 3076, 2409, 985, 1563, 1673, 1831, 2682, 1052, 3387, 2565, 1963, 781]. stages [3410, 26]. Stagewise [2185]. staggered [3201]. staggering [3265]. Standard [1158, 3303, 1710, 2073, 264, 3062, 212, 2162, 161, 338, 250, 1582, 2030, 2826, 2627, 226, 382, 192, 3355, 443, 1358, 3627, 928, 2526, 729, 695, 3425, 1584, 94, 1257, 1196, 1480, 481, 1306, 2203]. standardised [2639]. standardization [1959, 1552, 3489]. standardized [3070, 1065, 590, 2875, 3267]. Standardizing [1004]. standards [1394, 2212]. standing [860]. Staquet [1380]. STARD [2823]. STARD-BLCM [2823]. Start [2002, 2061, 2450]. Start-up [2002]. State [3203, 1290, 1966, 1637, 2087, 189, 122, 1449, 1531, 2419, 2802, 803, 1388, 119, 730, 3410, 93, 2957, 1551, 1969, 2023, 1223, 1049, 331, 2652, 2139, 2472, 121, 2599]. State-level [3203]. State-space [1290, 2802, 3410]. Statement [875]. states [2218, 1970, 164, 2608, 65, 1892, 1927, 1538, 904, 240, 164]. stationary [2067]. Statist [3185]. statistic [1539, 840, 3354, 1864, 190, 1276, 1077, 1865, 657, 581, 1245, 1374, 980, 2251, 2322, 1244, 1587, 2538, 1576, 1866]. Statistical [964, 1036, 3447, 107, 2272, 1060, 1815, 1030, 2463, 1867, 3332, 1031, 2634, 143, 496, 3499, 3394, 82, 601, 1784, 2661, 1718, 2736, 2155, 2646, 3267, 3540, 47, 1150, 827, 3273, 2657, 442, 2182, 2004, 3620, 1899, 2142, 2812, 3274, 1158, 1778, 819, 2048, 2236, 1754, 461, 1134, 1149, 1405, 690, 1271, 42, 3609, 598, 3405, 2619, 292, 889, 2874, 2159, 2141, 2136, 1093, 1782, 689, 310, 2904, 749, 877, 2798, 2902, 906, 577, 1076, 2857, 550, 3133, 495, 412, 3059]. statistician [742]. statisticians [719, 2591, 1043, 432]. Statistics [460, 3656, 1594, 1381, 431, 152, 328, 20, 2801, 153, 1560, 1536, 325, 2951, 461, 688, 240, 2306, 22, 1825, 1921, 411, 1467, 367, 915, 760, 1598, 1588, 289, 109, 3163, 1169, 1522, 428, 567, 2620, 2653, 1499, 68, 701, 3648, 543, 2004, 584, 495, 208, 1497, 72, 461, 557, 459, 586, 150, 1127, 2669, 3, 2539, 347, 194, 287, 408, 1440, 175, 3649, 184, 3369]. status [2636, 1462, 1113, 124, 3414, 2880, 93, 3636, 2931, 830, 2050, 1330, 997, 1407, 3012, 3514, 1155]. stay [2187, 2572, 1187, 757]. stayer [2793, 2555]. STEIN [2863, 2683, 3558]. Stein-like [2683]. Stein-type [3558]. stem [932]. Step [1660, 3027, 3277, 269, 3461, 2391]. Step-up [1660, 269]. Stephen [289]. STEPP [2400]. Stepped [1805, 2836, 2272, 3245, 3342, 2999, 2478, 3584, 2828, 3107, 2766, 2455]. Stepped-wedge [1805, 3342, 2999, 3584, 2766]. Stepwise [1794, 3375, 408, 166, 409]. stereotype [2473]. Steyerberg [647, 1592]. Stian [3059, 1522]. still [2891, 1136]. stimulation [2925, 2121]. Stochastic [1678, 3534, 501, 67, 552, 2018, 1848, 1979, 1376, 455, 293]. stop [3551, 3549, 3548]. stopping [1216, 1357, 2545, 103, 2668, 3434]. stories [1457]. strains [563]. STRAND [3379]. strata [1327, 916]. Strategies [447, 3167, 624, 3428, 3477, 982, 705, 901, 1461, 2593, 2161, 601, 677, 3524, 144, 3165, 877, 2954]. strategy [786, 160, 1623, 3244, 2194, 1760, 148, 619, 1701, 2832, 2526, 805, 3033, 1639, 3210]. Stratification [2726, 318, 1814, 579, 1195, 1752, 3462, 1142, 3256, 2003, 714, 79, 1801, 2775]. stratified [2655, 3443, 1555, 771, 330, 1685, 667, 2103, 1624, 1252, 784, 2030, 173, 1714, 1996, 2526, 13, 3330, 468, 399, 3243, 765, 3490, 3456, 1776]. STRATOS [1789]. stratum [772, 1335, 3589]. stream [1836]. Strength [154]. strengthen [1328]. Strengthening [390, 1789]. strep [2454]. Streptococcus [753]. stroke [1331, 1056, 58, 618, 1690]. strong [1013, 748]. structural [1651, 1541, 3008, 1219, 389, 1514, 186, 3169, 1897, 3363, 1495, 1799, 2025, 979, 1494, 2503, 3364, 2707, 2672, 259, 1509, 1110, 677, 3475, 3393, 2450, 1632, 49, 775, 3214, 2867, 1265, 1889, 1981, 1845, 658, 801, 1855, 1182]. structural-nested [1514]. Structure [3443, 1897, 3047, 2952, 1802, 3395, 2245, 3005, 3153, 1540, 53, 2831, 2289, 1114, 3138, 1933, 3162, 1365, 802, 2718, 1564, 3575]. Structured [2472, 2177, 293, 1742]. structures [2484, 2961, 2601, 3253, 2852, 1644, 182, 2967]. Stuart [1320]. Studentized [2627]. Studies [1789, 560, 165, 116, 178, 3501, 216, 2053, 375, 2876, 1350, 447, 147, 1323, 23, 1970, 1781, 3220, 3415, 1818, 77, 28, 444, 1361, 833, 2083, 2323, 653, 101, 124, 1580, 3523, 1254, 2397, 284, 2519, 2000, 264, 3477, 3333, 1208, 2717, 1423, 107, 849, 819, 823, 2339, 2844, 3114, 1526, 2700, 1536, 2367, 2271, 1650, 3551, 102, 1195, 2100, 2272, 1967, 566, 2780, 1141, 2034, 1752, 1098, 174, 557, 2480, 2960, 3528, 1808, 3529, 2869, 3512, 3131, 146, 238, 507, 514, 2404, 1401, 1399, 2304, 3561, 1504, 1932, 756, 22, 2340, 3463, 1301, 183, 490, 2710, 3288, 2823]. studies [2457, 1611, 1405, 1867, 698, 2119, 863, 1026, 1166, 3578, 252, 733, 87, 813, 1111, 1551, 1461, 3222, 3198, 3398, 2511, 916, 48, 3619, 1830, 250, 1183, 2226, 2634, 2984, 620, 3509, 1896, 472, 3122, 907, 662, 2923, 3100, 748, 3405, 642, 1356, 1378, 2505, 64, 722, 2175, 2064, 1931, 1148, 508, 2635, 2088, 3013, 137, 492, 2686, 2874, 1978, 2352, 1123, 2338, 228, 3507, 1773, 1968, 1938, 2337, 1000, 1713, 1953, 2376, 741, 268, 619, 2212, 2152, 3549, 2425, 1419, 3548, 717, 1444, 1012, 1590, 1800, 1347, 3356, 2859, 721, 521, 1389, 144, 2773, 1042, 1875, 352, 3595, 2875, 2825]. studies [734, 518, 1653, 2794, 2598, 3270, 1253, 1343, 3033, 2786, 2402, 828, 789, 2440, 394, 1244, 3294, 1400, 1917, 2063, 380, 1328, 1425, 2946, 3210, 513, 1903, 2922, 3652]. Study [3391, 1886, 3201, 3385, 142, 270, 2439, 2421, 1836, 1600, 143, 583, 931, 2875, 802, 618, 3117, 426, 1053, 1709, 3513, 970, 3371, 3008, 321, 2364, 1618, 1173, 2072, 1621, 114, 1882, 1272, 2466, 3168, 841, 842, 2343, 1304, 3178, 2098, 448, 2215, 962, 3155, 3159, 141, 2146, 494, 1188, 3225, 1355, 768, 808, 574, 519, 1677, 1979, 2021, 240, 2498, 1475, 1246, 945, 2042, 1023, 538, 3093, 1874, 254, 1702, 674, 1082, 1796, 1107, 89, 3060, 30, 2154, 1652, 622, 977, 3001, 3484, 901, 1249, 2381, 2442, 1116, 1431, 2574, 2726, 2047, 2617, 1998, 2485, 3609, 750, 2619, 2086, 1049]. study [3499, 1200, 3030, 26, 340, 1751, 95, 2874, 2077, 595, 3289, 1245, 2122, 2387, 1285, 1735, 2078, 3145, 2661, 1114, 997, 1128, 443, 3537, 475, 779, 49, 1609, 1389, 1868, 767, 3467, 2104, 2904, 554, 3330, 2723, 351, 2827, 1857, 2229, 782, 258, 2039, 2671, 3349, 1094, 801, 1679, 3543, 497, 2857, 691, 1150, 3115, 1258, 141, 49, 297]. study-level [2343]. study-specific [2661]. studying [263, 2155, 2899]. sub [390, 2040, 2907, 821]. sub-phenotypes [2907]. sub-population [821]. sub-Saharan [390, 2040]. subclasses [2495]. subclassification [3174]. Subcopula [2843]. Subcopula-based [2843]. subdistribution [2589, 90, 1728, 498, 1286]. subgraph [2499]. subgraph-level [2499]. Subgroup [2225, 578, 2914, 3480, 553, 2733, 3532, 3049, 3064, 2657, 2941, 1542, 1488, 2894, 1420, 1165, 988, 1191, 2367, 2732, 2662, 2022, 2573, 1204, 3589, 3045, 350, 3165, 3162, 3567, 3436, 3439, 3554, 3133, 2934]. Subgroups [3302, 391, 2988, 3288, 3566, 1918, 2486, 3440, 3436, 868]. subject [1381, 3074, 2767, 3205, 2200, 1770, 2966, 910, 2869, 3561, 1611, 249, 549, 2790, 954, 2244, 3315, 3141, 2859, 976, 342, 157, 3594, 831, 2943, 2865, 2600, 2231, 458, 829, 1721, 1776, 537]. subject-specific [1770, 342, 537]. subjective [227]. subjects [1561, 217, 1696, 2104, 2309, 2788]. sublingual [3201]. submissions [2332]. submodels [2929]. subpopulation [3227, 1976, 1687, 2405, 2400, 966]. subpopulations [553, 1927]. subsample [675]. subsampling [3000]. subset [1073, 3157, 1677, 1168, 2497, 484]. substance [628, 529, 964, 479, 1767, 2622]. substantial [2210]. substate [3380]. substitution [818, 1721]. substructure [3605]. subtype [2155]. subtypes [930, 2554]. succeed [1867]. Success [3211, 673, 1809, 3386]. successful [2736]. successive [1449]. sufficiency [3587]. Sufficient [824, 536, 1664, 1647]. suggests [672]. suicide [901, 3611]. suitability [2342]. sum [3070, 2307, 1507, 593, 3537, 1625, 1499]. sum-based [2307]. summaries [3107]. Summarising [1929, 1938]. summarize [544, 387]. summarized [2249, 257]. Summary [299, 2898, 2688, 1580, 2000, 1536, 90, 2480, 136, 3010, 3336, 3298, 774, 3120, 1169, 2798, 960, 3053, 871]. Sun [459].

Sungho [584]. Sungim [3656]. super [3382, 2707]. Superchain [1040]. superimposed [1836]. superiority [759, 3264, 214, 1919, 78, 2620, 2786, 1192]. Supervised [745, 3030]. supplementary [505, 3119]. Supplementation [2628]. support [34, 2151]. suppression [232]. surface [1703, 2489, 3488]. surfaces [3206, 3066]. surge [3054]. surgeons [2749]. surgeries [2036]. surgery [3495, 2388, 1662, 1772]. surgical [1188, 2515, 775, 2906]. surrogacy [2873, 536]. surrogate [3285, 2641, 3238, 2973, 3035, 335, 3237, 2180, 3485, 1351, 208, 874, 279, 1289, 1640, 1322, 758, 2846, 2228, 2687, 4, 1154, 700, 3386, 2179, 2375, 3461, 2859, 2775, 950, 2197]. Surrogates [1912, 1817, 2809, 3344, 847]. Surveillance [1949, 3562, 368, 403, 596, 1864, 1866, 2307, 1980, 1759, 3422, 803, 1597, 1475, 366, 378, 1865, 1643, 522, 2350, 1095, 478, 1598, 2276, 2532, 2211, 2942, 3408, 2251, 3219]. Survey [1681, 2512, 3117, 188, 417, 1104, 1105, 1959, 3380, 903, 2164, 637, 2310, 3087, 1921, 59, 1833, 1125, 3203, 371, 2159, 2532, 61, 57, 443, 1432, 205, 2502, 2830, 1474, 1880, 374, 1054, 30, 344, 1264]. surveys [905, 3017, 1308, 485, 2757, 1222, 103, 3586, 1120, 3540, 626]. Survival [1844, 2178, 2927, 3217, 787, 3285, 1517, 1929, 731, 1550, 2755, 1386, 2484, 2517, 955, 1087, 1496, 2070, 2551, 2630, 3316, 3491, 987, 2201, 3444, 3339, 34, 243, 1539, 1618, 1781, 1971, 1731, 1692, 323, 527, 1726, 3455, 841, 842, 1122, 1580, 3523, 3652, 2291, 2534, 998, 1307, 1779, 1674, 2188, 2919, 702, 172, 1906, 3447, 528, 1518, 1946, 1873, 2317, 2696, 1547, 2789, 206, 1018, 3570, 1871, 1167, 1238, 3587, 1064, 1944, 1948, 3197, 2016, 3379, 516, 1408, 2856, 41, 1552, 3397, 85, 3446, 800, 2436, 3097, 816, 2324, 3596, 1523, 663, 3303, 438, 2316, 730, 2880]. survival [1838, 2914, 2163, 254, 2205, 1353, 30, 70, 3057, 91, 1921, 31, 92, 1390, 58, 773, 746, 1111, 1969, 2917, 3198, 3017, 2866, 1826, 2920, 3188, 1159, 3065, 1177, 1707, 3278, 1210, 450, 3189, 3509, 2372, 3515, 2800, 3556, 1467, 464, 1297, 2041, 1509, 306, 215, 2086, 1180, 173, 292, 1633, 3643, 1256, 1724, 760, 2510, 954, 86, 1947, 3193, 1485, 707, 1071, 2032, 3507, 1945, 535, 2293, 774, 2010, 997, 2152, 272, 2248, 2709, 3108, 1227, 524, 843, 2318, 50, 2390, 2065, 3103, 386, 2418, 580, 2699, 1846, 2666, 342, 3338, 486, 2393, 428, 1185, 2054, 3323, 293, 1742, 2623, 3614]. survival [3409, 3313, 1997, 3427, 2786, 2483, 2369, 1673, 32, 343, 2653, 3112, 1639, 2135, 420, 1120, 1917, 3305, 3301, 327, 3387, 3292, 2993, 1847, 2020, 3072, 984, 1737, 781, 1887, 3625]. survival-cure [1633]. survival-sacrifice [3509]. survivor [1659, 1150]. survivors [274, 205]. survivorship [2613]. susceptibility [2970, 2280, 2321]. suspected [2498]. Sussex [289]. sustained [3201]. Swedish [1258]. sweet [1869]. Swiss [270, 583]. switch [2672, 3177, 3651]. switching [2210, 3202, 2399]. Symmetric [3622, 2293]. symmetrized [345]. symmetry [418, 895, 1708]. Symposium [1127]. symptoms [2025, 693]. synchronicity [3130]. Syndrome [696, 2813, 1655, 1096, 1449]. syndromic [803, 478]. synergy [1377]. synthesis [138, 855, 2467, 2998, 588, 154, 2874, 357, 3311, 1128, 1583, 2229, 2878]. synthesizing [3437]. synthetic [2911, 1308]. system [3422, 688, 1069, 1949, 370, 3563, 2736]. Systematic [1991, 659, 807]. systematically [1920, 2457, 1359]. systems [2427, 3034, 366, 1095, 1185, 2538].

T [3185, 432, 459, 1865, 347, 1144, 2216]. table [981]. Tables [2202, 3648, 1828, 3443, 3055, 2468, 407, 1519, 3555, 3650, 406, 3059, 2843]. Taguri [1266]. tail [2360]. tailed [3648, 3555, 3650]. Tailoring [1140]. taking [935]. Talbot [1889]. Tango [1106]. target [2988, 1612, 2048, 2236, 640, 1586, 3286, 3621]. Targeted [2401, 2611, 3110, 1579, 3470, 2031, 936, 2006, 2838, 1848, 2432, 2975, 3598, 2416, 3456, 2954, 3624]. Targeting [399, 468]. targets [1798]. task [2309]. task-related [2309]. taxonomy [2884, 2331, 2885]. Taylor [1440]. TB [563]. TDT [346]. TDT-type [346]. team [2333]. techniques [1140, 823, 2154]. technologies [192]. technology [848, 2393]. telephone [3586]. tell [3491]. temperature [2802, 3410]. temporal [2247, 233, 803, 2514, 3575, 1847]. temporally [1274, 3399]. Ten [1648]. tensor [2944, 2980]. tensor-based [2944]. Term [1264, 2364, 2758, 3114, 1894, 1468, 3397, 1689, 3036, 622, 3101, 98, 2014, 115]. terminal [2347, 3495, 2102, 3178, 3561, 1111, 2917, 1582, 732, 3645, 3124, 2680, 2455, 561, 2616, 589]. Terminating [2104]. termination [674, 1082, 1200]. terms [587]. ternary [2218]. Test [463, 2128, 437, 2433, 2129, 491, 3095, 1057, 1059, 2898, 2347, 2738, 2138, 2589, 633, 83, 3604, 1555, 346, 1463, 1534, 1827, 2571, 3070, 3423, 1535, 124, 1698, 838, 264, 3457, 3487, 2924, 2717, 819, 1171, 3204, 2473, 1547, 1754, 3009, 3096, 1829, 3209, 1654, 2780, 3403, 45, 3290, 1186, 3097, 1924, 212, 2319, 2880, 3591, 3655, 675, 551, 161, 338, 2130, 1549, 593, 1327, 3060, 556, 14, 2166, 895, 1708, 33, 1461, 2381, 3154, 3483, 250, 2574, 1582, 817, 1819, 1137, 758, 2627, 1574, 2150, 3102, 2064, 1377, 2294, 137, 1004, 2782, 707, 1892, 1938, 2337, 995, 2754, 617, 3475]. test [1860, 392, 2375, 2211, 3138, 251, 2699, 627, 3012, 2054, 301, 1859, 729, 3226, 258, 56, 3146, 2805, 576, 1033, 3305, 3350, 1129, 332, 1286, 3407, 943, 739, 2325, 2127]. test/estimation [3097]. Testing [1621, 759, 326, 2300, 1686, 3487, 2759, 2468, 1513, 1771, 1136, 2043, 3065, 820, 26, 1714, 1117, 591, 933, 2418, 1596, 1425, 624, 1205, 160, 2213, 2296, 2518, 902, 2768, 3150, 403, 699, 1466, 2970, 462, 2219, 1894, 108, 27, 214, 1544, 3626, 1097, 3599, 1807, 2324, 473, 2196, 207, 2539, 269, 1391, 1660, 3572, 3636, 1532, 1783, 3048, 464, 642, 656, 2161, 78, 3607, 2387, 3520, 779, 1014, 1436, 251, 222, 796, 797, 1638, 1753, 2905, 831, 805, 2049, 2502, 2929, 1615, 906, 668, 2286, 2540, 3032, 3301, 497, 1963, 1363, 1893, 1054, 104]. Testlet [205]. Tests [2066, 1906, 2751, 2046, 3479, 3648, 3658, 1710, 1618, 3220, 2182, 2961, 83, 771, 1820, 1899, 2499, 2142, 124, 2384, 18, 1603, 941, 3136, 3264, 2421, 1138, 2595, 156, 1186, 1925, 2989, 645, 1575, 2498, 2468, 3466, 675, 3462, 3241, 1077, 1956, 1016, 2381, 1072, 2426, 3127, 2574, 1007, 3519, 3077, 1896, 281, 273, 1467, 1394, 3166, 2030, 1373, 2, 64, 3481, 217, 173, 1622, 2686, 3555, 3650, 1738, 3419, 282, 3607, 305, 2141, 2314, 1888, 1178, 2168, 265, 2690, 1042, 931, 3152, 2392, 782, 1145, 1742, 2928, 2598, 2766, 3425, 2374, 3078, 3622, 668, 1196, 184, 481, 710, 1306, 2203, 374, 2788, 943]. tests [739, 1192]. textbook [289, 445]. Theil [345]. their [141, 2944, 1486, 3015, 396, 3172, 1070, 480, 1435, 1857, 3421, 3649]. Theophile [2846]. theorem [2407]. theoretic [938, 990, 1153, 3484, 1279, 781]. theoretical [935]. Theory [3277, 205, 1321, 1319, 1048, 1044, 2801, 1759, 740, 688, 2057, 106, 1068, 2968, 2591, 581, 662, 1320, 1046, 1045, 1546]. therapeutic [2566, 688, 1501, 3346, 2606]. therapies [2006, 2021, 3196, 1590, 3624, 2122]. therapists [1005]. therapy [2389, 661, 254, 523, 2088, 677, 1782, 2450, 1590, 532, 1422, 11, 794, 2646, 1917]. There [3351, 2848, 3182, 2030, 631, 3605]. thickness [1600]. things [2901]. Thinking [1789]. thorough [1183]. those [1496]. thoughts [1321]. threatening [2323]. Three [214, 2436, 2981, 1514, 1449, 635, 3590, 300, 1536, 487, 2149, 2019, 341, 1051, 5, 348, 863, 901, 347, 896, 3516, 2132, 2826, 2627, 285, 1756, 3355, 1358, 2333, 928, 439, 2391]. three-arm [487, 2149, 341, 5, 348, 863, 347, 2132, 2826, 2627, 3355, 1358, 928]. three-armed [1514]. three-class [300, 285]. three-decade [901]. three-dimensional [2019]. three-factor [439]. three-level [896, 3516]. Three-part [2981]. Three-sided [214]. three-state [1449]. three-step [2391]. Threshold [92, 2470, 477, 2676, 494, 3157, 3591, 3655, 3410, 3151, 3017, 539, 2452, 2698, 2179, 727, 726, 2493, 370, 3053, 1451]. threshold-free [3053]. threshold-selection [494]. thresholds [3120]. throat [2454]. throughput [3452, 1390, 1872, 2619, 978, 1900]. tied [3175]. Ties [1181, 416]. Tim [3505]. Time [1651, 3069, 1767, 2437, 2246, 300, 2480, 3001, 440, 1124, 594, 1152, 1156, 3123, 1742, 3476, 3560, 728, 266, 3370, 955, 1087, 1496, 3417, 1197, 1350, 2664, 3339, 23, 3199, 1934, 112, 2178, 1393, 3517, 2182, 460, 294, 1404, 2389, 2787, 2469, 3307, 1342, 2676, 1726, 1993, 2066, 2744, 2804, 2839, 3050, 3168, 2516, 1895, 2964, 2937, 3178, 1229, 1923, 2915, 3385, 2870, 1126, 1008, 2471, 3477, 1199, 2225, 81, 2200, 1247, 1300, 435, 3266, 2454, 2048, 2236, 2802, 1817, 2809, 233, 3109, 1871, 1167, 1238, 384, 2290, 2067, 1417, 270, 909, 3379, 2421, 713, 1759, 3275, 888, 1478, 336]. time [979, 1836, 3397, 3042, 1433, 3097, 2747, 1990, 582, 2553, 2273, 3015, 1388, 119, 1061, 2503, 818, 2546, 3182, 2992, 1226, 3200, 1549, 1016, 1695, 1740, 1339, 1777, 2007, 387, 3578, 204, 92, 1411, 87, 1551, 1477, 1837, 2348, 2672, 2654, 2818, 2609, 2852, 3249, 3592, 99, 2866, 876, 1826, 2920, 3541, 379, 2274, 1573, 837, 349, 1446, 3011, 2047, 1530, 817, 2036, 213, 3515, 1723, 2522, 2485, 517, 2041, 2350, 2452, 3372, 1579, 1842, 3025, 3564, 2635, 3251, 3223, 3324, 429, 1416, 2510, 1751, 2864, 1598, 515, 1588, 3167, 2276, 1268, 2990, 2052, 1071, 1406, 700, 1739, 3647, 3353, 451, 1176]. time [2181, 2275, 1273, 535, 2908, 2754, 2441, 2010, 3393, 997, 15, 826, 839, 1397, 392, 1221, 230, 2375, 1080, 1632, 295, 3039, 3075, 530, 3461, 524, 1006, 50, 3531, 3022, 3282, 2659, 525, 3467, 2251, 342, 1879, 2393, 634, 11, 2827, 422, 3594, 1009, 704, 2280, 3614, 1997, 2001, 343, 3148, 1639, 2135, 1368, 191, 960, 1567, 2326, 2756, 3242, 3158, 1917, 1831, 2622, 2076, 747, 1081, 3053, 2991, 1641, 2673, 3147, 2185, 3525]. Time-dependent [300, 2480, 440, 1742, 1934, 112, 2178, 1393, 3517, 2744, 2516, 1229, 1126, 2471, 1167, 270, 713, 336, 979, 1433, 2503, 1740, 387, 87, 1477, 2654, 2274, 1573, 3564, 2052, 1176, 2181, 230, 3075, 1879, 634, 1639, 2135, 1917, 2076, 1641, 3525]. time-invariant [2066, 3200, 1739]. time-lag [3242]. time-related [1826]. time-series [1751]. time-shift [2393]. Time-to-event [3001, 594, 1152, 266, 1496, 3417, 2664, 3517, 2182, 2389, 2787, 2469, 3307, 1993, 1923, 2870, 2225, 1300, 435, 1817, 2809, 1417, 2421, 888, 3097, 582, 2553, 2273, 2546, 1226, 1549, 1016, 1837, 2818, 3592, 2920, 1530, 213, 3025, 515, 1406, 700, 3647, 3353, 451, 2754, 15, 839, 2375, 1080, 295, 3039, 530, 3022, 3282, 2659, 2827, 422, 3594, 1009, 1997, 2001, 3148, 3158]. time-to-first-event [1268, 2990]. time-variant [3275]. Time-varying [1651, 1767, 2437, 2246, 1156, 955, 1087, 1350, 23, 294, 3069, 1726, 3477, 2454, 909, 3182, 2992, 1695, 92, 1411, 3001, 2348, 3249, 2866, 1579, 1842, 1416, 1071, 3393, 826, 392, 1221, 1632, 525, 3594, 2280, 343, 2756, 1831, 2622, 1081, 2185]. timeliness [2498]. Timely [376]. times [955, 1087, 1393, 3201, 3523, 3652, 941, 3606, 3097, 93, 58, 773, 746, 1225, 3175, 1223, 1071, 1352, 2152, 1006, 121, 1741, 486, 94, 1181, 865, 1847, 1852, 456]. timescale [2052]. timing [192]. tipping [3309]. tissue [1505, 3364, 1657]. Title [2079, 2282, 2297, 2311, 2327, 2344, 2361, 2377, 2394, 2410, 2428, 2445, 2460, 2474, 2092, 2490, 2506, 2523, 2541, 2108, 2124, 2143, 2156, 2170, 2191, 2206, 2221, 2237, 2252, 2267, 2559]. Tiwari [194]. Tobit [1214, 1291, 2870, 2496, 2519, 604]. TOC [2081, 2284, 2299, 2313, 2329, 2346, 2363, 2379, 2396, 2412, 2430, 2446, 2462, 2476, 2492, 2508, 2525, 2543, 2094, 2110, 2126, 2145, 2158, 2172, 2193, 2208, 2223, 2239, 2254, 2269]. today [857]. tolerability [297]. tolerance [2287, 3642, 2535]. tolerated [2583, 2580, 2494]. toll [967]. Tom [1648]. Tomasz [2846]. tonsil [340]. Too [2466, 2588]. tool [1420, 2150, 1261, 1163, 2065]. toolkit [1557]. tools [2342, 1329, 3586]. top [3222]. top-ranked [3222]. topical [547]. topics [2795, 2004, 3079, 3134]. total [3449, 371, 2028, 3367]. totality [3178]. totals [375, 1516]. toxicities [1206, 515]. Toxicity [75, 2741, 1396, 2584, 514, 2863, 110, 509, 274, 1039, 1782, 1609, 2579, 2568, 950]. Toxicity-dependent [2741]. Toxicity-evaluation [75]. toxicological [1491]. trace [1351]. trachomatis [1123]. Tracing [3100, 351]. tracking [1368]. tract [2204, 1189, 1601]. tradeoff [2898]. Traditional [1379, 692, 688, 690, 696, 694, 695, 693, 691]. train [2556]. trains [2121, 1421]. trait [1119, 3310, 1041, 2794, 3002, 710, 1425]. trait-based [1425]. traits [836, 3087, 2166, 416, 1093, 316, 3210]. trajectories [3505, 3503, 528, 3402, 2069, 3398, 965, 3247, 3504, 2782, 2622]. trajectory [3459, 3322, 3458, 3215, 1589]. transcriptome [3360]. transfers [2722]. transformation [1281, 2365, 2288, 1211, 1017, 1283, 1908, 2324, 1298, 2746, 1225, 3362, 3581, 2735, 3368, 1282, 1712, 589]. transformations [1280, 415]. transformative [837]. transformed [2536]. Transforming [1144]. transfused [932]. transfusion [1325, 2813, 2086]. transient [3636]. transition [2701, 1631, 2919, 1680, 2957, 919, 2137, 1739, 532, 1264, 66, 681]. transition-specific [2919]. transitions [3329]. translation [2159]. translational [107, 571]. transmission [3004, 2952, 3544, 753, 206, 916, 2743, 210, 1879]. transmitted [1994, 917]. transplant [3397, 1200, 683, 1732, 1887]. transplantation [2182, 300, 932]. transtheoretical [2047]. trauma [2086]. treat [1891, 2560, 1417, 88, 1464, 1960, 213]. treated [3429, 168, 1652, 486]. Treatment [2934, 787, 1062, 786, 391, 1561, 3451, 2925, 2053, 2551, 2973, 3035, 2401, 1197, 335, 2178, 1430, 2210, 923, 2389, 1762, 2787, 1361, 3150, 3428, 3429, 1132, 1882, 936, 1025, 2851, 3178, 547, 1891, 96, 2431, 208, 1199, 2225, 947, 1073, 1420, 637, 941, 697, 466, 2367, 661, 2264, 1871, 1001, 1060, 270, 2147, 1384, 2527, 3157, 2999, 467, 1992, 1324, 2618, 1572, 2037, 1711, 1074, 1661, 1331, 2133, 2042, 2316, 588, 3038, 3584, 254, 3309, 3288, 1513, 705, 1640, 2694, 1395, 1810, 1360, 1151, 1999, 400, 1516, 2194, 3036, 1143, 1652, 3484, 3420, 1600, 2449, 2935, 1418, 187, 1461, 2258, 2593]. treatment [2672, 2693, 3202, 2511, 553, 1960, 2043, 2285, 3191, 3636, 1918, 2486, 2274, 1284, 2226, 2726, 3110, 3589, 2617, 213, 1984, 464, 3166, 662, 255, 2041, 3609, 573, 3101, 3401, 192, 1075, 1675, 2161, 245, 2228, 3063, 3061, 331, 159, 3440, 4, 1133, 1562, 2052, 1773, 1853, 2625, 1005, 2417, 601, 677, 583, 535, 1285, 1735, 3145, 2441, 2010, 3228, 924, 594, 1152, 3627, 1785, 1218, 3315, 3600, 3630, 2242, 1607, 3075, 530, 524, 3214, 2014, 3103, 3111, 60, 3615, 2526, 3460, 1801, 632, 3286, 1422, 3330, 2827, 2229, 782, 115, 1445, 743, 2668, 3421, 3434, 3649, 1856, 2639, 1935, 2351]. treatment [2400, 1577, 3160, 2369, 880, 2646, 1193, 3404, 2606, 3242, 3453, 1727, 1917, 3301, 2568, 497, 2257, 984, 2399, 1911, 3625, 246, 3407, 3143, 1733]. treatment-biomarker [2043]. treatment-covariate [2618, 2316, 1935]. treatment-dependent [3143]. treatment-effect-modifying [464]. treatment-induced [1801]. treatment-related [1856, 2639]. treatment-sensitive [3157]. treatment-specific [984]. treatment-subgroup [1420]. treatments [2655, 547, 697, 1141, 2034, 861, 1092, 2604, 863, 2134, 2096, 531, 1251, 631, 1178, 237, 3393, 15, 3474, 1632, 3567, 1936]. tree [563, 1918, 3440, 2152, 293, 1742, 2250]. tree-structured [293, 1742]. Trees [2318, 764, 2952, 1420, 2927, 3302, 1346, 2733, 2390, 3532, 3111, 3615, 3567, 3608]. trend [1172, 6, 330, 1693, 2781, 593, 2917, 194, 1010, 257, 374]. trend-renewal [1010]. trends [3476, 3370, 1768, 461, 1124, 1739, 1397, 655, 3436, 2514]. Trial [3258, 3260, 786, 391, 2655, 2106, 3279, 1982, 1514, 777, 2365, 3150, 3428, 759, 2006, 942, 1942, 2566, 2674, 1490, 947, 798, 1500, 882, 1547, 578, 466, 3344, 2628, 487, 2149, 3245, 3606, 3342, 3559, 136, 44, 467, 884, 341, 100, 3176, 2998, 2618, 465, 1711, 2082, 3246, 303, 1441, 3381, 1360, 2194, 3332, 690, 1099, 1101, 58, 3634, 3633, 1279, 876, 1624, 3542, 1699, 3261, 437, 2433, 2617, 213, 1100, 509, 464, 511, 1687, 373, 1075, 2864, 1263, 3193, 1562, 1039, 1793, 3257, 15, 1440, 1218, 24, 1632, 1616, 98, 2014, 767, 3469, 2775, 3343, 3211, 2335, 2828, 3357, 1558]. trial [1792, 743, 3421, 3649, 658, 3437, 606, 3259, 1655, 1917, 1791, 436, 1690, 2922, 768, 2325]. trial-level [947, 3344]. trials [3285, 2561, 3546, 1454, 2657, 875, 1561, 639, 1863, 2911, 1608, 3370, 2648, 3513, 1697, 2973, 3035, 1398, 1858, 2465, 2401, 3345, 1370, 2764, 3444, 1595, 2563, 2364, 1934, 638, 972, 1806, 244, 1678, 949, 1430, 2210, 2729, 633, 2787, 477, 1688, 862, 155, 2259, 291, 1862, 771, 780, 1310, 1612, 298, 39, 719, 2235, 3227, 434, 239, 2383, 25, 1977, 2308, 2581, 2728, 1325, 1923, 887, 2765, 2560, 872, 172, 1778, 2838, 2382, 462, 1379, 2890, 2884, 1073, 1396, 552, 860, 1941, 1165, 271, 938, 1518, 2862, 874, 3264, 988, 2367, 661, 1869, 3473, 579, 768, 1829]. trials [108, 1473, 2403, 3571, 27, 2547, 2570, 138, 3395, 461, 1269, 1910, 2147, 3559, 516, 888, 1700, 1795, 2987, 990, 1986, 1805, 2999, 680, 1334, 2021, 1186, 2478, 3265, 3097, 2783, 1324, 381, 1523, 1809, 866, 2302, 1074, 1331, 1061, 1224, 1401, 1092, 1194, 1399, 2133, 2604, 2061, 2241, 869, 667, 1091, 1336, 948, 2546, 154, 3038, 3486, 551, 150, 2463, 1515, 5, 348, 1962, 3309, 799, 1216, 1976, 3151, 2958, 2334, 2608, 1226, 2936, 1040, 1151, 991, 1549, 863, 790, 3306, 1777, 2007, 881, 873, 3448, 2449, 2935, 1249, 1656, 187, 347, 1031, 1811, 2224, 2593, 2837, 2693, 3308, 3631, 1448, 2256]. trials [791, 2912, 2863, 2573, 1131, 918, 1960, 2660, 2808, 8, 111, 867, 1907, 2285, 3073, 3516, 1252, 53, 641, 1116, 3278, 3464, 531, 1210, 1636, 311, 1878, 1581, 1904, 1984, 2800, 273, 3175, 1723, 2522, 75, 312, 769, 662, 2824, 2041, 3609, 534, 2586, 1373, 2447, 2132, 2826, 2627, 3617, 573, 1919, 177, 2161, 3046, 2582, 1714, 2331, 2885, 2562, 2578, 515, 1634, 3419, 1133, 1268, 1406, 700, 1207, 1790, 2122, 1005, 2227, 2417, 2564, 601, 3396, 630, 535, 109, 2754, 414, 3145, 2010, 148, 1736, 2107, 2892, 2432, 1358, 1357, 470, 1115, 2195, 2807, 1939, 1326, 2986, 2709, 1761, 1701, 2832]. trials [689, 2295, 879, 2861, 3533, 2242, 775, 2893, 3461, 1179, 220, 721, 712, 2065, 3103, 3111, 2659, 3165, 767, 2526, 2418, 580, 3460, 2775, 3384, 806, 983, 3049, 3107, 1672, 2583, 399, 3125, 518, 115, 3552, 1548, 2668, 3434, 1856, 2639, 704, 2039, 2671, 2766, 3346, 985, 665, 1563, 3313, 1997, 3187, 3137, 2620, 339, 2366, 2579, 2741, 181, 56, 499, 1673, 1936, 1459, 859, 1193, 1596, 1855, 3640, 865, 2606, 3242, 3158, 1400, 1917, 3305, 3301, 2650, 1052, 512, 944, 2416, 2567, 2565, 3072, 67, 380, 1737, 1985, 1893, 2399, 2494, 2605, 3133, 2934, 3383, 781, 1776, 871, 2695, 950, 1852, 1671]. trials [2799, 3089, 3088, 672, 3143, 1635, 107, 1191]. triangle [3656, 3442]. triangular [273]. triggering [3553]. trimmed [325]. trivalent [752]. Trivariate [2555, 752, 1925, 950]. truck [3541]. true [3206, 12, 733, 788]. truncate [2341]. truncated [1317, 2689, 2246, 2702, 306, 1084, 1442, 1080, 498]. truncation [3135, 954, 3507, 2217]. tuberculosis [2976, 904, 2743]. Tukey [1771]. tumor [239, 963, 2388, 3085, 2554, 1773, 1533]. tumors [735, 163]. tumour [1194, 2912, 1633, 1343]. tumours [935]. tuning [3026]. tunnel [400, 1516, 2075]. Turner [3649]. Tutorial [2657, 3019, 869, 2573, 3133, 2264, 2025, 3529, 3110, 1251, 3184]. Tweedie [3173]. Twenty [2111, 2113, 2112, 2115, 2114]. Twenty-five [2111, 2113, 2112, 2115, 2114]. Twice [1495]. Twice-weighted [1495]. twin [645, 1493, 793]. twins [115]. Two [3648, 2738, 987, 10, 3027, 972, 488, 3604, 1924, 866, 3463, 1962, 901, 3633, 2660, 2808, 3449, 2923, 3037, 1765, 3555, 3650, 222, 2392, 782, 1831, 1052, 3387, 3095, 1628, 1608, 922, 751, 1059, 2031, 2451, 3206, 147, 1323, 457, 1806, 1781, 949, 1430, 2688, 2988, 28, 2787, 2685, 1349, 2624, 315, 3150, 2259, 1132, 1862, 771, 495, 3055, 124, 942, 898, 2870, 1130, 798, 2219, 1518, 1908, 905, 3009, 3096, 1473, 1967, 740, 1665, 1334, 3446, 1324, 2989, 2273, 989, 3389, 303, 1838, 711, 675, 183, 3151, 674, 1082, 991, 1611, 2633, 400, 2740, 1549, 2194, 93, 657, 2356]. two [2693, 3572, 111, 3542, 1699, 2026, 2096, 531, 1569, 3048, 1998, 2030, 255, 496, 573, 3481, 2038, 1075, 2879, 402, 812, 1714, 3545, 1614, 631, 357, 3579, 1713, 2168, 148, 1725, 15, 2195, 779, 1939, 3533, 1750, 251, 611, 1435, 796, 797, 734, 2054, 1145, 1558, 2580, 743, 3421, 3649, 1951, 1794, 3450, 3270, 2409, 1666, 2620, 2786, 3287, 1673, 2706, 1437, 2682, 2565, 1963, 458, 1306, 781, 1606, 1545, 1429, 943, 739, 2807]. two-arm [147, 1323, 2740, 2194, 1699, 1075, 1714, 1939, 2620, 2786]. two-armed [2259]. two-component [183]. two-dimensional [2580]. two-fold [1666]. two-level [2259, 657]. Two-part [1924, 3449, 2870, 1750, 458]. two-phase [1908, 1473, 1998, 3579, 3450, 3270, 2706, 1437]. Two-sample [3604, 2660, 2808, 2392, 782, 3095, 2688, 315, 3009, 3096, 2807]. Two-stage [2738, 987, 972, 488, 866, 3463, 1962, 3633, 2923, 3037, 1765, 222, 1831, 1052, 3387, 949, 1430, 28, 2787, 2685, 1349, 2624, 3150, 1132, 1862, 771, 1130, 798, 2219, 1518, 3446, 2273, 989, 3389, 303, 711, 3151, 674, 1082, 991, 2633, 1549, 2693, 3572, 111, 2026, 3048, 2030, 255, 573, 3545, 148, 2195, 251, 796, 797, 734, 743, 3421, 3649, 1951, 1794, 2409, 1673, 2682, 2565, 1963, 781]. Two-step [3027]. Two-tailed [3648, 3555, 3650]. two-way [3055, 3542, 2038]. type [375, 1854, 1555, 2407, 476, 643, 1424, 2319, 154, 3539, 2660, 2808, 3558, 3048, 3183, 616, 3018, 1190, 2807, 757, 2257, 1854, 2315, 2044, 2942, 1147, 346]. Type-I [1854]. Type-II [2315]. types [1099, 2855, 3132].

U [3624]. U-BOIN [3624]. U.K [289]. U.S. [2997, 345, 390]. Ugandan [2916]. ulceration [532]. unavailable [221]. unbalanced [266, 467, 2044]. Unbiased [2417, 2653, 1858, 1216, 1742]. uncertain [3399, 532]. uncertainty [1841, 519, 3015, 22, 2119, 3058, 1840, 3314, 3101, 1524, 1362, 2260, 1616, 2913, 1937]. unconditional [2740]. unconfounded [1141, 2034]. uncover [2907]. Uncovering [1195, 1542]. Under-reported [3325, 2488, 3326]. underlying [591]. underpowered [970]. underreported [3565, 2930]. underrunning [1941]. Understanding [120, 2883, 2598, 3371, 1099, 1919, 787]. underused [3547]. undiagnosed [2820]. Unequal [3215, 968, 2696, 3245, 400, 698, 1516, 2740, 2417, 3145, 1725, 3460, 3367, 969]. unequally [685]. unethical [970]. unidirectional [3002, 3177, 3651]. Unified [3118, 1505, 174, 1391, 3138, 1185]. unifying [1898]. unimodal [2699]. unique [2089]. unit [2976, 3321]. United [904, 240, 164]. units [932, 1277]. univariate [2209, 1511, 1594]. University [107, 445]. unknown [2707, 2196, 3183, 796, 3143]. unmasked [639, 638]. unmasking [2266]. unmatched [1361, 102, 3284]. unmeasured [2176, 1959, 2380, 2796, 1654, 3446, 421, 1870, 2356, 671, 2769, 2918, 714, 2827, 835]. unobserved [3491, 3058]. unordered [2738]. Unplanned [975]. unpooled [62]. unrecognised [1599]. unrelated [284]. unresponsiveness [3201]. unspecified [1610]. unstratified [173]. unstructured [1213]. Unsupervised [1109]. unsynchronized [1221]. Untangling [3565]. unusual [1943]. Up-front [705]. updates [856]. updating [3029, 2493, 50, 2905]. uptake [2159, 606]. uranium [996]. urgency [2736, 683, 1732]. urinary [2204, 2854, 1189]. Urn [350]. USA [1315, 1293]. usage [2677]. Use [903, 1528, 909, 3246, 2485, 500, 2368, 1736, 2107, 628, 3459, 3322, 1709, 1496, 2070, 20, 479, 935, 1665, 800, 2611, 1441, 799, 2823, 1652, 503, 1072, 1255, 1509, 3401, 1282, 1305, 1978, 1263, 1245, 1938, 535, 247, 1440, 2904, 1889, 1981, 1845, 3099, 2766, 1213, 2798, 3210, 2303, 2117, 2106]. used [1496, 2498, 1074, 1249, 1607]. useful [2630, 2919, 2848, 1010, 2065]. usefulness [647, 322]. Using [3501, 1514, 2323, 284, 3114, 845, 1269, 1654, 2498, 2304, 1301, 1289, 1909, 2194, 913, 795, 3017, 1821, 2824, 3405, 3365, 2003, 2052, 2359, 3598, 333, 3214, 802, 1009, 205, 1343, 1264, 1641, 1372, 624, 1651, 1929, 1214, 2755, 1386, 2458, 754, 1439, 1709, 2053, 2551, 2973, 3035, 3371, 2876, 987, 2664, 3391, 920, 3329, 2590, 345, 389, 2210, 463, 3237, 188, 417, 1104, 2685, 2758, 135, 2624, 3429, 372, 1463, 2209, 1535, 298, 39, 2964, 3374, 1449, 239, 124, 716, 369, 2006, 1864, 1866, 1375, 12, 1674, 547, 2215, 128, 1891, 715, 105, 2431, 81, 2355, 1396, 552, 2717, 3637]. using [1568, 3465, 3544, 1470, 1206, 2406, 2415, 466, 785, 233, 3551, 1195, 487, 505, 330, 2683, 1995, 921, 2760, 574, 1060, 2407, 2570, 2413, 2025, 3275, 2089, 1681, 3422, 1752, 1409, 2612, 2956, 1408, 1965, 2310, 1552, 2443, 910, 2667, 3529, 3446, 19, 1925, 224, 1990, 1987, 146, 2324, 2842, 1809, 2572, 2971, 119, 737, 1734, 1194, 2133, 1954, 3106, 1447, 1834, 2090, 1920, 3364, 1825, 667, 3038, 3510, 2707, 992, 106, 164, 1693, 2817, 2710, 3186, 2972, 3309, 3656, 3526, 1513, 1771, 317, 1241, 1865, 1322, 1405, 3060, 556, 91, 3628, 1777, 2007, 690, 3433, 40, 977, 678, 1316, 2957, 3442, 1220]. using [65, 187, 696, 1111, 1833, 2654, 3592, 3044, 791, 2912, 3058, 1202, 2877, 3541, 1308, 2226, 2426, 3065, 349, 1287, 3054, 1530, 1392, 1783, 1251, 1330, 646, 385, 539, 1237, 285, 2137, 1592, 2751, 3183, 684, 215, 2064, 1201, 2294, 3499, 163, 1543, 3046, 2687, 137, 889, 2084, 2864, 1524, 3193, 978, 1949, 2537, 1176, 2532, 3396, 1273, 583, 3386, 2597, 1989, 924, 2432, 604, 408, 57, 443, 1782, 3549, 2975, 3092, 1346, 2414, 3548, 171, 3317, 3207, 1288, 1632, 1761, 1303, 3630, 1800, 1436, 3461, 2318, 79, 3492, 3532, 2867, 976, 1435, 1933, 3531, 222, 1823, 2699]. using [2749, 1453, 3107, 1672, 2583, 1432, 304, 1719, 115, 2489, 1951, 69, 484, 951, 1312, 1935, 2502, 2496, 3489, 3313, 1997, 3190, 3160, 2402, 356, 1348, 3287, 166, 409, 1459, 2001, 1102, 2653, 1855, 2637, 2980, 3580, 960, 1244, 1567, 1813, 2326, 2756, 3091, 3400, 3453, 1533, 2682, 3635, 3305, 327, 2391, 1480, 3367, 710, 550, 984, 1425, 2453, 3090, 1911, 3625, 2954, 691, 2712, 1733]. usual [278]. Utility [2105, 2447, 3382, 1869, 1050, 3078, 3624]. Utility-based [2105, 2447, 3624]. utilization [1494, 3566, 168, 1022]. utilizing [3512, 2037, 3153, 1921]. uveitis [3388].

V [150]. vaccination [905, 752, 3559, 2956]. vaccinations [388]. Vaccine [1915, 3004, 768, 752, 3528, 1924, 2468, 1643, 265, 2211, 2942, 2859, 767, 2948, 3346, 394, 1244, 2606, 2085]. vaccines [2566, 2503]. vaginosis [1631]. Valid [2022, 1347]. validate [1082]. validated [1079]. Validating [2941, 1834, 1765, 1938]. Validation [2775, 1258, 2953, 3028, 1886, 2469, 2098, 3637, 1623, 2974, 1355, 2777, 3197, 3446, 1950, 758, 1888, 2179, 3461, 1775, 2435]. validatory [2719]. Validity [2082, 2286, 2540, 1154, 712, 3226, 2798, 2539]. valuations [164]. Value [1534, 787, 1517, 2641, 2898, 83, 2768, 584, 1375, 921, 1759, 3438, 2443, 85, 3547, 1424, 1825, 3, 3044, 231, 1139, 741, 2493, 2185, 2899]. valued [548]. values [235, 2415, 2413, 2987, 2753, 1077, 1241, 2633, 2022, 2719, 2424, 600, 1230, 1237, 3396, 1419, 2414, 1617, 283, 1]. Vance [969, 639]. Vanguard [143]. variability [3391, 291, 849, 3392, 1001, 1507, 1657, 1033, 3554, 1034]. Variable [3047, 2204, 3445, 1272, 2320, 2839, 1295, 2720, 1728, 3583, 867, 2077, 2189, 1436, 3467, 944, 691, 810, 1747, 1570, 1619, 1854, 2201, 3415, 763, 1349, 529, 1731, 488, 2721, 897, 3527, 1686, 3457, 1988, 1905, 1654, 2165, 3171, 1915, 3303, 150, 2463, 738, 3016, 1909, 1796, 3420, 503, 696, 1837, 1870, 3222, 1660, 3572, 8, 1431, 2923, 2255, 2289, 3324, 2177, 1362, 2479, 3084, 2532, 1285, 1735, 268, 408, 826, 1785, 1163, 2533, 223, 333, 1616, 530, 1712, 1529, 251, 976, 3132, 1875, 3518, 1742, 1951, 3076, 3226, 1653, 3243, 3349, 3427, 166, 409, 3105, 2653, 3387, 2651]. variable-enriched [3243]. Variable-ratio [2077]. variables [2091, 3327, 1806, 135, 2249, 1977, 2970, 2287, 2683, 981, 2621, 982, 1990, 3510, 992, 847, 2442, 2877, 2486, 3223, 651, 736, 599, 714, 1347, 1529, 1435, 1902, 1717, 634, 2482, 626, 2231]. Variance [2551, 590, 765, 1459, 1214, 2630, 1318, 2516, 2343, 1304, 3169, 3215, 510, 2874, 679, 2661, 2174, 2487, 3521, 2656, 441, 2279, 3367, 1170, 2386, 3115]. variances [910, 2958, 2449, 2935, 245, 2243, 2417, 1725, 1610, 3143]. variant [3275, 2166, 3350]. variants [36, 1535, 1107, 880]. Variation [2679, 1678, 318, 841, 2534, 2048, 2236, 3041, 2436, 2163, 3094, 2174, 272, 3317, 2058, 3090]. Variational [2095, 3545]. variations [537]. various [3122, 1218, 2788]. varying [1651, 955, 1087, 1350, 23, 294, 1274, 155, 2259, 3069, 1726, 3477, 1500, 1767, 2437, 2246, 2454, 909, 1874, 2481, 3182, 2992, 1695, 92, 913, 1411, 3001, 2348, 3249, 3202, 2866, 3247, 682, 1579, 1842, 2635, 1416, 2510, 1751, 1071, 2010, 3393, 826, 392, 1221, 1527, 1632, 1156, 525, 3594, 2280, 343, 1757, 2756, 3490, 1831, 2622, 1081, 2993, 2453, 2185, 672]. varying-coefficient [3202, 1527]. varying-stage [1500]. vector [34]. Veierød [1522]. veil [883]. ventricular [2018]. Verbeke [506]. verification [491, 1698, 1764, 3398, 250, 3488, 943, 739]. verified [124, 3012]. verifying [2595]. versatile [1468, 2054]. versus [3199, 1497, 947, 3169, 3559, 2604, 705, 2923, 192, 2013, 3068, 2425, 1498, 3408, 2557, 2833, 1368]. Vertical [123, 206]. very [1700, 221, 1114, 276, 277, 2632]. via [810, 1824, 815, 2201, 1743, 1147, 1820, 122, 3385, 208, 3452, 1295, 2367, 1990, 1915, 940, 3005, 588, 608, 3539, 2464, 3086, 3480, 3170, 2726, 231, 3139, 319, 1667, 2452, 1020, 915, 2289, 1148, 2697, 4, 1362, 3104, 15, 3537, 3108, 2733, 144, 3348, 2775, 3049, 1175, 2068, 1732, 577, 3536, 537]. Vickers [1059]. victims [3443]. view [3517]. viewed [1748]. Violations [1761]. violence [3565]. viral [2389, 424, 800, 1024, 2652]. virological [11]. virtual [115]. visit [3234]. visual [2685, 3492]. visualisation [1261]. visualization [367]. Visualizing [1395, 248]. vital [1921]. vitamin [2271]. vitro [1530]. VLAD [2906]. VLAD-based [2906]. voice [857]. void [1014]. vol [3656]. volume [902, 963, 670]. volume-outcome [670]. Vossoughi [1058]. vs [442, 108, 130, 287, 129, 133, 132, 131]. vulnerability [1176].

W [639, 447, 432, 647, 969, 1592, 3059, 1265]. W2 [370]. Wages [1559, 1793]. waiting [93, 1223]. Walter [3649]. Walters [289]. Wang [459, 1974, 777]. ward [1112, 1203]. ward-specific [1112]. wards [1768]. warning [1069]. washout [3077]. waste [2518]. Wavelet [2703, 2067, 1273]. Wavelet-based [2703]. way [3055, 3612, 3542, 2038, 324, 2392]. WCUSUM [1625]. weak [444, 763, 1822, 2062]. wedge [3277, 3245, 3342, 2595, 1805, 2999, 2478, 3584, 2836, 2828, 3107, 2766, 2455]. Weibull [1539, 1669, 1969, 1885]. weight [3459, 3322, 1219, 3604, 2853, 1895, 3458, 2596, 519, 2707, 1220, 893, 1190]. weight-model [2707]. Weighted [2596, 941, 2966, 2737, 255, 1614, 1715, 1364, 3390, 668, 1651, 3027, 2589, 20, 1325, 1008, 1908, 1495, 224, 240, 1077, 87, 3196, 3191, 2426, 349, 3519, 2831, 217, 256, 2990, 299, 2376, 604, 393, 1940, 50, 1625, 3286, 1338, 157, 2054, 3450, 2322, 3622, 3400, 3305, 3319]. weighting [2053, 2551, 3429, 2051, 1497, 661, 3023, 3528, 130, 40, 1250, 2726, 231, 3189, 1569, 750, 1084, 1442, 287, 1444, 129, 133, 1498, 1389, 132, 304, 131, 1459, 1733]. weights [3415, 2210, 1964, 1799, 2707, 1026, 1144, 1713, 2441, 3393, 1889, 1981, 1845]. Weinberg [833, 326, 235, 252, 2502]. Wenbin [1974]. Wendler [445]. West [289, 511]. Western [689, 693]. Westreich [1265]. WGEEs [2376]. where [1696]. whether [2595]. Which [2923, 631, 2049]. while [2740, 1136, 889]. Whitehead [586]. Whitney [3658, 1499, 18, 3136, 3204, 3290, 1819, 627, 1859]. Who [2454, 2089, 1519]. Wide [1516, 284, 819, 494, 235, 1536, 2240, 733, 1356, 2060, 717, 1012, 1800, 521, 3210]. widespread [2070]. Wiegand [408]. Wilcoxon [3658, 18, 3136, 3204, 45, 3290, 593, 1708, 1819, 627, 1499]. wild [1424]. wild-type [1424]. Wiley [289]. William [969]. Wim [2846]. Win [3266, 2737]. WinBUGS [795]. window [310]. windows [2321, 3645, 394, 1244]. winner [1518]. wise [2315, 1013]. Withdrawn [3497]. within [996, 3209, 910, 2618, 3410, 1611, 2512, 916, 3483, 349, 1344, 1718, 1227, 3524, 2104, 3107, 1253, 1094, 68, 1635]. within-cluster [3209, 3483, 349]. within-cycle [3410]. within-study [1094]. within-subject [910]. within-trial [2618]. without [1709, 2465, 2401, 920, 1627, 3517, 2073, 716, 1698, 1502, 2162, 1867, 3127, 1582, 788, 2661, 3425, 1257, 1306]. wombling [2854]. Women [1886, 141, 3381, 825, 1917]. Won [584, 3]. Woods [153]. work [1619]. workers [3559]. Working [597, 2334, 1945, 469, 1944, 1948, 2245, 2718]. workplace [3337]. workshop [871]. world [694, 1579, 3101, 3437]. wound [532]. wrong [247]. Wu [412, 446].

X [459, 344]. X-ray [344]. xenograft [262, 239, 1633]. xii [411]. ximelagatran [1609]. xiv [506]. Xu [2539]. xvi [2591]. xvii [520]. xviii [398]. xx [446]. xxi [2846]. xxii [106]. xxiii [460].

Years [2115, 1385, 2819, 2111, 2900, 141, 3079, 2112, 2113, 939, 2114, 859]. Yin [1793]. York [2012]. Youden [3020, 3021, 2087, 1899, 2142, 539, 285, 1079, 2141, 299, 2871, 1755, 1479, 1510]. younger [939].

Z [2539]. Zealand [1277]. zero [628, 1541, 529, 964, 1555, 3142, 2588, 1103, 2320, 2994, 1304, 3161, 479, 423, 2959, 2939, 1716, 895, 2981, 1125, 1769, 3192, 2549, 319, 3071, 1565, 1020, 2504, 2505, 2762, 2233, 125, 1744, 350, 2024, 891, 1772, 1849, 1102, 3124, 2622]. zero- [423]. zero-and-one-inflated [2981]. zero-enriched [350]. zero-inflated [628, 529, 964, 1555, 3142, 2588, 1103, 2320, 2994, 3161, 479, 2959, 1716, 2981, 1769, 3192, 2549, 319, 3071, 1565, 1020, 2233, 1744, 2024, 891, 1772, 1849, 1102, 2622]. zero-modified [2504, 2505]. zeros [45, 2959, 1716, 3449, 3071, 2698, 1913, 3640, 1182]. Zhang [1974]. Ziv [2846]. Zohar [1792, 1791]. Zou [175].

References

Won:2009:COM

[1] Sungho Won, Nathan Morris, Qing Lu, and Robert C. Elston. Choosing an optimal method to combine P -values. Statistics in Medicine, 28(11): 1537–1553, May 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letters [3, 584].

Nam:2009:NIT

[2] Jun mo Nam and Deukwoo Kwon. Non-inferiority tests for clustered matched-pair data. Statistics in Medicine, 28(12):1668–1679, May 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [184].

Koziol:2009:CCO

[3] J. A. Koziol. Comments on ‘Choosing an optimal method to combine P - values’ by S. Won, N. Morris, Q. Liu, R. C. Elston, Statistics in Medicine 2009; 28:1537–1553. Statistics in Medicine, 28(24):3043–3045, October 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1].

Qu:2006:QIT

[4] Yongming Qu and Michael Case. Quantifying the indirect treatment effect via surrogate markers. Statistics in Medicine, 25(2):223–231, January 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [208].

Kieser:2007:PAT

[5] Meinhard Kieser and Tim Friede. Planning and analysis of three-arm non-inferiority trials with binary endpoints. Statistics in Medicine, 26 (2):253–273, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [347].

Clegg:2009:EAA

[6] Limin X. Clegg, Benjamin F. Hankey, Ram Tiwari, Eric J. Feuer, and Brenda K. Edwards. Estimating average annual per cent change in trend analysis. Statistics in Medicine, 28(29):3670–3682, December 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comment [194].

Zou:2008:CCL

[7] G. Y. Zou and A. Donner. Construction of confidence limits about effect measures: a general approach. Statistics in Medicine, 27(10):1693–1702, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [175].

Liu:2009:SBC

[8] Guanghan F. Liu, Kaifeng Lu, Robin Mogg, Madhuja Mallick, and Devan V. Mehrotra. Should baseline be a covariate or dependent variable in analyses of change from baseline in clinical trials? Statistics in Medicine, 28(20):2509–2530, September 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [150].

Neuenschwander:2009:NPP

[9] Beat Neuenschwander, Michael Branson, and David J. Spiegelhalter. A note on the power prior. Statistics in Medicine, 28(28):3562–3566, December 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [72].

Bauer:2010:SBT

[10] Peter Bauer, Franz Koenig, Werner Brannath, and Martin Posch. Selection and bias — two hostile brothers. Statistics in Medicine, 29(1): 1–13, January 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tournoud:2010:AAT

[11] M. Tournoud, J. F. Etard, R. Ecochard, and V. DeGruttola. Adherence to antiretroviral therapy, virological response, and time to resistance in the Dakar cohort. Statistics in Medicine, 29(1):14–32, January 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crager:2010:GIU

[12] Michael R. Crager. Gene identification using true discovery rate degree of association sets and estimates corrected for regression to the mean. Statistics in Medicine, 29(1):33–45, January 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2010:CMC

[13] Nian-Sheng Tang, Hui-Qiong Li, and Man-Lai Tang. A comparison of methods for the construction of confidence interval for relative risk in stratified matched-pair designs. Statistics in Medicine, 29(1):46–62, January 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lage-Castellanos:2010:FDR

[14] Agustín Lage-Castellanos, Eduardo Martínez-Montes, Juan A. Hernández-Cabrera, and Lídice Galán. False discovery rate and permutation test: an evaluation in ERP data analysis. Statistics in Medicine, 29(1):63–74, January 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saville:2010:RMC

[15] Benjamin R. Saville, Amy H. Herring, and Gary G. Koch. A robust method for comparing two treatments in a confirmatory clinical trial via multivariate time-to-event methods that jointly incorporate information from longitudinal and time-to-event data. Statistics in Medicine, 29(1): 75–85, January 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2010:CID

[16] Man-Lai Tang, Man-Ho Ling, Leevan Ling, and Guoliang Tian. Confidence intervals for a difference between proportions based on paired data. Statistics in Medicine, 29(1):86–96, January 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Novikov:2010:MAE

[17] I. Novikov, N. Fund, and L. S. Freedman. A modified approach to estimating sample size for simple logistic regression with one continuous covariate. Statistics in Medicine, 29(20):97–107, January 15, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Divine:2010:EDS

[18] George Divine, Alissa Kapke, Suzanne Havstad, and Christine L. M. Joseph. Exemplary data set sample size calculation for Wilcoxon–Mann– Whitney tests. Statistics in Medicine, 29(20):108–115, January 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holford:2010:IEM

[19] Theodore R. Holford, Keita Ebisu, Lisa A. McKay, Janneane F. Gent, Elizabeth W. Triche, Michael B. Bracken, and Brian P. Leaderer. Integrated exposure modeling: a model using GIS and GLM. Statistics in Medicine, 29(20):116–129, January 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chi:2010:SUW

[20] Yunchan Chi and Pei-Fang Su. The simultaneous use of weighted logrank and weighted Kaplan–Meier statistics with clustered right-censored data. Statistics in Medicine, 29(20):130–141, January 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2010:BSM

[21] Ji in Kim, Andrew B. Lawson, Suzanne McDermott, and C. Marjorie Aelion. Bayesian spatial modeling of disease risk in relation to multivariate environmental risk fields. Statistics in Medicine, 29(20):142–157, January 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Joo:2010:ERS

[22] Jungnam Joo, Minjung Kwak, Zehua Chen, and Gang Zheng. Efficiency robust statistics for genetic linkage and association studies under genetic model uncertainty. Statistics in Medicine, 29(20):158–180, January 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Basagana:2010:PSS

[23] X. Basagaña and D. Spiegelman. Power and sample size calculations for longitudinal studies comparing rates of change with a time-varying exposure. Statistics in Medicine, 29(2):181–192, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schouten:2010:SAS

[24] Hubert Schouten and Arnold Kester. A simple analysis of a simple crossover trial with a dichotomous outcome measure. Statistics in Medicine, 29(2):193–198, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clark:2010:BMA

[25] Allan B. Clark and Max O. Bachmann. Bayesian methods of analysis for cluster randomized trials with count outcome data. Statistics in Medicine, 29(2):199–209, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parker:2010:TQI

[26] Robert A. Parker. Testing for qualitative interactions between stages in an adaptive study. Statistics in Medicine, 29(2):210–218, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Glimm:2010:HTM

[27] Ekkehard Glimm, Willi Maurer, and Frank Bretz. Hierarchical testing of multiple endpoints in group-sequential trials. Statistics in Medicine, 29(2):219–228, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Browne:2010:OTS

[28] Ryan Browne, Stefan H. Steiner, and R. Jock MacKay. Optimal twostage reliability studies. Statistics in Medicine, 29(2):229–235, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2010:CRE

[29] Huazhen Lin, Danping Liu, and Xiao-Hua Zhou. A correlated randomeffects model for normal longitudinal data with nonignorable missingness. Statistics in Medicine, 29(2):236–247, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lagona:2010:MCC

[30] Francesco Lagona and Zhen Zhang. A missing composite covariate in survival analysis: a case study of the Chinese Longitudinal Health and Longevity Survey. Statistics in Medicine, 29(2):248–261, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawless:2010:EPE

[31] Jerald F. Lawless and Yan Yuan. Estimation of prediction error for survival models. Statistics in Medicine, 29(2):262–274, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wienke:2010:BSM

[32] A. Wienke, S. Ripatti, J. Palmgren, and A. Yashin. A bivariate survival model with compound Poisson frailty. Statistics in Medicine, 29(2):275– 283, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2010:CMS

[33] Lingling Li and Martin Kulldorff. A conditional maximized sequential probability ratio test for pharmacovigilance. Statistics in Medicine, 29 (2):284–295, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Belle:2010:ASL

[34] V. Van Belle, K. Pelckmans, J. A. K. Suykens, and S. Van Huffel. Additive survival least-squares support vector machines. Statistics in Medicine, 29(2):296–308, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bottai:2010:LQR

[35] Matteo Bottai, Bo Cai, and Robert E. McKeown. Logistic quantile regression for bounded outcomes. Statistics in Medicine, 29(2):309–317, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Calster:2010:IDI

[36] Ben Van Calster and Sabine Van Huffel. Integrated discrimination improvement and probability-sensitive AUC variants. Statistics in Medicine, 29(2):318–319, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nadarajah:2010:C

[37] Saralees Nadarajah. Correction. Statistics in Medicine, 29(2):320, January 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anderson:2010:FSF

[38] Keaven M. Anderson and Jason B. Clark. Fitting spending functions. Statistics in Medicine, 29(3):321–327, February 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheung:2010:EIE

[39] Yin Bun Cheung, Ying Xu, Sze Huey Tan, Felicity Cutts, and Paul Milligan. Estimation of intervention effects using first or multiple episodes in clinical trials: The Andersen–Gill model re-examined. Statistics in Medicine, 29(3):328–336, February 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2010:IPS

[40] Brian K. Lee, Justin Lessler, and Elizabeth A. Stuart. Improving propensity score weighting using machine learning. Statistics in Medicine, 29 (29):337–346, February 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2010:ABD

[41] Jianghua He, Daniel L. McGee, and Xufeng Niu. Application of the Bayesian dynamic survival model in medicine. Statistics in Medicine, 29 (29):347–360, February 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Long:2010:RSM

[42] Qi Long, W. Dana Flanders, Veronika Fedirko, and Roberd M. Bostick. Robust statistical methods for analysis of biomarkers measured with batch/experiment-specific errors. Statistics in Medicine, 29(29):361–370, February 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Szychowski:2010:PDC

[43] Jeff M. Szychowski, David L. Roth, Olivio J. Clay, and Mary S. Mittelman. Patient death as a censoring event or competing risk event in models of nursing home placement. Statistics in Medicine, 29(29):371– 381, February 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heo:2010:SSR

[44] Moonseong Heo, Yongman Kim, Xiaonan Xue, and Mimi Y. Kim. Sample size requirement to detect an intervention effect at the end of follow-up in a longitudinal cluster randomized trial. Statistics in Medicine, 29(29): 382–390, February 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hallstrom:2010:MWT

[45] Alfred P. Hallstrom. A modified Wilcoxon test for non-negative distributions with a clump of zeros. Statistics in Medicine, 29(29):391–400, February 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carstensen:2010:CMM

[46] Bendix Carstensen. Comparing methods of measurement: Extending the LoA by regression. Statistics in Medicine, 29(29):401–410, February 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zelterman:2010:SDS

[47] Daniel Zelterman, Alexander Tulupyev, Robert Heimer, and Nadia Abdala. Statistical design for a small serial dilution series. Statistics in Medicine, 29(29):411–420, February 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longford:2010:EES

[48] Nicholas T. Longford. Estimation of the effect size in meta-analysis with few studies. Statistics in Medicine, 29(4):421–430, February 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shortreed:2010:MDE

[49] Susan M. Shortreed and Andrew B. Forbes. Missing data in the exposure of interest and marginal structural models: a simulation study based on the Framingham Heart Study. Statistics in Medicine, 29(4):431–443, February 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steiner:2010:RAS

[50] Stefan H. Steiner and Mark Jones. Risk-adjusted survival time monitoring with an updating exponentially weighted moving average (EWMA) control chart. Statistics in Medicine, 29(4):444–454, February 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bayman:2010:DQI

[51] Emine Özgür Bayman, Kathryn Chaloner, and Mary Kathryn Cowles. Detecting qualitative interaction: a Bayesian approach. Statistics in Medicine, 29(4):455–463, February 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stoner:2010:OCE

[52] J. A. Stoner, B. G. Leroux, and M. Puumala. Optimal combination of estimating equations in the analysis of multilevel nested correlated data. Statistics in Medicine, 29(4):464–473, February 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2010:SCS

[53] Kaifeng Lu and Devan V. Mehrotra. Specification of covariance structure in longitudinal data analysis for randomized clinical trials. Statistics in Medicine, 29(4):474–488, February 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2010:BAJ

[54] Yang Xie, Wei Pan, Kyeong S. Jeong, Guanghua Xiao, and Arkady B. Khodursky. A Bayesian approach to joint modeling of protein-DNA binding, gene expression and sequence data. Statistics in Medicine, 29(4): 489–503, February 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheng:2010:RLD

[55] Jing Cheng, Lloyd J. Edwards, Mildred M. Maldonado-Molina, Kelli A. Komro, and Keith E. Muller. Real longitudinal data analysis for real people: Building a good enough mixed model. Statistics in Medicine, 29(4):504–520, February 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2010:CST

[56] John Whitehead, Michael Branson, and Susan Todd. A combined score test for binary and ordinal endpoints from clinical trials. Statistics in Medicine, 29(5):521–532, February 28, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schenker:2010:IAS

[57] Nathaniel Schenker, Trivellore E. Raghunathan, and Irina Bondarenko. Improving on analyses of self-reported data in a large-scale health survey by using information from an examination-based survey. Statistics in Medicine, 29(5):533–545, February 28, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2010:JML

[58] Ning Li, Robert M. Elashoff, Gang Li, and Jeffrey Saver. Joint modeling of longitudinal ordinal data and competing risks survival times and analysis of the NINDS rt-PA stroke trial. Statistics in Medicine, 29(5): 546–557, February 28, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2010:ANQ

[59] Yan Li, Barry I. Graubard, and Edward L. Korn. Application of nonparametric quantile regression to body mass index percentile curves from survey data. Statistics in Medicine, 29(5):558–572, February 28, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Subtil:2010:RNL

[60] F. Subtil and M. Rabilloud. Robust non-linear mixed modelling of longitudinal PSA levels after prostate cancer treatment. Statistics in Medicine, 29(5):573–587, February 28, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roy:2010:EAS

[61] Jason Roy and Walter F. Stewart. Estimation of age-specific incidence rates from cross-sectional survey data. Statistics in Medicine, 29(5):588– 596, February 28, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schisterman:2010:HPU

[62] Enrique F. Schisterman, Albert Vexler, Sunni L. Mumford, and Neil J. Perkins. Hybrid pooled-unpooled design for cost-efficient measurement of biomarkers. Statistics in Medicine, 29(5):597–613, February 28, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stallard:2010:C

[63] N. Stallard, T. Friede, M. Posch, F. Koenig, and W. Brannath. Correction. Statistics in Medicine, 29(5):614–615, February 28, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelson:2010:IRD

[64] Kerrie P. Nelson and Don Edwards. Improving the reliability of diagnostic tests in population-based agreement studies. Statistics in Medicine, 29(6):617–626, March 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2010:MSC

[65] Yimei Li, E. Paul Wileyto, and Daniel F. Heitjan. Modeling smoking cessation data with alternating states and a cure fraction using frailty models. Statistics in Medicine, 29(6):627–638, March 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2010:NMT

[66] Lei Yu, William S. Griffith, Suzanne L. Tyas, David A. Snowdon, and Richard J. Kryscio. A nonstationary Markov transition model for computing the relative risk of dementia before death. Statistics in Medicine, 29(6):639–648, March 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2010:SMP

[67] Xiaoxi Zhang and Qi Long. Stochastic modeling and prediction for accrual in clinical trials. Statistics in Medicine, 29(6):649–658, March 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2010:SBA

[68] Ma Yan, Gonzalez Della Valle Alejandro, Zhang Hui, and X. M. Tu. A U-statistics-based approach for modeling Cronbach coefficient alpha within a longitudinal data setting. Statistics in Medicine, 29(6):659–670, March 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2010:CBN

[69] Cuiling Wang and Charles B. Hall. Correction of bias from non-random missing longitudinal data using auxiliary information. Statistics in Medicine, 29(6):671–679, March 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lam:2010:MIA

[70] K. F. Lam, Ying Xu, and Tak-Lun Cheung. A multiple imputation approach for clustered interval-censored survival data. Statistics in Medicine, 29(6):680–693, March 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2010:CBM

[71] Richard J. Cook, Jerald F. Lawless, and Ker-Ai Lee. A copulabased mixed Poisson model for bivariate recurrent events under eventdependent censoring. Statistics in Medicine, 29(6):694–707, March 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gajewski:2010:CNP

[72] Byron J. Gajewski. Comments on ‘A note on the power prior’ by Neuenschwander B., Branson M. and Spiegelhalter D. J. Statistics in Medicine; DOI: 10.1002/sim.3722. Statistics in Medicine, 29(6):708–709, March 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [9].

Neuenschwander:2010:AR

[73] Beat Neuenschwander, Michael Branson, and David J. Spiegelhalter. Authors’ reply. Statistics in Medicine, 29(6):709–710, March 15, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hougaard:2010:P

[74] Philip Hougaard. Preface. Statistics in Medicine, 29(7-8):711, March– April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Messer:2010:TED

[75] Karen Messer, Loki Natarajan, Edward D. Ball, and Thomas A. Lane. Toxicity-evaluation designs for phase I/II cancer immunotherapy trials. Statistics in Medicine, 29(7-8):712–720, March–April 30–15, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2010:CMA

[76] Stephen Senn, Vladimir V. Anisimov, and Valerii V. Fedorov. Comparisons of minimization and Atkinson’s algorithm. Statistics in Medicine, 29(7-8):721–730, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bretz:2010:PCO

[77] Frank Bretz, Holger Dette, and Jose C. Pinheiro. Practical considerations for optimal designs in clinical dose finding studies. Statistics in Medicine, 29(7-8):731–742, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ohrn:2010:OGS

[78] Fredrik Öhrn and Christopher Jennison. Optimal group-sequential designs for simultaneous testing of superiority and non-inferiority. Statistics in Medicine, 29(7-8):743–759, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stampf:2010:ECI

[79] Susanne Stampf, Erika Graf, Claudia Schmoor, and Martin Schumacher. Estimators and confidence intervals for the marginal odds ratio using logistic regression and propensity score stratification. Statistics in Medicine, 29(7-8):760–769, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heinze:2010:BRS

[80] Georg Heinze and Rainer Puhr. Bias-reduced and separation-proof conditional logistic regression with small or sparse data sets. Statistics in Medicine, 29(7-8):770–777, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diya:2010:ERB

[81] L. Diya, E. Lesaffre, K. Van den Heede, W. Sermeus, and A. Vleugels. Establishing the relationship between nurse staffing and hospital mortality using a clustered discrete-time logistic model. Statistics in Medicine, 29(7-8):778–785, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ring:2010:SMH

[82] Arne Ring. Statistical models for heart rate correction of the QT interval. Statistics in Medicine, 29(7-8):786–796, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burman:2010:DTS

[83] Carl-Fredrik Burman and Vera Lisovskaja. The dual test: Safeguarding p-value combination tests for adaptive designs. Statistics in Medicine, 29(7-8):797–807, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Binder:2010:ALC

[84] H. Binder and W. Sauerbrei. Adding local components to global functions for continuous covariates in multivariable regression modeling. Statistics in Medicine, 29(7-8):808–817, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hielscher:2010:PVS

[85] T. Hielscher, M. Zucknick, W. Werft, and A. Benner. On the prognostic value of survival models with application to gene expression signatures. Statistics in Medicine, 29(7-8):818–829, March–April 30–15, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Porzelius:2010:GPE

[86] Christine Porzelius, Martin Schumacher, and Harald Binder. A general, prediction error-based criterion for selecting model complexity for highdimensional survival models. Statistics in Medicine, 29(7-8):830–838, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leffondre:2010:WCM

[87] Karen Leffondre, Willy Wynant, Zhirong Cao, Michal Abrahamowicz, Georg Heinze, and Jack Siemiatycki. A weighted Cox model for modelling time-dependent exposures in the analysis of case-control studies. Statistics in Medicine, 29(7-8):839–850, March–April 30–15, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laubender:2010:EAR

[88] R. P. Laubender and R. Bender. Estimating adjusted risk difference (RD) and number needed to treat (NNT) measures in the Cox regression model. Statistics in Medicine, 29(7-8):851–859, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laaksonen:2010:EPA

[89] M. A. Laaksonen, T. Härkänen, P. Knekt, E. Virtala, and H. Oja. Estimation of population attributable fraction (PAF) for disease occurrence in a cohort study design. Statistics in Medicine, 29(7-8):860–874, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grambauer:2010:PSH

[90] Nadine Grambauer, Martin Schumacher, and Jan Beyersmann. Proportional subdistribution hazards modeling offers a summary analysis, even if misspecified. Statistics in Medicine, 29(7-8):875–884, March–April 30– 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lambert:2010:ECP

[91] P. C. Lambert, P. W. Dickman, C. P. Nelson, and P. Royston. Estimating the crude probability of death due to cancer and other causes using relative survival models. Statistics in Medicine, 29(7-8):885–895, March– April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2010:TRS

[92] Mei-Ling Ting Lee, G. A. Whitmore, and Bernard A. Rosner. Threshold regression for survival data with time-varying covariates. Statistics in Medicine, 29(7-8):896–905, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lan:2010:CSO

[93] Ling Lan and Somnath Datta. Comparison of state occupation, entry, exit and waiting times in two or more groups based on current status data in a multistate model. Statistics in Medicine, 29(7-8):906–914, March– April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiao:2010:BBM

[94] Yongling Xiao and Michal Abrahamowicz. Bootstrap-based methods for estimating standard errors in Cox’s regression analyses of clustered event times. Statistics in Medicine, 29(7-8):915–923, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petersen:2010:SFM

[95] Liselotte Petersen, Thorkild I. A. Sørensen, and Per Kragh Andersen. A shared frailty model for case-cohort samples: Parent and offspring relations in an adoption study. Statistics in Medicine, 29(7-8):924–931, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dias:2010:CCM

[96] S. Dias, N. J. Welton, D. M. Caldwell, and A. E. Ades. Checking consistency in mixed treatment comparison meta-analysis. Statistics in Medicine, 29(7-8):932–944, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bowden:2010:MMS

[97] Jack Bowden, Dan Jackson, and Simon G. Thompson. Modelling multiple sources of dissemination bias in meta-analysis. Statistics in Medicine, 29(7-8):945–955, March–April 30–15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stallard:2010:CSP

[98] Nigel Stallard. A confirmatory seamless phase II/III clinical trial design incorporating short-term endpoint information. Statistics in Medicine, 29(9):959–971, April 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2010:SPM

[99] Xiaoyan Lin and Lianming Wang. A semiparametric probit model for case 2 interval-censored failure time data. Statistics in Medicine, 29(9): 972–981, April 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilton:2010:NTD

[100] Joan F. Hilton. Noninferiority trial designs for odds ratios and risk differences. Statistics in Medicine, 29(9):982–993, April 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chu:2010:BAE

[101] Rong Chu, Paul Gustafson, and Nhu Le. Bayesian adjustment for exposure misclassification in case-control studies. Statistics in Medicine, 29(9):994–1003, April 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gebregziabher:2010:PCL

[102] Mulugeta Gebregziabher, Paulo Guimaraes, Wendy Cozen, and David V. Conti. A polytomous conditional likelihood approach for combining matched and unmatched case-control studies. Statistics in Medicine, 29(9):1004–1013, April 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wagner:2010:NSR

[103] James Wagner and Trivellore E. Raghunathan. A new stopping rule for surveys. Statistics in Medicine, 29(9):1014–1024, April 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zimmerman:2010:SAA

[104] Dale L. Zimmerman, Jie Li, and Xiangming Fang. Spatial autocorrelation among automated geocoding errors and its effects on testing for disease clustering. Statistics in Medicine, 29(9):1025–1036, April 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Desquilbet:2010:DRA

[105] Loic Desquilbet and François Mariotti. Dose-response analyses using restricted cubic spline functions in public health research. Statistics in Medicine, 29(9):1037–1057, April 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kelly:2010:BRA

[106] Patrick Kelly. Book reviews: Adaptive design theory and implementation using SAS and R. Mark Chang, Chapman and Hall/CRC, Boca Raton, 2008. No. of pages: xxii + 418. Price: $89.95. ISBN 10: 1-58488-962-4, ISBN 13: 978-1-58488-962-5. Statistics in Medicine, 29(9):1058, April 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:2010:PUP

[107] Jonas H. Ellenberg, Rosemarie Mick, and Susan S. Ellenberg. Proceedings of the University of Pennsylvania Annual Conference on Statistical Issues in Clinical Trials: Early, translational and proof of concept studies: The ‘go/no go’ decisions. Statistics in Medicine, 29(10):1059–1060, May 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilbert:2010:SDI

[108] Peter B. Gilbert. Some design issues in phase 2b vs phase 3 prevention trials for testing efficacy of products or concepts. Statistics in Medicine, 29(10):1061–1071, May 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubinstein:2010:SPT

[109] Larry V. Rubinstein, Seth M. Steinberg, Shivaani Kummar, Robert Kinders, Ralph E. Parchment, Anthony J. Murgo, Joseph E. Tomaszewski, and James H. Doroshow. The statistics of phase 0 trials. Statistics in Medicine, 29(10):1072–1076, May 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mandrekar:2010:MBP

[110] Sumithra J. Mandrekar, Rui Qin, and Daniel J. Sargent. Modelbased phase I designs incorporating toxicity and efficacy for single and dual agent drug combinations: Methods and challenges. Statistics in Medicine, 29(10):1077–1083, May 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2010:STS

[111] Junfeng Liu, Yong Lin, and Weichung Joe Shih. On Simon’s two-stage design for single-arm phase IIA cancer clinical trials under beta-binomial distribution. Statistics in Medicine, 29(10):1084–1095, May 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Benda:2010:MBA

[112] Norbert Benda. Model-based approaches for time-dependent dose finding with repeated binary data. Statistics in Medicine, 29(10):1096–1106, May 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nie:2010:CAR

[113] Lei Nie and Guoxing Soon. A covariate-adjustment regression model approach to noninferiority margin definition. Statistics in Medicine, 29 (10):1107–1113, May 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buyze:2010:DSR

[114] Jozefien Buyze, Bart Van Rompaye, and Els Goetghebeur. Designing a sequentially randomized study with adherence enhancing interventions for diabetes patients. Statistics in Medicine, 29(10):1114–1126, May 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vittinghoff:2010:ELT

[115] Eric Vittinghoff, Charles E. McCulloch, Claudine Woo, and Steven R. Cummings. Estimating long-term effects of treatment from placebocontrolled trials with an extension period, using virtual twins. Statistics in Medicine, 29(10):1127–1136, May 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ambrosius:2010:PSR

[116] Walter T. Ambrosius and Jonathan D. Mahnken. Power for studies with random group sizes. Statistics in Medicine, 29(10):1137–1144, May 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friede:2010:BSS

[117] Tim Friede and Heinz Schmidli. Blinded sample size reestimation with count data: Methods and applications in multiple sclerosis. Statistics in Medicine, 29(10):1145–1156, May 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hayen:2010:BRM

[118] Andrew Hayen. Book reviews: Medical biostatistics (2nd edn). Abhaya Indrayan, Chapman and Hall/CRC, Boca Raton, 2008. No. of pages:

824. Price: $99.95. ISBN 10: 1-58488-887-3, ISBN 13: 978-1-58488-887-1. Statistics in Medicine, 29(10):1157, May 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Isaman:2010:EIE

[119] D. J. M. Isaman, Jacob Barhak, and W. Ye. Erratum: Indirect estimation of a discrete-state discrete-time model using secondary data analysis of regression data. Statistics in Medicine, 29(10):1158, May 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aalen:2010:UDP

[120] Odd O. Aalen. Understanding disease processes. Statistics in Medicine, 29(11):1159–1160, May 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sweeting:2010:MSM

[121] M. J. Sweeting, V. T. Farewell, and D. De Angelis. Multi-state Markov models for disease progression in the presence of informative examination times: an application to hepatitis C. Statistics in Medicine, 29(11): 1161–1174, May 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2010:AIC

[122] Baojiang Chen, Grace Y. Yi, and Richard J. Cook. Analysis of intervalcensored disease progression data via multi-state models under a nonignorable inspection process. Statistics in Medicine, 29(11):1175–1189, May 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nicolaie:2010:VMP

[123] M. A. Nicolaie, Hans C. van Houwelingen, and H. Putter. Vertical modeling: a pattern mixture approach for competing risks modeling. Statistics in Medicine, 29(11):1190–1205, May 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chu:2010:EDP

[124] Haitao Chu, Yijie Zhou, Stephen R. Cole, and Joseph G. Ibrahim. On the estimation of disease prevalence by latent class models for screening studies using two screening tests with categorical disease status verified in test positives only. Statistics in Medicine, 29(11):1206–1218, May 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2010:MCE

[125] Patrick Royston, Willi Sauerbrei, and Heiko Becher. Modelling continuous exposures with a ‘spike’ at zero: a new procedure based on fractional polynomials. Statistics in Medicine, 29(11):1219–1227, May 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenberg:2010:PHM

[126] Philip S. Rosenberg and William F. Anderson. Proportional hazards models and age–period–cohort analysis of cancer rates. Statistics in Medicine, 29(11):1228–1238, May 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

MacNab:2010:BSC

[127] Ying C. MacNab. On Bayesian shared component disease mapping and ecological regression with errors in covariates. Statistics in Medicine, 29(11):1239–1249, May 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Das:2010:AED

[128] Sourish Das, Ofer Harel, Dipak K. Dey, Jonathan Covault, and Henry R. Kranzler. Analysis of extreme drinking in patients with alcohol dependence using Pareto regression. Statistics in Medicine, 29(11):1250–1258, May 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shuster:2010:EVN

[129] Jonathan J. Shuster. Empirical vs natural weighting in random effects meta-analysis. Statistics in Medicine, 29(12):1259–1265, May 30, 2010.

CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [130, 131, 132, 287] and reply [133].

Laird:2010:CEV

[130] Nan Laird, Garrett Fitzmaurice, and Xiao Ding. Comments on ‘Empirical vs natural weighting in random effects meta-analysis’. Statistics in Medicine, 29(12):1266–1267, May 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [129, 133].

Waksman:2010:CEV

[131] Joel A. Waksman. Comments on ‘Empirical vs natural weighting in random effects meta-analysis’. Statistics in Medicine, 29(12):1268–1269, May 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [129, 133].

Thompson:2010:CEV

[132] Simon G. Thompson and Julian P. T. Higgins. Comments on ‘Empirical vs natural weighting in random effects meta-analysis’. Statistics in Medicine, 29(12):1270–1271, May 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [129, 133].

Shuster:2010:RDE

[133] Jonathan J. Shuster, Randy C. Hatton, Leslie Hendeles, and Almut G. Winterstein. Reply to discussion of ‘Empirical vs natural weighting in random effects meta-analysis’. Statistics in Medicine, 29(12):1272–1281, May 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [129, 130, 131, 132].

Jackson:2010:EDL

[134] Dan Jackson, Ian R. White, and Simon G. Thompson. Extending DerSimonian and Laird’s methodology to perform multivariate random effects meta-analyses. Statistics in Medicine, 29(12):1282–1297, May 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burgess:2010:BMM

[135] Stephen Burgess, Simon G. Thompson, and Crp Chd Genetics Collaboration. Bayesian methods for meta-analysis of causal relationships estimated using genetic instrumental variables. Statistics in Medicine, 29 (12):1298–1311, May 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hemming:2010:MRP

[136] K. Hemming, J. L. Hutton, M. G. Maguire, and A. G. Marson. Metaregression with partial information on summary trial or patient characteristics. Statistics in Medicine, 29(12):1312–1324, May 30, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paul:2010:BBM

[137] M. Paul, A. Riebler, L. M. Bachmann, H. Rue, and L. Held. Bayesian bivariate meta-analysis of diagnostic test studies using integrated nested Laplace approximations. Statistics in Medicine, 29(12):1325–1339, May 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Govan:2010:CEB

[138] L. Govan, A. E. Ades, C. J. Weir, N. J. Welton, and P. Langhorne. Controlling ecological bias in evidence synthesis of trials reporting on collapsed and overlapping covariate categories. Statistics in Medicine, 29(12):1340–1356, May 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Horton:2010:PMI

[139] Nicholas J. Horton, Ian R. White, and James Carpenter. The performance of multiple imputation for missing covariates relative to complete case analysis. Statistics in Medicine, 29(12):1357, May 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giorgi:2010:AR

[140] Roch Giorgi, Aurélien Belot, Jean Gaudart, Guy Launoy, and The French Network of Cancer Registries Francim. Authors’ reply. Statistics in Medicine, 29(12):1358, May 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:2010:PCP

[141] Jonas H. Ellenberg. Proceedings of ‘The challenges and promises of a follow-up study of a randomly selected cohort of 100 000 pre and post conception women and their offspring through 21 years of life: design, implementation and analysis issues of the National Children’s Study’. Statistics in Medicine, 29(13):1359, June 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:2010:NCS

[142] Jonas H. Ellenberg. The National Children’s Study (NCS): Establishment and protection of the inferential base. Statistics in Medicine, 29 (13):1360–1367, June 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Montaquila:2010:SPI

[143] Jill M. Montaquila, J. Michael Brick, and Lester R. Curtin. Statistical and practical issues in the design of a national probability sample of births for the Vanguard study of the national Children’s study. Statistics in Medicine, 29(13):1368–1376, June 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Strauss:2010:ICE

[144] Warren J. Strauss, Louise Ryan, Michele Morara, Nicole Iroz-Elardo, Mark Davis, Matthew Cupp, Marcia G. Nishioka, James Quackenboss, Warren Galke, Haluk Özkaynak, and Peter Scheidt. Improving costeffectiveness of epidemiological studies via designed missingness strategies. Statistics in Medicine, 29(13):1377–1387, June 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Little:2010:D

[145] Roderick J. Little. Discussion. Statistics in Medicine, 29(13):1388–1390, June 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2010:ARP

[146] Ying Huang and Margaret Sullivan Pepe. Assessing risk prediction models in case-control studies using semiparametric and nonparametric methods. Statistics in Medicine, 29(13):1391–1410, June 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bartolucci:2010:CLE

[147] Francesco Bartolucci. On the conditional logistic estimator in two-arm experimental studies with non-compliance and before-after binary outcomes. Statistics in Medicine, 29(13):1411–1429, June 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sambucini:2010:BPS

[148] Valeria Sambucini. A Bayesian predictive strategy for an adaptive twostage design in phase II clinical trials. Statistics in Medicine, 29(13):

1430–1442, June 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2010:ECP

[149] Bing Han and Nelson Lim. Estimating conditional proportion curves by regression residuals. Statistics in Medicine, 29(13):1443–1454, June 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kenward:2010:LES

[150] Michael G. Kenward, Ian R. White, and James R. Carpenter. Letter to the Editor: Should baseline be a covariate or dependent variable in analyses of change from baseline in clinical trials? by G. F. Liu, K. Lu, R. Mogg, M. Mallick and D. V. Mehrotra, Statistics in Medicine 2009; 28:2509–2530. Statistics in Medicine, 29(13):1455–1456, June 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [8].

Liu:2010:AR

[151] G. F. Liu, K. Lu, R. Mogg, M. Mallick, and D. V. Mehrotra. Authors’ reply. Statistics in Medicine, 29(13):1457, June 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berger:2010:MSB

[152] Vance W. Berger. Making statistics boring again. Statistics in Medicine, 29(13):1458, June 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farewell:2010:WRH

[153] Vern Farewell and Tony Johnson. Woods and Russell, Hill, and the emergence of medical statistics. Statistics in Medicine, 29(14):1459–1476, June 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2010:SEN

[154] Seung-Ho Kang and Yi Tsong. Strength of evidence of non-inferiority trials — the adjustment of the type I error rate in non-inferiority trials with the synthesis method. Statistics in Medicine, 29(14):1477–1487, June 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Candel:2010:SSA

[155] Math J. J. M. Candel and Gerard J. P. Van Breukelen. Sample size adjustments for varying cluster sizes in cluster randomized trials with binary outcomes analyzed with second-order PQL mixed logistic regression. Statistics in Medicine, 29(14):1488–1501, June 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hand:2010:EDT

[156] David J. Hand. Evaluating diagnostic tests: The area under the ROC curve and the balance of errors. Statistics in Medicine, 29(14):1502–1510, June 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troxel:2010:WCP

[157] Andrea B. Troxel, Stuart R. Lipsitz, Garrett M. Fitzmaurice, Joseph G. Ibrahim, Debajyoti Sinha, and Geert Molenberghs. A weighted combination of pseudo-likelihood estimators for longitudinal binary data subject to non-ignorable non-monotone missingness. Statistics in Medicine, 29 (14):1511–1521, June 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shardell:2010:PMM

[158] Michelle Shardell, Gregory E. Hicks, Ram R. Miller, Patricia Langenberg, and Jay Magaziner. Pattern-mixture models for analyzing normal outcome data with proxy respondents. Statistics in Medicine, 29(14): 1522–1538, June 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pullenayegum:2010:ATE

[159] Eleanor M. Pullenayegum and Richard J. Cook. The analysis of treatment effects for recurring episodic conditions. Statistics in Medicine, 29 (14):1539–1558, June 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alosh:2010:CAA

[160] Mohamed Alosh and Mohammad F. Huque. A consistency-adjusted alpha-adaptive strategy for sequential testing. Statistics in Medicine, 29(15):1559–1571, July 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kosinski:2010:SSC

[161] Andrzej S. Kosinski, Ying Chen, and Robert H. Lyles. Sample size calculations for evaluating a diagnostic test when the gold standard is missing at random. Statistics in Medicine, 29(15):1572–1579, July 10, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vanhatalo:2010:AID

[162] Jarno Vanhatalo, Ville Pietiläinen, and Aki Vehtari. Approximate inference for disease mapping with sparse Gaussian processes. Statistics in Medicine, 29(15):1580–1607, July 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ostrovnaya:2010:MSI

[163] Irina Ostrovnaya, Adam B. Olshen, Venkatraman E. Seshan, Irene Orlow, Donna G. Albertson, and Colin B. Begg. A metastasis or a second independent cancer? Evaluating the clonal origin of tumors using array copy number data. Statistics in Medicine, 29(15):1608–1621, July 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kharroubi:2010:CUS

[164] Samer A. Kharroubi, Anthony O’Hagan, and John E. Brazier. A comparison of United States and United Kingdom EQ–5D health states valuations using a nonparametric Bayesian method. Statistics in Medicine, 29(15):1622–1634, July 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Akacha:2010:IDA

[165] Mouna Akacha and Norbert Benda. The impact of dropouts on the analysis of dose-finding studies with recurrent event data. Statistics in Medicine, 29(15):1635–1646, July 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wiegand:2010:PUM

[166] Ryan E. Wiegand. Performance of using multiple stepwise algorithms for variable selection. Statistics in Medicine, 29(15):1647–1659, July 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [408].

Wannemuehler:2010:LBM

[167] Kathleen A. Wannemuehler, Robert H. Lyles, Amita K. Manatunga, Metrecia L. Terrell, and Michele Marcus. Likelihood-based methods for estimating the association between a health outcome and left- or intervalcensored longitudinal exposure data. Statistics in Medicine, 29(16):1661– 1672, July 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laska:2010:ETP

[168] Eugene M. Laska, Morris Meisner, Joseph Wanderling, and Carole Siegel. Estimating treated prevalence and service utilization rates: Assessing disparities in mental health. Statistics in Medicine, 29(16):1673–1680, July 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sen:2010:BAC

[169] Ananda Sen, Mousumi Banerjee, Yun Li, and Anne-Michelle Noone. A Bayesian approach to competing risks analysis with masked cause of death. Statistics in Medicine, 29(16):1681–1695, July 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Conigliani:2010:BMA

[170] Caterina Conigliani. A Bayesian model averaging approach with noninformative priors for cost-effectiveness analyses. Statistics in Medicine, 29(16):1696–1709, July 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schwartz:2010:JBA

[171] Scott L. Schwartz, Alan E. Gelfand, and Marie L. Miranda. Joint Bayesian analysis of birthweight and censored gestational age using finite mixture models. Statistics in Medicine, 29(16):1710–1723, July 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dejardin:2010:JMP

[172] David Dejardin, Emmanuel Lesaffre, and Geert Verbeke. Joint modeling of progression-free survival and death in advanced cancer clinical trials. Statistics in Medicine, 29(16):1724–1734, July 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oakes:2010:CSU

[173] David Oakes and Changyong Feng. Combining stratified and unstratified log-rank tests in paired survival data. Statistics in Medicine, 29(16): 1735–1745, July 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harbron:2010:FUA

[174] Chris Harbron. A flexible unified approach to the analysis of pre-clinical combination studies. Statistics in Medicine, 29(16):1746–1756, July 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Singer:2010:LEC

[175] Julia Singer. Letter to the Editor: Construction of confidence limits about effect measures: a general approach, by G. Y. Zou and A. Donner, Statistics in Medicine 2008; 27:1693–1702. Statistics in Medicine, 29(16): 1757–1759, July 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [7].

Zou:2010:CCL

[176] G. Y. Zou and Allan Donner. Construction of confidence limits about effect measures: a general approach — reply. Statistics in Medicine, 29 (16):1760, July 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ning:2010:RAR

[177] Jing Ning and Xuelin Huang. Response-adaptive randomization for clinical trials with adjustment for covariate imbalance. Statistics in Medicine, 29(17):1761–1768, July 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

An:2010:MIA

[178] Di An, Roderick J. A. Little, and James W. McNally. A multiple imputation approach to disclosure limitation for high-age individuals in longitudinal studies. Statistics in Medicine, 29(17):1769–1778, July 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mahabadi:2010:ILS

[179] S. Eftekhari Mahabadi and M. Ganjali. An index of local sensitivity to non-ignorability for multivariate longitudinal mixed data with potential non-random dropout. Statistics in Medicine, 29(17):1779–1792, July 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2010:SDL

[180] Y. H. Joshua Chen, David L. DeMets, and K. K. Gordon Lan. Some dropthe-loser designs for monitoring multiple doses. Statistics in Medicine, 29(17):1793–1807, July 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2010:BDF

[181] John Whitehead, Helene Thygesen, and Anne Whitehead. A Bayesian dose-finding procedure for phase I clinical trials based only on the assumption of monotonicity. Statistics in Medicine, 29(17):1808–1824, July 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simpson:2010:LEA

[182] Sean L. Simpson, Lloyd J. Edwards, Keith E. Muller, Pranab K. Sen, and Martin A. Styner. A linear exponent AR(1) family of correlation structures. Statistics in Medicine, 29(17):1825–1838, July 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2010:TCN

[183] Inyoung Kim, Noah D. Cohen, Allen Roussel, and Naisyin Wang. A two-component nonlinear mixed effects model for longitudinal data, with application to gastric emptying studies. Statistics in Medicine, 29(17): 1839–1856, July 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2010:CNI

[184] Zhao Yang and Xuezheng Sun. Comments on ‘Non-inferiority tests for clustered matched-pair data’ by J. Nam and D. Kwon, Statistics in Medicine 2009; 28:1668–1679. Statistics in Medicine, 29(17):1857–1858, July 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2].

Nam:2010:AR

[185] Jun mo Nam and Deukwoo Kwon. Authors’ reply. Statistics in Medicine, 29(17):1859–1860, July 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2010:BAM

[186] Jing-Heng Cai, Xin-Yuan Song, and Yih-Ing Hser. A Bayesian analysis of mixture structural equation models with non-ignorable missing responses and covariates. Statistics in Medicine, 29(18):1861–1874, August 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2010:PTE

[187] Y. Li and J. M. G. Taylor. Predicting treatment effects using biomarker data in a meta-analysis of clinical trials. Statistics in Medicine, 29 (18):1875–1889, August 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brumback:2010:EAC

[188] Babette A. Brumback, Amy B. Dailey, Zhulin He, Lyndia C. Brumback, and Melvin D. Livingston. Efforts to adjust for confounding by neighborhood using complex survey data. Statistics in Medicine, 29(18):1890– 1899, August 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Biau:2010:MMP

[189] David J. Biau and Raphaël Porcher. A method for monitoring a process from an out of control to an in control state: Application to the learning curve. Statistics in Medicine, 29(18):1900–1909, August 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2010:SSS

[190] Inkyung Jung, Martin Kulldorff, and Otukei John Richard. A spatial scan statistic for multinomial data. Statistics in Medicine, 29(18):1910– 1918, August 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiong:2010:BMA

[191] Xiaoqin Xiong and Joel A. Dubin. A binning method for analyzing mixed longitudinal data measured at distinct time points. Statistics in Medicine, 29(18):1919–1931, August 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Noubary:2010:AAT

[192] Farzad Noubary and Michael D. Hughes. Assessing agreement in the timing of treatment initiation determined by repeated measurements of novel versus gold standard technologies with application to the monitoring of CD4 counts in HIV-infected patients. Statistics in Medicine, 29 (18):1932–1946, August 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Muggeo:2010:FAC

[193] Vito M. R. Muggeo and Miriam Tagliavia. A flexible approach to the crossing hazards problem. Statistics in Medicine, 29(18):1947–1957, August 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Muggeo:2010:CEA

[194] Vito M. R. Muggeo. Comment on ‘Estimating average annual per cent change in trend analysis’ by Clegg LX, Hankey BF, Tiwari R, Feuer EJ, Edwards BK, Statistics in Medicine 2009; 28:3670–3682. Statistics in Medicine, 29(18):1958–1960, August 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [6].

Clegg:2010:AR

[195] Limin X. Clegg, Benjamin F. Hankey, Ram Tiwari, Eric J. Feuer, and Brenda K. Edwards. Authors’ reply. Statistics in Medicine, 29(18):1961, August 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tunis:2010:CER

[196] Sean R. Tunis, Joshua Benner, and Mark McClellan. Comparative effectiveness research: Policy context, methods development and research infrastructure. Statistics in Medicine, 29(19):1963–1976, August 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gatsonis:2010:PRC

[197] Constantine Gatsonis. The promise and realities of comparative effectiveness research. Statistics in Medicine, 29(19):1977–1981, August 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lauer:2010:HMI

[198] Michael S. Lauer. The historical and moral imperatives of comparative effectiveness research. Statistics in Medicine, 29(19):1982–1984, August 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Normand:2010:WE

[199] Sharon-Lise T. Normand and Barbara J. McNeil. What is evidence? Statistics in Medicine, 29(19):1985–1988, August 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morton:2010:CCE

[200] Sally C. Morton. Comments on ‘Comparative effectiveness research: Policy context, methods development and research infrastructure’. Statistics in Medicine, 29(19):1989–1990, August 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubin:2010:LCE

[201] Donald B. Rubin. On the limitations of comparative effectiveness research. Statistics in Medicine, 29(19):1991–1995, August 30, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tunis:2010:RCC

[202] Sean R. Tunis, Joshua Benner, and Mark McClellan. Response to comments on ‘comparative effectiveness research’. Statistics in Medicine, 29 (19):1996–1997, August 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baurley:2010:DCP

[203] James W. Baurley, David V. Conti, W. James Gauderman, and Duncan C. Thomas. Discovery of complex pathways from observational data. Statistics in Medicine, 29(19):1998–2011, August 30, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawson:2010:STL

[204] Andrew B. Lawson, Hae-Ryoung Song, Bo Cai, Md Monir Hossain, and Kun Huang. Space-time latent component modeling of geo-referenced health data. Statistics in Medicine, 29(19):2012–2027, August 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2010:UTR

[205] Xiaohui Wang, Su Baldwin, Howard Wainer, Eric T. Bradlow, Bryce B. Reeve, Ashley W. Smith, Keith M. Bellizzi, and Kathy B. Baumgartner. Using testlet response theory to analyze data from a survey of attitude change among breast cancer survivors. Statistics in Medicine, 29 (19):2028–2044, August 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Frydman:2010:EOS

[206] Halina Frydman and Michael Szarek. Estimation of overall survival in an ‘illness-death’ model with application to the vertical transmission of HIV-1. Statistics in Medicine, 29(19):2045–2054, August 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kordzakhia:2010:MBA

[207] George Kordzakhia, Ohidul Siddiqui, and Mohammad F. Huque. Method of balanced adjustment in testing co-primary endpoints. Statistics in Medicine, 29(19):2055–2066, August 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dibaj:2010:LEM

[208] Shiva Dibaj, Soghrat Faghihzadeh, and Shohreh Jalaie. Letter to the Editor: Mistake on Quantifying the indirect treatment effect via surrogate markers by Y. Qu and M. Case Statistics in Medicine 2006; 25:223–231. Statistics in Medicine, 29(19):2067, August 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [4].

Qu:2010:AR

[209] Yongming Qu and Michael Case. Authors’ reply. Statistics in Medicine, 29(19):2068, August 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

ONeill:2010:ISR

[210] Philip D. O’Neill. Introduction and snapshot review: Relating infectious disease transmission models to data. Statistics in Medicine, 29(20):2069– 2077, September 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2010:MAR

[211] Tianxi Cai, Layla Parast, and Louise Ryan. Meta-analysis for rare events. Statistics in Medicine, 29(20):2078–2089, September 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jafarzadeh:2010:BER

[212] Seyed Reza Jafarzadeh, Wesley O. Johnson, Jessica M. Utts, and Ian A. Gardner. Bayesian estimation of the receiver operating characteristic curve for a diagnostic test with a limit of detection in the absence of a gold standard. Statistics in Medicine, 29(20):2090–2106, September 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matsuyama:2010:CRI

[213] Yutaka Matsuyama. A comparison of the results of intent-to-treat, perprotocol, and g-estimation in the presence of non-random treatment changes in a time-to-event non-inferiority trial. Statistics in Medicine, 29(20):2107–2116, September 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goeman:2010:TSH

[214] Jelle J. Goeman, Aldo Solari, and Theo Stijnen. Three-sided hypothesis testing: Simultaneous testing of superiority, equivalence and inferiority. Statistics in Medicine, 29(20):2117–2125, September 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nielsen:2010:AMS

[215] Jan Nielsen and Erik T. Parner. Analyzing multivariate survival data using composite likelihood and flexible parametric modeling of the hazard functions. Statistics in Medicine, 29(20):2126–2136, September 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2010:PDP

[216] Peter C. Austin. The performance of different propensity-score methods for estimating differences in proportions (risk differences or absolute risk reductions) in observational studies. Statistics in Medicine, 29(20):2137– 2148, September 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nofuentes:2010:CWK

[217] José Antonio Roldán Nofuentes and Juan Dios Luna del Castillo. Comparison of weighted kappa coefficients of multiple binary diagnostic tests done on the same subjects. Statistics in Medicine, 29(20):2149–2165, September 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Litiere:2010:C

[218] S. Litière, A. Alonso, and G. Molenberghs. Correction. Statistics in Medicine, 29(20):2166–2168, September 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2010:C

[219] M. L. Tang, L. Ling, M. H. Ling, and G. L. Tian. Correction. Statistics in Medicine, 27(16):2168, September 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sozu:2010:SSD

[220] Takashi Sozu, Tomoyuki Sugimoto, and Toshimitsu Hamasaki. Sample size determination in clinical trials with multiple co-primary binary endpoints. Statistics in Medicine, 29(21):2169–2179, September 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pekoz:2010:AMA

[221] Erol A. Peköz, Michael Shwartz, Cindy L. Christiansen, and Dan Berlowitz. Approximate models for aggregate data when individual-level data sets are very large or unavailable. Statistics in Medicine, 29(21): 2180–2193, September 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sudbury:2010:TST

[222] Aidan Sudbury. Two-stage testing using selection schemes. Statistics in Medicine, 29(21):2194–2199, September 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shentu:2010:NDC

[223] Yue Shentu and Minge Xie. A note on dichotomization of continuous response variable in the presence of contamination and model misspecification. Statistics in Medicine, 29(21):2200–2214, September 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsu:2010:RWK

[224] Chiu-Hsieh Hsu and Jeremy M. G. Taylor. A robust weighted Kaplan– Meier approach for data with dependent censoring using linear combinations of prognostic covariates. Statistics in Medicine, 29(21):2215–2223, September 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gasparrini:2010:DLN

[225] A. Gasparrini, B. Armstrong, and M. G. Kenward. Distributed lag nonlinear models. Statistics in Medicine, 29(21):2224–2234, September 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Naito:2010:MSC

[226] K. Naito, J. Udagawa, and H. Otani. Multidimensional standard curve for the development process of human fetuses. Statistics in Medicine, 29 (21):2235–2245, September 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fiedler:2010:JMP

[227] James Fiedler, Alan H. Feiveson, Matthew J. Hayat, Zalman Vaksman, Jason L. Boyd, and Lakshmi Putcha. Joint modeling of performance and subjective reporting to assess sensitivity to drug-induced sleepiness. Statistics in Medicine, 30(5):2246–2259, September 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rajan:2010:JMM

[228] Kumar B. Rajan and Sue E. Leurgans. Joint modeling of missing data due to non-participation and death in longitudinal aging studies. Statistics in Medicine, 30(5):2260–2268, September 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [557].

Savu:2010:ERR

[229] Anamaria Savu, Qi Liu, and Yutaka Yasui. Estimation of relative risk and prevalence ratio. Statistics in Medicine, 29(22):2269–2281, September 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shardell:2010:SRM

[230] Michelle Shardell, Gregory E. Hicks, Ram R. Miller, and Jay Magaziner. Semiparametric regression models for repeated measures of mortal cohorts with non-monotone missing outcomes and time-dependent covariates. Statistics in Medicine, 29(22):2282–2296, September 30, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lyles:2010:SAM

[231] Robert H. Lyles and Ji Lin. Sensitivity analysis for misclassification in logistic regression via likelihood methods and predictive value weighting. Statistics in Medicine, 29(22):2297–2309, September 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2010:LCJ

[232] Jenny J. Zhang and Molin Wang. Latent class joint model of ovarian function suppression and DFS for premenopausal breast cancer patients. Statistics in Medicine, 29(22):2310–2324, September 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gangnon:2010:MST

[233] Ronald E. Gangnon. A model for space-time cluster detection using spatial clusters with flexible temporal risk patterns. Statistics in Medicine, 29(22):2325–2337, September 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2010:AME

[234] Wen Li and Yongming Qu. Adjustment for the measurement error in evaluating biomarkers. Statistics in Medicine, 29(22):2338–2346, September 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finner:2010:HLC

[235] Helmut Finner, Klaus Strassburger, Iris M. Heid, Christian Herder, Wolfgang Rathmann, Guido Giani, Thorsten Dickhaus, Peter Lichtner, Thomas Meitinger, H.-Erich Wichmann, Thomas Illig, and Christian Gieger. How to link call rate and p-values for Hardy-Weinberg equilibrium as measures of genome-wide SNP data quality. Statistics in Medicine, 29(22):2347–2358, September 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carayol:2010:EDA

[236] Jerome Carayol, Frédéric Tores, Inke R. König, Jörg Hager, and Andreas Ziegler. Evaluating diagnostic accuracy of genetic profiles in affected offspring families. Statistics in Medicine, 29(22):2359–2368, September 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salanti:2010:ENA

[237] Georgia Salanti, Sofia Dias, Nicky J. Welton, A. E. Ades, Vassilis Golfinopoulos, Maria Kyrgiou, Davide Mauri, and John P. A. Ioannidis. Evaluating novel agent effects in multiple-treatments meta-regression. Statistics in Medicine, 29(23):2369–2383, October 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2010:SNB

[238] Yangxin Huang and Getachew Dagne. Skew-normal Bayesian nonlinear mixed-effects models with application to AIDS studies. Statistics in Medicine, 29(23):2384–2398, October 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choudhury:2010:AMA

[239] Kingshuk Roy Choudhury, Ian Kasman, and Greg D. Plowman. Analysis of multi-arm tumor growth trials in xenograft animals using phase change adaptive piecewise quadratic models. Statistics in Medicine, 29(23):2399– 2409, October 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2010:CAW

[240] Lan Huang, Ram C. Tiwari, Linda W. Pickle, and Zhaohui Zou. Covariate adjusted weighted normal spatial scan statistics with applications to study geographic clustering of obesity and lung cancer mortality in the United States. Statistics in Medicine, 29(23):2410–2422, October 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davies:2010:AKE

[241] Tilman M. Davies and Martin L. Hazelton. Adaptive kernel estimation of spatial relative risk. Statistics in Medicine, 29(23):2423–2437, October 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whittemore:2010:EHR

[242] Alice S. Whittemore. Evaluating health risk models. Statistics in Medicine, 29(23):2438–2452, October 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Belot:2010:FMC

[243] Aurélien Belot, Michal Abrahamowicz, Laurent Remontet, and Roch Giorgi. Flexible modeling of competing risks in survival analysis. Statistics in Medicine, 29(23):2453–2468, October 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bogowicz:2010:NSM

[244] P. Bogowicz, E. Gombay, and G. Heo. Nonparametric sequential monitoring of longitudinal trials. Statistics in Medicine, 29(24):2469–2479, October 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ozan:2010:AIC

[245] M. Ozgu Ozan and John Stufken. Assessing the impact of carryover effects on the variances of estimators of treatment differences in crossover designs. Statistics in Medicine, 29(24):2480–2485, October 30, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2010:BMB

[246] Kelly H. Zou, Martin O. Carlsson, and Sheila A. Quinn. Beta-mapping and beta-regression for changes of ordinal-rating measurements on Likert scales: a comparison of the change scores among multiple treatment groups. Statistics in Medicine, 29(24):2486–2500, October 30, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sabo:2010:WCG

[247] R. T. Sabo and N. R. Chaganty. What can go wrong when ignoring correlation bounds in the use of generalized estimating equations. Statistics in Medicine, 29(24):2501–2507, October 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2010:VAD

[248] Patrick Royston and Douglas G. Altman. Visualizing and assessing discrimination in the logistic regression model. Statistics in Medicine, 29 (24):2508–2520, October 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2010:IEA

[249] Yanhong Li, John J. Koval, Allan Donner, and G. Y. Zou. Interval estimation for the area under the receiver operating characteristic curve when data are subject to error. Statistics in Medicine, 29(24):2521–2531, October 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2010:BAS

[250] Ying Lu, Nandini Dendukuri, Ian Schiller, and Lawrence Joseph. A Bayesian approach to simultaneously adjusting for verification and reference standard bias in diagnostic test studies. Statistics in Medicine, 29 (24):2532–2543, October 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sperrin:2010:DET

[251] M. Sperrin and T. Jaki. Direct effects testing: a two-stage procedure to test for effect size and variable importance for correlated binary predictors and a binary response. Statistics in Medicine, 29(24):2544–2556, October 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2010:EIA

[252] Wen-Chung Lee, Liang-Yi Wang, and K. F. Cheng. An easy-to-implement approach for analyzing case-control and case-only studies assuming gene-environment independence and Hardy-Weinberg equilibrium. Statistics in Medicine, 29(24):2557–2567, October 30, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skaik:2010:LE

[253] Younis Abed El-Wahhab M. Skaik. Letter to the Editor. Statistics in Medicine, 29(24):2568, October 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kennedy:2010:EST

[254] Edward H. Kennedy, Jeremy M. G. Taylor, Douglas E. Schaubel, and Scott Williams. The effect of salvage therapy on survival in a longitudinal study with treatment by indication. Statistics in Medicine, 29(25):2569– 2580, November 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miyahara:2010:WKM

[255] Sachiko Miyahara and Abdus S. Wahed. Weighted Kaplan–Meier estimators for two-stage treatment regimes. Statistics in Medicine, 29(25): 2581–2591, November 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qi:2010:CMI

[256] Lihong Qi, Ying-Fang Wang, and Yulei He. A comparison of multiple imputation and fully augmented weighted estimators for Cox regression with missing covariates. Statistics in Medicine, 29(25):2592–2604, November 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Takahashi:2010:AGE

[257] Kunihiko Takahashi and Toshiro Tango. Assignment of grouped exposure levels for trend estimation in a regression analysis of summarized data. Statistics in Medicine, 29(25):2605–2616, November 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2010:CCL

[258] Bushi Wang and Xinping Cui. Consonant closed likelihood ratio test procedures with application to dose-response study. Statistics in Medicine, 29(25):2617–2630, November 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McAssey:2010:SES

[259] Michael P. McAssey and Fushing Hsieh. Slope estimation in structural line-segment heteroscedastic measurement error models. Statistics in Medicine, 29(25):2631–2642, November 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bandyopadhyay:2010:LMM

[260] Dipankar Bandyopadhyay, Victor H. Lachos, Carlos A. Abanto-Valle, and Pulak Ghosh. Linear mixed models for skew-normal/independent bivariate responses with an application to periodontal disease. Statistics in Medicine, 29(25):2643–2655, November 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moore:2010:ECB

[261] Reneé H. Moore, Robert H. Lyles, and Amita K. Manatunga. Empirical constrained Bayes predictors accounting for non-detects among repeated measures. Statistics in Medicine, 29(25):2656–2668, November 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheng:2010:IEQ

[262] Cheng Cheng and Jianrong Wu. Interval estimation of quantile ratios applied to anti-cancer drug screening by xenograft experiments. Statistics in Medicine, 29(26):2669–2678, November 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rota:2010:REM

[263] Matteo Rota, Rino Bellocco, Lorenza Scotti, Irene Tramacere, Mazda Jenab, Giovanni Corrao, Carlo La Vecchia, Paolo Boffetta, and Vincenzo Bagnardi. Random-effects meta-regression models for studying nonlinear dose-response relationship, with an application to alcohol and esophageal squamous cell carcinoma. Statistics in Medicine, 29(26):2679– 2687, November 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dendukuri:2010:BSS

[264] Nandini Dendukuri, Patrick Bélisle, and Lawrence Joseph. Bayesian sample size for diagnostic test studies in the absence of a gold standard: Comparing identifiable with non-identifiable models. Statistics in Medicine, 29(26):2688–2697, November 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:2010:SGL

[265] Mei-Chiung Shih, Tze Leung Lai, Joseph F. Heyse, and Jie Chen. Sequential generalized likelihood ratio tests for vaccine safety evaluation. Statistics in Medicine, 29(26):2698–2708, November 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arunajadai:2010:NMH

[266] Srikesh G. Arunajadai. A nonlinear model for highly unbalanced repeated time-to-event data: Application to labor progression. Statistics in Medicine, 29(26):2709–2722, November 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacqmin-Gadda:2010:SPL

[267] Hélène Jacqmin-Gadda, Cécile Proust-Lima, and Hélène Amiéva. Semiparametric latent process model for longitudinal ordinal data: Application to cognitive decline. Statistics in Medicine, 29(26):2723–2731, November 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salway:2010:BLV

[268] Ruth Salway, Duncan Lee, Gavin Shaddick, and Stephen Walker. Bayesian latent variable modelling in studies of air pollution and health. Statistics in Medicine, 29(26):2732–2742, November 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kwong:2010:SSD

[269] Koon Shing Kwong, Siu Hung Cheung, and Miin-Jye Wen. Sample size determination in step-up testing procedures for multiple comparisons with a control. Statistics in Medicine, 29(26):2743–2756, November 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gran:2010:SCA

[270] Jon Michael Gran, Kjetil Røysland, Marcel Wolbers, Vanessa Didelez, Jonathan A. C. Sterne, Bruno Ledergerber, Hansjakob Furrer, Viktor von Wyl, and Odd O. Aalen. A sequential Cox approach for estimating the causal effect of treatment in the presence of time-dependent confounding applied to data from the Swiss HIV cohort study. Statistics in Medicine, 29(26):2757–2768, November 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Everson-Stewart:2010:BCN

[271] Siobhan Everson-Stewart and Scott S. Emerson. Bio-creep in noninferiority clinical trials. Statistics in Medicine, 29(27):2769–2780, November 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seppa:2010:CFM

[272] Karri Seppä, Timo Hakulinen, Hyon-Jung Kim, and Esa Läärä. Cure fraction model with random effects for regional variation in cancer survival. Statistics in Medicine, 29(27):2781–2793, November 30, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McWilliams:2010:ADG

[273] Thomas P. McWilliams. An algorithm for the design of group sequential triangular tests for single-arm clinical trials with a binary endpoint. Statistics in Medicine, 29(27):2794–2801, November 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pintilie:2010:ACC

[274] Melania Pintilie, Yan Bai, Lingsong Yun, and David C. Hodgson. The analysis of case cohort design in the presence of competing risks with application to estimate the risk of delayed cardiac toxicity among Hodgkin lymphoma survivors. Statistics in Medicine, 29(27):2802–2810, November 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Comulada:2010:SSP

[275] W. Scott Comulada and Robert E. Weiss. Sample size and power calculations for correlations between bivariate longitudinal data. Statistics in Medicine, 29(27):2811–2824, November 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skene:2010:AVSa

[276] Simon S. Skene and Michael G. Kenward. The analysis of very small samples of repeated measurements i: an adjusted sandwich estimator. Statistics in Medicine, 29(27):2825–2837, November 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skene:2010:AVSb

[277] Simon S. Skene and Michael G. Kenward. The analysis of very small samples of repeated measurements II: a modified Box correction. Statistics in Medicine, 29(27):2838–2856, November 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tooze:2010:MEM

[278] Janet A. Tooze, Victor Kipnis, Dennis W. Buckman, Raymond J. Carroll, Laurence S. Freedman, Patricia M. Guenther, Susan M. Krebs-Smith, Amy F. Subar, and Kevin W. Dodd. A mixed-effects model approach for estimating the distribution of usual intake of nutrients: The NCI method. Statistics in Medicine, 29(27):2857–2868, November 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2010:LBA

[279] Debashis Ghosh, Michael R. Elliott, and Jeremy M. G. Taylor. Links between analysis of surrogate endpoints and endogeneity. Statistics in Medicine, 29(28):2869–2879, December 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kong:2010:MFA

[280] Xiangrong Kong, Ronald H. Gray, Lawrence H. Moulton, Maria Wawer, and Mei-Cheng Wang. A modeling framework for the analysis of HPV incidence and persistence: a semi-parametric approach for clustered binary longitudinal data analysis. Statistics in Medicine, 29(28):2880–2889, December 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mascha:2010:FAP

[281] Edward J. Mascha and Peter B. Imrey. Factors affecting power of tests for multiple binary outcomes. Statistics in Medicine, 29(28):2890–2904, December 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qin:2010:BCM

[282] Jing Qin and Biao Zhang. Best combination of multiple diagnostic tests for screening purposes. Statistics in Medicine, 29(28):2905–2919, December 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2010:BEM

[283] Ian R. White and John B. Carlin. Bias and efficiency of multiple imputation compared with complete-case analysis for missing covariate values. Statistics in Medicine, 29(28):2920–2931, December 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crossett:2010:UAM

[284] Andrew Crossett, Brian P. Kent, Lambertus Klei, Steven Ringquist, Massimo Trucco, Kathryn Roeder, and Bernie Devlin. Using ancestry matching to combine family-based and unrelated samples for genome-wide association studies. Statistics in Medicine, 29(28):2932–2945, December 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nakas:2010:ACP

[285] Christos T. Nakas, Todd A. Alonzo, and Constantin T. Yiannoutsos. Accuracy and cut-off point selection in three-class classification problems using a generalization of the Youden index. Statistics in Medicine, 29 (28):2946–2955, December 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2010:RAC

[286] Yang-Jin Kim. Regression analysis of clustered interval-censored data with informative cluster size. Statistics in Medicine, 29(28):2956–2962, December 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rucker:2010:CEV

[287] Gerta Rücker, Guido Schwarzer, James Carpenter, and Martin Schumacher. Comments on ‘Empirical vs natural weighting in random effects meta-analysis’ by J. J. Shuster, Statistics in Medicine 2009; 26, Published online, DOI: 10.1002/sim.3607. Statistics in Medicine, 29(28): 2963–2965, December 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [129, 133].

Shuster:2010:AR

[288] Jonathan J. Shuster. Authors’ reply. Statistics in Medicine, 29(28):2965– 2966, December 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prescott:2010:BRM

[289] Gordon J. Prescott. Book review: Medical statistics: a textbook for the Health Sciences (4th edn) Michael J. Campbell, David Machin and Stephen J. Walters, Wiley, Chichester, West Sussex, U.K., 2007. No. of pages 331. Price: $21.99, EUR26.40. ISBN: 978-0-470-02519-2. Statistics in Medicine, 29(28):2967–2968, December 10, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Henmi:2010:CIR

[290] Masayuki Henmi and John B. Copas. Confidence intervals for random effects meta-analysis and robustness to publication bias. Statistics in Medicine, 29(29):2969–2983, December 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carter:2010:CSV

[291] Ben Carter. Cluster size variability and imbalance in cluster randomized controlled trials. Statistics in Medicine, 29(29):2984–2993, December 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Odondi:2010:PSM

[292] Lang’o Odondi and Roseanne McNamee. Performance of statistical methods for analysing survival data in the presence of non-random compliance. Statistics in Medicine, 29(29):2994–3003, December 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wallace:2010:SMI

[293] Meredith L. Wallace, Stewart J. Anderson, and Sati Mazumdar. A stochastic multiple imputation algorithm for missing covariate data in tree-structured survival analysis. Statistics in Medicine, 29(29):3004– 3016, December 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bradshaw:2010:BPH

[294] Patrick T. Bradshaw, Joseph G. Ibrahim, and Marilie D. Gammon. A Bayesian proportional hazards regression model with non-ignorably missing time-varying covariates. Statistics in Medicine, 29(29):3017–3029, December 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siannis:2010:OSP

[295] F. Siannis, J. K. Barrett, V. T. Farewell, and J. F. Tierney. Onestage parametric meta-analysis of time-to-event outcomes. Statistics in Medicine, 29(29):3030–3045, December 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stijnen:2010:REM

[296] Theo Stijnen, Taye H. Hamza, and Pinar Özdemir. Random effects metaanalysis of event outcome in the framework of the generalized linear mixed model with applications in sparse data. Statistics in Medicine, 29 (29):3046–3067, December 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ivanova:2010:AAD

[297] Anastasia Ivanova, Kenneth Liu, Ellen Snyder, and Duane Snavely. Addendum to ‘An adaptive design for identifying the dose with the best efficacy/tolerability profile with application to a crossover dose-finding study’. Statistics in Medicine, 29(29):3068, December 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheung:2010:EEI

[298] Yin Bun Cheung, Ying Xu, Sze Huey Tan, Felicity Cutts, and Paul Milligan. Erratum: Estimation of intervention effects using first or multiple episodes in clinical trials: The Andersen–Gill model re-examined. Statistics in Medicine, 29(29):3068, December 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rucker:2010:SRC

[299] Gerta Rücker and Martin Schumacher. Summary ROC curve based on a weighted Youden index for selecting an optimal cutpoint in meta-analysis of diagnostic accuracy. Statistics in Medicine, 29(30):3069–3078, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Foucher:2010:TDR

[300] Y. Foucher, M. Giral, J. P. Soulillou, and J. P. Daures. Time-dependent ROC analysis for a three-class prognostic with application to kidney transplantation. Statistics in Medicine, 29(30):3079–3087, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Verde:2010:MAD

[301] Pablo E. Verde. Meta-analysis of diagnostic test data: a bivariate Bayesian modeling approach. Statistics in Medicine, 29(30):3088–3102, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Garcia-Zattera:2010:CMM

[302] M. J. García-Zattera, T. Mutsvari, A. Jara, D. Declerck, and E. Lesaffre. Correcting for misclassification for a monotone disease process with an application in dental research. Statistics in Medicine, 29(30):3103–3117, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jovic:2010:EMA

[303] Gordana Jovic and John Whitehead. An exact method for analysis following a two-stage phase II cancer clinical trial. Statistics in Medicine, 29(30):3118–3125, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ukoumunne:2010:CAE

[304] Obioha C. Ukoumunne, Elizabeth Williamson, Andrew B. Forbes, Martin C. Gulliford, and John B. Carlin. Confounder-adjusted estimates of the risk difference using propensity score-based weighting. Statistics in Medicine, 29(30):3126–3136, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rebora:2010:RNP

[305] Paola Rebora, Stefania Galimberti, and Maria Grazia Valsecchi. Robust non-parametric one-sample tests for the analysis of recurrent events. Statistics in Medicine, 29(30):3137–3146, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moreira:2010:SES

[306] Carla Moreira and Jacobo de Uña-Álvarez. A semiparametric estimator of survival for doubly truncated data. Statistics in Medicine, 29(30): 3147–3159, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oirbeek:2010:AHI

[307] R. Van Oirbeek and E. Lesaffre. An application of Harrell’s C-index to PH frailty models. Statistics in Medicine, 29(30):3160–3171, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bakoyannis:2010:MCR

[308] Giorgos Bakoyannis, Fotios Siannis, and Giota Touloumi. Modelling competing risks data with missing cause of failure. Statistics in Medicine, 29(30):3172–3185, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sumi:2010:IRR

[309] Mariko Sumi and Toshiro Tango. Inference on the rate ratio of recurrent events for the matched pairs design. Statistics in Medicine, 29(30):3186– 3193, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sweeting:2010:EDW

[310] Michael J. Sweeting, Daniela De Angelis, John Parry, and Barbara Suligoi. Estimating the distribution of the window period for recent HIV infections: a comparison of statistical methods. Statistics in Medicine, 29(30):3194–3202, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Makubate:2010:PAC

[311] Boikanyo Makubate and Stephen Senn. Planning and analysis of crossover trials in infertility. Statistics in Medicine, 29(30):3203–3210, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Metcalfe:2010:ACT

[312] Chris Metcalfe. The analysis of cross-over trials with baseline measurements. Statistics in Medicine, 29(30):3211–3218, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lecoutre:2010:FPB

[313] Bruno Lecoutre, Gérard Derzko, and Khadija ElQasyr. Frequentist performance of Bayesian inference with response-adaptive designs. Statistics in Medicine, 29(30):3219–3231, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klingenberg:2010:SCB

[314] B. Klingenberg. Simultaneous confidence bounds for relative risks in multiple comparisons to control. Statistics in Medicine, 29(30):3232– 3244, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buyse:2010:GPC

[315] Marc Buyse. Generalized pairwise comparisons of prioritized outcomes in the two-sample problem. Statistics in Medicine, 29(30):3245–3257, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yip:2010:GAA

[316] Benjamin H. Yip, Tron Anders Moger, and Yudi Pawitan. Genetic analysis of age-at-onset traits based on case-control family data. Statistics in Medicine, 29(30):3258–3266, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Komarek:2010:DAU

[317] Arnost Komárek, Bettina E. Hansen, Edith M. M. Kuiper, Henk R. van Buuren, and Emmanuel Lesaffre. Discriminant analysis using a multivariate linear mixed model with a normal mixture in the random effects distribution. Statistics in Medicine, 29(30):3267–3283, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Caceres:2010:MCD

[318] Alejandro Caceres, Xavier Basagaña, and Juan R. Gonzalez. Multiple correspondence discriminant analysis: an application to detect stratification in copy number variation. Statistics in Medicine, 29(30):3284–3293, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Molas:2010:HMM

[319] Marek Molas and Emmanuel Lesaffre. Hurdle models for multilevel zeroinflated data via h-likelihood. Statistics in Medicine, 29(30):3294–3310, December 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2011:E

[320] Ralph D’Agostino, Vern Farewell, and Joel Greenhouse. Editorial. Statistics in Medicine, 30(1):iii, January 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barrett:2011:SCR

[321] Jessica K. Barrett, Fotios Siannis, and Vern T. Farewell. A semicompeting risks model for data with interval-censoring and informative observation: an application to the MRC cognitive function and ageing study. Statistics in Medicine, 30(1):1–10, January 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2011:ENR

[322] Michael J. Pencina, Ralph B. D’Agostino, Sr., and Ewout W. Steyerberg. Extensions of net reclassification improvement calculations to measure usefulness of new biomarkers. Statistics in Medicine, 30(1):11–21, January 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chambless:2011:SMA

[323] Lloyd E. Chambless, Christopher P. Cummiskey, and Gang Cui. Several methods to assess improvement in risk prediction models: Extension to survival analysis. Statistics in Medicine, 30(1):22–38, January 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saha:2011:IED

[324] Krishna K. Saha. Interval estimation of the over-dispersion parameter in the analysis of one-way layout of count data. Statistics in Medicine, 30(1):39–51, January 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gleiss:2011:ATS

[325] Andreas Gleiss, Fatima Sanchez-Cabo, Paul Perco, Dan Tong, and Georg Heinze. Adaptive trimmed t-statistics for identifying predominantly high expression in a microarray experiment. Statistics in Medicine, 30(1): 52–61, January 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Consonni:2011:THW

[326] Guido Consonni, Elías Moreno, and Sergio Venturini. Testing Hardy– Weinberg equilibrium: an objective Bayesian analysis. Statistics in Medicine, 30(1):62–74, January 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yavuz:2011:SES

[327] Aysun Çetinyürek Yavuz and Philippe Lambert. Smooth estimation of survival functions and hazard ratios from interval-censored data using Bayesian penalized B-splines. Statistics in Medicine, 30(1):75–90, January 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chan:2011:LLL

[328] Hock Peng Chan and I-Ping Tu. Log-linear, logistic model fitting and local score statistics for cluster detection with covariate adjustments. Statistics in Medicine, 30(1):91–100, January 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ray:2011:BAE

[329] J. Ray, Y. M. Marzouk, and H. N. Najm. A Bayesian approach for estimating bioterror attacks from patient data. Statistics in Medicine, 30(2):101–126, January 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2011:BAC

[330] Pulak Ghosh, Kaushik Ghosh, and Ram C. Tiwari. Bayesian approach to cancer-trend analysis using age-stratified Poisson regression models. Statistics in Medicine, 30(2):127–139, January 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Price:2011:PTE

[331] Malcolm J. Price, Nicky J. Welton, and A. E. Ades. Parameterization of treatment effects for meta-analysis in multi-state Markov models. Statistics in Medicine, 30(2):140–151, January 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2011:ELR

[332] Mai Zhou and Jong-Hyeon Jeong. Empirical likelihood ratio test for median and mean residual lifetime. Statistics in Medicine, 30(2):152– 159, January 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siddique:2011:ULV

[333] Juned Siddique, Catherine M. Crespi, Robert D. Gibbons, and Bonnie L. Green. Using latent variable modeling and multiple imputation to calibrate rater bias in diagnosis assessment. Statistics in Medicine, 30(2): 160–174, January 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deLeon:2011:CBR

[334] A. R. de Leon and B. Wu. Copula-based regression models for a bivariate mixed discrete and continuous outcome. Statistics in Medicine, 30(2): 175–185, January 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Banerjee:2011:ETD

[335] Buddhananda Banerjee and Atanu Biswas. Estimating treatment difference for binary responses in the presence of surrogate endpoints. Statistics in Medicine, 30(2):186–196, January 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hauptmann:2011:FMC

[336] Michael Hauptmann and David B. Richardson. Flexible modeling of the cumulative effects of time-dependent exposures on the hazard. Statistics in Medicine, 30(2):197, January 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sylvestre:2011:AR

[337] Marie-Pierre Sylvestre and Michal Abrahamowicz. Authors’ reply. Statistics in Medicine, 30(2):198–199, January 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kosinski:2011:SSC

[338] Andrzej S. Kosinski, Ying Chen, and Robert H. Lyles. Sample size calculations for evaluating a diagnostic test when the gold standard is missing at random. Statistics in Medicine, 30(2):200, January 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Westgate:2011:ISS

[339] Philip M. Westgate and Thomas M. Braun. Improving small-sample inference in group randomized trials with binary outcomes. Statistics in Medicine, 30(3):201–210, February 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peng:2011:MCM

[340] Yingwei Peng and Jeremy M. G. Taylor. Mixture cure model with random effects for the analysis of a multi-center tonsil cancer study. Statistics in Medicine, 30(3):211–223, February 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hida:2011:TAN

[341] Eisuke Hida and Toshiro Tango. On the three-arm non-inferiority trial including a placebo with a prespecified margin. Statistics in Medicine, 30(3):224–231, February 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tapsoba:2011:JMS

[342] Jean de Dieu Tapsoba, Shen-Ming Lee, and C. Y. Wang. Joint modeling of survival time and longitudinal data with subject-specific changepoints in the covariates. Statistics in Medicine, 30(3):232–249, February 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wong:2011:AMG

[343] May C. M. Wong, K. F. Lam, and Edward C. M. Lo. Analysis of multilevel grouped survival data with time-varying regression coefficients. Statistics in Medicine, 30(3):250–259, February 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schenker:2011:MIM

[344] Nathaniel Schenker, Lori G. Borrud, Vicki L. Burt, Lester R. Curtin, Katherine M. Flegal, Jeffery Hughes, Clifford L. Johnson, Anne C. Looker, and Lisa Mirel. Multiple imputation of missing dual-energy Xray absorptiometry data in the National Health and Nutrition Examination Survey. Statistics in Medicine, 30(3):260–276, February 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Borrell:2011:STI

[345] Luisa N. Borrell and Makram Talih. A symmetrized Theil index measure of health disparities: an example using dental caries in U.S. children and adolescents. Statistics in Medicine, 30(3):277–290, February 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2011:RTT

[346] J. H. Chen and K. F. Cheng. A robust TDT-type association test under informative parental missingness. Statistics in Medicine, 30(3):291–297, February 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2011:CPA

[347] Gang Li. Comments on ‘Planning and analysis of three-arm noninferiority trials with binary endpoints’ by M. Kieser and T. Friede, Statistics in Medicine 2007; 26:253–273. Statistics in Medicine, 30(3): 298–299, February 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [5].

Kieser:2011:PAT

[348] Meinhard Kieser and Tim Friede. Planning and analysis of three-arm non-inferiority trials with binary endpoints. Statistics in Medicine, 30 (3):300, February 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2011:ARG

[349] Xianghua Luo and Chiung-Yu Huang. Analysis of recurrent gap time data using the weighted risk-set method and the modified within-cluster resampling method. Statistics in Medicine, 30(4):301–311, February 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sivaganesan:2011:BSA

[350] S. Sivaganesan, Purushottam W. Laud, and Peter Müller. A Bayesian subgroup analysis with a zero-enriched Polya urn scheme. Statistics in Medicine, 30(4):312–323, February 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Titman:2011:ABD

[351] Andrew C. Titman, Gillian A. Lancaster, Katie Carmichael, and Diane Scutt. Accounting for bias due to a non-ignorable tracing mechanism in a retrospective breast cancer cohort study. Statistics in Medicine, 30 (4):324–334, February 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tchetgen:2011:DRE

[352] Eric J. Tchetgen Tchetgen and Andrea Rotnitzky. Double-robust estimation of an exposure-outcome odds ratio adjusting for confounding in cohort and case-control studies. Statistics in Medicine, 30(4):335–347, February 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2011:BRC

[353] Jenny X. Sun, Samiran Sinha, Suojin Wang, and Tapabrata Maiti. Bias reduction in conditional logistic regression. Statistics in Medicine, 30 (4):348–355, February 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siannis:2011:SAM

[354] Fotios Siannis. Sensitivity analysis for multiple right censoring processes: Investigating mortality in psoriatic arthritis. Statistics in Medicine, 30 (4):356–367, February 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hutton:2011:MBH

[355] J. L. Hutton and E. Stanghellini. Modelling bounded health scores with censored skew-normal distributions. Statistics in Medicine, 30(4):368– 376, February 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2011:MIU

[356] Ian R. White, Patrick Royston, and Angela M. Wood. Multiple imputation using chained equations: Issues and guidance for practice. Statistics in Medicine, 30(4):377–399, February 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2011:EEB

[357] Richard D. Riley. Erratum: an evaluation of bivariate random-effects meta-analysis for the joint synthesis of two correlated outcomes. Statistics in Medicine, 30(4):400, February 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rolka:2011:P

[358] Henry R. Rolka. Preface. Statistics in Medicine, 30(5):401–402, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fricker:2011:SMI

[359] Ronald D. Fricker, Jr. Some methodological issues in biosurveillance. Statistics in Medicine, 30(5):403–415, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rolka:2011:CSM

[360] Henry R. Rolka. Comments on ‘Some methodological issues in biosurveillance’. Statistics in Medicine, 30(5):416–419, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buckeridge:2011:CSM

[361] D. L. Buckeridge. Comments on ‘Some methodological issues in biosurveillance’. Statistics in Medicine, 30(5):420–422, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hanni:2011:CSM

[362] Krista D. Hanni. Comments on ‘Some methodological issues in biosurveillance’. Statistics in Medicine, 30(5):423–425, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burkom:2011:CSM

[363] Howard S. Burkom. Comments on ‘Some methodological issues in biosurveillance’. Statistics in Medicine, 30(5):426–429, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Woodall:2011:CSM

[364] William H. Woodall and Kwok-Leung Tsui. Comments on ‘Some methodological issues in biosurveillance’. Statistics in Medicine, 30(5):

430–433, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fricker:2011:RSM

[365] Ronald D. Fricker, Jr. Rejoinder: Some methodological issues in biosurveillance. Statistics in Medicine, 30(5):434–441, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Izadi:2011:ORS

[366] Masoumeh Izadi and David L. Buckeridge. Optimizing the response to surveillance alerts in automated surveillance systems. Statistics in Medicine, 30(5):442–454, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neill:2011:FBS

[367] Daniel B. Neill. Fast Bayesian scan statistics for multivariate event detection and visualization. Statistics in Medicine, 30(5):455–469, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burkom:2011:IAF

[368] Howard S. Burkom, Liane Ramac-Thomas, Steven Babin, Rekha Holtry, Zaruhi Mnatsakanyan, and Cynthia Yund. An integrated approach for fusion of environmental and human health data for disease surveillance. Statistics in Medicine, 30(5):470–479, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cohen:2011:MDB

[369] Steven A. Cohen, Kenneth K. H. Chui, and Elena N. Naumova. Measuring disease burden in the older population using the slope-intercept method for population log-linear estimation (SIMPLE). Statistics in Medicine, 30(5):480–488, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Szarka:2011:CEA

[370] John L. Szarka III, Linmin Gan, and William H. Woodall. Comparison of the early aberration reporting system (EARS) W2 methods to an adaptive threshold method. Statistics in Medicine, 30(5):489–504, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Molinari:2011:QBH

[371] NoelleAngelique M. Molinari, Kirk M. Wolter, Benjamin Skalland, Robert Montgomery, Meena Khare, Philip J. Smith, Martin L. Barron, Kennon Copeland, Kathleen Santos, and James A. Singleton. Quantifying bias in a health survey: Modeling total survey error in the national immunization survey. Statistics in Medicine, 30(5):505–514, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Caudill:2011:IIR

[372] Samuel P. Caudill. Important issues related to using pooled samples for environmental chemical biomonitoring. Statistics in Medicine, 30(5): 515–521, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murphy:2011:BHM

[373] T. E. Murphy, H. G. Allore, L. Leo-Summers, and B. P. Carlin. Bayesian hierarchical modeling for a non-randomized, longitudinal fall prevention trial with spatially correlated observations. Statistics in Medicine, 30 (5):522–530, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2011:PTC

[374] Zhen Zhao. Power of tests for comparing trend curves with application to national immunization survey (NIS). Statistics in Medicine, 30(5):531– 540, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barker:2011:ETI

[375] Laurie K. Barker, Susan O. Griffin, Seonghye Jeon, Shellie Kolavic Gray, and Brani Vidakovic. Ecological-type inference in matched-pair studies with fixed marginal totals. Statistics in Medicine, 30(5):541–548, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greene:2011:TDL

[376] Sharon K. Greene, Martin Kulldorff, Jie Huang, Richard J. Brand, Kenneth P. Kleinman, John Hsu, and Richard Platt. Timely detection of localized excess influenza activity in northern California across patient care, prescription, and laboratory data. Statistics in Medicine, 30(5): 549–559, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hedt:2011:HIE

[377] Bethany L. Hedt and Marcello Pagano. Health indicators: Eliminating bias from convenience sampling estimators. Statistics in Medicine, 30 (5):560–568, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2011:SSM

[378] Wei Jiang, Sung Won Han, Kwok-Leung Tsui, and William H. Woodall. Spatiotemporal surveillance methods in the presence of spatial correlation. Statistics in Medicine, 30(5):569–583, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lockwood:2011:SAT

[379] J. R. Lockwood, Daniel F. McCaffrey, Claude Messan Setodji, and Marc N. Elliott. Smoothing across time in repeated cross-sectional data. Statistics in Medicine, 30(5):584–594, February 28, 2011. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2011:BRD

[380] Zhiwei Zhang. Book review: Design and analysis of quality of life studies in clinical trials (2nd edn) Diane L. Fairclough, Chapman and Hall/ CRC, Boca Raton, 2010. No. of pages: 424. Price: $89.95. ISBN: 978-1-4200-6117-8. Statistics in Medicine, 30(5):595–596, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2011:ERB

[381] Peng Huang and Robert F. Woolson. Erratum: a rank-based sample size method for multiple outcomes in clinical trials. Statistics in Medicine, 30 (5):597, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Naito:2011:EMS

[382] K. Naito, J. Udagawa, and H. Otani. Erratum: Multidimensional standard curve for the development process of human fetuses. Statistics in Medicine, 27(16):597, February 28, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vandenHout:2011:SRC

[383] Ardo van den Hout, Graciela Muniz-Terrera, and Fiona E. Matthews. Smooth random change point models. Statistics in Medicine, 30(6):599– 610, March 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2011:JML

[384] Pulak Ghosh, Kaushik Ghosh, and Ram C. Tiwari. Joint modeling of longitudinal data and informative dropout time in the presence of multiple changepoints. Statistics in Medicine, 30(6):611–626, March 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mitra:2011:EPS

[385] Robin Mitra and Jerome P. Reiter. Estimating propensity scores with missing covariate data using general location mixture models. Statistics in Medicine, 30(6):627–641, March 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Subramanian:2011:ERM

[386] Jyothi Subramanian and Richard Simon. An evaluation of resampling methods for assessment of survival risk prediction in high-dimensional settings. Statistics in Medicine, 30(6):642–653, March 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lau:2011:PMM

[387] Bryan Lau, Stephen R. Cole, and Stephen J. Gange. Parametric mixture models to evaluate and summarize hazard ratios in the presence of competing risks with time-dependent hazards and delayed entry. Statistics in Medicine, 30(6):654–665, March 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuhnert:2011:MSC

[388] Ronny Kuhnert, Hartmut Hecker, Christina Poethko-Müller, Martin Schlaud, Mechtild Vennemann, Heather J. Whitaker, and C. Paddy Farrington. A modified self-controlled case series method to examine association between multidose vaccinations and death. Statistics in Medicine, 31(16):666–677, March 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bowden:2011:MRA

[389] Jack Bowden and Stijn Vansteelandt. Mendelian randomization analysis of case-control data using structural mean models. Statistics in Medicine, 31(16):678–694, March 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gezmu:2011:SBR

[390] Misrak Gezmu, Victor DeGruttola, Dennis Dixon, Max Essex, Elizabeth Halloran, Joseph Hogan, Anneke Grobler, Soyeon Kim, Jeanne McDermott, Rosemary McKaig, and James D. Neaton. Strengthening biostatistics resources in sub-Saharan Africa: Research collaborations through U.S. partnerships. Statistics in Medicine, 30(7):695–708, March 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altstein:2011:MET

[391] Lily L. Altstein, Gang Li, and Robert M. Elashoff. A method to estimate treatment efficacy among latent subgroups of a randomized clinical trial. Statistics in Medicine, 30(7):709–717, March 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schoenfeld:2011:TRB

[392] David A. Schoenfeld, Natasa Rajicic, Linda H. Ficociello, and Dianne M. Finkelstein. A test for the relationship between a time-varying marker and both recovery and progression with missing data. Statistics in Medicine, 30(7):718–724, March 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shu:2011:CWC

[393] Lianjie Shu, Wei Jiang, and Kwok-Leung Tsui. A comparison of weighted CUSUM procedures that account for monotone changes in population size. Statistics in Medicine, 30(7):725–741, March 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2011:IOR

[394] Stanley Xu, Lijing Zhang, Jennifer C. Nelson, Chan Zeng, John Mullooly, David McClure, and Jason Glanz. Identifying optimal risk windows for self-controlled case series studies of vaccine safety. Statistics in Medicine, 30(7):742–752, March 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2011:ENB

[395] Fengqing (Zoe) Zhang and Don Hong. Elastic net-based framework for imaging mass spectrometry data biomarker selection and classification. Statistics in Medicine, 30(7):753–768, March 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Masca:2011:PIT

[396] Nicholas Masca, Nuala A. Sheehan, and Martin D. Tobin. Pharmacogenetic interactions and their potential effects on genetic analyses of blood pressure. Statistics in Medicine, 30(7):769–783, March 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leoutsakos:2011:ISK

[397] Jeannie-Marie S. Leoutsakos, Karen Bandeen-Roche, Elizabeth Garrett-Mayer, and Peter P. Zandi. Incorporating scientific knowledge into phenotype development: Penalized latent class regression. Statistics in Medicine, 30(7):784–798, March 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Adams:2011:BRS

[398] Niall M. Adams. Book review: Scientific data mining: a practical perspective Chandrika Kamath, SIAM, 2009. No. of pages: xviii + 286. Price:/SIAM Member Price: $49.70. ISBN: 978-0-898716-75-7. Statistics in Medicine, 30(7):799, March 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tournoux-Facon:2011:TPE

[399] Caroline Tournoux-Facon, Yann De Rycke, and Pascale Tubert-Bitter. Targeting population entering phase III trials: a new stratified adaptive phase II design. Statistics in Medicine, 30(8):801–811, April 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuznetsova:2011:BTR

[400] Olga M. Kuznetsova and Yevgen Tymofyeyev. Brick tunnel randomization for unequal allocation to two or more treatment groups. Statistics in Medicine, 30(8):812–824, April 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blanchin:2011:CCR

[401] Myriam Blanchin, Jean-Benoit Hardouin, Tanguy Le Neel, Gildas Kubis, Claire Blanchard, Eric Mirallié, and Véronique Sébille. Comparison of CTT and Rasch-based approaches for the analysis of longitudinal patient reported outcomes. Statistics in Medicine, 30(8):825–838, April 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:2011:NPB

[402] Yi Pan, Michael Haber, and Huiman X. Barnhart. A new permutationbased method for assessing agreement between two observers making replicated binary readings. Statistics in Medicine, 30(8):839–853, April 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carnegie:2011:BCI

[403] Nicole Bohme Carnegie. Bootstrap confidence intervals and bias correction in the estimation of HIV incidence from surveillance data with testing for recent infection. Statistics in Medicine, 30(8):854–865, April 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troxler:2011:ELM

[404] Steven Troxler, Trent Lalonde, and Jeffrey R. Wilson. Exact logistic models for nested binary data. Statistics in Medicine, 30(8):866–876, April 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gustafson:2011:BIG

[405] Paul Gustafson and Igor Burstyn. Bayesian inference of geneenvironment interaction from incomplete data: What happens when information on environment is disjoint from data on gene and disease? Statistics in Medicine, 30(8):877–889, April 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Richardson:2011:ACT

[406] John T. E. Richardson. The analysis of 2 × 2 contingency tables — yet again. Statistics in Medicine, 30(8):890, April 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lydersen:2011:ACT

[407] Stian Lydersen, Petter Laake, and Morten Wang Fagerland. The analysis of 2 × 2 contingency tables — yet again: Author’s reply. Statistics in Medicine, 30(8):891–892, April 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sauerbrei:2011:CPU

[408] Willi Sauerbrei, Patrick Royston, and Martin Schumacher. Comments on ‘Performance of using multiple stepwise algorithms for variable selection’ by Ryan E. Wiegand, Statistics in Medicine 2010; 29:1647–1659. Statistics in Medicine, 30(8):892–894, April 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [166].

Wiegand:2011:RCP

[409] Ryan E. Wiegand. Reply to comments on ‘performance of using multiple stepwise algorithms for variable selection’. Statistics in Medicine, 30(8):

894–896, April 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thygesen:2011:EPM

[410] Lau Caspar Thygesen, Niels Keiding, and Nils Koch-Henriksen. Evaluating a parametric model to correct multiple sclerosis incidence for reporting delay. Statistics in Medicine, 30(8):896–898, April 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2011:BRH

[411] Zudi Lu. Book review: Handbook of spatial statistics Alan E. Gelfand, Peter J. Diggle, Montserrat Fuentes and Peter Guttorp (eds), Chapman and Hall/CRC, Boca Raton, 2010. No. of pages: xii + 607. ISBN: 978-1-4200-7287-7. Statistics in Medicine, 30(8):899–900, April 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moldovan:2011:BRS

[412] Max Moldovan. Book review: Statistical and computational pharmacogenomics Rongling Wu and Min Lin, Chapman and Hall/CRC, Boca Raton, 2008. No. of pages: 368. Price: $83.95. ISBN: 978-1-58488-828-4. Statistics in Medicine, 30(8):900–901, April 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Higgins:2011:SMR

[413] Julian P. T. Higgins, Anne Whitehead, and Mark Simmonds. Sequential methods for random-effects meta-analysis. Statistics in Medicine, 30(9): 903–921, April 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Russell:2011:GMA

[414] D. Russell, Z. S. J. Hoare, Rh. Whitaker, C. J. Whitaker, and I. T. Russell. Generalized method for adaptive randomization in clinical trials. Statistics in Medicine, 30(9):922–934, April 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hutmacher:2011:ETR

[415] Matthew M. Hutmacher, Jonathan L. French, Sriram Krishnaswami, and Sujatha Menon. Estimating transformations for repeated measures modeling of continuous bounded outcome data. Statistics in Medicine, 30(9): 935–949, April 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OMalley:2011:LAL

[416] A. James O’Malley and Nicholas A. Christakis. Longitudinal analysis of large social networks: Estimating the effect of health traits on changes in friendship ties. Statistics in Medicine, 30(9):950–964, April 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brumback:2011:ACN

[417] Babette A. Brumback and Zhulin He. Adjusting for confounding by neighborhood using complex survey data. Statistics in Medicine, 30(9): 965–972, April 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cresswell:2011:AJS

[418] Lynne Cresswell and Vern Farewell. Assessment of joint symmetry in arthritis. Statistics in Medicine, 30(9):973–983, April 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mercer:2011:ERN

[419] G. N. Mercer, K. Glass, and N. G. Becker. Effective reproduction numbers are commonly overestimated early in a disease outbreak. Statistics in Medicine, 30(9):984–994, April 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiang:2011:MCM

[420] Liming Xiang, Xiangmei Ma, and Kelvin K. W. Yau. Mixture cure model with random effects for clustered interval-censored survival data. Statistics in Medicine, 30(9):995–1006, April 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2011:BBU

[421] Wen-Chung Lee. Bounding the bias of unmeasured factors with confounding and effect-modifying potentials. Statistics in Medicine, 30(9): 1007–1017, April 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tseng:2011:ACE

[422] Chi hong Tseng and Weng Kee Wong. Analysis of a composite endpoint with longitudinal and time-to-event data. Statistics in Medicine, 30(9): 1018–1027, April 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finkelman:2011:ZIM

[423] Matthew D. Finkelman, Jennifer Greif Green, Michael J. Gruber, and Alan M. Zaslavsky. A zero- and K-inflated mixture model for health questionnaire data. Statistics in Medicine, 30(9):1028–1043, April 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guedj:2011:DEO

[424] Jeremie Guedj, Caroline Bazzoli, Avidan U. Neumann, and France Mentré. Design evaluation and optimization for models of hepatitis C viral dynamics. Statistics in Medicine, 30(10):1045–1056, May 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moraga:2011:DSD

[425] Paula Moraga and Francisco Montes. Detection of spatial disease clusters with LISA functions. Statistics in Medicine, 30(10):1057–1071, May 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Akacha:2011:MRC

[426] Mouna Akacha and Jane L. Hutton. Modelling the rate of change in a longitudinal study with missing data, adjusting for contact attempts. Statistics in Medicine, 30(10):1072–1089, May 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gail:2011:PEB

[427] Mitchell H. Gail. Personalized estimates of breast cancer risk in clinical practice and public health. Statistics in Medicine, 30(10):1090–1104, May 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Uno:2011:CSE

[428] Hajime Uno, Tianxi Cai, Michael J. Pencina, Ralph B. D’Agostino, and L. J. Wei. On the c-statistics for evaluating overall adequacy of risk prediction procedures with censored survival data. Statistics in Medicine, 30(10):1105–1117, May 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paul:2011:PAN

[429] M. Paul and L. Held. Predictive assessment of a non-linear random effects model for multivariate time series of infectious disease counts. Statistics in Medicine, 30(10):1118–1136, May 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2011:FMI

[430] Yulei He, Recai Yucel, and Trivellore E. Raghunathan. A functional multiple imputation approach to incomplete longitudinal data. Statistics in Medicine, 30(10):1137–1156, May 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Asimit:2011:RMS

[431] Jennifer L. Asimit, Irene L. Andrulis, and Shelley B. Bull. Regression models, scan statistics and reappearance probabilities to detect regions of association between gene expression and copy number. Statistics in Medicine, 30(10):1157–1178, May 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheng:2011:BRB

[432] Dunlei Cheng. Book review: Bayesian ideas and data analysis: an introduction for scientists and statisticians R. Christensen, W. Johnson, A. Branscum and T. E. Hanson, CRC Press, Boca Raton, 2010. No. of pages: ix + 470 pp. Price: $69.95/£44.77. ISBN: 978-1-4398-0354-7. Statistics in Medicine, 30(10):1179–1180, May 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Woods:2011:BRD

[433] David Woods. Book reviews: Design and analysis of experiments with SAS. John Lawson, Chapman and Hall/CRC, Boca Raton, 2010. No. of pages: 582. Price: $99.95. ISBN 10: 1-4200-6060-0, ISBN 13: 978-1-4200-6060-7. Statistics in Medicine, 30(10):1180–1181, May 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chootrakool:2011:MAS

[434] Hathaikan Chootrakool, Jian Qing Shi, and Rongxian Yue. Meta-analysis and sensitivity analysis for multi-arm trials with selection bias. Statistics in Medicine, 30(11):1183–1198, May 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Emerson:2011:EBA

[435] Sarah C. Emerson, Kyle D. Rudser, and Scott S. Emerson. Exploring the benefits of adaptive sequential designs in time-to-event endpoint settings. Statistics in Medicine, 30(11):1199–1217, May 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yuan:2011:BRA

[436] Ying Yuan, Xuelin Huang, and Suyu Liu. A Bayesian response-adaptive covariate-balanced randomization design with application to a leukemia clinical trial. Statistics in Medicine, 30(11):1218–1229, May 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2011:TNI

[437] Kung-Jong Lui and Kuang-Chao Chang. Test non-inferiority (and equivalence) based on the odds ratio under a simple crossover trial. Statistics in Medicine, 30(11):1230–1242, May 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2011:CAN

[438] Honghua Jiang, James Symanowski, Yongming Qu, Xiao Ni, and Yanping Wang. Covariate-adjusted non-parametric survival curve estimation. Statistics in Medicine, 30(11):1243–1253, May 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2011:ERT

[439] Luqiang Wang, Kevin John Keen, and Burt Holland. Estimation of reliability in a three-factor model. Statistics in Medicine, 30(11):1254– 1265, May 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2011:TDR

[440] Jialiang Li and Shuangge Ma. Time-dependent ROC analysis under diverse censoring patterns. Statistics in Medicine, 30(11):1266–1277, May 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2011:MRV

[441] Ming Wang and Qi Long. Modified robust variance estimator for generalized estimating equations with improved small-sample performance. Statistics in Medicine, 30(11):1278–1291, May 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2011:CPV

[442] Peter C. Austin. Comparing paired vs non-paired statistical methods of analyses when making inferences about absolute risk reductions in propensity-score matched samples. Statistics in Medicine, 30(11):1292– 1301, May 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schenker:2011:ESE

[443] Nathaniel Schenker, Van L. Parsons, Kimberly A. Lochner, Gloria Wheatcroft, and Elsie R. Pamuk. Estimating standard errors for life expectancies based on complex survey data with mortality follow-up: a case study using the national health interview survey linked mortality files. Statistics in Medicine, 30(11):1302–1311, May 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burgess:2011:BCE

[444] Stephen Burgess and Simon G. Thompson. Bias in causal estimates from Mendelian randomization studies with weak instruments. Statistics in Medicine, 30(11):1312–1323, May 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rogers:2011:BRO

[445] Wendy Rogers. Book reviews: The Oxford textbook of clinical research ethics Ezekiel Emanuel, Christine Grady, Robert Crouch, Reidar Lie, Franklin Miller, David Wendler, Oxford University Press, 2008. No. of pages: 848. Price: $150.00. ISBN 13: 978-0-19-516865-5, ISBN 10: 0-19-516865-8 Hardback. Statistics in Medicine, 30(11):1324–1325, May 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ormerod:2011:BRM

[446] John T. Ormerod. Book review: Mixed effects models for complex data Lang Wu, Chapman and Hall/CRC, Boca Raton, 2010. No. of pages: xx + 419. Price: $89.95. ISBN: 978-1-4200-7402-4. Statistics in Medicine, 30(11):1326–1327, May 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bartlett:2011:BRM

[447] Jonathan Bartlett. Book review: Missing data in longitudinal studies: Strategies for Bayesian modeling and sensitivity analysis M. J. Daniels and J. W. Hogan, Chapman and Hall/CRC, Boca Raton, 2008. No. of pages: 303. Price: $83.95. ISBN: 978-1-58488-609-9. Statistics in Medicine, 30(11):1327–1328, May 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cuffe:2011:IHC

[448] Robert L. Cuffe. The inclusion of historical control data may reduce the power of a confirmatory study. Statistics in Medicine, 30(12):1329–1338, May 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2011:MIB

[449] Lyrica Xiaohong Liu, Susan Murray, and Alex Tsodikov. Multiple imputation based on restricted mean model for censored data. Statistics in Medicine, 30(12):1339–1350, May 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mahboubi:2011:FME

[450] Amel Mahboubi, Michal Abrahamowicz, Roch Giorgi, Christine Binquet, Claire Bonithon-Kopp, and Catherine Quantin. Flexible modeling of the effects of continuous prognostic factors in relative survival. Statistics in Medicine, 30(12):1351–1365, May 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rizopoulos:2011:BSM

[451] Dimitris Rizopoulos and Pulak Ghosh. A Bayesian semiparametric multivariate joint model for multiple longitudinal outcomes and a time-to-event. Statistics in Medicine, 30(12):1366–1380, May 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wakefield:2011:BCE

[452] Jon Wakefield, Sebastien Haneuse, Adrian Dobra, and Elizabeth Teeple. Bayes computation for ecological inference. Statistics in Medicine, 30 (12):1381–1396, May 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2011:ISCa

[453] Richard J. Cook and Pierre-Jérôme Bergeron. Information in the sample covariate distribution in prevalent cohorts. Statistics in Medicine, 30(12): 1397–1409, May 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demler:2011:EIA

[454] Olga V. Demler, Michael J. Pencina, and Ralph B. D’Agostino, Sr. Equivalence of improvement in area under ROC curve and linear discriminant analysis coefficient under assumption of normality. Statistics in Medicine, 30(26):1410–1418, May 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tervonen:2011:SMM

[455] Tommi Tervonen, Gert van Valkenhoef, Erik Buskens, Hans L. Hillege, and Douwe Postmus. A stochastic multicriteria model for evidence-based decision making in drug benefit-risk analysis. Statistics in Medicine, 30 (26):1419–1428, May 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2011:RAL

[456] Liang Zhu, Jianguo Sun, Xingwei Tong, and Stanley Pounds. Regression analysis of longitudinal data with informative observation times and application to medical cost data. Statistics in Medicine, 30(26):1429–1440, May 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Behseta:2011:CTP

[457] Sam Behseta and Shojaeddin Chenouri. Comparison of two populations of curves with an application in neuronal data analysis. Statistics in Medicine, 30(26):1441–1454, May 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2011:TPM

[458] Z. Zhang, O. Y. Addo, J. H. Himes, M. L. Hediger, P. S. Albert, A. L. Gollenberg, P. A. Lee, and G. M. Buck Louis. A two-part model for reference curve estimation subject to a limit of detection. Statistics in Medicine, 30(26):1455–1465, May 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heinze:2011:CBR

[459] Georg Heinze. Comment on ‘Bias reduction in conditional logistic regression’ by J. X. Sun, S. Sinha, S. Wang and T. Maiti, Statistics in Medicine 2010; DOI: 10.1002/sim.4105. Statistics in Medicine, 30(26): 1466–1467, May 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boshnakov:2011:BRI

[460] Georgi N. Boshnakov. Book reviews: Introduction to time series modeling Genshiro, Kitagawa, Chapman and Hall/CRC Monographs on Statistics and Applied Probability, 2010. xxiii. No. of pages: 296. Hardback, Price: $79.95. ISBN: 978-1-58488-921-2, ISBN 10: 1-58488-921-7. Statistics in Medicine, 30(26):1468–1469, May 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grouin:2011:SIF

[461] Jean-Marie Grouin and Michael Campbell. Sixth International French Society of Statistics meeting on statistical methods in biopharmacy: Recent advances and trends in statistics as applied to clinical trials. Statistics in Medicine, 30(13):1471–1472, June 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dmitrienko:2011:MMT

[462] Alex Dmitrienko and Ajit C. Tamhane. Mixtures of multiple testing procedures for gatekeeping applications in clinical trials. Statistics in Medicine, 30(13):1473–1488, June 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bretz:2011:TPC

[463] Frank Bretz, Willi Maurer, and Gerhard Hommel. Test and power considerations for multiple endpoint analyses using sequentially rejective graphical procedures. Statistics in Medicine, 30(13):1489–1501, June 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Michiels:2011:MTT

[464] Stefan Michiels, Richard F. Potthoff, and Stephen L. George. Multiple testing of treatment-effect-modifying biomarkers in a randomized clinical trial with a survival endpoint. Statistics in Medicine, 30(13):1502–1518, June 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hung:2011:FDC

[465] H. M. James Hung, Sue-Jane Wang, and Robert O’Neill. Flexible design clinical trial methodology in regulatory applications. Statistics in Medicine, 30(13):1519–1527, June 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friede:2011:DSP

[466] T. Friede, N. Parsons, N. Stallard, S. Todd, E. Valdes Marquez, J. Chataway, and R. Nicholas. Designing a seamless phase II/III clinical trial using early outcomes for treatment selection: an application in multiple sclerosis. Statistics in Medicine, 30(13):1528–1540, June 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heritier:2011:ACT

[467] Stephane Heritier, Serigne N. Lô, and Caroline C. Morgan. An adaptive confirmatory trial with interim treatment selection: Practical experiences and unbalanced randomization. Statistics in Medicine, 30(13):1541–1554, June 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tournoux-Facon:2011:HNS

[468] Caroline Tournoux-Facon, Yann De Rycke, and Pascale Tubert-Bitter. How a new stratified adaptive phase II design could improve targeting population. Statistics in Medicine, 30(13):1555–1562, June 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Daimon:2011:PMA

[469] T. Daimon, S. Zohar, and J. O’Quigley. Posterior maximization and averaging for Bayesian working model choice in the continual reassessment method. Statistics in Medicine, 30(13):1563–1573, June 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seegers:2011:DFD

[470] V. Seegers, S. Chevret, and M. Resche-Rigon. Dose-finding design driven by efficacy in onco-hematology phase I/II trials. Statistics in Medicine, 30(13):1574–1583, June 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mauguen:2011:DFA

[471] A. Mauguen, M. C. Le Deley, and S. Zohar. Dose-finding approach for dose escalation with overdose control considering incomplete observations. Statistics in Medicine, 30(13):1584–1594, June 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marti:2011:MIA

[472] Helena Marti and Michel Chavance. Multiple imputation analysis of casecohort studies. Statistics in Medicine, 30(13):1595–1607, June 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Katsahian:2011:ETC

[473] Sandrine Katsahian and Christian Boudreau. Estimating and testing for center effects in competing risks. Statistics in Medicine, 30(13):1608– 1617, June 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neuenschwander:2011:PCP

[474] Beat Neuenschwander, Nicolas Rouyrre, Norbert Hollaender, Emmanuel Zuber, and Michael Branson. A proof of concept phase II non-inferiority criterion. Statistics in Medicine, 30(13):1618–1627, June 15, 2011. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seldrup:2011:RAD

[475] Jørgen Seldrup. Regulatory advice and drug development — a case study in negotiating with regulators. Statistics in Medicine, 30(13):1628–1635, June 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Graf:2011:MIT

[476] Alexandra C. Graf and Peter Bauer. Maximum inflation of the type 1 error rate when sample size and allocation rate are adapted in a preplanned interim look. Statistics in Medicine, 30(14):1637–1647, June 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brutti:2011:EST

[477] P. Brutti, S. Gubbiotti, and V. Sambucini. An extension of the single threshold design for monitoring efficacy and safety in phase II clinical trials. Statistics in Medicine, 30(14):1648–1664, June 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peter:2011:RFA

[478] William Peter, Amir H. Najmi, and Howard S. Burkom. Reducing false alarms in syndromic surveillance. Statistics in Medicine, 30(14):1665– 1677, June 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeSantis:2011:HMM

[479] Stacia M. DeSantis and Dipankar Bandyopadhyay. Hidden Markov models for zero-inflated Poisson counts with an application to substance use. Statistics in Medicine, 30(14):1678–1694, June 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roca-Pardinas:2011:ANA

[480] Javier Roca-Pardiñas, Carmen Cadarso-Suárez, Jose L. Pardo-Vazquez, Victor Leboran, Geert Molenberghs, Christel Faes, and Carlos Acuña. Assessing neural activity related to decision-making through flexible odds ratio curves and their derivatives. Statistics in Medicine, 30(14):1695– 1711, June 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2011:CMC

[481] Binbing Yu, Chuan Zhou, and Stefania Bandinelli. Combining multiple continuous tests for the diagnosis of kidney impairment in the absence of a gold standard. Statistics in Medicine, 30(14):1712–1721, June 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McIntyre:2011:RAD

[482] Julie McIntyre and Leonard A. Stefanski. Regression-assisted deconvolution. Statistics in Medicine, 30(14):1722–1734, June 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chervoneva:2011:CEL

[483] Inna Chervoneva and Mark Vishnyakov. Constrained S-estimators for linear mixed effects models with covariance components. Statistics in Medicine, 30(14):1735–1750, June 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2011:ADS

[484] Yuanjia Wang, Huaihou Chen, Theresa Schwartz, Naihua Duan, Angela Parcesepe, and Roberto Lewis-Fernández. Assessment of a disease screener by hierarchical all-subset selection using area under the receiver operating characteristic curves. Statistics in Medicine, 30(14):1751–1760, June 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schur:2011:MGD

[485] Nadine Schur, L. Gosoniu, G. Raso, J. Utzinger, and P. Vounatsou. Modelling the geographical distribution of co-infection risk from single-disease surveys. Statistics in Medicine, 30(14):1761–1776, June 30, 2011. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thall:2011:DRE

[486] Peter F. Thall, Diane D. Liu, Su G. Berrak, and Johannes E. Wolff. Defining and ranking effects of individual agents based on survival times of cancer patients treated with combination chemotherapies. Statistics in Medicine, 30(15):1777–1794, July 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2011:ANT

[487] Pulak Ghosh, Farouk Nathoo, Mithat Gönen, and Ram C. Tiwari. Assessing noninferiority in a three-arm trial using the Bayesian approach. Statistics in Medicine, 30(15):1795–1808, July 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2011:TSI

[488] Bing Cai, Dylan S. Small, and Thomas R. Ten Have. Two-stage instrumental variable methods for estimating the causal odds ratio: Analysis of bias. Statistics in Medicine, 30(15):1809–1824, July 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2011:JMA

[489] Zhen Chen, Bo Zhang, and Paul S. Albert. A joint modeling approach to data with informative cluster size: Robustness to the cluster size model. Statistics in Medicine, 30(15):1825–1836, July 10, 2011. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2011:SRM

[490] Inyoung Kim, Hae-Kwan Cheong, and Ho Kim. Semiparametric regression models for detecting effect modification in matched case-crossover studies. Statistics in Medicine, 30(15):1837–1851, July 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alonzo:2011:BEA

[491] Todd A. Alonzo, John T. Brinton, Brandy M. Ringham, and Deborah H. Glueck. Bias in estimating accuracy of a binary screening test with differential disease verification. Statistics in Medicine, 30(15):1852–1864, July 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pennell:2011:CDC

[492] Michael L. Pennell, Erinn M. Hade, David M. Murray, and Dale A. Rhoda. Cutoff designs for community-based intervention studies. Statistics in Medicine, 30(15):1865–1882, July 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2011:FEC

[493] Yuanjia Wang. Flexible estimation of covariance function by penalized spline with application to longitudinal family data. Statistics in Medicine, 30(21):1883–1897, July 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Faye:2011:FGW

[494] Laura L. Faye, Lei Sun, Apostolos Dimitromanolakis, and Shelley B. Bull. A flexible genome-wide bootstrap method that accounts for ranking and threshold-selection bias in GWAS interpretation and replication study design. Statistics in Medicine, 30(21):1898–1912, July 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2011:FRS

[495] Yong Chen and Sheng Luo. A few remarks on ‘Statistical distribution of the difference of two proportions’ by Nadarajah and Kotz, Statistics in Medicine 2007; 26 (18):3518–3523. Statistics in Medicine, 30(21): 1913–1915, July 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nadarajah:2011:RFR

[496] Saralees Nadarajah. Response to ‘A few remarks on Statistical distribution of the difference of two proportions’ by Chen and Luo. Statistics in Medicine, 30(21):1916, July 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yoon:2011:AMT

[497] Frank B. Yoon, Garrett M. Fitzmaurice, Stuart R. Lipsitz, Nicholas J. Horton, Nan M. Laird, and Sharon-Lise T. Normand. Alternative methods for testing treatment effects on the basis of multiple outcomes: Simulation and case study. Statistics in Medicine, 30(16):1917–1932, July 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2011:PHR

[498] Xu Zhang, Mei-Jie Zhang, and Jason Fine. A proportional hazards regression model for the subdistribution with right-censored and lefttruncated competing risks data. Statistics in Medicine, 30(16):1933– 1951, July 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2011:BPP

[499] John Whitehead, Helene Thygesen, and Anne Whitehead. Bayesian procedures for phase I/II clinical trials investigating the safety and efficacy of drug combinations. Statistics in Medicine, 30(16):1952–1970, July 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OMalley:2011:ECO

[500] A. James O’Malley, R. G. Frank, and S.-L. T. Normand. Estimating cost-offsets of new medications: Use of new antipsychotics and mental health costs for schizophrenia. Statistics in Medicine, 30(16):1971–1988, July 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finkelman:2011:SCH

[501] Matthew D. Finkelman, Yulei He, Wonsuk Kim, and Albert M. Lai. Stochastic curtailment of health questionnaires: a method to reduce respondent burden. Statistics in Medicine, 30(16):1989–2004, July 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2011:MMCa

[502] Chunling Liu, Aiyi Liu, and Susan Halabi. A min-max combination of biomarkers to improve diagnostic accuracy. Statistics in Medicine, 30 (16):2005–2014, July 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leng:2011:PVS

[503] Chenlei Leng, Hua Liang, and Neil Martinson. Parametric variable selection in generalized partially linear models with an application to assess condom use by HIV-infected patients. Statistics in Medicine, 30(16): 2015–2027, July 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2011:SAJ

[504] Tae Young Yang. A SATS algorithm for jointly identifying multiple differentially expressed gene sets. Statistics in Medicine, 30(16):2028– 2039, July 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2011:ASA

[505] Palash Ghosh and Anup Dewanji. Analysis of spontaneous adverse drug reaction (ADR) reports using supplementary information. Statistics in Medicine, 30(16):2040–2055, July 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robinson:2011:BRL

[506] Andrew Robinson. Book review: Longitudinal data analysis Garrett Fitzmaurice, Marie Davidian, Geert Verbeke and Geert Molenberghs, CRC Press, Boca Raton, 2008. No. of pages: xiv + 618. Price: $99.95. ISBN: 978-1-58488-658-7. Statistics in Medicine, 30(16):2056, July 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Iasonos:2011:CRR

[507] Alexia Iasonos and John O’Quigley. Continual reassessment and related designs in dose-finding studies. Statistics in Medicine, 30(17):2057–2061, July 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OQuigley:2011:EMB

[508] John O’Quigley and Mark Conaway. Extended model-based designs for more complex dose-finding studies. Statistics in Medicine, 30(17):2062– 2069, July 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meter:2011:POM

[509] Emily M. Van Meter, Elizabeth Garrett-Mayer, and Dipankar Bandyopadhyay. Proportional odds model for dose-finding clinical trial designs with ordinal toxicity grading. Statistics in Medicine, 30(17):2070–2080, July 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2011:CPV

[510] Shing M. Lee and Ying Kuen Cheung. Calibration of prior variance in the Bayesian continual reassessment method. Statistics in Medicine, 30(17): 2081–2089, July 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morita:2011:ACR

[511] Satoshi Morita. Application of the continual reassessment method to a phase I dose-finding trial in Japanese patients: East meets West. Statistics in Medicine, 30(17):2090–2097, July 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yuan:2011:BHD

[512] Ying Yuan and Guosheng Yin. Bayesian hybrid dose-finding design in phase I oncology clinical trials. Statistics in Medicine, 30(17):2098–2108, July 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zohar:2011:AMA

[513] Sarah Zohar, Sandrine Katsahian, and John O’Quigley. An approach to meta-analysis of dose-finding studies. Statistics in Medicine, 30(17): 2109–2116, July 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Iasonos:2011:EDT

[514] Alexia Iasonos and Irina Ostrovnaya. Estimating the dose-toxicity curve in completed phase I studies. Statistics in Medicine, 30(17):2117–2129, July 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Polley:2011:PMT

[515] Mei-Yin C. Polley. Practical modifications to the time-to-event continual reassessment method for phase I cancer trials with fast patient accrual and late-onset toxicities. Statistics in Medicine, 30(17):2130–2143, July 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ha:2011:FMS

[516] Il Do Ha, Richard Sylvester, Catherine Legrand, and Gilbert MacKenzie. Frailty modelling for survival data from multi-centre clinical trials. Statistics in Medicine, 30(17):2144–2159, July 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mitchell:2011:DTS

[517] C. E. Mitchell, M. G. Hudgens, C. C. King, S. Cu-Uvin, Y. Lo, A. Rompalo, J. Sobel, and J. S. Smith. Discrete-time semi-Markov modeling of human papillomavirus persistence. Statistics in Medicine, 30(17): 2160–2170, July 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsou:2011:CAE

[518] Hsiao-Hui Tsou, Tuan-Ying Chien, Jen pei Liu, and Chin-Fu Hsiao. A consistency approach to evaluation of bridging studies and multi-regional trials. Statistics in Medicine, 30(17):2171–2186, July 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gray:2011:HSM

[519] Simone C. Gray, Alan E. Gelfand, and Marie Lynn Miranda. Hierarchical spatial modeling of uncertainty in air pollution and birth weight study. Statistics in Medicine, 30(17):2187–2198, July 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heller:2011:BRP

[520] Gillian Z. Heller. Book review: Proportional hazards regression John O’Quigley, Springer, 2008. No. of pages: xvii + 542. Price: 56.99 GBP. ISBN: 978-0-387-68639-4. Statistics in Medicine, 30(17):2199, July 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steen:2011:PGW

[521] K. Van Steen. Perspectives on genome-wide multi-stage family-based association studies. Statistics in Medicine, 30(18):2201–2221, August 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Manjourides:2011:IPC

[522] Justin Manjourides and Marcello Pagano. Improving the power of chronic disease surveillance by incorporating residential history. Statistics in Medicine, 30(18):2222–2233, August 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lunn:2011:FDM

[523] D. J. Lunn, C. Wei, and R. Hovorka. Fitting dynamic models with forcing functions: Application to continuous glucose monitoring in insulin therapy. Statistics in Medicine, 30(18):2234–2250, August 15, 2011. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2011:EMT

[524] Xiao Song and Xiao-Hua Zhou. Evaluating markers for treatment selection based on survival time. Statistics in Medicine, 30(18):2251–2264, August 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2011:MRM

[525] Liuquan Sun, Xian Zhou, and Shaojun Guo. Marginal regression models with time-varying coefficients for recurrent event data. Statistics in Medicine, 30(18):2265–2277, August 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2011:RAC

[526] Ying Guo and Roderick J. Little. Regression analysis with covariates that have heteroscedastic measurement error. Statistics in Medicine, 30 (18):2278–2294, August 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2011:SSP

[527] Liddy M. Chen, Joseph G. Ibrahim, and Haitao Chu. Sample size and power determination in joint modeling of longitudinal and survival data.

Statistics in Medicine, 30(18):2295–2309, August 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Entink:2011:MMJ

[528] Rinke H. Klein Entink, Jean-Paul Fox, and Ardo van den Hout. A mixture model for the joint analysis of latent developmental trajectories and survival. Statistics in Medicine, 30(18):2310–2325, August 15, 2011. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buu:2011:NVS

[529] Anne Buu, Norman J. Johnson, Runze Li, and Xianming Tan. New variable selection methods for zero-inflated count data with applications to the substance abuse field. Statistics in Medicine, 30(18):2326–2340, August 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Snapinn:2011:CMT

[530] Steven Snapinn and Qi Jiang. On the clinical meaningfulness of a treatment’s effect on a time-to-event variable. Statistics in Medicine, 30 (19):2341–2348, August 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2011:CMR

[531] Yan Ma, Jason Roy, and Bess Marcus. Causal models for randomized trials with two active treatments and continuous compliance. Statistics in Medicine, 30(19):2349–2362, August 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Soares:2011:MEE

[532] Marta O. Soares, Laura Bojke, Jo Dumville, Cynthia Iglesias, Nicky Cullum, and Karl Claxton. Methods to elicit experts’ beliefs over uncertain quantities: application to a cost effectiveness transition model of negative pressure wound therapy for severe pressure ulceration. Statistics in Medicine, 30(19):2363–2380, August 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2011:SIC

[533] H. Zhao, Y. Cheng, and H. Bang. Some insight on censored cost estimators. Statistics in Medicine, 30(19):2381–2388, August 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moore:2011:REC

[534] Kelly L. Moore, Romain Neugebauer, Thamban Valappil, and Mark J. van der Laan. Robust extraction of covariate information to improve estimation efficiency in randomized trials. Statistics in Medicine, 30 (19):2389–2408, August 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2011:URM

[535] Patrick Royston and Mahesh K. B. Parmar. The use of restricted mean survival time to estimate the treatment effect in randomized clinical trials when the proportional hazards assumption is in doubt. Statistics in Medicine, 30(19):2409–2421, August 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2011:SCC

[536] Zhenguo Wu, Ping He, and Zhi Geng. Sufficient conditions for concluding surrogacy based on observed data. Statistics in Medicine, 30(19):2422– 2434, August 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2011:MSS

[537] Li Zhu, Fushing Hsieh, Juan Li, and Eric Chi. Modeling subject-specific phase-dependent effects and variations in longitudinal responses via a geometric Brownian motion process. Statistics in Medicine, 30(19):2435– 2450, August 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2011:SSP

[538] Dingfeng Jiang and Jacob J. Oleson. Simulation study of power and sample size for repeated measures with multinomial outcomes: an application to sound direction identification experiments (SDIE). Statistics in Medicine, 30(19):2451–2466, August 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Molanes-Lopez:2011:IYI

[539] Elisa M. Molanes-López and Emilio Letón. Inference of the Youden index and associated threshold using empirical likelihood for quantiles. Statistics in Medicine, 30(19):2467–2480, August 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2011:MMA

[540] Dan Jackson, Richard Riley, and Ian R. White. Multivariate metaanalysis: Potential and promise. Statistics in Medicine, 30(20):2481– 2498, September 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hedges:2011:CMM

[541] Larry V. Hedges. Comment on ‘multivariate meta-analysis: Potential and promise’. Statistics in Medicine, 30(20):2499, September 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:2011:MMA

[542] D. R. Cox. Multivariate meta-analysis: a comment. Statistics in Medicine, 30(20):2500–2501, September 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bland:2011:CMM

[543] J. Martin Bland. Comments on ‘Multivariate meta-analysis: Potential and promise’ by Jackson et al., Statistics in Medicine. Statistics in Medicine, 30(20):2502–2503, September 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gasparrini:2011:MMA

[544] A. Gasparrini and B. Armstrong. Multivariate meta-analysis: a method to summarize non-linear associations. Statistics in Medicine, 30(20): 2504–2506, September 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harbord:2011:CMM

[545] Roger M. Harbord. Commentary on ‘multivariate meta-analysis: potential and promise’. Statistics in Medicine, 30(20):2507–2508, September 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2011:RCM

[546] Dan Jackson, Ian R. White, and Richard D. Riley. Rejoinder to commentaries on ‘multivariate meta-analysis: Potential and promise’. Statistics in Medicine, 30(20):2509–2510, September 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dakin:2011:MTC

[547] Helen A. Dakin, Nicky J. Welton, A. E. Ades, Sarah Collins, Michelle Orme, and Steven Kelly. Mixed treatment comparison of repeated measurements of a continuous endpoint: an example using topical treatments for primary open-angle glaucoma and ocular hypertension. Statistics in Medicine, 30(20):2511–2535, September 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cozza:2011:DCD

[548] Valentina Cozza, Maria Rosario Guarracino, Lucia Maddalena, and Adone Baroni. Dynamic clustering detection through multi-valued descriptors of dermoscopic images. Statistics in Medicine, 30(20):2536– 2550, September 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

May:2011:MLE

[549] Ryan C. May, Joseph G. Ibrahim, and Haitao Chu. Maximum likelihood estimation in generalized linear models with multiple covariates subject to detection limits. Statistics in Medicine, 30(20):2551–2561, September 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2011:FGL

[550] Hui Zhang, Naiji Lu, Changyong Feng, Sally W. Thurston, Yinglin Xia, Liang Zhu, and Xin M. Tu. On fitting generalized linear mixed-effects models for binary responses using different statistical packages. Statistics in Medicine, 30(20):2562–2572, September 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kennes:2011:ISB

[551] Lieven Nils Kennes, Erhard Cramer, Ralf-Dieter Hilgers, and Nicole Heussen. The impact of selection bias on test decisions in randomized clinical trials. Statistics in Medicine, 30(21):2573–2581, September 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dubois:2011:MBA

[552] Anne Dubois, Marc Lavielle, Sandro Gsteiger, Etienne Pigeolet, and France Mentré. Model-based analyses of bioequivalence crossover trials using the stochastic approximation expectation maximisation algorithm. Statistics in Medicine, 30(21):2582–2600, September 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lipkovich:2011:SIB

[553] Ilya Lipkovich, Alex Dmitrienko, Jonathan Denne, and Gregory Enas. Subgroup identification based on differential effect search — a recursive partitioning method for establishing response to treatment in patient subpopulations. Statistics in Medicine, 30(21):2601–2621, September 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:2011:CSM

[554] Huaming Tan, David Gruben, Jonathan French, and Neal Thomas. A case study of model-based Bayesian dose response estimation. Statistics in Medicine, 30(21):2622–2633, September 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

An:2011:EMF

[555] Xinming An and Peter M. Bentler. Extended mixture factor analysis model with covariates for mixed binary and continuous responses. Statistics in Medicine, 30(21):2634–2647, September 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ladouceur:2011:MCD

[556] Martin Ladouceur, Elham Rahme, Patrick Bélisle, Allison N. Scott, Kevin Schwartzman, and Lawrence Joseph. Modeling continuous diagnostic test data using approximate Dirichlet process distributions. Statistics in Medicine, 30(21):2648–2662, September 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harel:2011:LEJ

[557] Ofer Harel and Hakan Demirtas. Letter to the Editor: Joint modeling of missing data due to non-participation and death in longitudinal aging studies by K. B. Rajan and S. E. Leurgans, Statistics in Medicine 2010; 29:2260–2268. Statistics in Medicine, 30(21):2663–2665, September 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [228].

Rajan:2011:AR

[558] Kumar B. Rajan and Sue E. Leurgans. Authors’ reply. Statistics in Medicine, 30(21):2666–2668, September 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pennell:2011:C

[559] Michael Pennell. Correction. Statistics in Medicine, 30(21):2669, September 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aiello:2011:ACI

[560] Fabio Aiello, Massimo Attanasio, and Fabio Tinè. Assessing covariate imbalance in meta-analysis studies. Statistics in Medicine, 30(22):2671– 2682, September 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhangsheng:2011:JMR

[561] Yu Zhangsheng and Lei Liu. A joint model of recurrent events and a terminal event with a nonparametric covariate function. Statistics in Medicine, 30(22):2683–2695, September 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gurka:2011:ABM

[562] Matthew J. Gurka, Lloyd J. Edwards, and Keith E. Muller. Avoiding bias in mixed model inference for fixed effects. Statistics in Medicine, 30(22): 2696–2707, September 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Izu:2011:BMF

[563] Alane Izu, Ted Cohen, Carole Mitnick, Megan Murray, and Victor De Gruttola. Bayesian methods for fitting mixture models that characterize branching tree processes: an application to development of resistant TB strains. Statistics in Medicine, 30(22):2708–2720, September 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neelon:2011:BGM

[564] Brian Neelon, Geeta K. Swamy, Lane F. Burgette, and Marie Lynn Miranda. A Bayesian growth mixture model to examine maternal hypertension and birth outcomes. Statistics in Medicine, 30(22):2721–2735, September 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cattle:2011:MIC

[565] Brian A. Cattle, Paul D. Baxter, Darren C. Greenwood, Christopher P. Gale, and Robert M. West. Multiple imputation for completion of a national clinical audit dataset. Statistics in Medicine, 30(22):2736–2753, September 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Govindarajulu:2011:FMA

[566] Usha S. Govindarajulu, Haiqun Lin, Kathryn L. Lunetta, and R. B. D’Agostino, Sr. Frailty models: Applications to biomedical and genetic studies. Statistics in Medicine, 30(22):2754–2764, September 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Viechtbauer:2011:BRG

[567] Wolfgang Viechtbauer. Book review: Graphics for statistics and data analysis with R. Kevin J. Keen, Chapman and Hall/CRC, Boca Raton, 2010. No. of pages: 489. Price: $69.95. ISBN: 978-1-58488-087-5. Statistics in Medicine, 30(22):2765–2766, September 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubio:2011:EMB

[568] Doris McGartland Rubio, Deborah J. del Junco, Rafia Bhore, Christopher J. Lindsell, Robert A. Oster, Knut M. Wittkowski, Leah J. Welty, Yi-Ju Li, Dave DeMets, and Biostatistics and Epidemiology and Research Design (BERD) Key Function Committee of the Clinical and Translational Science Awards (CTSA) Consortium. Evaluation metrics for biostatistical and epidemiological collaborations. Statistics in Medicine, 30(23):2767–2777, October 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kane:2011:EBC

[569] Cathleen Kane and William M. Trochim. The end of the beginning: a commentary on ‘evaluation metrics for biostatistical and epidemiological collaborations’. Statistics in Medicine, 31(6):2778–2782, October 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubio:2011:EBC

[570] Doris McGartland Rubio, Deborah J. del Junco, Rafia Bhore, Christopher J. Lindsell, Robert A. Oster, Knut M. Wittkowski, Leah J. Welty, Yi-Ju Li, and Dave DeMets Biostatistics and Epidemiology and Research Design (BERD) Key Function Committee of the Clinical and Translational Science Awards (CTSA) Consortium. The end of the beginning: a commentary on ‘evaluation metrics for biostatistical and epidemiological collaborations’: a rejoinder. Statistics in Medicine, 31(6):2783–2784, October 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gelfond:2011:PEA

[571] Jonathan A. L. Gelfond, Elizabeth Heitman, Brad H. Pollock, and Craig M. Klugman. Principles for the ethical analysis of clinical and translational research. Statistics in Medicine, 31(6):2785–2792, October 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ivanova:2011:OSS

[572] Anastasia Ivanova, Bahjat Qaqish, and David A. Schoenfeld. Optimality, sample size, and power calculations for the sequential parallel comparison design. Statistics in Medicine, 31(6):2793–2803, October 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neal:2011:IET

[573] Dan Neal, George Casella, Mark C. K. Yang, and Samuel S. Wu. Interval estimation in two-stage, drop-the-losers clinical trials with flexible treatment selection. Statistics in Medicine, 31(6):2804–2814, October 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gombay:2011:MBO

[574] Edit Gombay, Abdulkadir A. Hussein, and Stefan H. Steiner. Monitoring binary outcomes using risk-adjusted charts: a comparative study. Statistics in Medicine, 31(6):2815–2826, October 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Waller:2011:SOS

[575] Lance A. Waller, Aila Särkkä, Viktor Olsbo, Mari Myllymäki, Ioanna G. Panoutsopoulou, William R. Kennedy, and Gwen Wendelschafer-Crabb. Second-order spatial analysis of epidermal nerve fibers. Statistics in Medicine, 31(6):2827–2841, October 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2011:QTF

[576] Hongquan Xu, Qing Shen, Xiaowei Yang, and Steven Shoptaw. A quasi F -test for functional linear models with functional covariates and its application to longitudinal data. Statistics in Medicine, 31(6):2842–2853, October 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2011:CGS

[577] Yang Xu, Gustavo P. Sudre, Wei Wang, Douglas J. Weber, and Robert E. Kass. Characterizing global statistical significance of spatiotemporal hot spots in magnetoencephalography/electroencephalography source space via excursion algorithms. Statistics in Medicine, 31(6):2854–2866, October 15, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Foster:2011:SIR

[578] Jared C. Foster, Jeremy M. G. Taylor, and Stephen J. Ruberg. Subgroup identification from randomized clinical trial data. Statistics in Medicine, 30(24):2867–2880, October 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ganju:2011:BSC

[579] Jitendra Ganju and Kefei Zhou. The benefit of stratification in clinical trials revisited. Statistics in Medicine, 30(24):2881–2889, October 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sverdlov:2011:ORA

[580] Oleksandr Sverdlov, Yevgen Tymofyeyev, and Weng Kee Wong. Optimal response-adaptive randomized designs for multi-armed survival trials. Statistics in Medicine, 30(24):2890–2910, October 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2011:MMA

[581] Yan Ma and Madhu Mazumdar. Multivariate meta-analysis: a robust approach based on the theory of U-statistic. Statistics in Medicine, 30 (24):2911–2929, October 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2011:BIJ

[582] Yangxin Huang, Getachew Dagne, and Lang Wu. Bayesian inference on joint models of HIV dynamics for time-to-event and longitudinal data with skewness and covariate measurement errors. Statistics in Medicine, 30(24):2930–2946, October 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roysland:2011:ADI

[583] Kjetil Røysland, Jon Michael Gran, Bruno Ledergerber, Viktor von Wyl, James Young, and Odd O. Aalen. Analyzing direct and indirect effects of treatment using dynamic path analysis applied to data from the Swiss HIV cohort study. Statistics in Medicine, 30(24):2947–2958, October 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2011:CCO

[584] Zhongxue Chen and Saralees Nadarajah. Comments on ‘Choosing an optimal method to combine p-values’ by Sungho Won, Nathan Morris, Qing Lu and Robert C. Elston, Statistics in Medicine 2009; 28:1537– 1553. Statistics in Medicine, 30(24):2959–2961, October 30, 2011. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1].

Won:2011:AR

[585] Sungho Won, Nathan Morris, Qing Lu, and Robert C. Elston. Authors’ reply. Statistics in Medicine, 30(24):2962–2964, October 30, 2011. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Imberger:2011:CSM

[586] Georgina Imberger, Christian Gluud, and Jørn Wetterslev. Comments on ‘Sequential methods for random-effects meta-analysis’ by J. P. Higgins, A. Whitehead and M. Simmonds, Statistics in Medicine 2010; DOI: 10.1002/sim.4088. Statistics in Medicine, 30(24):2965–2966, October 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anzures-Cabrera:2011:EFM

[587] Judith Anzures-Cabrera, Ameet Sarpatwari, and Julian Pt Higgins. Expressing findings from meta-analyses of continuous outcomes in terms of risks. Statistics in Medicine, 30(25):2967–2985, November 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaizar:2011:ETE

[588] Eloise E. Kaizar. Estimating treatment effect via simple cross design synthesis. Statistics in Medicine, 30(25):2986–3009, November 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2011:STM

[589] Liang Zhu, Jianguo Sun, Deo Kumar Srivastava, Xingwei Tong, Wendy Leisenring, Hui Zhang, and Leslie L. Robison. Semiparametric transformation models for joint analysis of multivariate recurrent and terminal events. Statistics in Medicine, 30(25):3010–3023, November 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Talbot:2011:VEC

[590] Denis Talbot, Thierry Duchesne, Jacques Brisson, and Nathalie Vandal. Variance estimation and confidence intervals for the standardized mortality ratio with application to the assessment of a cancer screening program. Statistics in Medicine, 30(25):3024–3037, November 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prisman:2011:TEP

[591] Eliezer Z. Prisman, Amiram Gafni, and Antonio Finelli. Testing the evolution process of prostate-specific antigen in early stage prostate cancer: what is the proper underlying model? Statistics in Medicine, 30(25): 3038–3049, November 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:2011:BIC

[592] Richard H. Jones. Bayesian information criterion for longitudinal and clustered data. Statistics in Medicine, 30(25):3050–3056, November 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachin:2011:PSS

[593] John M. Lachin. Power and sample size evaluation for the Cochran– Mantel–Haenszel mean score (Wilcoxon rank sum) test and the Cochran– Armitage test for trend. Statistics in Medicine, 30(25):3057–3066, November 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scala:2011:TEA

[594] L. Di Scala and E. Glimm. Time-to-event analysis with treatment arm selection at interim. Statistics in Medicine, 30(26):3067–3081, November 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pullenayegum:2011:IRP

[595] Eleanor M. Pullenayegum. An informed reference prior for betweenstudy heterogeneity in meta-analyses of binary outcomes. Statistics in Medicine, 32(11):3082–3094, November 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Corberan-Vallet:2011:CPI

[596] Ana Corberán-Vallet and Andrew B. Lawson. Conditional predictive inference for online surveillance of spatial disease incidence. Statistics in Medicine, 32(11):3095–3116, November 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carey:2011:WCM

[597] Vincent J. Carey and You-Gan Wang. Working covariance model selection for generalized estimating equations. Statistics in Medicine, 32(11): 3117–3124, November 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morina:2011:SMH

[598] David Moriña, Pedro Puig, José Ríos, Anna Vilella, and Antoni Trilla. A statistical model for hospital admissions caused by seasonal diseases. Statistics in Medicine, 32(11):3125–3136, November 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Romdhani:2011:ABV

[599] Héla Romdhani and Lajmi Lakhal-Chaieb. On the association between variables with lower detection limits. Statistics in Medicine, 32(11):3137– 3148, November 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Long:2011:NMI

[600] Qi Long, Xiaoxi Zhang, and Chiu-Hsieh Hsu. Nonparametric multiple imputation for receiver operating characteristics analysis when some biomarker values are missing at random. Statistics in Medicine, 32(11): 3149–3161, November 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rohmel:2011:SSA

[601] Joachim Röhmel and Iris Pigeot. Statistical strategies for the analysis of clinical trials with an experimental treatment, an active control and placebo, and a prespecified fixed non-inferiority margin for the difference in means. Statistics in Medicine, 32(11):3162–3164, November 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hida:2011:RJR

[602] Eisuke Hida and Toshiro Tango. Response to Joachim Röhmel and Iris Pigeot. Statistics in Medicine, 32(11):3165, November 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2011:ISCb

[603] Richard J. Cook and Pierre-Jérôme Bergeron. Information in the sample covariate distribution in prevalent cohorts: Addendum. Statistics in Medicine, 32(11):3166, November 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sattar:2011:ANI

[604] Abdus Sattar, Lisa A. Weissfeld, and Geert Molenberghs. Analysis of non-ignorable missing and left-censored longitudinal data using a weighted random effects tobit model. Statistics in Medicine, 30(27): 3167–3180, November 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lorenz:2011:MAM

[605] Douglas J. Lorenz, Somnath Datta, and Susan J. Harkema. Marginal association measures for clustered data. Statistics in Medicine, 30(27): 3181–3191, November 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2011:AMO

[606] Ian R. White, Eleftheria Kalaitzaki, and Simon G. Thompson. Allowing for missing outcome data and incomplete uptake of randomised interventions, with application to an Internet-based alcohol trial. Statistics in Medicine, 30(27):3192–3207, November 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schildcrout:2011:ALH

[607] Jonathan S. Schildcrout, Sebastien Haneuse, Josh F. Peterson, Joshua C. Denny, Michael E. Matheny, Lemuel R. Waitman, and Randolph A. Miller. Analyses of longitudinal, hospital clinical laboratory data with application to blood glucose concentrations. Statistics in Medicine, 30 (27):3208–3220, November 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2011:ICI

[608] Minjung Lee and Jason P. Fine. Inference for cumulative incidence quantiles via parametric and nonparametric approaches. Statistics in Medicine, 30(27):3221–3235, November 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2011:PMC

[609] Zhaoxia Yu and Li Deng. Pseudosibship methods in the case-parents design. Statistics in Medicine, 30(27):3236–3251, November 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martella:2011:MMR

[610] F. Martella, J. K. Vermunt, M. Beekman, R. G. J. Westendorp, P. E. Slagboom, and J. J. Houwing-Duistermaat. A mixture model with random-effects components for classifying sibling pairs. Statistics in Medicine, 30(27):3252–3264, November 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stamey:2011:BAA

[611] J. D. Stamey, J. W. Seaman, and D. Young. A Bayesian approach to adjust for diagnostic misclassification between two mortality causes in Poisson regression. Statistics in Medicine, 30(27):3265, November 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2011:MMCb

[612] Chunling Liu, Aiyi Liu, and Susan Halabi. A min-max combination of biomarkers to improve diagnostic accuracy. Statistics in Medicine, 30 (27):3266, November 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehta:2011:AIS

[613] Cyrus R. Mehta and Stuart J. Pocock. Adaptive increase in sample size when interim results are promising: a practical guide with examples. Statistics in Medicine, 30(28):3267–3284, December 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Emerson:2011:CAI

[614] Scott S. Emerson, Gregory P. Levin, and Sarah C. Emerson. Comments on ‘Adaptive increase in sample size when interim results are promising: a practical guide with examples’. Statistics in Medicine, 30(28):3285– 3301, December 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehta:2011:ARC

[615] Cyrus R. Mehta and Stuart J. Pocock. Authors’ response to “Comment on adaptive increase in sample size when interim results are promising”. Statistics in Medicine, 30(28):3302–3303, December 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Noma:2011:CIR

[616] Hisashi Noma. Confidence intervals for a random-effects meta-analysis based on Bartlett-type corrections. Statistics in Medicine, 30(28):3304– 3312, December 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saeki:2011:NIT

[617] Hiroyuki Saeki and Toshiro Tango. Non-inferiority test and confidence interval for the difference in correlated proportions in diagnostic procedures based on multiple raters. Statistics in Medicine, 30(28):3313–3327, December 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2011:REC

[618] Qiang Xu, Myunghee Cho Paik, Tatjana Rundek, Mitchell S. V. Elkind, and Ralph L. Sacco. Reweighting estimators for Cox regression with missing covariate data: Analysis of insulin resistance and risk of stroke in the northern Manhattan study. Statistics in Medicine, 30(28):3328– 3340, December 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sauerbrei:2011:NSM

[619] Willi Sauerbrei and Patrick Royston. A new strategy for meta-analysis of continuous covariates in observational studies. Statistics in Medicine, 30 (28):3341–3360, December 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2011:IAM

[620] Shuangge Ma, Jian Huang, Fengrong Wei, Yang Xie, and Kuangnan Fang. Integrative analysis of multiple cancer prognosis studies with gene expression measurements. Statistics in Medicine, 30(28):3361–3371, December 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fleischer:2011:HRI

[621] Frank Fleischer, Birgit Gaschler-Markefski, and Erich Bluhmki. How is retrospective independent review influenced by investigator-introduced informative censoring: a quantitative approach. Statistics in Medicine, 30(29):3373–3386, December 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2011:STC

[622] Terry C. K. Lee, C. B. Dean, and Robert Semenciw. Short-term cancer mortality projections: a comparative study of prediction methods. Statistics in Medicine, 30(29):3387–3402, December 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nguyen:2011:SEH

[623] Thuan Nguyen and Jiming Jiang. Simple estimation of hidden correlation in repeated measures. Statistics in Medicine, 30(29):3403–3415, December 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ahmed:2011:ACC

[624] Anwar E. Ahmed, Donna K. McClish, and Christine M. Schubert. Accuracy and cost comparison in medical testing using sequential testing strategies. Statistics in Medicine, 30(29):3416–3430, December 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zeng:2011:BAR

[625] L. Zeng and S. Zhou. A Bayesian approach to risk-adjusted outcome monitoring in healthcare. Statistics in Medicine, 30(29):3431–3446, December 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yucel:2011:GBR

[626] Recai M. Yucel, Yulei He, and Alan M. Zaslavsky. Gaussian-based routines to impute categorical variables in health surveys. Statistics in Medicine, 30(29):3447–3460, December 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2011:SPE

[627] Yongqiang Tang. Size and power estimation for the Wilcoxon–Mann– Whitney test for ordered categorical data. Statistics in Medicine, 30 (29):3461–3470, December 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:2011:CHM

[628] Rachel MacKay Altman. Comment on ‘Hidden Markov models for zeroinflated Poisson counts with an application to substance use’ by S. M. DeSantis and D. Bandyopadhyay. Statistics in Medicine, 30(29):3471– 3472, December 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeSantis:2011:AR

[629] Stacia M. DeSantis. Author’s reply. Statistics in Medicine, 30(29):3473, December 20, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenkranz:2011:IRA

[630] Gerd K. Rosenkranz. The impact of randomization on the analysis of clinical trials. Statistics in Medicine, 30(30):3475–3487, December 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qu:2011:NCA

[631] Yongming Qu, Rong Liu, Alexei Dmitrienko, and Walter Offen. A new classification approach for comparing two active treatments when there is no prior projection on which one is better. Statistics in Medicine, 30 (30):3488–3495, December 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tamura:2011:ETE

[632] Roy N. Tamura, Xiaohong Huang, and Dennis D. Boos. Estimation of treatment effect for the sequential parallel design. Statistics in Medicine, 30(30):3496–3506, December 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brittain:2011:HRT

[633] Erica Brittain and Dean Follmann. A hierarchical rank test for crossover trials with censored data. Statistics in Medicine, 30(30):3507–3519, December 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tom:2011:IOT

[634] Brian D. M. Tom and Vernon T. Farewell. Intermittent observation of time-dependent explanatory variables: a multistate modelling approach. Statistics in Medicine, 30(30):3520–3531, December 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dong:2011:PNP

[635] Tuochuan Dong, Lili Tian, Alan Hutson, and Chengjie Xiong. Parametric and non-parametric confidence intervals of the probability of identifying early disease stage given sensitivity to full disease and specificity with three ordinal diagnostic groups. Statistics in Medicine, 30(30):3532–3545, December 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2011:AAR

[636] Chia-Cheng Chen and Huiman X. Barnhart. Assessing agreement with repeated measures for random observers. Statistics in Medicine, 30(30):

3546–3559, December 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Egleston:2011:IMD

[637] Brian L. Egleston, Suzanne M. Miller, and Neal J. Meropol. The impact of misclassification due to survey response fatigue on estimation and identifiability of treatment effects. Statistics in Medicine, 30(30):3560– 3572, December 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berger:2011:COE

[638] Vance W. Berger. Can objective endpoints be manipulated in unmasked trials? Statistics in Medicine, 30(30):3573–3574, December 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2011:ARV

[639] Anonymous. Author’s reply to Vance W. Berger: Can objective endpoints be manipulated in unmasked trials? Statistics in Medicine, 30 (30):3575, December 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2012:DIM

[640] Yulei He and Alan M. Zaslavsky. Diagnosing imputation models by applying target analyses to posterior replicates of completed data. Statistics in Medicine, 31(1):1–18, January 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2012:SSC

[641] Kaifeng Lu. Sample size calculations with multiplicity adjustment for longitudinal clinical trials with missing data. Statistics in Medicine, 31 (1):19–28, January 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moser:2012:ETR

[642] Barry Kurt Moser and Susan Halabi. Estimation and testing of the relative risk of disease in case-control studies with a set of k matched controls per case with known prevalence of disease. Statistics in Medicine, 31(1):29–44, January 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Herberich:2012:DTI

[643] Esther Herberich and Torsten Hothorn. Dunnett-type inference in the frailty Cox model with covariates. Statistics in Medicine, 31(1):45–55, January 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2012:GLM

[644] Qiang Zhang and Edward Haksing Ip. Generalized linear model for partially ordered data. Statistics in Medicine, 31(1):56–68, January 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2012:LBC

[645] Jason Kunming Huang, Alyssa Yanzhen Liu, and John Jen Tai. Likelihood-based concordance tests for analysis of ascertained twin pair multinomial data. Statistics in Medicine, 31(1):69–79, January 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miller:2012:LCM

[646] William E. Miller. A latent class method for the selection of prototypes using expert ratings. Statistics in Medicine, 31(1):80–92, January 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2012:CEN

[647] Nancy R. Cook and Nina P. Paynter. Comments on ‘Extensions of net reclassification improvement calculations to measure usefulness of new biomarkers’ by M. J. Pencina, R. B. D’Agostino, Sr. and E. W. Steyerberg. Statistics in Medicine, 31(1):93–95, January 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2012:AR

[648] Michael J. Pencina, Ralph B. D’Agostino, Sr., and Ewout W. Steyerberg. Authors’ reply. Statistics in Medicine, 31(1):96–97, January 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Glimm:2012:CAI

[649] Ekkehard Glimm. Comments on ‘Adaptive increase in sample size when interim results are promising: a practical guide with examples’ by C. R. Mehta and S. J. Pocock. Statistics in Medicine, 31(1):98–99, January 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehta:2012:AR

[650] Cyrus R. Mehta and Stuart J. Pocock. Authors’ reply. Statistics in Medicine, 31(1):99–100, January 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2012:NME

[651] Michael J. Pencina, Ralph B. D’Agostino, Sr., and Olga V. Demler. Novel metrics for evaluating improvement in discrimination: net reclassification and integrated discrimination improvement for normal variables and nested models. Statistics in Medicine, 31(2):101–113, January 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rapsomaniki:2012:FQN

[652] Eleni Rapsomaniki, Ian R. White, Angela M. Wood, Simon G. Thompson, and Emerging Risk Factors Collaboration. A framework for quantifying net benefits of alternative prognostic models. Statistics in Medicine, 31(2):114–130, January 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheng:2012:SSC

[653] Dunlei Cheng, Adam J. Branscum, and Wesley O. Johnson. Sample size calculations for ROC studies: parametric robustness and Bayesian nonparametrics. Statistics in Medicine, 31(2):131–142, January 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:2012:LMM

[654] Paul S. Albert. A linear mixed model for predicting a binary event from longitudinal data under random effects misspecification. Statistics in Medicine, 31(2):143–154, January 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shan:2012:EEA

[655] Guogen Shan, Changxing Ma, Alan D. Hutson, and Gregory E. Wilding. An efficient and exact approach for detecting trends with binary endpoints. Statistics in Medicine, 31(2):155–164, January 30, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Noma:2012:ODP

[656] Hisashi Noma and Shigeyuki Matsui. The optimal discovery procedure in multiple significance testing: an empirical Bayes approach. Statistics in Medicine, 31(2):165–176, January 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2012:SSS

[657] Xiao-Zhou Li, Jin-Feng Wang, Wei-Zhong Yang, Zhong-Jie Li, and Sheng-Jie Lai. A spatial scan statistic for nonisotropic two-level risk cluster. Statistics in Medicine, 31(2):177–187, January 30, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2012:BAS

[658] Yi-Fu Wang and Tsai-Hung Fan. Bayesian analysis of the structural equation models with application to a longitudinal myopia trial. Statistics in Medicine, 31(2):188–200, January 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hemming:2012:PSR

[659] Karla Hemming, Russell James Bowater, and Richard J. Lilford. Pooling systematic reviews of systematic reviews: a Bayesian panoramic metaanalysis. Statistics in Medicine, 31(3):201–216, February 10, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fedorov:2012:ODF

[660] Valerii Fedorov, Yuehui Wu, and Rongmei Zhang. Optimal dose-finding designs with correlated continuous and discrete responses. Statistics in Medicine, 31(3):217–234, February 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fujii:2012:EIB

[661] Yosuke Fujii, Masayuki Henmi, and Toshiharu Fujita. Evaluating the interaction between the therapy and the treatment in clinical trials by the propensity score weighting method. Statistics in Medicine, 31(3): 235–252, February 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Minas:2012:HPD

[662] Giorgos Minas, Fabio Rigat, Thomas E. Nichols, John A. D. Aston, and Nigel Stallard. A hybrid procedure for detecting global treatment effects in multivariate clinical trials: theory and applications to fMRI studies. Statistics in Medicine, 31(3):253–268, February 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huszti:2012:RSM

[663] Ella Huszti, Michal Abrahamowicz, Ahmadou Alioum, Christine Binquet, and Catherine Quantin. Relative survival multistate Markov model. Statistics in Medicine, 31(3):269–286, February 10, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Porta:2012:DMR

[664] Núria Porta, M. Luz Calle, Núria Malats, and Guadalupe Gómez. A dynamic model for the risk of bladder cancer progression. Statistics in Medicine, 31(3):287–300, February 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wason:2012:OMD

[665] James M. S. Wason, Adrian P. Mander, and Simon G. Thompson. Optimal multistage designs for randomised clinical trials with continuous outcomes. Statistics in Medicine, 31(4):301–312, February 20, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guolo:2012:HOL

[666] Annamaria Guolo. Higher-order likelihood inference in meta-analysis and meta-regression. Statistics in Medicine, 31(4):313–327, February 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kahan:2012:IAT

[667] Brennan C. Kahan and Tim P. Morris. Improper analysis of trials randomised using stratified blocks or minimisation. Statistics in Medicine, 31(4):328–340, February 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2012:WMT

[668] Changchun Xie. Weighted multiple testing correction for correlated tests. Statistics in Medicine, 31(4):341–352, February 20, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leon-Novelo:2012:OBM

[669] Luis Leon-Novelo, Elías Moreno, and George Casella. Objective Bayes model selection in probit models. Statistics in Medicine, 31(4):353–365, February 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

French:2012:GFE

[670] Benjamin French, Farhood Farjah, David R. Flum, and Patrick J. Heagerty. A general framework for estimating volume-outcome associations from longitudinal data. Statistics in Medicine, 31(4):366–382, February 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McCandless:2012:HPB

[671] Lawrence C. McCandless, Paul Gustafson, Adrian R. Levy, and Sylvia Richardson. Hierarchical priors for bias parameters in Bayesian sensitivity analysis for unmeasured confounding. Statistics in Medicine, 31 (4):383–396, February 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanBreukelen:2012:ELB

[672] Gerard J. P. van Breukelen and Math J. J. M. Candel. Efficiency loss because of varying cluster size in cluster randomized trials is smaller than literature suggests. Statistics in Medicine, 31(4):397–400, February 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2012:FJM

[673] Weili He, Xiting Cao, and Lu Xu. A framework for joint modeling and joint assessment of efficacy and safety endpoints for probability of success evaluation and optimal dose selection. Statistics in Medicine, 31(5):401– 419, February 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koopmeiners:2012:CEA

[674] Joseph S. Koopmeiners, Ziding Feng, and Margaret Sullivan Pepe. Conditional estimation after a two-stage diagnostic biomarker study that allows early termination for futility. Statistics in Medicine, 31(5):420– 435, February 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Katki:2012:EAD

[675] Hormuzd A. Katki, Yan Li, David W. Edelstein, and Philip E. Castle. Estimating the agreement and diagnostic accuracy of two diagnostic tests when one test is conducted on only a subsample of specimens. Statistics in Medicine, 31(5):436–448, February 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:2012:LMM

[676] Rachel MacKay Altman, A. John Petkau, Dean Vrecko, and Alex Smith. A longitudinal model for magnetic resonance imaging lesion count data in multiple sclerosis patients. Statistics in Medicine, 31(5):449–469, February 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosthoj:2012:EDT

[677] S. Rosthøj, N. Keiding, and K. Schmiegelow. Estimation of dynamic treatment strategies for maintenance therapy of children with acute lymphoblastic leukaemia: an application of history-adjusted marginal structural models. Statistics in Medicine, 31(5):470–488, February 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2012:PAR

[678] Minjung Lee, Kathleen A. Cronin, Mitchell H. Gail, and Eric J. Feuer. Predicting the absolute risk of dying from colorectal cancer and from other causes using population-based cancer registry data. Statistics in Medicine, 31(5):489–500, February 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pinsky:2012:EDA

[679] Paul F. Pinsky and B. Gallas. Enriched designs for assessing discriminatory performance — analysis of bias and variance. Statistics in Medicine, 31(6):501–515, March 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hirakawa:2012:ADF

[680] Akihiro Hirakawa. An adaptive dose-finding approach for correlated bivariate binary and continuous outcomes in phase I oncology trials. Statistics in Medicine, 31(6):516–532, March 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vandeKassteele:2012:ENT

[681] J. van de Kassteele, R. T. Hoogenveen, P. M. Engelfriet, P. H. M. van Baal, and H. C. Boshuizen. Estimating net transition probabilities from cross-sectional data with application to risk factors in chronic disease modeling. Statistics in Medicine, 31(6):533–543, March 15, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2012:FMV

[682] Zhaohua Lu and Xinyuan Song. Finite mixture varying coefficient models for analyzing longitudinal heterogeneous data. Statistics in Medicine, 31(6):544–560, March 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiang:2012:RMM

[683] Fang Xiang and Susan Murray. Restricted mean models for transplant benefit and urgency. Statistics in Medicine, 31(6):561–576, March 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nguyen:2012:PIE

[684] Viet Cuong Nguyen. Program impact evaluation using a matching method with panel data. Statistics in Medicine, 31(6):577–588, March 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Funatogawa:2012:ALM

[685] Ikuko Funatogawa and Takashi Funatogawa. An autoregressive linear mixed effects model for the analysis of unequally spaced longitudinal data with dose-modification. Statistics in Medicine, 31(6):589–599, March 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubio:2012:CEM

[686] Doris McGartland Rubio, Deborah J. del Junco, Rafia Bhore, Christopher J. Lindsell, Robert A. Oster, Knut M. Wittkowski, Leah J. Welty, Yi-Ju Li, Dave DeMets, and Biostatistics and Epidemiology and Research Design (BERD) Key Function Committee of the Clinical and Translational Science Awards (CTSA) Consortium. Correction: Evaluation metrics for biostatistical and epidemiological collaborations. Statistics in Medicine, 31(6):600, March 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2012:P

[687] Xiao-Hua Zhou and Laura Lee Johnson. Preface. Statistics in Medicine, 31(7):601, March 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2012:BTD

[688] Jingqing Hu and Baoyan Liu. The basic theory, diagnostic, and therapeutic system of traditional Chinese medicine and the challenges they bring to statistics. Statistics in Medicine, 31(7):602–605, March 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shuai:2012:IDS

[689] Ping Shuai, Xiao-Hua Zhou, Lixing Lao, and Xiaosong Li. Issues of design and statistical analysis in controlled clinical acupuncture trials: an analysis of English-language reports from Western journals. Statistics in Medicine, 31(7):606–618, March 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lao:2012:ETC

[690] Lixing Lao, Yi Huang, Chiguang Feng, Brian M. Berman, and Ming T. Tan. Evaluating traditional Chinese medicine using modern clinical trial design and statistical methodology: application to a randomized controlled acupuncture trial. Statistics in Medicine, 31(7):619–627, March 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2012:VSU

[691] X. H. Zhou, B. Chen, Y. M. Xie, F. Tian, H. Liu, and X. Liang. Variable selection using the optimal ROC curve: an application to a traditional Chinese medicine study on osteoporosis disease. Statistics in Medicine, 31(7):628–635, March 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2012:DHF

[692] Ping He, Ke Deng, Zhihai Liu, Delin Liu, Jun S. Liu, and Zhi Geng. Discovering herbal functional groups of traditional Chinese medicine. Statistics in Medicine, 31(7):636–642, March 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2012:BDR

[693] X. H. Zhou, S. L. Li, F. Tian, B. J. Cai, Y. M. Xie, Y. Pei, S. Kang, M. Fan, and J. P. Li. Building a disease risk model of osteoporosis based on traditional Chinese medicine symptoms and Western medicine risk factors. Statistics in Medicine, 31(7):643–652, March 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2012:DPA

[694] Baoyan Liu, Xuezhong Zhou, Yinhui Wang, Jingqing Hu, Liyun He, Runshun Zhang, Shibo Chen, and Yufeng Guo. Data processing and analysis in real-world traditional Chinese medicine clinical data: challenges and approaches. Statistics in Medicine, 31(7):653–660, March 30, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2012:REM

[695] Zheyu Wang and Xiao-Hua Zhou. Random effects models for assessing diagnostic accuracy of traditional Chinese doctors in absence of a gold standard. Statistics in Medicine, 31(7):661–671, March 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2012:SET

[696] Yang Li, Danhui Yi, Huiyun Zhang, and Yichen Qin. Syndrome evaluation in traditional Chinese medicine using second-order latent variable model. Statistics in Medicine, 31(7):672–680, March 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:2012:GPS

[697] Ping Feng, Xiao-Hua Zhou, Qing-Ming Zou, Ming-Yu Fan, and Xiao-Song Li. Generalized propensity score for estimating the average treatment effect of multiple treatments. Statistics in Medicine, 31(7):681–697, March 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuznetsova:2012:PAR

[698] Olga M. Kuznetsova and Yevgen Tymofyeyev. Preserving the allocation ratio at every allocation with biased coin randomization and minimization in studies with unequal allocation. Statistics in Medicine, 31(8): 701–723, April 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chi:2012:GHT

[699] Yueh-Yun Chi, Matthew Gribbin, Yvonne Lamers, Jesse F. Gregory III, and Keith E. Muller. Global hypothesis testing for high-dimensional repeated measures outcomes. Statistics in Medicine, 31(8):724–742, April 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Renfro:2012:BAM

[700] Lindsay A. Renfro, Qian Shi, Daniel J. Sargent, and Bradley P. Carlin. Bayesian adjusted R^{2} for the meta-analytic evaluation of surrogate timeto-event endpoints in clinical trials. Statistics in Medicine, 31(8):743–761, April 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2012:SSS

[701] Tonglin Zhang, Zuoyi Zhang, and Ge Lin. Spatial scan statistics with overdispersion. Statistics in Medicine, 31(8):762–774, April 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Danieli:2012:ENS

[702] Coraline Danieli, Laurent Remontet, Nadine Bossard, Laurent Roche, and Aurélien Belot. Estimating net survival: the importance of allowing for informative censoring. Statistics in Medicine, 31(8):775–786, April 13, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2012:SSD

[703] Youna Hu and Peter X.-K. Song. Sample size determination for quadratic inference functions in longitudinal design with dichotomous outcomes. Statistics in Medicine, 31(8):787–800, April 13, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2012:PAT

[704] Jianming Wang, Chunlei Ke, Qi Jiang, Charlie Zhang, and Steven Snapinn. Predicting analysis time in event-driven clinical trials with eventreporting lag. Statistics in Medicine, 31(9):801–811, April 30, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ko:2012:FVS

[705] Jin H. Ko and Abdus S. Wahed. Up-front versus sequential randomizations for inference on adaptive treatment strategies. Statistics in Medicine, 31(9):812–830, April 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hong:2012:PPA

[706] Shengyan Hong and Li Shi. Predictive power to assist phase 3 go/no go decision based on phase 2 data on a different endpoint. Statistics in Medicine, 31(9):831–843, April 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rahbar:2012:NTE

[707] Mohammad H. Rahbar, Zhongxue Chen, Sangchoon Jeon, Joseph C. Gardiner, and Jing Ning. A nonparametric test for equality of survival medians. Statistics in Medicine, 31(9):844–854, April 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Swihart:2012:MEP

[708] Bruce J. Swihart, Brian S. Caffo, Ciprian M. Crainiceanu, and Naresh M. Punjabi. Mixed effect Poisson log-linear models for clinical and epidemiological sleep hypnogram data. Statistics in Medicine, 31(9):855–870, April 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maruotti:2012:MNH

[709] Antonello Maruotti and Roberto Rocci. A mixed non-homogeneous hidden Markov model for categorical data, with application to alcohol consumption. Statistics in Medicine, 31(9):871–886, April 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yuan:2012:DCC

[710] Ao Yuan, Jinfeng Xu, Qingqi Yue, and Gang Zheng. Detecting casecontrol expression quantitative trait loci using locally most powerful or maximin robust rank tests. Statistics in Medicine, 31(9):887–900, April 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karuri:2012:TSB

[711] Stella Wanjugu Karuri and Richard Simon. A two-stage Bayesian design for co-development of new drugs and companion diagnostics. Statistics in Medicine, 31(10):901–914, May 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stephens:2012:AGE

[712] Alisa J. Stephens, Eric J. Tchetgen Tchetgen, and Victor De Gruttola. Augmented generalized estimating equations for improving efficiency and validity of estimation in cluster randomized trials by leveraging clusterlevel and individual-level covariates. Statistics in Medicine, 31(10):915– 930, May 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guerra:2012:ABL

[713] Matthew W. Guerra, Justine Shults, Jay Amsterdam, and Thomas Ten-Have. The analysis of binary longitudinal data with time-dependent covariates. Statistics in Medicine, 31(10):931–948, May 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schwartz:2012:SAU

[714] Scott Schwartz, Fan Li, and Jerome P. Reiter. Sensitivity analysis for unmeasured confounding in principal stratification settings with binary variables. Statistics in Medicine, 31(10):949–962, May 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Deng:2012:EIR

[715] Lisha Deng, Peter J. Diggle, and John Cheesbrough. Estimating incidence rates using exact or interval-censored data with an application to hospital-acquired infections. Statistics in Medicine, 31(10):963–977, May 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cleries:2012:BAP

[716] Ramon Clèries, Josepa Ribes, Maria Buxo, Alberto Ameijide, Rafael Marcos-Gragera, Jaume Galceran, José Miguel Martínez, and Yutaka Yasui. Bayesian approach to predicting cancer incidence for an area without cancer registration by using cancer incidence data from nearby areas. Statistics in Medicine, 31(10):978–987, May 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shahbaba:2012:PAM

[717] Babak Shahbaba, Catherine M. Shachaf, and Zhaoxia Yu. A pathway analysis method for genome-wide association studies. Statistics in Medicine, 31(10):988–1000, May 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lesaffre:2012:P

[718] Emmanuel Lesaffre and Thomas Gerds. Preface. Statistics in Medicine, 31(11-12):1001, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chevret:2012:BAC

[719] Sylvie Chevret. Bayesian adaptive clinical trials: a dream for statisticians only? Statistics in Medicine, 31(11-12):1002–1013, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abrahamowicz:2012:CAM

[720] Michal Abrahamowicz, Marie-Eve Beauchamp, and Marie-Pierre Sylvestre. Comparison of alternative models for linking drug exposure with adverse effects. Statistics in Medicine, 31(11-12):1014–1030, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stallard:2012:OSS

[721] Nigel Stallard. Optimal sample sizes for phase II clinical trials and pilot studies. Statistics in Medicine, 31(11-12):1031–1042, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nguyen:2012:DEO

[722] Thu Thuy Nguyen, Caroline Bazzoli, and France Mentré. Design evaluation and optimisation in crossover pharmacokinetic studies analysed by nonlinear mixed effects models. Statistics in Medicine, 31(11-12):1043– 1058, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jaki:2012:NCE

[723] Thomas Jaki and Martin J. Wolfsegger. Non-compartmental estimation of pharmacokinetic parameters for flexible sampling designs. Statistics in Medicine, 31(11-12):1059–1073, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2012:IIF

[724] Per Kragh Andersen and Niels Keiding. Interpretability and importance of functionals in competing risks and multistate models. Statistics in Medicine, 31(11-12):1074–1088, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koller:2012:CRC

[725] Michael T. Koller, Heike Raatz, Ewout W. Steyerberg, and Marcel Wolbers. Competing risks and the clinical community: irrelevance or ignorance? Statistics in Medicine, 31(11-12):1089–1097, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skaltsa:2012:OTE

[726] Konstantina Skaltsa, Lluís Jover, David Fuster, and Josep Lluís Carrasco. Optimum threshold estimation based on cost function in a multistate diagnostic setting. Statistics in Medicine, 31(11-12):1098–1109, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shinkins:2012:LEO

[727] B. Shinkins, R. Stevens, and R. Perera. Letter to the Editor: ‘Optimum threshold estimation based on cost function in a multistate diagnostic setting’ by K. Skaltsa, L. Jover, D. Fuster and J. L. Carrasco. Statistics in Medicine, 31(11-12):1110–1112, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Antolini:2012:PBM

[728] Laura Antolini and Maria Grazia Valsecchi. Performance of binary markers for censored failure time outcome: nonparametric approach based on proportions. Statistics in Medicine, 31(11-12):1113–1128, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2012:EDE

[729] S. D. Walter, P. Macaskill, Sarah J. Lord, and L. Irwig. Effect of dependent errors in the assessment of diagnostic or screening test accuracy when the reference standard is imperfect. Statistics in Medicine, 31(11-12):1129–1138, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Joly:2012:ESD

[730] Pierre Joly, Thomas A. Gerds, Vibeke Qvist, Daniel Commenges, and Niels Keiding. Estimating survival of dental fillings on the basis of interval-censored data and multi-state models. Statistics in Medicine, 31(11-12):1139–1149, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ambler:2012:EPS

[731] G. Ambler, S. Seaman, and R. Z. Omar. An evaluation of penalised survival methods for developing prognostic models with rare events. Statistics in Medicine, 31(11-12):1150–1161, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mazroui:2012:GJF

[732] Yassin Mazroui, Simone Mathoulin-Pélissier, Pierre Soubeyran, and Virginie Rondeau. General joint frailty model for recurrent event data with a dependent terminal event: Application to follicular lymphoma data. Statistics in Medicine, 31(11-12):1162–1176, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2012:ENT

[733] Woojoo Lee, A. Gusnanto, A. Salim, P. Magnusson, Xueling Sim, E. S. Tai, and Y. Pawitan. Estimating the number of true discoveries in genome-wide association studies. Statistics in Medicine, 31(11-12):1177– 1189, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsonaka:2012:TSM

[734] Roula Tsonaka, Annette H. M. van der Helm-van Mil, and Jeanine J. Houwing-Duistermaat. A two-stage mixed-effects model approach for gene-set analyses in candidate gene studies. Statistics in Medicine, 31 (11-12):1190–1202, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dazard:2012:LSB

[735] Jean-Eudes Dazard, J. Sunil Rao, and Sanford Markowitz. Local sparse bump hunting reveals molecular heterogeneity of colon tumors. Statistics in Medicine, 31(11-12):1203–1220, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rockova:2012:HBF

[736] Veronika Rockova, Emmanuel Lesaffre, Jolanda Luime, and Bob Löwenberg. Hierarchical Bayesian formulations for selecting variables in regression models. Statistics in Medicine, 31(11-12):1221–1237, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2012:IIO

[737] Dan Jackson, Ian R. White, and James Carpenter. Identifying influential observations in Bayesian models by using Markov chain Monte Carlo. Statistics in Medicine, 31(11-12):1238–1248, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kifley:2012:MLV

[738] Annette Kifley, Gillian Z. Heller, Ken J. Beath, David Bulger, Jun Ma, and Val Gebski. Multilevel latent variable models for global healthrelated quality of life assessment. Statistics in Medicine, 31(11-12):1249– 1264, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanGeloven:2012:HDD

[739] Nan van Geloven, Kimiko A. Broeze, Brent C. Opmeer, Ben Willem Mol, and Aeilko H. Zwinderman. How to deal with double partial verification when evaluating two index tests in relation to a reference test? Statistics in Medicine, 31(11-12):1265–1276, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hardouin:2012:TPS

[740] Jean-Benoit Hardouin, Sarah Amri, Mohand-Larbi Feddag, and Véronique Sébille. Towards power and sample size calculations for the comparison of two groups of patients with item response theory models. Statistics in Medicine, 31(11-12):1277–1290, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salim:2012:VRP

[741] Agus Salim, Qian Yang, and Marie Reilly. The value of reusing prior nested case-control data in new studies with different outcome. Statistics in Medicine, 31(11-12):1291–1302, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gerlinger:2012:CWC

[742] Christoph Gerlinger, Lutz Edler, Tim Friede, Meinhard Kieser, Christos T. Nakas, Martin Schumacher, Jørgen Seldrup, and Norbert Victor. Considerations on what constitutes a ‘qualified statistician’ in regulatory guidelines. Statistics in Medicine, 31(11-12):1303–1305, May 20–30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2012:OAP

[743] S. D. Walter, R. M. Turner, P. Macaskill, K. J. McCaffery, and L. Irwig. Optimal allocation of participants for the estimation of selection, preference and treatment effects in the two-stage randomised trial design. Statistics in Medicine, 31(13):1307–1322, June 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2012:ESM

[744] Ziqi Chen, Ning-Zhong Shi, Wei Gao, and Man-Lai Tang. Efficient semiparametric mean-association estimation for longitudinal binary responses. Statistics in Medicine, 31(13):1323–1341, June 15, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Desantis:2012:SBL

[745] Stacia M. Desantis, E. Andrés Houseman, Brent A. Coull, Catherine L. Nutt, and Rebecca A. Betensky. Supervised Bayesian latent class models for high-dimensional data. Statistics in Medicine, 31(13):1342–1360, June 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2012:PES

[746] Yan Li, Mitchell H. Gail, Dale L. Preston, Barry I. Graubard, and Jay H. Lubin. Piecewise exponential survival times and analysis of case-cohort data. Statistics in Medicine, 31(13):1361–1368, June 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2012:EQL

[747] Lili Yu, Ruifeng Yu, Liang Liu, and Ding-Geng Chen. Extended quasilikelihood with fractional polynomials in the frame of the accelerated failure time model. Statistics in Medicine, 31(13):1369–1379, June 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moore:2012:CIE

[748] Kelly L. Moore, Romain Neugebauer, Mark J. van der Laan, and Ira B. Tager. Causal inference in epidemiological studies with strong confounding. Statistics in Medicine, 31(13):1380–1404, June 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turner:2012:BRS

[749] Robin Turner. Book review: SAS and R: Data management, statistical analysis, and graphics. Statistics in Medicine, 31(13):1405–1406, June 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moreno:2012:GWR

[750] Santiago G. Moreno, Alex J. Sutton, John R. Thompson, A. E. Ades, Keith R. Abrams, and Nicola J. Cooper. A generalized weighting regression-derived meta-analysis estimator robust to small-study effects and heterogeneity. Statistics in Medicine, 31(14):1407–1417, June 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bagos:2012:CTC

[751] Pantelis G. Bagos. On the covariance of two correlated log-odds ratios. Statistics in Medicine, 31(14):1418–1431, June 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goeyvaerts:2012:EVC

[752] Nele Goeyvaerts, Niel Hens, Heidi Theeten, Marc Aerts, Pierre Van Damme, and Philippe Beutels. Estimating vaccination coverage for the trivalent measles–mumps–rubella vaccine from trivariate serological data. Statistics in Medicine, 31(14):1432–1449, June 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Erasto:2012:MTM

[753] Panu Erästö, Fabian Hoti, and Kari Auranen. Modeling transmission of multitype infectious agents: application to carriage of Streptococcus pneumoniae. Statistics in Medicine, 31(14):1450–1463, June 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Archer:2012:PCR

[754] K. J. Archer and A. A. A. Williams. L_{1} penalized continuation ratio models for ordinal response prediction using high-dimensional datasets. Statistics in Medicine, 31(14):1464–1474, June 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Milanzi:2012:IOH

[755] Elasma Milanzi, Ariel Alonso, and Geert Molenberghs. Ignoring overdispersion in hierarchical loglinear models: Possible problems and solutions. Statistics in Medicine, 31(14):1475–1482, June 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:2012:CPP

[756] E. M. Jones, J. R. Thompson, V. Didelez, and N. A. Sheehan. On the choice of parameterisation and priors for the Bayesian analyses of Mendelian randomisation studies. Statistics in Medicine, 31(14):1483– 1501, June 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2012:MHL

[757] Xiaoqin Tang, Zhehui Luo, and Joseph C. Gardiner. Modeling hospital length of stay by Coxian phase-type regression with heterogeneity. Statistics in Medicine, 31(14):1502–1516, June 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miao:2012:EBN

[758] Xiaopeng Miao, Yong-Cheng Wang, and Ashis Gangopadhyay. An entropy-based nonparametric test for the validation of surrogate endpoints. Statistics in Medicine, 31(14):1517–1530, June 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2012:TSI

[759] Y. H. Joshua Chen and Cong Chen. Testing superiority at interim analyses in a non-inferiority trial. Statistics in Medicine, 31(15):1531–1542, July 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2012:QDF

[760] Michael J. Pencina, Ralph B. D’Agostino, Sr., and Linye Song. Quantifying discrimination of Framingham risk functions with different survival C statistics. Statistics in Medicine, 31(15):1543–1553, July 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bunouf:2012:AIB

[761] P. Bunouf, J-M. Grouin, and G. Molenberghs. Analysis of an incomplete binary outcome derived from frequently recorded longitudinal continuous data: application to daily pain evaluation. Statistics in Medicine, 31(15): 1554–1571, July 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Waernbaum:2012:MMR

[762] Ingeborg Waernbaum. Model misspecification and robustness in causal inference: comparing matching with doubly robust estimation. Statistics in Medicine, 31(15):1572–1581, July 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burgess:2012:IBC

[763] Stephen Burgess and Simon G. Thompson. Improving bias and coverage in instrumental variable analysis with weak instruments for continuous and binary outcomes. Statistics in Medicine, 31(15):1582–1600, July 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Banerjee:2012:IRT

[764] Mousumi Banerjee, Ying Ding, and Anne-Michelle Noone. Identifying representative trees from ensembles. Statistics in Medicine, 31(15):1601– 1616, July 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williamson:2012:VES

[765] E. J. Williamson, R. Morley, A. Lucas, and J. R. Carpenter. Variance estimation for stratified propensity score estimators. Statistics in Medicine, 31(15):1617–1632, July 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2012:BSR

[766] Inyoung Kim, Herbert Pang, and Hongyu Zhao. Bayesian semiparametric regression models for evaluating pathway effects on continuous and binary clinical outcomes. Statistics in Medicine, 31(15):1633–1651, July 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2012:CSS

[767] Xiao Sun, Xiaoming Li, and Joshua Chen. Comments on ‘Sample size for equivalence trials: a case study from a vaccine lot consistency trial’ by J. Ganju, A. Izu and A. Anemona. Statistics in Medicine, 31(15): 1652–1653, July 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ganju:2012:CSS

[768] Jitendra Ganju and Allen Izu. Comments on ‘Sample size for equivalence trials: a case study from a vaccine lot consistency trial’. Statistics in Medicine, 31(15):1654, July 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mijoule:2012:MPR

[769] Guillaume Mijoule, Stéphanie Savy, and Nicolas Savy. Models for patients’ recruitment in clinical trials and sensitivity analysis. Statistics in Medicine, 31(16):1655–1674, July 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ho:2012:DBC

[770] Weang Kee Ho, John N. S. Matthews, Robin Henderson, Daniel Farewell, and Lauren R. Rodgers. Dropouts in the AB/BA crossover design. Statistics in Medicine, 31(16):1675–1687, July 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:2012:ITS

[771] Myron N. Chang, Jonathan J. Shuster, and Wei Hou. Improved two-stage tests for stratified phase II cancer clinical trials. Statistics in Medicine, 31(16):1688–1698, July 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaiser:2012:IRM

[772] Lee D. Kaiser. Inefficiency of randomization methods that balance on stratum margins and improvements with permuted blocks and a sequential method. Statistics in Medicine, 31(16):1699–1706, July 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2012:JAB

[773] Ning Li, Robert M. Elashoff, Gang Li, and Chi-Hong Tseng. Joint analysis of bivariate longitudinal ordinal outcomes and competing risks survival times with nonparametric distributions for random effects. Statistics in Medicine, 31(16):1707–1721, July 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rudser:2012:DFI

[774] Kyle D. Rudser, Michael L. LeBlanc, and Scott S. Emerson. Distributionfree inference on contrasts of arbitrary summary measures of survival. Statistics in Medicine, 31(16):1722–1737, July 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sitlani:2012:LSM

[775] Colleen M. Sitlani, Patrick J. Heagerty, Emily A. Blood, and Tor D. Tosteson. Longitudinal structural mixed models for the analysis of surgical trials with noncompliance. Statistics in Medicine, 31(16):1738–1760, July 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosen:2012:ORI

[776] Sophia Rosen and Ori Davidov. Order-restricted inference for multivariate longitudinal data with applications to the natural history of hearing loss. Statistics in Medicine, 31(16):1761–1773, July 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2012:RBP

[777] Gengqian Cai and Tianhui Zhou. A remark on ‘Bayesian predictive approach to interim monitoring in clinical trials’ by A. Dmitrienko and M-D. Wang. Statistics in Medicine, 31(16):1774–1776, July 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lau:2012:CC

[778] Bryan Lau, Stephen Gange, and Stephen R. Cole. Clarification and correction. Statistics in Medicine, 31(16):1777–1778, July 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shao:2012:MTC

[779] Jun Shao, Sheng Zhang, Jiwei Zhao, and Alan Chiang. Multiple testing for a combination drug with two study endpoints. Statistics in Medicine, 31(17):1779–1790, July 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2012:HBM

[780] Xun Chen, Zhaoling Meng, and Ji Zhang. Handling of baseline measurements in the analysis of crossover trials. Statistics in Medicine, 31(17):

1791–1803, July 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2012:BDT

[781] Lili Zhao, Jeremy M. G. Taylor, and Scott M. Schuetze. Bayesian decision theoretic two-stage design in phase II clinical trials with survival endpoint. Statistics in Medicine, 31(17):1804–1820, July 30, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vexler:2012:TSD

[782] Albert Vexler, Wan-Min Tsai, Gregory Gurevich, and Jihnhee Yu. Two-sample density-based empirical likelihood ratio tests based on paired data, with application to a treatment study of attention-deficit/ hyperactivity disorder and severe mood dysregulation. Statistics in Medicine, 31(17):1821–1837, July 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2012:MIL

[783] MinJae Lee, Lan Kong, and Lisa Weissfeld. Multiple imputation for leftcensored biomarker data based on Gibbs sampling method. Statistics in Medicine, 31(17):1838–1848, July 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehrotra:2012:EAS

[784] Devan V. Mehrotra, Shu-Chih Su, and Xiaoming Li. An efficient alternative to the stratified Cox model analysis. Statistics in Medicine, 31(17): 1849–1856, July 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friedrich:2012:RGM

[785] Jan O. Friedrich, Neill K. J. Adhikari, and Joseph Beyene. Ratio of geometric means to analyze continuous outcomes in meta-analysis: comparison to mean differences and ratio of arithmetic means using empiric data and simulation. Statistics in Medicine, 31(17):1857–1886, July 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Almirall:2012:DPS

[786] Daniel Almirall, Scott N. Compton, Meredith Gunlicks-Stoessel, Naihua Duan, and Susan A. Murphy. Designing a pilot sequential multiple assignment randomized trial for developing an adaptive treatment strategy. Statistics in Medicine, 31(17):1887–1902, July 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aalen:2012:ALU

[787] Odd O. Aalen. Armitage Lecture 2010: Understanding treatment effects: the value of integrating longitudinal data and survival analysis. Statistics in Medicine, 31(18):1903–1917, August 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehrotra:2012:FCA

[788] Devan V. Mehrotra and Adeniyi J. Adewale. Flagging clinical adverse experiences: reducing false discoveries without materially compromising power for detecting true signals. Statistics in Medicine, 31(18):1918– 1930, August 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2012:NPI

[789] John Whitehead, Helene Thygesen, Thomas Jaki, Scot Davies, Sarah Halford, Helen Turner, Natalie Cook, and Duncan Jodrell. A novel phase I/IIa design for early phase oncology studies and its application in the evaluation of MK-0752 in pancreatic cancer. Statistics in Medicine, 31 (18):1931–1943, August 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lai:2012:SDP

[790] Tze Leung Lai, Philip W. Lavori, and Mei-Chiung Shih. Sequential design of phase II-III cancer trials. Statistics in Medicine, 31(18):1944– 1960, August 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2012:BCC

[791] Yunzhi Lin and Zheng Su. Balancing continuous and categorical baseline covariates in sequential clinical trials using the area between empirical cumulative distribution functions. Statistics in Medicine, 31(18):1961– 1971, August 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:2012:BNI

[792] Taeyoung Park, Jong-Hyeon Jeong, and Jae Won Lee. Bayesian nonparametric inference on quantile residual life function: Application to breast cancer data. Statistics in Medicine, 31(18):1972–1985, August 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stirnemann:2012:IPB

[793] J. J. Stirnemann, A. Samson, and J. C. Thalabard. Individual predictions based on nonlinear mixed modeling: application to prenatal twin growth. Statistics in Medicine, 31(18):1986–1999, August 15, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Westreich:2012:PFE

[794] Daniel Westreich, Stephen R. Cole, Jessica G. Young, Frank Palella, Phyllis C. Tien, Lawrence Kingsley, Stephen J. Gange, and Miguel A. Hernán. The parametric g-formula to estimate the effect of highly active antiretroviral therapy on incident AIDS or death. Statistics in Medicine, 31(18):2000–2009, August 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2012:URW

[795] Yuelin Li and Ray Baser. Using R and WinBUGS to fit a generalized partial credit model for developing and evaluating patient-reported outcomes assessments. Statistics in Medicine, 31(18):2010–2026, August 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tamhane:2012:AETa

[796] Ajit C. Tamhane, Yi Wu, and Cyrus R. Mehta. Adaptive extensions of a two-stage group sequential procedure for testing primary and secondary endpoints (i): unknown correlation between the endpoints. Statistics in Medicine, 31(19):2027–2040, August 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tamhane:2012:AETb

[797] Ajit C. Tamhane, Yi Wu, and Cyrus R. Mehta. Adaptive extensions of a two-stage group sequential procedure for testing primary and secondary endpoints (II): sample size re-estimation. Statistics in Medicine, 31(19):2041–2054, August 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dong:2012:BFT

[798] Gaohong Dong, Weichung Joe Shih, Dirk Moore, Hui Quan, and Stephen Marcella. A Bayesian-frequentist two-stage single-arm phase II clinical trial design. Statistics in Medicine, 31(19):2055–2067, August 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kimani:2012:PGA

[799] Peter K. Kimani, Ekkehard Glimm, Willi Maurer, Jane L. Hutton, and Nigel Stallard. Practical guidelines for adaptive seamless phase II/III clinical trials that use Bayesian methods. Statistics in Medicine, 31 (19):2068–2085, August 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holte:2012:EUL

[800] S. E. Holte, T. W. Randolph, J. Ding, J. Tien, R. S. McClelland, J. M. Baeten, and J. Overbaugh. Efficient use of longitudinal CD4 counts and viral load measures in survival analysis. Statistics in Medicine, 31 (19):2086–2097, August 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Westreich:2012:SSF

[801] Daniel Westreich, Stephen R. Cole, Enrique F. Schisterman, and Robert W. Platt. A simulation study of finite-sample properties of marginal structural Cox proportional hazards models. Statistics in Medicine, 31(19):2098–2109, August 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Viroli:2012:UFM

[802] C. Viroli. Using factor mixture analysis to model heterogeneity, cognitive structure, and determinants of dementia: an application to the aging, demographics, and memory study. Statistics in Medicine, 31(19):2110– 2122, August 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heaton:2012:STA

[803] Matthew J. Heaton, David L. Banks, Jian Zou, Alan F. Karr, Gauri Datta, James Lynch, and Francisco Vera. A spatio-temporal absorbing state model for disease and syndromic surveillance. Statistics in Medicine, 31(19):2123–2136, August 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gaio:2012:RMM

[804] A. Rita Gaio, Joaquim Pinto da Costa, Ana Cristina Santos, Elisabete Ramos, and Carla Lopes. A restricted mixture model for dietary pattern analysis in small samples. Statistics in Medicine, 31(19):2137–2150, August 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2012:NPT

[805] Bushi Wang and Xinping Cui. A new partition testing strategy for multiple endpoints. Statistics in Medicine, 31(20):2151–2168, September 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teerenstra:2012:SSS

[806] Steven Teerenstra, Sandra Eldridge, Maud Graff, Esther de Hoop, and George F. Borm. A simple sample size formula for analysis of covariance in cluster randomized trials. Statistics in Medicine, 31(20):2169–2178, September 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kirkham:2012:MMA

[807] Jamie J. Kirkham, Richard D. Riley, and Paula R. Williamson. A multivariate meta-analysis approach for reducing the impact of outcome reporting bias in systematic reviews. Statistics in Medicine, 31(20):2179– 2195, September 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2012:EPI

[808] M. Ghosh, J. Song, J. J. Forster, R. Mitra, and B. Mukherjee. On the equivalence of posterior inference based on retrospective and prospective likelihoods: application to a case-control study of colorectal cancer. Statistics in Medicine, 31(20):2196–2208, September 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Westgate:2012:ECS

[809] Philip M. Westgate and Thomas M. Braun. The effect of cluster size imbalance and covariates on the estimation performance of quadratic inference functions. Statistics in Medicine, 31(20):2209–2222, September 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Androulakis:2012:EVS

[810] E. Androulakis, C. Koukouvinos, and F. Vonta. Estimation and variable selection via frailty models with penalized likelihood. Statistics in Medicine, 31(20):2223–2239, September 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teunis:2012:BDE

[811] P. F. M. Teunis, J. C. H. van Eijkeren, C. W. Ang, Y. T. H. P. van Duynhoven, J. B. Simonsen, M. A. Strid, and W. van Pelt. Biomarker dynamics: estimating infection rates from serological data. Statistics in Medicine, 31(20):2240–2248, September 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:2012:NPB

[812] Yi Pan, Michael Haber, Jingjing Gao, and Huiman X. Barnhart. A new permutation-based method for assessing agreement between two observers making replicated quantitative readings. Statistics in Medicine, 31(20):2249–2261, September 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leon:2012:PPS

[813] Andrew C. Leon, Donald Hedeker, Chunshan Li, and Hakan Demirtas. Performance of a propensity score adjustment in longitudinal studies with covariate-dependent representation. Statistics in Medicine, 31(20): 2262–2274, September 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2012:EIC

[814] Xianhong Xie, Xiaonan Xue, Stephen J. Gange, Howard D. Strickler, Mimi Y. Kim, and Wihs Hpv Study Group. Estimation and inference on correlations between biomarkers with repeated measures and leftcensoring due to minimum detection levels. Statistics in Medicine, 31 (21):2275–2289, September 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Avalos:2012:AME

[815] Marta Avalos, Yves Grandvalet, Nuria Duran Adroher, Ludivine Orriols, and Emmanuel Lagarde. Analysis of multiple exposures in the casecrossover design via sparse conditional likelihood. Statistics in Medicine, 31(21):2290–2302, September 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2012:NMR

[816] Bo Hu, Liang Li, Xiaofeng Wang, and Tom Greene. Nonparametric multistate representations of survival and longitudinal data with measurement error. Statistics in Medicine, 31(21):2303–2317, September 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Malizia:2012:EJN

[817] Nicholas Malizia and Elizabeth A. Mack. Enhancing the Jacquez k nearest neighbor test for space-time interaction. Statistics in Medicine, 31 (21):2318–2334, September 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnson:2012:SES

[818] Lynn M. Johnson and Robert L. Strawderman. A smoothing expectation and substitution algorithm for the semiparametric accelerated failure time frailty model. Statistics in Medicine, 31(21):2335–2358, September 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Emily:2012:INS

[819] M. Emily. IndOR: a new statistical procedure to test for SNP-SNP epistasis in genome-wide association studies. Statistics in Medicine, 31 (21):2359–2373, September 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nysen:2012:TGF

[820] Ruth Nysen, Marc Aerts, and Christel Faes. Testing goodness of fit of parametric models for censored data. Statistics in Medicine, 31(21): 2374–2385, September 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xia:2012:BME

[821] Michelle Xia and Paul Gustafson. A Bayesian method for estimating prevalence in the presence of a hidden sub-population. Statistics in Medicine, 31(21):2386–2398, September 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schisterman:2012:I

[822] Enrique F. Schisterman and Paul S. Albert. Introduction. Statistics in Medicine, 31(22):2399, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Erickson:2012:MMB

[823] Heidi S. Erickson. Measuring molecular biomarkers in epidemiologic studies: laboratory techniques and biospecimen considerations. Statistics in Medicine, 31(22):2400–2413, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pfeiffer:2012:SDR

[824] Ruth M. Pfeiffer, Liliana Forzani, and Efstathia Bura. Sufficient dimension reduction for longitudinally measured predictors. Statistics in Medicine, 31(22):2414–2427, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roy:2012:BOR

[825] Anindya Roy, Michelle Danaher, Sunni L. Mumford, and Zhen Chen. A Bayesian order-restricted model for hormonal dynamics during menstrual cycles of healthy women. Statistics in Medicine, 31(22):2428–2440, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schildcrout:2012:ODS

[826] Jonathan S. Schildcrout, Sunni L. Mumford, Zhen Chen, Patrick J. Heagerty, and Paul J. Rathouz. Outcome-dependent sampling for longitudinal binary response data based on a time-varying auxiliary variable. Statistics in Medicine, 31(22):2441–2456, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:2012:NSM

[827] Paul S. Albert and Enrique F. Schisterman. Novel statistical methodology for analyzing longitudinal biomarker data. Statistics in Medicine, 31(22):2457–2460, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitcomb:2012:ASE

[828] Brian W. Whitcomb, Neil J. Perkins, Zhiwei Zhang, Aijun Ye, and Robert H. Lyles. Assessment of skewed exposure in case-control studies with pooling. Statistics in Medicine, 31(22):2461–2472, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2012:LRA

[829] Z. Zhang, A. Liu, R. H. Lyles, and B. Mukherjee. Logistic regression analysis of biomarker data subject to pooling and dichotomization. Statistics in Medicine, 31(22):2473–2484, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lyles:2012:LBM

[830] Robert H. Lyles, Li Tang, Ji Lin, Zhiwei Zhang, and Bhramar Mukherjee. Likelihood-based methods for regression analysis with binary exposure status assessed by pooling. Statistics in Medicine, 31(22):2485–2497, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vexler:2012:ETB

[831] Albert Vexler, Wan-Min Tsai, and Yaakov Malinovsky. Estimation and testing based on data subject to measurement errors: from parametric to non-parametric likelihood methods. Statistics in Medicine, 31(22): 2498–2512, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schisterman:2012:BR

[832] Enrique F. Schisterman and Paul S. Albert. The biomarker revolution. Statistics in Medicine, 31(22):2513–2515, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2012:EDG

[833] Jinbo Chen, Guolian Kang, Tyler VanderWeele, Cuilin Zhang, and Bhramar Mukherjee. Efficient designs of gene-environment interaction studies: implications of Hardy-Weinberg equilibrium and gene-environment independence. Statistics in Medicine, 31(22):2516–2530, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mukherjee:2012:PIA

[834] Bhramar Mukherjee, Yi-An Ko, Tyler VanderWeele, Anindya Roy, Sung Kyun Park, and Jinbo Chen. Principal interactions analysis for repeated measures data: application to gene-gene and gene-environment interactions. Statistics in Medicine, 31(22):2531–2551, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanderWeele:2012:SAI

[835] Tyler J. VanderWeele, Bhramar Mukherjee, and Jinbo Chen. Sensitivity analysis for interactions under unmeasured confounding. Statistics in Medicine, 31(22):2552–2564, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fan:2012:DGG

[836] Ruzong Fan, Paul S. Albert, and Enrique F. Schisterman. A discussion of gene-gene and gene-environment interactions and longitudinal genetic analysis of complex traits. Statistics in Medicine, 31(22):2565– 2568, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Louis:2012:ETT

[837] Germaine M. Buck Louis and Rajeshwari Sundaram. Exposome: time for transformative research. Statistics in Medicine, 31(22):2569–2575, September 28, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demler:2012:MDT

[838] Olga V. Demler, Michael J. Pencina, and Ralph B. D’Agostino, Sr. Misuse of DeLong test to compare AUCs for nested models. Statistics in Medicine, 31(23):2577–2587, October 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmid:2012:CEE

[839] Matthias Schmid and Sergej Potapov. A comparison of estimators to evaluate the discriminatory power of time-to-event models. Statistics in Medicine, 31(23):2588–2609, October 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Calster:2012:ESN

[840] Ben Van Calster, Vanya Van Belle, Yvonne Vergouwe, Dirk Timmerman, Sabine Van Huffel, and Ewout W. Steyerberg. Extending the c-statistic to nominal polytomous outcomes: the Polytomous Discrimination Index. Statistics in Medicine, 31(23):2610–2626, October 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choodari-Oskooei:2012:SSPa

[841] Babak Choodari-Oskooei, Patrick Royston, and Mahesh K. B. Parmar. A simulation study of predictive ability measures in a survival model i: Explained variation measures. Statistics in Medicine, 31(23):2627– 2643, October 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choodari-Oskooei:2012:SSPb

[842] B. Choodari-Oskooei, P. Royston, and Mahesh K. B. Parmar. A simulation study of predictive ability measures in a survival model II: explained randomness and predictive accuracy. Statistics in Medicine, 31 (23):2644–2659, October 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2012:NRO

[843] Xiao Song, Xiao-Hua Zhou, and Shuangge Ma. Nonparametric receiver operating characteristic-based evaluation for survival outcomes.

Statistics in Medicine, 31(23):2660–2675, October 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2012:CAC

[844] Xinhua Liu. Classification accuracy and cut point selection. Statistics in Medicine, 31(23):2676–2686, October 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gail:2012:UMR

[845] Mitchell H. Gail. Using multiple risk models with preventive interventions. Statistics in Medicine, 31(23):2687–2696, October 15, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Debray:2012:APP

[846] Thomas P. A. Debray, Hendrik Koffijberg, Yvonne Vergouwe, Karel G. M. Moons, and Ewout W. Steyerberg. Aggregating published prediction models with individual participant data: a comparison of different approaches. Statistics in Medicine, 31(23):2697–2712, October 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kipnis:2012:RCM

[847] Victor Kipnis, Douglas Midthune, Laurence S. Freedman, and Raymond J. Carroll. Regression calibration with more surrogates than mismeasured variables. Statistics in Medicine, 31(23):2713–2732, October 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Postmus:2012:MEH

[848] Douwe Postmus, Gimon de Graaf, Hans L. Hillege, Ewout W. Steyerberg, and Erik Buskens. A method for the early health technology assessment of novel biomarker measurement in primary prevention programs. Statistics in Medicine, 31(23):2733–2744, October 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elliott:2012:ABV

[849] Michael R. Elliott, Mary D. Sammel, and Jessica Faul. Associations between variability of risk factors and health outcomes in longitudinal studies. Statistics in Medicine, 31(23):2745–2756, October 15, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2012:SIH

[850] Anonymous. Special issue in honor of Jerome Cornfield on the centennial of his birth: Introduction. Statistics in Medicine, 31(24):2757–2759, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cornfield:2012:PR

[851] Jerome Cornfield. Principles of research. Statistics in Medicine, 31 (24):2760–2768, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:2012:CPR

[852] Peter Armitage. Comments on ‘Principles of research’ by Jerome Cornfield. Statistics in Medicine, 31(24):2769, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:2012:CPJ

[853] D. R. Cox. Comment on paper by J. Cornfield. Statistics in Medicine, 31(24):2770, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gail:2012:CPR

[854] Mitchell H. Gail. Comment on ‘Principles of research’ by Jerome Cornfield. Statistics in Medicine, 31(24):2771–2772, October 30, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:2012:CRR

[855] Sander Greenland. Cornfield, risk relativism, and research synthesis. Statistics in Medicine, 31(24):2773–2777, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubin:2012:PUC

[856] Donald B. Rubin. Potential updates to Cornfield’s 1959 ‘principles of research’. Statistics in Medicine, 31(24):2778–2779, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vandenbroucke:2012:VPL

[857] Jan P. Vandenbroucke. A voice from the past, lessons for today. Statistics in Medicine, 31(24):2780–2781, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenhouse:2012:BBE

[858] Joel B. Greenhouse. On becoming a Bayesian: Early correspondences between J. Cornfield and L. J. Savage. Statistics in Medicine, 31(24):2782– 2790, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittes:2012:JCC

[859] Janet Wittes. Jerome Cornfield’s contributions to early large randomized clinical trials and some reminiscences from the years of the slippery doorknobs. Statistics in Medicine, 31(24):2791–2797, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:2012:CIC

[860] Susan S. Ellenberg. Current issues in clinical trials: standing on the shoulders of Jerome Cornfield. Statistics in Medicine, 31(24):2798–2804, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2012:AMS

[861] Xuelin Huang and Jing Ning. Analysis of multi-stage treatments for recurrent diseases. Statistics in Medicine, 31(24):2805–2821, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Caille:2012:DCO

[862] Agnès Caille, Clémence Leyrat, and Bruno Giraudeau. Dichotomizing a continuous outcome in cluster randomized trials: impact on power. Statistics in Medicine, 31(24):2822–2832, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kwong:2012:ETA

[863] Koon Shing Kwong, Siu Hung Cheung, Anthony J. Hayter, and Miin-Jye Wen. Extension of three-arm non-inferiority studies to trials with multiple new treatments. Statistics in Medicine, 31(24):2833–2843, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2012:GSD

[864] Xiaoru Wu and Lu Cui. Group sequential and discretized sample size reestimation designs: a comparison of flexibility. Statistics in Medicine, 31 (24):2844–2857, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2012:EBJ

[865] Xinyi Xu, Michael L. Pennell, Bo Lu, and David M. Murray. Efficient Bayesian joint models for group randomized trials with multiple observation times and multiple outcomes. Statistics in Medicine, 31(24):2858– 2871, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ivanova:2012:TSD

[866] Anastasia Ivanova, Changfu Xiao, and Yevgen Tymofyeyev. Two-stage designs for phase 2 dose-finding trials. Statistics in Medicine, 31(24): 2872–2881, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2012:VSS

[867] Xiang Liu, Yingwei Peng, Dongsheng Tu, and Hua Liang. Variable selection in semiparametric cure models based on penalized likelihood, with application to breast cancer clinical trials. Statistics in Medicine, 31 (24):2882–2891, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tu:2012:MCD

[868] Yi-Hsuan Tu and Jason C. Hsu. Multiple comparisons of drug efficacy between subgroups defined by genetic polymorphisms. Statistics in Medicine, 31(24):2892–2903, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Julious:2012:TBS

[869] Steven A. Julious and Michael J. Campbell. Tutorial in biostatistics: sample sizes for parallel group clinical trials with binary data. Statistics in Medicine, 31(24):2904–2936, October 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Waclawiw:2012:P

[870] Myron A. Waclawiw, Colin O. Wu, and Song Yang. Preface. Statistics in Medicine, 31(25):2937, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zheng:2012:NCT

[871] Gang Zheng, Colin O. Wu, Song Yang, Myron A. Waclawiw, David L. DeMets, and Nancy L. Geller. NHLBI clinical trials workshop: an executive summary. Statistics in Medicine, 31(25):2938–2943, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeMets:2012:CDC

[872] David L. DeMets. Current development in clinical trials: issues old and new. Statistics in Medicine, 31(25):2944–2954, November 10, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2012:BCT

[873] J. Jack Lee and Caleb T. Chu. Bayesian clinical trials in action. Statistics in Medicine, 31(25):2955–2972, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fleming:2012:BSE

[874] Thomas R. Fleming and John H. Powers. Biomarkers and surrogate endpoints in clinical trials. Statistics in Medicine, 31(25):2973–2984, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:2012:IRR

[875] Douglas G. Altman, David Moher, and Kenneth F. Schulz. Improving the reporting of randomised trials: the CONSORT statement and beyond. Statistics in Medicine, 31(25):2985–2997, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2012:SDT

[876] Tao Liu and Daniel F. Heitjan. Sensitivity of the discrete-time Kaplan– Meier estimate to nonignorable censoring: Application in a clinical trial. Statistics in Medicine, 31(25):2998–3010, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2012:PAS

[877] Sue-Jane Wang, H. M. James Hung, and Robert O’Neill. Paradigms for adaptive statistical information designs: practical experiences and strategies. Statistics in Medicine, 31(25):3011–3023, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proschan:2012:SFC

[878] Michael A. Proschan and K. K. Gordon Lan. Spending functions and continuous-monitoring boundaries. Statistics in Medicine, 31(25):3024– 3030, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simon:2012:CTP

[879] Richard Simon. Clinical trials for predictive medicine. Statistics in Medicine, 31(25):3031–3040, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Witte:2012:RGV

[880] John S. Witte. Rare genetic variants and treatment response: sample size and analysis issues. Statistics in Medicine, 31(25):3041–3050, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lauer:2012:CHD

[881] Michael S. Lauer. Commentary: How the debate about comparative effectiveness research should impact the future of clinical trials. Statistics in Medicine, 31(25):3051–3053, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:2012:CHD

[882] Jonas H. Ellenberg. Commentary on ‘How the debate about comparative effectiveness research (CER) should impact the future of clinical trials’ by Michael S. Lauer. Statistics in Medicine, 31(25):3054–3056, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Basu:2012:PCC

[883] Anirban Basu. Patient-centered or ‘central’ patient: Raising the veil of ignorance over randomization. Statistics in Medicine, 31(25):3057–3059, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hernan:2012:CLH

[884] Miguel A. Hernán. Comments on Lauer’s ‘how the debate about comparative effectiveness research should impact the future of clinical trials’. Statistics in Medicine, 31(25):3060–3061, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kunz:2012:CER

[885] Lauren M. Kunz, Robert W. Yeh, and Sharon-Lise T. Normand. Comparative effectiveness research: does one size fit all? Statistics in Medicine, 31(25):3062–3065, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lauer:2012:RC

[886] Michael S. Lauer. Responses to commentaries. Statistics in Medicine, 31(25):3066–3067, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2012:FCT

[887] Ralph D’Agostino, Sr., David DeMets, William Friedewald, Steven Goodman, John Witte, and Nancy L. Geller. The future of clinical trials: a panel discussion. Statistics in Medicine, 31(25):3068–3072, November 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hampson:2012:IPF

[888] Lisa V. Hampson and Chris Metcalfe. Incorporating prognostic factors into causal estimators: a comparison of methods for randomised controlled trials with a time-to-event outcome. Statistics in Medicine, 31 (26):3073–3088, November 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peacock:2012:DCD

[889] J. L. Peacock, O. Sauzet, S. M. Ewings, and S. M. Kerry. Dichotomising continuous data while retaining statistical power using a distributional approach. Statistics in Medicine, 31(26):3089–3103, November 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Svensson:2012:DRA

[890] Elisabeth Svensson. Different ranking approaches defining association and agreement measures of paired ordinal data. Statistics in Medicine, 31(26):3104–3117, November 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2012:EME

[891] Wei Wang and Jeffrey M. Albert. Estimation of mediation effects for zero-inflated regression models. Statistics in Medicine, 31(26):3118–3132, November 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longford:2012:HLD

[892] Nicholas T. Longford. Handling the limit of detection by extrapolation. Statistics in Medicine, 31(26):3133–3146, November 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Macdonald-Wallis:2012:MMS

[893] Corrie Macdonald-Wallis, Debbie A. Lawlor, Tom Palmer, and Kate Tilling. Multivariate multilevel spline models for parallel growth processes: application to weight and mean arterial pressure in pregnancy. Statistics in Medicine, 31(26):3147–3164, November 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shi:2012:MEG

[894] J. Q. Shi, B. Wang, E. J. Will, and R. M. West. Mixed-effects Gaussian process functional regression models with application to dose-response curve prediction. Statistics in Medicine, 31(26):3165–3177, November 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laska:2012:MST

[895] Eugene Laska, Morris Meisner, and Joseph Wanderling. A maximally selected test of symmetry about zero. Statistics in Medicine, 31(26): 3178–3191, November 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2012:TLM

[896] Xue Li and Donald Hedeker. A three-level mixed-effects location scale model with an application to ecological momentary assessment data. Statistics in Medicine, 31(26):3192–3210, November 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2012:LVM

[897] Baojiang Chen and Xiao-Hua Zhou. A latent-variable marginal method for multi-level incomplete binary data. Statistics in Medicine, 31(26): 3211–3222, November 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crainiceanu:2012:BBI

[898] Ciprian M. Crainiceanu, Ana-Maria Staicu, Shubankar Ray, and Naresh Punjabi. Bootstrap-based inference on the difference in the means of two correlated functional processes. Statistics in Medicine, 31(26):3223– 3240, November 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Danaher:2012:EGE

[899] M. R. Danaher, E. F. Schisterman, A. Roy, and P. S. Albert. Estimation of gene-environment interaction by pooling biospecimens. Statistics in Medicine, 31(26):3241–3252, November 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hedeker:2012:ACL

[900] Donald Hedeker, Hakan Demirtas, and Robert D. Gibbons. Andrew C. Leon, Ph.D. (1951–2012). Statistics in Medicine, 31(27):3253–3254, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leon:2012:TPS

[901] Andrew C. Leon, Hakan Demirtas, Chunshan Li, and Donald Hedeker. Two propensity score-based strategies for a three-decade observational study: investigating psychotropic medications and suicide risk. Statistics in Medicine, 31(27):3255–3260, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bilder:2012:PTP

[902] Christopher R. Bilder and Joshua M. Tebbs. Pooled-testing procedures for screening high volume clinical specimens in heterogeneous populations. Statistics in Medicine, 31(27):3261–3268, November 30, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Caudill:2012:UPS

[903] Samuel P. Caudill. Use of pooled samples from the national health and nutrition examination survey. Statistics in Medicine, 31(27):3269–3277, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2012:BAD

[904] Michael P. Chen, Nong Shang, Carla A. Winston, and Jose E. Becerra. A Bayesian analysis of the 2009 decline in tuberculosis morbidity in the United States. Statistics in Medicine, 31(27):3278–3284, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Furlow-Parmley:2012:CET

[905] Carolyn Furlow-Parmley, James A. Singleton, Barbara Bardenheier, and Leah Bryan. Combining estimates from two surveys: an example from monitoring 2009 influenza A (H1N1) pandemic vaccination. Statistics in Medicine, 31(27):3285–3294, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Winston:2012:CIS

[906] Carla A. Winston, Andrew N. Hill, Michael P. Chen, Nong Shang, and José E. Becerra. Confidence intervals and statistical testing for ratio measures of percent change. Statistics in Medicine, 31(27):3295–3298, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miller:2012:SME

[907] Kelly M. Miller and Stephen W. Looney. A simple method for estimating the odds ratio in matched case-control studies with incomplete paired data. Statistics in Medicine, 31(27):3299–3312, November 30, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murphy:2012:DOM

[908] T. E. Murphy, G. McAvay, N. J. Carriero, C. P. Gross, M. E. Tinetti, H. G. Allore, and H. Lin. Deaths observed in Medicare beneficiaries: average attributable fraction and its longitudinal extension for many diseases. Statistics in Medicine, 31(27):3313–3319, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Griffin:2012:UAT

[909] Beth Ann Griffin, Garnet L. Anderson, Regina A. Shih, and Eric A. Whitsel. Use of alternative time scales in Cox proportional hazard models: implications for time-varying environmental exposures. Statistics in Medicine, 31(27):3320–3327, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hedeker:2012:MBS

[910] Donald Hedeker, Robin J. Mermelstein, and Hakan Demirtas. Modeling between-subject and within-subject variances in ecological momentary assessment data using mixed-effects location scale models. Statistics in Medicine, 31(27):3328–3336, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demirtas:2012:SMP

[911] Hakan Demirtas, Donald Hedeker, and Robin J. Mermelstein. Simulation of massive public health data by power polynomials. Statistics in Medicine, 31(27):3337–3346, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bar:2012:AHR

[912] Haim Y. Bar and Dean R. Lillard. Accounting for heaping in retrospectively reported event data — a mixture-model approach. Statistics in Medicine, 31(27):3347–3365, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2012:USM

[913] Duncan Lee. Using spline models to estimate the varying health risks from air pollution across Scotland. Statistics in Medicine, 31(27):3366– 3378, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dray:2012:IEM

[914] Alyssa Dray and Talithia Williams. An incidence estimation model for multi-stage diseases with differential mortality. Statistics in Medicine, 31(27):3379–3392, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ng:2012:ICP

[915] Shu Kay Ng, Libby Holden, and Jing Sun. Identifying comorbidity patterns of health conditions via cluster analysis of pairwise concordance statistics. Statistics in Medicine, 31(27):3393–3405, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Long:2012:CCR

[916] Dustin M. Long and Michael G. Hudgens. Comparing competing risk outcomes within principal strata, with application to studies of motherto-child transmission of HIV. Statistics in Medicine, 31(27):3406–3418, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walker:2012:RSM

[917] Robert Walker, Carolyn Nickson, Jie-Bin Lew, Megan Smith, and Karen Canfell. A revision of sexual mixing matrices in models of sexually transmitted infection. Statistics in Medicine, 31(27):3419–3432, November 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Little:2012:DCC

[918] R. J. Little, M. L. Cohen, K. Dickersin, S. S. Emerson, J. T. Farrar, J. D. Neaton, W. Shih, J. P. Siegel, and H. Stern. The design and conduct of clinical trials to limit missing data. Statistics in Medicine, 31(28):3433– 3443, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liao:2012:TMQ

[919] Kaijun Liao, Derek R. Freres, and Andrea B. Troxel. A transition model for quality-of-life data with non-ignorable non-monotone missing data. Statistics in Medicine, 31(28):3444–3466, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Batistatou:2012:PBC

[920] Evridiki Batistatou and Roseanne McNamee. Performance of biascorrection methods for exposure measurement error using repeated measurements with and without missing data. Statistics in Medicine, 31(28): 3467–3480, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldstein:2012:ARL

[921] Harvey Goldstein, Katie Harron, and Angie Wade. The analysis of record-linked data using multiple imputation with data value priors. Statistics in Medicine, 31(28):3481–3493, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Atkinson:2012:BLT

[922] Anthony C. Atkinson. Bias and loss: the two sides of a biased coin. Statistics in Medicine, 31(28):3494–3503, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boyd:2012:ETE

[923] Adam P. Boyd, John M. Kittelson, and Daniel L. Gillen. Estimation of treatment effect under non-proportional hazards and conditionally independent censoring. Statistics in Medicine, 31(28):3504–3515, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saramago:2012:MTC

[924] Pedro Saramago, Alex J. Sutton, Nicola J. Cooper, and Andrea Manca. Mixed treatment comparisons using aggregate and individual participant level data. Statistics in Medicine, 31(28):3516–3536, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lebowitsch:2012:GMD

[925] Jonathan Lebowitsch, Yan Ge, Benjamin Young, and Feifang Hu. Generalized multidimensional dynamic allocation method. Statistics in Medicine, 31(28):3537–3544, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2012:MCR

[926] Longyang Wu and Richard J. Cook. Misspecification of Cox regression models with composite endpoints. Statistics in Medicine, 31(28):3545– 3562, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilding:2012:NLM

[927] Gregory E. Wilding, Rameela Chandrasekhar, and Alan D. Hutson. A new linear model-based approach for inferences about the mean area under the curve. Statistics in Medicine, 31(28):3563–3578, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stucke:2012:GAS

[928] Kathrin Stucke and Meinhard Kieser. A general approach for sample size calculation for the three-arm ‘gold standard’ non-inferiority design. Statistics in Medicine, 31(28):3579–3596, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2012:MAM

[929] Ming-Hui Chen, Joseph G. Ibrahim, Arvind K. Shah, Jianxin Lin, and Hui Yao. Meta-analysis methods and models with applications in evaluation of cholesterol-lowering drugs. Statistics in Medicine, 31(28):3597– 3616, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2012:PSM

[930] Shi Li, Bhramar Mukherjee, and Stuart Batterman. Point source modeling of matched case-control data with multiple disease subtypes. Statistics in Medicine, 31(28):3617–3637, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2012:HTC

[931] Liansheng Larry Tang, Aiyi Liu, Enrique F. Schisterman, Xiao-Hua Zhou, and Catherine Chun ling Liu. Homogeneity tests of clustered diagnostic markers with applications to the BioCycle study. Statistics in Medicine, 31(28):3638–3648, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heller:2012:MAB

[932] Gillian Z. Heller and Lindsay C. Dunlop. A modelling approach for blood units transfused after stem cell transplantation. Statistics in Medicine, 31(28):3649–3655, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Solari:2012:TGF

[933] Aldo Solari, Saskia le Cessie, and Jelle J. Goeman. Testing goodness of fit in regression: a general approach for specified alternatives. Statistics in Medicine, 31(28):3656–3666, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vanbelle:2012:HMA

[934] Sophie Vanbelle, Timothy Mutsvari, Dominique Declerck, and Emmanuel Lesaffre. Hierarchical modeling of agreement. Statistics in Medicine, 31(28):3667–3680, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Frederiksen:2012:EAM

[935] Kirsten Frederiksen, Isabelle Deltour, and Joachim Schüz. Estimating associations of mobile phone use and brain tumours taking into account laterality: a comparison and theoretical evaluation of applied methods. Statistics in Medicine, 31(28):3681–3692, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2012:BME

[936] Wei Chen, Debashis Ghosh, Trivellore E. Raghunathan, Maxim Norkin, Daniel J. Sargent, and Gerold Bepler. On Bayesian methods of exploring qualitative interactions for targeted treatment. Statistics in Medicine, 31 (28):3693–3707, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheon:2012:IRE

[937] Kyeongmi Cheon, Paul S. Albert, and Zhiwei Zhang. The impact of random-effect misspecification on percentile estimation for longitudinal growth data. Statistics in Medicine, 31(28):3708–3718, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fan:2012:SBD

[938] Shenghua Kelly Fan, Ying Lu, and You-Gan Wang. A simple Bayesian decision-theoretic design for dose-finding trials. Statistics in Medicine, 31 (28):3719–3730, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Valberg:2012:FMA

[939] Morten Valberg, Tom Grotmol, Steinar Tretli, Marit B. Veierød, Susan S. Devesa, and Odd O. Aalen. Frailty modeling of age-incidence curves of osteosarcoma and Ewing sarcoma among individuals younger than 40 years. Statistics in Medicine, 31(28):3731–3747, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2012:CCD

[940] Yunda Huang, Ying Huang, Zoe Moodie, Sue Li, and Steve Self. Comparing and combining data across multiple sources via integration of pairedsample data to correct for measurement error. Statistics in Medicine, 31 (28):3748–3759, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fay:2012:WLT

[941] Michael P. Fay and Joanna H. Shih. Weighted logrank tests for interval censored data when assessment times depend on treatment. Statistics in Medicine, 31(28):3760–3772, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Conde:2012:CST

[942] D. Conde, M. A. Fernández, C. Rueda, and B. Salvador. Classification of samples into two or more ordered populations with application to a cancer trial. Statistics in Medicine, 31(28):3773–3786, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanGeloven:2012:CPD

[943] Nan van Geloven, Kimiko A. Broeze, Brent C. Opmeer, Ben Willem Mol, and Aeilko H. Zwinderman. Correction: PHow to deal with double partial verification when evaluating two index tests in relation to a reference test? Statistics in Medicine, 31(28):3787–3788, December 10, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yuan:2012:VSC

[944] Shuai Yuan, Hao Helen Zhang, and Marie Davidian. Variable selection for covariate-adjusted semiparametric inference in randomized clinical trials.

Statistics in Medicine, 31(29):3789–3804, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2012:QIB

[945] Dan Jackson, Ian R. White, and Richard D. Riley. Quantifying the impact of between-study heterogeneity in multivariate meta-analyses. Statistics in Medicine, 31(29):3805–3820, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gasparrini:2012:MMA

[946] A. Gasparrini, B. Armstrong, and M. G. Kenward. Multivariate metaanalysis for non-linear and other multi-parameter associations. Statistics in Medicine, 31(29):3821–3839, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donegan:2012:ACA

[947] Sarah Donegan, Paula Williamson, Umberto D’Alessandro, and Catrin Tudur Smith. Assessing the consistency assumption by exploring treatment by covariate interactions in mixed treatment comparison metaanalysis: individual patient-level covariates versus aggregate trial-level covariates. Statistics in Medicine, 31(29):3840–3857, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaiser:2012:DRR

[948] Lee D. Kaiser. Dynamic randomization and a randomization model for clinical trials data. Statistics in Medicine, 31(29):3858–3873, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bowden:2012:ICD

[949] Jack Bowden and James Wason. Identifying combined design and analysis procedures in two-stage trials with a binary end point. Statistics in Medicine, 31(29):3874–3884, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhong:2012:TCR

[950] Wei Zhong, Joseph S. Koopmeiners, and Bradley P. Carlin. A trivariate continual reassessment method for phase I/II trials of toxicity, efficacy, and surrogate efficacy. Statistics in Medicine, 31(29):3885–3895, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2012:CLA

[951] Ming Wang, W. Dana Flanders, Roberd M. Bostick, and Qi Long. A conditional likelihood approach for regression analysis using biomarkers measured with batch-specific error. Statistics in Medicine, 31(29):3896– 3906, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Delzell:2012:KPO

[952] Darcie A. P. Delzell, Richard F. Gunst, William R. Schucany, Patrick S. Carmack, Qihua Lin, Jeffrey S. Spence, and Robert W. Haley. Key properties of D-optimal designs for event-related functional MRI experiments with application to nonlinear models. Statistics in Medicine, 31 (29):3907–3920, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gerds:2012:ARR

[953] Thomas A. Gerds, Thomas H. Scheike, and Per K. Andersen. Absolute risk regression for competing risks: interpretation, link functions, and prediction. Statistics in Medicine, 31(29):3921–3930, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Piccorelli:2012:JMR

[954] Annalisa V. Piccorelli and Mark D. Schluchter. Jointly modeling the relationship between longitudinal and survival data subject to left truncation with applications to cystic fibrosis. Statistics in Medicine, 31(29): 3931–3945, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2012:GST

[955] Peter C. Austin. Generating survival times to simulate Cox proportional hazards models with time-varying covariates. Statistics in Medicine, 31 (29):3946–3958, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2012:SSC

[956] Songfeng Wang, Jiajia Zhang, and Wenbin Lu. Sample size calculation for the proportional hazards cure model. Statistics in Medicine, 32(6): 3959–3971, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2012:SSF

[957] G. Y. Zou. Sample size formulas for estimating intraclass correlation coefficients with precision and assurance. Statistics in Medicine, 32(6): 3972–3981, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saha:2012:PLB

[958] Krishna K. Saha. Profile likelihood-based confidence interval of the intraclass correlation for binary outcome data sampled from clusters. Statistics in Medicine, 32(6):3982–4002, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Westgate:2012:BCC

[959] Philip M. Westgate. A bias-corrected covariance estimate for improved inference with quadratic inference functions. Statistics in Medicine, 32 (6):4003–4022, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2012:ESP

[960] Ying Xu, Yin Bun Cheung, K. F. Lam, and Paul Milligan. Estimation of summary protective efficacy using a frailty mixture model for recurrent event time data. Statistics in Medicine, 32(6):4023–4039, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2012:SCD

[961] Inkyung Jung and Hana Lee. Spatial cluster detection for ordinal outcome data. Statistics in Medicine, 32(6):4040–4048, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Das:2012:RAJ

[962] Kalyan Das and Arindom Chakraborty. Robust analysis in joint models: an application to a study on muscular dystrophy. Statistics in Medicine, 32(6):4049–4060, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choudhury:2012:CLS

[963] Kingshuk Roy Choudhury, Finbarr O’Sullivan, Ian Kasman, and Greg D. Plowman. A comparison of least squares and conditional maximum likelihood estimators under volume endpoint censoring in tumor growth experiments. Statistics in Medicine, 32(6):4061–4073, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buu:2012:SML

[964] Anne Buu, Runze Li, Xianming Tan, and Robert A. Zucker. Statistical models for longitudinal zero-inflated count data with applications to the substance abuse field. Statistics in Medicine, 32(6):4074–4086, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2012:PLT

[965] Xian Liu and Charles C. Engel. Predicting longitudinal trajectories of health probabilities with random-effects multinomial logit regression. Statistics in Medicine, 32(6):4087–4101, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2012:ASR

[966] Bin Zhu, David B. Dunson, and Allison E. Ashley-Koch. Adverse subpopulation regression for multivariate outcomes with high-dimensional predictors. Statistics in Medicine, 32(6):4102–4113, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Furlan:2012:PMF

[967] Claudia Furlan and Cinzia Mortarino. Pleural mesothelioma: forecasts of the death toll in the area of Casale Monferrato, Italy. Statistics in Medicine, 32(6):4114–4134, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palys:2012:UAA

[968] Kaitlin E. Palys, Vance W. Berger, and William C. Grant. Unequal allocation and allocation concealment. Statistics in Medicine, 32(6):4135– 4136, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuznetsova:2012:RUA

[969] Olga M. Kuznetsova and Yevgen Tymofyeyev. Response to ‘Unequal allocation and allocation concealment’ by Kaitlin E. Palys, Vance W. Berger, and William Grant. Statistics in Medicine, 32(6):4137, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bacchetti:2012:BUD

[970] Peter Bacchetti, Charles McCulloch, and Mark R. Segal. Being ‘underpowered’ does not make a study unethical. Statistics in Medicine, 32 (6):4138–4139, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gelfond:2012:PEO

[971] Jonathan Adam Gelfond, Elizabeth Heitman, Brad H. Pollock, and Craig H. Klugman. Power, ethics, and obligation. Statistics in Medicine, 32(6):4140–4141, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berger:2012:TSR

[972] Vance W. Berger. Two-stage randomized trials: outstanding issues. Statistics in Medicine, 32(6):4142, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tamura:2012:AR

[973] Roy N. Tamura, Xiaohong Huang, and Dennis Boos. Authors’ reply. Statistics in Medicine, 32(6):4143–4144, December 20, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2012:P

[974] KyungMann Kim and Simon G. Thompson. Preface. Statistics in Medicine, 31(30):4145, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Posch:2012:UAB

[975] Martin Posch and Michael A. Proschan. Unplanned adaptations before breaking the blind. Statistics in Medicine, 31(30):4146–4153, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stirnemann:2012:DEB

[976] J. J. Stirnemann, F. Comte, and A. Samson. Density estimation of a biomedical variable subject to measurement error using an auxiliary set of replicate observations. Statistics in Medicine, 31(30):4154–4163, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2012:CMI

[977] Katherine J. Lee, John C. Galati, Julie A. Simpson, and John B. Carlin. Comparison of methods for imputing ordinal data using multivariate normal imputation: a case study of non-linear effects in a large cohort study. Statistics in Medicine, 31(30):4164–4174, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pu:2012:AHT

[978] Minya Pu, Tomoko Hayashi, Howard Cottam, Joseph Mulvaney, Michelle Arkin, Maripat Corr, Dennis Carson, and Karen Messer. Analysis of high-throughput screening assays using cluster enrichment. Statistics in Medicine, 31(30):4175–4189, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Havercroft:2012:SMS

[979] W. G. Havercroft and V. Didelez. Simulating from marginal structural models with time-dependent confounding. Statistics in Medicine, 31(30): 4190–4206, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:2012:FSS

[980] Toshiro Tango and Kunihiko Takahashi. A flexible spatial scan statistic with a restricted likelihood ratio for detecting disease clusters. Statistics in Medicine, 31(30):4207–4218, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Graffeo:2012:IAL

[981] Nathalie Grafféo, Valérie Jooste, and Roch Giorgi. The impact of additional life-table variables on excess mortality estimates. Statistics in Medicine, 31(30):4219–4230, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hof:2012:MAD

[982] M. H. P. Hof and A. H. Zwinderman. Methods for analyzing data from probabilistic linkage strategies based on partially identifying variables. Statistics in Medicine, 31(30):4231–4242, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teramukai:2012:BPS

[983] Satoshi Teramukai, Takashi Daimon, and Sarah Zohar. A Bayesian predictive sample size selection design for single-arm exploratory clinical trials. Statistics in Medicine, 31(30):4243–4254, December 30, 2012. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2012:CTS

[984] Min Zhang and Douglas E. Schaubel. Contrasting treatment-specific survival using double-robust estimators. Statistics in Medicine, 31(30): 4255–4268, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wason:2012:ODM

[985] James M. S. Wason and Thomas Jaki. Optimal design of multi-arm multi-stage trials. Statistics in Medicine, 31(30):4269–4279, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2012:GDQ

[986] Stephen Senn, Nick Holford, and Hans Hockey. The ghosts of departed quantities: approaches to dealing with observations below the limit of quantitation. Statistics in Medicine, 31(30):4280–4295, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barrett:2012:TSM

[987] Jessica K. Barrett, Vern T. Farewell, Fotios Siannis, Jayne Tierney, and Julian P. T. Higgins. Two-stage meta-analysis of survival data from individual participants using percentile ratios. Statistics in Medicine, 31 (30):4296–4308, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friede:2012:CEF

[988] T. Friede, N. Parsons, and N. Stallard. A conditional error function approach for subgroup selection in adaptive clinical trials. Statistics in Medicine, 31(30):4309–4320, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Iasonos:2012:IPS

[989] Alexia Iasonos and John O’Quigley. Interplay of priors and skeletons in two-stage continual reassessment method. Statistics in Medicine, 32(14): 4321–4336, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hee:2012:DSD

[990] Siew Wan Hee and Nigel Stallard. Designing a series of decision-theoretic phase II trials in a small population. Statistics in Medicine, 32(14):4337– 4351, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kunz:2012:ESE

[991] Cornelia Ursula Kunz and Meinhard Kieser. Estimation of secondary endpoints in two-stage phase II oncology trials. Statistics in Medicine, 32(14):4352–4368, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karlsson:2012:MBP

[992] Magdalena Kauczynska Karlsson, Anders Lönneborg, and Solve Sæbø. Microarray-based prediction of Parkinson’s disease using clinical data as additional response variables. Statistics in Medicine, 32(14):4369–4381, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williamson:2012:DRE

[993] E. J. Williamson, A. Forbes, and R. Wolfe. Doubly robust estimators of causal exposure effects with missing data in the outcome, exposure or a confounder. Statistics in Medicine, 32(14):4382–4400, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ryan:2012:EAM

[994] Patrick B. Ryan, David Madigan, Paul E. Stang, J. Marc Overhage, Judith A. Racoosin, and Abraham G. Hartzema. Empirical assessment of methods for risk identification in healthcare data: results from the experiments of the observational medical outcomes partnership. Statistics in Medicine, 32(14):4401–4415, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rodriguez-Girondo:2012:NTM

[995] Mar Rodríguez-Girondo and Jacobo de Uña-Álvarez. A nonparametric test for Markovianity in the illness–death model. Statistics in Medicine, 32(14):4416–4427, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Allodji:2012:PFM

[996] Rodrigue S. Allodji, Anne C. M. Thiébaut, Klervi Leuraud, Estelle Rage, Stéphane Henry, Dominique Laurier, and Jacques Bénichou. The performance of functional methods for correcting non-Gaussian measurement error within Poisson regression: corrected excess risk of lung cancer mortality in relation to radon exposure among French uranium miners. Statistics in Medicine, 32(14):4428–4443, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Savignoni:2012:EEO

[997] Alexia Savignoni, David Hajage, Pascale Tubert-Bitter, and Yann De Rycke. Effect of an event occurring over time and confounded by health status: estimation and interpretation. a study based on survival data simulations with application on breast cancer. Statistics in Medicine, 32 (14):4444–4455, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crowther:2012:FPJ

[998] Michael J. Crowther, Keith R. Abrams, and Paul C. Lambert. Flexible parametric joint modelling of longitudinal and survival data. Statistics in Medicine, 32(14):4456–4471, December 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nietert:2013:CRB

[999] Paul J. Nietert, Amy E. Wahlquist, and Teri Lynn Herbert. Characteristics of recent biostatistical methods adopted by researchers publishing in general/internal medicine journals. Statistics in Medicine, 32(1):1–10, January 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roloff:2013:PFS

[1000] Verena Roloff, Julian P. T. Higgins, and Alex J. Sutton. Planning future studies based on the conditional power of a meta-analysis. Statistics in Medicine, 32(1):11–24, January 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giovane:2013:NMA

[1001] Cinzia Del Giovane, Laura Vacchi, Dimitris Mavridis, Graziella Filippini, and Georgia Salanti. Network meta-analysis models to account for variability in treatment definitions: application to dose effects. Statistics in Medicine, 32(1):25–39, January 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guolo:2013:FMB

[1002] A. Guolo. Flexibly modeling the baseline risk in meta-analysis. Statistics in Medicine, 32(1):40–50, January 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mavridis:2013:FBA

[1003] Dimitris Mavridis, Alex Sutton, Andrea Cipriani, and Georgia Salanti. A fully Bayesian application of the Copas selection model for publication bias extended to network meta-analysis. Statistics in Medicine, 32(1): 51–66, January 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paul:2013:SPH

[1004] Prabasaj Paul, Michael L. Pennell, and Stanley Lemeshow. Standardizing the power of the Hosmer–Lemeshow goodness of fit test in large data sets. Statistics in Medicine, 32(1):67–80, January 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roberts:2013:DAN

[1005] Chris Roberts and Rebecca Walwyn. Design and analysis of nonpharmacological treatment trials with multiple therapists per patient. Statistics in Medicine, 32(1):81–98, January 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sperrin:2013:MTE

[1006] Matthew Sperrin and Iain Buchan. Modelling time to event with observations made at arbitrary times. Statistics in Medicine, 32(1):99–109, January 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lumley:2013:PLR

[1007] Thomas Lumley and Alastair Scott. Partial likelihood ratio tests for the Cox model under complex sampling. Statistics in Medicine, 32(1): 110–123, January 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Darlington:2013:EWP

[1008] G. A. Darlington and S. N. Dixon. Event-weighted proportional hazards modelling for recurrent gap time data. Statistics in Medicine, 32(1): 124–130, January 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wages:2013:UTE

[1009] Nolan A. Wages, Mark R. Conaway, and John O’Quigley. Using the time-to-event continual reassessment method in the presence of partial orders. Statistics in Medicine, 32(1):131–141, January 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pietzner:2013:TRP

[1010] Diana Pietzner and Andreas Wienke. The trend-renewal process: a useful model for medical recurrence data. Statistics in Medicine, 32(1):142–152, January 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2013:NFD

[1011] Yingchun Zhou and Nell Sedransk. A new functional data-based biomarker for monitoring cardiovascular behavior. Statistics in Medicine, 32(1):153–164, January 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sikorska:2013:FLM

[1012] Karolina Sikorska, Fernando Rivadeneira, Patrick J. F. Groenen, Albert Hofman, André G. Uitterlinden, Paul H. C. Eilers, and Emmanuel Lesaffre. Fast linear mixed model computations for genome-wide association studies with longitudinal data. Statistics in Medicine, 32(1):165–180, January 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2013:EAA

[1013] Haihong Li, Abdul J. Sankoh, and Ralph B. D’Agostino, Sr. Extension of adaptive alpha allocation methods for strong control of the familywise error rate. Statistics in Medicine, 32(2):181–195, January 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shaw:2013:NVC

[1014] Pamela A. Shaw and Michael A. Proschan. Null but not void: considerations for hypothesis testing. Statistics in Medicine, 32(2):196–205, January 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hanin:2013:OSS

[1015] Leonid Hanin and Lyudmila Pavlova. Optimal screening schedules for prevention of metastatic cancer. Statistics in Medicine, 32(2):206–219, January 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachin:2013:PMH

[1016] John M. Lachin. Power of the Mantel–Haenszel and other tests for discrete or grouped time-to-event data under a chained binomial model. Statistics in Medicine, 32(2):220–229, January 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:2013:LTA

[1017] Changyong Feng, Hongyue Wang, Naiji Lu, and Xin M. Tu. Log transformation: application and interpretation in biomedical research. Statistics in Medicine, 32(2):230–239, January 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fu:2013:JMP

[1018] Haoda Fu, Yanping Wang, Jingyi Liu, Pandurang M. Kulkarni, and Allen S. Melemed. Joint modeling of progression-free survival and overall survival by a Bayesian normal induced copula estimation model. Statistics in Medicine, 32(2):240–254, January 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pashova:2013:BDG

[1019] H. Pashova, M. LeBlanc, and C. Kooperberg. Boosting for detection of gene-environment interactions. Statistics in Medicine, 32(2):255–266, January 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Musal:2013:BSM

[1020] Muzaffer Musal and Tevfik Aktekin. Bayesian spatial modeling of HIV mortality via zero-inflated Poisson models. Statistics in Medicine, 32(2): 267–281, January 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dalton:2013:FRB

[1021] Jarrod E. Dalton. Flexible recalibration of binary clinical prediction models. Statistics in Medicine, 32(2):282–289, January 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nathoo:2013:SES

[1022] Farouk S. Nathoo and Pulak Ghosh. Skew-elliptical spatial random effect modeling for areal data with application to mapping health utilization rates. Statistics in Medicine, 32(2):290–306, January 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jaynes:2013:AFF

[1023] Jessica Jaynes, Xianting Ding, Hongquan Xu, Weng Kee Wong, and Chih-Ming Ho. Application of fractional factorial designs to study drug combinations. Statistics in Medicine, 32(2):307–318, January 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2013:SMV

[1024] Yangxin Huang. Segmental modeling of viral load changes for HIV longitudinal data with skewness and detection limits. Statistics in Medicine, 32(2):319–334, January 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2013:PLE

[1025] Jinsong Chen, Lei Liu, Bankole A. Johnson, and John O’Quigley. Penalized likelihood estimation for semiparametric mixed models, with application to alcohol treatment research. Statistics in Medicine, 32(2): 335–346, January 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Landsman:2013:EAC

[1026] V. Landsman and B. I. Graubard. Efficient analysis of case-control studies with sample weights. Statistics in Medicine, 32(2):347–360, January 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:2013:ABI

[1027] Shein-Chung Chow. Assessing biosimilarity and interchangeability of biosimilar products. Statistics in Medicine, 32(3):361–363, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:2013:CFD

[1028] Shein-Chung Chow, Laszlo Endrenyi, and Peter A. Lachenbruch. Comments on the FDA draft guidance on biosimilar products. Statistics in Medicine, 32(3):364–369, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:2013:SCA

[1029] Shein-Chung Chow, Jun Wang, Laszlo Endrenyi, and Peter A. Lachenbruch. Scientific considerations for assessing biosimilar products. Statistics in Medicine, 32(3):370–381, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2013:SAB

[1030] Seung-Ho Kang and Shein-Chung Chow. Statistical assessment of biosimilarity based on relative distance between follow-on biologics. Statistics in Medicine, 32(3):382–392, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2013:SCB

[1031] Yulan Li, Qing Liu, Patricia Wood, and Anandhi Johri. Statistical considerations in biosimilar clinical efficacy trials with asymmetrical margins. Statistics in Medicine, 32(3):393–405, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsieh:2013:EBI

[1032] Tsung-Cheng Hsieh, Shein-Chung Chow, Lan-Yan Yang, and Eric Chi. The evaluation of biosimilarity index based on reproducibility probability for assessing follow-on biologics. Statistics in Medicine, 32(3):406–414, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2013:ATH

[1033] Jun Yang, Nan Zhang, Shein-Chung Chow, and Eric Chi. An adapted F - test for homogeneity of variability in follow-on biological products. Statistics in Medicine, 32(3):415–423, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2013:IVC

[1034] Nan Zhang, Jun Yang, Shein-Chung Chow, Laszlo Endrenyi, and Eric Chi. Impact of variability on the choice of biosimilarity limits in assessing follow-on biologics. Statistics in Medicine, 32(3):424–433, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Endrenyi:2013:IBD

[1035] Laszlo Endrenyi, Chiann Chang, Shein-Chung Chow, and Laszlo Tothfalusi. On the interchangeability of biologic drug products. Statistics in Medicine, 32(3):434–441, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:2013:SMA

[1036] Shein-Chung Chow, Lan-Yan Yang, Aijing Starr, and Shih-Ting Chiu. Statistical methods for assessing interchangeability of biosimilars. Statistics in Medicine, 32(3):442–448, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2013:APL

[1037] Jr-Rung Lin, Shein-Chung Chow, Chih-Hsi Chang, Ya-Ching Lin, and Jen pei Liu. Application of the parallel line assay to assessment of biosimilar products based on binary endpoints. Statistics in Medicine, 32(18): 449–461, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liao:2013:CCQ

[1038] Jason J. Z. Liao and Patrick F. Darken. Comparability of critical quality attributes for establishing biosimilarity. Statistics in Medicine, 32(18): 462–469, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ray:2013:FBP

[1039] Herman E. Ray and Shesh N. Rai. Flexible bivariate phase II clinical trial design incorporating toxicity and response on different schedules. Statistics in Medicine, 32(18):470–485, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kordzakhia:2013:SPC

[1040] George Kordzakhia and Alex Dmitrienko. Superchain procedures in clinical trials with multiple objectives. Statistics in Medicine, 32(18):486– 508, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Das:2013:DSB

[1041] Kiranmoy Das, Jiahan Li, Guifang Fu, Zhong Wang, Runze Li, and Rongling Wu. Dynamic semiparametric Bayesian models for genetic mapping of complex trait with irregular longitudinal data. Statistics in Medicine, 32(18):509–523, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Szymczak:2013:ALR

[1042] Silke Szymczak, Markus O. Scheinhardt, Tanja Zeller, Philipp S. Wild, Stefan Blankenberg, and Andreas Ziegler. Adaptive linear rank tests for eQTL studies. Statistics in Medicine, 32(18):524–537, February 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OMalley:2013:ASN

[1043] A. James O’Malley. The analysis of social network data: an exciting frontier for statisticians. Statistics in Medicine, 32(4):539–555, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Christakis:2013:SCT

[1044] Nicholas A. Christakis and James H. Fowler. Social contagion theory: examining dynamic social networks and human behavior. Statistics in Medicine, 32(4):556–577, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wasserman:2013:CSC

[1045] Stanley Wasserman. Comments on ‘Social contagion theory: examining dynamic social networks and human behavior’ by Nicholas Christakis and James Fowler. Statistics in Medicine, 32(4):578–580, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomas:2013:SCH

[1046] A. C. Thomas. The social contagion hypothesis: comment on ‘social contagion theory: examining dynamic social networks and human behavior’. Statistics in Medicine, 32(4):581–590, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanderWeele:2013:IIM

[1047] Tyler J. VanderWeele. Inference for influence over multiple degrees of separation on a social network. Statistics in Medicine, 32(4):591–596, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Christakis:2013:RCS

[1048] Nicholas A. Christakis and James H. Fowler. Rejoinder to commentaries on social contagion theory. Statistics in Medicine, 32(4):597–599, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OKeeffe:2013:MDM

[1049] Aidan G. O’Keeffe, Brian D. M. Tom, and Vernon T. Farewell. Mixture distributions in multi-state modelling: Some considerations in a study of psoriatic arthritis. Statistics in Medicine, 32(4):600–619, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2013:GFM

[1050] Liansheng Tang and Xiao-Hua Zhou. A general framework of marker design with optimal allocation to assess clinical utility. Statistics in Medicine, 32(4):620–630, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2013:LCB

[1051] Le Kang, Chengjie Xiong, Paul Crane, and Lili Tian. Linear combinations of biomarkers to improve diagnostic accuracy with three ordinal diagnostic categories. Statistics in Medicine, 32(4):631–643, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yin:2013:TSD

[1052] Guosheng Yin, Shurong Zheng, and Jiajing Xu. Two-stage dose finding for cytostatic agents in phase I oncology trials. Statistics in Medicine, 32(4):644–660, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2013:PAR

[1053] Peter C. Austin and Ewout W. Steyerberg. Predictive accuracy of risk factors and markers: a simulation study of the effect of novel markers on different performance measures for logistic regression models. Statistics in Medicine, 32(4):661–672, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zheng:2013:DRT

[1054] Hao W. Zheng, Babette A. Brumback, Xiaomin Lu, Erin D. Bouldin, Michael B. Cannell, and Elena M. Andresen. Doubly robust testing and estimation of model-adjusted effect-measure modification with complex survey data. Statistics in Medicine, 32(4):673–684, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAngelo:2013:ACI

[1055] Gina M. D’Angelo and Lisa A. Weissfeld. Application of copulas to improve covariance estimation for partial least squares. Statistics in Medicine, 32(4):685–696, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kapetanakis:2013:SMM

[1056] Venediktos Kapetanakis, Fiona E. Matthews, and Ardo van den Hout. A semi-Markov model for stroke with piecewise-constant hazards in the presence of left, right and interval censoring. Statistics in Medicine, 32 (4):697–713, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2013:CDF

[1057] Anonymous. Comments on ‘A distribution-free test of constant mean in linear mixed effects models’. Statistics in Medicine, 32(4):714–716, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lim:2013:RLE

[1058] Johan Lim and Xinlei Wang. Response to letter to the Editor by Dr. Vossoughi. Statistics in Medicine, 32(4):717, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2013:RSD

[1059] Stuart G. Baker. Remarks on ‘A simple decision analytic solution to the comparison of two binary diagnostic tests’ by Vickers et al. Statistics in Medicine, 32(4):718, February 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gomez:2013:SCW

[1060] Guadalupe Gómez and Stephen W. Lagakos. Statistical considerations when using a composite endpoint for comparing treatment groups. Statistics in Medicine, 32(5):719–738, February 28, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jahn-Eimermacher:2013:SSC

[1061] Antje Jahn-Eimermacher, Katharina Ingel, and Astrid Schneider. Sample size in cluster-randomized trials with time to event as the primary endpoint. Statistics in Medicine, 32(5):739–751, February 28, 2013. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Achana:2013:EMI

[1062] Felix A. Achana, Nicola J. Cooper, Sofia Dias, Guobing Lu, Stephen J. C. Rice, Denise Kendrick, and Alex J. Sutton. Extending methods for investigating the relationship between treatment effect and baseline risk from pairwise meta-analysis to network meta-analysis. Statistics in Medicine, 32(5):752–771, February 28, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mohammed:2013:DCC

[1063] Sandra M. Mohammed, Lorien S. Dalrymple, Damla Sentürk, and Danh V. Nguyen. Design considerations for case series models with exposure onset measurement error. Statistics in Medicine, 32(5):772–786, February 28, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gotte:2013:SSP

[1064] Heiko Götte and Isabella Zwiener. Sample size planning for survival prediction with focus on high-dimensional data. Statistics in Medicine, 32(5):787–807, February 28, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kovalchik:2013:GBR

[1065] Stephanie A. Kovalchik, Ravi Varadhan, Barbara Fetterman, Nancy E. Poitras, Sholom Wacholder, and Hormuzd A. Katki. A general binomial regression model to estimate standardized risk differences from binary response data. Statistics in Medicine, 32(5):808–821, February 28, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kor:2013:MAC

[1066] Chew-Teng Kor, Kuang-Fu Cheng, and Yi-Hau Chen. A method for analyzing clustered interval-censored data based on Cox’s model. Statistics in Medicine, 32(5):822–832, February 28, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2013:EMM

[1067] Haocheng Li and Grace Y. Yi. Estimation methods for marginal and association parameters for longitudinal binary data with nonignorable missing observations. Statistics in Medicine, 32(5):833–848, February 28, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longford:2013:SAD

[1068] Nicholas T. Longford. Screening as an application of decision theory. Statistics in Medicine, 32(5):849–863, February 28, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lowe:2013:DEW

[1069] Rachel Lowe, Trevor C. Bailey, David B. Stephenson, Tim E. Jupp, Richard J. Graham, Christovam Barcellos, and Marilia Sá Carvalho. The development of an early warning system for climate-sensitive disease risk with a focus on dengue epidemics in southeast Brazil. Statistics in Medicine, 32(5):864–883, February 28, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proschan:2013:HME

[1070] Michael Proschan, Charles E. Ford, Jeffrey A. Cutler, James F. Graumlich, Valory Pavlik, William C. Cushman, Barry R. Davis, Michael H. Alderman, David Gordon, Curt D. Furberg, Stanley S. Franklin, Samuel S. Blumenthal, Richard S. Castaldo, Richard A. Preston, and Allhat Collaborative Research Group. How much effect of different antihypertensive medications on cardiovascular outcomes is attributable to their effects on blood pressure? Statistics in Medicine, 32(5):884–897, February 28, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reinikainen:2013:CGS

[1071] Jaakko Reinikainen and Juha Karvanen. Comment on ‘generating survival times to simulate Cox proportional hazards models with timevarying covariates’. Statistics in Medicine, 32(5):898, February 28, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Logan:2013:UGS

[1072] Brent R. Logan and Mei-Jie Zhang. The use of group sequential designs with common competing risks tests. Statistics in Medicine, 32(6):899– 913, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donegan:2013:CIP

[1073] Sarah Donegan, Paula Williamson, Umberto D’Alessandro, Paul Garner, and Catrin Tudur Smith. Combining individual patient data and aggregate data in mixed treatment comparison meta-analysis: Individual patient data may be beneficial if only for a subset of trials. Statistics in Medicine, 32(6):914–930, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2013:SAF

[1074] Dan Jackson, Rose Baker, and Jack Bowden. A sensitivity analysis framework for the treatment effect measure used in the meta-analysis of comparative binary data from randomised controlled trials. Statistics in Medicine, 32(6):931–940, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nunney:2013:ETE

[1075] Ian Nunney, Allan Clark, and Lee Shepstone. Estimating treatment effects in a two-arm parallel trial of a continuous outcome. Statistics in Medicine, 32(6):941–955, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2013:IBC

[1076] Zhiwu Yan. The impact of baseline covariates on the efficiency of statistical analyses of crossover designs. Statistics in Medicine, 32(6):956–963, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kosinski:2013:WGS

[1077] Andrzej S. Kosinski. A weighted generalized score statistic for comparison of predictive values of diagnostic tests. Statistics in Medicine, 32(6): 964–977, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stock:2013:RPM

[1078] Shannon Stock and Victor DeGruttola. Recursive partitioning for monotone missing at random longitudinal markers. Statistics in Medicine, 32(6):978–994, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nakas:2013:GYI

[1079] Christos T. Nakas, John C. Dalrymple-Alford, Tim J. Anderson, and Todd A. Alonzo. Generalization of Youden index for multiple-class classification problems applied to the assessment of externally validated cognition in Parkinson disease screening. Statistics in Medicine, 32(6):995– 1003, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2013:RIC

[1080] Hua Shen and Richard J. Cook. Regression with incomplete covariates and left-truncated time-to-event data. Statistics in Medicine, 32(6):1004– 1015, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2013:SRE

[1081] Zhangsheng Yu, Lei Liu, Dawn M. Bravata, Linda S. Williams, and Robert S. Tepper. A semiparametric recurrent events model with timevarying coefficients. Statistics in Medicine, 32(6):1016–1026, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koopmeiners:2013:ETT

[1082] Joseph S. Koopmeiners and Rachel Isaksson Vogel. Early termination of a two-stage study to develop and validate a panel of biomarkers. Statistics in Medicine, 32(6):1027–1037, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2013:BRC

[1083] Lili Yang and Sujuan Gao. Bivariate random change point models for longitudinal outcomes. Statistics in Medicine, 32(6):1038–1053, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pullenayegum:2013:DRE

[1084] Eleanor M. Pullenayegum and Brian M. Feldman. Doubly robust estimation, optimally truncated inverse-intensity weighting and incrementbased methods for the analysis of irregularly observed longitudinal data. Statistics in Medicine, 32(6):1054–1072, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gagne:2013:CEA

[1085] Joshua J. Gagne and Sebastian Schneeweiss. Comment on ‘empirical assessment of methods for risk identification in healthcare data: results from the experiments of the observational medical outcomes partnership’. Statistics in Medicine, 33(3):1073–1074, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ryan:2013:RCE

[1086] Patrick B. Ryan, David Madigan, Paul E. Stang, J. Marc Overhage, Judith A. Racoosin, and Abraham G. Hartzema. Response to comment on ‘empirical assessment of methods for risk identification in healthcare data’. Statistics in Medicine, 33(3):1075–1077, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2013:CGS

[1087] Peter C. Austin. Correction: ‘Generating survival times to simulate Cox proportional hazards models with time-varying covariates’. Statistics in Medicine, 33(3):1078, March 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dmitrienko:2013:KMI

[1088] Alex Dmitrienko, Ralph B. D’Agostino, Sr., and Mohammad F. Huque. Key multiplicity issues in clinical drug development. Statistics in Medicine, 32(7):1079–1111, March 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ye:2013:GSH

[1089] Yining Ye, Ai Li, Lingyun Liu, and Bin Yao. A group sequential Holm procedure with multiple primary endpoints. Statistics in Medicine, 32 (7):1112–1124, March 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fontana:2013:LMM

[1090] Andrea Fontana, Massimiliano Copetti, Gianluigi Mazzoccoli, Theodore Kypraios, and Fabio Pellegrini. A linear mixed model approach to compare the evolution of multiple biological rhythms. Statistics in Medicine, 32(7):1125–1135, March 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kahan:2013:AMT

[1091] Brennan C. Kahan and Tim P. Morris. Analysis of multicentre trials with continuous outcomes: when and how should we account for centre effects? Statistics in Medicine, 32(7):1136–1149, March 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jaki:2013:CCE

[1092] T. Jaki and D. Magirr. Considerations on covariates and endpoints in multi-arm multi-stage clinical trials selecting all promising treatments. Statistics in Medicine, 32(7):1150–1163, March 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Satagopan:2013:ERS

[1093] Jaya M. Satagopan and Robert C. Elston. Evaluation of removable statistical interaction for binary traits. Statistics in Medicine, 32(7):1164– 1190, March 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:2013:EWS

[1094] Yinghui Wei and Julian Pt Higgins. Estimating within-study covariances in multivariate meta-analysis with multiple outcomes. Statistics in Medicine, 32(7):1191–1205, March 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Noufaily:2013:IAO

[1095] Angela Noufaily, Doyo G. Enki, Paddy Farrington, Paul Garthwaite, Nick Andrews, and André Charlett. An improved algorithm for outbreak detection in multiple surveillance systems. Statistics in Medicine, 32(7): 1206–1222, March 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanGeloven:2013:EDH

[1096] N. van Geloven, I. Martin, P. Damman, R. J. de Winter, J. G. Tijssen, and H. P. Lopuhaä. Estimation of a decreasing hazard of patients with acute coronary syndrome. Statistics in Medicine, 32(7):1223–1238, March 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2013:STA

[1097] Pei He, Liang Fang, and Zheng Su. A sequential testing approach to detecting multiple change points in the proportional hazards model. Statistics in Medicine, 32(7):1239–1245, March 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harbord:2013:SBS

[1098] Roger M. Harbord, Vanessa Didelez, Tom M. Palmer, Sha Meng, Jonathan A. C. Sterne, and Nuala A. Sheehan. Severity of bias of a simple estimator of the causal odds ratio in Mendelian randomization studies. Statistics in Medicine, 32(7):1246–1258, March 30, 2013. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Levin:2013:ACT

[1099] Gregory P. Levin, Sarah C. Emerson, and Scott S. Emerson. Adaptive clinical trial designs with pre-specified rules for modifying the sample size: understanding efficient types of adaptation. Statistics in Medicine, 32(8):1259–1275, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehta:2013:ACT

[1100] Cyrus R. Mehta. Adaptive clinical trial designs with pre-specified rules for modifying the sample size: a different perspective. Statistics in Medicine, 32(8):1276–1279, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Levin:2013:ARA

[1101] Gregory P. Levin, Sarah C. Emerson, and Scott S. Emerson. Authors’ response to ‘adaptive clinical trial designs with pre-specified rules for modifying the sample size: a different perspective’. Statistics in Medicine, 32(8):1280–1282, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wong:2013:MZI

[1102] Kin-Yau Wong and K. F. Lam. Modeling zero-inflated count data using a covariate-dependent random effect model. Statistics in Medicine, 32(8): 1283–1293, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2013:LIM

[1103] Xue-Dong Chen, Ying-Zi Fu, and Xue-Ren Wang. Local influence measure of zero-inflated generalized Poisson mixture regression models. Statistics in Medicine, 32(8):1294–1312, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brumback:2013:ACN

[1104] Babette A. Brumback, Hao W. Zheng, and Amy B. Dailey. Adjusting for confounding by neighborhood using generalized linear mixed models and complex survey data. Statistics in Medicine, 32(8):1313–1324, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brumback:2013:CPM

[1105] Babette A. Brumback, Zhuangyu Cai, Zhulin He, Hao W. Zheng, and Amy B. Dailey. Conditional pseudolikelihood methods for clustered ordinal, multinomial, or count outcomes with complex survey data. Statistics in Medicine, 32(8):1325–1335, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2013:NII

[1106] Zhao Yang, Xuezheng Sun, and James W. Hardin. A non-iterative implementation of Tango’s score confidence interval for a paired difference of proportions. Statistics in Medicine, 32(8):1336–1342, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuk:2013:FCD

[1107] Anthony Y. C. Kuk, Xiang Li, and Jinfeng Xu. A fast collapsed data method for estimating haplotype frequencies from pooled genotype data with applications to the study of rare variants. Statistics in Medicine, 32(8):1343–1360, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mun:2013:DRM

[1108] Jungwon Mun and Mary J. Lindstrom. Diagnostics for repeated measurements in linear mixed effects models. Statistics in Medicine, 32(8): 1361–1375, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abraham:2013:UCM

[1109] Christophe Abraham, Nicolas Molinari, and Rémi Servien. Unsupervised clustering of multivariate circular data. Statistics in Medicine, 32(8):

1376–1382, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Platt:2013:ICM

[1110] Robert W. Platt, M. Alan Brookhart, Stephen R. Cole, Daniel Westreich, and Enrique F. Schisterman. An information criterion for marginal structural models. Statistics in Medicine, 32(8):1383–1393, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2013:JMQ

[1111] Zhigang Li, Tor D. Tosteson, and Marie A. Bakitas. Joint modeling quality of life and survival using a terminal decline model in palliative care studies. Statistics in Medicine, 32(8):1394–1406, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aldrin:2013:ARH

[1112] Magne Aldrin, Ragnhild Raastad, Ingunn Fride Tvete, Dag Berild, Arnoldo Frigessi, Truls Leegaard, Dominique L. Monnet, Mette Walberg, and Fredrik Müller. Antibiotic resistance in hospitals: a wardspecific random effect model in a low antibiotic consumption environment. Statistics in Medicine, 32(8):1407–1418, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choi:2013:PCI

[1113] Byeong Yeob Choi, Jason P. Fine, and M. Alan Brookhart. Practicable confidence intervals for current status data. Statistics in Medicine, 32(8): 1419–1428, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sauzet:2013:MHS

[1114] O. Sauzet, K. C. Wright, L. Marston, P. Brocklehurst, and J. L. Peacock. Modelling the hierarchical structure in datasets with very small clusters: a simulation study to explore the effect of the proportion of clusters when the outcome is continuous. Statistics in Medicine, 32(8):1429–1438, April 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2013:SMR

[1115] Stephen Senn. Seven myths of randomisation in clinical trials. Statistics in Medicine, 32(9):1439–1450, April 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2013:ESD

[1116] Bo Lu and Constantine Gatsonis. Efficiency of study designs in diagnostic randomized clinical trials. Statistics in Medicine, 32(9):1451–1466, April 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pepe:2013:TIP

[1117] Margaret Sullivan Pepe, Kathleen F. Kerr, Gary Longton, and Zheyu Wang. Testing for improvement in prediction model performance. Statistics in Medicine, 32(9):1467–1482, April 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seshan:2013:CRC

[1118] Venkatraman E. Seshan, Mithat Gönen, and Colin B. Begg. Comparing ROC curves derived from regression models. Statistics in Medicine, 32 (9):1483–1493, April 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2013:BCS

[1119] Hua Yun Chen, Rick Kittles, and Wei Zhang. Bias correction to secondary trait analysis with case-control design. Statistics in Medicine, 32(9):1494–1508, April 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2013:JMM

[1120] Huiping Xu, Joanne Daggy, Danni Yu, Bruce A. Craig, and Laura Sands. Joint modeling of medical cost and survival in complex sample surveys. Statistics in Medicine, 32(9):1509–1523, April 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eng:2013:PIM

[1121] Kevin H. Eng, Sijian Wang, William H. Bradley, Janet S. Rader, and Christina Kendziorski. Pathway index models for construction of patientspecific risk profiles. Statistics in Medicine, 32(9):1524–1535, April 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cobre:2013:MBC

[1122] Juliana Cobre, Gleici S. Castro Perdoná, Fernanda M. Peria, and Francisco Louzada. A mechanistic breast cancer survival modelling through the axillary lymph node chain. Statistics in Medicine, 32(9):1536–1546, April 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Price:2013:MEM

[1123] Malcolm J. Price, A. E. Ades, Daniela De Angelis, Nicky J. Welton, John Macleod, Kate Soldan, Katy Turner, Ian Simms, and Paddy J. Horner. Mixture-of-exponentials models to explain heterogeneity in studies of the duration of Chlamydia trachomatis infection. Statistics in Medicine, 32 (9):1547–1560, April 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Majer:2013:TTF

[1124] Istvan M. Majer, Johan P. Mackenbach, and Pieter H. M. van Baal. Time trends and forecasts of body mass index from repeated cross-sectional data: a different approach. Statistics in Medicine, 32(9):1561–1571, April 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2013:MHS

[1125] Ting Hsiang Lin and Min-Hsiao Tsai. Modeling health survey data with excessive zero and K responses. Statistics in Medicine, 32(9):1572–1583, April 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Daniel:2013:MDT

[1126] R. M. Daniel, S. N. Cousens, B. L. De Stavola, M. G. Kenward, and J. A. C. Sterne. Methods for dealing with time-dependent confounding. Statistics in Medicine, 32(9):1584–1618, April 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koenig:2013:JEI

[1127] Franz Koenig and Frank Bretz. Joint EMA, ISBS, and DR-IBS International Symposium on Biopharmaceutical Statistics: Bridging drug development from research to marketing. Statistics in Medicine, 32(10): 1619–1620, May 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scala:2013:CSI

[1128] L. Di Scala, J. Kerman, and B. Neuenschwander. Collection, synthesis, and interpretation of evidence: a proof-of-concept study in COPD. Statistics in Medicine, 32(10):1621–1634, May 10, 2013. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2013:GJT

[1129] Ying Zhang and Paul Cabilio. A generalized Jonckheere test against ordered alternatives for repeated measures in randomized blocks. Statistics in Medicine, 32(10):1635–1645, May 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dette:2013:ETS

[1130] Holger Dette, Björn Bornkamp, and Frank Bretz. On the efficiency of two-stage response-adaptive designs. Statistics in Medicine, 32(10):1646– 1660, May 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lisovskaja:2013:CDP

[1131] Vera Lisovskaja and Carl-Fredrik Burman. On the choice of doses for phase III clinical trials. Statistics in Medicine, 32(10):1661–1676, May 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carreras:2013:SET

[1132] Máximo Carreras and Werner Brannath. Shrinkage estimation in twostage adaptive designs with midtrial treatment selection. Statistics in Medicine, 32(10):1677–1690, May 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quan:2013:ESE

[1133] Hui Quan, Mingyu Li, Weichung Joe Shih, Soo Peter Ouyang, Joshua Chen, Ji Zhang, and Peng-Liang Zhao. Empirical shrinkage estimator for consistency assessment of treatment effects in multi-regional clinical trials. Statistics in Medicine, 32(10):1691–1706, May 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kieser:2013:ASS

[1134] Meinhard Kieser, Tim Friede, and Matthias Gondan. Assessment of statistical significance and clinical relevance. Statistics in Medicine, 32 (10):1707–1719, May 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hasler:2013:SCI

[1135] M. Hasler and L. A. Hothorn. Simultaneous confidence intervals on multivariate non-inferiority. Statistics in Medicine, 32(10):1720–1729, May 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2013:TEH

[1136] Jianjun (David) Li. Testing each hypothesis marginally at alpha while still controlling FWER: how and when. Statistics in Medicine, 32(10): 1730–1738, May 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maurer:2013:MOP

[1137] Willi Maurer and Frank Bretz. Memory and other properties of multiple test procedures generated by entangled graphs. Statistics in Medicine, 32(10):1739–1753, May 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gutjahr:2013:MTC

[1138] G. Gutjahr and W. Brannath. Multiple tests of cost-effectiveness angles. Statistics in Medicine, 32(10):1754–1762, May 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Patel:2013:MMM

[1139] Nitin R. Patel, Suresh Ankolekar, Zoran Antonijevic, and Natasa Rajicic. A mathematical model for maximizing the value of phase 3 drug development portfolios incorporating budget constraints and risk. Statistics in Medicine, 32(10):1763–1777, May 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Binder:2013:TSM

[1140] H. Binder, A. Benner, L. Bullinger, and M. Schumacher. Tailoring sparse multivariable regression techniques for prognostic single-nucleotide polymorphism signatures. Statistics in Medicine, 32(10):1778–1791, May 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gutman:2013:REC

[1141] R. Gutman and D. B. Rubin. Robust estimation of causal effects of binary treatments in unconfounded studies with dichotomous outcomes. Statistics in Medicine, 32(11):1795–1814, May 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Odondi:2013:AOM

[1142] Lang’o Odondi and Roseanne McNamee. Applying optimal model selection in principal stratification for causal inference. Statistics in Medicine, 32(11):1815–1828, May 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Landsman:2013:EAE

[1143] V. Landsman and R. M. Pfeiffer. On estimating average effects for multiple treatment groups. Statistics in Medicine, 32(11):1829–1841, May 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Malloy:2013:TMR

[1144] Michael J. Malloy, Luke A. Prendergast, and Robert G. Staudte. Transforming the model t: random effects meta-analysis with stable weights. Statistics in Medicine, 32(11):1842–1864, May 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vickers:2013:SDA

[1145] Andrew J. Vickers, Angel M. Cronin, and Mithat Gönen. A simple decision analytic solution to the comparison of two binary diagnostic tests. Statistics in Medicine, 32(11):1865–1876, May 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bansal:2013:WDC

[1146] Aasthaa Bansal and Margaret Sullivan Pepe. When does combining markers improve classification performance and what are implications for practice? Statistics in Medicine, 32(11):1877–1892, May 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:2013:MRT

[1147] Yuan chin Ivan Chang. Maximizing an ROC-type measure via linear combination of markers when the gold reference is continuous. Statistics in Medicine, 32(11):1893–1903, May 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Noma:2013:EBR

[1148] Hisashi Noma and Shigeyuki Matsui. Empirical Bayes ranking and selection methods via semiparametric hierarchical mixture models in microarray studies. Statistics in Medicine, 32(11):1904–1916, May 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korostil:2013:AMC

[1149] Igor A. Korostil, Gareth W. Peters, Julien Cornebise, and David G. Regan. Adaptive Markov chain Monte Carlo forward projection for statistical analysis in epidemic modelling of human papillomavirus. Statistics in Medicine, 32(11):1917–1953, May 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2013:SAM

[1150] Liang Zhu, Xingwei Tong, Hui Zhao, Jianguo Sun, Deo Kumar Srivastava, Wendy Leisenring, and Leslie L. Robison. Statistical analysis of mixed recurrent event data with application to cancer survivor study. Statistics in Medicine, 32(11):1954–1963, May 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kraemer:2013:DCC

[1151] Helena Chmura Kraemer. Discovering, comparing, and combining moderators of treatment on outcome after randomized clinical trials: a parametric approach. Statistics in Medicine, 32(11):1964–1973, May 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scala:2013:ENT

[1152] Lilla Di Scala and Ekkehard Glimm. Erratum note to ‘time-to-event analysis with treatment arm selection at interim’. Statistics in Medicine, 32(11):1974, May 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2013:BDT

[1153] Fei Jiang, J. Jack Lee, and Peter Müller. A Bayesian decisiontheoretic sequential response-adaptive randomization design. Statistics in Medicine, 32(12):1975–1994, May 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qu:2013:ESM

[1154] Yongming Qu. Evaluation of a surrogate marker: validity and efficiency. Statistics in Medicine, 32(12):1995–2000, May 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wen:2013:AAO

[1155] Chi-Chung Wen and Yi-Hau Chen. Assessing age-at-onset risk factors with incomplete covariate current status data under proportional odds models. Statistics in Medicine, 32(12):2001–2012, May 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2013:TVC

[1156] Xiao Song and Ching-Yun Wang. Time-varying coefficient proportional hazards model with missing covariates. Statistics in Medicine, 32(12): 2013–2030, May 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nicolaie:2013:DPL

[1157] M. A. Nicolaie, J. C. van Houwelingen, T. M. de Witte, and H. Putter. Dynamic prediction by landmarking in competing risks. Statistics in Medicine, 32(12):2031–2047, May 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demidenko:2013:SEI

[1158] Eugene Demidenko, Benjamin B. Williams, Ann Barry Flood, and Harold M. Swartz. Standard error of inverse prediction for dose-response relationship: approximate and exact statistical inference. Statistics in Medicine, 32(12):2048–2061, May 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lowsky:2013:NNS

[1159] D. J. Lowsky, Y. Ding, D. K. K. Lee, C. E. McCulloch, L. F. Ross, J. R. Thistlethwaite, and S. A. Zenios. A K-nearest neighbors survival probability prediction method. Statistics in Medicine, 32(12):2062–2069, May 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brinks:2013:DAS

[1160] Ralph Brinks, Sandra Landwehr, Andrea Icks, Michael Koch, and Guido Giani. Deriving age-specific incidence from prevalence with an ordinary differential equation. Statistics in Medicine, 32(12):2070–2078, May 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mukherjee:2013:EBC

[1161] Partha Sarathi Mukherjee and Peihua Qiu. Efficient bias correction for magnetic resonance image denoising. Statistics in Medicine, 32(12):2079– 2096, May 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2013:IIG

[1162] Guangyu Zhang, Nathaniel Schenker, Jennifer D. Parker, and Dan Liao. Identifying implausible gestational ages in preterm babies with Bayesian mixture models. Statistics in Medicine, 32(12):2097–2113, May 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shahbaba:2013:BNV

[1163] Babak Shahbaba and Wesley O. Johnson. Bayesian nonparametric variable selection as an exploratory tool for discovering differentially expressed genes. Statistics in Medicine, 32(12):2114–2126, May 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2013:NBR

[1164] Hokeun Sun and Shuang Wang. Network-based regularization for matched case-control analysis of high-dimensional DNA methylation data. Statistics in Medicine, 32(12):2127–2139, May 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Esserman:2013:SSE

[1165] Denise Esserman, Yingqi Zhao, Yiyun Tang, and Jianwen Cai. Sample size estimation in educational intervention trials with subgroup heterogeneity in only one arm. Statistics in Medicine, 32(12):2140–2154, May 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawless:2013:ALD

[1166] Jerald F. Lawless. Armitage Lecture 2011: The design and analysis of life history studies. Statistics in Medicine, 32(13):2155–2172, June 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gerds:2013:ETD

[1167] Thomas A. Gerds, Michael W. Kattan, Martin Schumacher, and Changhong Yu. Estimating a time-dependent concordance index for survival prediction models with covariate dependent censoring. Statistics in Medicine, 32(13):2173–2184, June 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neill:2013:FSS

[1168] Daniel B. Neill, Edward Mcfowland III, and Huanian Zheng. Fast subset scan for multivariate event detection. Statistics in Medicine, 32(13): 2185–2208, June 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2013:NRS

[1169] Liansheng Larry Tang, Aiyi Liu, Zhen Chen, Enrique F. Schisterman, Bo Zhang, and Zhuang Miao. Nonparametric ROC summary statistics for correlated diagnostic marker data. Statistics in Medicine, 32(13): 2209–2220, June 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2013:APV

[1170] Jin Zhang, Thomas M. Braun, and Jeremy M. G. Taylor. Adaptive prior variance calibration in the Bayesian continual reassessment method. Statistics in Medicine, 32(13):2221–2234, June 15, 2013. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fagerland:2013:GFT

[1171] Morten W. Fagerland and David W. Hosmer. A goodness-of-fit test for the proportional odds regression model. Statistics in Medicine, 32(13): 2235–2249, June 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Capuano:2013:TOM

[1172] Ana W. Capuano and Jeffrey D. Dawson. The trend odds model for ordinal data. Statistics in Medicine, 32(13):2250–2261, June 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Binder:2013:CBS

[1173] Harald Binder, Willi Sauerbrei, and Patrick Royston. Comparison between splines and fractional polynomials for multivariable model building with continuous covariates: a simulation study with continuous response. Statistics in Medicine, 32(13):2262–2277, June 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ning:2013:ELB

[1174] J. Ning, J. Qin, M. Asgharian, and Y. Shen. Empirical likelihood-based confidence intervals for length-biased data. Statistics in Medicine, 32 (13):2278–2291, June 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2013:ECD

[1175] Xinlei Wang, Miao Zang, and Guanghua Xiao. Epigenetic change detection and pattern recognition via Bayesian hierarchical hidden Markov models. Statistics in Medicine, 32(13):2292–2307, June 15, 2013. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roberts:2013:BMT

[1176] Eric M. Roberts and Paul B. English. Bayesian modeling of timedependent vulnerability to environmental hazards: an example using autism and pesticide data. Statistics in Medicine, 32(13):2308–2319, June 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2013:BMM

[1177] Sheng Luo, Min Yi, Xuelin Huang, and Kelly K. Hunt. A Bayesian model for misclassified binary outcomes and correlated survival data with applications to breast cancer. Statistics in Medicine, 32(13):2320–2334, June 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rohmel:2013:INI

[1178] Joachim Röhmel and Meinhard Kieser. Investigations on non-inferiority-the food and drug administration draft guidance on treatments for nosocomial pneumonia as a case for exact tests for binomial proportions. Statistics in Medicine, 32(14):2335–2348, June 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Soon:2013:AOE

[1179] Guoxing Soon, Zhiwei Zhang, Yi Tsong, and Lei Nie. Assessing overall evidence from noninferiority trials with shared historical data. Statistics in Medicine, 32(14):2349–2363, June 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Niu:2013:SMM

[1180] Yi Niu and Yingwei Peng. A semiparametric marginal mixture cure model for clustered survival data. Statistics in Medicine, 32(14):2364– 2373, June 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xin:2013:TBE

[1181] X. Xin, J. Horrocks, and G. A. Darlington. Ties between event times and jump times in the Cox model. Statistics in Medicine, 32(14):2374–2389, June 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2013:DFM

[1182] Q. Yu, R. Chen, W. Tang, H. He, R. Gallop, P. Crits-Christoph, J. Hu, and X. M. Tu. Distribution-free models for longitudinal count responses with overdispersion and structural zeros. Statistics in Medicine, 32(14): 2390–2405, June 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2013:ERA

[1183] Kaifeng Lu. An efficient and robust analysis of covariance model for baseline adjustment in parallel-group thorough QT/QTc studies. Statistics in Medicine, 32(14):2406–2418, June 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neuhaus:2013:ECE

[1184] John M. Neuhaus, Charles E. McCulloch, and Ross Boylan. Estimation of covariate effects in generalized linear mixed models with a misspecified distribution of random intercepts and slopes. Statistics in Medicine, 32 (14):2419–2429, June 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Uno:2013:UIP

[1185] Hajime Uno, Lu Tian, Tianxi Cai, Isaac S. Kohane, and L. J. Wei. A unified inference procedure for a class of measures to assess improvement in risk prediction systems with survival data. Statistics in Medicine, 32 (14):2430–2442, June 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hooper:2013:CDT

[1186] R. Hooper, K. Díaz-Ordaz, A. Takeda, and K. Khan. Comparing diagnostic tests: trials in people with discordant test results. Statistics in Medicine, 32(14):2443–2456, June 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Locatelli:2013:RPI

[1187] Isabella Locatelli and Alfio Marazzi. Robust parametric indirect estimates of the expected cost of a hospital stay with covariates and censored data. Statistics in Medicine, 32(14):2457–2466, June 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fluss:2013:CSB

[1188] Ronen Fluss, Micha Mandel, Laurence S. Freedman, Inbal Salz Weiss, Anat Ekka Zohar, Ziona Haklai, Ethel-Sherry Gordon, and Elisheva Simchen. Correction of sampling bias in a cross-sectional study of postsurgical complications. Statistics in Medicine, 32(14):2467–2478, June 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2013:RRE

[1189] Dalei Yu and Kelvin K. W. Yau. Robust REML estimation for kcomponent Poisson mixture with random effects: application to the epilepsy seizure count data and urinary tract infections data. Statistics in Medicine, 32(14):2479–2499, June 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Persson:2013:EMC

[1190] Emma Persson and Ingeborg Waernbaum. Estimating a marginal causal odds ratio in a case-control design: analyzing the effect of low birth weight on the risk of type 1 diabetes mellitus. Statistics in Medicine, 32(14):2500–2512, June 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friede:2013:CCE

[1191] Tim Friede, Nick Parsons, and Nigel Stallard. Correction: ‘A conditional error function approach for subgroup selection in adaptive clinical trials’. Statistics in Medicine, 32(14):2513–2514, June 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vandeVen:2013:NIS

[1192] Peter M. van de Ven and Johannes Berkhof. Non-inferiority and superiority tests for partially matched data when a constraint is placed on the number of new tests. Statistics in Medicine, 32(15):2515–2528, July 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2013:ITS

[1193] Yujun Wu and Peng-Liang Zhao. Interim treatment selection with a flexible selection margin in clinical trials. Statistics in Medicine, 32(15): 2529–2543, July 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jaki:2013:DEC

[1194] Thomas Jaki, Valérie André, Ting-Li Su, and John Whitehead. Designing exploratory cancer trials using change in tumour size as primary endpoint. Statistics in Medicine, 32(15):2544–2554, July 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Geneletti:2013:USB

[1195] S. Geneletti, N. Best, M. B. Toledano, P. Elliott, and S. Richardson. Uncovering selection bias in case-control studies using Bayesian poststratification. Statistics in Medicine, 32(15):2555–2570, July 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2013:EAD

[1196] Huiping Xu, Michael A. Black, and Bruce A. Craig. Evaluating accuracy of diagnostic tests with intermediate results in the absence of a gold standard. Statistics in Medicine, 32(15):2571–2584, July 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bamia:2013:SCA

[1197] Christina Bamia, Ian R. White, and Michael G. Kenward. Some consequences of assuming simple patterns for the treatment effect over time in a linear mixed model. Statistics in Medicine, 32(15):2585–2594, July 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Botella-Rocamora:2013:SMA

[1198] P. Botella-Rocamora, A. López-Quílez, and M. A. Martinez-Beneito. Spatial moving average risk smoothing. Statistics in Medicine, 32(15): 2595–2612, July 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ding:2013:BIM

[1199] Ying Ding and Haoda Fu. Bayesian indirect and mixed treatment comparisons across longitudinal time points. Statistics in Medicine, 32(15): 2613–2628, July 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ouyang:2013:BAR

[1200] Bichun Ouyang, Debajyoti Sinha, Elizabeth H. Slate, and Adrian B. Van Bakel. Bayesian analysis of recurrent event with dependent termination: an application to a heart transplant study. Statistics in Medicine, 32(15): 2629–2642, July 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nunes:2013:NIE

[1201] Baltazar Nunes, Isabel Natário, and M. Lucília Carvalho. Nowcasting influenza epidemics using non-homogeneous hidden Markov models. Statistics in Medicine, 32(15):2643–2660, July 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2013:PPA

[1202] Kaibo Liu, Shabnam Jain, and Jianjun Shi. Physician performance assessment using a composite quality index. Statistics in Medicine, 32(15): 2661–2680, July 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2013:EIW

[1203] Wei Luo, Jiguo Cao, Marcus Gallagher, and Janet Wiles. Estimating the intensity of ward admission and its effect on emergency department access block. Statistics in Medicine, 32(15):2681–2694, July 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Magnusson:2013:GSE

[1204] Baldur P. Magnusson and Bruce W. Turnbull. Group sequential enrichment design incorporating subgroup selection. Statistics in Medicine, 32(16):2695–2714, July 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ahmed:2013:RCS

[1205] Anwar E. Ahmed, Christine M. Schubert, and Donna K. McClish. Reducing cost in sequential testing: a limit of indifference approach. Statistics in Medicine, 32(16):2715–2727, July 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ezzalfani:2013:DFD

[1206] Monia Ezzalfani, Sarah Zohar, Rui Qin, Sumithra J. Mandrekar, and Marie-Cécile Le Deley. Dose-finding designs using a novel quasicontinuous endpoint for multiple toxicities. Statistics in Medicine, 32 (16):2728–2746, July 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2013:MAR

[1207] Richard D. Riley, Iram Kauser, Martin Bland, Lutgarde Thijs, Jan A. Staessen, Jiguang Wang, Francois Gueyffier, and Jonathan J. Deeks. Meta-analysis of randomised trials with a continuous outcome according to baseline imbalance and availability of individual participant data. Statistics in Medicine, 32(16):2747–2766, July 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Doros:2013:RMM

[1208] Gheorghe Doros, Michael Pencina, Denis Rybin, Allison Meisner, and Maurizio Fava. A repeated measures model for analysis of continuous outcomes in sequential parallel comparison design studies. Statistics in Medicine, 32(16):2767–2789, July 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heggeseth:2013:ICM

[1209] Brianna C. Heggeseth and Nicholas P. Jewell. The impact of covariance misspecification in multivariate Gaussian mixtures on estimation and inference: an application to longitudinal modeling. Statistics in Medicine, 32(16):2790–2803, July 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

MacKenzie:2013:ICP

[1210] Gilbert MacKenzie and Defen Peng. Interval-censored parametric regression survival models and the analysis of longitudinal trials. Statistics in Medicine, 32(16):2804–2822, July 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fagerland:2013:CIO

[1211] Morten W. Fagerland and Robert G. Newcombe. Confidence intervals for odds ratio and relative risk based on the inverse hyperbolic sine transformation. Statistics in Medicine, 32(16):2823–2836, July 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2013:PDP

[1212] Peter C. Austin. The performance of different propensity score methods for estimating marginal hazard ratios. Statistics in Medicine, 32(16): 2837–2849, July 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Westgate:2013:BCC

[1213] Philip M. Westgate. A bias correction for covariance estimators to improve inference with generalized estimating equations that use an unstructured correlation matrix. Statistics in Medicine, 32(16):2850–2858, July 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2013:CCC

[1214] Andreas Andersen, Christine S. Benn, Mathias J. Jørgensen, and Henrik Ravn. Censored correlated cytokine concentrations: multivariate tobit regression using clustered variance estimation. Statistics in Medicine, 32(16):2859–2874, July 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petkova:2013:IMR

[1215] Eva Petkova, Thaddeus Tarpey, Lei Huang, and Liping Deng. Interpreting meta-regression: application to recent controversies in antidepressants’ efficacy. Statistics in Medicine, 32(17):2875–2892, July 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kimani:2013:CUE

[1216] Peter K. Kimani, Susan Todd, and Nigel Stallard. Conditionally unbiased estimation in phase II/III clinical trials with early stopping for futility. Statistics in Medicine, 32(17):2893–2910, July 30, 2013. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:2013:BMM

[1217] Yinghui Wei and Julian P. T. Higgins. Bayesian multivariate metaanalysis with multiple outcomes. Statistics in Medicine, 32(17):2911– 2934, July 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmitz:2013:IDV

[1218] Susanne Schmitz, Roisin Adams, and Cathal Walsh. Incorporating data from various trial designs into a mixed treatment comparison model. Statistics in Medicine, 32(17):2935–2949, July 30, 2013. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bollen:2013:GAB

[1219] Kenneth A. Bollen, Mark D. Noble, and Linda S. Adair. Are gestational age, birth weight, and birth length indicators of favorable fetal growth conditions? A structural equation analysis of Filipino infants. Statistics in Medicine, 32(17):2950–2961, July 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lehre:2013:DBW

[1220] Anne-Catherine Lehre, Petter Laake, and Joseph Andrew Sexton. Differences in birth weight by sex using adjusted quantile distance functions. Statistics in Medicine, 32(17):2962–2970, July 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senturk:2013:MTV

[1221] Damla Sentürk, Lorien S. Dalrymple, Sandra M. Mohammed, George A. Kaysen, and Danh V. Nguyen. Modeling time-varying effects with generalized and unsynchronized longitudinal data. Statistics in Medicine, 32(17):2971–2987, July 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Verhagen:2013:LMH

[1222] Josine Verhagen and Jean-Paul Fox. Longitudinal measurement in health-related surveys. a Bayesian joint growth model for multivariate ordinal responses. Statistics in Medicine, 32(17):2988–3005, July 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mostajabi:2013:NRS

[1223] Farida Mostajabi and Somnath Datta. Nonparametric regression of state occupation, entry, exit, and waiting times with multistate right-censored data. Statistics in Medicine, 32(17):3006–3019, July 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jaki:2013:AMD

[1224] Thomas Jaki, Alice Parry, Katherine Winter, and Ian Hastings. Analysing malaria drug trials on a per-individual or per-clone basis: a comparison of methods. Statistics in Medicine, 32(17):3020–3038, July 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2013:STM

[1225] Ni Li, Hui Zhao, and Jianguo Sun. Semiparametric transformation models for panel count data with correlated observation and follow-up times. Statistics in Medicine, 32(17):3039–3054, July 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kombrink:2013:DSA

[1226] Karola Kombrink, Axel Munk, and Tim Friede. Design and semiparametric analysis of non-inferiority trials with active and placebo control for censored time-to-event data. Statistics in Medicine, 32(18):3055–3066, August 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sjolander:2013:BWM

[1227] Arvid Sjölander, Paul Lichtenstein, Henrik Larsson, and Yudi Pawitan. Between-within models for survival analysis. Statistics in Medicine, 32 (18):3067–3076, August 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuk:2013:MSC

[1228] Deborah Kuk and Ravi Varadhan. Model selection in competing risks regression. Statistics in Medicine, 32(18):3077–3088, August 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cortese:2013:CPA

[1229] Giuliana Cortese, Thomas A. Gerds, and Per K. Andersen. Comparing predictions among competing risks models with time-dependent covariates. Statistics in Medicine, 32(18):3089–3101, August 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mogensen:2013:RFA

[1230] Ulla B. Mogensen and Thomas A. Gerds. A random forest approach for competing risks based on pseudo-values. Statistics in Medicine, 32 (18):3102–3114, August 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dai:2013:MMS

[1231] Hongying Dai, Richard Charnigo, Carrie A. Vyhlidal, Bridgette L. Jones, and Madhusudan Bhandary. Mixed modeling and sample size calculations for identifying housekeeping genes. Statistics in Medicine, 32 (18):3115–3125, August 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2013:SOR

[1232] Hua Yun Chen, Muredach P. Reilly, and Mingyao Li. Semiparametric odds ratio model for case-control and matched case-control designs. Statistics in Medicine, 32(18):3126–3142, August 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sabbe:2013:ELL

[1233] N. Sabbe, O. Thas, and J-P. Ottoy. EMLasso: logistic lasso with missing data. Statistics in Medicine, 32(18):3143–3157, August 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Debray:2013:FDI

[1234] Thomas P. A. Debray, Karel G. M. Moons, Ikhlaaq Ahmed, Hendrik Koffijberg, and Richard David Riley. A framework for developing, implementing, and evaluating clinical prediction models in an individual participant data meta-analysis. Statistics in Medicine, 32(18):3158–3180, August 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parsons:2013:POM

[1235] Nick R. Parsons. Proportional-odds models for repeated composite and long ordinal outcome scales. Statistics in Medicine, 32(18):3181–3191, August 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2013:PCO

[1236] Yueqiong Lin, Siu Hung Cheung, Wai-Yin Poon, and Tong-Yu Lu. Pairwise comparisons with ordered categorical data. Statistics in Medicine, 32(18):3192–3205, August 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moreno-Betancur:2013:RMC

[1237] Margarita Moreno-Betancur and Aurélien Latouche. Regression modeling of the cumulative incidence function with missing causes of failure using pseudo-values. Statistics in Medicine, 32(18):3206–3223, August 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Getachew:2013:CTS

[1238] Yehenew Getachew, Paul Janssen, Delenasaw Yewhalaw, Niko Speybroeck, and Luc Duchateau. Coping with time and space in modelling malaria incidence: a comparison of survival and count regression models. Statistics in Medicine, 32(18):3224–3233, August 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2013:C

[1239] Aijing Zhang. Correction. Statistics in Medicine, 32(18):3234, August 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ning:2013:RRR

[1240] Yang Ning and Mark Woodward. Relative risk regression models with inverse polynomials. Statistics in Medicine, 32(19):3235–3246, August 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuan:2013:SRM

[1241] Pei Fen Kuan and Bo Huang. A simple and robust method for partially matched samples using the p-values pooling approach. Statistics in Medicine, 32(19):3247–3259, August 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Westgate:2013:IQI

[1242] Philip M. Westgate and Thomas M. Braun. An improved quadratic inference function for parameter estimation in the analysis of correlated data. Statistics in Medicine, 32(19):3260–3273, August 30, 2013. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2013:AMN

[1243] Yunlong Xie and Dale L. Zimmerman. Antedependence models for nonstationary categorical longitudinal data with ignorable missingness: likelihood-based inference. Statistics in Medicine, 32(19):3274–3289, August 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2013:SSI

[1244] Stanley Xu, Simon J. Hambidge, David L. McClure, Matthew F. Daley, and Jason M. Glanz. A scan statistic for identifying optimal risk windows in vaccine safety studies using self-controlled case series design. Statistics in Medicine, 32(19):3290–3299, August 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Read:2013:SUG

[1245] S. Read, P. A. Bath, P. Willett, and R. Maheswaran. A study on the use of Gumbel approximation with the Bernoulli spatial scan statistic. Statistics in Medicine, 32(19):3300–3313, August 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ip:2013:MPH

[1246] E. H. Ip, Q. Zhang, R. Schwartz, J. Tooze, X. Leng, H. Han, and D. A. Williamson. Multi-profile hidden Markov model for mood, dietary intake, and physical activity in an intervention study of childhood obesity. Statistics in Medicine, 32(19):3314–3331, August 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Draisma:2013:NCT

[1247] Gerrit Draisma and Joost van Rosmalen. A note on the catch-up time method for estimating lead or sojourn time in prostate cancer screening. Statistics in Medicine, 32(19):3332–3341, August 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Langrock:2013:CHM

[1248] Roland Langrock, Bruce J. Swihart, Brian S. Caffo, Naresh M. Punjabi, and Ciprian M. Crainiceanu. Combining hidden Markov models for comparing the dynamics of multiple sleep electroencephalograms. Statistics in Medicine, 32(19):3342–3356, August 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leyrat:2013:PSU

[1249] C. Leyrat, A. Caille, A. Donner, and B. Giraudeau. Propensity scores used for analysis of cluster randomized trials with selection bias: a simulation study. Statistics in Medicine, 32(19):3357–3372, August 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2013:PSW

[1250] Fan Li, Alan M. Zaslavsky, and Mary Beth Landrum. Propensity score weighting with multilevel data. Statistics in Medicine, 32(19):3373–3387, August 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McCaffrey:2013:TPS

[1251] Daniel F. McCaffrey, Beth Ann Griffin, Daniel Almirall, Mary Ellen Slaughter, Rajeev Ramchand, and Lane F. Burgette. A tutorial on propensity score estimation for multiple treatments using generalized boosted models. Statistics in Medicine, 32(19):3388–3414, August 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lloyd:2013:ACL

[1252] Chris J. Lloyd. Accurate confidence limits for stratified clinical trials. Statistics in Medicine, 32(20):3415–3423, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wason:2013:PMA

[1253] James M. S. Wason, Thomas Jaki, and Nigel Stallard. Planning multiarm screening studies within the context of a drug development program. Statistics in Medicine, 32(20):3424–3435, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coubard:2013:MPM

[1254] Olivier A. Coubard. A method for processing multivariate data in medical studies. Statistics in Medicine, 32(20):3436–3448, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2013:UPA

[1255] Hua Ma, Andriy I. Bandos, Howard E. Rockette, and David Gur. On use of partial area under the ROC curve for evaluation of diagnostic performance. Statistics in Medicine, 32(20):3449–3458, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parast:2013:LRP

[1256] Layla Parast and Tianxi Cai. Landmark risk prediction of residual life for breast cancer survival. Statistics in Medicine, 32(20):3459–3471, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2013:CRE

[1257] Yunlong Xie, Zhen Chen, and Paul S. Albert. A crossed random effects modeling approach for estimating diagnostic accuracy from ordinal ratings without a gold standard. Statistics in Medicine, 32(20):3472–3485, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanHouwelingen:2013:VCR

[1258] Hans C. van Houwelingen, Paola Rebora, and Marie Reilly. Validation, calibration and refitting of a familial breast cancer model in sisterships: a case study in the Swedish sisters data. Statistics in Medicine, 32(20): 3486–3500, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2013:IHP

[1259] Xiaobo Guo, Dungang Liu, Canhong Wen, Mingguang He, and Xueqin Wang. Incorporating heterogeneous parent-child environmental effects in biometrical genetic models. Statistics in Medicine, 32(20):3501–3508, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2013:IAM

[1260] Jin Liu, Jian Huang, and Shuangge Ma. Integrative analysis of multiple cancer genomic datasets under the heterogeneity model. Statistics in Medicine, 32(20):3509–3521, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sansom:2013:CAC

[1261] P. Sansom, V. R. Copley, F. C. Naik, S. Leach, and I. M. Hall. A caseassociation cluster detection and visualisation tool with an application to Legionnaires’ disease. Statistics in Medicine, 32(20):3522–3538, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Uddin:2013:MMP

[1262] Shahadat Uddin, Jafar Hamra, and Liaquat Hossain. Mapping and modeling of physician collaboration network. Statistics in Medicine, 32(20): 3539–3551, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pressler:2013:UPS

[1263] Taylor R. Pressler and Eloise E. Kaizar. The use of propensity scores and observational data to estimate randomized controlled trial generalizability bias. Statistics in Medicine, 32(20):3552–3568, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2013:UGB

[1264] Toby A. White and Elena A. Erosheva. Using group-based latent class transition models to analyze chronic disability data from the National Long-Term Care Survey 1984–2004. Statistics in Medicine, 32(20):3569– 3589, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taguri:2013:CIC

[1265] Masataka Taguri and Yutaka Matsuyama. Comments on ‘An information criterion for marginal structural models’ by R. W. Platt, M. A. Brookhart, S. R. Cole, D. Westreich, and E. F. Schisterman. Statistics in Medicine, 32(20):3590–3591, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Platt:2013:RTM

[1266] Robert W. Platt, M. Alan Brookhart, Stephen R. Cole, Daniel Westreich, and Enrique F. Schisterman. Reply to Taguri and Matsuyama. Statistics in Medicine, 32(20):3592–3593, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2013:C

[1267] Shuangge Ma. Correction. Statistics in Medicine, 32(20):3594, September 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rauch:2013:PEC

[1268] G. Rauch and J. Beyersmann. Planning and evaluating clinical trials with composite time-to-first-event endpoints in a competing risk framework. Statistics in Medicine, 32(21):3595–3608, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gsteiger:2013:UHC

[1269] Sandro Gsteiger, Beat Neuenschwander, Francois Mercier, and Heinz Schmidli. Using historical control information for the design and analysis of clinical trials with overdispersed count data. Statistics in Medicine, 32(21):3609–3622, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schneider:2013:RMB

[1270] Simon Schneider, Heinz Schmidli, and Tim Friede. Robustness of methods for blinded sample size re-estimation with overdispersed count data. Statistics in Medicine, 32(21):3623–3635, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2013:SSE

[1271] Wen Li, Yongming Qu, and Pandurang M. Kulkarni. Sensitivity in statistical evaluation of biomarkers. Statistics in Medicine, 32(21):3636–3645, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2013:VSM

[1272] Qixuan Chen and Sijian Wang. Variable selection for multiply-imputed data with application to dioxin exposure study. Statistics in Medicine, 32(21):3646–3659, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roislien:2013:FEA

[1273] Jo Røislien and Brita Winje. Feature extraction across individual time series observations with spikes using wavelet principal component analysis. Statistics in Medicine, 32(21):3660–3669, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2013:BSM

[1274] Bo Cai, Andrew B. Lawson, Md. Monir Hossain, Jungsoon Choi, Russell S. Kirby, and Jihong Liu. Bayesian semiparametric model with spatially-temporally varying coefficients selection. Statistics in Medicine, 32(21):3670–3685, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xing:2013:CBH

[1275] Li Xing, Igor Burstyn, David B. Richardson, and Paul Gustafson. A comparison of Bayesian hierarchical modeling with group-based exposure assessment in occupational epidemiology. Statistics in Medicine, 32(21):

3686–3699, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2013:KSC

[1276] Chaeryon Kang, Bahjat Qaqish, Jane Monaco, Stacey L. Sheridan, and Jianwen Cai. Kappa statistic for clustered dichotomous responses from physicians and patients. Statistics in Medicine, 32(21):3700–3719, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kasza:2013:EPA

[1277] J. Kasza, J. L. Moran, P. J. Solomon, and ANZICS Centre for Outcome and Resource Evaluation (CORE) of the Australian and New Zealand Intensive Care Society (ANZICS). Evaluating the performance of Australian and New Zealand intensive care units in 2009 and 2010. Statistics in Medicine, 32(21):3720–3736, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wigger:2013:MHF

[1278] Leonore Wigger, Julia E. Vogt, and Volker Roth. Malaria haplotype frequency estimation. Statistics in Medicine, 32(21):3737–3751, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lipsky:2013:RAD

[1279] Ari M. Lipsky and Roger J. Lewis. Response-adaptive decision-theoretic trial design: operating characteristics and ethics. Statistics in Medicine, 32(21):3752–3765, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bland:2013:DLT

[1280] J. M. Bland, D. G. Altman, and F. J. Rohlf. In defence of logarithmic transformations. Statistics in Medicine, 32(21):3766–3768, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alexander:2013:CLT

[1281] Neal Alexander and Karim Anaya-Izquierdo. Comments on ‘Log transformation: application and interpretation in biomedical research’. Statistics in Medicine, 32(21):3768–3769, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nieboer:2013:LTB

[1282] Daan Nieboer, Ewout W. Steyerberg, Sabita Soedamah-Muthu, and Yvonne Vergouwe. Log transformation in biomedical research: (mis)use for covariates. Statistics in Medicine, 32(21):3770–3771, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:2013:RCL

[1283] Changyong Feng, Hongyue Wang, Naiji Lu, and Xin M. Tu. Response to comments on ‘log transformation: Application and interpretation in biomedical research’. Statistics in Medicine, 32(21):3772–3774, September 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Loux:2013:SFE

[1284] Travis M. Loux. A simple, flexible, and effective covariate-adaptive treatment allocation procedure. Statistics in Medicine, 32(22):3775–3787, September 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2013:ITC

[1285] Patrick Royston and Willi Sauerbrei. Interaction of treatment with a continuous variable: simulation study of significance level for several methods of analysis. Statistics in Medicine, 32(22):3788–3803, September 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2013:GFT

[1286] Bingqing Zhou, Jason Fine, and Glen Laird. Goodness-of-fit test for proportional subdistribution hazards model. Statistics in Medicine, 32 (22):3804–3811, September 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2013:RBI

[1287] Sheng Luo, Junsheng Ma, and Karl D. Kieburtz. Robust Bayesian inference for multivariate longitudinal data by using normal/independent distributions. Statistics in Medicine, 32(22):3812–3828, September 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shinohara:2013:EPM

[1288] Russell T. Shinohara, Anand K. Narayan, Kelvin Hong, Hyun S. Kim, Josef Coresh, Michael B. Streiff, and Constantine E. Frangakis. Estimating parsimonious models of longitudinal causal effects using regressions on propensity scores. Statistics in Medicine, 32(22):3829–3837, September 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keogh:2013:USB

[1289] Ruth H. Keogh, Ian R. White, and Sheila A. Rodwell. Using surrogate biomarkers to improve measurement error models in nutritional epidemiology. Statistics in Medicine, 32(22):3838–3861, September 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Regnier:2013:SSS

[1290] Eva D. Regnier and Steven M. Shechter. State-space size considerations for disease-progression models. Statistics in Medicine, 32(22):3862–3880, September 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dagne:2013:BSM

[1291] Getachew A. Dagne and Yangxin Huang. Bayesian semiparametric mixture tobit models with left censoring, skewness, and covariate measurement errors. Statistics in Medicine, 32(22):3881–3898, September 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Das:2013:BSM

[1292] Kiranmoy Das, Runze Li, Subhajit Sengupta, and Rongling Wu. A Bayesian semiparametric model for bivariate sparse longitudinal data. Statistics in Medicine, 32(22):3899–3910, September 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2013:PCL

[1293] Binbing Yu. Predicting county-level cancer incidence rates and counts in the USA. Statistics in Medicine, 32(22):3911–3925, September 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bujkiewicz:2013:MMA

[1294] Sylwia Bujkiewicz, John R. Thompson, Alex J. Sutton, Nicola J. Cooper, Mark J. Harrison, Deborah P. M. Symmons, and Keith R. Abrams. Multivariate meta-analysis of mixed outcomes: a Bayesian approach. Statistics in Medicine, 32(22):3926–3943, September 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Foster:2013:VSM

[1295] Jared C. Foster, Jeremy M. G. Taylor, and Bin Nan. Variable selection in monotone single-index models via the adaptive LASSO. Statistics in Medicine, 32(22):3944–3954, September 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2013:ARR

[1296] Richard J. Cook. Authors’ redress on ‘robustness of methods for blinded sample size re-estimation with overdispersed count data’. Statistics in Medicine, 32(22):3955–3957, September 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mittal:2013:LSP

[1297] Sushil Mittal, David Madigan, Jerry Q. Cheng, and Randall S. Burd. Large-scale parametric survival analysis. Statistics in Medicine, 32(23): 3955–3971, October 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2013:BIM

[1298] Sungduk Kim, Ming-Hui Chen, Joseph G. Ibrahim, Arvind K. Shah, and Jianxin Lin. Bayesian inference for multivariate meta-analysis Box– Cox transformation models for individual patient data with applications to evaluation of cholesterol-lowering drugs. Statistics in Medicine, 32 (23):3972–3990, October 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:2013:EIA

[1299] Ping Gao, Lingyun Liu, and Cyrus Mehta. Exact inference for adaptive group sequential designs. Statistics in Medicine, 32(23):3991–4005, October 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eberly:2013:EPB

[1300] Lynn E. Eberly, James S. Hodges, Kay Savik, Olga Gurvich, Donna Z. Bliss, and Christine Mueller. Extending the Peters–Belson approach for assessing disparities to right censored time-to-event outcomes. Statistics in Medicine, 32(23):4006–4020, October 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keogh:2013:UFC

[1301] Ruth H. Keogh and Ian R. White. Using full-cohort data in nested case-control and case-cohort studies by multiple imputation. Statistics in Medicine, 32(23):4021–4043, October 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tulupyev:2013:BPR

[1302] Alexander Tulupyev, Alena Suvorova, Jennifer Sousa, and Daniel Zelterman. Beta prime regression with application to risky behavior frequency screening. Statistics in Medicine, 32(23):4044–4056, October 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siewert:2013:MCU

[1303] Elizabeth Siewert and Katerina J. Kechris. Modeling considerations for using expression data from multiple species. Statistics in Medicine, 32 (23):4057–4070, October 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chung:2013:AZB

[1304] Yeojin Chung, Sophia Rabe-Hesketh, and In-Hee Choi. Avoiding zero between-study variance estimates in random-effects meta-analysis. Statistics in Medicine, 32(23):4071–4089, October 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oman:2013:SEU

[1305] Samuel D. Oman. Shrinkage estimation and the use of additional information when calibrating in the presence of random effects. Statistics in Medicine, 33(4):4090–4101, October 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2013:BAC

[1306] Jingyang Zhang, Kathryn Chaloner, James H. McLinden, and Jack T. Stapleton. Bayesian analysis and classification of two enzyme-linked immunosorbent assay tests without a gold standard. Statistics in Medicine, 33(4):4102–4117, October 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crowther:2013:SBP

[1307] Michael J. Crowther and Paul C. Lambert. Simulating biologically plausible complex survival data. Statistics in Medicine, 33(4):4118–4134, October 15, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Loong:2013:DCU

[1308] Bronwyn Loong, Alan M. Zaslavsky, Yulei He, and David P. Harrington. Disclosure control using partially synthetic data for large-scale health surveys, with applications to CanCORS. Statistics in Medicine, 32 (24):4139–4161, October 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Amatya:2013:SSD

[1309] Anup Amatya, Dulal Bhaumik, and Robert D. Gibbons. Sample size determination for clustered count data. Statistics in Medicine, 32(24):4162– 4179, October 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2013:BMI

[1310] Qingxia Chen, Ming-Hui Chen, David Ohlssen, and Joseph G. Ibrahim. Bayesian modeling and inference for clinical trials with partial retrieved data following dropout. Statistics in Medicine, 32(24):4180–4195, October 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demler:2013:ICP

[1311] Olga V. Demler, Michael J. Pencina, and Ralph B. D’Agostino, Sr. Impact of correlation on predictive ability of biomarkers. Statistics in Medicine, 32(24):4196–4210, October 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2013:ECM

[1312] Wei Wang, Suchitra Nelson, and Jeffrey M. Albert. Estimation of causal mediation effects for a dichotomous outcome in multiple-mediator models using the mediation formula. Statistics in Medicine, 32(24):4211–4228, October 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

An:2013:LFL

[1313] Xinming An, Qing Yang, and Peter M. Bentler. A latent factor linear mixed model for high-dimensional longitudinal data analysis. Statistics in Medicine, 32(24):4229–4239, October 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davies:2013:MDM

[1314] Tilman M. Davies, Jon Cornwall, and Philip W. Sheard. Modelling dichotomously marked muscle fibre configurations. Statistics in Medicine, 32(24):4240–4258, October 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DiRienzo:2013:PCM

[1315] A. Gregory DiRienzo. Parsimonious conditional-mean model selection with multiple covariates: an analysis of infant mortality in the USA. Statistics in Medicine, 32(24):4259–4274, October 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2013:CIB

[1316] Keunbaik Lee and Michael J. Daniels. Causal inference for bivariate longitudinal quality of life data in presence of death by using global odds ratios. Statistics in Medicine, 32(24):4275–4284, October 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2013:LAL

[1317] Chen-Hsin Chen, Yuh-Chyuan Tsay, Ya-Chi Wu, and Cheng-Fang Horng. Logistic-AFT location-scale mixture regression models with nonsusceptibility for left-truncated and general interval-censored data. Statistics in Medicine, 32(24):4285–4305, October 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2013:FMM

[1318] Jinsong Chen, Lei Liu, Daowen Zhang, and Ya-Chen T. Shih. A flexible model for the mean and variance functions, with application to medical cost data. Statistics in Medicine, 32(24):4306–4318, October 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2013:CIP

[1319] Stuart G. Baker. Causal inference, probability theory, and graphical insights. Statistics in Medicine, 32(25):4319–4330, November 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pearl:2013:CCI

[1320] Judea Pearl. Comment on ‘Causal inference, probability theory, and graphical insights’ by Stuart G. Baker. Statistics in Medicine, 33(11): 4331–4333, November 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2013:ATC

[1321] Stuart G. Baker. Additional thoughts on causal inference, probability theory, and graphical insights. Statistics in Medicine, 33(11):4334–4337, November 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuroki:2013:SBC

[1322] Manabu Kuroki. Sharp bounds on causal effects using a surrogate endpoint. Statistics in Medicine, 33(11):4338–4347, November 10, 2013. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bartolucci:2013:CIP

[1323] Francesco Bartolucci and Alessio Farcomeni. Causal inference in paired two-arm experimental studies under noncompliance with application to prognosis of myocardial infarction. Statistics in Medicine, 33(11):4348– 4366, November 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hou:2013:DRP

[1324] Wei Hou, Myron N. Chang, Sin-Ho Jung, and Yang Li. Designs for randomized phase II clinical trials with two treatment arms. Statistics in Medicine, 33(11):4367–4379, November 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2013:IPW

[1325] Richard J. Cook, Ker-Ai Lee, Meaghan Cuerden, and Cecilia A. Cotton. Inverse probability weighted estimating equations for randomized trials in transfusion medicine. Statistics in Medicine, 33(11):4380–4399, November 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shi:2013:EOC

[1326] Yun Shi and Guosheng Yin. Escalation with overdose control for phase I drug-combination trials. Statistics in Medicine, 33(11):4400–4412, November 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachin:2013:SSP

[1327] John M. Lachin. Sample size and power for a logrank test and Cox proportional hazards model with multiple groups and strata, or a quantitative covariate with multiple strata. Statistics in Medicine, 33(11): 4413–4425, November 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2013:BAS

[1328] Jing Zhang, Stephen R. Cole, David B. Richardson, and Haitao Chu. A Bayesian approach to strengthen inference for case-control studies with multiple error-prone exposure assessments. Statistics in Medicine, 33 (11):4426–4437, November 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murray:2013:GTM

[1329] K. Murray, S. Heritier, and S. Müller. Graphical tools for model selection in generalized linear models. Statistics in Medicine, 33(11):4438–4451, November 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McMahan:2013:RAC

[1330] Christopher S. McMahan, Lianming Wang, and Joshua M. Tebbs. Regression analysis for current status data using the EM algorithm. Statistics in Medicine, 33(11):4452–4466, November 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jafari:2013:WCR

[1331] Nahid Jafari, John Hearne, and Leonid Churilov. Why caution is recommended with post-hoc individual patient matching for estimation of treatment effect in parallel-group randomized controlled trials: The case of acute stroke trials. Statistics in Medicine, 33(11):4467–4481, November 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Won:2013:GFR

[1332] Sungho Won and Christoph Lange. A general framework for robust and efficient association analysis in family-based designs: quantitative and dichotomous phenotypes. Statistics in Medicine, 33(11):4482–4498, November 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burgess:2013:CMI

[1333] Stephen Burgess, Ian R. White, Matthieu Resche-Rigon, and Angela M. Wood. Combining multiple imputation and meta-analysis with individual participant data. Statistics in Medicine, 32(26):4499–4514, November 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hirakawa:2013:DFA

[1334] Akihiro Hirakawa, Chikuma Hamada, and Shigeyuki Matsui. A dosefinding approach based on shrunken predictive probability for combinations of two agents in phase I trials. Statistics in Medicine, 32(26): 4515–4525, November 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2013:RBP

[1335] X. Lu, D. V. Mehrotra, and B. E. Shepherd. Rank-based principal stratum sensitivity analyses. Statistics in Medicine, 32(26):4526–4539, November 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kahan:2013:BRF

[1336] Brennan C. Kahan. Bias in randomised factorial trials. Statistics in Medicine, 32(26):4540–4549, November 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Capanu:2013:AEM

[1337] Marinela Capanu, Mithat Gönen, and Colin B. Begg. An assessment of estimation methods for generalized linear mixed models with binary outcomes. Statistics in Medicine, 32(26):4550–4566, November 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tchetgen:2013:IOR

[1338] Eric J. Tchetgen Tchetgen. Inverse odds ratio-weighted estimation for causal mediation analysis. Statistics in Medicine, 32(26):4567–4580, November 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lange:2013:FIC

[1339] Jane M. Lange and Vladimir N. Minin. Fitting and interpreting continuous-time latent Markov models for panel data. Statistics in Medicine, 32(26):4581–4595, November 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellis:2013:PII

[1340] Jules L. Ellis. Probability interpretations of intraclass reliabilities. Statistics in Medicine, 32(26):4596–4608, November 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2013:MSS

[1341] Chen-Yen Lin and Susan Halabi. On model specification and selection of the Cox proportional hazards model. Statistics in Medicine, 32(26): 4609–4623, November 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carlson:2013:CMA

[1342] N. E. Carlson, K. W. Horton, and G. K. Grunwald. A comparison of methods for analyzing time series of pulsatile hormone data. Statistics in Medicine, 32(26):4624–4638, November 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wason:2013:UCD

[1343] James M. S. Wason and Shaun R. Seaman. Using continuous data on tumour measurements to improve inference in phase II cancer studies. Statistics in Medicine, 32(26):4639–4650, November 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mbougua:2013:NMI

[1344] Jules Brice Tchatchueng Mbougua, Christian Laurent, Ibra Ndoye, Eric Delaporte, Henri Gwet, and Nicolas Molinari. Nonlinear multiple imputation for continuous covariate within semiparametric Cox model: application to HIV data in Senegal. Statistics in Medicine, 32(26):4651–4665, November 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sperrin:2013:CRB

[1345] Matthew Sperrin, Lawrence Bardwell, Jamie C. Sergeant, Susan Astley, and Iain Buchan. Correcting for rater bias in scores on a continuous scale, with application to breast density. Statistics in Medicine, 32(26): 4666–4678, November 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schorgendorfer:2013:RAU

[1346] Angela Schörgendorfer and Adam J. Branscum. Regression analysis using dependent Polya trees. Statistics in Medicine, 32(27):4679–4695, November 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sjolander:2013:IMV

[1347] Arvid Sjölander and Sander Greenland. Ignoring the matching variables in cohort studies — when is it valid and why? Statistics in Medicine, 32(27):4696–4708, November 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2013:AME

[1348] Matthew T. White and Sharon X. Xie. Adjustment for measurement error in evaluating diagnostic biomarkers by using an internal reliability sample. Statistics in Medicine, 32(27):4709–4725, November 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burgess:2013:IOR

[1349] Stephen Burgess and Crp Chd Genetics Collaboration. Identifying the odds ratio estimated by a two-stage instrumental variable analysis with a logistic regression model. Statistics in Medicine, 32(27):4726–4747, November 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barrera-Gomez:2013:OCN

[1350] Jose Barrera-Gómez, Donna Spiegelman, and Xavier Basagaña. Optimal combination of number of participants and number of repeated measurements in longitudinal studies with time-varying exposure. Statistics in Medicine, 32(27):4748–4762, November 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2013:NEM

[1351] Baojiang Chen and Jing Qin. A new estimation with minimum trace of asymptotic covariance matrix for incomplete longitudinal data with a surrogate process. Statistics in Medicine, 32(27):4763–4780, November 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schemper:2013:ECB

[1352] Michael Schemper, Alexandra Kaider, Samo Wakounig, and Georg Heinze. Estimating the correlation of bivariate failure times under censoring. Statistics in Medicine, 32(27):4781–4790, November 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kodre:2013:ICR

[1353] Anamarija Rebolj Kodre and Maja Pohar Perme. Informative censoring in relative survival. Statistics in Medicine, 32(27):4791–4802, November 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mauguen:2013:CMS

[1354] Audrey Mauguen, Sandra Collette, Jean-Pierre Pignon, and Virginie Rondeau. Concordance measures in shared frailty models: application to clustered data in cancer prognosis. Statistics in Medicine, 32(27):4803– 4820, November 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fox:2013:MRI

[1355] J.-P. Fox, M. Avetisyan, and J. van der Palen. Mixture randomized item-response modeling: a smoking behavior validation study. Statistics in Medicine, 32(27):4821–4837, November 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagashima:2013:ERM

[1356] Kengo Nagashima, Yasunori Sato, Hisashi Noma, and Chikuma Hamada. An efficient and robust method for analyzing population pharmacokinetic data in genome-wide pharmacogenomic studies: a generalized estimating equation approach. Statistics in Medicine, 32(27):4838–4858, November 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schou:2013:MAC

[1357] I. Manjula Schou and Ian C. Marschner. Meta-analysis of clinical trials with early stopping: an investigation of potential bias. Statistics in Medicine, 32(28):4859–4874, December 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schlomer:2013:GSD

[1358] Patrick Schlömer and Werner Brannath. Group sequential designs for three-arm ‘gold standard’ non-inferiority trials with fixed margin. Statistics in Medicine, 32(28):4875–4889, December 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Resche-Rigon:2013:MIH

[1359] Matthieu Resche-Rigon, Ian R. White, Jonathan W. Bartlett, Sanne A. E. Peters, Simon G. Thompson, and on behalf of the Prog-Imt Study Group. Multiple imputation for handling systematically missing confounders in meta-analysis of individual participant data. Statistics in Medicine, 32(28):4890–4905, December 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kovalchik:2013:AHT

[1360] Stephanie A. Kovalchik, Ravi Varadhan, and Carlos O. Weiss. Assessing heterogeneity of treatment effect in a clinical trial with the proportional interactions model. Statistics in Medicine, 32(28):4906–4923, December 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Byun:2013:HMC

[1361] Jinyoung Byun, Dejian Lai, Sheng Luo, Jan Risser, Betty Tung, and Robert J. Hardy. A hybrid method in combining treatment effects from matched and unmatched studies. Statistics in Medicine, 32(28):4924– 4937, December 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quintana:2013:IVS

[1362] M. A. Quintana and D. V. Conti. Integrative variable selection via Bayesian model uncertainty. Statistics in Medicine, 32(28):4938–4953, December 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2013:RAM

[1363] Boan Zhang, Christopher R. Bilder, and Joshua M. Tebbs. Regression analysis for multiple-disease group testing data. Statistics in Medicine, 32(28):4954–4966, December 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sherwood:2013:WQR

[1364] Ben Sherwood, Lan Wang, and Xiao-Hua Zhou. Weighted quantile regression for analyzing health care cost data with missing covariates. Statistics in Medicine, 32(28):4967–4979, December 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tong:2013:GSH

[1365] Xingwei Tong, Liang Zhu, Chenlei Leng, Wendy Leisenring, and Leslie L. Robison. A general semiparametric hazards regression model: efficient estimation and structure selection. Statistics in Medicine, 32(28):4980– 4994, December 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

May:2013:CPM

[1366] Ryan C. May, Haitao Chu, Joseph G. Ibrahim, Michael G. Hudgens, Abigail C. Lees, and David M. Margolis. Change-point models to estimate the limit of detection. Statistics in Medicine, 32(28):4995–5007, December 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pozzi:2013:BAD

[1367] L. Pozzi, H. Schmidli, M. Gasparini, and A. Racine-Poon. A Bayesian adaptive dose selection procedure with an overdispersed count endpoint.

Statistics in Medicine, 32(28):5008–5027, December 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xanthakis:2013:MMV

[1368] Vanessa Xanthakis, Lisa M. Sullivan, and Ramachandran S. Vasan. Multilevel modeling versus cross-sectional analysis for assessing the longitudinal tracking of cardiovascular risk factors over time. Statistics in Medicine, 32(28):5028–5038, December 10, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Begg:2013:CMF

[1369] Colin B. Begg, Emily C. Zabor, Jonine L. Bernstein, Leslie Bernstein, Michael F. Press, and Venkatraman E. Seshan. A conceptual and methodological framework for investigating etiologic heterogeneity. Statistics in Medicine, 32(29):5039–5052, December 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bandyopadhyay:2013:CAR

[1370] Uttam Bandyopadhyay and Rahul Bhattacharya. A covariate-adjusted response-adaptive allocation in clinical trials for a general class of responses. Statistics in Medicine, 32(29):5053–5061, December 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heinze:2013:CIA

[1371] Georg Heinze, Meinhard Ploner, and Jan Beyea. Confidence intervals after multiple imputation: combining profile likelihood information from logistic regressions. Statistics in Medicine, 32(29):5062–5076, December 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2013:UCI

[1372] G. Y. Zou and Li Yue. Using confidence intervals to compare several correlated areas under the receiver operating characteristic curves. Statistics in Medicine, 32(29):5077–5090, December 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mori:2013:LCM

[1373] Joji Mori, Yutaka Kano, Masahiro Yoshizaki, and Satoru Fukinbara. Latent class models for medical diagnostic tests in multicenter trials. Statistics in Medicine, 32(29):5091–5105, December 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosychuk:2013:SSS

[1374] Rhonda J. Rosychuk and Hsing-Ming Chang. A spatial scan statistic for compound Poisson data. Statistics in Medicine, 32(29):5106–5118, December 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cottone:2013:AVP

[1375] Francesco Cottone, Fabio Efficace, Giovanni Apolone, and Gary S. Collins. The added value of propensity score matching when using healthrelated quality of life reference data. Statistics in Medicine, 32(29):5119– 5132, December 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2013:JAS

[1376] Sheng Luo. Joint analysis of stochastic processes with application to smoking patterns and insomnia. Statistics in Medicine, 32(29):5133– 5144, December 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Novick:2013:STS

[1377] Steven J. Novick. A simple test for synergy for a small number of combinations. Statistics in Medicine, 32(29):5145–5155, December 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nawarathna:2013:MAM

[1378] Lakshika S. Nawarathna and Pankaj K. Choudhary. Measuring agreement in method comparison studies with heteroscedastic measurements. Statistics in Medicine, 32(29):5156–5171, December 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dmitrienko:2013:TMA

[1379] Alex Dmitrienko and Ralph D’Agostino, Sr. Traditional multiplicity adjustment methods in clinical trials. Statistics in Medicine, 32(29): 5172–5218, December 20, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duez:2013:PMS

[1380] Nicole Duez, Dominique Eeckhoudt, Anne Kirkpatrick, Gabriel Solbu, Stefan Suciu, Richard Sylvester, Martine Van Glabbeke, Marc Buyse, Otilia Dalesio, Josette Renard, and Nicole Rotmensz. Professor Maurice Staquet, MD, MS (1930–2013). Statistics in Medicine, 32(30):5219–5220, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aalen:2013:MSS

[1381] Odd O. Aalen, Ørnulf Borgan, and Jan Terje Kvaløy. Medical statistics — a subject of increasing breadth and importance. Statistics in Medicine, 32(30):5221, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sorensen:2013:IMA

[1382] Helle Sørensen, Jeff Goldsmith, and Laura M. Sangalli. An introduction with medical applications to functional data analysis. Statistics in Medicine, 32(30):5222–5240, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mutsvari:2013:MMS

[1383] Timothy Mutsvari, Dipankar Bandyopadhyay, Dominique Declerck, and Emmanuel Lesaffre. A multilevel model for spatially correlated binary data in the presence of misclassification: an application in oral health research. Statistics in Medicine, 32(30):5241–5259, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haneuse:2013:EEI

[1384] S. Haneuse and A. Rotnitzky. Estimation of the effect of interventions that modify the received treatment. Statistics in Medicine, 32(30):5260– 5277, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2013:DNL

[1385] P. K. Andersen. Decomposition of number of life years lost according to causes of death. Statistics in Medicine, 32(30):5278–5285, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersson:2013:ELE

[1386] Therese M-L Andersson, Paul W. Dickman, Sandra Eloranta, Mats Lambe, and Paul C. Lambert. Estimating the loss in expectation of life due to cancer using flexible parametric survival models. Statistics in Medicine, 32(30):5286–5300, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rodriguez-Girondo:2013:MBN

[1387] Mar Rodríguez-Girondo, Thomas Kneib, Carmen Cadarso-Suárez, and Emad Abu-Assi. Model building in nonproportional hazard regression. Statistics in Medicine, 32(30):5301–5314, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Iacobelli:2013:MTS

[1388] Simona Iacobelli and Bendix Carstensen. Multiple time scales in multistate models. Statistics in Medicine, 32(30):5315–5327, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stoer:2013:IPW

[1389] Nathalie C. Støer and Sven Ove Samuelsen. Inverse probability weighting in nested case-control studies with additional matching — a simulation study. Statistics in Medicine, 32(30):5328–5339, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2013:SPL

[1390] Donghwan Lee, Youngjo Lee, Yudi Pawitan, and Woojoo Lee. Sparse partial least-squares regression for high-throughput survival data analysis. Statistics in Medicine, 32(30):5340–5352, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2013:UME

[1391] Myeongjee Lee, Paola Rebora, Maria Grazia Valsecchi, Kamila Czene, and Marie Reilly. A unified model for estimating and testing familial aggregation. Statistics in Medicine, 32(30):5353–5365, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mauguen:2013:DPR

[1392] Audrey Mauguen, Bernard Rachet, Simone Mathoulin-Pélissier, Gaetan MacGrogan, Alexandre Laurent, and Virginie Rondeau. Dynamic prediction of risk of death using history of cancer recurrences in joint frailty models. Statistics in Medicine, 32(30):5366–5380, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blanche:2013:ECT

[1393] Paul Blanche, Jean-François Dartigues, and Hélène Jacqmin-Gadda. Estimating and comparing time-dependent areas under receiver operating characteristic curves for censored event times with competing risks.

Statistics in Medicine, 32(30):5381–5397, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Menten:2013:BMA

[1394] J. Menten, M. Boelaert, and E. Lesaffre. Bayesian meta-analysis of diagnostic tests allowing for imperfect reference standards. Statistics in Medicine, 32(30):5398–5413, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Konig:2013:VFE

[1395] Jochem König, Ulrike Krahn, and Harald Binder. Visualizing the flow of evidence in network meta-analysis and characterizing mixed treatment comparisons. Statistics in Medicine, 32(30):5414–5429, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Doussau:2013:DFD

[1396] Adélaïde Doussau, Rodolphe Thiébaut, and Xavier Paoletti. Dosefinding design using mixed-effect proportional odds model for longitudinal graded toxicity data in phase I oncology clinical trials. Statistics in Medicine, 32(30):5430–5447, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schneider:2013:BSS

[1397] S. Schneider, H. Schmidli, and T. Friede. Blinded sample size reestimation for recurrent event data with time trends. Statistics in Medicine, 32(30):5448–5457, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bakhshi:2013:SIP

[1398] Andisheh Bakhshi, Stephen Senn, and Alan Phillips. Some issues in predicting patient recruitment in multi-centre clinical trials. Statistics in Medicine, 32(30):5458–5468, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jaki:2013:EBC

[1399] Thomas Jaki, Philip Pallmann, and Martin J. Wolfsegger. Estimation in AB/BA crossover trials with application to bioequivalence studies with incomplete and complete data designs. Statistics in Medicine, 32(30): 5469–5483, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2013:CEB

[1400] Zhiwu Yan. Comments on ‘Estimation in AB/BA crossover trials with application to bioequivalence studies with incomplete and complete data designs’. Statistics in Medicine, 32(30):5484–5486, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jaki:2013:ARC

[1401] Thomas Jaki, Philip Pallmann, and Martin J. Wolfsegger. Authors’ reply to Comments on ‘Estimation in AB/BA crossover trials with application to bioequivalence studies with incomplete and complete data designs’. Statistics in Medicine, 32(30):5487–5488, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lesaffre:2013:C

[1402] Emmanuel Lesaffre. Correction. Statistics in Medicine, 32(30):5489, December 30, 2013. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2014:EN

[1403] Anonymous. Editors’ note. Statistics in Medicine, 33(1):1, January 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Braun:2014:ABD

[1404] Julia Braun, Leonhard Held, Bruno Ledergerber, and the Swiss Hiv Cohort Study. Accounting for baseline differences and measurement error in the analysis of change over time. Statistics in Medicine, 33(1):2–16, January 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuss:2014:MAD

[1405] Oliver Kuss, Annika Hoyer, and Alexander Solms. Meta-analysis for diagnostic accuracy studies: a new statistical model using beta-binomial distributions and bivariate copulas. Statistics in Medicine, 33(1):17–30, January 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ren:2014:ACP

[1406] Shijie Ren and Jeremy E. Oakley. Assurance calculations for planning clinical trials with time-to-event outcomes. Statistics in Medicine, 33(1): 31–45, January 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Su:2014:MRA

[1407] Pei-Fang Su and Yunchan Chi. Marginal regression approach for additive hazards models with clustered current status data. Statistics in Medicine, 33(1):46–58, January 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2014:ISM

[1408] Gang Han, Michael J. Schell, and Jongphil Kim. Improved survival modeling in cancer research using a reduced piecewise exponential approach. Statistics in Medicine, 33(1):59–73, January 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hade:2014:BAU

[1409] Erinn M. Hade and Bo Lu. Bias associated with using the estimated propensity score as a regression covariate. Statistics in Medicine, 33(1): 74–87, January 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morris:2014:MII

[1410] Tim P. Morris, Ian R. White, Patrick Royston, Shaun R. Seaman, and Angela M. Wood. Multiple imputation for an incomplete covariate that is a ratio. Statistics in Medicine, 33(1):88–104, January 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2014:FPA

[1411] Katherine J. Lee and John B. Carlin. Fractional polynomial adjustment for time-varying covariates in a self-controlled case series analysis. Statistics in Medicine, 33(1):105–116, January 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carrasco:2014:DAC

[1412] Josep L. Carrasco, Alejandro Caceres, Georgia Escaramis, and Lluis Jover. Distinguishability and agreement with continuous data. Statistics in Medicine, 33(1):117–128, January 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schomaker:2014:NIL

[1413] M. Schomaker, T. Gsponer, J. Estill, M. Fox, and A. Boulle. Nonignorable loss to follow-up: correcting mortality estimates based on additional outcome ascertainment. Statistics in Medicine, 33(1):129–142, January 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2014:AMF

[1414] P. Wu, X. M. Tu, and J. Kowalski. On assessing model fit for distributionfree longitudinal models under missing data. Statistics in Medicine, 33 (1):143–157, January 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sala:2014:PPE

[1415] Sara Sala, Piero Quatto, Paola Valsasina, Federica Agosta, and Massimo Filippi. pFDR and pFNR estimation for brain networks construction. Statistics in Medicine, 33(1):158–169, January 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perperoglou:2014:CMD

[1416] Aris Perperoglou. Cox models with dynamic ridge penalties on timevarying effects of the covariates. Statistics in Medicine, 33(1):170–180, January 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gouskova:2014:NNT

[1417] Natalia A. Gouskova, Suprateek Kundu, Peter B. Imrey, and Jason P. Fine. Number needed to treat for time-to-event data with competing risks. Statistics in Medicine, 33(2):181–192, January 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leuchs:2014:ETE

[1418] Ann-Kristin Leuchs, Jörg Zinserling, Gabriele Schlosser-Weber, Manfred Berres, Markus Neuhäuser, and Norbert Benda. Estimation of the treatment effect in the presence of non-compliance and missing data. Statistics in Medicine, 33(2):193–208, January 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schuemie:2014:IOS

[1419] Martijn J. Schuemie, Patrick B. Ryan, William DuMouchel, Marc A. Suchard, and David Madigan. Interpreting observational studies: why empirical calibration is needed to correct p-values. Statistics in Medicine, 33(2):209–218, January 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dusseldorp:2014:QIT

[1420] Elise Dusseldorp and Iven Van Mechelen. Qualitative interaction trees: a tool to identify qualitative treatment-subgroup interactions. Statistics in Medicine, 33(2):219–237, January 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ramezan:2014:MAN

[1421] Reza Ramezan, Paul Marriott, and Shojaeddin Chenouri. Multiscale analysis of neural spike trains. Statistics in Medicine, 33(2):238–256, January 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taylor:2014:CME

[1422] Jeremy M. G. Taylor, Jincheng Shen, Edward H. Kennedy, Lu Wang, and Douglas E. Schaubel. Comparison of methods for estimating the effect of salvage therapy in prostate cancer when treatment is given by indication. Statistics in Medicine, 33(2):257–274, January 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elie:2014:MCE

[1423] Caroline Elie, Paul Landais, and Yann De Rycke. A model combining excess and relative mortality for population-based studies. Statistics in Medicine, 33(2):275–288, January 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jaspers:2014:EWT

[1424] Stijn Jaspers, Marc Aerts, Geert Verbeke, and Pierre-Alexandre Beloeil. Estimation of the wild-type minimum inhibitory concentration value distribution. Statistics in Medicine, 33(2):289–303, January 30, 2014. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2014:TGE

[1425] Weiming Zhang, Carl D. Langefeld, Gary K. Grunwald, and Tasha E. Fingerlin. Testing gene-environment interactions in family-based association studies using trait-based ascertained samples. Statistics in Medicine, 33(2):304–318, January 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roislien:2014:NPP

[1426] Jo Røislien and Eigil Samset. A non-parametric permutation method for assessing agreement for distance matrix observations. Statistics in Medicine, 33(2):319–329, January 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hillis:2014:MMA

[1427] Stephen L. Hillis. A marginal-mean ANOVA approach for analyzing multireader multicase radiological imaging data. Statistics in Medicine, 33(2):330–360, January 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2014:AIC

[1428] Zhehui Luo, Joshua Breslau, Joseph C. Gardiner, Qiaoling Chen, and Naomi Breslau. Assessing interchangeability at cluster levels with multiple-informant data. Statistics in Medicine, 33(3):361–375, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2014:SSC

[1429] Haiyuan Zhu and Hassan Lakkis. Sample size calculation for comparing two negative binomial rates. Statistics in Medicine, 33(3):376–387, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bowden:2014:EBE

[1430] Jack Bowden, Werner Brannath, and Ekkehard Glimm. Empirical Bayes estimation of the selected treatment mean for two-stage drop-the-loser trials: a meta-analytic approach. Statistics in Medicine, 33(3):388–400, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2014:CSV

[1431] Feihan Lu and Eva Petkova. A comparative study of variable selection methods in the context of developing psychiatric screening instruments. Statistics in Medicine, 33(3):401–421, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turner:2014:IEU

[1432] Elizabeth L. Turner, Michael J. Sweeting, Robert J. Lindfield, and Daniela DeAngelis. Incidence estimation using a single cross-sectional age-specific prevalence survey with differential mortality. Statistics in Medicine, 33(3):422–435, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hendry:2014:DGC

[1433] David J. Hendry. Data generation for the Cox proportional hazards model with time-dependent covariates: a method for medical researchers.

Statistics in Medicine, 33(3):436–454, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boonstra:2014:PSB

[1434] Philip S. Boonstra, Irina Bondarenko, Sung Kyun Park, Pantel S. Vokonas, and Bhramar Mukherjee. Propensity score-based diagnostics for categorical response regression models. Statistics in Medicine, 33(3): 455–469, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Strand:2014:RCM

[1435] Matthew Strand, Stefan Sillau, Gary K. Grunwald, and Nathan Rabinovitch. Regression calibration for models with two predictor variables measured with error and their interaction, using instrumental variables and longitudinal data. Statistics in Medicine, 33(3):470–487, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smits:2014:VLT

[1436] Niels Smits and Matthew D. Finkelman. Variable length testing using the ordinal regression model. Statistics in Medicine, 33(3):488–499, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2014:OTP

[1437] Huiping Xu, Siu L. Hui, and Shaun Grannis. Optimal two-phase sampling design for comparing accuracies of two binary classification rules. Statistics in Medicine, 33(3):500–513, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:2014:SNJ

[1438] Dai Feng, Richard Baumgartner, and Vladimir Svetnik. A short note on jackknifing the concordance correlation coefficient. Statistics in Medicine, 33(3):514–516, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2014:GAI

[1439] Peter C. Austin and Ewout W. Steyerberg. Graphical assessment of internal and external calibration of logistic regression models by using loess smoothers. Statistics in Medicine, 33(3):517–535, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmidt:2014:CUP

[1440] Amand F. Schmidt, Arno W. Hoes, and Rolf H. H. Groenwold. Comments on ‘The use of propensity scores and observational data to estimate randomized controlled trial generalizability bias’ by Taylor R. Pressler and Eloise E. Kaizar, Statistics in Medicine 2013. Statistics in Medicine, 33(3):536–537, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaizar:2014:RCU

[1441] Eloise E. Kaizar and Taylor Pressler Vydra. Response to comments on ‘the use of propensity scores and observational data to estimate randomized controlled trial generalizability bias’. Statistics in Medicine, 33(3): 538–539, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pullenayegum:2014:CDR

[1442] Eleanor M. Pullenayegum. Correction for “Doubly robust estimation, optimally truncated inverse-intensity weighting and increment-based methods for the analysis of irregularly observed longitudinal data” by Eleanor Pullenayegum and Brian Feldman. Statistics in Medicine, 33(3):540, February 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2014:RMM

[1443] Daniel Jackson and Richard D. Riley. A refined method for multivariate meta-analysis and meta-regression. Statistics in Medicine, 33(4):541– 554, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2014:IPW

[1444] Changyu Shen, Xiaochun Li, and Lingling Li. Inverse probability weighting for covariate adjustment in randomized studies. Statistics in Medicine, 33(17):555–568, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wages:2014:PDC

[1445] Nolan A. Wages, John O’Quigley, and Mark R. Conaway. Phase I design for completely or partially ordered treatment schedules. Statistics in Medicine, 33(17):569–579, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2014:BAJ

[1446] Sheng Luo. A Bayesian approach to joint analysis of multivariate longitudinal data and parametric accelerated failure time. Statistics in Medicine, 33(17):580–594, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2014:APC

[1447] Bei Jiang and Keumhee C. Carriere. Age–period-cohort models using smoothing splines: a generalized additive model approach. Statistics in Medicine, 33(17):595–606, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liang:2014:MOR

[1448] Yuanyuan Liang, Yin Li, Jing Wang, and Keumhee C. Carriere. Multipleobjective response-adaptive repeated measurement designs in clinical trials for binary responses. Statistics in Medicine, 33(17):607–617, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choi:2014:ESC

[1449] Yun-Hee Choi, Laurent Briollais, Jane Green, Patrick Parfrey, and Karen Kopciuk. Estimating successive cancer risks in Lynch Syndrome families using a progressive three-state model. Statistics in Medicine, 33(17):618– 638, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghebremichael-Weldeselassie:2014:SCC

[1450] Yonas Ghebremichael-Weldeselassie, Heather J. Whitaker, and C. Paddy Farrington. Self-controlled case series method with smooth age effect. Statistics in Medicine, 33(17):639–649, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2014:BRT

[1451] Lili Zhao, Dai Feng, Emily L. Bellile, and Jeremy M. G. Taylor. Bayesian random threshold estimation in a Cox proportional hazards cure model. Statistics in Medicine, 33(17):650–661, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baek:2014:HMI

[1452] Jonggyu Baek, Brisa N. Sánchez, and Emma V. Sanchez-Vaznaugh. Hierarchical multiple informants models: examining food environment contributions to the childhood obesity epidemic. Statistics in Medicine, 33 (17):662–674, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tapsoba:2014:EEE

[1453] Jean de Dieu Tapsoba, Shen-Ming Lee, and Ching-Yun Wang. Expected estimating equation using calibration data for generalized linear models with a mixture of Berkson and classical errors in covariates. Statistics in Medicine, 33(17):675–692, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alosh:2014:AMA

[1454] Mohamed Alosh, Frank Bretz, and Mohammad Huque. Advanced multiplicity adjustment methods in clinical trials. Statistics in Medicine, 33 (17):693–713, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanKlaveren:2014:ICM

[1455] David van Klaveren, Ewout W. Steyerberg, and Yvonne Vergouwe. Interpretation of concordance measures for clustered data. Statistics in Medicine, 33(17):714–716, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mauguen:2014:RIC

[1456] Audrey Mauguen, Sandra Collette, Jean-Pierre Pignon, and Virginie Rondeau. Reply to ‘interpretation of concordance measures for clustered data’. Statistics in Medicine, 33(17):717–718, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lendrem:2014:BRN

[1457] Dennis Lendrem. Book review: The Norm Chronicles: stories and numbers about danger by Blastland, Michael and Spiegelhalter, David, Profile Books Ltd, London, 2013. No. of pages: 288. Price: £12.99. ISBN: 978-1-84668-620-7. Statistics in Medicine, 33(17):719–720, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chung:2014:C

[1458] Yeojin Chung. Correction. Statistics in Medicine, 33(17):720, February 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williamson:2014:VRR

[1459] Elizabeth J. Williamson, Andrew Forbes, and Ian R. White. Variance reduction in randomised trials by inverse probability weighting using the propensity score. Statistics in Medicine, 33(5):721–737, February 28, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Iosif:2014:MRC

[1460] Ana-Maria Iosif and Allan R. Sampson. A model for repeated clustered data with informative cluster sizes. Statistics in Medicine, 33(5):738– 759, February 28, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2014:GLT

[1461] Zhiguo Li, Marcia Valenstein, Paul Pfeiffer, and Dara Ganoczy. A global logrank test for adaptive treatment strategies based on observational studies. Statistics in Medicine, 33(5):760–771, February 28, 2014. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2014:RAM

[1462] Chyong-Mei Chen, James Cheng-Chung Wei, Chao-Min Hsu, and Ming-Yung Lee. Regression analysis of multivariate current status data with dependent censoring: application to ankylosing spondylitis data. Statistics in Medicine, 33(5):772–785, February 28, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2014:NAA

[1463] Yuh-Ing Chen and Chi-Shen Huang. New approach to assess bioequivalence parameters using generalized gamma mixed-effect model (modelbased asymptotic bioequivalence test). Statistics in Medicine, 33(5):786– 797, February 28, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laubender:2014:NCA

[1464] Ruediger P. Laubender and Ralf Bender. A note on calculating asymptotic confidence intervals for the adjusted risk difference and number needed to treat in the Cox regression model. Statistics in Medicine, 33 (5):798–810, February 28, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laubender:2014:C

[1465] Ruediger P. Laubender. Correction. Statistics in Medicine, 33(5):810– 811, February 28, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chi:2014:PCO

[1466] Yueh-Yun Chi, Matthew J. Gribbin, Jacqueline L. Johnson, and Keith E. Muller. Power calculation for overall hypothesis testing with highdimensional commensurate outcomes. Statistics in Medicine, 33(5):812– 827, February 28, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mendolia:2014:CSA

[1467] Franco Mendolia, John P. Klein, Effie W. Petersdorf, Mari Malkki, and Tao Wang. Comparison of statistics in association tests of genetic markers for survival outcomes. Statistics in Medicine, 33(5):828–844, February 28, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Havulinna:2014:BAP

[1468] Aki S. Havulinna. Bayesian age-period-cohort models with versatile interactions and long-term predictions: mortality and population in Finland 1878–2050. Statistics in Medicine, 33(5):845–856, February 28, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:2014:OMP

[1469] John N. S. Matthews. An optimal multi-period crossover design for an application in paediatric nephrology. Statistics in Medicine, 33(5):857– 866, February 28, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Epifanio:2014:HSA

[1470] Irene Epifanio and Noelia Ventura-Campos. Hippocampal shape analysis in Alzheimer’s disease using functional data analysis. Statistics in Medicine, 33(5):867–880, February 28, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gasparrini:2014:MEL

[1471] Antonio Gasparrini. Modeling exposure-lag-response associations with distributed lag non-linear models. Statistics in Medicine, 33(5):881–899, February 28, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gasparrini:2014:C

[1472] A. Gasparrini. Correction. Statistics in Medicine, 33(5):900, February 28, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilbert:2014:OAC

[1473] Peter B. Gilbert, Xuesong Yu, and Andrea Rotnitzky. Optimal auxiliarycovariate-based two-phase sampling design for semiparametric efficient estimation of a mean or mean difference, with application to clinical trials. Statistics in Medicine, 33(6):901–917, March 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2014:FNR

[1474] Qin Wu and Man-Lai Tang. Flexible non-randomized response models for survey with sensitive question. Statistics in Medicine, 33(6):918–929, March 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Inoue:2014:MGP

[1475] Lurdes Y. T. Inoue, Bruce J. Trock, Alan W. Partin, Herbert B. Carter, and Ruth Etzioni. Modeling grade progression in an active surveillance study. Statistics in Medicine, 33(6):930–939, March 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:2014:BAD

[1476] A. Lawrence Gould and Xiaohua Douglas Zhang. Bayesian adaptive determination of the sample size required to assure acceptably low adverse event risk. Statistics in Medicine, 33(6):940–957, March 15, 2014. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2014:SSE

[1477] H. Li and C. Gatsonis. Sample size estimation for time-dependent receiver operating characteristic. Statistics in Medicine, 33(6):958–970, March 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harun:2014:BRE

[1478] Nusrat Harun and Bo Cai. Bayesian random effects selection in mixed accelerated failure time model for interval-censored data. Statistics in Medicine, 33(6):971–984, March 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yin:2014:JCR

[1479] Jingjing Yin and Lili Tian. Joint confidence region estimation for area under ROC curve and Youden index. Statistics in Medicine, 33(6):985– 1000, March 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Young:2014:SKC

[1480] Jessica G. Young and Eric J. Tchetgen Tchetgen. Simulation from a known Cox MSM using standard parametric models for the g-formula. Statistics in Medicine, 33(6):1001–1014, March 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lai:2014:IMC

[1481] Yinglei Lai and Paul S. Albert. Identifying multiple change points in a linear mixed effects model. Statistics in Medicine, 33(6):1015–1028, March 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sommer:2014:SRE

[1482] Julia C. Sommer, Jan Gertheiss, and Volker J. Schmid. Spatially regularized estimation for the analysis of dynamic contrast-enhanced magnetic resonance imaging data. Statistics in Medicine, 33(6):1029–1041, March 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:2014:CPH

[1483] H. Campbell and C. B. Dean. The consequences of proportional hazards based model selection. Statistics in Medicine, 33(6):1042–1056, March 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2014:CAM

[1484] Peter C. Austin. A comparison of 12 algorithms for matching on the propensity score. Statistics in Medicine, 33(6):1057–1069, March 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quaresma:2014:FPP

[1485] M. Quaresma, M. P. Coleman, and B. Rachet. Funnel plots for population-based cancer survival: principles, methods and applications. Statistics in Medicine, 33(6):1070–1080, March 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2014:ACH

[1486] Yulei He, Frederic Selck, and Sharon-Lise T. Normand. On the accuracy of classifying hospitals on their performance measures. Statistics in Medicine, 33(7):1081–1103, March 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rauch:2014:OCC

[1487] Geraldine Rauch, Antje Jahn-Eimermacher, Werner Brannath, and Meinhard Kieser. Opportunities and challenges of combined effect measures based on prioritized outcomes. Statistics in Medicine, 33(7):1104– 1120, March 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ding:2014:ISC

[1488] Peng Ding and Zhi Geng. Identifiability of subgroup causal effects in randomized experiments with nonignorable missing covariates. Statistics in Medicine, 33(7):1121–1133, March 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2014:AMP

[1489] Kaifeng Lu. An analytic method for the placebo-based pattern-mixture model. Statistics in Medicine, 33(7):1134–1145, March 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Darssan:2014:IAE

[1490] Darsy Darssan, Mery H. Thompson, and Anthony N. Pettitt. Incorporating adverse event relatedness into dose-finding clinical trial designs. Statistics in Medicine, 33(7):1146–1161, March 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hwang:2014:SBJ

[1491] Beom Seuk Hwang and Michael L. Pennell. Semiparametric Bayesian joint modeling of a binary and continuous outcome with applications in toxicological risk assessment. Statistics in Medicine, 33(7):1162–1175, March 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Azmon:2014:ERN

[1492] Amin Azmon, Christel Faes, and Niel Hens. On the estimation of the reproduction number based on misreported epidemic data. Statistics in Medicine, 33(7):1176–1192, March 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scheike:2014:ETC

[1493] Thomas H. Scheike, Klaus K. Holst, and Jacob B. Hjelmborg. Estimating twin concordance for bivariate competing risks twin data. Statistics in Medicine, 33(7):1193–1204, March 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heroux:2014:MSM

[1494] Julie Héroux, Erica E. M. Moodie, Erin Strumpf, Natalie Coyle, Pierre Tousignant, and Mamadou Diop. Marginal structural models for skewed outcomes: identifying causal relationships in health care utilization. Statistics in Medicine, 33(7):1205–1221, March 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldfeld:2014:TWM

[1495] K. S. Goldfeld. Twice-weighted multiple interval estimation of a marginal structural model to analyze cost-effectiveness. Statistics in Medicine, 33 (7):1222–1241, March 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2014:UPS

[1496] Peter C. Austin. The use of propensity score methods with survival or time-to-event outcomes: reporting measures of effect similar to those used in randomized experiments. Statistics in Medicine, 33(7):1242–1258, March 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Doi:2014:LEE

[1497] Suhail A. R. Doi. Letter to the Editor: Empirical versus natural weighting in random effects meta-analysis by J. J. Shuster, Statistics in Medicine 2010; 29 (12):1259–65. Statistics in Medicine, 33(7):1259, March 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shuster:2014:EVN

[1498] Jonathan J. Shuster. Empirical versus natural weighting in random effects meta-analysis. Statistics in Medicine, 33(7):1260, March 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2014:CIM

[1499] P. Wu, Y. Han, T. Chen, and X. M. Tu. Causal inference for Mann– Whitney–Wilcoxon rank sum and other nonparametric statistics. Statistics in Medicine, 33(8):1261–1271, April 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dong:2014:VSA

[1500] Gaohong Dong. A varying-stage adaptive phase II/III clinical trial design. Statistics in Medicine, 33(8):1272–1287, April 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2014:MLQ

[1501] Wanjie Sun, Michael D. Larsen, and John M. Lachin. Methods for a longitudinal quantitative outcome with a multivariate Gaussian distribution multi-dimensionally censored by therapeutic intervention. Statistics in Medicine, 33(8):1288–1306, April 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dong:2014:CML

[1502] Ting Dong, Catherine Chunling Liu, Emanuel F. Petricoin, and Liansheng Larry Tang. Combining markers with and without the limit of detection. Statistics in Medicine, 33(8):1307–1320, April 15, 2014. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xi:2014:GMP

[1503] Dong Xi and Ajit C. Tamhane. A general multistage procedure for k-out-of-n gatekeeping. Statistics in Medicine, 33(8):1321–1335, April 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jennen-Steinmetz:2014:SSD

[1504] Christine Jennen-Steinmetz. Sample size determination for studies designed to estimate covariate-dependent reference quantile curves. Statistics in Medicine, 33(8):1336–1348, April 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:2014:UBH

[1505] Dai Feng, Dong Liang, and Luke Tierney. A unified Bayesian hierarchical model for MRI tissue classification. Statistics in Medicine, 33(8):1349– 1368, April 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2014:NMA

[1506] Yanyuan Ma and Yuanjia Wang. Nonparametric modeling and analysis of association between Huntington’s disease onset and CAG repeats. Statistics in Medicine, 33(8):1369–1382, April 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Krafty:2014:NSA

[1507] Robert T. Krafty, Mengyuan Zhao, Daniel J. Buysse, Julian F. Thayer, and Martica Hall. Nonparametric spectral analysis of heart rate variability through penalized sum of squares. Statistics in Medicine, 33(8): 1383–1394, April 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2014:BMH

[1508] Yue Zhang and Kiros Berhane. Bayesian mixed hidden Markov models: a multi-level approach to modeling categorical outcomes with differential misclassification. Statistics in Medicine, 33(8):1395–1408, April 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moodie:2014:MSM

[1509] Erica E. M. Moodie, David A. Stephens, and Marina B. Klein. A marginal structural model for multiple-outcome survival data:assessing the impact of injection drug use on several causes of death in the Canadian Co-infection Cohort. Statistics in Medicine, 33(8):1409–1425, April 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yin:2014:OLC

[1510] Jingjing Yin and Lili Tian. Optimal linear combinations of multiple diagnostic biomarkers based on Youden index. Statistics in Medicine, 33(8):1426–1440, April 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Trikalinos:2014:ECU

[1511] Thomas A. Trikalinos, David C. Hoaglin, and Christopher H. Schmid. An empirical comparison of univariate and multivariate meta-analyses for categorical outcomes. Statistics in Medicine, 33(9):1441–1459, April 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ravva:2014:LAM

[1512] Patanjali Ravva, Mats O. Karlsson, and Jonathan L. French. A linearization approach for the model-based analysis of combined aggregate and individual patient data. Statistics in Medicine, 33(9):1460–1476, April 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kitsche:2014:TQI

[1513] Andreas Kitsche and Ludwig A. Hothorn. Testing for qualitative interaction using ratios of treatment differences. Statistics in Medicine, 33(9): 1477–1489, April 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brumback:2014:USN

[1514] Babette A. Brumback, Zhulin He, Mansi Prasad, Matthew C. Freeman, and Richard Rheingans. Using structural-nested models to estimate the effect of cluster-level adherence on individual-level outcomes with a threearmed cluster-randomized trial. Statistics in Medicine, 33(9):1490–1502, April 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kidwell:2014:IMR

[1515] Kelley M. Kidwell, Jin H. Ko, and Abdus S. Wahed. Inference for the median residual life function in sequential multiple assignment randomized trials. Statistics in Medicine, 33(9):1503–1513, April 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuznetsova:2014:WBT

[1516] Olga M. Kuznetsova and Yevgen Tymofyeyev. Wide brick tunnel randomization — an unequal allocation procedure that limits the imbalance in treatment totals. Statistics in Medicine, 33(9):1514–1530, April 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ahn:2014:PVA

[1517] Kwang Woo Ahn and Franco Mendolia. Pseudo-value approach for comparing survival medians for dependent data. Statistics in Medicine, 33(9): 1531–1538, April 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fang:2014:MDT

[1518] Fang Fang, Yong Lin, Weichung J. Shih, Yulin Li, Jay Yang, and Xiaosha Zhang. Methods of designing two-stage winner trials with survival outcomes. Statistics in Medicine, 33(9):1539–1563, April 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Overstall:2014:ICT

[1519] Antony M. Overstall, Ruth King, Sheila M. Bird, Sharon J. Hutchinson, and Gordon Hay. Incomplete contingency tables with censored cells with application to estimating the number of people who inject drugs in Scotland. Statistics in Medicine, 33(9):1564–1579, April 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weidemann:2014:BPI

[1520] Felix Weidemann, Manuel Dehnert, Judith Koch, Ole Wichmann, and Michael Höhle. Bayesian parameter inference for dynamic infectious disease modelling: rotavirus in Germany. Statistics in Medicine, 33(9): 1580–1599, April 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2014:BSM

[1521] Ming-Hui Chen, Joseph G. Ibrahim, H. Amy Xia, Thomas Liu, and Violeta Hennessey. Bayesian sequential meta-analysis design in evaluating cardiovascular risk in a new antidiabetic drug development program. Statistics in Medicine, 33(9):1600–1618, April 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teixeira-Pinto:2014:BRM

[1522] Armando Teixeira-Pinto. Book review: Medical statistics in clinical and epidemiological research by Veierød, Marit B., Lydersen, Stian and Laake, Petter, Gyldendal Norsk Forlag, Oslo, 2012. No. of pages: 714. Price: 575 Norwegian Kroner. ISBN: 978-82-05-39959-4. Statistics in Medicine, 33(9):1619–1620, April 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hudson:2014:SMM

[1523] Harold M. Hudson, Serigne N. Lô, R. John Simes, Andrew M. Tonkin, and Stephane Heritier. Semiparametric methods for multistate survival models in randomised trials. Statistics in Medicine, 33(10):1621–1645, May 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pinheiro:2014:MBD

[1524] José Pinheiro, Björn Bornkamp, Ekkehard Glimm, and Frank Bretz. Model-based dose finding under model uncertainty using general parametric models. Statistics in Medicine, 33(10):1646–1661, May 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ceyhan:2014:SID

[1525] Elvan Ceyhan. Segregation indices for disease clustering. Statistics in Medicine, 33(10):1662–1684, May 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Franklin:2014:MCB

[1526] Jessica M. Franklin, Jeremy A. Rassen, Diana Ackermann, Dorothee B. Bartels, and Sebastian Schneeweiss. Metrics for covariate balance in cohort studies of causal effects. Statistics in Medicine, 33(10):1685–1699, May 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shao:2014:SVC

[1527] Fang Shao, Jialiang Li, Shuangge Ma, and Mei-Ling Ting Lee. Semiparametric varying-coefficient model for interval censored data with a cured proportion. Statistics in Medicine, 33(10):1700–1712, May 10, 2014. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2014:UEL

[1528] Baojiang Chen and Jing Qin. Use of empirical likelihood to calibrate auxiliary information in partly linear monotone regression models. Statistics in Medicine, 33(10):1713–1722, May 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2014:LVM

[1529] Xinyuan Song, Zhaohua Lu, and Xiangnan Feng. Latent variable models with nonparametric interaction effects of latent variables. Statistics in Medicine, 33(10):1723–1737, May 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maity:2014:AVF

[1530] Arnab Maity, Paige L. Williams, Louise Ryan, Stacey A. Missmer, Brent A. Coull, and Russ Hauser. Analysis of in vitro fertilization data with multiple outcomes using discrete time-to-event analysis. Statistics in Medicine, 33(10):1738–1749, May 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Conlon:2014:MSM

[1531] A. S. C. Conlon, J. M. G. Taylor, and D. J. Sargent. Multi-state models for colon cancer recurrence and death with a cured fraction. Statistics in Medicine, 33(10):1750–1766, May 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mahiane:2014:MMC

[1532] Severin Guy Mahiane, Agnès Fiamma, and Bertran Auvert. Mixture models for calibrating the BED for HIV incidence testing. Statistics in Medicine, 33(10):1767–1783, May 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2014:DCT

[1533] Jichen Yang, Xinlei Wang, Minsoo Kim, Yang Xie, and Guanghua Xiao. Detection of candidate tumor driver genes using a fully integrated Bayesian approach. Statistics in Medicine, 33(10):1784–1800, May 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2014:VIM

[1534] Maggie Hong Chen and Andrew R. Willan. Value of information methods for assessing a new diagnostic test. Statistics in Medicine, 33(11):1801– 1815, May 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheng:2014:PAT

[1535] K. F. Cheng, J. Y. Lee, W. Zheng, and C. Li. A powerful association test of multiple genetic variants using a random-effects model. Statistics in Medicine, 33(11):1816–1827, May 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freytag:2014:CTS

[1536] Saskia Freytag and Heike Bickeböller. Comparison of three summary statistics for ranking genes in genome-wide association studies. Statistics in Medicine, 33(11):1828–1841, May 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salim:2014:MLM

[1537] Agus Salim, Ma Xiangmei, Li Jialiang, and Marie Reilly. A maximum likelihood method for secondary analysis of nested case-control data. Statistics in Medicine, 33(11):1842–1852, May 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2014:CIR

[1538] Shunpu Zhang, Jun Luo, Li Zhu, David G. Stinchcomb, Dave Campbell, Ginger Carter, Scott Gilkeson, and Eric J. Feuer. Confidence intervals for ranks of age-adjusted rates across states or counties. Statistics in Medicine, 33(11):1853–1866, May 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bhatt:2014:SSS

[1539] Vijaya Bhatt and Neeraj Tiwari. A spatial scan statistic for survival data based on Weibull distribution. Statistics in Medicine, 33(11):1867–1876, May 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2014:MMG

[1540] Baoyue Li, Luk Bruyneel, and Emmanuel Lesaffre. A multivariate multilevel Gaussian model with a mixed effects structure in the mean and covariance part. Statistics in Medicine, 33(11):1877–1899, May 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baio:2014:BMC

[1541] Gianluca Baio. Bayesian models for cost-effectiveness analysis in the presence of structural zero costs. Statistics in Medicine, 33(11):1900– 1913, May 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Biswas:2014:RPU

[1542] Swati Biswas, Banu Arun, and Giovanni Parmigiani. Reclassification of predictions for uncovering subgroup specific improvement. Statistics in Medicine, 33(11):1914–1927, May 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paiva:2014:ICD

[1543] Thais Paiva, Avishek Chakraborty, Jerry Reiter, and Alan Gelfand. Imputation of confidential data sets with spatial locations using disease mapping models. Statistics in Medicine, 33(11):1928–1945, May 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goeman:2014:MHT

[1544] Jelle J. Goeman and Aldo Solari. Multiple hypothesis testing in genomics. Statistics in Medicine, 33(11):1946–1978, May 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2014:CSS

[1545] H. Zhu and H. Lakkis. Comments on ‘Sample size calculation for comparing two negative binomial rates’. Statistics in Medicine, 33(11):1979, May 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zwahlen:2014:CCI

[1546] Marcel Zwahlen. Correction for ‘causal inference, probability theory, and graphical insights’. Statistics in Medicine, 33(11):1980, May 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finkelstein:2014:JTP

[1547] Dianne M. Finkelstein and David A. Schoenfeld. A joint test for progression and survival with interval-censored data from a cancer clinical trial. Statistics in Medicine, 33(12):1981–1989, May 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wages:2014:PIA

[1548] Nolan A. Wages and Mark R. Conaway. Phase I/II adaptive design for drug combination oncology trials. Statistics in Medicine, 33(12): 1990–2003, May 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kwak:2014:PIC

[1549] Minjung Kwak and Sin-Ho Jung. Phase II clinical trials with time-to-event endpoints: optimal two-stage designs with one-sample log-rank test. Statistics in Medicine, 33(12):2004–2016, May 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

An:2014:CPD

[1550] Ming-Wen An, Constantine E. Frangakis, and Constantin T. Yiannoutsos. Choosing profile double-sampling designs for survival estimation with application to President’s emergency plan for AIDS relief evaluation. Statistics in Medicine, 33(12):2017–2029, May 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2014:CST

[1551] Xiaodong Li, Jingchen Liu, Naihua Duan, Huiping Jiang, Ragy Girgis, and Jeffrey Lieberman. Cumulative sojourn time in longitudinal studies: a sequential imputation method to handle missing health state data due to dropout. Statistics in Medicine, 33(12):2030–2047, May 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2014:MCC

[1552] Kevin He and Douglas E. Schaubel. Methods for comparing centerspecific survival outcomes using direct standardization. Statistics in Medicine, 33(12):2048–2061, May 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guolo:2014:SAM

[1553] A. Guolo. The SIMEX approach to measurement error correction in meta-analysis with baseline risk as covariate. Statistics in Medicine, 33 (12):2062–2076, May 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hosseini-Nasab:2014:FAG

[1554] Mohammad Hosseini-Nasab and Zahra Mirzaei K. Functional analysis of glaucoma data. Statistics in Medicine, 33(12):2077–2102, May 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cao:2014:STT

[1555] Guanqun Cao, Wei-Wen Hsu, and David Todem. A score-type test for heterogeneity in zero-inflated models in a stratified population. Statistics in Medicine, 33(12):2103–2114, May 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bryan:2014:DRM

[1556] Matthew Bryan and Patrick J. Heagerty. Direct regression models for longitudinal rates of change. Statistics in Medicine, 33(12):2115–2136, May 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keogh:2014:TME

[1557] Ruth H. Keogh and Ian R. White. A toolkit for measurement error correction, with a focus on nutritional epidemiology. Statistics in Medicine, 33(12):2137–2155, May 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wages:2014:CDF

[1558] Nolan A. Wages, Mark R. Conaway, and John O’Quigley. Comments on ‘A dose-finding approach based on shrunken predictive probability for combinations of two agents in phase I trials’ by Akihiro Hirakawa, Chikuma Hamada, and Shigeyuki Matsui. Statistics in Medicine, 33(12): 2156–2158, May 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hirakawa:2014:RLE

[1559] Akihiro Hirakawa and Shigeyuki Matsui. Response to letter to the Editor by Dr Wages et al. Statistics in Medicine, 33(12):2159–2160, May 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farewell:2014:MGE

[1560] Vern Farewell and Tony Johnson. Major Greenwood’s early career and the first departments of medical statistics. Statistics in Medicine, 33(13): 2161–2177, June 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andridge:2014:AMI

[1561] Rebecca R. Andridge, Abigail B. Shoben, Keith E. Muller, and David M. Murray. Analytic methods for individually randomized group treatment trials and group-randomized trials when subjects belong to multiple groups. Statistics in Medicine, 33(13):2178–2190, June 15, 2014. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quan:2014:MRC

[1562] Hui Quan, Xuezhou Mao, Joshua Chen, Weichung Joe Shih, Soo Peter Ouyang, Ji Zhang, Peng-Liang Zhao, and Bruce Binkowitz. Multiregional clinical trial design and consistency assessment of treatment effects. Statistics in Medicine, 33(13):2191–2205, June 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wason:2014:CBA

[1563] James M. S. Wason and Lorenzo Trippa. A comparison of Bayesian adaptive randomization and multi-stage designs for multi-arm clinical trials. Statistics in Medicine, 33(13):2206–2221, June 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Westgate:2014:ICS

[1564] Philip M. Westgate. Improving the correlation structure selection approach for generalized estimating equations and balanced longitudinal data. Statistics in Medicine, 33(13):2222–2237, June 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morgan:2014:HFM

[1565] Charity J. Morgan, Mark F. Lenzenweger, Donald B. Rubin, and Deborah L. Levy. A hierarchical finite mixture model that accommodates zero-inflated counts, non-independence, and heterogeneity. Statistics in Medicine, 33(13):2238–2250, June 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2014:MIH

[1566] Ren He and Thomas Belin. Multiple imputation for high-dimensional mixed incomplete continuous and binary data. Statistics in Medicine, 33(13):2251–2262, June 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2014:EIE

[1567] Ying Xu, K. F. Lam, and Yin Bun Cheung. Estimation of intervention effects using recurrent event time data in the presence of event dependence and a cured fraction. Statistics in Medicine, 33(13):2263–2274, June 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Efthimiou:2014:AMM

[1568] Orestis Efthimiou, Dimitris Mavridis, Andrea Cipriani, Stefan Leucht, Pantelis Bagos, and Georgia Salanti. An approach for modelling multiple correlated outcomes in a network of interventions using odds ratios. Statistics in Medicine, 35(16):2275–2287, June 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maruo:2014:CIB

[1569] Kazushi Maruo and Norisuke Kawai. Confidence intervals based on some weighting functions for the difference of two binomial proportions. Statistics in Medicine, 35(16):2288–2296, June 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baiocchi:2014:IVM

[1570] Michael Baiocchi, Jing Cheng, and Dylan S. Small. Instrumental variable methods for causal inference. Statistics in Medicine, 35(16):2297–2340, June 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Debray:2014:MAA

[1571] Thomas P. A. Debray, Hendrik Koffijberg, Daan Nieboer, Yvonne Vergouwe, Ewout W. Steyerberg, and Karel G. M. Moons. Meta-analysis and aggregation of multiple published prediction models. Statistics in Medicine, 33(14):2341–2362, June 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2014:OID

[1572] Xuelin Huang, Jing Ning, and Abdus S. Wahed. Optimization of individualized dynamic treatment regimes for recurrent diseases. Statistics in Medicine, 33(14):2363–2378, June 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lorent:2014:NTD

[1573] Marine Lorent, Magali Giral, and Yohann Foucher. Net time-dependent ROC curves: a solution for evaluating the accuracy of a marker to predict disease-related mortality. Statistics in Medicine, 33(14):2379–2389, June 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nattino:2014:NCT

[1574] Giovanni Nattino, Stefano Finazzi, and Guido Bertolini. A new calibration test and a reappraisal of the calibration belt for the assessment of prediction models based on dichotomous outcomes. Statistics in Medicine, 33(14):2390–2407, June 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2014:LRB

[1575] Lan Huang, Jyoti Zalkikar, and Ram Tiwari. Likelihood ratio based tests for longitudinal drug safety data. Statistics in Medicine, 33(14): 2408–2424, June 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2014:NKS

[1576] Ming Zhou and Zhao Yang. A note on the kappa statistic for clustered dichotomous data. Statistics in Medicine, 33(14):2425–2448, June 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Warren:2014:HNM

[1577] Fiona C. Warren, Keith R. Abrams, and Alex J. Sutton. Hierarchical network meta-analysis models to address sparsity of events and differing treatment classifications with regard to adverse outcomes. Statistics in Medicine, 33(14):2449–2466, June 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2014:LAP

[1578] Hong Zhu. Likelihood approaches for proportional likelihood ratio model with right-censored data. Statistics in Medicine, 33(14):2467–2479, June 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neugebauer:2014:TLR

[1579] Romain Neugebauer, Julie A. Schmittdiel, and Mark J. van der Laan. Targeted learning in real-world comparative effectiveness research with time-varying interventions. Statistics in Medicine, 33(14):2480–2520, June 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Combescure:2014:MAS

[1580] Christophe Combescure, Yohann Foucher, and Daniel Jackson. Metaanalysis of single-arm survival studies: a distribution-free approach for estimating summary survival curves with random effects. Statistics in Medicine, 33(15):2521–2537, July 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Manju:2014:SSC

[1581] Md. Abu Manju, Math J. J. M. Candel, and Martijn P. F. Berger. Sample size calculation in cost-effectiveness cluster randomized trials: optimal and maximin approaches. Statistics in Medicine, 33(15):2538–2553, July 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2014:JMD

[1582] Sheng Luo, Xiao Su, Stacia M. DeSantis, Xuelin Huang, Min Yi, and Kelly K. Hunt. Joint model for a diagnostic test without a gold standard in the presence of a dependent terminal event. Statistics in Medicine, 33(15):2554–2566, July 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sheng:2014:NSA

[1583] Elisa Sheng, Xiao Hua Zhou, Hua Chen, Guizhou Hu, and Ashlee Duncan. A new synthesis analysis method for building logistic regression prediction models. Statistics in Medicine, 33(15):2567–2576, July 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xanthakis:2014:AIP

[1584] Vanessa Xanthakis, Lisa M. Sullivan, Ramachandran S. Vasan, Emelia J. Benjamin, Joseph M. Massaro, Ralph B. D’Agostino, Sr., and Michael J. Pencina. Assessing the incremental predictive performance of novel biomarkers over standard predictors. Statistics in Medicine, 33(15):2577– 2584, July 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoorde:2014:ACM

[1585] Kirsten Van Hoorde, Yvonne Vergouwe, Dirk Timmerman, Sabine Van Huffel, Ewout W. Steyerberg, and Ben Van Calster. Assessing calibration of multinomial risk prediction models. Statistics in Medicine, 33(15):

2585–2596, July 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mandal:2014:SIT

[1586] Saumen Mandal and Atanu Biswas. Shift-invariant target in allocation problems. Statistics in Medicine, 33(15):2597–2611, July 10, 2014. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2014:KSC

[1587] Zhao Yang and Ming Zhou. Kappa statistic for clustered matched-pair data. Statistics in Medicine, 33(15):2612–2633, July 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prates:2014:RRE

[1588] Marcos O. Prates, Martin Kulldorff, and Renato M. Assunção. Relative risk estimates from spatial and space-time scan statistics: are they biased? Statistics in Medicine, 33(15):2634–2644, July 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2014:DTC

[1589] Haiqun Lin, Ling Han, Peter N. Peduzzi, Terrence E. Murphy, Thomas M. Gill, and Heather G. Allore. A dynamic trajectory class model for intensive longitudinal categorical outcome. Statistics in Medicine, 33(15):2645–2664, July 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sinclair:2014:BAD

[1590] Karen Sinclair and Anne Whitehead. A Bayesian approach to dosefinding studies for cancer therapies: incorporating later cycles of therapy. Statistics in Medicine, 33(15):2665–2680, July 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2014:CAB

[1591] Jung Ae Lee, Kevin K. Dobbin, and Jeongyoun Ahn. Covariance adjustment for batch effect in gene expression data. Statistics in Medicine, 33(15):2681–2695, July 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nattino:2014:CGA

[1592] Giovanni Nattino, Stefano Finazzi, and Guido Bertolini. Comments on ‘Graphical assessment of internal and external calibration of logistic regression models by using loess smoothers’ by Peter C. Austin and Ewout W. Steyerberg. Statistics in Medicine, 33(15):2696–2698, July 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2014:BCI

[1593] Peter C. Austin and Ewout W. Steyerberg. Bootstrap confidence intervals for loess-based calibration curves. Statistics in Medicine, 33(15): 2699–2700, July 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2014:SQH

[1594] Yan Zhou and Nandini Dendukuri. Statistics for quantifying heterogeneity in univariate and bivariate meta-analyses of binary data: The case of meta-analyses of diagnostic accuracy. Statistics in Medicine, 33(16): 2701–2717, July 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bartroff:2014:NAD

[1595] Jay Bartroff, Tze Leung Lai, and Balasubramanian Narasimhan. A new approach to designing phase I-II cancer trials for cytotoxic chemotherapies. Statistics in Medicine, 33(16):2718–2735, July 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2014:TEA

[1596] Samuel S. Wu, Yi-Hsuan Tu, and Ying He. Testing for efficacy in adaptive clinical trials with enrichment. Statistics in Medicine, 33(16):2736–2745, July 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hund:2014:ECL

[1597] Lauren Hund and Marcello Pagano. Extending cluster lot quality assurance sampling designs for surveillance programs. Statistics in Medicine, 33(16):2746–2757, July 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Piroutek:2014:STP

[1598] Aline Piroutek, Renato Assunção, and Thaís Paiva. Space-time prospective surveillance based on Knox local statistics. Statistics in Medicine, 33(16):2758–2773, July 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sfikas:2014:QUR

[1599] Nikolaos Sfikas, David Greenhalgh, Wenwen Huo, Janet Mortimer, and Chris. Robertson. Quantifying unrecognised replication present in reports of HIV diagnoses. Statistics in Medicine, 33(16):2774–2796, July 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lefebvre:2014:EBA

[1600] Geneviève Lefebvre, Joseph A. Delaney, and Robyn L. McClelland. Extending the Bayesian adjustment for confounding algorithm to binary treatment covariates to estimate the effect of smoking on carotid intimamedia thickness: the multi-ethnic study of atherosclerosis. Statistics in Medicine, 33(16):2797–2813, July 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2014:SMI

[1601] Peng Zhang, Peichao Peng, Lu Wang, and Yu Kang. G scores: a method for identifying disease-causing pathogens with application to lower respiratory tract infections. Statistics in Medicine, 33(16):2814–2829, July 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2014:BAN

[1602] Xiaosun Lu and Yangxin Huang. Bayesian analysis of nonlinear mixedeffects mixture models for longitudinal data with heterogeneity and skewness. Statistics in Medicine, 33(16):2830–2849, July 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fagerland:2014:RTC

[1603] Morten W. Fagerland, Stian Lydersen, and Petter Laake. Recommended tests and confidence intervals for paired binomial proportions. Statistics in Medicine, 33(16):2850–2875, July 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2014:NRR

[1604] Stephen Senn. A note regarding ‘random effects’. Statistics in Medicine, 33(16):2876–2877, July 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kahan:2014:NRR

[1605] Brennan C. Kahan and Tim P. Morris. A note regarding ‘random effects’ — authors’ response. Statistics in Medicine, 33(16):2878–2879, July 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2014:ARC

[1606] Haiyuan Zhu and Hassan Lakkis. Authors’ reply to comments on ‘Sample size calculation for comparing two negative binomial rates’. Statistics in Medicine, 33(16):2880, July 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sitlani:2014:ALD

[1607] Colleen M. Sitlani and Patrick J. Heagerty. Analyzing longitudinal data to characterize the accuracy of markers used to select treatment. Statistics in Medicine, 33(17):2881–2896, July 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Asakura:2014:SSD

[1608] Koko Asakura, Toshimitsu Hamasaki, Tomoyuki Sugimoto, Kenichi Hayashi, Scott R. Evans, and Takashi Sozu. Sample size determination in group-sequential clinical trials with two co-primary endpoints. Statistics in Medicine, 33(17):2897–2913, July 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Southworth:2014:PPL

[1609] Harry Southworth. Predicting potential liver toxicity from phase 2 data: a case study with ximelagatran. Statistics in Medicine, 33(17):2914–2923, July 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2014:IBL

[1610] Siyan Xu, Steven Y. Hua, Ronald Menton, Kerry Barker, Sandeep Menon, and Ralph B. D’Agostino. Inference of bioequivalence for log-normal distributed data with unspecified variances. Statistics in Medicine, 33(17):2924–2938, July 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kunz:2014:RAF

[1611] Michael Kunz. On responder analyses in the framework of within subject comparisons — considerations and two case studies. Statistics in Medicine, 33(17):2939–2952, July 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2014:SED

[1612] Yeh-Fong Chen, Xiangmin Zhang, Roy N. Tamura, and Chiung M. Chen. A sequential enriched design for target patient population in psychiatric clinical trials. Statistics in Medicine, 33(17):2953–2967, July 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klingenberg:2014:NIC

[1613] Bernhard Klingenberg. A new and improved confidence interval for the Mantel–Haenszel risk difference. Statistics in Medicine, 33(17):2968– 2983, July 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pradhan:2014:WPL

[1614] Vivek Pradhan, Krishna K. Saha, Tathagata Banerjee, and John C. Evans. Weighted profile likelihood-based confidence interval for the difference between two proportions with paired binomial data. Statistics in Medicine, 33(17):2984–2997, July 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilding:2014:EAT

[1615] Gregory E. Wilding, Joseph D. Consiglio, and Guogen Shan. Exact approaches for testing hypotheses based on the intra-class kappa coefficient. Statistics in Medicine, 33(17):2998–3012, July 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siddique:2014:BVM

[1616] Juned Siddique, Ofer Harel, Catherine M. Crespi, and Donald Hedeker. Binary variable multiple-model multiple imputation to address missing data mechanism uncertainty: application to a smoking cessation trial. Statistics in Medicine, 33(17):3013–3028, July 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2014:MNM

[1617] Wan-Lun Wang and Tsung-I Lin. Multivariate t nonlinear mixed-effects models for multi-outcome longitudinal data with missing values. Statistics in Medicine, 33(17):3029–3046, July 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Biard:2014:PTC

[1618] L. Biard, R. Porcher, and M. Resche-Rigon. Permutation tests for centre effect on survival endpoints with application in an acute myeloid leukaemia multicentre study. Statistics in Medicine, 33(17):3047–3057, July 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2014:IVM

[1619] Stuart G. Baker. Instrumental variable methods for causal inference: early work and recent developments. Statistics in Medicine, 33(17):3058– 3059, July 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2014:C

[1620] Dan Jackson. Correction. Statistics in Medicine, 33(17):3060, July 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burkner:2014:TPB

[1621] Paul-Christian Bürkner and Philipp Doebler. Testing for publication bias in diagnostic meta-analysis: a simulation study. Statistics in Medicine, 33(18):3061–3077, August 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parhat:2014:CMC

[1622] Parwen Parhat, William F. Rosenberger, and Guoqing Diao. Conditional Monte Carlo randomization tests for regression models. Statistics in Medicine, 33(18):3078–3088, August 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eng:2014:RRM

[1623] Kevin H. Eng. Randomized reverse marker strategy design for prospective biomarker validation. Statistics in Medicine, 33(18):3089–3099, August 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2014:AMS

[1624] Chunling Liu, Aiyi Liu, Jiang Hu, Vivian Yuan, and Susan Halabi. Adjusting for misclassification in a stratified biomarker clinical trial. Statistics in Medicine, 33(18):3100–3113, August 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2014:WCS

[1625] Rena Jie Sun, John D. Kalbfleisch, and Douglas E. Schaubel. A weighted cumulative sum (WCUSUM) to monitor medical outcomes with dependent censoring. Statistics in Medicine, 33(18):3114–3129, August 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kliethermes:2014:BAF

[1626] Stephanie Kliethermes and Jacob Oleson. A Bayesian approach to functional mixed-effects modeling for longitudinal data with binomial outcomes. Statistics in Medicine, 33(18):3130–3146, August 15, 2014. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Belot:2014:JFM

[1627] Aurélien Belot, Virginie Rondeau, Laurent Remontet, Roch Giorgi, and the Censur working survival group. A joint frailty model to estimate the recurrence process and the disease-specific mortality process without needing the cause of death. Statistics in Medicine, 33(18):3147–3166, August 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andrinopoulou:2014:JMT

[1628] Eleni-Rosalina Andrinopoulou, Dimitris Rizopoulos, Johanna J. M. Takkenberg, and Emmanuel Lesaffre. Joint modeling of two longitudinal outcomes and competing risk data. Statistics in Medicine, 33(18):3167– 3178, August 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gong:2014:AGF

[1629] Gail Gong, Anne S. Quante, Mary Beth Terry, and Alice S. Whittemore. Assessing the goodness of fit of personal risk models. Statistics in Medicine, 33(18):3179–3190, August 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gerds:2014:CPR

[1630] Thomas A. Gerds, Per K. Andersen, and Michael W. Kattan. Calibration plots for risk prediction models in the presence of competing risks. Statistics in Medicine, 33(18):3191–3203, August 15, 2014. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheon:2014:MTM

[1631] Kyeongmi Cheon, Marie E. Thoma, Xiangrong Kong, and Paul S. Albert. A mixture of transition models for heterogeneous longitudinal ordinal data: with applications to longitudinal bacterial vaginosis data. Statistics in Medicine, 33(18):3204–3213, August 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shinozaki:2014:ECD

[1632] Tomohiro Shinozaki, Yutaka Matsuyama, and Yasuo Ohashi. Estimation of controlled direct effects in time-varying treatments using structural nested mean models: application to a primary prevention trial for coronary events with pravastatin. Statistics in Medicine, 33(18):3214–3228, August 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:2014:JLS

[1633] Jianxin Pan, Yanchun Bao, Hongsheng Dai, and Hong-Bin Fang. Joint longitudinal and survival-cure models in tumour xenograft experiments. Statistics in Medicine, 33(18):3229–3240, August 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proschan:2014:MPD

[1634] Michael A. Proschan and Lori E. Dodd. A modest proposal for dropping poor arms in clinical trials. Statistics in Medicine, 33(19):3241–3252, August 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanSchie:2014:RES

[1635] S. van Schie and M. Moerbeek. Re-estimating sample size in cluster randomised trials with active recruitment within clusters. Statistics in Medicine, 33(19):3253–3268, August 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Magirr:2014:FSD

[1636] D. Magirr, N. Stallard, and T. Jaki. Flexible sequential designs for multiarm clinical trials. Statistics in Medicine, 33(19):3269–3279, August 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Batterton:2014:CIA

[1637] Katherine A. Batterton and Christine M. Schubert. Confidence intervals around Bayes cost in multi-state diagnostic settings to estimate optimal performance. Statistics in Medicine, 33(19):3280–3299, August 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsai:2014:AAS

[1638] Chen-An Tsai, Chih-Yang Huang, and Jen pei Liu. An approximate approach to sample size determination in bioequivalence testing with multiple pharmacokinetic responses. Statistics in Medicine, 33(19):3300– 3317, August 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wynant:2014:IMB

[1639] Willy Wynant and Michal Abrahamowicz. Impact of the model-building strategy on inference about nonlinear and time-dependent covariate effects in survival analysis. Statistics in Medicine, 33(19):3318–3337, August 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kobayashi:2014:NPM

[1640] Fumiaki Kobayashi and Manabu Kuroki. A new proportion measure of the treatment effect captured by candidate surrogate endpoints. Statistics in Medicine, 33(19):3338–3353, August 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2014:UMA

[1641] Yi Zhou, John Lefante, Janet Rice, and Shande Chen. Using modified approaches on marginal regression analysis of longitudinal data with time-dependent covariates. Statistics in Medicine, 33(19):3354–3364, August 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brannath:2014:NCP

[1642] Werner Brannath and Sylvia Schmidt. A new class of powerful and informative simultaneous confidence intervals. Statistics in Medicine, 33 (19):3365–3386, August 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2014:CSB

[1643] Rongxia Li, Brock Stewart, Eric Weintraub, and Michael M. McNeil. Continuous sequential boundaries for vaccine safety surveillance. Statistics in Medicine, 33(19):3387–3397, August 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shults:2014:NID

[1644] Justine Shults and Matthew W. Guerra. A note on implementation of decaying product correlation structures for quasi-least squares. Statistics in Medicine, 33(19):3398–3404, August 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilden:2014:NEN

[1645] Jørgen Hilden and Thomas A. Gerds. A note on the evaluation of novel biomarkers: do not rely on integrated discrimination improvement and net reclassification index. Statistics in Medicine, 33(19):3405–3414, August 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leening:2014:NRI

[1646] Maarten J. G. Leening, Ewout W. Steyerberg, Ben Van Calster, Ralph B. D’Agostino, Sr., and Michael J. Pencina. Net reclassification improvement and integrated discrimination improvement require calibrated models: relevance from a marker and model perspective. Statistics in Medicine, 33(19):3415–3418, August 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gerds:2014:CMS

[1647] Thomas A. Gerds and Jørgen Hilden. Calibration of models is not sufficient to justify NRI. Statistics in Medicine, 33(19):3419–3420, August 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Small:2014:TTH

[1648] Dylan S. Small, Marshall M. Joffe, Kevin G. Lynch, Jason A. Roy, and A. Russell Localio. Tom ten Have’s contributions to causal inference and biostatistics: review and future research directions. Statistics in Medicine, 33(20):3421–3433, September 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leiby:2014:BMG

[1649] Benjamin E. Leiby, Thomas R. Ten Have, Kevin G. Lynch, and Mary D. Sammel. Bayesian multivariate growth curve latent class models for mixed outcomes. Statistics in Medicine, 33(20):3434–3452, September 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:2014:JMC

[1650] Xin Gao, Gregory K. Brown, and Michael R. Elliott. Joint modeling compliance and outcome for causal analysis in longitudinal studies. Statistics in Medicine, 33(20):3453–3465, September 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Almirall:2014:TVE

[1651] Daniel Almirall, Beth Ann Griffin, Daniel F. McCaffrey, Rajeev Ramchand, Robert A. Yuen, and Susan A. Murphy. Time-varying effect moderation using the structural nested mean model: estimation using inverse-weighted regression with residuals. Statistics in Medicine, 33 (20):3466–3487, September 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leacy:2014:JUP

[1652] Finbarr P. Leacy and Elizabeth A. Stuart. On the joint use of propensity and prognostic scores in estimation of the average treatment effect on the treated: a simulation study. Statistics in Medicine, 33(20):3488– 3508, September 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2014:CIL

[1653] Chen-Pin Wang, Booil Jo, and C. Hendricks Brown. Causal inference in longitudinal comparative effectiveness studies with repeated measures of a continuous intermediate variable. Statistics in Medicine, 33(20):3509– 3527, September 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2014:UIV

[1654] Zijian Guo, Jing Cheng, Scott A. Lorch, and Dylan S. Small. Using an instrumental variable to test for unmeasured confounding. Statistics in Medicine, 33(20):3528–3546, September 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2014:EEI

[1655] Wei Yang, Kathleen J. Propert, and J. Richard Landis. Estimating the efficacy of an interstitial cystitis/painful bladder syndrome medication in a randomized trial with both non-adherence and loss to follow-up. Statistics in Medicine, 33(20):3547–3555, September 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leyrat:2014:PSM

[1656] Clémence Leyrat, Agnès Caille, Allan Donner, and Bruno Giraudeau. Propensity score methods for estimating relative risks in cluster randomized trials with low-incidence binary outcomes and selection bias. Statistics in Medicine, 33(20):3556–3575, September 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2014:PAG

[1657] Hao Xu, Bertrand Thirion, and Stéphanie Allassonnière. Probabilistic atlas and geometric variability estimation to drive tissue segmentation. Statistics in Medicine, 33(20):3576–3599, September 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Small:2014:C

[1658] Dylan Small. Correction. Statistics in Medicine, 33(20):3600, September 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tchetgen:2014:IES

[1659] Eric J. Tchetgen Tchetgen. Identification and estimation of survivor average causal effects. Statistics in Medicine, 33(21):3601–3628, September 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2014:STP

[1660] Yueqiong Lin, Koon Shing Kwong, Siu Hung Cheung, and Wai-Yin Poon. Step-up testing procedure for multiple comparisons with a control for a latent variable model with ordered categorical responses. Statistics in Medicine, 33(21):3629–3638, September 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2014:DTI

[1661] Dan Jackson, Jessica K. Barrett, Stephen Rice, Ian R. White, and Julian P. T. Higgins. A design-by-treatment interaction model for network meta-analysis with random inconsistency effects. Statistics in Medicine, 33(21):3639–3654, September 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindhagen:2014:LAF

[1662] Lars Lindhagen, Bahman Darkahi, Gabriel Sandblom, and Lars Berglund. Level-adjusted funnel plots based on predicted marginal expectations: an application to prophylactic antibiotics in gallstone surgery. Statistics in Medicine, 33(21):3655–3675, September 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moreno:2014:OBM

[1663] E. Moreno, F. J. Vázquez-Polo, and M. A. Negrín. Objective Bayesian meta-analysis for sparse discrete data. Statistics in Medicine, 33(21):

3676–3692, September 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2014:SDR

[1664] Xiaobing Zhao and Xian Zhou. Sufficient dimension reduction on the mean and rate functions of recurrent events. Statistics in Medicine, 33 (21):3693–3709, September 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2014:CIT

[1665] Yulei He, Mary Beth Landrum, and Alan M. Zaslavsky. Combining information from two data sources with misreporting and incompleteness to assess hospice-use among cancer patients: a multiple imputation approach. Statistics in Medicine, 33(21):3710–3724, September 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Welch:2014:ETF

[1666] Catherine A. Welch, Irene Petersen, Jonathan W. Bartlett, Ian R. White, Louise Marston, Richard W. Morris, Irwin Nazareth, Kate Walters, and James Carpenter. Evaluation of two-fold fully conditional specification multiple imputation for longitudinal electronic health record data. Statistics in Medicine, 33(21):3725–3737, September 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mollica:2014:EPM

[1667] Cristina Mollica and Luca Tardella. Epitope profiling via mixture modeling of ranked data. Statistics in Medicine, 33(21):3738–3758, September 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Galvis:2014:AMB

[1668] Diana M. Galvis, Dipankar Bandyopadhyay, and Victor H. Lachos. Augmented mixed beta regression models for periodontal proportion data. Statistics in Medicine, 33(21):3759–3771, September 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:2014:CGG

[1669] Christopher Cox and Matthew Matheson. A comparison of the generalized gamma and exponentiated Weibull distributions. Statistics in Medicine, 33(21):3772–3780, September 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kapur:2014:SSD

[1670] Kush Kapur, Runa Bhaumik, X. Charlene Tang, Kwan Hur, Domenic J. Reda, and Dulal K. Bhaumik. Sample size determination for longitudinal designs with binary response. Statistics in Medicine, 33(22):3781–3800, September 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2014:IBS

[1671] Jing Zhou, Adeniyi Adewale, Yue Shentu, Jiajun Liu, and Keaven Anderson. Information-based sample size re-estimation in group sequential design for longitudinal trials. Statistics in Medicine, 33(22):3801–3814, September 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tighiouart:2014:DFD

[1672] Mourad Tighiouart, Steven Piantadosi, and André Rogatko. Dose finding with drug combinations in cancer phase I clinical trials using conditional escalation with overdose control. Statistics in Medicine, 33(22):3815– 3829, September 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2014:OST

[1673] John Whitehead. One-stage and two-stage designs for phase II clinical trials with survival endpoints. Statistics in Medicine, 33(22):3830–3843, September 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crowther:2014:MME

[1674] Michael J. Crowther, Maxime P. Look, and Richard D. Riley. Multilevel mixed effects parametric survival models using adaptive Gauss-Hermite quadrature with application to recurrent events and individual participant data meta-analysis. Statistics in Medicine, 33(22):3844–3858, September 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OGorman:2014:RCC

[1675] Thomas W. O’Gorman. Regaining confidence in confidence intervals for the mean treatment effect. Statistics in Medicine, 33(22):3859–3868, September 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paul:2014:SSG

[1676] Sudhir Paul and Xuemao Zhang. Small sample GEE estimation of regression parameters for longitudinal data. Statistics in Medicine, 33(22): 3869–3881, September 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hendry:2014:ASC

[1677] G. Hendry, D. North, T. Zewotir, and R. N. Naidoo. The application of subset correspondence analysis to address the problem of missing data in a study on asthma severity in childhood. Statistics in Medicine, 33(22): 3882–3893, September 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boren:2014:SVN

[1678] David Boren, Patrick S. Sullivan, Chris Beyrer, Stefan D. Baral, Linda-Gail Bekker, and Ron Brookmeyer. Stochastic variation in network epidemic models: implications for the design of community level HIV prevention trials. Statistics in Medicine, 33(22):3894–3904, September 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2014:CIC

[1679] Pan Wu, Douglas Gunzler, Naiji Lu, Tian Chen, Peter Wymen, and Xin M. Tu. Causal inference for community-based multi-layered intervention study. Statistics in Medicine, 33(22):3905–3918, September 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hess:2014:CRD

[1680] Wolfgang Hess, Larissa Schwarzkopf, Matthias Hunger, and Rolf Holle. Competing-risks duration models with correlated random effects: an application to dementia patients’ transition histories. Statistics in Medicine, 33(22):3919–3931, September 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ha:2014:MRS

[1681] Neung Soo Ha, Partha Lahiri, and Van Parsons. Methods and results for small area estimation using smoking data from the 2008 national health interview survey. Statistics in Medicine, 33(22):3932–3945, September 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2014:HIS

[1682] Stuart G. Baker, Ewoud Schuit, Ewout W. Steyerberg, Michael J. Pencina, Andew Vickers, Karel G. M. Moons, Ben W. J. Mol, and Karen S. Lindeman. How to interpret a small increase in AUC with an additional risk prediction marker: decision analysis comes through. Statistics in Medicine, 33(22):3946–3959, September 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2014:C

[1683] Stuart G. Baker. Correction. Statistics in Medicine, 34(5):3960, September 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McLain:2014:SMG

[1684] Alexander C. McLain, Rajeshwari Sundaram, Marie Thoma, and Germaine M. Buck Louis. Semiparametric modeling of grouped current duration data with preferential reporting. Statistics in Medicine, 33(23):3961– 3972, October 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2014:SSP

[1685] Wenrong Hu, Jianwen Cai, and Donglin Zeng. Sample size/power calculation for stratified case-cohort design. Statistics in Medicine, 33(23): 3973–3985, October 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dai:2014:TCI

[1686] James Y. Dai, Kwun Chuen Gary Chan, and Li Hsu. Testing concordance of instrumental variable effects in generalized linear models with application to Mendelian randomization. Statistics in Medicine, 33(23):3986– 4007, October 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morita:2014:BBB

[1687] Satoshi Morita, Hideharu Yamamoto, and Yasuo Sugitani. Biomarkerbased Bayesian randomized phase II clinical trial design to identify a sensitive patient subpopulation. Statistics in Medicine, 33(23):4008–4016, October 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2014:BDP

[1688] Chunyan Cai, Suyu Liu, and Ying Yuan. A Bayesian design for phase II clinical trials with delayed responses based on multiple imputation. Statistics in Medicine, 33(23):4017–4028, October 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2014:OAR

[1689] Mi-Ok Kim, Chunyan Liu, Feifang Hu, and J. Jack Lee. Outcomeadaptive randomization for a delayed outcome with a short-term predictor: imputation-based designs. Statistics in Medicine, 33(23):4029– 4042, October 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2014:MCD

[1690] Wenle Zhao and Valerie Durkalski. Managing competing demands in the implementation of response-adaptive randomization in a large multicenter phase III acute stroke trial. Statistics in Medicine, 33(23):4043–4052, October 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vansteelandt:2014:RAP

[1691] S. Vansteelandt and R. M. Daniel. On regression adjustment for the propensity score. Statistics in Medicine, 33(23):4053–4072, October 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carlin:2014:AAS

[1692] Caroline S. Carlin and Craig A. Solid. An approach to addressing selection bias in survival analysis. Statistics in Medicine, 33(23):4073–4086, October 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2014:CTD

[1693] Hyune-Ju Kim, Jun Luo, Jeankyung Kim, Huann-Sheng Chen, and Eric J. Feuer. Clustering of trend data using joinpoint regression models. Statistics in Medicine, 33(23):4087–4103, October 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Young:2014:CCP

[1694] Derek S. Young and Terence M. Mills. Choosing a coverage probability for forecasting the incidence of cancer. Statistics in Medicine, 33 (23):4104–4115, October 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lagona:2014:LTV

[1695] Francesco Lagona, Dmitri Jdanov, and Maria Shkolnikova. Latent timevarying factors in longitudinal analysis: a linear mixed hidden Markov model for heart rates. Statistics in Medicine, 33(23):4116–4134, October 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2014:NRL

[1696] Stephen Senn and Edith Nijedeka Jude-Eze. A note regarding Lee’s checks for minimum numbers of subjects where relative risks have been calculated. Statistics in Medicine, 33(23):4135–4138, October 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2014:BER

[1697] Stuart G. Baker. Biomarker evaluation in randomized trials: addressing different research questions. Statistics in Medicine, 33(23):4139–4140, October 15, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collins:2014:EDT

[1698] John Collins and Minh Huynh. Estimation of diagnostic test accuracy without full verification: a review of latent class methods. Statistics in Medicine, 33(24):4141–4169, October 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liublinska:2014:SAP

[1699] Victoria Liublinska and Donald B. Rubin. Sensitivity analysis for a partially missing binary outcome in a two-arm randomized clinical trial. Statistics in Medicine, 33(24):4170–4185, October 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hampson:2014:BMD

[1700] Lisa V. Hampson, John Whitehead, Despina Eleftheriou, and Paul Brogan. Bayesian methods for the design and interpretation of clinical trials in very rare diseases. Statistics in Medicine, 33(24):4186–4201, October 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shortreed:2014:MIS

[1701] Susan M. Shortreed, Eric Laber, T. Scott Stroup, Joelle Pineau, and Susan A. Murphy. A multiple imputation strategy for sequential multiple assignment randomized trials. Statistics in Medicine, 33(24):4202–4214, October 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2014:ASO

[1702] Ryung S. Kim and Robert C. Kaplan. Analysis of secondary outcomes in nested case-control study designs. Statistics in Medicine, 36(23):4215– 4226, October 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ning:2014:AHB

[1703] Shaoyang Ning, Hongquan Xu, Ibrahim Al-Shyoukh, Jiaying Feng, and Ren Sun. An application of a Hill-based response surface model for a drug combination experiment on lung cancer. Statistics in Medicine, 36 (23):4227–4236, October 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kose:2014:ELP

[1704] T. Köse, M. Orman, F. Ikiz, M. F. Baksh, J. Gallagher, and D. Böhning. Extending the Lincoln–Petersen estimator for multiple identifications in one source. Statistics in Medicine, 36(23):4237–4249, October 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2669].

Daggy:2014:ELC

[1705] Joanne Daggy, Huiping Xu, Siu Hui, and Shaun Grannis. Evaluating latent class models with conditional dependence in record linkage. Statistics in Medicine, 36(9):4250–4265, October 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jankowski:2014:RSA

[1706] Hanna Jankowski, Xiang Ji, and Larissa Stanberry. A random set approach to confidence regions with applications to the effective dose with combinations of agents. Statistics in Medicine, 36(9):4266–4278, October 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2014:BHM

[1707] Sheng Luo and Jue Wang. Bayesian hierarchical model for multiple repeated measures and survival data: an application to Parkinson’s disease. Statistics in Medicine, 36(9):4279–4291, October 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laska:2014:EDM

[1708] Eugene Laska, Morris Meisner, and Joseph Wanderling. Exact distribution of a maximally selected Wilcoxon and a new hybrid test of symmetry. Statistics in Medicine, 36(9):4292–4305, October 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2014:UBW

[1709] Peter C. Austin and Dylan S. Small. The use of bootstrapping when using propensity-score matching without replacement: a simulation study. Statistics in Medicine, 36(9):4306–4319, October 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2014:LCM

[1710] Stuart G. Baker. A latent class method for diagnostic tests: the new, reference, gold standard problem. Statistics in Medicine, 36(9):4320, October 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huque:2014:HNT

[1711] Mohammad F. Huque, Thamban Valappil, and Guoxing (Greg) Soon. Hierarchical nested trial design (HNTD) for demonstrating treatment efficacy of new antibacterial drugs in patient populations with emerging bacterial resistance. Statistics in Medicine, 33(25):4321–4336, November 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Snavely:2014:LVT

[1712] Anna C. Snavely, David P. Harrington, and Yi Li. A latent variable transformation model approach for exploring dysphagia. Statistics in Medicine, 33(25):4337–4352, November 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rucker:2014:RDR

[1713] Gerta Rücker and Guido Schwarzer. Reduce dimension or reduce weights? Comparing two approaches to multi-arm studies in network meta-analysis. Statistics in Medicine, 33(25):4353–4369, November 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pei:2014:THP

[1714] Yanbo Pei, Guo-Liang Tian, and Man-Lai Tang. Testing homogeneity of proportion ratios for stratified correlated bilateral data in two-arm randomized clinical trials. Statistics in Medicine, 33(25):4370–4386, November 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senturk:2014:WHR

[1715] Damla Sentürk, Lorien S. Dalrymple, Yi Mu, and Danh V. Nguyen. Weighted hurdle regression method for joint modeling of cardiovascular events likelihood and rate in the US dialysis population. Statistics in Medicine, 33(25):4387–4401, November 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kassahun:2014:MMH

[1716] Wondwosen Kassahun, Thomas Neyens, Geert Molenberghs, Christel Faes, and Geert Verbeke. Marginalized multilevel hurdle and zeroinflated models for overdispersed and correlated count data with excess zeros. Statistics in Medicine, 33(25):4402–4419, November 10, 2014. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2014:ECE

[1717] Lu Tian, Ramón A. Durazo-Arvizu, Gary Myers, Steve Brooks, Kurtis Sarafin, and Christopher T. Sempos. The estimation of calibration equations for variables with heteroscedastic measurement errors. Statistics in Medicine, 33(25):4420–4436, November 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shardell:2014:SAM

[1718] Michelle Shardell and Gregory E. Hicks. Statistical analysis with missing exposure data measured by proxy respondents: a misclassification problem within a missing-data problem. Statistics in Medicine, 33(25): 4437–4452, November 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ventura:2014:ECI

[1719] Leonardo Ventura and Maura Mezzetti. Estimating cancer incidence using a Bayesian back-calculation approach. Statistics in Medicine, 33 (25):4453–4468, November 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tekwe:2014:MIM

[1720] Carmen D. Tekwe, Randy L. Carter, Harry M. Cullings, and Raymond J. Carroll. Multiple indicators, multiple causes measurement error models. Statistics in Medicine, 33(25):4469–4481, November 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2014:PES

[1721] Zhiwei Zhang, Chunling Liu, Sungduk Kim, and Aiyi Liu. Prevalence estimation subject to misclassification: the mis-substitution bias and some remedies. Statistics in Medicine, 33(25):4482–4500, November 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2014:CFA

[1722] Zhiwei Jiang, Ling Wang, Chanjuan Li, Jielai Xia, and William Wang. CP function: an alpha spending function based on conditional power. Statistics in Medicine, 33(26):4501–4514, November 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehta:2014:BDP

[1723] Cyrus Mehta, Helmut Schäfer, Hanna Daniel, and Sebastian Irle. Biomarker driven population enrichment for adaptive oncology trials with time to event endpoints. Statistics in Medicine, 33(26):4515–4531, November 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:2014:MSD

[1724] Ka Young Park and Peihua Qiu. Model selection and diagnostics for joint modeling of survival and longitudinal data with crossing hazard rate functions. Statistics in Medicine, 35(28):4532–4546, November 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sauzet:2014:EDO

[1725] O. Sauzet and J. L. Peacock. Estimating dichotomised outcomes in two groups with unequal variances: a distributional approach. Statistics in Medicine, 35(28):4547–4559, November 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2014:JML

[1726] Qingxia Chen, Ryan C. May, Joseph G. Ibrahim, Haitao Chu, and Stephen R. Cole. Joint modeling of longitudinal and survival data with missing and left-censored time-varying covariates. Statistics in Medicine, 35(28):4560–4576, November 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2014:CAP

[1727] Wei Yang, Marshall M. Joffe, Sean Hennessy, and Harold I. Feldman. Covariance adjustment on propensity parameters for continuous treatment in linear models. Statistics in Medicine, 35(28):4577–4589, November 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ha:2014:VSS

[1728] Il Do Ha, Minjung Lee, Seungyoung Oh, Jong-Hyeon Jeong, Richard Sylvester, and Youngjo Lee. Variable selection in subdistribution hazard frailty models with competing risks data. Statistics in Medicine, 35(28): 4590–4604, November 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2014:MIM

[1729] Minjung Lee, James J. Dignam, and Junhee Han. Multiple imputation methods for nonparametric inference on cumulative incidence with missing cause of failure. Statistics in Medicine, 35(28):4605–4626, November 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jurgens:2014:BGA

[1730] Verena Jürgens, Silvia Ess, Thomas Cerny, and Penelope Vounatsou. A Bayesian generalized age-period-cohort power model for cancer projections. Statistics in Medicine, 35(28):4627–4636, November 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cabras:2014:COB

[1731] Stefano Cabras, Maria Eugenia Castellanos, and Silvia Perra. Comparison of objective Bayes factors for variable selection in parametric regression models for survival analysis. Statistics in Medicine, 35(28): 4637–4654, November 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiang:2014:ATU

[1732] Fang Xiang, Susan Murray, and Xiaohong Liu. Analysis of transplant urgency and benefit via multiple imputation. Statistics in Medicine, 35 (28):4655–4670, November 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanGeloven:2014:CDC

[1733] N. van Geloven, R. B. Geskus, B. W. Mol, and A. H. Zwinderman. Correcting for the dependent competing risk of treatment using inverse probability of censoring weighting and copulas in the estimation of natural conception chances. Statistics in Medicine, 35(28):4671–4680, November 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2014:RIC

[1734] Dan Jackson, Ian R. White, Shaun Seaman, Hannah Evans, Kathy Baisley, and James Carpenter. Relaxing the independent censoring assumption in the Cox proportional hazards model using multiple imputation. Statistics in Medicine, 33(27):4681–4694, November 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2014:ITC

[1735] Patrick Royston and Willi Sauerbrei. Interaction of treatment with a continuous variable: simulation study of power for several methods of analysis. Statistics in Medicine, 33(27):4695–4708, November 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sankoh:2014:UCE

[1736] Abdul J. Sankoh, Haihong Li, and Ralph B. D’Agostino, Sr. Use of composite endpoints in clinical trials. Statistics in Medicine, 33(27): 4709–4714, November 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2014:AMF

[1737] Danjie Zhang, Ming-Hui Chen, Joseph G. Ibrahim, Mark E. Boye, Ping Wang, and Wei Shen. Assessing model fit in joint models of longitudinal and survival data with applications to cancer clinical trials. Statistics in Medicine, 33(27):4715–4733, November 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proschan:2014:CBP

[1738] Michael Proschan, Ekkehard Glimm, and Martin Posch. Connections between permutation and t-tests: relevance to adaptive methods. Statistics in Medicine, 33(27):4734–4742, November 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rikhtehgaran:2014:ETI

[1739] Reyhaneh Rikhtehgaran, Iraj Kazemi, and Geert Verbeke. Effects of time-invariant covariates on the estimation of longitudinal trends for transition mixed models. Statistics in Medicine, 33(27):4743–4755, November 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lalonde:2014:GLR

[1740] Trent L. Lalonde, Jeffrey R. Wilson, and Jianqiong Yin. GMM logistic regression models for longitudinal data with time-dependent covariates and extended classifications. Statistics in Medicine, 33(27):4756–4769, November 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:2014:RML

[1741] Kay See Tan, Benjamin French, and Andrea B. Troxel. Regression modeling of longitudinal data with outcome-dependent observation times: extensions and comparative evaluation. Statistics in Medicine, 33(27): 4770–4789, November 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wallace:2014:TDT

[1742] M. L. Wallace. Time-dependent tree-structured survival analysis with unbiased variable selection through permutation tests. Statistics in Medicine, 33(27):4790–4804, November 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chambers:2014:DMN

[1743] Ray Chambers, Emanuela Dreassi, and Nicola Salvati. Disease mapping via negative binomial regression M-quantiles. Statistics in Medicine, 33 (27):4805–4824, November 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senturk:2014:FLM

[1744] Damla Sentürk, Lorien S. Dalrymple, and Danh V. Nguyen. Functional linear models for zero-inflated count data with application to modeling hospitalizations in patients on dialysis. Statistics in Medicine, 33(27): 4825–4840, November 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaciroti:2014:BSA

[1745] Niko A. Kaciroti and Trivellore Raghunathan. Bayesian sensitivity analysis of incomplete data: bridging pattern-mixture and selection models. Statistics in Medicine, 33(27):4841–4857, November 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

MacNab:2014:C

[1746] Ying C. MacNab. Correction. Statistics in Medicine, 33(27):4858, November 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baiocchi:2014:CIV

[1747] Michael Baiocchi, Jing Cheng, and Dylan Small. Correction to “Instrumental variable methods for causal inference”. Statistics in Medicine, 26(24):4859–4860, November 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brockhaus:2014:POR

[1748] A. Catharina Brockhaus, Ralf Bender, and Guido Skipka. The Peto odds ratio viewed as a new effect measure. Statistics in Medicine, 33(28):4861– 4874, December 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Valeri:2014:MAW

[1749] Linda Valeri, Xihong Lin, and Tyler J. VanderWeele. Mediation analysis when a continuous mediator is measured with error and the outcome follows a generalized linear model. Statistics in Medicine, 33(28):4875– 4890, December 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:2014:MTP

[1750] Valerie A. Smith, John S. Preisser, Brian Neelon, and Matthew L. Maciejewski. A marginalized two-part model for semicontinuous data. Statistics in Medicine, 33(28):4891–4903, December 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perrakis:2014:CSP

[1751] Konstantinos Perrakis, Alexandros Gryparis, Joel Schwartz, Alain Le Tertre, Klea Katsouyanni, Francesco Forastiere, Massimo Stafoggia, and Evangelia Samoli. Controlling for seasonal patterns and time varying confounders in time-series epidemiological models: a simulation study.

Statistics in Medicine, 33(28):4904–4918, December 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hackstadt:2014:IEI

[1752] Amber J. Hackstadt, Elizabeth C. Matsui, D’Ann L. Williams, Gregory B. Diette, Patrick N. Breysse, Arlene M. Butz, and Roger D. Peng. Inference for environmental intervention studies using principal stratification. Statistics in Medicine, 33(28):4919–4933, December 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ueki:2014:CDF

[1753] Masao Ueki. On the choice of degrees of freedom for testing gene-gene interactions. Statistics in Medicine, 33(28):4934–4948, December 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fujita:2014:NPS

[1754] André Fujita, Daniel Y. Takahashi, Alexandre G. Patriota, and João R. Sato. A non-parametric statistical test to compare clusters with applications in functional magnetic resonance imaging data. Statistics in Medicine, 33(28):4949–4962, December 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2014:MFE

[1755] Tu Xu, Junhui Wang, and Yixin Fang. A model-free estimation for the covariate-adjusted Youden index and its associated cut-point. Statistics in Medicine, 33(28):4963–4974, December 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2014:WEN

[1756] Karol M. Pencina, Michael J. Pencina, and Ralph B. D’Agostino, Sr. What to expect from net reclassification improvement with three categories. Statistics in Medicine, 33(28):4975–4987, December 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2014:IAG

[1757] Cen Wu, Yuehua Cui, and Shuangge Ma. Integrative analysis of geneenvironment interactions under a multi-response partially linear varying coefficient model. Statistics in Medicine, 33(28):4988–4998, December 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2014:EBG

[1758] Xiang Li, Anthony Y. C. Kuk, and Jinfeng Xu. Empirical Bayes Gaussian likelihood estimation of exposure distributions from pooled samples in human biomonitoring. Statistics in Medicine, 33(28):4999–5014, December 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guillou:2014:EVT

[1759] A. Guillou, M. Kratz, and Y. Le Strat. An extreme value theory approach for the early detection of time clusters. a simulation-based assessment and an illustration to the surveillance of salmonella. Statistics in Medicine, 33(28):5015–5027, December 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mitchell:2014:HED

[1760] Emily M. Mitchell, Robert H. Lyles, Amita K. Manatunga, Neil J. Perkins, and Enrique F. Schisterman. A highly efficient design strategy for regression with outcome pooling. Statistics in Medicine, 33(28): 5028–5040, December 10, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shoben:2014:VII

[1761] Abigail B. Shoben and Scott S. Emerson. Violations of the independent increment assumption when using generalized estimating equation in longitudinal group sequential trials. Statistics in Medicine, 33(29): 5041–5056, December 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brinkley:2014:DRE

[1762] Jason Brinkley. A doubly robust estimator for the attributable benefit of a treatment regime. Statistics in Medicine, 33(29):5057–5073, December 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kleinman:2014:BSM

[1763] Alan Kleinman. The basic science and mathematics of random mutation and natural selection. Statistics in Medicine, 33(29):5074–5080, December 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gu:2014:BRC

[1764] Jiezhun Gu, Subhashis Ghosal, and David E. Kleiner. Bayesian ROC curve estimation under verification bias. Statistics in Medicine, 33(29): 5081–5096, December 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Polley:2014:TSA

[1765] Mei-Yin C. Polley, Eric C. Polley, Erich P. Huang, Boris Freidlin, and Richard Simon. Two-stage adaptive cutoff design for building and validating a prognostic biomarker signature. Statistics in Medicine, 33(29): 5097–5110, December 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2014:BAF

[1766] Baoguang Han, Menggang Yu, James J. Dignam, and Paul J. Rathouz. Bayesian approach for flexible modeling of semicompeting risks data. Statistics in Medicine, 33(29):5111–5125, December 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dziak:2014:TVE

[1767] John J. Dziak, Runze Li, Marc A. Zimmerman, and Anne Buu. Timevarying effect models for ordinal responses with applications in substance abuse research. Statistics in Medicine, 33(29):5126–5137, December 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Congdon:2014:MCS

[1768] Peter Congdon. Modelling changes in small area disability free life expectancy: trends in London wards between 2001 and 2011. Statistics in Medicine, 33(29):5138–5150, December 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Long:2014:MZI

[1769] D. Leann Long, John S. Preisser, Amy H. Herring, and Carol E. Golin. A marginalized zero-inflated Poisson regression model with overall exposure effects. Statistics in Medicine, 33(29):5151–5165, December 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elmi:2014:EBC

[1770] Angelo Elmi, Sarah J. Ratcliffe, and Wensheng Guo. The estimation of branching curves in the presence of subject-specific random effects.

Statistics in Medicine, 33(29):5166–5176, December 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ko:2014:TDA

[1771] Yi-An Ko, Bhramar Mukherjee, Jennifer A. Smith, Sung Kyun Park, Sharon L. R. Kardia, Matthew A. Allison, Pantel S. Vokonas, Jinbo Chen, and Ana V. Diez-Roux. Testing departure from additivity in Tukey’s model using shrinkage: application to a longitudinal setting. Statistics in Medicine, 33(29):5177–5191, December 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2014:ERZ

[1772] Zhu Wang, Shuangge Ma, Ching-Yun Wang, Michael Zappitelli, Prasad Devarajan, and Chirag Parikh. EM for regularized zero-inflated regression models with applications to postoperative morbidity after cardiac surgery in children. Statistics in Medicine, 33(29):5192–5208, December 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Revzin:2014:CQR

[1773] Ella Revzin, Dibyen Majumdar, and Gilbert W. Bassett, Jr. Conditional quantile regression models of melanoma tumor growth curves for assessing treatment effect in small sample studies. Statistics in Medicine, 33 (29):5209–5220, December 20, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friede:2014:GEI

[1774] Tim Friede, Robin Henderson, and Philip Hougaard. Guest Editors’ introduction. Statistics in Medicine, 33(30):5221–5222, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanHouwelingen:2014:MBV

[1775] Hans C. van Houwelingen. From model building to validation and back: a plea for robustness. Statistics in Medicine, 33(30):5223–5238, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2014:BAS

[1776] Wenle Zhao. A better alternative to stratified permuted block design for subject randomization in clinical trials. Statistics in Medicine, 33(30): 5239–5248, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lange:2014:ODC

[1777] Markus R. Lange and Heinz Schmidli. Optimal design of clinical trials with biologics using dose-time-response models. Statistics in Medicine, 33(30):5249–5264, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Desmet:2014:LME

[1778] L. Desmet, D. Venet, E. Doffagne, C. Timmermans, T. Burzykowski, C. Legrand, and M. Buyse. Linear mixed-effects models for central statistical monitoring of multicenter clinical trials. Statistics in Medicine, 33(30):5265–5279, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crowther:2014:GFP

[1779] Michael J. Crowther and Paul C. Lambert. A general framework for parametric survival analysis. Statistics in Medicine, 33(30):5280–5297, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lapidus:2014:AAA

[1780] Nathanael Lapidus, Sylvie Chevret, and Matthieu Resche-Rigon. Assessing assay agreement estimation for multiple left-censored data: a multiple imputation approach. Statistics in Medicine, 33(30):5298–5309, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bin:2014:IPA

[1781] Riccardo De Bin, Willi Sauerbrei, and Anne-Laure Boulesteix. Investigating the prediction ability of survival models based on both clinical and omics data: two case studies. Statistics in Medicine, 33(30):5310– 5329, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schipper:2014:PDS

[1782] Matthew J. Schipper, Jeremy M. G. Taylor, Randy TenHaken, Martha M. Matuzak, Feng-Ming Kong, and Theodore S. Lawrence. Personalized dose selection in radiation therapy using statistical models for toxicity and efficacy with dose and biomarkers as covariates. Statistics in Medicine, 33(30):5330–5339, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maurer:2014:NTF

[1783] Willi Maurer and Frank Bretz. A note on testing families of hypotheses using graphical procedures. Statistics in Medicine, 33(30):5340–5346, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saadati:2014:SCH

[1784] Maral Saadati and Axel Benner. Statistical challenges of highdimensional methylation data. Statistics in Medicine, 38(13):5347–5357, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmidtmann:2014:CVS

[1785] I. Schmidtmann, A. Elsäßer, A. Weinmann, and H. Binder. Coupled variable selection for regression modeling of complex treatment patterns in a clinical cancer registry. Statistics in Medicine, 38(13):5358–5370, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seaman:2014:RMH

[1786] Shaun Seaman, Menelaos Pavlou, and Andrew Copas. Review of methods for handling confounding by cluster and informative cluster size in clustered data. Statistics in Medicine, 38(13):5371–5387, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salim:2014:AIP

[1787] Agus Salim, Xiangmei Ma, Katja Fall, Ove Andrén, and Marie Reilly. Analysis of incidence and prognosis from ‘extreme’ case-control designs. Statistics in Medicine, 38(13):5388–5398, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mavridis:2014:SMA

[1788] Dimitris Mavridis, Nicky J. Welton, Alex Sutton, and Georgia Salanti. A selection model for accounting for publication bias in a full network metaanalysis. Statistics in Medicine, 38(13):5399–5412, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sauerbrei:2014:SAT

[1789] Willi Sauerbrei, Michal Abrahamowicz, Douglas G. Altman, Saskia le Cessie, James Carpenter, and on behalf of the Stratos initiative. STRengthening analytical thinking for observational studies: the STRATOS initiative. Statistics in Medicine, 38(13):5413–5432, December 30, 2014. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riviere:2015:CDD

[1790] M.-K. Riviere, F. Dubois, and S. Zohar. Competing designs for drug combination in phase I dose-finding clinical trials. Statistics in Medicine, 34 (1):1–12, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yin:2015:CCD

[1791] Guosheng Yin and Ruitao Lin. Comments on ‘Competing designs for drug combination in phase I dose-finding clinical trials’ by M-K. Riviere, F. Dubois, and S. Zohar. Statistics in Medicine, 34(1):13–17, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wages:2015:CCD

[1792] Nolan A. Wages. Comments on ‘Competing designs for drug combination in phase I dose-finding clinical trials’ by M-K. Riviere, F. Dubois, S. Zohar. Statistics in Medicine, 34(1):18–22, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riviere:2015:RCC

[1793] Marie-Karelle Riviere, Frédéric Dubois, and Sarah Zohar. Response to comments on ‘Competing designs for drug combination in phase I dosefinding clinical trials’ by G. Yin, R. Lin and N. Wages. Statistics in Medicine, 34(1):23–26, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wan:2015:STS

[1794] Hong Wan, Susan S. Ellenberg, and Keaven M. Anderson. Stepwise two-stage sample size adaptation. Statistics in Medicine, 34(1):27–38, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hampson:2015:ODC

[1795] Lisa V. Hampson and Christopher Jennison. Optimizing the data combination rule for seamless phase II/III clinical trials. Statistics in Medicine, 34(1):39–58, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kujala:2015:CSN

[1796] M. Kujala and J. Nevalainen. A case study of normalization, missing data and variable selection methods in lipidomics. Statistics in Medicine, 34 (1):59–73, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hof:2015:MMA

[1797] M. H. P. Hof and A. H. Zwinderman. A mixture model for the analysis of data derived from record linkage. Statistics in Medicine, 34(1):74–92, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hubbard:2015:CAC

[1798] Rebecca A. Hubbard, Rhondee Benjamin-Johnson, Tracy Onega, Rebecca Smith-Bindman, Weiwei Zhu, and Joshua J. Fenton. Classification accuracy of claims-based methods for identifying providers failing to meet performance targets. Statistics in Medicine, 34(1):93–105, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gruber:2015:ELI

[1799] Susan Gruber, Roger W. Logan, Inmaculada Jarrín, Susana Monge, and Miguel A. Hernán. Ensemble learning of inverse probability weights for marginal structural modeling in large observational datasets. Statistics in Medicine, 34(1):106–117, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sitlani:2015:GEE

[1800] Colleen M. Sitlani, Kenneth M. Rice, Thomas Lumley, Barbara McKnight, L. Adrienne Cupples, Christy L. Avery, Raymond Noordam, Bruno H. C. Stricker, Eric A. Whitsel, and Bruce M. Psaty. Generalized estimating equations for genome-wide association studies using longitudinal phenotype data. Statistics in Medicine, 34(1):118–130, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taguri:2015:PSA

[1801] Masataka Taguri and Yasutaka Chiba. A principal stratification approach for evaluating natural direct and indirect effects in the presence of treatment-induced intermediate confounding. Statistics in Medicine, 34(1):131–144, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

George:2015:SSP

[1802] Brandon George and Inmaculada Aban. Selecting a separable parametric spatiotemporal covariance structure for longitudinal imaging data. Statistics in Medicine, 34(1):145–161, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2015:IMM

[1803] Yen-Tsung Huang. Integrative modeling of multi-platform genomic data under the framework of mediation analysis. Statistics in Medicine, 34(1): 162–178, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collins:2015:AR

[1804] John Collins and Minh Huynh. Authors’ reply. Statistics in Medicine, 34(1):179–180, January 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hemming:2015:SWC

[1805] Karla Hemming, Richard Lilford, and Alan J. Girling. Stepped-wedge cluster randomised controlled trials: a generic framework including parallel and multiple-level designs. Statistics in Medicine, 34(2):181–196, January 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bersimis:2015:FDP

[1806] S. Bersimis, A. Sachlas, and T. Papaioannou. Flexible designs for phase II comparative clinical trials involving two response variables. Statistics in Medicine, 34(2):197–214, January 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hua:2015:MAT

[1807] Steven Y. Hua, Siyan Xu, and Ralph B. D’Agostino, Sr. Multiplicity adjustments in testing for bioequivalence. Statistics in Medicine, 34(2): 215–231, January 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Helms:2015:SBP

[1808] Hans-Joachim Helms, Norbert Benda, Jörg Zinserling, Thomas Kneib, and Tim Friede. Spline-based procedures for dose-finding studies with active control. Statistics in Medicine, 34(2):232–248, January 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ibrahim:2015:BPS

[1809] Joseph G. Ibrahim, Ming-Hui Chen, Mani Lakshminarayanan, Guanghan F. Liu, and Joseph F. Heyse. Bayesian probability of success for clinical trials using historical data. Statistics in Medicine, 34(2):249– 264, January 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:2015:ATB

[1810] Edward L. Korn, James J. Dignam, and Boris Freidlin. Assessing treatment benefit with competing risks not affected by the randomized treatment. Statistics in Medicine, 34(2):265–280, January 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2015:SSP

[1811] Peng Li and David T. Redden. Small sample performance of biascorrected sandwich estimators for cluster-randomized trials with binary outcomes. Statistics in Medicine, 34(2):281–296, January 30, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosner:2015:OAR

[1812] Bernard Rosner, Sara Hendrickson, and Walter Willett. Optimal allocation of resources in a biomarker setting. Statistics in Medicine, 34(2): 297–306, January 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2015:EIE

[1813] Ying Xu, K. F. Lam, Benjamin J. Cowling, and Yin Bun Cheung. Estimation of intervention effect using paired interval-censored data with clumping below lower detection limit. Statistics in Medicine, 34(2):307– 316, January 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2015:PAP

[1814] Gong Chen, Hua Zhong, Anton Belousov, and Viswanath Devanarayan. A PRIM approach to predictive-signature development for patient stratification. Statistics in Medicine, 34(2):317–342, January 30, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jin:2015:SMD

[1815] Zhi-Chao Jin, Xiao-Hua Zhou, and Jia He. Statistical methods for dealing with publication bias in meta-analysis. Statistics in Medicine, 34(2):

343–360, January 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2015:APM

[1816] Yong Chen, Chuan Hong, and Richard D. Riley. An alternative pseudolikelihood method for multivariate random-effects meta-analysis. Statistics in Medicine, 34(3):361–380, February 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gabriel:2015:CCB

[1817] Erin E. Gabriel, Michael C. Sachs, and Peter B. Gilbert. Comparing and combining biomarkers as principal surrogates for time-to-event clinical endpoints. Statistics in Medicine, 34(3):381–395, February 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brentnall:2015:CIM

[1818] Adam R. Brentnall, Jack Cuzick, John Field, and Stephen W. Duffy. A concordance index for matched case-control studies with applications in cancer risk. Statistics in Medicine, 36(21):396–405, February 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matsouaka:2015:PSS

[1819] Roland A. Matsouaka and Rebecca A. Betensky. Power and sample size calculations for the Wilcoxon–Mann–Whitney test in the presence of death-censored observations. Statistics in Medicine, 36(21):406–431, February 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chaurasia:2015:PTM

[1820] Ashok Chaurasia and Ofer Harel. Partial F -tests with multiply imputed data in the linear regression framework via coefficient of determination. Statistics in Medicine, 36(21):432–443, February 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lipsitz:2015:UJE

[1821] Stuart R. Lipsitz, Garrett M. Fitzmaurice, Alex Arriaga, Debajyoti Sinha, and Atul A. Gawande. Using the jackknife for estimation in log link Bernoulli regression models. Statistics in Medicine, 36(21):444–453, February 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davies:2015:MWI

[1822] Neil M. Davies, Stephanie von Hinke Kessler Scholder, Helmut Farbmacher, Stephen Burgess, Frank Windmeijer, and George Davey Smith. The many weak instruments problem and Mendelian randomization. Statistics in Medicine, 36(21):454–468, February 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2015:BCD

[1823] Bingrui Sun and Brajendra Sutradhar. Bivariate categorical data analysis using normal linear conditional multinomial probability model. Statistics in Medicine, 36(21):469–486, February 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arima:2015:ISB

[1824] Serena Arima. Item selection via Bayesian IRT models. Statistics in Medicine, 36(21):487–503, February 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2015:VAS

[1825] Inkyung Jung and Goeun Park. P -value approximations for spatial scan statistics using extreme value distributions. Statistics in Medicine, 36 (21):504–514, February 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2015:OST

[1826] Xinhua Liu and Zhezhen Jin. Optimal survival time-related cut-point with censored data. Statistics in Medicine, 36(21):515–524, February 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2015:LRT

[1827] Yi-Fan Chen, Jonathan G. Yabes, Maria M. Brooks, Sonia Singh, and Lisa A. Weissfeld. A likelihood ratio test for nested proportions. Statistics in Medicine, 36(21):525–538, February 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2015:NMI

[1828] Anonymous. A note on modelling incomplete contingency tables with censored cells. Statistics in Medicine, 36(21):539–540, February 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gares:2015:OTS

[1829] Valérie Garès, Sandrine Andrieu, Jean-François Dupuy, and Nicolas Savy. An omnibus test for several hazard alternatives in prevention randomized controlled clinical trials. Statistics in Medicine, 34(4):541–557, February 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Loux:2015:RMP

[1830] Travis M. Loux. Randomization, matching, and propensity scores in the design and analysis of experimental studies with measured baseline covariates. Statistics in Medicine, 34(4):558–570, February 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2015:TSM

[1831] Hanyu Yang, James A. Cranford, Runze Li, and Anne Buu. Two-stage model for time-varying effects of discrete longitudinal covariates with applications in analysis of daily process data. Statistics in Medicine, 34 (4):571–581, February 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Piepho:2015:MIN

[1832] Hans-Peter Piepho, Laurence V. Madden, and Emlyn R. Williams. Multiplicative interaction in network meta-analysis. Statistics in Medicine, 34(4):582–594, February 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2015:EPB

[1833] Y. Li, B. I. Graubard, P. Huang, and J. L. Gastwirth. Extension of the Peters–Belson method to estimate health disparities among multiple groups using logistic regression with survey data. Statistics in Medicine, 34(4):595–612, February 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2015:MVB

[1834] Yu Jiang, Steve Simon, Matthew S. Mayo, and Byron J. Gajewski. Modeling and validating Bayesian accrual models on clinical data and simulations using adaptive priors. Statistics in Medicine, 34(4):613–629, February 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kapur:2015:BME

[1835] Kush Kapur, Xue Li, Emily A. Blood, and Donald Hedeker. Bayesian mixed-effects location and scale models for multivariate longitudinal outcomes: an application to ecological momentary assessment data. Statistics in Medicine, 34(4):630–651, February 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2015:MTF

[1836] Xin He, G. A. Whitmore, Geok Yan Loo, Marc C. Hochberg, and Mei-Ling Ting Lee. A model for time to fracture with a shock stream superimposed on progressive degradation: the study of osteoporotic fractures. Statistics in Medicine, 34(4):652–663, February 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2015:BAI

[1837] Gang Li and Xuyang Lu. A Bayesian approach for instrumental variable analysis with censored time-to-event outcome. Statistics in Medicine, 34 (4):664–684, February 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2015:CTC

[1838] Le Kang, Weijie Chen, Nicholas A. Petrick, and Brandon D. Gallas. Comparing two correlated C indices with right-censored survival outcome: a one-shot nonparametric approach. Statistics in Medicine, 34 (4):685–703, February 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2015:ACC

[1839] Hsiang-Chun Chen and Thomas E. Wehrly. Assessing correlation of clustered mixed outcomes from a multivariate generalized linear mixed model. Statistics in Medicine, 34(4):704–720, February 20, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mavridis:2015:AUD

[1840] Dimitris Mavridis, Ian R. White, Julian P. T. Higgins, Andrea Cipriani, and Georgia Salanti. Allowing for uncertainty due to missing continuous outcome data in pairwise and network meta-analysis. Statistics in Medicine, 34(5):721–741, February 28, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dimitrakopoulou:2015:AUD

[1841] Vasiliki Dimitrakopoulou, Orestis Efthimiou, Stefan Leucht, and Georgia Salanti. Accounting for uncertainty due to ‘last observation carried forward’ outcome imputation in a meta-analysis model. Statistics in Medicine, 34(5):742–752, February 28, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neugebauer:2015:HDP

[1842] Romain Neugebauer, Julie A. Schmittdiel, Zheng Zhu, Jeremy A. Rassen, John D. Seeger, and Sebastian Schneeweiss. High-dimensional propensity score algorithm in comparative effectiveness research with time-varying interventions. Statistics in Medicine, 34(5):753–781, February 28, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2015:PCD

[1843] Kaifeng Lu. Power calculations for delta-adjusted pattern-mixture models. Statistics in Medicine, 34(5):782–795, February 28, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Benaglia:2015:SEP

[1844] Tatiana Benaglia, Christopher H. Jackson, and Linda D. Sharples. Survival extrapolation in the presence of cause specific hazards. Statistics in Medicine, 34(5):796–811, February 28, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Talbot:2015:CNC

[1845] Denis Talbot, Juli Atherton, Amanda M. Rossi, Simon L. Bacon, and Geneviève Lefebvre. A cautionary note concerning the use of stabilized weights in marginal structural models. Statistics in Medicine, 34(5):812– 823, February 28, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2015:SBI

[1846] An-Min Tang and Nian-Sheng Tang. Semiparametric Bayesian inference on skew-normal joint modeling of multivariate longitudinal and survival data. Statistics in Medicine, 34(5):824–843, February 28, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zeng:2015:ATA

[1847] Donglin Zeng, Emil Cornea, Jun Dong, Jean Pan, and Joseph G. Ibrahim. Assessing temporal agreement between central and local progression-free survival times. Statistics in Medicine, 34(5):844–858, February 28, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2015:BDFa

[1848] Beibei Guo and Yisheng Li. Bayesian dose-finding designs for combination of molecularly targeted agents assuming partial stochastic ordering. Statistics in Medicine, 34(5):859–875, February 28, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2015:GPL

[1849] Xiaoguang Wang, Jun Zhang, Liang Yu, and Guosheng Yin. Generalized partially linear single-index model for zero-inflated count data. Statistics in Medicine, 34(5):876–886, February 28, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Speiser:2015:RFC

[1850] Jaime Lynn Speiser, Valerie L. Durkalski, and William M. Lee. Random forest classification of etiologies for an orphan disease. Statistics in Medicine, 34(5):887–899, February 28, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2015:HIS

[1851] Stuart G. Baker, Ewoud Schuit, Ewout W. Steyerberg, Michael J. Pencina, Andrew Vickers, Karel G. M. Moons, Ben W. J. Mol, and Karen S. Lindeman. How to interpret a small increase in AUC with an additional risk prediction marker: decision analysis comes through. Statistics in Medicine, 34(5):900, February 28, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhong:2015:SSR

[1852] Yujie Zhong and Richard J. Cook. Sample size and robust marginal methods for cluster-randomized trials with censored event times. Statistics in Medicine, 34(6):901–923, March 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rigdon:2015:RIT

[1853] Joseph Rigdon and Michael G. Hudgens. Randomization inference for treatment effects on a binary outcome. Statistics in Medicine, 34(6): 924–935, March 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barnwell-Menard:2015:ECM

[1854] Jean-Louis Barnwell-Ménard, Qing Li, and Alan A. Cohen. Effects of categorization method, regression type, and variable distribution on the inflation of Type-I error rate when categorizing a confounding variable. Statistics in Medicine, 34(6):936–949, March 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2015:CNC

[1855] Ying Wu, Liang Zhao, Yan Hou, Kang Li, and Xiaohua Zhou. Correcting for non-compliance in randomized non-inferiority trials with active and placebo control using structural models. Statistics in Medicine, 34(6): 950–965, March 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walwyn:2015:MAA

[1856] Rebecca Walwyn and Chris Roberts. Meta-analysis of absolute mean differences from randomised trials with treatment-related clustering associated with care providers. Statistics in Medicine, 34(6):966–983, March 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turner:2015:PDB

[1857] Rebecca M. Turner, Dan Jackson, Yinghui Wei, Simon G. Thompson, and Julian P. T. Higgins. Predictive distributions for between-study heterogeneity and simple methods for their application in Bayesian metaanalysis. Statistics in Medicine, 34(6):984–998, March 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Balzer:2015:APM

[1858] Laura B. Balzer, Maya L. Petersen, Mark J. van der Laan, and the Search Consortium. Adaptive pair-matching in randomized trials with unbiased and efficient effect estimation. Statistics in Medicine, 34(6): 999–1011, March 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vermeulen:2015:IPM

[1859] Karel Vermeulen, Olivier Thas, and Stijn Vansteelandt. Increasing the power of the Mann–Whitney test in randomized experiments through flexible covariate adjustment. Statistics in Medicine, 34(6):1012–1030, March 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmidt:2015:SSC

[1860] René Schmidt, Robert Kwiecien, Andreas Faldum, Frank Berthold, Barbara Hero, and Sandra Ligges. Sample size calculation for the one-sample log-rank test. Statistics in Medicine, 34(6):1031–1040, March 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Muthen:2015:GMM

[1861] Bengt Muthén and Tihomir Asparouhov. Growth mixture modeling with non-normal distributions. Statistics in Medicine, 34(6):1041–1058, March 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cellamare:2015:RTS

[1862] Matteo Cellamare and Valeria Sambucini. A randomized two-stage design for phase II clinical trials based on a Bayesian predictive approach. Statistics in Medicine, 34(6):1059–1078, March 15, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2015:IFH

[1863] Anonymous. Incorporation of family history in clinical trials. Statistics in Medicine, 34(6):1079–1080, March 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Correa:2015:CLP

[1864] Thais R. Correa, Renato M. Assunção, and Marcelo A. Costa. A critical look at prospective surveillance using a scan statistic. Statistics in Medicine, 34(7):1081–1093, March 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kulldorff:2015:CCL

[1865] Martin Kulldorff and Ken Kleinman. Comments on ‘A critical look at prospective surveillance using a scan statistic’ by T. Correa, M. Costa, and R. Assunção. Statistics in Medicine, 34(7):1094–1095, March 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Correa:2015:RCC

[1866] Thais R. Correa, Renato M. Assunção, and Marcelo A. Costa. Response to commentary on ‘A critical look at prospective surveillance using a scan statistic’. Statistics in Medicine, 34(7):1096, March 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuss:2015:SMM

[1867] O. Kuss. Statistical methods for meta-analyses including information from studies without any events-add nothing to nothing and succeed nevertheless. Statistics in Medicine, 34(7):1097–1116, March 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stratton:2015:SSC

[1868] Kelly G. Stratton, Andrea J. Cook, Lisa A. Jackson, and Jennifer C. Nelson. Simulation study comparing exposure matching with regression adjustment in an observational safety setting with group sequential monitoring. Statistics in Medicine, 34(7):1117–1133, March 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gajewski:2015:BEC

[1869] Byron J. Gajewski, Scott M. Berry, Melanie Quintana, Mamatha Pasnoor, Mazen Dimachkie, Laura Herbelin, and Richard Barohn. Building efficient comparative effectiveness trials through adaptive designs, utility functions, and accrual rate optimization: finding the sweet spot. Statistics in Medicine, 34(7):1134–1149, March 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2015:PBI

[1870] Yun Li, Yoonseok Lee, Robert A. Wolfe, Hal Morgenstern, Jinyao Zhang, Friedrich K. Port, and Bruce M. Robinson. On a preference-based instrumental variable approach in reducing unmeasured confounding-by-indication. Statistics in Medicine, 34(7):1150–1168, March 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Geng:2015:OTR

[1871] Yuan Geng, Hao Helen Zhang, and Wenbin Lu. On optimal treatment regimes selection for mean survival time. Statistics in Medicine, 34(7): 1169–1184, March 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lim:2015:RRR

[1872] Changwon Lim. Robust ridge regression estimators for nonlinear models with applications to high throughput screening assay data. Statistics in Medicine, 34(7):1185–1198, March 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farcomeni:2015:LQR

[1873] Alessio Farcomeni and Sara Viviani. Longitudinal quantile regression in the presence of informative dropout through longitudinal-survival joint modeling. Statistics in Medicine, 34(7):1199–1213, March 30, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Junior:2015:PPA

[1874] Jony Arrais Pinto Junior, Dani Gamerman, Marina Silva Paez, and Regina Helena Fonseca Alves. Point pattern analysis with spatially varying covariate effects, applied to the study of cerebrovascular deaths. Statistics in Medicine, 34(7):1214–1226, March 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tao:2015:LVM

[1875] Yebin Tao, Brisa N. Sánchez, and Bhramar Mukherjee. Latent variable models for gene-environment interactions in longitudinal studies with multiple correlated exposures. Statistics in Medicine, 34(7):1227–1241, March 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nawarathna:2015:HME

[1876] Lakshika S. Nawarathna and Pankaj K. Choudhary. A heteroscedastic measurement error model for method comparison data with replicate measurements. Statistics in Medicine, 34(7):1242–1258, March 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gelman:2015:BFR

[1877] Andrew Gelman. Bayesian and frequentist regression methods. Statistics in Medicine, 34(7):1259–1260, March 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mander:2015:PIB

[1878] Adrian P. Mander and Michael J. Sweeting. A product of independent beta probabilities dose escalation design for dual-agent phase I trials. Statistics in Medicine, 34(8):1261–1276, April 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teeple:2015:ATD

[1879] Elizabeth A. Teeple and Elizabeth R. Brown. Adjusting for timedependent sensitivity in an illness-death model, with application to mother-to-child transmission of HIV. Statistics in Medicine, 34(8):1277– 1292, April 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yao:2015:ERC

[1880] Wenliang Yao, Zhaohai Li, and Barry I. Graubard. Estimation of ROC curve with complex survey data. Statistics in Medicine, 34(8):1293–1303, April 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williams:2015:ORN

[1881] James D. Williams, Jeffrey B. Birch, and Abdel-Salam G. Abdel-Salam. Outlier robust nonlinear mixed model estimation. Statistics in Medicine, 34(8):1304–1316, April 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carreras:2015:ASD

[1882] Máximo Carreras, Georg Gutjahr, and Werner Brannath. Adaptive seamless designs with interim treatment selection: a case study in oncology. Statistics in Medicine, 34(8):1317–1333, April 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rompaye:2015:EHP

[1883] Bart Van Rompaye, Marie Eriksson, and Els Goetghebeur. Evaluating hospital performance based on excess cause-specific incidence. Statistics in Medicine, 34(8):1334–1350, April 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Willemsen:2015:MBM

[1884] Sten P. Willemsen, Paul H. C. Eilers, Régine P. M. Steegers-Theunissen, and Emmanuel Lesaffre. A multivariate Bayesian model for embryonic growth. Statistics in Medicine, 34(8):1351–1365, April 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ortega:2015:PSB

[1885] Edwin M. M. Ortega, Gauss M. Cordeiro, Ana K. Campelo, Michael W. Kattan, and Vicente G. Cancho. A power series beta Weibull regression model for predicting breast carcinoma. Statistics in Medicine, 34(8): 1366–1388, April 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buonaccorsi:2015:SRC

[1886] John P. Buonaccorsi, Ingvild Dalen, Petter Laake, Anette Hjartåker, Dagrun Engeset, and Magne Thoresen. Sensitivity of regression calibration to non-perfect validation data with application to the Norwegian women and cancer study. Statistics in Medicine, 34(8):1389–1403, April 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2015:DPM

[1887] Lili Zhao, Jingchunzi Shi, Tempie H. Shearon, and Yi Li. A Dirichlet process mixture model for survival outcome data: assessing nationwide kidney transplant centers. Statistics in Medicine, 34(8):1404–1416, April 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robertson:2015:CBS

[1888] David S. Robertson, A. Toby Prevost, and Jack Bowden. Correcting for bias in the selection and validation of informative diagnostic tests. Statistics in Medicine, 34(8):1417–1437, April 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taguri:2015:CCN

[1889] Masataka Taguri. Comments on ‘A cautionary note concerning the use of stabilized weights in marginal structural models’ by D. Talbot, J. Atherton, A. M. Rossi, S. L. Bacon, and G. Lefebvre. Statistics in Medicine, 34(8):1438–1439, April 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lagona:2015:C

[1890] Francesco Lagona. Correction. Statistics in Medicine, 34(8):1440, April 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dawson:2015:DII

[1891] Ree Dawson and Philip W. Lavori. Design and inference for the intentto-treat principle using adaptive treatment. Statistics in Medicine, 34(9): 1441–1453, April 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ramchandani:2015:MIR

[1892] Ritesh Ramchandani, Dianne M. Finkelstein, and David A. Schoenfeld. A model-informed rank test for right-censored data with intermediate states. Statistics in Medicine, 34(9):1454–1466, April 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2015:OOH

[1893] Zhiwei Zhang, Chenguang Wang, and James F. Troendle. Optimizing the order of hypotheses in serial testing of multiple endpoints in clinical trials. Statistics in Medicine, 34(9):1467–1482, April 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fong:2015:CPT

[1894] Youyi Fong, Chongzhi Di, and Sallie Permar. Change point testing in logistic regression models with interaction term. Statistics in Medicine, 34(9):1483–1494, April 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chiou:2015:RBE

[1895] S. Chiou, S. Kang, and J. Yan. Rank-based estimating equations with general weight for accelerated failure time models: an induced smoothing approach. Statistics in Medicine, 34(9):1495–1510, April 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marozzi:2015:MMT

[1896] Marco Marozzi. Multivariate multidistance tests for high-dimensional low sample size case-control studies. Statistics in Medicine, 34(9):1511–1526, April 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:2015:SDS

[1897] Xiang-Nan Feng, Guo-Chang Wang, Yi-Fan Wang, and Xin-Yuan Song. Structure detection of semiparametric structural equation models with Bayesian adaptive group lasso. Statistics in Medicine, 34(9):1527–1547, April 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Botella-Rocamora:2015:UMF

[1898] P. Botella-Rocamora, M. A. Martinez-Beneito, and S. Banerjee. A unifying modeling framework for highly multivariate disease mapping. Statistics in Medicine, 34(9):1548–1559, April 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2015:EST

[1899] Fangyao Chen, Yuqiang Xue, Ming T. Tan, and Pingyan Chen. Efficient statistical tests to compare Youden index: accounting for contingency correlation. Statistics in Medicine, 34(9):1560–1576, April 30, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2015:PMP

[1900] Hong Zhang, Jinfeng Xu, Ning Jiang, Xiaohua Hu, and Zewei Luo. PLNseq: a multivariate Poisson lognormal distribution for highthroughput matched RNA-sequencing read count data. Statistics in Medicine, 34(9):1577–1589, April 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alonso:2015:NMA

[1901] Ariel Alonso, Elasma Milanzi, Geert Molenberghs, Christophe Buyck, and Luc Bijnens. A new modeling approach for quantifying expert opinion in the drug discovery process. Statistics in Medicine, 34(9):1590– 1604, April 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2015:BRD

[1902] Li Tang, Robert H. Lyles, Caroline C. King, David D. Celentano, and Yungtai Lo. Binary regression with differentially misclassified response and exposure variables. Statistics in Medicine, 34(9):1605–1620, April 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2015:MSR

[1903] Baiming Zou, Bo Jin, Gary G. Koch, Haibo Zhou, Stephen E. Borst, Sandeep Menon, and Jonathan J. Shuster. On model selections for repeated measurement data in clinical studies. Statistics in Medicine, 34 (10):1621–1633, May 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maruo:2015:IMD

[1904] K. Maruo, N. Isogawa, and M. Gosho. Inference of median difference based on the Box–Cox model in randomized clinical trials. Statistics in Medicine, 34(10):1634–1644, May 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2015:PRP

[1905] Debashis Ghosh, Yeying Zhu, and Donna L. Coffman. Penalized regression procedures for variable selection in the potential outcomes framework. Statistics in Medicine, 34(10):1645–1658, May 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demler:2015:TCG

[1906] Olga V. Demler, Nina P. Paynter, and Nancy R. Cook. Tests of calibration and goodness-of-fit in the survival setting. Statistics in Medicine, 34(10):1659–1680, May 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2015:BCR

[1907] Suyu Liu, Haitao Pan, Jielai Xia, Qin Huang, and Ying Yuan. Bridging continual reassessment method for phase I clinical trials in different ethnic populations. Statistics in Medicine, 34(10):1681–1694, May 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fong:2015:CWE

[1908] Youyi Fong and Peter Gilbert. Calibration weighted estimation of semiparametric transformation models for two-phase sampling. Statistics in Medicine, 34(10):1695–1707, May 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kong:2015:UDC

[1909] Jing Kong, Sijian Wang, and Grace Wahba. Using distance covariance for improved variable selection with application to learning genetic risk models. Statistics in Medicine, 34(10):1708–1720, May 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2015:BDFb

[1910] Beibei Guo and Ying Yuan. A Bayesian dose-finding design for phase I/II clinical trials with nonignorable dropouts. Statistics in Medicine, 34(10):1721–1732, May 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2015:ETE

[1911] Lili Zhao, Dai Feng, Brian Neelon, and Marc Buyse. Evaluation of treatment efficacy using a Bayesian mixture piecewise linear model of longitudinal biomarkers. Statistics in Medicine, 34(10):1733–1746, May 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Long:2015:SPR

[1912] Dustin M. Long, Michael G. Hudgens, and Chih-Da Wu. Surrogates of protection in repeated low-dose challenge experiments. Statistics in Medicine, 34(10):1747–1760, May 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rodrigues-Motta:2015:MEM

[1913] Mariana Rodrigues-Motta, Diana Milena Galvis Soto, Victor H. Lachos, Filidor Vilca, Valéria Troncoso Baltar, Eliseu Verly Junior, Regina Mara Fisberg, and Dirce Maria Lobo Marchioni. A mixed-effect model for positive responses augmented by zeros. Statistics in Medicine, 34(10): 1761–1778, May 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ferrante:2015:BCF

[1914] Luigi Ferrante, Edlira Skrami, Rosaria Gesuita, and Roberto Cameriere. Bayesian calibration for forensic age estimation. Statistics in Medicine, 34(10):1779–1790, May 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2015:VDN

[1915] Zonghui Hu, Dean A. Follmann, and Kazutoyo Miura. Vaccine design via nonnegative lasso-based variable selection. Statistics in Medicine, 34(10):1791–1798, May 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brodersen:2015:PRO

[1916] John Brodersen. Patient-reported outcomes: measurement, implementation and interpretation. Statistics in Medicine, 34(10):1799–1800, May 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2015:APT

[1917] Song Yang and Ross L. Prentice. Assessing potentially time-dependent treatment effect from clinical trials and observational studies for survival data, with applications to the Women’s health initiative combined hormone therapy trial. Statistics in Medicine, 34(11):1801–1817, May 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Loh:2015:RTA

[1918] Wei-Yin Loh, Xu He, and Michael Man. A regression tree approach to identifying subgroups with differential treatment effects. Statistics in Medicine, 34(11):1818–1833, May 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nicholas:2015:ICA

[1919] Katherine Nicholas, Sharon D. Yeatts, Wenle Zhao, Jody Ciolino, Keith Borg, and Valerie Durkalski. The impact of covariate adjustment at randomization and analysis for binary outcomes: understanding differences between superiority and noninferiority trials. Statistics in Medicine, 34 (11):1834–1840, May 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jolani:2015:ISM

[1920] Shahab Jolani, Thomas P. A. Debray, Hendrik Koffijberg, Stef van Buuren, and Karel G. M. Moons. Imputation of systematically missing predictors in an individual participant data meta-analysis: a generalized approach using MICE. Statistics in Medicine, 34(11):1841–1863, May 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Landsman:2015:ESP

[1921] V. Landsman, W. Y. W. Lou, and B. I. Graubard. Estimating survival probabilities by exposure levels: utilizing vital statistics and complex survey data with mortality follow-up. Statistics in Medicine, 34(11): 1864–1875, May 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2015:EMB

[1922] Soeun Kim, Catherine A. Sugar, and Thomas R. Belin. Evaluating model-based imputation methods for missing covariates in regression models with interactions. Statistics in Medicine, 34(11):1876–1888, May 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cotterill:2015:BMS

[1923] Amy Cotterill and John Whitehead. Bayesian methods for setting sample sizes and choosing allocation ratios in phase II clinical trials with time-to-event endpoints. Statistics in Medicine, 34(11):1889–1903, May 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2015:TPT

[1924] Zonghui Hu and Michael Proschan. Two-part test of vaccine effect. Statistics in Medicine, 34(11):1904–1911, May 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoyer:2015:MAD

[1925] A. Hoyer and O. Kuss. Meta-analysis of diagnostic tests accounting for disease prevalence: a new model using trivariate copulas. Statistics in Medicine, 34(11):1912–1924, May 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2015:AIM

[1926] Ji Lin and Robert H. Lyles. Accounting for informatively missing data in logistic regression by means of reassessment sampling. Statistics in Medicine, 34(11):1925–1939, May 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Voulgaraki:2015:EDR

[1927] Anastasia Voulgaraki, Rong Wei, and Benjamin Kedem. Estimation of death rates in US states with small subpopulations. Statistics in Medicine, 34(11):1940–1952, May 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oh:2015:APC

[1928] Cheongeun Oh and Theodore R. Holford. Age–period–cohort approaches to back-calculation of cancer incidence rate. Statistics in Medicine, 34 (11):1953–1964, May 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alvarez-Iglesias:2015:SCS

[1929] Alberto Alvarez-Iglesias, John Newell, Carl Scarrott, and John Hinde. Summarising censored survival data using the mean residual life function. Statistics in Medicine, 34(11):1965–1976, May 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kleinman:2015:RRE

[1930] Alan Kleinman. Random recombination and evolution of drug resistance. Statistics in Medicine, 34(11):1977–1980, May 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nolen:2015:ARL

[1931] Tracy L. Nolen, Michael G. Hudgens, Pranab K. Senb, and Gary G. Koch. Analysis of repeated low-dose challenge studies. Statistics in Medicine, 34(12):1981–1992, May 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jensen:2015:EMO

[1932] Signe M. Jensen, Christian B. Pipper, and Christian Ritz. Evaluation of multi-outcome longitudinal studies. Statistics in Medicine, 34(12): 1993–2003, May 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Su:2015:BMC

[1933] Li Su and Michael J. Daniels. Bayesian modeling of the covariance structure for irregular longitudinal data using the partial autocorrelation function. Statistics in Medicine, 34(12):2004–2018, May 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becque:2015:CML

[1934] Taeko Becque, Ian R. White, and Mark Haggard. A causal model for longitudinal randomised trials with time-dependent non-compliance. Statistics in Medicine, 34(12):2019–2034, May 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2015:DTC

[1935] Rui Wang, David A. Schoenfeld, Bettina Hoeppner, and A. Eden Evins. Detecting treatment-covariate interactions using permutation methods. Statistics in Medicine, 34(12):2035–2047, May 30, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2015:BSS

[1936] John Whitehead, Faye Cleary, and Amanda Turner. Bayesian sample sizes for exploratory clinical trials comparing multiple experimental treatments with a control. Statistics in Medicine, 34(12):2048–2061, May 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turner:2015:BFA

[1937] N. L. Turner, S. Dias, A. E. Ades, and N. J. Welton. A Bayesian framework to account for uncertainty due to missing binary outcome data in pairwise meta-analysis. Statistics in Medicine, 34(12):2062–2080, May 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2015:SVT

[1938] Richard D. Riley, Ikhlaaq Ahmed, Thomas P. A. Debray, Brian H. Willis, J. Pieter Noordzij, Julian P. T. Higgins, and Jonathan J. Deeks. Summarising and validating test accuracy results across multiple studies for use in clinical practice. Statistics in Medicine, 34(13):2081–2103, June 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2015:BSM

[1939] Weining Shen, Jing Ning, and Ying Yuan. Bayesian sequential monitoring design for two-arm randomized clinical trials with noncompliance. Statistics in Medicine, 34(13):2104–2115, June 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stanley:2015:NFN

[1940] T. D. Stanley and Hristos Doucouliagos. Neither fixed nor random: weighted least squares meta-analysis. Statistics in Medicine, 34(13): 2116–2127, June 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Englert:2015:MPH

[1941] Stefan Englert and Meinhard Kieser. Methods for proper handling of overrunning and underrunning in phase II designs for oncology trials. Statistics in Medicine, 34(13):2128–2137, June 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cotterill:2015:PDD

[1942] Amy Cotterill, Daniel Lorand, Jixian Wang, and Thomas Jaki. A practical design for a dual-agent dose-escalation trial that incorporates pharmacokinetic data. Statistics in Medicine, 34(13):2138–2164, June 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kondo:2015:FME

[1943] Yumi Kondo, Yinshan Zhao, and John Petkau. A flexible mixed-effect negative binomial regression model for detecting unusual increases in MRI lesion counts in individual multiple sclerosis patients. Statistics in Medicine, 34(13):2165–2180, June 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:2015:JMS

[1944] A. Lawrence Gould, Mark Ernest Boye, Michael J. Crowther, Joseph G. Ibrahim, George Quartey, Sandrine Micallef, and Frederic Y. Bois. Joint modeling of survival and longitudinal non-survival data: current methods and issues. report of the DIA Bayesian joint modeling working group. Statistics in Medicine, 34(14):2181–2195, June 30, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rizopoulos:2015:CJM

[1945] Dimitris Rizopoulos. Comments on ‘Joint modeling of survival and longitudinal non-survival data: current methods and issues. Report of the DIA Bayesian Joint Modeling Working Group’. Statistics in Medicine, 34(14):2196–2197, June 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farcomeni:2015:DJM

[1946] Alessio Farcomeni, Bhuvanesh Pareek, and Pulak Ghosh. Discussion on ‘joint modeling of survival and longitudinal non-survival data’ by Gould et al. Statistics in Medicine, 34(14):2198–2199, June 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Price:2015:CJM

[1947] Dionne L. Price and Yan Wang. Commentary on ‘joint modeling of survival and longitudinal non-survival data: current methods and issues’. Statistics in Medicine, 34(14):2200–2201, June 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:2015:RDJ

[1948] A. Lawrence Gould, Mark Ernest Boye, Michael J. Crowther, Joseph G. Ibrahim, George Quartey, Sandrine Micallef, and Frederic Y. Bois. Responses to discussants of ‘joint modeling of survival and longitudinal nonsurvival data: current methods and issues. report of the DIA Bayesian joint modeling working group’. Statistics in Medicine, 34(14):2202–2203, June 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qiu:2015:SCD

[1949] Peihua Qiu and Dongdong Xiang. Surveillance of cardiovascular diseases using a multivariate dynamic screening system. Statistics in Medicine, 34(14):2204–2221, June 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2015:ICD

[1950] Meijuan Li, Tinghui Yu, and Yun-Fu Hu. The impact of companion diagnostic device measurement performance on clinical validation of personalized medicine. Statistics in Medicine, 34(14):2222–2234, June 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wan:2015:BEC

[1951] Fei Wan, Dylan Small, Justin E. Bekelman, and Nandita Mitra. Bias in estimating the causal hazard ratio when using two-stage instrumental variable methods. Statistics in Medicine, 34(14):2235–2265, June 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sattar:2015:FMP

[1952] Abdus Sattar, Sanjoy K. Sinha, Xiao-Feng Wang, and Yehua Li. Frailty models for pneumonia to death with a left-censored covariate. Statistics in Medicine, 34(14):2266–2280, June 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rybin:2015:PNR

[1953] Denis Rybin, Gheorghe Doros, Michael J. Pencina, and Maurizio Fava. Placebo non-response measure in sequential parallel comparison design studies. Statistics in Medicine, 34(15):2281–2293, July 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jansen:2015:NMA

[1954] J. P. Jansen, M. C. Vieira, and S. Cope. Network meta-analysis of longitudinal data using fractional polynomials. Statistics in Medicine, 34(15): 2294–2311, July 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McGinley:2015:NMF

[1955] James S. McGinley, Patrick J. Curran, and Donald Hedeker. A novel modeling framework for ordinal data defined by collapsed counts. Statistics in Medicine, 34(15):2312–2324, July 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kwak:2015:COC

[1956] Minjung Kwak, Sang-Won Um, and Sin-Ho Jung. Comparison of operational characteristics for binary tests with clustered data. Statistics in Medicine, 34(15):2325–2333, July 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Geneletti:2015:BRD

[1957] Sara Geneletti, Aidan G. O’Keeffe, Linda D. Sharples, Sylvia Richardson, and Gianluca Baio. Bayesian regression discontinuity designs: incorporating clinical knowledge in the causal analysis of primary care data. Statistics in Medicine, 34(15):2334–2352, July 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Khudyakov:2015:ICM

[1958] Polyna Khudyakov, Malka Gorfine, David Zucker, and Donna Spiegelman. The impact of covariate measurement error on risk prediction. Statistics in Medicine, 34(15):2353–2367, July 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2015:MBS

[1959] Zhuangyu Cai and Babette A. Brumback. Model-based standardization to adjust for unmeasured cluster-level confounders with complex survey data. Statistics in Medicine, 34(15):2368–2380, July 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Little:2015:ITA

[1960] Roderick Little and Shan Kang. Intention-to-treat analysis with treatment discontinuation and missing data in clinical trials. Statistics in Medicine, 34(16):2381–2390, July 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2015:RPM

[1961] Chengcheng Liu, Sarah J. Ratcliffe, and Wensheng Guo. A random pattern mixture model for ordinal outcomes with informative dropouts. Statistics in Medicine, 34(16):2391–2402, July 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kieser:2015:TSD

[1962] Meinhard Kieser and Geraldine Rauch. Two-stage designs for cross-over bioequivalence trials. Statistics in Medicine, 34(16):2403–2416, July 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zeng:2015:HTT

[1963] Donglin Zeng, Fei Gao, Kuolung Hu, Catherine Jia, and Joseph G. Ibrahim. Hypothesis testing for two-stage designs with over or under enrollment. Statistics in Medicine, 34(16):2417–2426, July 20, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2015:PCR

[1964] Tianle Chen, Yanyuan Ma, and Yuanjia Wang. Predicting cumulative risk of disease onset by redistributing weights. Statistics in Medicine, 34(16):2427–2443, July 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hardouin:2015:PSS

[1965] Jean-Benoit Hardouin, Myriam Blanchin, Mohand-Larbi Feddag, Tanguy Le Néel, Bastien Perrot, and Véronique Sébille. Power and sample size determination for group comparison of patient-reported outcomes using polytomous Rasch models. Statistics in Medicine, 34(16):2444–2455, July 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thom:2015:SSM

[1966] Howard H. Z. Thom, Christopher H. Jackson, Daniel Commenges, and Linda D. Sharples. State selection in Markov models for panel data with application to psoriatic arthritis. Statistics in Medicine, 34(16): 2456–2475, July 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gonnermann:2015:NSY

[1967] Andrea Gonnermann, Theodor Framke, Anika Großhennig, and Armin Koch. No solution yet for combining two independent studies in the presence of heterogeneity. Statistics in Medicine, 34(16):2476–2480, July 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2015:MMA

[1968] Richard D. Riley, Eleni G. Elia, Gemma Malin, Karla Hemming, and Malcolm P. Price. Multivariate meta-analysis of prognostic factor studies with multiple cut-points and/or methods of measurement. Statistics in Medicine, 34(17):2481–2496, July 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2015:WMS

[1969] Yimei Li and Qiang Zhang. A Weibull multi-state model for the dependence of progression-free survival and overall survival. Statistics in Medicine, 34(17):2497–2513, July 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Betensky:2015:CSA

[1970] Rebecca A. Betensky, Jackie Szymonifka, Eudocia Q. Lee, Catherine L. Nutt, and Tracy T. Batchelor. Computationally simple analysis of matched, outcome-based studies of ordinal disease states. Statistics in Medicine, 34(17):2514–2527, July 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Branders:2015:BHR

[1971] Samuel Branders and Pierre Dupont. A balanced hazard ratio for risk group evaluation from survival data. Statistics in Medicine, 34(17):2528– 2543, July 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mitchell:2015:PFS

[1972] Emily M. Mitchell, Robert H. Lyles, and Enrique F. Schisterman. Positing, fitting, and selecting regression models for pooled biomarker data. Statistics in Medicine, 34(17):2544–2558, July 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neelon:2015:SQR

[1973] Brian Neelon, Fan Li, Lane F. Burgette, and Sara E. Benjamin Neelon. A spatiotemporal quantile regression model for emergency department expenditures. Statistics in Medicine, 34(17):2559–2575, July 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2015:CSS

[1974] Jianrong Wu. Comments on ‘Sample size calculation for the proportional hazards cure model’ by Songfeng Wang, Jiajia Zhang and Wenbin Lu. Statistics in Medicine, 34(17):2576–2577, July 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2015:ARC

[1975] Songfeng Wang, Jiajia Zhang, and Wenbin Lu. Authors’ reply to comments on ‘sample size calculation for the proportional hazards cure model’. Statistics in Medicine, 34(17):2578–2580, July 30, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kimani:2015:EAS

[1976] Peter K. Kimani, Susan Todd, and Nigel Stallard. Estimation after subpopulation selection in adaptive seamless trials. Statistics in Medicine, 34(18):2581–2601, August 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colantuoni:2015:LPB

[1977] Elizabeth Colantuoni and Michael Rosenblum. Leveraging prognostic baseline variables to gain precision in randomized trials. Statistics in Medicine, 34(18):2602–2617, August 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pfeiffer:2015:UMS

[1978] R. M. Pfeiffer and R. Riedl. On the use and misuse of scalar scores of confounders in design and analysis of observational studies. Statistics in Medicine, 36(27):2618–2635, August 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hermann:2015:FSG

[1979] Philipp Hermann, Tomás Mrkvicka, Torsten Mattfeldt, Mária Minárová, Katerina Helisová, Orietta Nicolis, Fabian Wartner, and Milan Stehlík. Fractal and stochastic geometry inference for breast cancer: a case study with random fractal models and Quermass-interaction process. Statistics in Medicine, 36(27):2636–2661, August 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Good:2015:PMC

[1980] Norm M. Good, Krithika Suresh, Graeme P. Young, Trevor J. Lockett, Finlay A. Macrae, and Jeremy M. G. Taylor. A prediction model for colon cancer surveillance data. Statistics in Medicine, 36(27):2662–2675, August 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Talbot:2015:ARC

[1981] Denis Talbot, Juli Atherton, Geneviève Lefebvre, Amanda M. Rossi, and Simon L. Bacon. Authors’ reply to comments on “A cautionary note concerning the use of stabilized weights in marginal structural models”. Statistics in Medicine, 36(27):2676–2677, August 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bratton:2015:CMP

[1982] Daniel J. Bratton, Babak Choodari-Oskooei, Patrick P. J. Phillips, Matthew R. Sydes, and Mahesh K. B. Parmar. Comments on ‘A modest proposal for dropping poor arms in clinical trials’ by Proschan and Dodd. Statistics in Medicine, 36(27):2678–2679, August 15, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marozzi:2015:C

[1983] Marco Marozzi. Correction. Statistics in Medicine, 36(27):2680, August 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:2015:ITE

[1984] J. N. S. Matthews and N. H. Badi. Inconsistent treatment estimates from mis-specified logistic regression analyses of randomized trials. Statistics in Medicine, 34(19):2681–2694, August 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2015:DOT

[1985] Jing Zhang, Haoda Fu, and Bradley P. Carlin. Detecting outlying trials in network meta-analysis. Statistics in Medicine, 36(27):2695–2707, August 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heimann:2015:EAI

[1986] Günter Heimann, Mark Von Tress, and Mauro Gasparini. Exact and asymptotic inference in clinical trials with small event rates under inverse sampling. Statistics in Medicine, 36(27):2708–2724, August 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2015:SDF

[1987] Na Hu, Lan Huang, and Ram C. Tiwari. Signal detection in FDA AERS database using Dirichlet process. Statistics in Medicine, 36(27):2725– 2742, August 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duan:2015:EVI

[1988] Ran Duan and Haoda Fu. Estimate variable importance for recurrent event outcomes with an application to identify hypoglycemia risk factors. Statistics in Medicine, 36(27):2743–2754, August 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salmeron:2015:RMC

[1989] Diego Salmerón, Juan A. Cano, and María D. Chirlaque. Reducing Monte Carlo error in the Bayesian estimation of risk ratios using logbinomial regression models. Statistics in Medicine, 36(27):2755–2767, August 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsu:2015:AAF

[1990] Chiu-Hsieh Hsu, Jeremy M. G. Taylor, and Chengcheng Hu. Analysis of accelerated failure time data with dependent censoring using auxiliary variables via nonparametric multiple imputation. Statistics in Medicine, 36(27):2768–2780, August 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kicinski:2015:PBM

[1991] Michal Kicinski, David A. Springate, and Evangelos Kontopantelis. Publication bias in meta-analyses from the Cochrane database of systematic reviews. Statistics in Medicine, 34(20):2781–2793, September 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hong:2015:IIP

[1992] Hwanhee Hong, Haoda Fu, Karen L. Price, and Bradley P. Carlin. Incorporation of individual-patient data in network meta-analysis for multiple continuous endpoints, with application to diabetes treatment. Statistics in Medicine, 34(20):2794–2819, September 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2015:BMS

[1993] Jiaqing Chen and Yangxin Huang. A Bayesian mixture of semiparametric mixed-effects joint models for skewed-longitudinal and time-to-event data. Statistics in Medicine, 34(20):2820–2843, September 10, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2015:STI

[1994] Zhuokai Li, Hai Liu, and Wanzhu Tu. A sexually transmitted infection screening algorithm based on semiparametric regression models. Statistics in Medicine, 34(20):2844–2857, September 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goeyvaerts:2015:MDA

[1995] Nele Goeyvaerts, Elke Leuridan, Christel Faes, Pierre Van Damme, and Niel Hens. Multi-disease analysis of maternal antibody decay using nonlinear mixed models accounting for censoring. Statistics in Medicine, 34 (20):2858–2871, September 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pomann:2015:SSC

[1996] Gina-Maria Pomann, Elizabeth M. Sweeney, Daniel S. Reich, Ana-Maria Staicu, and Russell T. Shinohara. Scan-stratified case-control sampling for modeling blood-brain barrier integrity in multiple sclerosis. Statistics in Medicine, 34(20):2872–2880, September 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:2015:MAT

[1997] Yinghui Wei, Patrick Royston, Jayne F. Tierney, and Mahesh K. B. Parmar. Meta-analysis of time-to-event outcomes from randomized trials using restricted mean survival time: application to individual participant data. Statistics in Medicine, 34(21):2881–2898, September 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McIsaac:2015:AST

[1998] Michael A. McIsaac and Richard J. Cook. Adaptive sampling in twophase designs: a biomarker study for progression in arthritis. Statistics in Medicine, 34(21):2899–2912, September 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kunz:2015:MAR

[1999] Lauren M. Kunz, Sharon-Lise T. Normand, and Art Sedrakyan. Metaanalysis of rate ratios with differential follow-up by treatment arm: inferring comparative effectiveness of medical devices. Statistics in Medicine, 34(21):2913–2925, September 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DelGrecoM:2015:DPM

[2000] Fabiola Del Greco M., Cosetta Minelli, Nuala A. Sheehan, and John R. Thompson. Detecting pleiotropy in Mendelian randomisation studies with summary data and a continuous outcome. Statistics in Medicine, 34(21):2926–2940, September 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wolfson:2015:NBM

[2001] Julian Wolfson, Sunayan Bandyopadhyay, Mohamed Elidrisi, Gabriela Vazquez-Benitez, David M. Vock, Donald Musgrove, Gediminas Adomavicius, Paul E. Johnson, and Patrick J. O’Connor. A naive Bayes machine learning approach to risk prediction using censored, time-to-event data. Statistics in Medicine, 34(21):2941–2957, September 20, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haines:2015:SDR

[2002] Linda M. Haines and Hassan Sadiq. Start-up designs for responseadaptive randomization procedures with sequential estimation. Statistics in Medicine, 34(21):2958–2970, September 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qu:2015:ULC

[2003] Pingping Qu, Bart Barlogie, and John Crowley. Using a latent class model to refine risk stratification in multiple myeloma. Statistics in Medicine, 34(21):2971–2980, September 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bretz:2015:SIF

[2004] Frank Bretz. Seventh international French Society of Statistics meeting on statistical methods in biopharmacy: emerging topics for statistical methodology in clinical drug development. Statistics in Medicine, 34 (22):2981–2982, September 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paoletti:2015:DFL

[2005] Xavier Paoletti, Adélaïde Doussau, Monia Ezzalfani, Elisa Rizzo, and Rodolphe Thiébaut. Dose finding with longitudinal data: simpler models, richer outcomes. Statistics in Medicine, 34(22):2983–2998, September 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colin:2015:TUF

[2006] Pierre Colin, Sandrine Micallef, Maud Delattre, Pierre Mancini, and Eric Parent. Towards using a full spectrum of early clinical trial data: a retrospective analysis to compare potential longitudinal categorical models for molecular targeted therapies in oncology. Statistics in Medicine, 34(22): 2999–3016, September 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lange:2015:ACT

[2007] Markus R. Lange and Heinz Schmidli. Analysis of clinical trials with biologics using dose-time-response models. Statistics in Medicine, 34 (22):3017–3028, September 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teramukai:2015:EBP

[2008] Satoshi Teramukai, Takashi Daimon, and Sarah Zohar. An extension of Bayesian predictive sample size selection designs for monitoring efficacy and safety. Statistics in Medicine, 34(22):3029–3039, September 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duke:2015:SBG

[2009] Susan P. Duke, Fabrice Bancken, Brenda Crowe, Mat Soukup, Taxiarchis Botsis, and Richard Forshee. Seeing is believing: good graphic design principles for medical research. Statistics in Medicine, 34(22):3040–3059, September 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Safarkhani:2015:ODL

[2010] Maryam Safarkhani and Mirjam Moerbeek. Optimal designs in longitudinal trials with varying treatment effects and discrete-time survival endpoints. Statistics in Medicine, 34(22):3060–3074, September 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Inan:2015:CAC

[2011] Gul Inan. Comments on ‘Assessing correlation of clustered mixed outcomes from a multivariate generalized linear mixed model’. Statistics in Medicine, 34(22):3075–3078, September 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Azizi:2015:BRL

[2012] Lamiae Azizi. Book review: Lyle D. Broemeling, Chapman and Hall/ CRC Biostatistics Series: Boca Raton, London, New York, 2014. No. of Pages: 454. Price: $85.46, 59.99GBP. ISBN: 978-1-4665-6497-8. Statistics in Medicine, 34(22):3079–3080, September 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OKeefe:2015:IPV

[2013] Christine M. O’Keefe and Donald B. Rubin. Individual privacy versus public good: protecting confidentiality in health research. Statistics in Medicine, 34(23):3081–3103, October 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stallard:2015:FSS

[2014] Nigel Stallard, Cornelia Ursula Kunz, Susan Todd, Nicholas Parsons, and Tim Friede. Flexible selection of a single treatment incorporating short-term endpoint information in a phase II/III clinical trial. Statistics in Medicine, 34(23):3104–3115, October 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelson:2015:MAB

[2015] Kerrie P. Nelson and Don Edwards. Measures of agreement between many raters for ordinal classifications. Statistics in Medicine, 34(23):

3116–3132, October 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:2015:PBR

[2016] Sander Greenland and Mohammad Ali Mansournia. Penalization, bias reduction, and default priors in logistic and related categorical and survival regressions. Statistics in Medicine, 34(23):3133–3143, October 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goudie:2015:RME

[2017] Robert J. B. Goudie, Roman Hovorka, Helen R. Murphy, and David Lunn. Rapid model exploration for complex hierarchical data: application to pharmacokinetics of insulin aspart. Statistics in Medicine, 34 (23):3144–3158, October 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gundersen:2015:MVF

[2018] Kenneth Gundersen, Jan Terje Kvaløy, Trygve Eftestøl, and Jo Kramer-Johansen. Modelling ventricular fibrillation coarseness during cardiopulmonary resuscitation by mixed effects stochastic differential equations. Statistics in Medicine, 34(23):3159–3169, October 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grell:2015:TDP

[2019] Kathrine Grell, Peter J. Diggle, Kirsten Frederiksen, Joachim Schüz, Elisabeth Cardis, and Per K. Andersen. A three-dimensional point process model for the spatial distribution of disease occurrence in relation to an exposure source. Statistics in Medicine, 34(23):3170–3180, October 15, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zeng:2015:BPF

[2020] Leilei Zeng, Richard J. Cook, Lan Wen, and Audrey Boruvka. Bias in progression-free survival analysis due to intermittent assessment of progression. Statistics in Medicine, 34(24):3181–3193, October 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hirakawa:2015:CSA

[2021] Akihiro Hirakawa, Nolan A. Wages, Hiroyuki Sato, and Shigeyuki Matsui. A comparative study of adaptive dose-finding designs for phase I oncology trials of combination therapies. Statistics in Medicine, 34 (24):3194–3213, October 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2015:VRB

[2022] Joseph J. Lee and Donald B. Rubin. Valid randomization-based p-values for partially post hoc subgroup analyses. Statistics in Medicine, 34(24): 3214–3222, October 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2015:MDH

[2023] Ziyue Liu and Wensheng Guo. Modeling diurnal hormone profiles by hierarchical state space models. Statistics in Medicine, 34(24):3223–3234, October 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2015:PPD

[2024] Wan Tang, Naiji Lu, Tian Chen, Wenjuan Wang, Douglas David Gunzler, Yu Han, and Xin M. Tu. On performance of parametric and distribution-free models for zero-inflated and over-dispersed count responses. Statistics in Medicine, 34(24):3235–3245, October 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gunzler:2015:TSE

[2025] Douglas D. Gunzler and Nathan Morris. A tutorial on structural equation modeling for analysis of overlapping symptoms in co-occurring conditions using MPlus. Statistics in Medicine, 34(24):3246–3280, October 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longford:2015:CTS

[2026] Nicholas T. Longford. Classification in two-stage screening. Statistics in Medicine, 34(25):3281–3297, November 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morris:2015:CFP

[2027] Tim P. Morris, Ian R. White, James R. Carpenter, Simon J. Stanworth, and Patrick Royston. Combining fractional polynomial model building with multiple imputation. Statistics in Medicine, 34(25):3298–3317, November 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perez-Jaume:2015:NPA

[2028] Sara Perez-Jaume and Josep L. Carrasco. A non-parametric approach to estimate the total deviation index for non-normal data. Statistics in Medicine, 34(25):3318–3335, November 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2015:DPC

[2029] Xiang Zhang and William H. Woodall. Dynamic probability control limits for risk-adjusted Bernoulli CUSUM charts. Statistics in Medicine, 34(25):3336–3348, November 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miller:2015:EIP

[2030] Ezer Miller, Amit Huppert, Ilya Novikov, Alon Warburg, Asrat Hailu, Ibrahim Abbasi, and Laurence S. Freedman. Estimation of infection prevalence and sensitivity in a stratified two-stage sampling design employing highly specific diagnostic tests when there is no gold standard. Statistics in Medicine, 34(25):3349–3361, November 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bansal:2015:BMD

[2031] Naveen K. Bansal, Hongmei Jiang, and Prachi Pradeep. A Bayesian methodology for detecting targeted genes under two related experiments. Statistics in Medicine, 34(25):3362–3375, November 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Remontet:2015:FOP

[2032] Laurent Remontet, Nadine Bossard, Jean Iwaz, Jacques Estève, and Aurelien Belot. Framework and optimisation procedure for flexible parametric survival models. Statistics in Medicine, 34(25):3376–3377, November 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crowther:2015:RLE

[2033] Michael J. Crowther and Paul C. Lambert. Reply to letter to the Editor by remontet et al. Statistics in Medicine, 34(25):3378–3380, November 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gutman:2015:ECE

[2034] Roee Gutman and Donald B. Rubin. Estimation of causal effects of binary treatments in unconfounded studies. Statistics in Medicine, 34 (26):3381–3398, November 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siddique:2015:MIH

[2035] Juned Siddique, Jerome P. Reiter, Ahnalee Brincks, Robert D. Gibbons, Catherine M. Crespi, and C. Hendricks Brown. Multiple imputation for harmonizing longitudinal non-commensurate measures in individual participant data meta-analysis. Statistics in Medicine, 34(26):3399–3414, November 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mandel:2015:AMC

[2036] Micha Mandel. Analyzing multiple cross-sectional samples with application to hospitalization time after surgeries. Statistics in Medicine, 34 (26):3415–3423, November 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2015:OMS

[2037] Xuelin Huang, Sangbum Choi, Lu Wang, and Peter F. Thall. Optimization of multi-stage dynamic treatment regimes utilizing accumulated data. Statistics in Medicine, 34(26):3424–3443, November 20, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ng:2015:TWC

[2038] S. K. Ng. A two-way clustering framework to identify disparities in multimorbidity patterns of mental and physical health conditions among Australians. Statistics in Medicine, 34(26):3444–3460, November 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2015:MCC

[2039] Sue-Jane Wang, Frank Bretz, Alex Dmitrienko, Jason Hsu, H. M. James Hung, Gary Koch, Willi Maurer, Walt Offen, and Robert O’Neill. Multiplicity in confirmatory clinical trials: a case study with discussion from a JSM panel. Statistics in Medicine, 34(26):3461–3480, November 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Machekano:2015:ACB

[2040] Rhoderick Machekano, Taryn Young, Simbarashe Rusakaniko, Patrick Musonda, Ben Sartorius, Jim Todd, Greg Fegan, Lehana Thabane, and Usuf Chikte. The Africa Center for Biostatistical Excellence: a proposal for enhancing biostatistics capacity for sub-Saharan Africa. Statistics in Medicine, 34(27):3481–3489, November 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moerbeek:2015:OTA

[2041] Mirjam Moerbeek and Weng-Kee Wong. Optimal treatment allocation for placebo-treatment comparisons in trials with discrete-time survival endpoints. Statistics in Medicine, 34(27):3490–3502, November 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Janes:2015:DSE

[2042] Holly Janes, Marshall D. Brown, and Margaret S. Pepe. Designing a study to evaluate the benefit of a biomarker for selecting patient treatment. Statistics in Medicine, 34(27):3503–3515, November 30, 2015. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2015:TTB

[2043] Yicong Liu, Wenyu Jiang, and Bingshu E. Chen. Testing for treatmentbiomarker interaction based on local partial-likelihood. Statistics in Medicine, 34(27):3516–3530, November 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnson:2015:RCA

[2044] Jacqueline L. Johnson, Sarah M. Kreidler, Diane J. Catellier, David M. Murray, Keith E. Muller, and Deborah H. Glueck. Recommendations for choosing an analysis method that controls Type I error for unbalanced cluster sample designs with Gaussian outcomes. Statistics in Medicine, 34(27):3531–3545, November 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baayen:2015:CBN

[2045] C. Baayen and P. Hougaard. Confidence bounds for nonlinear doseresponse relationships. Statistics in Medicine, 34(27):3546–3562, November 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2015:TDO

[2046] Yingsi Yang and Man Yu Wong. Tests for detecting overdispersion in models with measurement error in covariates. Statistics in Medicine, 34 (27):3563–3576, November 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2015:ATM

[2047] Junsheng Ma, Wenyaw Chan, Chu-Lin Tsai, Momiao Xiong, and Barbara C. Tilley. Analysis of transtheoretical model of health behavioral changes in a nutrition intervention study — a continuous time Markov chain model with Bayesian approach. Statistics in Medicine, 34(27): 3577–3589, November 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:2015:SMM

[2048] Laurence S. Freedman, Douglas Midthune, Kevin W. Dodd, Raymond J. Carroll, and Victor Kipnis. A statistical model for measurement error that incorporates variation over time in the target measure, with application to nutritional epidemiology. Statistics in Medicine, 34(27):3590– 3605, November 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2015:GRF

[2049] Dewei Wang, Christopher S. McMahan, and Colin M. Gallagher. A general regression framework for group testing data, which incorporates pool dilution effects. Statistics in Medicine, 35(10):3606–3621, November 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lyles:2015:NSB

[2050] Robert H. Lyles, Lawrence L. Kupper, Huiman X. Barnhart, and Sandra L. Martin. Numeric score-based conditional and overall change-in-status indices for ordered categorical data. Statistics in Medicine, 35 (10):3622–3636, November 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2015:ISD

[2051] Qixuan Chen, Andrew Gelman, Melissa Tracy, Fran H. Norris, and Sandro Galea. Incorporating the sampling design in weighting adjustments for panel attrition. Statistics in Medicine, 34(28):3637–3647, December 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rebora:2015:UMT

[2052] Paola Rebora, Stefania Galimberti, and Maria Grazia Valsecchi. Using multiple timescale models for the evaluation of a time-dependent treatment. Statistics in Medicine, 34(28):3648–3660, December 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2015:MTB

[2053] Peter C. Austin and Elizabeth A. Stuart. Moving towards best practice when using inverse probability of treatment weighting (IPTW) using the propensity score to estimate causal treatment effects in observational studies. Statistics in Medicine, 34(28):3661–3679, December 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Uno:2015:VTE

[2054] Hajime Uno, Lu Tian, Brian Claggett, and L. J. Wei. A versatile test for equality of two survival functions based on weighted differences of Kaplan–Meier curves. Statistics in Medicine, 34(28):3680–3695, December 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clemencon:2015:CIS

[2055] Stéphan Clémençon, Anthony Cousien, Miraine Dávila Felipe, and Viet Chi Tran. On computer-intensive simulation and estimation methods for rare-event analysis in epidemic models. Statistics in Medicine, 34 (28):3696–3713, December 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Magnusdottir:2015:SEP

[2056] Bergrun T. Magnusdottir and Hans Nyquist. Simultaneous estimation of parameters in the bivariate Emax model. Statistics in Medicine, 34(28): 3714–3723, December 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ibrahim:2015:PPT

[2057] Joseph G. Ibrahim, Ming-Hui Chen, Yeongjin Gwon, and Fang Chen. The power prior: theory and applications. Statistics in Medicine, 34 (28):3724–3749, December 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wirth:2015:CLA

[2058] Kathleen E. Wirth, Denis Agniel, Christopher D. Barr, Matthew D. Austin, and Victor DeGruttola. A composite likelihood approach for estimating HIV prevalence in the presence of spatial variation. Statistics in Medicine, 34(28):3750–3759, December 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Antognini:2015:EOC

[2059] Alessandro Baldi Antognini, William F. Rosenberger, Yang Wang, and Maroussa Zagoraiou. Exact optimum coin bias in Efron’s randomization procedure. Statistics in Medicine, 34(28):3760–3768, December 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reed:2015:GGW

[2060] Eric Reed, Sara Nunez, David Kulp, Jing Qian, Muredach P. Reilly, and Andrea S. Foulkes. A guide to genome-wide association analysis and post-analytic interrogation. Statistics in Medicine, 34(28):3769–3792, December 10, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jennison:2015:ASS

[2061] Christopher Jennison and Bruce W. Turnbull. Adaptive sample size modification in clinical trials: start small then ask for more? Statistics in Medicine, 34(29):3793–3810, December 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2015:CLN

[2062] Li Yan, Lili Tian, and Song Liu. Combining large number of weak biomarkers based on AUC. Statistics in Medicine, 34(29):3811–3830, December 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zapf:2015:NMA

[2063] Antonia Zapf, Annika Hoyer, Katharina Kramer, and Oliver Kuss. Nonparametric meta-analysis for diagnostic accuracy studies. Statistics in Medicine, 34(29):3831–3841, December 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nikoloulopoulos:2015:MEM

[2064] Aristidis K. Nikoloulopoulos. A mixed effect model for bivariate metaanalysis of diagnostic test accuracy studies using a copula representation of the random effects distribution. Statistics in Medicine, 34(29):3842– 3865, December 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Strohmaier:2015:DPA

[2065] Susanne Strohmaier, Kjetil Røysland, Rune Hoff, Ørnulf Borgan, Terje R. Pedersen, and Odd O. Aalen. Dynamic path analysis — a useful tool to investigate mediation processes in clinical survival trials. Statistics in Medicine, 34(29):3866–3887, December 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2015:TTI

[2066] Su Chen and Thomas M. Braun. Tests for time-invariant correlation of longitudinally measured biomarkers. Statistics in Medicine, 34(29):3888– 3900, December 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gott:2015:EPL

[2067] Aimee N. Gott, Idris A. Eckley, and John A. D. Aston. Estimating the population local wavelet spectrum with application to nonstationary functional magnetic resonance imaging time series. Statistics in Medicine, 34(29):3901–3915, December 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Withanage:2015:JEM

[2068] Niroshan Withanage, Alexander R. de Leon, and Christopher J. Rudnisky. Joint estimation of multiple disease-specific sensitivities and specificities via crossed random effects models for correlated reader-based diagnostic data: application of data cloning. Statistics in Medicine, 34 (29):3916–3928, December 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lesperance:2015:JMI

[2069] Mary Louise Lesperance, Veronica Sabelnykova, Farouk Salim Nathoo, Francis Lau, and Michael G. Downing. A joint model for intervalcensored functional decline trajectories under informative observation. Statistics in Medicine, 34(29):3929–3948, December 20, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2015:OFM

[2070] Peter C. Austin and Elizabeth A. Stuart. Optimal full matching for survival outcomes: a method that merits more widespread use. Statistics in Medicine, 34(30):3949–3967, December 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qiu:2015:GFS

[2071] Zhiying Qiu, Wenge Guo, and Gavin Lynch. On generalized fixed sequence procedures for controlling the FWER. Statistics in Medicine, 34 (30):3968–3983, December 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brownstein:2015:PEC

[2072] Naomi C. Brownstein, Jianwen Cai, Gary D. Slade, and Eric Bair. Parameter estimation in Cox models with missing failure indicators and the OPPERA study. Statistics in Medicine, 34(30):3984–3996, December 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Branscum:2015:FRM

[2073] Adam J. Branscum, Wesley O. Johnson, Timothy E. Hanson, and Andre T. Baron. Flexible regression models for ROC and risk analysis, with or without a gold standard. Statistics in Medicine, 34(30):3997–4015, December 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2015:PRS

[2074] Cen Wu, Xingjie Shi, Yuehua Cui, and Shuangge Ma. A penalized robust semiparametric approach for gene-environment interactions. Statistics in Medicine, 34(30):4016–4030, December 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuznetsova:2015:BTR

[2075] Olga M. Kuznetsova. Brick tunnel randomization and the momentum of the probability mass. Statistics in Medicine, 34(30):4031–4056, December 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yee:2015:NIT

[2076] Laura M. Yee and Kwun Chuen Gary Chan. Nonparametric inference for time-dependent incremental cost-effectiveness ratios. Statistics in Medicine, 34(30):4057–4069, December 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pimentel:2015:VRM

[2077] Samuel D. Pimentel, Frank Yoon, and Luke Keele. Variable-ratio matching with fine balance in a study of the peer health exchange. Statistics in Medicine, 34(30):4070–4082, December 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rua:2015:CPD

[2078] Sandra M. Hurtado Rúa, Madhu Mazumdar, and Robert L. Strawderman. The choice of prior distribution for a covariance matrix in multivariate meta-analysis: a simulation study. Statistics in Medicine, 34(30): 4083–4104, December 30, 2015. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITa

[2079] Anonymous. Issue information — title page. Statistics in Medicine, 35 (1):1, January 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIa

[2080] Anonymous. Issue information — info page. Statistics in Medicine, 35 (1):2, January 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITb

[2081] Anonymous. Issue information — TOC. Statistics in Medicine, 35(1): 3, January 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huque:2016:VHP

[2082] Mohammad F. Huque. Validity of the Hochberg procedure revisited for clinical trial applications. Statistics in Medicine, 35(1):5–20, January 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2016:MAS

[2083] Yong Chen, Chuan Hong, Yang Ning, and Xiao Su. Meta-analysis of studies with bivariate binary outcomes: a marginal beta-binomial model approach. Statistics in Medicine, 35(1):21–40, January 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petersen:2016:SPE

[2084] Jørgen Holm Petersen. Semi-parametric estimation of random effects in a logistic regression model using conditional inference. Statistics in Medicine, 35(1):41–52, January 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2016:BAE

[2085] Jincheng Zhou, Haitao Chu, Michael G. Hudgens, and M. Elizabeth Halloran. A Bayesian approach to estimating causal vaccine effects on binary post-infection outcomes. Statistics in Medicine, 35(1):53–64, January 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ning:2016:JLC

[2086] Jing Ning, Mohammad H. Rahbar, Sangbum Choi, Chuan Hong, Jin Piao, Deborah J. del Junco, Erin E. Fox, Elaheh Rahbar, and John B. Holcomb. A joint latent class analysis for adjusting survival bias with application to a trauma transfusion study. Statistics in Medicine, 35(1): 65–77, January 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Batterton:2016:NFI

[2087] Katherine A. Batterton and Christine M. Schubert. A nonparametric fiducial interval for the Youden index in multi-state diagnostic settings. Statistics in Medicine, 35(1):78–96, January 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paddock:2016:BRS

[2088] Susan M. Paddock, Thomas J. Leininger, and Sarah B. Hunter. Bayesian restricted spatial regression for examining session features and patient outcomes in open-enrollment group therapy studies. Statistics in Medicine, 35(1):97–114, January 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gutman:2016:EAF

[2089] R. Gutman, C. J. Sammartino, T. C. Green, and B. T. Montague. Error adjustments for file linking methods using encrypted unique client identifier (eUCI) with application to recently released prisoners who are HIV+. Statistics in Medicine, 35(1):115–129, January 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2016:ICS

[2090] H. Jiang, P. E. Brown, and S. D. Walter. Inference on cancer screening exam accuracy using population-level administrative data. Statistics in Medicine, 35(1):130–146, January 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2016:LCI

[2091] Stuart G. Baker, Barnett S. Kramer, and Karen S. Lindeman. Latent class instrumental variables: a clinical and biostatistical perspective. Statistics in Medicine, 35(1):147–160, January 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITc

[2092] Anonymous. Issue information — title page. Statistics in Medicine, 35 (2):161, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIb

[2093] Anonymous. Issue information — info page. Statistics in Medicine, 35 (2):162, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITd

[2094] Anonymous. Issue information — TOC. Statistics in Medicine, 35(2): 163, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2016:VMF

[2095] Cathy Yuen Yi Lee and Matt P. Wand. Variational methods for fitting complex Bayesian mixed effects models to health data. Statistics in Medicine, 35(2):165–188, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2016:CTT

[2096] Tong-Yu Lu, Wai-Yin Poon, and Siu Hung Cheung. Comparison of two treatments with skewed ordinal responses. Statistics in Medicine, 35(2): 189–201, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsu:2016:OLC

[2097] Man-Jen Hsu and Yi-Hau Chen. Optimal linear combination of biomarkers for multi-category diagnosis. Statistics in Medicine, 35(2):202–213, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collins:2016:SSC

[2098] Gary S. Collins, Emmanuel O. Ogundimu, and Douglas G. Altman. Sample size considerations for the external validation of a multivariable prognostic model: a resampling study. Statistics in Medicine, 35(2):214–226, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Varewyck:2016:PIC

[2099] Machteld Varewyck, Stijn Vansteelandt, Marie Eriksson, and Els Goetghebeur. On the practice of ignoring center-patient interactions in evaluating hospital performance. Statistics in Medicine, 35(2):227–238, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giraudeau:2016:SSC

[2100] Bruno Giraudeau, Julian P. T. Higgins, Elsa Tavernier, and Ludovic Trinquart. Sample size calculation for meta-epidemiological studies. Statistics in Medicine, 35(2):239–250, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Christian:2016:HLI

[2101] Nicholas J. Christian, Il Do Ha, and Jong-Hyeon Jeong. Hierarchical likelihood inference on clustered competing risks data. Statistics in Medicine, 35(2):251–267, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2016:PAM

[2102] Chyong-Mei Chen, Pao-Sheng Shen, and Ya-Wen Chuang. The partly Aalen’s model for recurrent event data with a dependent terminal event. Statistics in Medicine, 35(2):268–281, January 30, 2016. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2016:IEE

[2103] Soyoung Kim, Jianwen Cai, and David Couper. Improving the efficiency of estimation in the additive hazards model for stratified case-cohort design with multiple diseases. Statistics in Medicine, 35(2):282–293, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sutradhar:2016:TOW

[2104] Rinku Sutradhar, Nancy N. Baxter, and Peter C. Austin. Terminating observation within matched pairs of subjects in a matched cohort analysis: a Monte Carlo simulation study. Statistics in Medicine, 35(2): 294–304, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kirchner:2016:UBO

[2105] Marietta Kirchner, Meinhard Kieser, Heiko Götte, and Armin Schüler. Utility-based optimization of phase II/III programs. Statistics in Medicine, 35(2):305–316, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:CUC

[2106] Anonymous. Comments on ‘Use of composite endpoints in clinical trials’ by Abdul J. Sankoh, Haihong Li and Ralph B. D’Agostino, Sr. Statistics in Medicine, 35(2):317–318, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sankoh:2016:UCE

[2107] Abdul J. Sankoh, Haihong Li, and Ralph B. D’Agostino, Sr. Use of composite endpoints in clinical trials. Statistics in Medicine, 35(2):319– 320, January 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITe

[2108] Anonymous. Issue information — title page. Statistics in Medicine, 35 (3):321, February 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIc

[2109] Anonymous. Issue information — info page. Statistics in Medicine, 35 (3):322, February 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITf

[2110] Anonymous. Issue information — TOC. Statistics in Medicine, 35(3): 323, February 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bauer:2016:TFY

[2111] Peter Bauer, Frank Bretz, Vladimir Dragalin, Franz König, and Gernot Wassmer. Twenty-five years of confirmatory adaptive designs: opportunities and pitfalls. Statistics in Medicine, 35(3):325–347, February 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hung:2016:CTF

[2112] H. M. J. Hung, Sue-Jane Wang, and John Lawrence. Comments on ‘Twenty-five years of confirmatory adaptive designs: opportunities and pitfalls’. Statistics in Medicine, 35(3):348–349, February 10, 2016. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehta:2016:ORE

[2113] Cyrus Mehta and Lingyun Liu. An objective re-evaluation of adaptive sample size re-estimation: commentary on ‘twenty-five years of confirmatory adaptive designs’. Statistics in Medicine, 35(3):350–358, February 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vollmar:2016:CTF

[2114] Joachim Vollmar. Comments on ‘Twenty-five years of confirmatory adaptive designs: opportunities and pitfalls’. Statistics in Medicine, 35(3): 359–361, February 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maurer:2016:CTF

[2115] Willi Maurer. Comments on ‘Twenty-five Years of Confirmatory Adaptive Designs: Opportunities and Pitfalls’. Statistics in Medicine, 35(3): 362–363, February 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bauer:2016:ARC

[2116] Peter Bauer, Frank Bretz, Vladimir Dragalin, Franz König, and Gernot Wassmer. Authors’ response to comments. Statistics in Medicine, 35 (3):364–367, February 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vandeWiel:2016:BPU

[2117] Mark A. van de Wiel, Tonje G. Lien, Wina Verlaat, Wessel N. van Wieringen, and Saskia M. Wilting. Better prediction by use of co-data: adaptive group-regularized ridge regression. Statistics in Medicine, 35(3):368–381, February 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proust-Lima:2016:JMR

[2118] Cécile Proust-Lima, Jean-François Dartigues, and Hélène Jacqmin-Gadda. Joint modeling of repeated multivariate cognitive measures and competing risks of dementia and death: a latent process and latent class approach. Statistics in Medicine, 35(3):382–398, February 10, 2016. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kwon:2016:BDR

[2119] Deukwoo Kwon, F. Owen Hoffman, Brian E. Moroz, and Steven L. Simon. Bayesian dose-response analysis for epidemiological studies with complex uncertainty in dose estimation. Statistics in Medicine, 35(3): 399–423, February 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Serfling:2016:PMN

[2120] Robert Serfling and Gerald Ogola. Probability modeling of the number of positive cores in a prostate cancer biopsy session, with applications. Statistics in Medicine, 35(3):424–454, February 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2016:PPR

[2121] X. Luo, S. Gee, V. Sohal, and D. Small. A point-process response model for spike trains from single neurons in neural circuits under optogenetic stimulation. Statistics in Medicine, 35(3):455–474, February 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riviere:2016:CCS

[2122] Marie-Karelle Riviere and Sarah Zohar. Comments on ‘A comparative study of adaptive dose-finding designs for phase I oncology trials of combination therapies’. Statistics in Medicine, 35(3):475–478, February 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hirakawa:2016:AR

[2123] Akihiro Hirakawa and Hiroyuki Sato. Authors’ reply. Statistics in Medicine, 35(3):479–480, February 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITg

[2124] Anonymous. Issue information — title page. Statistics in Medicine, 35 (4):481, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIId

[2125] Anonymous. Issue information — info page. Statistics in Medicine, 35 (4):482, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITh

[2126] Anonymous. Issue information — TOC. Statistics in Medicine, 35(4): 483, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoaglin:2016:MAC

[2127] David C. Hoaglin. Misunderstandings about Q and ‘Cochran’s Q test’ in meta-analysis. Statistics in Medicine, 35(4):485–495, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hedges:2016:CMA

[2128] Larry V. Hedges. Comment on ‘Misunderstandings about Q and “Cochran’s Q Test” in meta analysis’. Statistics in Medicine, 35(4):496– 497, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shuster:2016:DMA

[2129] Jonathan J. Shuster. Discussion of ‘Misunderstandings about Q and “Cochran’s Q Test” in meta-analysis’. Statistics in Medicine, 35(4):498– 500, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kulinskaya:2016:CMA

[2130] Elena Kulinskaya and Michael B. Dollinger. Commentary on ‘Misunderstandings about Q and “Cochran’s Q test” in meta analysis’. Statistics in Medicine, 35(4):501–502, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoaglin:2016:RD

[2131] David C. Hoaglin. Reply to discussants. Statistics in Medicine, 35(4): 503–504, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mutze:2016:DAT

[2132] Tobias Mütze, Axel Munk, and Tim Friede. Design and analysis of threearm trials with negative binomially distributed endpoints. Statistics in Medicine, 35(4):505–521, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jaki:2016:DMA

[2133] Thomas Jaki and Lisa V. Hampson. Designing multi-arm multi-stage clinical trials using a risk-benefit criterion for treatment selection. Statistics in Medicine, 35(4):522–533, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Linden:2016:ECE

[2134] Ariel Linden, S. Derya Uysal, Andrew Ryan, and John L. Adams. Estimating causal effects for multivalued treatments: a comparison of approaches. Statistics in Medicine, 35(4):534–552, February 20, 2016. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wynant:2016:FES

[2135] Willy Wynant and Michal Abrahamowicz. Flexible estimation of survival curves conditional on non-linear and time-dependent predictor effects. Statistics in Medicine, 35(4):553–565, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosa:2016:GDS

[2136] Patricio S. La Rosa, Terrence L. Brooks, Elena Deych, Berkley Shands, Fred Prior, Linda J. Larson-Prior, and William D. Shannon. Gibbs distribution for statistical analysis of graphical data with a sample application to fcMRI brain images. Statistics in Medicine, 35(4):566–580, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nasserinejad:2016:PHB

[2137] Kazem Nasserinejad, Joost van Rosmalen, Wim de Kort, Dimitris Rizopoulos, and Emmanuel Lesaffre. Prediction of hemoglobin in blood donors using a latent class mixed-effects transition model. Statistics in Medicine, 35(4):581–594, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barrado:2016:DDT

[2138] Leandro García Barrado, Els Coart, Tomasz Burzykowski, and for the Alzheimer’s Disease Neuroimaging Initiative. Development of a diagnostic test based on multiple continuous biomarkers with an imperfect reference test. Statistics in Medicine, 35(4):595–608, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rotolo:2016:INF

[2139] Federico Rotolo, Virginie Rondeau, and Catherine Legrand. Incorporation of nested frailties into semiparametric multi-state models. Statistics in Medicine, 35(4):609–621, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pelle:2016:LLM

[2140] E. Pelle, D. J. Hessen, and P. G. M. van der Heijden. A log-linear multidimensional Rasch model for capture-recapture. Statistics in Medicine, 35(4):622–634, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reiser:2016:EST

[2141] Benjamin Reiser and Christos T. Nakas. On ‘efficient statistical tests to compare Youden index: accounting for contingency correlation’. Statistics in Medicine, 35(4):635–636, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2016:RCE

[2142] Fangyao Chen, Yuqiang Xue, Ming T. Tan, and Pingyan Chen. Response to comment on ‘efficient statistical tests to compare Youden index: accounting for contingency correlation’. Statistics in Medicine, 35 (4):637–640, February 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITi

[2143] Anonymous. Issue information — title page. Statistics in Medicine, 35 (5):641, February 28, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIe

[2144] Anonymous. Issue information — info page. Statistics in Medicine, 35 (5):642, February 28, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITj

[2145] Anonymous. Issue information — TOC. Statistics in Medicine, 35(5): 643, February 28, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farewell:2016:MGBa

[2146] Vern Farewell and Tony Johnson. Major Greenwood (1880–1949): a biographical and bibliographical study. Statistics in Medicine, 35(5): 645–670, February 28, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2016:ELT

[2147] Hua Guo, Joshua Chen, and Hui Quan. Evaluation of local treatment effect by borrowing information from similar countries in multi-regional clinical trials. Statistics in Medicine, 35(5):671–684, February 28, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berger:2016:CMR

[2148] Vance W. Berger, Klejda Bejleri, and Rebecca Agnor. Comparing MTI randomization procedures to blocked randomization. Statistics in Medicine, 35(5):685–694, February 28, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2016:BAA

[2149] Samiran Ghosh, Santu Ghosh, and Ram C. Tiwari. Bayesian approach for assessing non-inferiority in a three-arm trial with pre-specified margin. Statistics in Medicine, 35(5):695–708, February 28, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nattino:2016:NTG

[2150] Giovanni Nattino, Stefano Finazzi, and Guido Bertolini. A new test and graphical tool to assess the goodness of fit of logistic regression models. Statistics in Medicine, 35(5):709–720, February 28, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Porter:2016:SEM

[2151] Aaron T. Porter and Jacob J. Oleson. A spatial epidemic model for disease spread over a heterogeneous spatial support. Statistics in Medicine, 35(5):721–733, February 28, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmid:2016:STM

[2152] Matthias Schmid, Helmut Küchenhoff, Achim Hoerauf, and Gerhard Tutz. A survival tree method for the analysis of discrete event times in clinical and epidemiological studies. Statistics in Medicine, 35(5):734– 751, February 28, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Porcher:2016:PPS

[2153] Raphaël Porcher, Clémence Leyrat, Gabriel Baron, Bruno Giraudeau, and Isabelle Boutron. Performance of principal scores to estimate the marginal compliers causal effect of an intervention. Statistics in Medicine, 35(5):752–767, February 28, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Landwehr:2016:CSP

[2154] Sandra Landwehr and Ralph Brinks. A comparative study of prevalencebased incidence estimation techniques with application to dementia data in Germany. Statistics in Medicine, 35(5):768–781, February 28, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2016:SMS

[2155] Molin Wang, Donna Spiegelman, Aya Kuchiba, Paul Lochhead, Sehee Kim, Andrew T. Chan, Elizabeth M. Poole, Rulla Tamimi, Shelley S.

Tworoger, Edward Giovannucci, Bernard Rosner, and Shuji Ogino. Statistical methods for studying disease subtype heterogeneity. Statistics in Medicine, 35(5):782–800, February 28, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITk

[2156] Anonymous. Issue information — title page. Statistics in Medicine, 35 (6):801, March 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIf

[2157] Anonymous. Issue information — info page. Statistics in Medicine, 35 (6):802, March 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITl

[2158] Anonymous. Issue information — TOC. Statistics in Medicine, 35(6): 803, March 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pullenayegum:2016:KTB

[2159] Eleanor M. Pullenayegum, Robert W. Platt, Melanie Barwick, Brian M. Feldman, Martin Offringa, and Lehana Thabane. Knowledge translation in biostatistics: a survey of current practices, preferences, and barriers to the dissemination and uptake of new statistical methods. Statistics in Medicine, 35(6):805–818, March 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2016:EDL

[2160] Dan Jackson, Martin Law, Jessica K. Barrett, Rebecca Turner, Julian P. T. Higgins, Georgia Salanti, and Ian R. White. Extending Der-Simonian and Laird’s methodology to perform network meta-analyses with random inconsistency effects. Statistics in Medicine, 35(6):819–839, March 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ogbagaber:2016:DSR

[2161] Semhar B. Ogbagaber, Jordan Karp, and Abdus S. Wahed. Design of sequentially randomized trials for testing adaptive treatment strategies. Statistics in Medicine, 35(6):840–858, March 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jafarzadeh:2016:BMI

[2162] S. Reza Jafarzadeh, Wesley O. Johnson, and Ian A. Gardner. Bayesian modeling and inference for diagnostic accuracy and probability of disease based on multiple diagnostic biomarkers with and without a perfect reference standard. Statistics in Medicine, 35(6):859–876, March 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kejzar:2016:NBM

[2163] Natasa Kejzar, Delphine Maucort-Boulch, and Janez Stare. A note on bias of measures of explained variation for survival data. Statistics in Medicine, 35(6):877–882, March 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2016:FMC

[2164] Jinsong Chen, Lei Liu, Ya-Chen T. Shih, Daowen Zhang, and Thomas A. Severini. A flexible model for correlated medical costs, with application to medical expenditure panel survey data. Statistics in Medicine, 35(6): 883–894, March 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hossain:2016:CEP

[2165] Abu Hossain, Robert Rigby, Mikis Stasinopoulos, and Marco Enea. Centile estimation for a proportion response variable. Statistics in Medicine, 35(6):895–904, March 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lakhal-Chaieb:2016:RVA

[2166] Lajmi Lakhal-Chaieb, Karim Oualkacha, Brent J. Richards, and Celia M. T. Greenwood. A rare variant association test in family-based designs and non-normal quantitative traits. Statistics in Medicine, 35(6):905– 921, March 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Graf:2016:PML

[2167] Alexandra Christine Graf, Georg Gutjahr, and Werner Brannath. Precision of maximum likelihood estimation in adaptive designs. Statistics in Medicine, 35(6):922–941, March 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sahoo:2016:GTP

[2168] Shyamsundar Sahoo and Debasis Sengupta. On graphical tests for proportionality of hazards in two samples. Statistics in Medicine, 35(6):

942–956, March 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2016:ECIa

[2169] Xinran Li and Peng Ding. Exact confidence intervals for the average causal effect on a binary outcome. Statistics in Medicine, 35(6):957–960, March 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITm

[2170] Anonymous. Issue information — title page. Statistics in Medicine, 35 (7):961, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIg

[2171] Anonymous. Issue information — info page. Statistics in Medicine, 35 (7):962, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITn

[2172] Anonymous. Issue information — TOC. Statistics in Medicine, 35(7): 963, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heinze:2016:P

[2173] Georg Heinze, Stefan Michiels, and Martin Posch. Preface. Statistics in Medicine, 35(7):965, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2016:MVV

[2174] Stephen Senn. Mastering variation: variance components and personalised medicine. Statistics in Medicine, 35(7):966–977, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nikolakopoulou:2016:PFS

[2175] Adriani Nikolakopoulou, Dimitris Mavridis, and Georgia Salanti. Planning future studies based on the precision of network meta-analysis results. Statistics in Medicine, 35(7):978–1000, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abrahamowicz:2016:MCA

[2176] Michal Abrahamowicz, Lise M. Bjerre, Marie-Eve Beauchamp, Jacques LeLorier, and Rebecca Burne. The missing cause approach to unmeasured confounding in pharmacoepidemiology. Statistics in Medicine, 35 (7):1001–1016, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peterson:2016:JBV

[2177] Christine B. Peterson, Francesco C. Stingo, and Marina Vannucci. Joint Bayesian variable and graph selection for regression models with networkstructured predictors. Statistics in Medicine, 35(7):1017–1031, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernasconi:2016:SPT

[2178] Davide Paolo Bernasconi, Paola Rebora, Simona Iacobelli, Maria Grazia Valsecchi, and Laura Antolini. Survival probabilities with timedependent treatment indicator: quantities and non-parametric estimators. Statistics in Medicine, 35(7):1032–1048, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schurmann:2016:DST

[2179] Christoph Schürmann and Wiebke Sieben. Differences in surrogate threshold effect estimates between original and simplified correlationbased validation approaches. Statistics in Medicine, 35(7):1049–1062, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bujkiewicz:2016:BMA

[2180] Sylwia Bujkiewicz, John R. Thompson, Richard D. Riley, and Keith R. Abrams. Bayesian meta-analytical methods to incorporate multiple surrogate endpoints in drug development process. Statistics in Medicine, 35(7):1063–1089, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rodriguez-Alvarez:2016:NET

[2181] María Xosé Rodríguez-Álvarez, Luís Meira-Machado, Emad Abu-Assi, and Sergio Raposeiras-Roubín. Nonparametric estimation of timedependent ROC curves conditional on a continuous covariate. Statistics in Medicine, 35(7):1090–1102, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Borgne:2016:CPD

[2182] Florent Le Borgne, Bruno Giraudeau, Anne Héléne Querard, Magali Giral, and Yohann Foucher. Comparisons of the performance of different statistical tests for time-to-event analysis with confounding factors: practical illustrations in kidney transplantation. Statistics in Medicine, 35 (7):1103–1116, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gron:2016:MPR

[2183] Randi Grøn, Thomas A. Gerds, and Per K. Andersen. Misspecified Poisson regression models for large-scale registry data: inference for ‘large n and small p’. Statistics in Medicine, 35(7):1117–1129, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stupnik:2016:ADR

[2184] Tomaz Stupnik and Maja Pohar Perme. Analyzing disease recurrence with missing at risk information. Statistics in Medicine, 35(7):1130–1143, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zoller:2016:SPV

[2185] Daniela Zöller, Irene Schmidtmann, Arndt Weinmann, Thomas A. Gerds, and Harald Binder. Stagewise pseudo-value regression for time-varying effects on the cumulative incidence. Statistics in Medicine, 35(7):1144– 1158, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pavlou:2016:REP

[2186] Menelaos Pavlou, Gareth Ambler, Shaun Seaman, Maria De Iorio, and Rumana Z. Omar. Review and evaluation of penalised regression methods for risk prediction in low-dimensional data with few events. Statistics in Medicine, 35(7):1159–1177, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grand:2016:RME

[2187] Mia Klinten Grand and Hein Putter. Regression models for expected length of stay. Statistics in Medicine, 35(7):1178–1192, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crowther:2016:JML

[2188] Michael J. Crowther, Therese M.-L Andersson, Paul C. Lambert, Keith R. Abrams, and Keith Humphreys. Joint modelling of longitudinal and survival data: incorporating delayed entry and an assessment of model misspecification. Statistics in Medicine, 35(7):1193–1209, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scolas:2016:VSF

[2189] Sylvie Scolas, Anouar El Ghouch, Catherine Legrand, and Abderrahim Oulhaj. Variable selection in a flexible parametric mixture cure model with interval-censored data. Statistics in Medicine, 35(7):1210–1225, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nevo:2016:CMC

[2190] Daniel Nevo, Micha Mandel, Eliana Ein-Mor, Ori Shen, Avraham Ben Chetrit, Etty Daniel-Spiegel, and Simcha Yagel. A comparison of methods for construction of fetal reference charts. Statistics in Medicine, 35 (7):1226–1240, March 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITo

[2191] Anonymous. Issue information — title page. Statistics in Medicine, 35 (8):1241, April 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIh

[2192] Anonymous. Issue information — info page. Statistics in Medicine, 35 (8):1242, April 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITp

[2193] Anonymous. Issue information — TOC. Statistics in Medicine, 35(8): 1243, April 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laber:2016:UPD

[2194] Eric B. Laber, Ying-Qi Zhao, Todd Regh, Marie Davidian, Anastasios Tsiatis, Joseph B. Stanford, Donglin Zeng, Rui Song, and Michael R. Kosorok. Using pilot data to size a two-arm randomized trial to find a nearly optimal personalized treatment strategy. Statistics in Medicine, 35(8):1245–1256, April 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shan:2016:OAT

[2195] Guogen Shan, Gregory E. Wilding, Alan D. Hutson, and Shawn Gerstenberger. Optimal adaptive two-stage designs for early phase II clinical trials. Statistics in Medicine, 35(8):1257–1266, April 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koopmeiners:2016:GST

[2196] Joseph S. Koopmeiners and Ziding Feng. Group sequential testing of the predictive accuracy of a continuous biomarker with unknown prevalence. Statistics in Medicine, 35(8):1267–1280, April 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

derElst:2016:ERB

[2197] Wim Van der Elst, Geert Molenberghs, and Ariel Alonso. Exploring the relationship between the causal-inference and meta-analytic paradigms for the evaluation of surrogate endpoints. Statistics in Medicine, 35(8): 1281–1298, April 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2016:PCA

[2198] Lili Yang, Menggang Yu, and Sujuan Gao. Prediction of coronary artery disease risk based on multiple longitudinal biomarkers. Statistics in Medicine, 35(8):1299–1314, April 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yin:2016:MIA

[2199] Xiaoyan Yin, Daniel Levy, Christine Willinger, Aram Adourian, and Martin G. Larson. Multiple imputation and analysis for high-dimensional incomplete proteomics data. Statistics in Medicine, 35(8):1315–1326, April 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dong:2016:AFT

[2200] Xinxin Dong, Lan Kong, and Abdus S. Wahed. Accelerated failure time model for case-cohort design with longitudinal covariates subject to measurement error and detection limits. Statistics in Medicine, 35(8): 1327–1339, April 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barrett:2016:CDR

[2201] James E. Barrett and Anthony C. C. Coolen. Covariate dimension reduction for survival data via the Gaussian process latent variable model. Statistics in Medicine, 35(8):1340–1353, April 15, 2016. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Busing:2016:TNC

[2202] F. M. T. A. Busing, B. Weaver, and S. Dubois. 2 × 2 tables: a note on Campbell’s recommendation. Statistics in Medicine, 35(8):1354–1358, April 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2016:SCR

[2203] Jingyang Zhang, Ying Zhang, Kathryn Chaloner, and Jack T. Stapleton. A sequential classification rule based on multiple quantitative tests in the absence of a gold standard. Statistics in Medicine, 35(8):1359–1372, April 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barcella:2016:VSC

[2204] William Barcella, Maria De Iorio, Gianluca Baio, and James Malone-Lee. Variable selection in covariate dependent random partition models: an application to urinary tract infection. Statistics in Medicine, 35(8): 1373–1389, April 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2016:PSS

[2205] Hae-Young Kim, John M. Williamson, and Hung-Mo Lin. Power and sample size calculations for interval-censored survival analysis. Statistics in Medicine, 35(8):1390–1400, April 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITq

[2206] Anonymous. Issue information — title page. Statistics in Medicine, 35 (9):1401, April 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIi

[2207] Anonymous. Issue information — info page. Statistics in Medicine, 35 (9):1402, April 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITr

[2208] Anonymous. Issue information — TOC. Statistics in Medicine, 35(9): 1403, April 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2016:ICE

[2209] Yong Chen, Yi Cai, Chuan Hong, and Dan Jackson. Inference for correlated effect sizes using multiple univariate meta-analyses. Statistics in Medicine, 35(9):1405–1422, April 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bowden:2016:GPP

[2210] Jack Bowden, Shaun Seaman, Xin Huang, and Ian R. White. Gaining power and precision by using model-based weights in the analysis of late stage cancer trials with substantial treatment switching. Statistics in Medicine, 35(9):1423–1440, April 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Silva:2016:CSM

[2211] Ivair R. Silva. Composite sequential Monte Carlo test for post-market vaccine safety surveillance. Statistics in Medicine, 35(9):1441–1453, April 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schiller:2016:BDC

[2212] Ian Schiller, Maarten van Smeden, Alula Hadgu, Michael Libman, Johannes B. Reitsma, and Nandini Dendukuri. Bias due to composite reference standards in diagnostic accuracy studies. Statistics in Medicine, 35(9):1454–1470, April 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

An:2016:BHM

[2213] Qian An, Jian Kang, Ruiguang Song, and H. Irene Hall. A Bayesian hierarchical model with novel prior specifications for estimating HIV testing rates. Statistics in Medicine, 35(9):1471–1487, April 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ganguli:2016:DDG

[2214] B. Ganguli, S. Sen Roy, M. Naskar, E. J. Malloy, and E. A. Eisen. Deletion diagnostics for the generalised linear mixed model with independent random effects. Statistics in Medicine, 35(9):1488–1501, April 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Danaher:2016:EIE

[2215] Michelle R. Danaher, Paul S. Albert, Aninyda Roy, and Enrique F. Schisterman. Estimation of interaction effects using pooled biospecimens in a case-control study. Statistics in Medicine, 35(9):1502–1513, April 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stirrup:2016:FBM

[2216] Oliver T. Stirrup, Abdel G. Babiker, James R. Carpenter, and Andrew J. Copas. Fractional Brownian motion and multivariate-t models for longitudinal biomedical data, with application to CD4 counts in HIV-positive patients. Statistics in Medicine, 35(9):1514–1532, April 30, 2016. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vakulenko-Lagun:2016:CEA

[2217] Bella Vakulenko-Lagun and Micha Mandel. Comparing estimation approaches for the illness-death model under left truncation and right censoring. Statistics in Medicine, 35(9):1533–1548, April 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Benoit:2016:HMM

[2218] Julia S. Benoit, Wenyaw Chan, Sheng Luo, Hung-Wen Yeh, and Rachelle Doody. A hidden Markov model approach to analyze longitudinal ternary outcomes when some observed states are possibly misclassified. Statistics in Medicine, 35(9):1549–1557, April 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Englert:2016:CHT

[2219] Stefan Englert and Meinhard Kieser. Comments on ‘Hypothesis testing for two-stage designs with over or under enrollment’. Statistics in Medicine, 35(9):1558–1559, April 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zeng:2016:RC

[2220] Donglin Zeng, Fei Gao, Joseph G. Ibrahim, Kuolung Hu, and Catherine Jia. Reply to comments. Statistics in Medicine, 35(9):1560, April 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITs

[2221] Anonymous. Issue information — title page. Statistics in Medicine, 35 (10):1561, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIj

[2222] Anonymous. Issue information — info page. Statistics in Medicine, 35 (10):1562, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITt

[2223] Anonymous. Issue information — TOC. Statistics in Medicine, 35(10): 1563, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2016:ECR

[2224] Fan Li, Yuliya Lokhnygina, David M. Murray, Patrick J. Heagerty, and Elizabeth R. DeLong. An evaluation of constrained randomization for the design and analysis of group-randomized trials. Statistics in Medicine, 35(10):1565–1579, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ding:2016:SMI

[2225] Ying Ding, Hui-Min Lin, and Jason C. Hsu. Subgroup mixable inference on treatment efficacy in mixture populations, with an application to time-to-event outcomes. Statistics in Medicine, 35(10):1580–1594, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2016:CDT

[2226] Xi Lu, Inbal Nahum-Shani, Connie Kasari, Kevin G. Lynch, David W. Oslin, William E. Pelham, Gregory Fabiano, and Daniel Almirall. Comparing dynamic treatment regimes using repeated-measures outcomes: modeling considerations in SMART studies. Statistics in Medicine, 35 (10):1595–1615, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roberts:2016:DAT

[2227] Chris Roberts, Evridiki Batistatou, and Stephen A. Roberts. Design and analysis of trials with a partially nested design and a binary outcome measure. Statistics in Medicine, 35(10):1616–1636, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parast:2016:REP

[2228] Layla Parast, Mary M. McDermott, and Lu Tian. Robust estimation of the proportion of treatment effect explained by surrogate marker information. Statistics in Medicine, 35(10):1637–1653, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Verde:2016:BES

[2229] Pablo E. Verde, Christian Ohmann, Stephan Morbach, and Andrea Icks. Bayesian evidence synthesis for exploring generalizability of treatment effects: a case study of combining randomized and non-randomized results in diabetes. Statistics in Medicine, 35(10):1654–1675, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2016:MAM

[2230] Ching-Yun Wang, Jean De Dieu Tapsoba, Catherine Duggan, Kristin L. Campbell, and Anne McTiernan. Methods to adjust for misclassification in the quantiles for the generalized linear model with measurement error in continuous exposures. Statistics in Medicine, 35(10):1676–1688, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yue:2016:BIG

[2231] Yu Ryan Yue and Xiao-Feng Wang. Bayesian inference for generalized linear mixed models with predictors subject to detection limits: an approach that leverages information from auxiliary variables. Statistics in Medicine, 35(10):1689–1705, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2016:CEGa

[2232] Ming Wang, Lan Kong, Zheng Li, and Lijun Zhang. Covariance estimators for generalized estimating equations (GEE) in longitudinal analysis with small samples. Statistics in Medicine, 35(10):1706–1721, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Preisser:2016:MZI

[2233] John S. Preisser, Kalyan Das, D. Leann Long, and Kimon Divaris. Marginalized zero-inflated negative binomial regression with application to dental caries. Statistics in Medicine, 35(10):1722–1735, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2016:BAI

[2234] Wenle Zhao. A better alternative to the inferior permuted block design is not necessarily complex. Statistics in Medicine, 35(10):1736–1738, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chiba:2016:NEC

[2235] Yasutaka Chiba. A note on exact confidence interval for causal effects on a binary outcome in randomized trials. Statistics in Medicine, 35(10): 1739–1741, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:2016:CSM

[2236] Laurence S. Freedman. Correction — a statistical model for measurement error that incorporates variation over time in the target measure, with application to nutritional epidemiology. Statistics in Medicine, 35(10): 1742–1744, May 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITu

[2237] Anonymous. Issue information — title page. Statistics in Medicine, 35 (11):1745, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIk

[2238] Anonymous. Issue information — info page. Statistics in Medicine, 35 (11):1746, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITv

[2239] Anonymous. Issue information — TOC. Statistics in Medicine, 35(11): 1747, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ioannidis:2016:EWE

[2240] John P. A. Ioannidis. Exposure-wide epidemiology: revisiting Bradford Hill. Statistics in Medicine, 35(11):1749–1762, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Judkins:2016:ROL

[2241] David R. Judkins and Kristin E. Porter. Robustness of ordinary least squares in randomized clinical trials. Statistics in Medicine, 35(11):1763– 1773, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sit:2016:DAC

[2242] Tony Sit, Mengling Liu, Michael Shnaidman, and Zhiliang Ying. Design and analysis of clinical trials in the presence of delayed treatment effect. Statistics in Medicine, 35(11):1774–1779, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prendergast:2016:MAR

[2243] Luke A. Prendergast and Robert G. Staudte. Meta-analysis of ratios of sample variances. Statistics in Medicine, 35(11):1780–1799, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pullenayegum:2016:MOA

[2244] Eleanor M. Pullenayegum. Multiple outputation for the analysis of longitudinal data subject to irregular observation. Statistics in Medicine, 35(11):1800–1818, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jaman:2016:DBC

[2245] Ajmery Jaman, Mahbub A. H. M. Latif, Wasimul Bari, and Abdus S. Wahed. A determinant-based criterion for working correlation structure selection in generalized estimating equations. Statistics in Medicine, 35 (11):1819–1833, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Estes:2016:TVE

[2246] Jason P. Estes, Danh V. Nguyen, Lorien S. Dalrymple, Yi Mu, and Damla Sentürk. Time-varying effect modeling with longitudinal data truncated by death: conditional models, interpretations, and inference. Statistics in Medicine, 35(11):1834–1847, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bauer:2016:BPS

[2247] Cici Bauer, Jon Wakefield, Håvard Rue, Steve Self, Zijian Feng, and Yu Wang. Bayesian penalized spline models for the analysis of spatiotemporal count data. Statistics in Medicine, 35(11):1848–1865, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seppa:2016:CNS

[2248] Karri Seppä, Timo Hakulinen, Esa Läärä, and Janne Pitkäniemi. Comparing net survival estimators of cancer patients. Statistics in Medicine, 35(11):1866–1879, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burgess:2016:CIM

[2249] Stephen Burgess, Frank Dudbridge, and Simon G. Thompson. Combining information on multiple instrumental variables in Mendelian randomization: comparison of allele score and summarized data methods. Statistics in Medicine, 35(11):1880–1906, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2016:ROC

[2250] Dan Wang, Kristopher Attwood, and Lili Tian. Receiver operating characteristic analysis under tree orderings of disease classes. Statistics in Medicine, 35(11):1907–1926, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:2016:RDS

[2251] Toshiro Tango. On the recent debate on the space-time scan statistic for prospective surveillance. Statistics in Medicine, 35(11):1927–1928, May 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITw

[2252] Anonymous. Issue information — title page. Statistics in Medicine, 35 (12):1929, May 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIl

[2253] Anonymous. Issue information — info page. Statistics in Medicine, 35 (12):1930, May 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITx

[2254] Anonymous. Issue information — TOC. Statistics in Medicine, 35(12): 1931, May 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Munoz:2016:CRE

[2255] Joel Muñoz, Daniel Alcaide, and Jordi Ocaña. Consumer’s risk in the EMA and FDA regulatory approaches for bioequivalence in highly variable drugs. Statistics in Medicine, 35(12):1933–1943, May 30, 2016. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liao:2016:SSC

[2256] Peng Liao, Predrag Klasnja, Ambuj Tewari, and Susan A. Murphy. Sample size calculations for micro-randomized trials in mHealth. Statistics in Medicine, 35(12):1944–1971, May 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zebrowska:2016:MTE

[2257] Magdalena Zebrowska, Martin Posch, and Dominic Magirr. Maximum type I error rate inflation from sample size reassessment when investigators are blind to treatment labels. Statistics in Medicine, 35(12): 1972–1984, May 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2016:PSD

[2258] Jiaqi Li, Elizabeth Handorf, Justin Bekelman, and Nandita Mitra. Propensity score and doubly robust methods for estimating the effect of treatment on censored cost. Statistics in Medicine, 35(12):1985–1999, May 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Candel:2016:REL

[2259] Math J. J. M. Candel and Gerard J. P. Van Breukelen. Repairing the efficiency loss due to varying cluster sizes in two-level two-armed randomized trials with heterogeneous clustering. Statistics in Medicine, 35(12): 2000–2015, May 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2016:UIP

[2260] Changyu Shen and Xiaochun Li. On the uncertainty of individual prediction because of sampling predictors. Statistics in Medicine, 35(12): 2016–2030, May 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hillis:2016:EBL

[2261] Stephen L. Hillis. Equivalence of binormal likelihood-ratio and bi-chi-squared ROC curve models. Statistics in Medicine, 35(12):2031–2057, May 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moss:2016:CCP

[2262] Angela Moss, E. Juarez-Colunga, Farouk Nathoo, Brandie Wagner, and Scott Sagel. A comparison of change point models with application to longitudinal lung function measurements in children with cystic fibrosis. Statistics in Medicine, 35(12):2058–2073, May 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arpino:2016:PSM

[2263] Bruno Arpino and Massimo Cannas. Propensity score matching with clustered data. an application to the estimation of the impact of caesarean section on the Apgar score. Statistics in Medicine, 35(12):2074– 2091, May 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gajic-Veljanoski:2016:TBB

[2264] Olga Gajic-Veljanoski, Angela M. Cheung, Ahmed M. Bayoumi, and George Tomlinson. A tutorial on Bayesian bivariate meta-analysis of mixed binary-continuous outcomes with missing treatment effects. Statistics in Medicine, 35(12):2092–2108, May 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crisp:2016:LPB

[2265] Adam Crisp. On the limitations of permuted blocked randomization. Statistics in Medicine, 35(12):2109–2110, May 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berger:2016:DUS

[2266] Vance W. Berger and Rebecca Agnor. Delayed unmasking and selection bias. Statistics in Medicine, 35(12):2111–2112, May 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITy

[2267] Anonymous. Issue information — title page. Statistics in Medicine, 35 (13):2113, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIm

[2268] Anonymous. Issue information — info page. Statistics in Medicine, 35 (13):2114, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITz

[2269] Anonymous. Issue information — TOC. Statistics in Medicine, 35(13): 2115, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hlavin:2016:EEE

[2270] Gerald Hlavin, Franz Koenig, Christoph Male, Martin Posch, and Peter Bauer. Evidence, eminence and extrapolation. Statistics in Medicine, 35(13):2117–2132, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gail:2016:CSA

[2271] Mitchell H. Gail, Jincao Wu, Molin Wang, Shiaw-Shyuan Yaun, Nancy R. Cook, A. Heather Eliassen, Marjorie L. McCullough, Kai Yu, Anne Zeleniuch-Jacquotte, Stephanie A. Smith-Warner, Regina G. Ziegler, and Raymond J. Carroll. Calibration and seasonal adjustment for matched case-control studies of vitamin D and cancer. Statistics in Medicine, 35 (13):2133–2148, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Girling:2016:SEO

[2272] Alan J. Girling and Karla Hemming. Statistical efficiency and optimal design for stepped cluster studies under linear mixed effects models. Statistics in Medicine, 35(13):2149–2166, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2016:TSA

[2273] Xuelin Huang, Fangrong Yan, Jing Ning, Ziding Feng, Sangbum Choi, and Jorge Cortes. A two-stage approach for dynamic prediction of timeto-event distributions. Statistics in Medicine, 35(13):2167–2182, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lok:2016:EPC

[2274] Judith J. Lok and Michael D. Hughes. Evaluating predictors of competing risk outcomes when censoring depends on time-dependent covariates, with application to safety and efficacy of HIV treatment. Statistics in Medicine, 35(13):2183–2194, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rogers:2016:ARE

[2275] Jennifer K. Rogers, Alex Yaroshinsky, Stuart J. Pocock, David Stokar, and Janice Pogoda. Analysis of recurrent events with an associated informative dropout time: Application of the joint frailty model. Statistics in Medicine, 35(13):2195–2205, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rao:2016:RTS

[2276] Yao Rao and Brendan McCabe. Real-time surveillance for abnormal events: the case of influenza outbreaks. Statistics in Medicine, 35(13): 2206–2220, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ertefaie:2016:QLR

[2277] Ashkan Ertefaie, Susan Shortreed, and Bibhas Chakraborty. Q-learning residual analysis: application to the effectiveness of sequences of antipsychotic medications for patients with schizophrenia. Statistics in Medicine, 35(13):2221–2234, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2016:IRP

[2278] Binghui Liu, Xiaotong Shen, and Wei Pan. Integrative and regularized principal component analysis of multiple sources of data. Statistics in Medicine, 35(13):2235–2250, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2016:EAV

[2279] Jihnhee Yu, Luge Yang, Albert Vexler, and Alan D. Hutson. Easy and accurate variance estimation of the nonparametric estimator of the partial area under the ROC curve and its application. Statistics in Medicine, 35(13):2251–2282, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2016:QRL

[2280] Molin Wang, Xiaomei Liao, Francine Laden, and Donna Spiegelman. Quantifying risk over the life course — latency, age-related susceptibility, and other time-varying exposure metrics. Statistics in Medicine, 35(13): 2283–2295, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2016:ECIb

[2281] Xinran Li and Peng Ding. Exact confidence intervals for the average causal effect on a binary outcome. Statistics in Medicine, 35(13):2296, June 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITba

[2282] Anonymous. Issue information — title page. Statistics in Medicine, 35 (14):2297, June 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIn

[2283] Anonymous. Issue information — info page. Statistics in Medicine, 35 (14):2298, June 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbb

[2284] Anonymous. Issue information — TOC. Statistics in Medicine, 35(14): 2299, June 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2016:ACT

[2285] Jung-Tzu Liu, Hsiao-Hui Tsou, K. K. Gordon Lan, Chi-Tian Chen, Yi-Hsuan Lai, Wan-Jung Chang, Chyng-Shyan Tzeng, and Chin-Fu Hsiao. Assessing the consistency of the treatment effect under the discrete random effects model in multiregional clinical trials. Statistics in Medicine, 35(14):2301–2314, June 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2016:VPCa

[2286] Zhenzhen Xu, Michael Proschan, and Shiowjen Lee. Validity and power considerations on hypothesis testing under minimization. Statistics in Medicine, 35(14):2315–2327, June 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Francq:2016:HRP

[2287] Bernard G. Francq and Bernadette Govaerts. How to regress and predict in a Bland–Altman plot? Review and contribution based on tolerance intervals and correlated-errors-in-variables models. Statistics in Medicine, 35(14):2328–2358, June 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2016:HTC

[2288] Chyong-Mei Chen and Chen-Hsin Chen. Heteroscedastic transformation cure regression models. Statistics in Medicine, 35(14):2359–2376, June 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nikoloulopoulos:2016:CSV

[2289] Aristidis K. Nikoloulopoulos. Correlation structure and variable selection in generalized estimating equations via composite likelihood information criteria. Statistics in Medicine, 35(14):2377–2390, June 30, 2016. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goicoa:2016:AST

[2290] T. Goicoa, M. D. Ugarte, J. Etxeberria, and A. F. Militino. Age-space-time CAR models in Bayesian disease mapping. Statistics in Medicine, 35(14):2391–2405, June 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:2016:PDM

[2291] Trevor F. Cox and Gabriela Czanner. A practical divergence measure for survival distributions that can be estimated from Kaplan–Meier curves. Statistics in Medicine, 35(14):2406–2421, June 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilani:2016:SCR

[2292] Owais Gilani, Lisa A. McKay, Timothy G. Gregoire, Yongtao Guan, Brian P. Leaderer, and Theodore R. Holford. Spatiotemporal calibration and resolution refinement of output from deterministic models. Statistics in Medicine, 35(14):2422–2440, June 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubio:2016:BLR

[2293] Francisco J. Rubio and Marc G. Genton. Bayesian linear regression with skew-symmetric error distributions with applications to survival analysis. Statistics in Medicine, 35(14):2441–2454, June 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ogura:2016:IEM

[2294] Toru Ogura and Takemi Yanagimoto. Improving and extending the Mc-Nemar test using the Bayesian method. Statistics in Medicine, 35(14): 2455–2466, June 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shuster:2016:LER

[2295] Jonathan J. Shuster and Michael A. Walker. Low-event-rate metaanalyses of clinical trials: implementing good practices. Statistics in Medicine, 35(14):2467–2478, June 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:CMA

[2296] Anonymous. Comments on ‘Multiplicity adjustments in testing for bioequivalence’. Statistics in Medicine, 35(14):2479–2480, June 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbc

[2297] Anonymous. Issue information — title page. Statistics in Medicine, 35 (15):2481, July 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIo

[2298] Anonymous. Issue information — info page. Statistics in Medicine, 35 (15):2482, July 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbd

[2299] Anonymous. Issue information — TOC. Statistics in Medicine, 35(15): 2483, July 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dagne:2016:TMN

[2300] Getachew A. Dagne, C. Hendricks Brown, George Howe, Sheppard G. Kellam, and Lei Liu. Testing moderation in network meta-analysis with individual participant data. Statistics in Medicine, 35(15):2485–2502, July 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wiksten:2016:HKM

[2301] Anna Wiksten, Gerta Rücker, and Guido Schwarzer. Hartung–Knapp method is not always conservative compared with fixed-effect metaanalysis. Statistics in Medicine, 35(15):2503–2515, July 10, 2016. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ivanova:2016:RED

[2302] Anastasia Ivanova, Yunfei Wang, and Matthew C. Foster. The rapid enrollment design for Phase I clinical trials. Statistics in Medicine, 35 (15):2516–2524, July 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deSouza:2016:UAD

[2303] Roberto Molina de Souza, Jorge Alberto Achcar, Edson Zangiacomi Martinez, and Josmar Mazucheli. The use of asymmetric distributions in average bioequivalence. Statistics in Medicine, 35(15):2525–2542, July 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jaynes:2016:UBF

[2304] Jessica Jaynes, Weng-Kee Wong, and Hongquan Xu. Using blocked fractional factorial designs to construct discrete choice experiments for healthcare studies. Statistics in Medicine, 35(15):2543–2560, July 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ternes:2016:EEL

[2305] Nils Ternès, Federico Rotolo, and Stefan Michiels. Empirical extensions of the lasso penalty to reduce the false discovery rate in high-dimensional Cox regression models. Statistics in Medicine, 35(15):2561–2573, July 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2016:BRS

[2306] Zhipeng Huang, Jialiang Li, Ching-Yu Cheng, Carol Cheung, and Tien-Yin Wong. Bayesian reclassification statistics for assessing improvements in diagnostic accuracy. Statistics in Medicine, 35(15):2574–2592, July 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dassanayake:2016:ICS

[2307] Sesha Dassanayake and Joshua P. French. An improved cumulative sumbased procedure for prospective disease surveillance for count data in multiple regions. Statistics in Medicine, 35(15):2593–2608, July 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coley:2016:EEH

[2308] Rebecca Yates Coley and Elizabeth R. Brown. Estimating effectiveness in HIV prevention trials with a Bayesian hierarchical compound Poisson frailty model. Statistics in Medicine, 35(15):2609–2634, July 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ye:2016:FCT

[2309] Jun Ye, Yehua Li, Nicole A. Lazar, David J. Schaeffer, and Jennifer E. McDowell. Finding common task-related regions in fMRI data from multiple subjects by periodogram clustering and clustering ensemble. Statistics in Medicine, 35(15):2635–2651, July 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hass:2016:MME

[2310] Zachary Hass, Michael Levine, Laura P. Sands, Jeffrey Ting, and Huiping Xu. The modeling of medical expenditure data from a longitudinal survey using the generalized method of moments (GMM) approach. Statistics in Medicine, 35(15):2652–2664, July 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbe

[2311] Anonymous. Issue information — title page. Statistics in Medicine, 35 (16):2665, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIp

[2312] Anonymous. Issue information — info page. Statistics in Medicine, 35 (16):2666, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbf

[2313] Anonymous. Issue information — TOC. Statistics in Medicine, 35(16): 2667, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ristl:2016:FTC

[2314] Robin Ristl, Florian Frommlet, Armin Koch, and Martin Posch. Fallback tests for co-primary endpoints. Statistics in Medicine, 35(16):2669–2686, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Delorme:2016:TIG

[2315] Phillipe Delorme, Pierre Lafaye de Micheaux, Benoit Liquet, and Jérémie Riou. Type-II generalized family-wise error rate formulas with application to sample size determination. Statistics in Medicine, 35(16):2687– 2714, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2016:ITC

[2316] Shan Jiang, Bingshu Chen, and Dongshengn Tu. Inference on treatmentcovariate interaction based on a nonparametric measure of treatment effects and censored survival data. Statistics in Medicine, 35(16):2715– 2725, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fay:2016:FSP

[2317] Michael P. Fay and Erica H. Brittain. Finite sample pointwise confidence intervals for a survival distribution with right-censored data. Statistics in Medicine, 35(16):2726–2740, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sparapani:2016:NSA

[2318] Rodney A. Sparapani, Brent R. Logan, Robert E. McCulloch, and Purushottam W. Laud. Nonparametric survival analysis using Bayesian additive regression trees (BART). Statistics in Medicine, 35(16):2741–2753, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jelizarow:2016:CAT

[2319] Monika Jelizarow, Ulrich Mansmann, and Jelle J. Goeman. A Cochran– Armitage-type and a score-free global test for multivariate ordinal data. Statistics in Medicine, 35(16):2754–2769, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2016:VSD

[2320] Tian Chen, Pan Wu, Wan Tang, Hui Zhang, Changyong Feng, Jeanne Kowalski, and Xin M. Tu. Variable selection for distribution-free models for longitudinal zero-inflated count responses. Statistics in Medicine, 35 (16):2770–2785, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Warren:2016:BMP

[2321] Joshua L. Warren, Jeanette A. Stingone, Amy H. Herring, Thomas J. Luben, Montserrat Fuentes, Arthur S. Aylsworth, Peter H. Langlois, Lorenzo D. Botto, Adolfo Correa, Andrew F. Olshan, and National Birth Defects Prevention Study. Bayesian multinomial probit modeling of daily windows of susceptibility for maternal PM_{2.5} exposure and congenital heart defects. Statistics in Medicine, 35(16):2786–2801, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:2016:WSA

[2322] Changshuai Wei, Robert C. Elston, and Qing Lu. A weighted U statistic for association analyses considering genetic heterogeneity. Statistics in Medicine, 35(16):2802–2814, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2016:UFM

[2323] Lu Chen, Clarice R. Weinberg, and Jinbo Chen. Using family members to augment genetic case-control studies of a life-threatening disease. Statistics in Medicine, 35(16):2815–2830, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2016:IGT

[2324] Yen-Tsung Huang, Tianxi Cai, and Eunhee Kim. Integrative genomic testing of cancer survival using semiparametric linear transformation models. Statistics in Medicine, 35(16):2831–2844, July 20, 2016. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2016:CTN

[2325] Kung-Jong Lui. Comments on ‘Test non-inferiority (and equivalence) based on the odds ratio under a simple crossover trial’. Statistics in Medicine, 35(16):2845–2846, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2016:CEI

[2326] Ying Xu, K. F. Lam, and Yin Bun Cheung. Correction — estimation of intervention effects using recurrent event time data in the presence of event dependence and a cured fraction. Statistics in Medicine, 35(16): 2847–2848, July 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbg

[2327] Anonymous. Issue information — title page. Statistics in Medicine, 35 (17):2849, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIq

[2328] Anonymous. Issue information — info page. Statistics in Medicine, 35 (17):2850, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbh

[2329] Anonymous. Issue information — TOC. Statistics in Medicine, 35(17): 2851, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

LaVange:2016:RPM

[2330] Lisa M. LaVange and Thomas Permutt. A regulatory perspective on missing data in the aftermath of the NRC report. Statistics in Medicine, 35(17):2853–2864, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Permutt:2016:TER

[2331] Thomas Permutt. A taxonomy of estimands for regulatory clinical trials with discontinuations. Statistics in Medicine, 35(17):2865–2875, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Permutt:2016:SAM

[2332] Thomas Permutt. Sensitivity analysis for missing data in regulatory submissions. Statistics in Medicine, 35(17):2876–2879, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:2016:CTP

[2333] Weichung Joe Shih. Comments on the three papers by the FDA/CDER research team on the regulatory perspective of the missing data problem. Statistics in Medicine, 35(17):2880–2886, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koch:2016:CMD

[2334] Gary G. Koch and Laura Elizabeth Wiener. Commentary for the missing data working Group’s perspective for regulatory clinical trials, estimands, and sensitivity analyses. Statistics in Medicine, 35(17):2887–2893, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomas:2016:ACT

[2335] Neal Thomas, Ofer Harel, and Roderick J. A. Little. Analyzing clinical trial outcomes based on incomplete daily diary reports. Statistics in Medicine, 35(17):2894–2906, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Galimard:2016:MIA

[2336] Jacques-Emmanuel Galimard, Sylvie Chevret, Camelia Protopopescu, and Matthieu Resche-Rigon. A multiple imputation approach for MNAR mechanisms compatible with Heckman’s model. Statistics in Medicine, 35(17):2907–2920, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ringham:2016:MTP

[2337] Brandy M. Ringham, Sarah M. Kreidler, Keith E. Muller, and Deborah H. Glueck. Multivariate test power approximations for balanced linear mixed models in studies with missing data. Statistics in Medicine, 35(17):2921–2937, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quartagno:2016:MII

[2338] M. Quartagno and J. R. Carpenter. Multiple imputation for IPD metaanalysis: allowing for heterogeneity and studies with missing covariates. Statistics in Medicine, 35(17):2938–2954, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Erler:2016:DMC

[2339] Nicole S. Erler, Dimitris Rizopoulos, Joost van Rosmalen, Vincent W. V. Jaddoe, Oscar H. Franco, and Emmanuel M. E. H. Lesaffre. Dealing with missing covariates in epidemiologic studies: a comparison between multiple imputation and a full Bayesian approach. Statistics in Medicine, 35(17):2955–2974, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2016:FAH

[2340] Jinhyouk Jung, Ofer Harel, and Sangwook Kang. Fitting additive hazards models for case-cohort studies: a multiple imputation approach. Statistics in Medicine, 35(17):2975–2990, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2016:JMI

[2341] Bo Hu, Liang Li, and Tom Greene. Joint multiple imputation for longitudinal outcomes and clinical events that truncate longitudinal followup. Statistics in Medicine, 35(17):2991–3006, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bondarenko:2016:GND

[2342] Irina Bondarenko and Trivellore Raghunathan. Graphical and numerical diagnostic tools to assess suitability of multiple imputations and imputation models. Statistics in Medicine, 35(17):3007–3020, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chowdhry:2016:MAM

[2343] Amit K. Chowdhry, Robert H. Dworkin, and Michael P. McDermott. Meta-analysis with missing study-level sample variance data. Statistics in Medicine, 35(17):3021–3032, July 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbi

[2344] Anonymous. Issue information — title page. Statistics in Medicine, 35 (18):3033, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIr

[2345] Anonymous. Issue information — info page. Statistics in Medicine, 35 (18):3034, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbj

[2346] Anonymous. Issue information — TOC. Statistics in Medicine, 35(18): 3035, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Balan:2016:STA

[2347] Theodor-Adrian Balan, Stephanie E. Boonk, Maarten H. Vermeer, and Hein Putter. Score test for association between recurrent events and a terminal event. Statistics in Medicine, 35(18):3037–3048, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2016:RED

[2348] Shanshan Li, Yifei Sun, Chiung-Yu Huang, Dean A. Follmann, and Richard Krause. Recurrent event data analysis with intermittently observed time-varying covariates. Statistics in Medicine, 35(18):3049–3065, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Charvat:2016:MEH

[2349] Hadrien Charvat, Laurent Remontet, Nadine Bossard, Laurent Roche, Olivier Dejardin, Bernard Rachet, Guy Launoy, Aurélien Belot, and the Censur Working Survival Group. A multilevel excess hazard model to estimate net survival on hierarchical data allowing for non-linear and nonproportional effects of covariates. Statistics in Medicine, 35(18):3066– 3084, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morrison:2016:LPM

[2350] Kathryn T. Morrison, Gavin Shaddick, Sarah B. Henderson, and David L. Buckeridge. A latent process model for forecasting multiple time series in environmental public health surveillance. Statistics in Medicine, 35(18):3085–3100, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2016:DRE

[2351] Xuan Wang, Lauren A. Beste, Marissa M. Maier, and Xiao-Hua Zhou. Double robust estimator of average causal treatment effect for censored medical cost data. Statistics in Medicine, 35(18):3101–3116, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pirikahu:2016:BMC

[2352] Sarah Pirikahu, Geoffrey Jones, Martin L. Hazelton, and Cord Heuer. Bayesian methods of confidence interval construction for the population attributable risk from cross-sectional studies. Statistics in Medicine, 35 (18):3117–3130, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roy:2016:CDA

[2353] Surupa Roy, Subrata Rana, and Kalyan Das. Clustered data analysis under miscategorized ordinal outcomes and missing covariates. Statistics in Medicine, 35(18):3131–3152, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barnhart:2016:AAR

[2354] Huiman X. Barnhart. Assessing agreement with relative area under the coverage probability curve. Statistics in Medicine, 35(18):3153–3165, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donoghoe:2016:EAR

[2355] Mark W. Donoghoe and Ian C. Marschner. Estimation of adjusted rate differences using additive negative binomial regression. Statistics in Medicine, 35(18):3166–3178, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2016:AUC

[2356] Li Li, Babette A. Brumback, Thomas A. Weppelmann, J. Glenn Morris, Jr., and Afsar Ali. Adjusting for unmeasured confounding due to either of two crossed factors with a logistic regression model. Statistics in Medicine, 35(18):3179–3188, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neyens:2016:BCM

[2357] Thomas Neyens, Andrew B. Lawson, Russell S. Kirby, and Christel Faes. The bivariate combined model for spatial data analysis. Statistics in Medicine, 35(18):3189–3202, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2016:NER

[2358] Donghwan Lee, Andrea Ganna, Yudi Pawitan, and Woojoo Lee. Nonparametric estimation of the rediscovery rate. Statistics in Medicine, 35 (18):3203–3212, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rivera:2016:UEH

[2359] C. L. Rivera and T. Lumley. Using the entire history in the analysis of nested case cohort samples. Statistics in Medicine, 35(18):3213–3228, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kedem:2016:IES

[2360] Benjamin Kedem, Lemeng Pan, Wen Zhou, and Carlos A. Coelho. Interval estimation of small tail probabilities — applications in food safety. Statistics in Medicine, 35(18):3229–3240, August 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbk

[2361] Anonymous. Issue information — title page. Statistics in Medicine, 35 (19):3241, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIs

[2362] Anonymous. Issue information — info page. Statistics in Medicine, 35 (19):3242, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbl

[2363] Anonymous. Issue information — TOC. Statistics in Medicine, 35(19): 3243, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bayar:2016:NIE

[2364] Mohamed Amine Bayar, Gwénaël Le Teuff, Stefan Michiels, Daniel J. Sargent, and Marie-Cécile Le Deley. New insights into the evaluation of randomized controlled trials for rare diseases over a long-term research horizon: a simulation study. Statistics in Medicine, 35(19):3245–3258, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2016:MRB

[2365] Xiaoyu Cai, Huiyun Li, and Aiyi Liu. A marginal rank-based inverse normal transformation approach to comparing multiple clinical trial endpoints. Statistics in Medicine, 35(19):3259–3271, August 30, 2016. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Westgate:2016:ICC

[2366] Philip M. Westgate. Intra-cluster correlation selection for cluster randomized trials. Statistics in Medicine, 35(19):3272–3284, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fu:2016:EOT

[2367] Haoda Fu, Jin Zhou, and Douglas E. Faries. Estimating optimal treatment regimes via subgroup identification in randomized control trials and observational studies. Statistics in Medicine, 35(19):3285–3302, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ott:2016:BPC

[2368] Miles Q. Ott, Joseph W. Hogan, Krista J. Gile, Crystal Linkletter, and Nancy P. Barnett. Bayesian peer calibration with application to alcohol use. Statistics in Medicine, 35(19):3303–3318, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wey:2016:ERM

[2369] Andrew Wey, David M. Vock, John Connett, and Kyle Rudser. Estimating restricted mean treatment effects with stacked survival models. Statistics in Medicine, 35(19):3319–3332, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2016:RPR

[2370] Tian Chen, Jeanne Kowalski, Rui Chen, Pan Wu, Hui Zhang, Changyong Feng, and Xin M. Tu. Rank-preserving regression: a more robust rank regression model against outliers. Statistics in Medicine, 35(19):3333– 3346, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hagar:2016:MCM

[2371] Yolanda C. Hagar, Danielle J. Harvey, and Laurel A. Beckett. A multivariate cure model for left-censored and right-censored data with application to colorectal cancer screening patterns. Statistics in Medicine, 35 (19):3347–3367, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martins:2016:BJM

[2372] Rui Martins, Giovani L. Silva, and Valeska Andreozzi. Bayesian joint modeling of longitudinal and spatial survival AIDS data. Statistics in Medicine, 35(19):3368–3384, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2016:OFS

[2373] Lanju Zhang, Lu Cui, and Bo Yang. Optimal flexible sample size design with robust power. Statistics in Medicine, 35(19):3385–3396, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2016:RST

[2374] Mixia Wu, Yu Shu, Zhaohai Li, and Aiyi Liu. Repeated significance tests of linear combinations of sensitivity and specificity of a diagnostic biomarker. Statistics in Medicine, 35(19):3397–3412, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shaw:2016:RTB

[2375] Pamela A. Shaw and Michael P. Fay. A rank test for bivariate time-to-event outcomes when one event is a surrogate. Statistics in Medicine, 35 (19):3413–3423, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salazar:2016:SGE

[2376] Alejandro Salazar, Begoña Ojeda, María Dueñas, Fernando Fernández, and Inmaculada Failde. Simple generalized estimating equations (GEEs) and weighted generalized estimating equations (WGEEs) in longitudinal studies with dropouts: guidelines and implementation in R. Statistics in Medicine, 35(19):3424–3448, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbm

[2377] Anonymous. Issue information — title page. Statistics in Medicine, 35 (20):3449, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIt

[2378] Anonymous. Issue information — info page. Statistics in Medicine, 35 (20):3450, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbn

[2379] Anonymous. Issue information — TOC. Statistics in Medicine, 35(20): 3451, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dorie:2016:FIF

[2380] Vincent Dorie, Masataka Harada, Nicole Bohme Carnegie, and Jennifer Hill. A flexible, interpretable framework for assessing sensitivity to unmeasured confounding. Statistics in Medicine, 35(20):3453–3470, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2016:LRS

[2381] Xiaochun Li, Huilin Li, Man Jin, and Judith D. Goldberg. Likelihood ratio and score tests to test the non-inferiority (or equivalence) of the odds ratio in a crossover study with binary outcomes. Statistics in Medicine, 35(20):3471–3481, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DiazOrdaz:2016:MIM

[2382] K. DiazOrdaz, M. G. Kenward, M. Gomes, and R. Grieve. Multiple imputation methods for bivariate outcomes in cluster randomised trials. Statistics in Medicine, 35(20):3482–3496, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chu:2016:ADM

[2383] Yiyi Chu, Haitao Pan, and Ying Yuan. Adaptive dose modification for phase I clinical trials. Statistics in Medicine, 35(20):3497–3508, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dimou:2016:MMM

[2384] Niki L. Dimou, Maria Adam, and Pantelis G. Bagos. A multivariate method for meta-analysis and comparison of diagnostic tests. Statistics in Medicine, 35(20):3509–3523, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2016:HBA

[2385] Hong Zhao, James S. Hodges, Haijun Ma, Qi Jiang, and Bradley P. Carlin. Hierarchical Bayesian approaches for detecting inconsistency in network meta-analysis. Statistics in Medicine, 35(20):3524–3536, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2016:VEC

[2386] Baiming Zou, Fei Zou, Jonathan J. Shuster, Patrick J. Tighe, Gary G. Koch, and Haibo Zhou. On variance estimate for covariate adjustment by propensity score analysis. Statistics in Medicine, 35(20):3537–3548, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rodriguez-Girondo:2016:MTM

[2387] Mar Rodríguez-Girondo and Jacobo de Uña-Álvarez. Methods for testing the Markov condition in the illness-death model: a comparative study. Statistics in Medicine, 35(20):3549–3562, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Francq:2016:DMB

[2388] Bernard G. Francq and Olivier Cartiaux. Delta method and bootstrap in linear mixed models to estimate a proportion when no event is observed: application to intralesional resection in bone tumor surgery. Statistics in Medicine, 35(20):3563–3582, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brilleman:2016:JLH

[2389] Samuel L. Brilleman, Michael J. Crowther, Margaret T. May, Mark Gompels, and Keith R. Abrams. Joint longitudinal hurdle and time-to-event models: an application related to viral load and duration of the first treatment regimen in patients with HIV initiating therapy. Statistics in Medicine, 35(20):3583–3594, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steingrimsson:2016:DRS

[2390] Jon Arni Steingrimsson, Liqun Diao, Annette M. Molinaro, and Robert L. Strawderman. Doubly robust survival trees. Statistics in Medicine, 35(20):3595–3612, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ye:2016:TSE

[2391] Lei Ye, Ada O. Youk, Susan M. Sereika, Stewart J. Anderson, and Lora E. Burke. A three-step estimation procedure using local polynomial smoothing for inconsistently sampled longitudinal data. Statistics in Medicine, 35(20):3613–3622, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thulin:2016:TST

[2392] M. Thulin. Two-sample tests and one-way MANOVA for multivariate biomarker data with nondetects. Statistics in Medicine, 35(20):3623– 3644, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Titman:2016:ETS

[2393] Andrew C. Titman. Estimation of time-shift models with application to survival calibration in health technology assessment. Statistics in Medicine, 35(20):3645–3656, September 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbo

[2394] Anonymous. Issue information — title page. Statistics in Medicine, 35 (21):3657, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIu

[2395] Anonymous. Issue information — info page. Statistics in Medicine, 35 (21):3658, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbp

[2396] Anonymous. Issue information — TOC. Statistics in Medicine, 35(21): 3659, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crippa:2016:NMB

[2397] Alessio Crippa, Polyna Khudyakov, Molin Wang, Nicola Orsini, and Donna Spiegelman. A new measure of between-studies heterogeneity in meta-analysis. Statistics in Medicine, 35(21):3661–3675, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gomes:2016:HIC

[2398] Manuel Gomes, Laura Hatfield, and Sharon-Lise Normand. Handling incomplete correlated continuous and binary outcomes in meta-analysis of individual participant data. Statistics in Medicine, 35(21):3676–3689, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2016:CTE

[2399] Jin Zhang and Cong Chen. Correcting treatment effect for treatment switching in randomized oncology trials with a modified iterative parametric estimation method. Statistics in Medicine, 35(21):3690–3703, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2016:MSS

[2400] Xin Victoria Wang, Bernard Cole, Marco Bonetti, and Richard D. Gelber. Meta-STEPP: subpopulation treatment effect pattern plot for individual patient data meta-analysis. Statistics in Medicine, 35(21):3704– 3716, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Balzer:2016:TEI

[2401] Laura B. Balzer, Maya L. Petersen, Mark J. van der Laan, and the Search Collaboration. Targeted estimation and inference for the sample average treatment effect in trials with and without pair-matching. Statistics in Medicine, 35(21):3717–3732, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Westgate:2016:IPS

[2402] Philip M. Westgate and Woodrow W. Burchett. Improving power in small-sample longitudinal studies when using generalized estimating equations. Statistics in Medicine, 35(21):3733–3744, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilbert:2016:PSS

[2403] Peter B. Gilbert, Holly E. Janes, and Yunda Huang. Power/sample size calculations for assessing correlates of risk in clinical efficacy trials. Statistics in Medicine, 35(21):3745–3759, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Iasonos:2016:DMP

[2404] Alexia Iasonos, Nolan A. Wages, Mark R. Conaway, Ken Cheung, Ying Yuan, and John O’Quigley. Dimension of model parameter space and operating characteristics in adaptive dose-finding studies. Statistics in Medicine, 35(21):3760–3775, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenblum:2016:GSD

[2405] Michael Rosenblum, Brandon Luber, Richard E. Thompson, and Daniel Hanley. Group sequential designs with prospectively planned rules for subpopulation enrichment. Statistics in Medicine, 35(21):3776–3791, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fong:2016:CBL

[2406] Youyi Fong, Shuxin Yin, and Ying Huang. Combining biomarkers linearly and nonlinearly for classification using the area under the ROC curve. Statistics in Medicine, 35(21):3792–3809, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gordon:2016:CAM

[2407] Andrew S. Gordon, Adele H. Marshall, and Karen J. Cairns. A conditional approach for modelling patient readmissions to hospital using a mixture of Coxian phase-type distributions incorporating Bayes’ theorem. Statistics in Medicine, 35(21):3810–3826, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

MacNab:2016:LMC

[2408] Ying C. MacNab. Linear models of coregionalization for multivariate lattice data: a general framework for coregionalized multivariate CAR models. Statistics in Medicine, 35(21):3827–3850, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Warasi:2016:EPM

[2409] Md S. Warasi, Joshua M. Tebbs, Christopher S. McMahan, and Christopher R. Bilder. Estimating the prevalence of multiple diseases from two-stage hierarchical pooling. Statistics in Medicine, 35(21):3851–3864, September 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbq

[2410] Anonymous. Issue information — title page. Statistics in Medicine, 35 (22):3865, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIv

[2411] Anonymous. Issue information — info page. Statistics in Medicine, 35 (22):3866, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbr

[2412] Anonymous. Issue information — TOC. Statistics in Medicine, 35(22): 3867, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gruber:2016:LEC

[2413] Susan Gruber and Eric Tchetgen Tchetgen. Limitations of empirical calibration of p-values using observational data. Statistics in Medicine, 35(22):3869–3882, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schuemie:2016:REC

[2414] Martijn J. Schuemie, George Hripcsak, Patrick B. Ryan, David Madigan, and Marc A. Suchard. Robust empirical calibration of p-values using observational data. Statistics in Medicine, 35(22):3883–3888, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Franklin:2016:VDM

[2415] Jessica M. Franklin. p-values and decision-making: discussion of ‘Limitations of empirical calibration of p-values using observational data’. Statistics in Medicine, 35(22):3889–3891, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yuan:2016:MPB

[2416] Ying Yuan, Beibei Guo, Mark Munsell, Karen Lu, and Amir Jazaeri. MIDAS: a practical Bayesian design for platform trials with molecularly targeted agents. Statistics in Medicine, 35(22):3892–3906, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robertson:2016:UES

[2417] David S. Robertson, A. Toby Prevost, and Jack Bowden. Unbiased estimation in seamless phase II/III trials with unequal treatment effect variances and hypothesis-driven selection rules. Statistics in Medicine, 35(22):3907–3922, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2016:TEB

[2418] Yan Sun, Wenting Wu, and Daniel Sargent. Testing of evaluation bias for progression free survival endpoint in oncology clinical trials. Statistics in Medicine, 35(22):3923–3932, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ferrer:2016:JML

[2419] Loïc Ferrer, Virginie Rondeau, James Dignam, Tom Pickles, Hélène Jacqmin-Gadda, and Cécile Proust-Lima. Joint modelling of longitudinal and multi-state processes: application to clinical progressions in prostate cancer. Statistics in Medicine, 35(22):3933–3948, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dankowski:2016:CRF

[2420] Theresa Dankowski and Andreas Ziegler. Calibrating random forests for probability estimation. Statistics in Medicine, 35(22):3949–3960, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gu:2016:SDN

[2421] Xiangdong Gu and Raji Balasubramanian. Study design for nonrecurring, time-to-event outcomes in the presence of error-prone diagnostic tests or self-reports. Statistics in Medicine, 35(22):3961–3975, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bebu:2016:PMC

[2422] Ionut Bebu, Thomas Mathew, and John M. Lachin. Probabilistic measures of cost-effectiveness. Statistics in Medicine, 35(22):3976–3986, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buonaccorsi:2016:CBM

[2423] John Buonaccorsi, Agnieszka Prochenka, Magne Thoresen, and Rafal Ploski. Correcting for binomial measurement error in predictors in regression with application to analysis of DNA methylation rates by bisulfite sequencing. Statistics in Medicine, 35(22):3987–4007, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lok:2016:DEC

[2424] Judith J. Lok. Defining and estimating causal direct and indirect effects when setting the mediator to specific values is not feasible. Statistics in Medicine, 35(22):4008–4020, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schorning:2016:MSV

[2425] Kirsten Schorning, Björn Bornkamp, Frank Bretz, and Holger Dette. Model selection versus model averaging in dose finding studies. Statistics in Medicine, 35(22):4021–4040, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2016:GAU

[2426] Kaifeng Lu. Graphical approaches using a Bonferroni mixture of weighted Simes tests. Statistics in Medicine, 35(22):4041–4055, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2016:DPB

[2427] Peter C. Austin, Douglas S. Lee, Ralph B. D’Agostino, and Jason P. Fine. Developing points-based risk-scoring systems in the presence of competing risks. Statistics in Medicine, 35(22):4056–4072, September 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbs

[2428] Anonymous. Issue information — title page. Statistics in Medicine, 35(23):4073, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIw

[2429] Anonymous. Issue information — info page. Statistics in Medicine, 35(23):4074, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbt

[2430] Anonymous. Issue information — TOC. Statistics in Medicine, 35(23): 4075, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diaz:2016:MIB

[2431] Francisco J. Diaz. Measuring the individual benefit of a medical or behavioral treatment using generalized linear mixed-effects models. Statistics in Medicine, 35(23):4077–4092, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sato:2016:ADF

[2432] Hiroyuki Sato, Akihiro Hirakawa, and Chikuma Hamada. An adaptive dose-finding method using a change-point model for molecularly targeted agents in phase I trials. Statistics in Medicine, 35(23):4093–4109, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2016:TEB

[2433] Kung-Jong Lui and Kuang-Chao Chang. Test equality in binary data for a 4×4 crossover trial under a Latin-square design. Statistics in Medicine, 35(23):4110–4123, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collins:2016:QID

[2434] Gary S. Collins, Emmanuel O. Ogundimu, Jonathan A. Cook, Yannick Le Manach, and Douglas G. Altman. Quantifying the impact of different approaches for handling continuous predictors on the performance of a prognostic model. Statistics in Medicine, 35(23):4124–4135, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanKlaveren:2016:NCM

[2435] David van Klaveren, Mithat Gönen, Ewout W. Steyerberg, and Yvonne Vergouwe. A new concordance measure for risk prediction models in external validation settings. Statistics in Medicine, 35(23):4136–4152, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Honerkamp-Smith:2016:TME

[2436] Gordon Honerkamp-Smith and Ronghui Xu. Three measures of explained variation for correlated survival data under the proportional hazards mixed-effects model. Statistics in Medicine, 35(23):4153–4165, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eshaghi:2016:TVC

[2437] Ehsan Eshaghi, Hossein Baghishani, and Davood Shahsavani. Timevarying coefficients models for recurrent event data when different varying coefficients admit different degrees of smoothness: application to heart disease modeling. Statistics in Medicine, 35(23):4166–4182, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Balan:2016:ACF

[2438] Theodor A. Balan, Marianne A. Jonker, Paul C. Johannesma, and Hein Putter. Ascertainment correction in frailty models for recurrent events data. Statistics in Medicine, 35(23):4183–4201, October 15, 2016. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grigorova:2016:CPA

[2439] D. Grigorova and R. Gueorguieva. Correlated probit analysis of repeatedly measured ordinal and continuous outcomes with application to the health and retirement study. Statistics in Medicine, 35(23):4202–4225, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xing:2016:RAS

[2440] Chuanhua Xing, Janice M. McCarthy, Josée Dupuis, L. Adrienne Cupples, James B. Meigs, Xihong Lin, and Andrew S. Allen. Robust analysis of secondary phenotypes in case-control genetic association studies. Statistics in Medicine, 35(23):4226–4237, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saarela:2016:FPA

[2441] Olli Saarela and Zhihui (Amy) Liu. A flexible parametric approach for estimating continuous-time inverse probability of treatment and censoring weights. Statistics in Medicine, 35(23):4238–4251, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2016:BAM

[2442] Juxin Liu, Paul Gustafson, and Dezheng Huo. Bayesian adjustment for the misclassification in both dependent and independent variables with application to a breast cancer study. Statistics in Medicine, 35(23):4252– 4263, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heath:2016:EEV

[2443] Anna Heath, Ioanna Manolopoulou, and Gianluca Baio. Estimating the expected value of partial perfect information in health economic evaluations using integrated nested Laplace approximation. Statistics in Medicine, 35(23):4264–4280, October 15, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIx

[2444] Anonymous. Issue information — info page. Statistics in Medicine, 35(24):4281, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbu

[2445] Anonymous. Issue information — title page. Statistics in Medicine, 35(24):4282, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbv

[2446] Anonymous. Issue information — TOC. Statistics in Medicine, 35(24): 4283, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murray:2016:UBD

[2447] Thomas A. Murray, Peter F. Thall, and Ying Yuan. Utility-based designs for randomized comparative trials with categorical outcomes. Statistics in Medicine, 35(24):4285–4305, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2016:CEE

[2448] Ye Shen, Hui Huang, and Yongtao Guan. A conditional estimating equation approach for recurrent event data with additional longitudinal information. Statistics in Medicine, 35(24):4306–4319, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lemme:2016:ETA

[2449] Francesca Lemme, Gerard J. P. van Breukelen, and Martijn P. F. Berger. Efficient treatment allocation in 2 × 2 cluster randomized trials, when costs and variances are heterogeneous. Statistics in Medicine, 35(24):

4320–4334, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shepherd:2016:CRM

[2450] Bryan E. Shepherd, Qi Liu, Nathaniel Mercaldo, Cathy A. Jenkins, Bryan Lau, Stephen R. Cole, Michael S. Saag, and Timothy R. Sterling. Comparing results from multiple imputation and dynamic marginal structural models for estimating when to start antiretroviral therapy. Statistics in Medicine, 35(24):4335–4351, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bantis:2016:CTC

[2451] Leonidas E. Bantis and Ziding Feng. Comparison of two correlated ROC curves at a given specificity or sensitivity level. Statistics in Medicine, 35 (24):4352–4367, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mulatya:2016:ETE

[2452] Caroline M. Mulatya, Alexander C. McLain, Bo Cai, James W. Hardin, and Paul S. Albert. Estimating time to event characteristics via longitudinal threshold regression models — an application to cervical dilation progression. Statistics in Medicine, 35(24):4368–4379, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2016:BPM

[2453] Fengqing Zhang, Wenxin Jiang, Patrick Wong, and Ji-Ping Wang. A Bayesian probit model with spatially varying coefficients for brain decoding using fMRI data. Statistics in Medicine, 35(24):4380–4397, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Follmann:2016:WRG

[2454] Dean Follmann, Chiung-Yu Huang, and Erin Gabriel. Who really gets strep sore throat? Confounding and effect modification of a time-varying exposure on recurrent events. Statistics in Medicine, 35(24):4398–4412, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhan:2016:ATE

[2455] Zhuozhao Zhan, Geertruida H. de Bock, Theo Wiggers, and Edwin van den Heuvel. The analysis of terminal endpoint events in stepped wedge designs. Statistics in Medicine, 35(24):4413–4426, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersson:2016:DED

[2456] Claes Andersson, Peter Guttorp, and Aila Särkkä. Discovering early diabetic neuropathy from epidermal nerve fiber patterns. Statistics in Medicine, 35(24):4427–4442, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kovacic:2016:GMA

[2457] Jelena Kovacić and Veda Marija Varnai. A graphical model approach to systematically missing data in meta-analysis of observational studies. Statistics in Medicine, 35(24):4443–4458, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aralis:2016:MCU

[2458] Hilary J. Aralis, Pamina M. Gorbach, and Ron Brookmeyer. Measuring concurrency using a joint multistate and point process model for retrospective sexual history data. Statistics in Medicine, 35(24):4459–4473, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2016:IDD

[2459] Xiaolu Zhu and Annie Qu. Individualizing drug dosage with longitudinal data. Statistics in Medicine, 35(24):4474–4488, October 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbw

[2460] Anonymous. Issue information — title page. Statistics in Medicine, 35 (25):4489, November 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIy

[2461] Anonymous. Issue information — info page. Statistics in Medicine, 35 (25):4490, November 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbx

[2462] Anonymous. Issue information — TOC. Statistics in Medicine, 35(25): 4491, November 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keogh:2016:SIR

[2463] Ruth H. Keogh, Raymond J. Carroll, Janet A. Tooze, Sharon I. Kirkpatrick, and Laurence S. Freedman. Statistical issues related to dietary intake as the response variable in intervention trials. Statistics in Medicine, 35(25):4493–4508, November 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2016:OBA

[2464] Dandan Li and Siva Sivaganesan. An objective Bayesian analysis of a crossover design via model selection and model averaging. Statistics in Medicine, 35(25):4509–4527, November 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Balzer:2016:APS

[2465] Laura B. Balzer, Mark J. van der Laan, Maya L. Petersen, and the Search Collaboration. Adaptive pre-specification in randomized trials with and without pair-matching. Statistics in Medicine, 35(25):4528– 4545, November 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2016:TMC

[2466] Qingxia Chen, Hui Nian, Yuwei Zhu, H. Keipp Talbot, Marie R. Griffin, and Frank E. Harrell, Jr. Too many covariates and too few cases? — a comparative study. Statistics in Medicine, 35(25):4546–4558, November 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hattori:2016:EPC

[2467] Satoshi Hattori and Xiao-Hua Zhou. Evaluation of predictive capacities of biomarkers based on research synthesis. Statistics in Medicine, 35 (25):4559–4572, November 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Iannario:2016:TCI

[2468] Maria Iannario and Joseph B. Lang. Testing conditional independence in sets of I × J tables by means of moment and correlation score tests with application to HPV vaccine. Statistics in Medicine, 35(25):4573– 4587, November 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burne:2016:MRB

[2469] Rebecca M. Burne and Michal Abrahamowicz. Martingale residual-based method to control for confounders measured only in a validation sample in time-to-event analysis. Statistics in Medicine, 35(25):4588–4606, November 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernhardt:2016:FCR

[2470] Paul W. Bernhardt. A flexible cure rate model with dependent censoring and a known cure threshold. Statistics in Medicine, 35(25):4607–4623, November 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Deng:2016:ECM

[2471] Dianliang Deng. Estimating the cumulative mean function for history process with time-dependent covariates and censoring mechanism. Statistics in Medicine, 35(25):4624–4636, November 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sennhenn-Reulen:2016:SFL

[2472] Holger Sennhenn-Reulen and Thomas Kneib. Structured fusion lasso penalized multi-state models. Statistics in Medicine, 35(25):4637–4659, November 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fernandez:2016:GFT

[2473] Daniel Fernández and Ivy Liu. A goodness-of-fit test for the ordered stereotype model. Statistics in Medicine, 35(25):4660–4696, November 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITby

[2474] Anonymous. Issue information — title page. Statistics in Medicine, 35 (26):4697, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIz

[2475] Anonymous. Issue information — info page. Statistics in Medicine, 35 (26):4698, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITbz

[2476] Anonymous. Issue information — TOC. Statistics in Medicine, 35(26): 4699, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beesley:2016:MIM

[2477] Lauren J. Beesley, Jonathan W. Bartlett, Gregory T. Wolf, and Jeremy M. G. Taylor. Multiple imputation of missing covariates for the Cox proportional hazards cure model. Statistics in Medicine, 35(26):4701– 4717, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hooper:2016:SSC

[2478] Richard Hooper, Steven Teerenstra, Esther de Hoop, and Sandra Eldridge. Sample size calculation for stepped wedge and other longitudinal cluster randomised trials. Statistics in Medicine, 35(26):4718–4728, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ren:2016:LLV

[2479] Chunfeng Ren and Yongyun Shin. Longitudinal latent variable models given incompletely observed biomarkers and covariates. Statistics in Medicine, 35(26):4729–4745, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hattori:2016:TDS

[2480] Satoshi Hattori and Xiao-Hua Zhou. Time-dependent summary receiver operating characteristics for meta-analysis of prognostic studies. Statistics in Medicine, 35(26):4746–4763, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ke:2016:SVC

[2481] Yuan Ke, Bo Fu, and Wenyang Zhang. Semi-varying coefficient multinomial logistic regression for disease progression risk prediction. Statistics in Medicine, 35(26):4764–4778, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Valeri:2016:ECC

[2482] Linda Valeri and Brent A. Coull. Estimating causal contrasts involving intermediate variables in the presence of selection bias. Statistics in Medicine, 35(26):4779–4793, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wen:2016:BMJ

[2483] Sijin Wen, Xuelin Huang, Ralph F. Frankowski, Janice N. Cormier, and Peter Pisters. A Bayesian multivariate joint frailty model for disease recurrences and survival. Statistics in Medicine, 35(26):4794–4812, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andrinopoulou:2016:BSA

[2484] Eleni-Rosalina Andrinopoulou and Dimitris Rizopoulos. Bayesian shrinkage approach for a joint model of longitudinal and survival outcomes assuming different association structures. Statistics in Medicine, 35(26): 4813–4823, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mi:2016:ULM

[2485] Xiaojuan Mi, Bradley G. Hammill, Lesley H. Curtis, Edward Chia-Cheng Lai, and Soko Setoguchi. Use of the landmark method to address immortal person-time bias in comparative effectiveness research: a simulation study. Statistics in Medicine, 35(26):4824–4836, November 20, 2016. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Loh:2016:ISD

[2486] Wei-Yin Loh, Haoda Fu, Michael Man, Victoria Champion, and Menggang Yu. Identification of subgroups with differential treatment effects for longitudinal and multiresponse variables. Statistics in Medicine, 35 (26):4837–4855, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sidik:2016:CVE

[2487] Kurex Sidik and Jeffrey N. Jonkman. A comparison of the variance estimation methods for heteroscedastic nonlinear models. Statistics in Medicine, 35(26):4856–4874, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fernandez-Fontelo:2016:URD

[2488] Amanda Fernández-Fontelo, Alejandra Cabaña, Pedro Puig, and David Moriña. Under-reported data analysis with INAR-hidden Markov chains. Statistics in Medicine, 35(26):4875–4890, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vuollo:2016:BRA

[2489] Ville Vuollo, Lasse Holmström, Henri Aarnivala, Virpi Harila, Tuomo Heikkinen, Pertti Pirttiniemi, and Arja Marita Valkama. Book review: Analyzing infant head flatness and asymmetry using kernel density estimation of directional surface data from a craniofacial 3D model. Statistics in Medicine, 35(26):4891–4904, November 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITca

[2490] Anonymous. Issue information — title page. Statistics in Medicine, 35 (27):4905, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIba

[2491] Anonymous. Issue information — info page. Statistics in Medicine, 35 (27):4906, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITcb

[2492] Anonymous. Issue information — TOC. Statistics in Medicine, 35(27): 4907, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spencer:2016:ADU

[2493] Amy V. Spencer, Chris Harbron, Adrian Mander, James Wason, and Ian Peers. An adaptive design for updating the threshold value of a continuous biomarker. Statistics in Medicine, 35(27):4909–4923, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2016:PBD

[2494] Liangcai Zhang and Ying Yuan. A practical Bayesian design to identify the maximum tolerated dose contour for drug combination trials. Statistics in Medicine, 35(27):4924–4936, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rudolph:2016:OCP

[2495] Kara E. Rudolph, K. Ellicott Colson, Elizabeth A. Stuart, and Jennifer Ahern. Optimally combining propensity score subclasses. Statistics in Medicine, 35(27):4937–4947, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2016:EOE

[2496] Wei Wang and Michael E. Griswold. Estimating overall exposure effects for the clustered and censored outcome using random effect tobit regression models. Statistics in Medicine, 35(27):4948–4960, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Resa:2016:ESM

[2497] María de los Angeles Resa and José R. Zubizarreta. Evaluation of subset matching methods and forms of covariate balance. Statistics in Medicine, 35(27):4961–4979, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hubbard:2016:USM

[2498] R. A. Hubbard, J. Lange, Y. Zhang, B. A. Salim, J. R. Stroud, and L. Y. T. Inoue. Using semi-Markov processes to study timeliness and tests used in the diagnostic evaluation of suspected breast cancer. Statistics in Medicine, 35(27):4980–4993, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2016:ESL

[2499] Huaihou Chen, Bingxin Zhao, Eric C. Porges, Ronald A. Cohen, and Natalie C. Ebner. Edgewise and subgraph-level tests for brain networks. Statistics in Medicine, 35(27):4994–5008, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ren:2016:AEL

[2500] Kaili Ren, Christopher A. Drummond, Pamela S. Brewster, Steven T. Haller, Jiang Tian, Christopher J. Cooper, and Biao Zhang. An alternative empirical likelihood method in missing response problems and causal inference. Statistics in Medicine, 35(27):5009–5028, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2016:NPP

[2501] Dandan Xu, Arkendu Chatterjee, and Michael Daniels. A note on posterior predictive checks to assess model fit for incomplete data. Statistics in Medicine, 35(27):5029–5039, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2016:CLA

[2502] Lingxiao Wang, Barry I. Graubard, and Yan Li. A composite likelihood approach in testing for Hardy Weinberg equilibrium using family-based genetic survey data. Statistics in Medicine, 35(27):5040–5050, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jensen:2016:MSM

[2503] Aksel K. G. Jensen, Henrik Ravn, Signe Sørup, and Per Andersen. A marginal structural model for recurrent events in the presence of timedependent confounding: non-specific effects of vaccines on child hospitalisations. Statistics in Medicine, 35(27):5051–5069, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neelon:2016:MZMa

[2504] Brian Neelon, A. James O’Malley, and Valerie A. Smith. Modeling zeromodified count and semicontinuous data in health services research. Part 1: background and overview. Statistics in Medicine, 35(27):5070–5093, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neelon:2016:MZMb

[2505] Brian Neelon, A. James O’Malley, and Valerie A. Smith. Modeling zeromodified count and semicontinuous data in health services research. Part 2: case studies. Statistics in Medicine, 35(27):5094–5112, November 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITcc

[2506] Anonymous. Issue information — title page. Statistics in Medicine, 35(28):5113, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIbb

[2507] Anonymous. Issue information — info page. Statistics in Medicine, 35 (28):5114, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITcd

[2508] Anonymous. Issue information — TOC. Statistics in Medicine, 35(28): 5115, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bryan:2016:MAL

[2509] Matthew Bryan and Patrick J. Heagerty. Multivariate analysis of longitudinal rates of change. Statistics in Medicine, 35(28):5117–5134, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perperoglou:2016:SCR

[2510] Aris Perperoglou. A special case of reduced rank models for identification and modelling of time varying effects in survival analysis. Statistics in Medicine, 35(28):5135–5148, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2016:PER

[2511] Nan Xuan Lin and William Edward Henley. Prior event rate ratio adjustment for hidden confounding in observational studies of treatment effectiveness: a pairwise Cox likelihood approach. Statistics in Medicine, 35(28):5149–5169, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2016:MIC

[2512] Benmei Liu, Mandi Yu, Barry I. Graubard, Richard P. Troiano, and Nathaniel Schenker. Multiple imputation of completely missing repeated measures data within person from a complex sample: application to accelerometer data in the national health and nutrition examination survey. Statistics in Medicine, 35(28):5170–5188, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yaesoubi:2016:ICE

[2513] Reza Yaesoubi and Ted Cohen. Identifying cost-effective dynamic policies to control epidemics. Statistics in Medicine, 35(28):5189–5209, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanBuren:2016:IIS

[2514] John VanBuren, Jacob J. Oleson, Gideon K. D. Zamba, and Michael Wall. Integrating independent spatio-temporal replications to assess population trends in disease spread. Statistics in Medicine, 35(28):5210– 5221, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Papachristofi:2016:EEM

[2515] Olympia Papachristofi, Andrew Klein, and Linda Sharples. Evaluation of the effects of multiple providers in complex surgical interventions. Statistics in Medicine, 35(28):5222–5246, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chiang:2016:BVR

[2516] Chin-Tsang Chiang, Ming-Yueh Huang, and Shao-Hsuan Wang. Bias and variance reduction in nonparametric estimation of time-dependent accuracy measures. Statistics in Medicine, 35(28):5247–5266, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armero:2016:BJO

[2517] C. Armero, C. Forné, M. Rué, A. Forte, H. Perpiñán, G. Gómez, and M. Baré. Bayesian joint ordinal and survival modeling for breast cancer risk assessment. Statistics in Medicine, 35(28):5267–5282, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aprahamian:2016:RRW

[2518] Hrayer Aprahamian, Douglas R. Bish, and Ebru K. Bish. Residual risk and waste in donated blood with pooled nucleic acid testing. Statistics in Medicine, 35(28):5283–5301, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dagne:2016:BBC

[2519] Getachew A. Dagne. Bayesian bent-cable growth mixture tobit models for longitudinal data with skewness and detection limit: application to AIDS studies. Statistics in Medicine, 35(28):5302–5314, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

daSilva:2016:CCE

[2520] José Luiz P. da Silva and Enrico A. Colosimo. Comments on ‘Covariance estimators for generalized estimating equations (GEE) in longitudinal analysis with small samples’. Statistics in Medicine, 35(28):5315–5317, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2016:CEGb

[2521] Ming Wang, Lan Kong, Zheng Li, and Lijun Zhang. Covariance estimators for generalized estimating equations (GEE) in longitudinal analysis with small samples. Statistics in Medicine, 35(28):5318–5319, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehta:2016:BDP

[2522] Cyrus Mehta, Helmut Schäfer, Hanna Daniel, and Sebastian Irle. Biomarker-driven population enrichment for adaptive oncology trials with time to event endpoints. Statistics in Medicine, 35(28):5320, December 10, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITce

[2523] Anonymous. Issue information — title page. Statistics in Medicine, 35(29):5321, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIbc

[2524] Anonymous. Issue information — info page. Statistics in Medicine, 35 (29):5322, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITcf

[2525] Anonymous. Issue information — TOC. Statistics in Medicine, 35(29): 5323, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2016:CST

[2526] Hong Sun, Frank Bretz, Oke Gerke, and Werner Vach. Comparing a stratified treatment strategy with the standard treatment in randomized clinical trials. Statistics in Medicine, 35(29):5325–5337, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harrar:2016:ATE

[2527] Solomon Harrar, Anup Amatya, and Leonid Kalachev. Assessing treatment efficacy in the presence of diagnostic errors. Statistics in Medicine, 35(29):5338–5355, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spence:2016:BAS

[2528] Graeme T. Spence, David Steinsaltz, and Thomas R. Fanshawe. A Bayesian approach to sequential meta-analysis. Statistics in Medicine, 35 (29):5356–5375, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Held:2016:OBM

[2529] Leonhard Held, Isaac Gravestock, and Daniel Sabanés Bové. Objective Bayesian model selection for Cox regression. Statistics in Medicine, 35 (29):5376–5390, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kleinman:2016:MRM

[2530] Alan Kleinman. The mathematics of random mutation and natural selection for multiple simultaneous selection pressures and the evolution of antimicrobial drug resistance. Statistics in Medicine, 35(29):5391–5400, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qin:2016:REP

[2531] Guoyou Qin, Jiajia Zhang, Zhongyi Zhu, and Wing Fung. Robust estimation of partially linear models for longitudinal data with dropouts and measurement error. Statistics in Medicine, 35(29):5401–5416, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roberts:2016:AMV

[2532] Eric M. Roberts and Paul B. English. Analysis of multiple-variable missing-not-at-random survey data for child lead surveillance using NHANES. Statistics in Medicine, 35(29):5417–5429, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shardell:2016:IVA

[2533] Michelle Shardell and Luigi Ferrucci. Instrumental variable analysis of multiplicative models with potentially invalid instruments. Statistics in Medicine, 36(3):5430–5447, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cramb:2016:FPA

[2534] Susanna M. Cramb, Kerrie L. Mengersen, Paul C. Lambert, Louise M. Ryan, and Peter D. Baade. A flexible parametric approach to examining spatial variation in relative survival. Statistics in Medicine, 36(3):5448– 5463, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhai:2016:CDD

[2535] Shuyan Zhai, Thomas Mathew, and Yi Huang. Comparison of drug dissolution profiles: a proposal based on tolerance limits. Statistics in Medicine, 36(3):5464–5476, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mitchell:2016:ERR

[2536] Emily M. Mitchell, Torie C. Plowden, and Enrique F. Schisterman. Estimating relative risk of a log-transformed exposure measured in pools. Statistics in Medicine, 36(3):5477–5494, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rhodes:2016:IIP

[2537] Kirsty M. Rhodes, Rebecca M. Turner, Ian R. White, Dan Jackson, David J. Spiegelhalter, and Julian P. T. Higgins. Implementing informative priors for heterogeneity in meta-analysis using meta-regression and pseudo data. Statistics in Medicine, 36(3):5495–5511, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yuan:2016:NCS

[2538] Zhongshang Yuan, Jiadong Ji, Tao Zhang, Yi Liu, Xiaoshuai Zhang, Wei Chen, and Fuzhong Xue. A novel chi-square statistic for detecting group differences between pathways in systems epidemiology. Statistics in Medicine, 36(3):5512–5524, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuznetsova:2016:CVP

[2539] Olga M. Kuznetsova and Yevgen Tymofyeyev. Comments on ‘Validity and power considerations on hypothesis testing under minimization’: by Z. Xu, M. Proschan, and S. Lee, Statistics in Medicine 2016. Statistics in Medicine, 36(3):5525–5526, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2016:VPCb

[2540] Zhenzhen Xu, Michael Proschan, and Shiowjen Lee. Validity and power considerations on hypothesis testing under minimization. Statistics in Medicine, 36(3):5527–5528, December 20, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITcg

[2541] Anonymous. Issue information — title page. Statistics in Medicine, 35(30):5529, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IIIbd

[2542] Anonymous. Issue information — info page. Statistics in Medicine, 35 (30):5530, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2016:IITch

[2543] Anonymous. Issue information — TOC. Statistics in Medicine, 35(30): 5531, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farewell:2016:MGBb

[2544] Vern Farewell and Tony Johnson. Major Greenwood (1880–1949): the biography. Statistics in Medicine, 35(30):5533–5535, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Urach:2016:MAG

[2545] S. Urach and M. Posch. Multi-arm group sequential designs with a simultaneous stopping rule. Statistics in Medicine, 35(30):5536–5550, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kalia:2016:EIC

[2546] Sumeet Kalia, Neil Klar, and Allan Donner. On the estimation of intracluster correlation for time-to-event outcomes in cluster randomized trials. Statistics in Medicine, 35(30):5551–5560, December 30, 2016. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:2016:CCM

[2547] A. Lawrence Gould. Control charts for monitoring accumulating adverse event count frequencies from single and multiple blinded trials. Statistics in Medicine, 35(30):5561–5578, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:2016:SAM

[2548] Weihua Gao, Donald Hedeker, Robin Mermelstein, and Hui Xie. A scalable approach to measuring the impact of nonignorable nonresponse with an EMA application. Statistics in Medicine, 35(30):5579–5602, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mian:2016:EZI

[2549] Rajibul Mian and Sudhir Paul. Estimation for zero-inflated overdispersed count data model with missing response. Statistics in Medicine, 35(30):5603–5624, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jemielita:2016:IPC

[2550] Thomas Jemielita, Mary Putt, and Devan Mehrotra. Improved power in crossover designs through linear combinations of baselines. Statistics in Medicine, 35(30):5625–5641, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2016:VEW

[2551] Peter C. Austin. Variance estimation when using inverse probability of treatment weighting (IPTW) with survival analysis. Statistics in Medicine, 35(30):5642–5655, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:2016:MAR

[2552] Yi Pan and Huiman X. Barnhart. Methods for assessing the reliability of quality of life based on SF-36. Statistics in Medicine, 35(30):5656– 5665, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2016:BQR

[2553] Yangxin Huang and Jiaqing Chen. Bayesian quantile regression-based nonlinear mixed-effects joint models for time-to-event and longitudinal data with multiple features. Statistics in Medicine, 35(30):5666–5685, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nevo:2016:AME

[2554] Daniel Nevo, David M. Zucker, Rulla M. Tamimi, and Molin Wang. Accounting for measurement error in biomarker data and misclassification of subtypes in the analysis of tumor data. Statistics in Medicine, 35(30): 5686–5700, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yiu:2016:TMS

[2555] Sean Yiu, Brian D. M. Tom, and Vernon T. Farewell. Trivariate moverstayer counting process models for investigating joint damage in psoriatic arthritis. Statistics in Medicine, 35(30):5701–5716, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ramezan:2016:SPR

[2556] Reza Ramezan, Paul Marriott, and Shojaeddin Chenouri. Skellam process with resetting: a neural spike train model. Statistics in Medicine, 35 (30):5717–5729, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tamhane:2016:POR

[2557] Ashutosh R. Tamhane, Andrew O. Westfall, Greer A. Burkholder, and Gary R. Cutter. Prevalence odds ratio versus prevalence ratio: choice comes with consequences. Statistics in Medicine, 35(30):5730–5735, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Samworth:2016:HBD

[2558] Richard J. Samworth. Handbook of big data. Statistics in Medicine, 36(23):5736, December 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIT

[2559] Anonymous. Issue information — title page. Statistics in Medicine, 36 (1):1–2, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Day:2017:EEC

[2560] Simon Day. Editorial: Estimands in clinical trials — a challenge to intention to treat? Statistics in Medicine, 36(1):3–4, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Akacha:2017:ECT

[2561] Mouna Akacha, Frank Bretz, and Stephen Ruberg. Estimands in clinical trials — broadening the perspective. Statistics in Medicine, 36(1):5–19, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Permutt:2017:CEC

[2562] Thomas Permutt. Comments on ‘Estimands in clinical trials — broadening the perspective’. Statistics in Medicine, 36(1):20–21, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bauer:2017:CEC

[2563] Peter Bauer. Comments on ‘Estimands in clinical trials — broadening the perspective’. Statistics in Medicine, 36(1):22–23, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rockhold:2017:CEC

[2564] Frank Rockhold. Comments on ‘Estimands in clinical trials — broadening the perspective’. Statistics in Medicine, 36(1):24–26, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zang:2017:RTS

[2565] Yong Zang and J. Jack Lee. A robust two-stage design identifying the optimal biological dose for phase I/II clinical trials. Statistics in Medicine, 36(1):27–42, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cunanan:2017:BAP

[2566] Kristen M. Cunanan and Joseph S. Koopmeiners. A Bayesian adaptive phase I-II trial design for optimizing the schedule of therapeutic cancer vaccines. Statistics in Medicine, 36(1):43–53, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zang:2017:OSE

[2567] Yong Zang and Ying Yuan. Optimal sequential enrichment designs for phase II clinical trials. Statistics in Medicine, 36(1):54–66, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yin:2017:BDF

[2568] Jun Yin, Rui Qin, Monia Ezzalfani, Daniel J. Sargent, and Sumithra J. Mandrekar. A Bayesian dose-finding design incorporating toxicity data from multiple treatment cycles. Statistics in Medicine, 36(1):67–80, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Solomon:2017:PML

[2569] Ghideon Solomon and Lisa Weissfeld. Pseudo maximum likelihood approach for the analysis of multivariate left-censored longitudinal data. Statistics in Medicine, 36(1):81–91, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:2017:MPA

[2570] A. Lawrence Gould and William B. Wang. Monitoring potential adverse event rate differences using data from blinded trials: the canary in the coal mine. Statistics in Medicine, 36(1):92–104, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2017:SRM

[2571] Yong Chen, Yulun Liu, Haitao Chu, Mei-Ling Ting Lee, and Christopher H. Schmid. A simple and robust method for multivariate metaanalysis of diagnostic test accuracy. Statistics in Medicine, 36(1):105– 121, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ickowicz:2017:MHL

[2572] Adrien Ickowicz and Ross Sparks. Modelling hospital length of stay using convolutive mixtures distributions. Statistics in Medicine, 36(1): 122–135, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lipkovich:2017:TBD

[2573] Ilya Lipkovich, Alex Dmitrienko, and Ralph B. D’Agostino, Sr. Tutorial in biostatistics: data-driven subgroup identification and analysis in clinical trials. Statistics in Medicine, 36(1):136–196, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2017:CLR

[2574] Kung-Jong Lui. Comments on ‘Likelihood ratio and score tests to test the non-inferiority (or equivalence) of the odds ratio in a crossover study with binary outcomes’. Statistics in Medicine, 36(1):197–198, January 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIa

[2575] Anonymous. Issue information. Statistics in Medicine, 36(2):199–200, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Iasonos:2017:EPD

[2576] Alexia Iasonos and John O’Quigley. Early phase dose finding methodology. Statistics in Medicine, 36(2):201–203, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Iasonos:2017:SMP

[2577] Alexia Iasonos and John O’Quigley. Sequential monitoring of Phase I dose expansion cohorts. Statistics in Medicine, 36(2):204–214, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petroni:2017:IAM

[2578] Gina R. Petroni, Nolan A. Wages, Gautier Paux, and Frédéric Dubois. Implementation of adaptive methods in early-phase clinical trials. Statistics in Medicine, 36(2):215–224, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wheeler:2017:MSA

[2579] Graham M. Wheeler, Michael J. Sweeting, Adrian P. Mander, Shing M. Lee, and Ying Kuen K. Cheung. Modelling semi-attributable toxicity in dual-agent phase I trials with non-concurrent drug administration. Statistics in Medicine, 36(2):225–241, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wages:2017:IMT

[2580] Nolan A. Wages. Identifying a maximum tolerated contour in twodimensional dose finding. Statistics in Medicine, 36(2):242–253, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Conaway:2017:DPTa

[2581] Mark R. Conaway and Nolan A. Wages. Designs for phase I trials in ordered groups. Statistics in Medicine, 36(2):254–265, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:2017:DMS

[2582] Haitao Pan and Ying Yuan. A default method to specify skeletons for Bayesian model averaging continual reassessment method for phase I clinical trials. Statistics in Medicine, 36(2):266–279, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tighiouart:2017:BAD

[2583] Mourad Tighiouart, Quanlin Li, and André Rogatko. A Bayesian adaptive design for estimating the maximum tolerated dose curve using drug combinations in cancer phase I clinical trials. Statistics in Medicine, 36 (2):280–290, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Horton:2017:PTP

[2584] Bethany Jablonski Horton, Nolan A. Wages, and Mark R. Conaway. Performance of toxicity probability interval based designs in contrast to the continual reassessment method. Statistics in Medicine, 36(2):291– 300, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Partlett:2017:REM

[2585] Christopher Partlett and Richard D. Riley. Random effects metaanalysis: Coverage performance of 95% confidence and prediction intervals following REML estimation. Statistics in Medicine, 36(2):301–317, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morgan:2017:CAA

[2586] Katy E. Morgan, Andrew B. Forbes, Ruth H. Keogh, Vipul Jairath, and Brennan C. Kahan. Choosing appropriate analysis methods for cluster randomised cross-over trials with a binary outcome. Statistics in Medicine, 36(2):318–333, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mahiane:2017:SPI

[2587] Severin Guy Mahiané and Oliver Laeyendecker. Segmented polynomials for incidence rate estimation from prevalence data. Statistics in Medicine, 36(2):334–344, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chebon:2017:MZI

[2588] Sammy Chebon, Christel Faes, Frank Cools, and Helena Geys. Models for zero-inflated, correlated count data with extra heterogeneity: when is it too complex? Statistics in Medicine, 36(2):345–361, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boher:2017:GFT

[2589] Jean-Marie Boher, Thomas Filleron, Roch Giorgi, Andrew Kramar, and Richard J. Cook. Goodness-of-fit test for monotone proportional subdistribution hazards assumptions based on weighted residuals. Statistics in Medicine, 36(2):362–377, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bodnar:2017:BER

[2590] Olha Bodnar, Alfred Link, Barbora Arendacká, Antonio Possolo, and Clemens Elster. Bayesian estimation in random effects meta-analysis using a non-informative prior. Statistics in Medicine, 36(2):378–399, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lumley:2017:BRE

[2591] Thomas Lumley. Book review: Essentials of probability theory for statisticians . Michael A. Proschan and Pamela A. Shaw. Chapman and Hall/ CRC, Boca Rato, 2016. No. of pages: xvi + 328 pages. Price: 40.59 GBP. ISBN: 978-1-4987-0419-9 (hardcover). Statistics in Medicine, 36 (2):400, January 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIb

[2592] Anonymous. Issue information. Statistics in Medicine, 36(3):401–402, February 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2017:CAT

[2593] Zhiguo Li. Comparison of adaptive treatment strategies based on longitudinal outcomes in sequential multiple assignment randomized trials. Statistics in Medicine, 36(3):403–415, February 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2017:IDM

[2594] Fei Chen, Gang Li, and K. K. Gordon Lan. Inconsistency and dropminimum data analysis. Statistics in Medicine, 36(3):416–425, February 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hampson:2017:AIW

[2595] Lisa V. Hampson, Roland Fisch, Linh M. Van, and Thomas Jaki. Asymmetric inner wedge group sequential tests with applications to verifying whether effective drug concentrations are similar in adults and children. Statistics in Medicine, 36(3):426–441, February 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duc:2017:WAC

[2596] Anh Nguyen Duc and Marcel Wolbers. Weighted analysis of composite endpoints with simultaneous inference for flexible weight constraints.

Statistics in Medicine, 36(3):442–454, February 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salim:2017:CRP

[2597] Agus Salim, Bénédicte Delcoigne, Krystyn Villaflores, Woon-Puay Koh, Jian-Min Yuan, Rob M. van Dam, and Marie Reilly. Comparisons of risk prediction methods using nested case-control data. Statistics in Medicine, 36(3):455–465, February 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2017:UEC

[2598] Zhuoyu Wang, Nandini Dendukuri, and Lawrence Joseph. Understanding the effects of conditional dependence in research studies involving imperfect diagnostic tests. Statistics in Medicine, 36(3):466–480, February 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vonCube:2017:CCA

[2599] Maja von Cube, Martin Schumacher, Mercedes Palomar-Martinez, Pedro Olaechea-Astigarraga, Francisco Alvarez-Lerma, and Martin Wolkewitz. A case-cohort approach for multi-state models in hospital epidemiology. Statistics in Medicine, 36(3):481–495, February 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yiu:2017:JMA

[2600] Sean Yiu and Brian Tom. A joint modelling approach for multistate processes subject to resolution and under intermittent observations. Statistics in Medicine, 36(3):496–508, February 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2017:PSM

[2601] Yuan Huang, Jin Liu, Huangdi Yi, Ben-Chang Shia, and Shuangge Ma. Promoting similarity of model sparsity structures in integrative analysis of cancer genetic data. Statistics in Medicine, 36(3):509–559, February 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roberts:2017:C

[2602] Eric M. Roberts. Correction. Statistics in Medicine, 36(3):560, February 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIc

[2603] Anonymous. Issue information. Statistics in Medicine, 36(4):561–562, February 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jaki:2017:FVM

[2604] Thomas Jaki and Despina Vasileiou. Factorial versus multi-arm multistage designs for clinical trials with multiple treatments. Statistics in Medicine, 36(4):563–580, February 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2017:ISS

[2605] Song Zhang, Jing Cao, and Chul Ahn. Inference and sample size calculation for clinical trials with incomplete observations of paired binary outcomes. Statistics in Medicine, 36(4):581–591, February 20, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2017:DTC

[2606] Zhenzhen Xu, Boguang Zhen, Yongsoek Park, and Bin Zhu. Designing therapeutic cancer vaccine trials with delayed treatment effect. Statistics in Medicine, 36(4):592–605, February 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2017:MIP

[2607] Katherine J. Lee and John B. Carlin. Multiple imputation in the presence of non-normal data. Statistics in Medicine, 36(4):606–617, February 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koerner:2017:CPD

[2608] F. Spencer Koerner, John R. Anderson, Jon M. Fincham, and Robert E. Kass. Change-point detection of cognitive states across multiple trials in functional neuroimaging. Statistics in Medicine, 36(4):618–642, February 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2017:FMR

[2609] Xiaoqi Li, Yong Chen, and Ruosha Li. A frailty model for recurrent events during alternating restraint and non-restraint time periods. Statistics in Medicine, 36(4):643–654, February 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wan:2017:MSI

[2610] Yubing Wan, Susmita Datta, J. Jack Lee, and Maiying Kong. Monotonic single-index models to assess drug interactions. Statistics in Medicine, 36(4):655–670, February 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jo:2017:TUG

[2611] Booil Jo, Robert L. Findling, Chen-Pin Wang, Trevor J. Hastie, Eric A. Youngstrom, L. Eugene Arnold, Mary A. Fristad, and Sarah McCue Horwitz. Targeted use of growth mixture modeling: a learning perspective. Statistics in Medicine, 36(4):671–686, February 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hajage:2017:ECM

[2612] David Hajage, Yann De Rycke, Guillaume Chauvet, and Florence Tubach. Estimation of conditional and marginal odds ratios using the prognostic score. Statistics in Medicine, 36(4):687–716, February 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2017:CSA

[2613] Linbo Wang and Thomas Richardson. On the concordant survivorship assumption. Statistics in Medicine, 36(4):717–720, February 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IId

[2614] Anonymous. Issue information. Statistics in Medicine, 36(5):721–722, February 28, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holzhauer:2017:MAA

[2615] Björn Holzhauer. Meta-analysis of aggregate data on medical events. Statistics in Medicine, 36(5):723–737, February 28, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2017:ASC

[2616] Renke Zhou, Hong Zhu, Melissa Bondy, and Jing Ning. Analyzing semi-competing risks data with missing cause of informative terminal event. Statistics in Medicine, 36(5):738–753, February 28, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mason:2017:DBA

[2617] Alexina J. Mason, Juan Gonzalez-Maffe, Killian Quinn, Nicki Doyle, Ken Legg, Peter Norsworthy, Roy Trevelion, Alan Winston, and Deborah Ashby. Developing a Bayesian adaptive design for a phase I clinical trial: a case study for a novel HIV treatment. Statistics in Medicine, 36(5): 754–771, February 28, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hua:2017:OSI

[2618] Hairui Hua, Danielle L. Burke, Michael J. Crowther, Joie Ensor, Catrin Tudur Smith, and Richard D. Riley. One-stage individual participant data meta-analysis models: estimation of treatment-covariate interactions must avoid ecological bias by separating out within-trial and across-trial information. Statistics in Medicine, 36(5):772–789, February 28, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nicholson:2017:NSR

[2619] George Nicholson and Chris Holmes. A note on statistical repeatability and study design for high-throughput assays. Statistics in Medicine, 36 (5):790–798, February 28, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wellek:2017:SBA

[2620] Stefan Wellek. A U-statistics based approach to sample size planning of two-arm trials with discrete outcome criterion aiming to establish either superiority or noninferiority. Statistics in Medicine, 36(5):799–812, February 28, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2017:APM

[2621] Haijin He, Jingheng Cai, Xinyuan Song, and Liuquan Sun. Analysis of proportional mean residual life model with latent variables. Statistics in Medicine, 36(5):813–826, February 28, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2017:TVE

[2622] Songshan Yang, James A. Cranford, Jennifer M. Jester, Runze Li, Robert A. Zucker, and Anne Buu. A time-varying effect model for examining group differences in trajectories of zero-inflated count outcomes with applications in substance abuse research. Statistics in Medicine, 36 (5):827–837, February 28, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wan:2017:SSD

[2623] Fei Wan. Simulating survival data with predefined censoring rates for proportional hazards models. Statistics in Medicine, 36(5):838–854, February 28, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burke:2017:MAU

[2624] Danielle L. Burke, Joie Ensor, and Richard D. Riley. Meta-analysis using individual participant data: one-stage and two-stage approaches, and why they may differ. Statistics in Medicine, 36(5):855–875, February 28, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rigdon:2017:RCR

[2625] Joseph Rigdon, Wen Wei Loh, and Michael G. Hudgens. Response to comment on ‘randomization inference for treatment effects on a binary outcome’. Statistics in Medicine, 36(5):876–880, February 28, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIe

[2626] Anonymous. Issue information. Statistics in Medicine, 36(6):881–882, March 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mutze:2017:SPT

[2627] Tobias Mütze, Frank Konietschke, Axel Munk, and Tim Friede. A Studentized permutation test for three-arm trials in the ‘gold standard’ design. Statistics in Medicine, 36(6):883–898, March 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Galwey:2017:SCT

[2628] N. W. Galwey. Supplementation of a clinical trial by historical control data: is the prospect of dynamic borrowing an illusion? Statistics in Medicine, 36(6):899–916, March 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Noma:2017:QIE

[2629] Hisashi Noma, Shiro Tanaka, Shigeyuki Matsui, Andrea Cipriani, and Toshi A. Furukawa. Quantifying indirect evidence in network metaanalysis. Statistics in Medicine, 36(6):917–927, March 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2017:MHR

[2630] Peter C. Austin, Philippe Wagner, and Juan Merlo. The median hazard ratio: a useful measure of variance and general contextual effects in multilevel survival analysis. Statistics in Medicine, 36(6):928–938, March 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vandenberghe:2017:BPM

[2631] S. Vandenberghe, S. Vansteelandt, and T. Loeys. Boosting the precision of mediation analyses of randomised experiments through covariate adjustment. Statistics in Medicine, 36(6):939–957, March 15, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Westgate:2017:AVS

[2632] Philip M. Westgate and Woodrow W. Burchett. On the analysis of very small samples of Gaussian repeated measurements: an alternative approach. Statistics in Medicine, 36(6):958–970, March 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kunzmann:2017:PEV

[2633] K. Kunzmann and M. Kieser. Point estimation and p-values in phase II adaptive two-stage designs with a binary endpoint. Statistics in Medicine, 36(6):971–984, March 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2017:SID

[2634] Tsui-Shan Lu, Matthew P. Longnecker, and Haibo Zhou. Statistical inferences for data from studies conducted with an aggregated multivariate outcome-dependent sample design. Statistics in Medicine, 36(6):985–997, March 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ortega-Villa:2017:STV

[2635] Ana Maria Ortega-Villa, Inyoung Kim, and H. Kim. Semiparametric time varying coefficient model for matched case-crossover studies. Statistics in Medicine, 36(6):998–1013, March 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andridge:2017:IER

[2636] Rebecca Andridge, Anne-Michelle Noone, and Nadia Howlader. Imputing estrogen receptor (ER) status in a population-based cancer registry: a sensitivity analysis. Statistics in Medicine, 36(6):1014–1028, March 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2017:LBB

[2637] Hongqian Wu, Ying Zhang, and Jeffrey D. Long. Longitudinal betabinomial modeling using GEE for overdispersed binomial data. Statistics in Medicine, 36(6):1029–1040, March 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIf

[2638] Anonymous. Issue information. Statistics in Medicine, 36(7):1041–1042, March 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walwyn:2017:MAS

[2639] Rebecca Walwyn and Chris Roberts. Meta-analysis of standardised mean differences from randomised trials with treatment-related clustering associated with care providers. Statistics in Medicine, 36(7):1043–1067, March 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chiang:2017:ASS

[2640] Chieh Chiang and Chin-Fu Hsiao. An approach for sample size determination of average bioequivalence based on interval estimation. Statistics in Medicine, 36(7):1068–1082, March 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alonso:2017:ASP

[2641] Ariel Alonso, Wim Van der Elst, and Paul Meyvisch. Assessing a surrogate predictive value: a causal inference approach. Statistics in Medicine, 36(7):1083–1098, March 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2017:EPD

[2642] Ying Liu, Yuanjia Wang, Chaorui Huang, and Donglin Zeng. Estimating personalized diagnostic rules depending on individualized characteristics. Statistics in Medicine, 36(7):1099–1117, March 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2017:CDS

[2643] Junho Lee, Ronald E. Gangnon, and Jun Zhu. Cluster detection of spatial regression coefficients. Statistics in Medicine, 36(7):1118–1133, March 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grill:2017:CAI

[2644] Sonja Grill, Donna P. Ankerst, Mitchell H. Gail, Nilanjan Chatterjee, and Ruth M. Pfeiffer. Comparison of approaches for incorporating new information into existing risk prediction models. Statistics in Medicine, 36(7):1134–1156, March 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2017:EMA

[2645] Jie Zhou, Jiajia Zhang, and Wenbin Lu. An expectation maximization algorithm for fitting the generalized odds-rate model to interval censored data. Statistics in Medicine, 36(7):1157–1171, March 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wolsztynski:2017:SAT

[2646] E. Wolsztynski, F. O’Sullivan, J. O’Sullivan, and J. F. Eary. Statistical assessment of treatment response in a cancer patient based on pretherapy and post-therapy FDG-PET scans. Statistics in Medicine, 36 (7):1172–1200, March 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIg

[2647] Anonymous. Issue information. Statistics in Medicine, 36(8):1201–1202, April 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2017:ACR

[2648] Peter C. Austin and Jason P. Fine. Accounting for competing risks in randomized controlled trials: a review and recommendations for improvement. Statistics in Medicine, 36(8):1203–1209, April 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Efthimiou:2017:CRN

[2649] Orestis Efthimiou, Dimitris Mavridis, Thomas P. A. Debray, Myrto Samara, Mark Belger, George C. M. Siontis, Stefan Leucht, Georgia Salanti, and GetReal Work Package 4. Combining randomized and nonrandomized evidence in network meta-analysis. Statistics in Medicine, 36(8):1210–1226, April 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yelland:2017:SSC

[2650] Lisa N. Yelland, Thomas R. Sullivan, David J. Price, and Katherine J. Lee. Sample size calculations for randomised trials including both independent and paired data. Statistics in Medicine, 36(8):1227–1239, April 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zheng:2017:IVC

[2651] Cheng Zheng, Ran Dai, Parameswaran N. Hari, and Mei-Jie Zhang. Instrumental variable with competing risk model. Statistics in Medicine, 36(8):1240–1255, April 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rad:2017:ESO

[2652] N. Nazeri Rad and J. F. Lawless. Estimation of state occupancy probabilities in multistate models with dependent intermittent observation, with application to HIV viral rebounds. Statistics in Medicine, 36(8): 1256–1271, April 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wright:2017:USV

[2653] Marvin N. Wright, Theresa Dankowski, and Andreas Ziegler. Unbiased split variable selection for random survival forests using maximally selected rank statistics. Statistics in Medicine, 36(8):1272–1284, April 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2017:ETD

[2654] Shanshan Li. Estimating time-dependent ROC curves using data under prevalent sampling. Statistics in Medicine, 36(8):1285–1301, April 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anisimov:2017:IPC

[2655] Vladimir V. Anisimov, Wai Y. Yeung, and D. Stephen Coad. Imbalance properties of centre-stratified permuted-block and complete randomisation for several treatments in a clinical trial. Statistics in Medicine, 36(8): 1302–1318, April 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsai:2017:CCC

[2656] Miao-Yu Tsai. Concordance correlation coefficients estimated by generalized estimating equations and variance components for longitudinal repeated measurements. Statistics in Medicine, 36(8):1319–1333, April 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alosh:2017:TSC

[2657] Mohamed Alosh, Mohammad F. Huque, Frank Bretz, and Ralph B. D’Agostino, Sr. Tutorial on statistical considerations on subgroup analysis in confirmatory clinical trials. Statistics in Medicine, 36(8):1334–1360, April 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIh

[2658] Anonymous. Issue information. Statistics in Medicine, 36(9):1361–1362, April 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sugimoto:2017:SCT

[2659] Tomoyuki Sugimoto, Toshimitsu Hamasaki, Scott R. Evans, and Takashi Sozu. Sizing clinical trials when comparing bivariate time-to-event outcomes. Statistics in Medicine, 36(9):1363–1382, April 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Litwin:2017:TSBa

[2660] Samuel Litwin, Stanley Basickes, and Eric A. Ross. Two-sample binary phase 2 trials with low type I error and low sample size. Statistics in Medicine, 36(9):1383–1394, April 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sangnawakij:2017:SME

[2661] Patarawan Sangnawakij, Dankmar Böhning, Stephen Adams, Michael Stanton, and Heinz Holling. Statistical methodology for estimating the mean difference in a meta-analysis without study-specific variance information. Statistics in Medicine, 36(9):1395–1413, April 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2017:PSI

[2662] Xin Huang, Yan Sun, Paul Trow, Saptarshi Chatterjee, Arunava Chakravartty, Lu Tian, and Viswanath Devanarayan. Patient subgroup identification for clinical drug development. Statistics in Medicine, 36(9):

1414–1428, April 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lambert:2017:FPM

[2663] Paul C. Lambert, Sally R. Wilkes, and Michael J. Crowther. Flexible parametric modelling of the cause-specific cumulative incidence function. Statistics in Medicine, 36(9):1429–1446, April 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barrett:2017:DPU

[2664] Jessica Barrett and Li Su. Dynamic predictions using flexible joint models of longitudinal and time-to-event data. Statistics in Medicine, 36(9): 1447–1460, April 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Krall:2017:HMA

[2665] Jenna R. Krall, Amber J. Hackstadt, and Roger D. Peng. A hierarchical modeling approach to estimate regional acute health effects of particulate matter sources. Statistics in Medicine, 36(9):1461–1475, April 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2017:IAS

[2666] An-Min Tang, Nian-Sheng Tang, and Hongtu Zhu. Influence analysis for skew-normal semiparametric joint models of multivariate longitudinal and multivariate survival data. Statistics in Medicine, 36(9):1476–1490, April 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Henn:2017:DSC

[2667] Lisa L. Henn, John Hughes, Eleena Iisakka, Jutta Ellermann, Shabnam Mortazavi, Connor Ziegler, Mikko J. Nissi, and Patrick Morgan. Disease severity classification using quantitative magnetic resonance imaging data of cartilage in femoroacetabular impingement. Statistics in Medicine, 36(9):1491–1505, April 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2017:QBE

[2668] S. D. Walter, H. Han, M. Briel, and G. H. Guyatt. Quantifying the bias in the estimated treatment effect in randomized trials having interim analyses and a rule for early stopping for futility. Statistics in Medicine, 36(9):1506–1518, April 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kose:2017:LEC

[2669] T. Köse, M. Orman, F. Ikiz, J. Gallagher, and D. Böhning. Letter to the Editor: Correction on Extending the Lincoln–Petersen estimator for multiple identifications in one source. Statistics in Medicine 2014; 33:4237– 4249. Statistics in Medicine, 36(9):1519–1520, April 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1704].

Anonymous:2017:IIi

[2670] Anonymous. Issue information. Statistics in Medicine, 36(10):1521–1522, May 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2017:GAC

[2671] Wei Wang, Ying Ma, Yangxin Huang, and Henian Chen. Generalizability analysis for clinical trials: a simulation study. Statistics in Medicine, 36 (10):1523–1531, May 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2017:ERP

[2672] Lingling Li, Shijie Tang, and Liewen Jiang. On an enhanced rankpreserving structural failure time model to handle treatment switch, crossover, and dropout. Statistics in Medicine, 36(10):1532–1547, May 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2017:EPF

[2673] Yayuan Zhu, Jerald F. Lawless, and Cecilia A. Cotton. Estimation of parametric failure time distributions based on interval-censored data with irregular dependent follow-up. Statistics in Medicine, 36(10):1548– 1567, May 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cunanan:2017:EBT

[2674] Kristen M. Cunanan, Alexia Iasonos, Ronglai Shen, Colin B. Begg, and Mithat Gönen. An efficient basket trial design. Statistics in Medicine, 36(10):1568–1579, May 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stanley:2017:FPR

[2675] T. D. Stanley, Hristos Doucouliagos, and John P. A. Ioannidis. Finding the power to reduce publication bias. Statistics in Medicine, 36(10):

1580–1598, May 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cederkvist:2017:ITA

[2676] Luise Cederkvist, Klaus K. Holst, Klaus K. Andersen, David V. Glidden, Kirsten Frederiksen, Susanne K. Kjær, and Thomas H. Scheike. Incorporation of the time aspect into the liability-threshold model for case-control-family data. Statistics in Medicine, 36(10):1599–1618, May 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Son:2017:CGB

[2677] Junbo Son, Patricia Flatley Brennan, and Shiyu Zhou. Correlated gamma-based hidden Markov model for the smart asthma management based on rescue inhaler usage. Statistics in Medicine, 36(10):1619–1637, May 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thurman:2017:DSD

[2678] Andrew L. Thurman, Jiwoong Choi, Sanghun Choi, Ching-Long Lin, Eric A. Hoffman, Chang Hyun Lee, and Kung-Sik Chan. Detection of smoothly distributed spatial outliers, with applications to identifying the distribution of parenchymal hyperinflation following an airway challenge in asthmatics. Statistics in Medicine, 36(10):1638–1654, May 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longford:2017:VRN

[2679] Nicholas T. Longford. Variation of the rates of necrotising enterocolitis in the neonatal networks in England. Statistics in Medicine, 36(10): 1655–1668, May 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2017:RAM

[2680] Guanglei Yu, Liang Zhu, Yang Li, Jianguo Sun, and Leslie L. Robison. Regression analysis of mixed panel count data with dependent terminal events. Statistics in Medicine, 36(10):1669–1680, May 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIj

[2681] Anonymous. Issue information. Statistics in Medicine, 36(11):1681–1682, May 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2017:IER

[2682] Xue Yang and Yong Zhou. Improve efficiency and reduce bias of Cox regression models for two-stage randomization designs using auxiliary covariates. Statistics in Medicine, 36(11):1683–1695, May 20, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ginestet:2017:DRM

[2683] Cedric E. Ginestet, Richard Emsley, and Sabine Landau. Dose-response modeling in mental health using Stein-like estimators with instrumental variables. Statistics in Medicine, 36(11):1696–1714, May 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bakbergenuly:2017:BBM

[2684] Ilyas Bakbergenuly and Elena Kulinskaya. Beta-binomial model for meta-analysis of odds ratios. Statistics in Medicine, 36(11):1715–1734, May 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bryan:2017:BHM

[2685] Susan R. Bryan, Paul H. C. Eilers, Joost van Rosmalen, Dimitris Rizopoulos, Koenraad A. Vermeer, Hans G. Lemij, and Emmanuel M. E. H. Lesaffre. Bayesian hierarchical modeling of longitudinal glaucomatous visual fields using a two-stage approach. Statistics in Medicine, 36(11):1735–1753, May 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perera:2017:BSS

[2686] Yamuni Perera, Mingchen Ren, Joyce Raymond B. Punzalan, Christopher J. Rudnisky, and Alexander R. de Leon. Binocular sensitivity and specificity of screening tests in cross-sectional diagnostic studies of paired organs. Statistics in Medicine, 36(11):1754–1766, May 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parast:2017:ESM

[2687] Layla Parast, Tianxi Cai, and Lu Tian. Evaluating surrogate marker information using censored data. Statistics in Medicine, 36(11):1767– 1782, May 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bowden:2017:FIP

[2688] Jack Bowden, Fabiola Del Greco M., Cosetta Minelli, George Davey Smith, Nuala Sheehan, and John Thompson. A framework for the investigation of pleiotropy in two-sample summary data Mendelian randomization. Statistics in Medicine, 36(11):1783–1802, May 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cortese:2017:RMR

[2689] Giuliana Cortese, Stine A. Holmboe, and Thomas H. Scheike. Regression models for the restricted residual mean life for right-censored and lefttruncated data. Statistics in Medicine, 36(11):1803–1822, May 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Su:2017:PSS

[2690] Pei-Fang Su, Chia-Hua Chung, Yu-Wen Wang, Yunchan Chi, and Ying-Ju Chang. Power and sample size calculation for paired recurrent events data based on robust nonparametric tests. Statistics in Medicine, 36(11): 1823–1838, May 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robinson:2017:BRE

[2691] Andrew Robinson. Book review: Extending R. John Chambers, Chapman and Hall/CRC, Boca Raton, 2016. No. of pages: 364. Price: 44.99 GBP. ISBN: 978-1-4987-7571-7. Statistics in Medicine, 36(11):1839– 1840, May 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIk

[2692] Anonymous. Issue information. Statistics in Medicine, 36(12):1841–1842, May 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2017:ETE

[2693] Wen Li, Cong Chen, Xiaoyun Li, and Robert A. Beckman. Estimation of treatment effect in two-stage confirmatory oncology trials of personalized medicines. Statistics in Medicine, 36(12):1843–1861, May 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kondo:2017:ITR

[2694] Yumi Kondo, Yinshan Zhao, and John Petkau. Identification of treatment responders based on multiple longitudinal outcomes with applications to multiple sclerosis patients. Statistics in Medicine, 36(12): 1862–1883, May 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zheng:2017:GPL

[2695] Xueying Zheng, Guoyou Qin, and Dongsheng Tu. A generalized partially linear mean-covariance regression model for longitudinal proportional data, with applications to the analysis of quality of life data from cancer clinical trials. Statistics in Medicine, 36(12):1884–1894, May 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:2017:NCS

[2696] Yanqin Feng, Ran Duan, and Jianguo Sun. Nonparametric comparison of survival functions based on interval-censored data with unequal censoring. Statistics in Medicine, 36(12):1895–1906, May 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Novelo:2017:BAQ

[2697] Luis G. León Novelo, Andrew Womack, Hongxiao Zhu, and Xiaowei Wu. A Bayesian analysis of quantal bioassay experiments incorporating historical controls via Bayes factors. Statistics in Medicine, 36(12): 1907–1923, May 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rice:2017:SPL

[2698] John D. Rice and Alex Tsodikov. Semiparametric profile likelihood estimation for continuous outcomes with excess zeros in a random-threshold damage-resistance model. Statistics in Medicine, 36(12):1924–1935, May 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tahir:2017:CHS

[2699] M. Ramzan Tahir, Quang X. Tran, and Mikhail S. Nikulin. Comparison of hypertabastic survival model with other unimodal hazard rate functions using a goodness-of-fit test. Statistics in Medicine, 36(12): 1936–1945, May 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Franklin:2017:CPP

[2700] Jessica M. Franklin, Wesley Eddings, Peter C. Austin, Elizabeth A. Stuart, and Sebastian Schneeweiss. Comparing the performance of propensity score methods in healthcare database studies with rare outcomes. Statistics in Medicine, 36(12):1946–1963, May 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Azarang:2017:DMR

[2701] Leyla Azarang, Thomas Scheike, and Jacobo de Uña-Álvarez. Direct modeling of regression effects for transition probabilities in the progressive illness-death model. Statistics in Medicine, 36(12):1964–1976, May 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Im:2017:CES

[2702] Jongho Im, Eunyong Ahn, Namseon Beck, Jae Kwang Kim, and Taesung Park. Correlation estimation with singly truncated bivariate data. Statistics in Medicine, 36(12):1977–1988, May 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roberts:2017:WBS

[2703] T. Roberts, M. Newell, W. Auffermann, and B. Vidakovic. Wavelet-based scaling indices for breast cancer diagnostics. Statistics in Medicine, 36 (12):1989–2000, May 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIl

[2704] Anonymous. Issue information. Statistics in Medicine, 36(13):2001–2002, June 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rathnayake:2017:SMA

[2705] Lasitha N. Rathnayake and Pankaj K. Choudhary. Semiparametric modeling and analysis of longitudinal method comparison data. Statistics in Medicine, 36(13):2003–2015, June 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2017:TPM

[2706] Ying Wu and Richard J. Cook. A two-phase model for chronic disease processes under intermittent inspection. Statistics in Medicine, 36(13): 2016–2031, June 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karim:2017:EIP

[2707] Mohammad Ehsanul Karim, Robert W. Platt, and The Beams study group. Estimating inverse probability weights using super learner when weight-model specification is unknown in a marginal structural Cox model context. Statistics in Medicine, 36(13):2032–2047, June 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sanchez:2017:ELL

[2708] Brisa N. Sánchez, Sehee Kim, and Mary D. Sammel. Estimators for longitudinal latent exposure models: examining measurement model assumptions. Statistics in Medicine, 36(13):2048–2066, June 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shimura:2017:CCB

[2709] Masashi Shimura, Masahiko Gosho, and Akihiro Hirakawa. Comparison of conditional bias-adjusted estimators for interim analysis in clinical trials with survival data. Statistics in Medicine, 36(13):2067–2080, June 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2017:SCC

[2710] Wonji Kim, Dandi Qiao, Michael H. Cho, Soo Heon Kwak, Kyong Soo Park, Edwin K. Silverman, Pak Sham, and Sungho Won. Selecting cases and controls for DNA sequencing studies using family histories of disease. Statistics in Medicine, 36(13):2081–2099, June 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rowley:2017:LCM

[2711] M. Rowley, H. Garmo, M. Van Hemelrijck, W. Wulaningsih, B. Grundmark, B. Zethelius, N. Hammar, G. Walldius, M. Inoue, L. Holmberg, and A. C. C. Coolen. A latent class model for competing risks. Statistics in Medicine, 36(13):2100–2119, June 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

laCruz:2017:PPO

[2712] Rolando De la Cruz, Claudio Fuentes, Cristian Meza, Dae-Jin Lee, and Ana Arribas-Gil. Predicting pregnancy outcomes using longitudinal information: a penalized splines mixed-effects model approach. Statistics in Medicine, 36(13):2120–2134, June 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2017:CFR

[2713] Yongqiang Tang. Closed-form REML estimators and sample size determination for mixed effects models for repeated measures under monotone missingness. Statistics in Medicine, 36(13):2135–2147, June 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zawistowski:2017:CRA

[2714] Matthew Zawistowski, Jeremy B. Sussman, Timothy P. Hofer, Douglas Bentley, Rodney A. Hayward, and Wyndy L. Wiitala. Corrected ROC analysis for misclassified binary outcomes. Statistics in Medicine, 36(13): 2148–2160, June 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIm

[2715] Anonymous. Issue information. Statistics in Medicine, 36(14):2161–2162, June 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosner:2017:ERC

[2716] Bernard Rosner and Robert J. Glynn. Estimation of rank correlation for clustered data. Statistics in Medicine, 36(14):2163–2186, June 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dwivedi:2017:ASS

[2717] Alok Kumar Dwivedi, Indika Mallawaarachchi, and Luis A. Alvarado. Analysis of small sample size studies using nonparametric bootstrap test with pooled resampling method. Statistics in Medicine, 36(14):2187– 2205, June 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2017:SWC

[2718] You-Gan Wang and Liya Fu. Selection of working correlation structure in generalized estimating equations. Statistics in Medicine, 36(14):2206– 2219, June 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2017:ECV

[2719] Longhai Li, Cindy X. Feng, and Shi Qiu. Estimating cross-validatory predictive p-values with integrated importance sampling for disease mapping models. Statistics in Medicine, 36(14):2220–2236, June 30, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldsmith:2017:VSF

[2720] Jeff Goldsmith and Joseph E. Schwartz. Variable selection in the functional linear concurrent model. Statistics in Medicine, 36(14):2237–2250, June 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2017:BSV

[2721] Bo Cai and Dipankar Bandyopadhyay. Bayesian semiparametric variable selection with applications to periodontal data. Statistics in Medicine, 36(14):2251–2264, June 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vu:2017:REM

[2722] Duy Vu, Alessandro Lomi, Daniele Mascia, and Francesca Pallotti. Relational event models for longitudinal network data with an application to interhospital patient transfers. Statistics in Medicine, 36(14):2265–2287, June 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tissier:2017:SPA

[2723] Renaud Tissier, Roula Tsonaka, Simon P. Mooijaart, Eline Slagboom, and Jeanine J. Houwing-Duistermaat. Secondary phenotype analysis in ascertained family designs: application to the Leiden longevity study. Statistics in Medicine, 36(14):2288–2301, June 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Puhr:2017:FLR

[2724] Rainer Puhr, Georg Heinze, Mariana Nold, Lara Lusa, and Angelika Geroldinger. Firth’s logistic regression with rare events: accurate effect estimates and predictions? Statistics in Medicine, 36(14):2302–2317, June 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ormerod:2017:BRE

[2725] John T. Ormerod. Book review: Extending the linear model with R: generalized linear, mixed effects and nonparametric regression models, second edition by Julian J. Faraway, Chapman and Hall/CRC, Boca Raton, 2016. No. of pages: 399. Price: £63.99 (book + eBook); £44.79 (eBook). ISBN 978-1-4987-2096-0. Statistics in Medicine, 36(14):2318– 2319, June 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lunceford:2017:SWP

[2726] Jared K. Lunceford. Stratification and weighting via the propensity score in estimation of causal treatment effects: a comparative study. Statistics in Medicine, 36(14):2320, June 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIn

[2727] Anonymous. Issue information. Statistics in Medicine, 36(15):2321–2322, July 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Conaway:2017:DPTb

[2728] Mark R. Conaway. A design for phase I trials in completely or partially ordered groups. Statistics in Medicine, 36(15):2323–2332, July 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brentnall:2017:EET

[2729] Adam R. Brentnall, Peter Sasieni, and Jack Cuzick. Estimating efficacy in trials with selective crossover. Statistics in Medicine, 36(15):2333– 2346, July 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2017:CBC

[2730] Soyoung Kim and Ying Huang. Combining biomarkers for classification with covariate adjustment. Statistics in Medicine, 36(15):2347–2362, July 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2017:GFR

[2731] Yan Liu, Christopher McMahan, and Colin Gallagher. A general framework for the regression analysis of pooled biomarker assessments. Statistics in Medicine, 36(15):2363–2377, July 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gotte:2017:SBA

[2732] Heiko Götte, Marietta Kirchner, Martin Oliver Sailer, and Meinhard Kieser. Simulation-based adjustment after exploratory biomarker subgroup selection in phase II. Statistics in Medicine, 36(15):2378–2390, July 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sivaganesan:2017:SFB

[2733] Siva. Sivaganesan, Peter Müller, and Bin Huang. Subgroup finding via Bayesian additive regression trees. Statistics in Medicine, 36(15):2391– 2403, July 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Persson:2017:EMC

[2734] Emma Persson, Ingeborg Waernbaum, and Torbjörn Lind. Estimating marginal causal effects in a secondary analysis of case-control data. Statistics in Medicine, 36(15):2404–2419, July 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maruo:2017:IIM

[2735] K. Maruo, Y. Yamaguchi, H. Noma, and M. Gosho. Interpretable inference on the mixed effect model with the Box–Cox transformation. Statistics in Medicine, 36(15):2420–2434, July 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tayob:2017:SCS

[2736] Nabihah Tayob and Susan Murray. Statistical consequences of a successful lung allocation system — recovering information and reducing bias in models for urgency. Statistics in Medicine, 36(15):2435–2451, July 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2017:WWL

[2737] Xiaodong Luo, Junshan Qiu, Steven Bai, and Hong Tian. Weighted win loss approach for analyzing prioritized outcomes. Statistics in Medicine, 36(15):2452–2465, July 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Banerjee:2017:TST

[2738] Tathagata Banerjee, Gaurangadeb Chattopadhyay, and Kaustav Banerjee. Two-stage test of means of unordered pairs. Statistics in Medicine, 36(15):2466–2480, July 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIo

[2739] Anonymous. Issue information. Statistics in Medicine, 36(16):2481–2482, July 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuznetsova:2017:AET

[2740] Olga M. Kuznetsova and Victoria Plamadeala Johnson. Approaches to expanding the two-arm biased coin randomization to unequal allocation while preserving the unconditional allocation ratio. Statistics in Medicine, 36(16):2483–2498, July 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wheeler:2017:TDF

[2741] Graham M. Wheeler, Michael J. Sweeting, and Adrian P. Mander. Toxicity-dependent feasibility bounds for the escalation with overdose control approach in phase I cancer trials. Statistics in Medicine, 36(16): 2499–2513, July 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldstein:2017:SAR

[2742] Harvey Goldstein, Katie Harron, and Mario Cortina-Borja. A scaling approach to record linkage. Statistics in Medicine, 36(16):2514–2521, July 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McIntosh:2017:EBG

[2743] Avery I. McIntosh, Gheorghe Doros, Edward C. Jones-López, Mary Gaeddert, Helen E. Jenkins, Patricia Marques-Rodrigues, Jerrold J. Ellner, Reynaldo Dietze, and Laura F. White. Extensions to Bayesian generalized linear mixed effects models for household tuberculosis transmission. Statistics in Medicine, 36(16):2522–2532, July 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2017:IMM

[2744] I-Chen Chen and Philip M. Westgate. Improved methods for the marginal analysis of longitudinal data in the presence of time-dependent covariates. Statistics in Medicine, 36(16):2533–2546, July 20, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2017:DPC

[2745] Xiang Zhang, Justin B. Loda, and William H. Woodall. Dynamic probability control limits for risk-adjusted CUSUM charts based on multiresponses. Statistics in Medicine, 36(16):2547–2558, July 20, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koutras:2017:FFT

[2746] M. V. Koutras and F. S. Milienos. A flexible family of transformation cure rate models. Statistics in Medicine, 36(16):2559–2575, July 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Horton:2017:PBA

[2747] K. W. Horton, N. E. Carlson, G. K. Grunwald, M. J. Mulvahill, and A. J. Polotsky. A population-based approach to analyzing pulses in time series of hormone data. Statistics in Medicine, 36(16):2576–2589, July 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OBrien:2017:MML

[2748] Robert M. O’Brien. Mixed models, linear dependency, and identification in age-period-cohort models. Statistics in Medicine, 36(16):2590–2600, July 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2017:CPS

[2749] Xu Tang and Fah F. Gan. Comparing performance of surgeons using risk-adjusted procedures. Statistics in Medicine, 36(16):2601–2613, July 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2017:RBI

[2750] Tao Lu. Retracted: Bayesian inference on mixed-effects location scale models with skew-t distribution and mismeasured covariates for longitudinal data. Statistics in Medicine, 36(16):2614–2629, July 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See retraction [3185].

Nevalainen:2017:TIC

[2751] Jaakko Nevalainen, Hannu Oja, and Somnath Datta. Tests for informative cluster size using a novel balanced bootstrap scheme. Statistics in Medicine, 37(25):2630–2640, July 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIp

[2752] Anonymous. Issue information. Statistics in Medicine, 36(17):2641–2642, July 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heritier:2017:AVA

[2753] Stephane Heritier, Chris J. Lloyd, and Serigne N. Lô. Accurate p-values for adaptive designs with binary endpoints. Statistics in Medicine, 36 (17):2643–2655, July 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ruckbeil:2017:AIS

[2754] Marcia Viviane Rückbeil, Ralf-Dieter Hilgers, and Nicole Heussen. Assessing the impact of selection bias on test decisions in trials with a time-to-event outcome. Statistics in Medicine, 36(17):2656–2668, July 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2017:CIS

[2755] Per K. Andersen, Elisavet Syriopoulou, and Erik T. Parner. Causal inference in survival analysis using pseudo-observations. Statistics in Medicine, 36(17):2669–2681, July 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2017:SET

[2756] Jiajun Xu, K. F. Lam, Feng Chen, Paul Milligan, and Yin Bun Cheung. Semiparametric estimation of time-varying intervention effects using recurrent event data. Statistics in Medicine, 36(17):2682–2696, July 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sternberg:2017:MIE

[2757] Maya Sternberg. Multiple imputation to evaluate the impact of an assay change in national surveys. Statistics in Medicine, 36(17):2697–2719, July 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Budgeon:2017:CLD

[2758] C. A. Budgeon, K. Murray, B. A. Turlach, S. Baker, V. L. Villemagne, S. C. Burnham, and for the Alzheimer’s Disease Neuroimaging Initiative. Constructing longitudinal disease progression curves using sparse, shortterm individual data with an application to Alzheimer’s disease. Statistics in Medicine, 36(17):2720–2734, July 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hazelton:2017:TCS

[2759] Martin L. Hazelton. Testing for changes in spatial relative risk. Statistics in Medicine, 36(17):2735–2749, July 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldstein:2017:CRP

[2760] Benjamin A. Goldstein, Gina Maria Pomann, Wolfgang C. Winkelmayer, and Michael J. Pencina. A comparison of risk prediction methods using repeated observations: an application to electronic health records for hemodialysis. Statistics in Medicine, 36(17):2750–2763, July 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Govindarajulu:2017:LCEa

[2761] Usha S. Govindarajulu, Marco Stillo, David Goldfarb, Michael E. Matheny, and Frederic S. Resnic. Learning curve estimation in medical devices and procedures: hierarchical modeling. Statistics in Medicine, 36(17):2764–2785, July 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parveen:2017:CCD

[2762] Nabila Parveen, Erica Moodie, and Bluma Brenner. Correcting covariate-dependent measurement error with non-zero mean. Statistics in Medicine, 36(27):2786–2800, July 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIq

[2763] Anonymous. Issue information. Statistics in Medicine, 36(18):2801–2802, August 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barrett:2017:IAC

[2764] James E. Barrett. Information-adaptive clinical trials with selective recruitment and binary outcomes. Statistics in Medicine, 36(18):2803– 2813, August 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dakin:2017:EEF

[2765] Helen Dakin and Alastair Gray. Economic evaluation of factorial randomised controlled trials: challenges, methods and recommendations. Statistics in Medicine, 36(18):2814–2830, August 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2017:UPT

[2766] Rui Wang and Victor De Gruttola. The use of permutation tests for the analysis of parallel and stepped-wedge cluster-randomized trials. Statistics in Medicine, 36(18):2831–2843, August 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coolen-Maturi:2017:PIB

[2767] Tahani Coolen-Maturi. Predictive inference for best linear combination of biomarkers subject to limits of detection. Statistics in Medicine, 36 (18):2844–2874, August 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cabras:2017:VCM

[2768] Stefano Cabras and Maria Eugenia Castellanos. P -value calibration in multiple hypotheses testing. Statistics in Medicine, 36(18):2875–2886, August 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McCandless:2017:CBM

[2769] Lawrence C. McCandless and Paul Gustafson. A comparison of Bayesian and Monte Carlo sensitivity analysis for unmeasured confounding. Statistics in Medicine, 36(18):2887–2901, August 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Caimo:2017:BER

[2770] Alberto Caimo, Francesca Pallotti, and Alessandro Lomi. Bayesian exponential random graph modelling of interhospital patient referral networks. Statistics in Medicine, 36(18):2902–2920, August 15, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duc:2017:SSN

[2771] Anh Nguyen Duc and Marcel Wolbers. Smooth semi-nonparametric (SNP) estimation of the cumulative incidence function. Statistics in Medicine, 36(18):2921–2934, August 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2017:EED

[2772] Yuchen Yang, Brent J. Shelton, Thomas T. Tucker, Li Li, Richard Kryscio, and Li Chen. Estimation of exposure distribution adjusting for association between exposure level and detection limit. Statistics in Medicine, 36(18):2935–2946, August 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2017:EEO

[2773] Zhichao Sun, Bhramar Mukherjee, Jason P. Estes, Pantel S. Vokonas, and Sung Kyun Park. Exposure enriched outcome dependent designs for longitudinal studies of gene-environment interaction. Statistics in Medicine, 36(18):2947–2960, August 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIr

[2774] Anonymous. Issue information. Statistics in Medicine, 36(19):2961–2962, August 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tanaka:2017:VSE

[2775] Shiro Tanaka, Yutaka Matsuyama, and Yasuo Ohashi. Validation of surrogate endpoints in cancer clinical trials via principal stratification with an application to a prostate cancer trial. Statistics in Medicine, 36 (19):2963–2977, August 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zheng:2017:ABI

[2776] Jiayin Zheng, Shein-Chung Chow, and Mengdie Yuan. On assessing bioequivalence and interchangeability between generics based on indirect comparisons. Statistics in Medicine, 36(19):2978–2993, August 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:2017:PMM

[2777] Fei Gao, Jun Dong, Donglin Zeng, Alan Rong, and Joseph G. Ibrahim. Pattern mixture models for clinical validation of biomarkers in the presence of missing data. Statistics in Medicine, 36(19):2994–3004, August 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2017:NIM

[2778] Joanna J. J. Wang, Mark Bartlett, and Louise Ryan. Non-ignorable missingness in logistic regression. Statistics in Medicine, 36(19):3005– 3021, August 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghebremichael-Weldeselassie:2017:SBS

[2779] Yonas Ghebremichael-Weldeselassie, Heather J. Whitaker, and C. Paddy Farrington. Spline-based self-controlled case series method. Statistics in Medicine, 36(19):3022–3038, August 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2017:BBM

[2780] Jingyi Guo, Andrea Riebler, and Håvard Rue. Bayesian bivariate metaanalysis of diagnostic test studies with interpretable priors. Statistics in Medicine, 36(19):3039–3058, August 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2017:ICI

[2781] Hyune-Ju Kim, Jun Luo, Huann-Sheng Chen, Don Green, Dennis Buckman, Jeffrey Byrne, and Eric J. Feuer. Improved confidence interval for average annual percent change in trend analysis. Statistics in Medicine, 36(19):3059–3074, August 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qian:2017:STG

[2782] Jing Qian, Sara Nunez, Soohyun Kim, Muredach P. Reilly, and Andrea S. Foulkes. A score test for genetic class-level association with nonlinear biomarker trajectories. Statistics in Medicine, 36(19):3075–3091, August 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hossain:2017:MBO

[2783] Anower Hossain, Karla DiazOrdaz, and Jonathan W. Bartlett. Missing binary outcomes under covariate-dependent missingness in cluster randomised trials. Statistics in Medicine, 36(19):3092–3109, August 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2017:ACL

[2784] Fei He, Armando Teixeira-Pinto, and Jaroslaw Harezlak. Autoregressive and cross-lagged model for bivariate non-commensurate outcomes. Statistics in Medicine, 36(19):3110–3120, August 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIs

[2785] Anonymous. Issue information. Statistics in Medicine, 36(20):3121–3122, September 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wellek:2017:SSP

[2786] Stefan Wellek. Sample size planning of two-arm superiority and noninferiority survival studies with discrete follow-up. Statistics in Medicine, 36(20):3123–3136, September 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bruckner:2017:EMA

[2787] Matthias Brückner, Andrew Titman, and Thomas Jaki. Estimation in multi-arm two-stage trials with treatment selection and time-to-event endpoint. Statistics in Medicine, 36(20):3137–3153, September 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

daSilva:2017:BAM

[2788] Hélio Doyle Pereira da Silva, Carlos Ascaso, Alessandra Queiroga Gonçalves, Patricia Puccinelli Orlandi, and Rosa Abellana. A Bayesian approach to model the conditional correlation between several diagnostic tests and various replicated subjects measurements. Statistics in Medicine, 36(20):3154–3170, September 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Flandre:2017:APA

[2789] Philippe Flandre, Reena Deutsch, and John O’Quigley. Accuracy of predictive ability measures for survival models. Statistics in Medicine, 36(20):3171–3180, September 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelson:2017:AIR

[2790] Kerrie P. Nelson, Aya A. Mitani, and Don Edwards. Assessing the influence of rater and subject characteristics on measures of agreement for ordinal ratings. Statistics in Medicine, 36(20):3181–3199, September 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Allen:2017:PGM

[2791] Chelsea McCarty Allen, Sandra D. Griffith, Saul Shiffman, and Daniel F. Heitjan. Proximity and gravity: modeling heaped self-reports. Statistics in Medicine, 36(20):3200–3215, September 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Malesios:2017:BEM

[2792] Chrisovalantis Malesios, Nikolaos Demiris, Konstantinos Kalogeropoulos, and Ioannis Ntzoufras. Bayesian epidemic models for spatially aggregated count data. Statistics in Medicine, 36(20):3216–3230, September 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yi:2017:APD

[2793] Grace Y. Yi, Wenqing He, and Feng He. Analysis of panel data under hidden mover-stayer models. Statistics in Medicine, 36(20):3231–3243, September 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2017:MLT

[2794] Jue Wang and Sheng Luo. Multidimensional latent trait linear mixed model: an application in clinical studies with multivariate longitudinal outcomes. Statistics in Medicine, 36(20):3244–3256, September 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2017:IAT

[2795] Peter C. Austin and Juan Merlo. Intermediate and advanced topics in multilevel logistic regression analysis. Statistics in Medicine, 36(20): 3257–3277, September 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:2017:CCB

[2796] Sander Greenland. A commentary on ‘A comparison of Bayesian and Monte Carlo sensitivity analysis for unmeasured confounding’. Statistics in Medicine, 36(20):3278–3280, September 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIu

[2797] Anonymous. Issue information. Statistics in Medicine, 36(21):3281–3282, September 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Willis:2017:MSV

[2798] Brian H. Willis and Richard D. Riley. Measuring the statistical validity of summary meta-analysis and meta-regression results for use in clinical practice. Statistics in Medicine, 36(21):3283–3301, September 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2017:BBO

[2799] Heng Zhou, J. Jack Lee, and Ying Yuan. BOP2: Bayesian optimal design for phase II clinical trials with simple and complex endpoints. Statistics in Medicine, 36(21):3302–3314, September 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McNamee:2017:HSB

[2800] Roseanne McNamee. How serious is bias in effect estimation in randomised trials with survival data given risk heterogeneity and informative censoring? Statistics in Medicine, 36(21):3315–3333, September 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demler:2017:ADN

[2801] Olga V. Demler, Michael J. Pencina, Nancy R. Cook, and Ralph B. D’Agostino, Sr. Asymptotic distribution of ΔAUC, NRIs, and IDI based on theory of U-statistics. Statistics in Medicine, 36(21):3334– 3360, September 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fukaya:2017:FMB

[2802] Keiichi Fukaya, Ai Kawamori, Yutaka Osada, Masumi Kitazawa, and Makio Ishiguro. The forecasting of menstruation based on a state-space modeling of basal body temperature time series. Statistics in Medicine, 36(21):3361–3379, September 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2017:CMQ

[2803] Chi-Chuan Yang, Yi-Hau Chen, and Hsing-Yi Chang. Composite marginal quantile regression analysis for longitudinal adolescent body mass index data. Statistics in Medicine, 36(21):3380–3397, September 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2017:SRA

[2804] Chyong-Mei Chen and Pao sheng Shen. Semiparametric regression analysis of failure time data with dependent interval censoring. Statistics in Medicine, 36(21):3398–3411, September 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Witte:2017:EAN

[2805] Birgit I. Witte, Johannes Berkhof, and Marianne A. Jonker. An EM algorithm for nonparametric estimation of the cumulative incidence function from repeated imperfect test results. Statistics in Medicine, 36(21):3412– 3421, September 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagashima:2017:ICF

[2806] Kengo Nagashima and Yasunori Sato. Information criteria for Firth’s penalized partial likelihood approach in Cox regression models. Statistics in Medicine, 36(21):3422–3436, September 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shan:2017:CTS

[2807] Guogen Shan. Comments on ‘Two-sample binary phase 2 trials with low type I error and low sample size’. Statistics in Medicine, 36(21):3437– 3438, September 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Litwin:2017:TSBb

[2808] Samuel Litwin, Eric Ross, and Stanley Basickes. Two-sample binary phase 2 trials with low type I error and low sample size. Statistics in Medicine, 36(21):3439, September 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gabriel:2017:CCB

[2809] Erin E. Gabriel, Michael C. Sachs, and Peter B. Gilbert. Comparing and combining biomarkers as principal surrogates for time-to-event clinical endpoints. Statistics in Medicine, 36(21):3440, September 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIv

[2810] Anonymous. Issue information. Statistics in Medicine, 36(22):3441–3442, September 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Held:2017:PFI

[2811] Leonhard Held, Sebastian Meyer, and Johannes Bracher. Probabilistic forecasting in infectious disease epidemiology: the 13th Armitage Lecture. Statistics in Medicine, 36(22):3443–3460, September 30, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chung:2017:SFB

[2812] Dongjun Chung, Andrew Lawson, and W. Jim Zheng. A statistical framework for biomedical literature mining. Statistics in Medicine, 36 (22):3461–3474, September 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diao:2017:MEC

[2813] Guoqing Diao, Donglin Zeng, Kuolung Hu, and Joseph G. Ibrahim. Modeling event count data in the presence of informative dropout with application to bleeding and transfusion events in myelodysplastic syndrome. Statistics in Medicine, 36(22):3475–3494, September 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2017:DPM

[2814] Jiehuan Sun, Jose D. Herazo-Maya, Naftali Kaminski, Hongyu Zhao, and Joshua L. Warren. A Dirichlet process mixture model for clustering longitudinal gene expression data. Statistics in Medicine, 36(22):3495– 3506, September 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kunkel:2017:CEM

[2815] Deborah Kunkel and Eloise E. Kaizar. A comparison of existing methods for multiple imputation in individual participant data meta-analysis.

Statistics in Medicine, 36(22):3507–3532, September 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2017:EGR

[2816] Annie J. Lee, Karen Marder, Roy N. Alcalay, Helen Mejia-Santana, Avi Orr-Urtreger, Nir Giladi, Susan Bressman, and Yuanjia Wang. Estimation of genetic risk function with covariates in the presence of missing genotypes. Statistics in Medicine, 36(22):3533–3546, September 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2017:DGE

[2817] Gwangsu Kim, Chao-Qiang Lai, Donna K. Arnett, Laurence D. Parnell, Jose M. Ordovas, Yongdai Kim, and Joungyoun Kim. Detection of gene-environment interactions in a family-based population using SCAD. Statistics in Medicine, 36(22):3547–3559, September 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2017:FJM

[2818] Kan Li and Sheng Luo. Functional joint model for longitudinal and time-to-event data: an application to Alzheimer’s disease. Statistics in Medicine, 37(3):3560–3572, September 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2017:LYL

[2819] Per Kragh Andersen. Life years lost among patients with a given disease. Statistics in Medicine, 37(3):3573–3582, September 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheung:2017:MMU

[2820] Li C. Cheung, Qing Pan, Noorie Hyun, Mark Schiffman, Barbara Fetterman, Philip E. Castle, Thomas Lorey, and Hormuzd A. Katki. Mixture models for undiagnosed prevalent disease and interval-censored incident disease: applications to a cohort assembled from electronic health records. Statistics in Medicine, 37(3):3583–3595, September 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lunardon:2017:CSS

[2821] N. Lunardon and D. Scharfstein. Comment on ‘small sample GEE estimation of regression parameters for longitudinal data’. Statistics in Medicine, 37(3):3596–3600, September 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIw

[2822] Anonymous. Issue information. Statistics in Medicine, 36(23):3601–3602, October 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kostoulas:2017:RGD

[2823] Polychronis Kostoulas, Søren S. Nielsen, Adam J. Branscum, Wesley O. Johnson, Nandini Dendukuri, Navneet K. Dhand, Nils Toft, and Ian A. Gardner. Reporting guidelines for diagnostic accuracy studies that use Bayesian latent class models (STARD-BLCM). Statistics in Medicine, 36 (23):3603–3604, October 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Minois:2017:UPG

[2824] Nathan Minois, Valérie Lauwers-Cances, Stéphanie Savy, Michel Attal, Sandrine Andrieu, Vladimir Anisimov, and Nicolas Savy. Using Poissongamma model to evaluate the duration of recruitment process when historical trials are available. Statistics in Medicine, 36(23):3605–3620, October 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tipton:2017:FMA

[2825] Elizabeth Tipton and Jonathan Shuster. A framework for the metaanalysis of Bland–Altman studies based on a limits of agreement approach. Statistics in Medicine, 36(23):3621–3635, October 15, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mutze:2017:BSS

[2826] Tobias Mütze and Tim Friede. Blinded sample size re-estimation in three-arm trials with ‘gold standard’ design. Statistics in Medicine, 36 (23):3636–3653, October 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troendle:2017:HCU

[2827] James Troendle, Eric Leifer, Zhiwei Zhang, Song Yang, and Heather Tewes. How to control for unmeasured confounding in an observational time-to-event study with exposure incidence information: the treatment choice Cox model. Statistics in Medicine, 36(23):3654–3669, October 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:2017:BIM

[2828] Jennifer A. Thompson, Katherine L. Fielding, Calum Davey, Alexander M. Aiken, James R. Hargreaves, and Richard J. Hayes. Bias and inference from misspecified mixed-effect models in stepped wedge trial analysis. Statistics in Medicine, 36(23):3670–3682, October 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bakoyannis:2017:SRC

[2829] Giorgos Bakoyannis, Menggang Yu, and Constantin T. Yiannoutsos. Semiparametric regression on cumulative incidence function with interval-censored competing risks data. Statistics in Medicine, 36(23): 3683–3707, October 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Watjou:2017:SSA

[2830] K. Watjou, C. Faes, A. Lawson, R. S. Kirby, M. Aregay, R. Carroll, and Y. Vandendijck. Spatial small area smoothing models for handling survey data with nonresponse. Statistics in Medicine, 36(23):3708–3745, October 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mauff:2017:EAS

[2831] Katya Mauff, Ewout W. Steyerberg, Giel Nijpels, Amber A. W. A. van der Heijden, and Dimitris Rizopoulos. Extension of the association structure in joint models to include weighted cumulative effects. Statistics in Medicine, 36(23):3746–3759, October 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shortreed:2017:MIS

[2832] Susan M. Shortreed, Eric Laber, T. Scott Stroup, Joelle Pineau, and Susan A. Murphy. A multiple imputation strategy for sequential multiple assignment randomized trials. Statistics in Medicine, 36(23):3760, October 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tamhane:2017:POR

[2833] Ashutosh R. Tamhane, Andrew O. Westfall, Greer A. Burkholder, and Gary R. Cutter. Prevalence odds ratio versus prevalence ratio: choice comes with consequences. Statistics in Medicine, 33(24):3760, October 15, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIx

[2834] Anonymous. Issue information. Statistics in Medicine, 36(24):3761–3762, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simons-Morton:2017:DSA

[2835] Bruce Simons-Morton. Driving in search of analyses. Statistics in Medicine, 36(24):3763–3771, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:2017:SWD

[2836] J. N. S. Matthews and A. B. Forbes. Stepped wedge designs: insights from a design of experiments perspective. Statistics in Medicine, 36 (24):3772–3790, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2017:ECR

[2837] Fan Li, Elizabeth L. Turner, Patrick J. Heagerty, David M. Murray, William M. Vollmer, and Elizabeth R. DeLong. An evaluation of constrained randomization for the design and analysis of group-randomized trials with binary outcomes. Statistics in Medicine, 36(24):3791–3806, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diaz:2017:DRI

[2838] Iván Díaz and Mark J. van der Laan. Doubly robust inference for targeted minimum loss-based estimation in randomized trials with missing outcome data. Statistics in Medicine, 36(24):3807–3819, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2017:VSJ

[2839] Yuqi Chen and Yuedong Wang. Variable selection for joint models of multivariate longitudinal measurements and event time data. Statistics in Medicine, 36(24):3820–3829, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bantis:2017:ESR

[2840] Leonidas E. Bantis, Qingxiang Yan, John V. Tsimikas, and Ziding Feng. Estimation of smooth ROC curves for biomarkers with limits of detection. Statistics in Medicine, 36(24):3830–3843, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dosne:2017:MAR

[2841] A. G. Dosne, M. Bergstrand, M. O. Karlsson, D. Renard, and G. Heimann. Model averaging for robust assessment of QT prolongation by concentration-response analysis. Statistics in Medicine, 36(24):3844– 3857, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hughes:2017:DCU

[2842] David M. Hughes, Arnost Komárek, Laura J. Bonnett, Gabriela Czanner, and Marta García-Fiñana. Dynamic classification using credible intervals in longitudinal discriminant analysis. Statistics in Medicine, 36 (24):3858–3874, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:2017:SBM

[2843] Zheng Wei and Daeyoung Kim. Subcopula-based measure of asymmetric association for contingency tables. Statistics in Medicine, 36(24):3875– 3894, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Estes:2017:MAG

[2844] Jason P. Estes, John D. Rice, Shi Li, Heather M. Stringham, Michael Boehnke, and Bhramar Mukherjee. Meta-analysis of gene-environment interaction exploiting gene-environment independence across multiple case-control studies. Statistics in Medicine, 36(24):3895–3909, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shuster:2017:LCC

[2845] Jonathan J. Shuster. Linear combinations come alive in crossover designs. Statistics in Medicine, 36(24):3910–3918, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Olivier:2017:BRA

[2846] Jake Olivier. Book review: Applied surrogate endpoint evaluation methods with SAS and R. Ariel Alonso, Theophile Bigirumurame, Tomasz Burzykowski, Marc Buyse, Geert Molenberghs, Leacky Muchene, Nolen Joy Perualila, Ziv Shkedy, Wim Van der Elst, Chapman and Hall/CRC Biostatics Series, 2017, xxi + 373, 57.99 GBP, ISBN 13: 978-1-4822-4936-1 (Hardback). Statistics in Medicine, 36(24):3919–3920, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIy

[2847] Anonymous. Issue information. Statistics in Medicine, 36(25):3921–3922, November 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2017:HKM

[2848] Dan Jackson, Martin Law, Gerta Rücker, and Guido Schwarzer. The Hartung–Knapp modification for random-effects meta-analysis: a useful refinement but are there any residual concerns? Statistics in Medicine, 36(25):3923–3934, November 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kunzmann:2017:PEA

[2849] Kevin Kunzmann, Laura Benner, and Meinhard Kieser. Point estimation in adaptive enrichment designs. Statistics in Medicine, 36(25):3935–3947, November 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2017:IEC

[2850] Wei Li and Xiao-Hua Zhou. Identifiability and estimation of causal mediation effects with missing data. Statistics in Medicine, 36(25):3948–3965, November 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chiba:2017:SNB

[2851] Yasutaka Chiba. Sharp nonparametric bounds and randomization inference for treatment effects on an ordinal outcome. Statistics in Medicine, 36(25):3966–3975, November 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2017:SBA

[2852] Zhaonan Li, Xinyi Xu, and Junshan Shen. Semiparametric Bayesian analysis of accelerated failure time models with cluster structures. Statistics in Medicine, 36(25):3976–3989, November 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carles:2017:JBW

[2853] Sophie Carles, Marie-Aline Charles, Barbara Heude, Ismaïl Ahmed, Jérémie Botton, and the Eden mother-child study group. Joint Bayesian weight and height postnatal growth model to study the effects of maternal smoking during pregnancy. Statistics in Medicine, 36(25):3990–4006, November 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guhaniyogi:2017:BNA

[2854] Rajarshi Guhaniyogi. Bayesian nonparametric areal wombling for smallscale maps with an application to urinary bladder cancer data from Connecticut. Statistics in Medicine, 36(25):4007–4027, November 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2017:JME

[2855] Haocheng Li, Yukun Zhang, Raymond J. Carroll, Sarah Kozey Keadle, Joshua N. Sampson, and Charles E. Matthews. A joint modeling and estimation method for multivariate longitudinal data with mixed types of responses to analyze physical activity data generated by accelerometers. Statistics in Medicine, 36(25):4028–4040, November 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2017:CCI

[2856] Xiaoxia Han, Yilong Zhang, and Yongzhao Shao. On comparing 2 correlated C indices with censored survival data. Statistics in Medicine, 36 (25):4041–4049, November 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zabor:2017:CSM

[2857] Emily C. Zabor and Colin B. Begg. A comparison of statistical methods for the study of etiologic heterogeneity. Statistics in Medicine, 36(25): 4050–4060, November 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jeyarajah:2017:ELB

[2858] Jenny Jeyarajah and Gengsheng Qin. Empirical likelihood-based confidence intervals for mean medical cost with censored data. Statistics in Medicine, 36(25):4061–4070, November 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spieker:2017:MAB

[2859] Andrew J. Spieker and Ying Huang. A method to address betweensubject heterogeneity for identification of principal surrogate markers in repeated low-dose challenge HIV vaccine studies. Statistics in Medicine, 36(25):4071–4080, November 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIz

[2860] Anonymous. Issue information. Statistics in Medicine, 36(26):4081–4082, November 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simon:2017:IMA

[2861] Richard Simon and Noah Simon. Inference for multimarker adaptive enrichment trials. Statistics in Medicine, 36(26):4083–4093, November 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fiero:2017:PMM

[2862] Mallorie H. Fiero, Chiu-Hsieh Hsu, and Melanie L. Bell. A patternmixture model approach for handling missing continuous outcome data in longitudinal cluster randomized trials. Statistics in Medicine, 36(26): 4094–4105, November 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2017:SST

[2863] Ruitao Lin and Guosheng Yin. STEIN: a simple toxicity and efficacy interval design for seamless phase I/II clinical trials. Statistics in Medicine, 36(26):4106–4120, November 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Phadnis:2017:CTD

[2864] Milind A. Phadnis, James B. Wetmore, and Matthew S. Mayo. A clinical trial design using the concept of proportional time using the generalized gamma ratio distribution. Statistics in Medicine, 36(26):4121–4140, November 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yin:2017:SMC

[2865] Lixuan Yin, Guoqing Diao, and Aiyi Liu. A semiparametric method for comparing the discriminatory ability of biomarkers subject to limit of detection. Statistics in Medicine, 36(26):4141–4152, November 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2017:MAS

[2866] Sheng-Hsuan Lin, Jessica G. Young, Roger Logan, and Tyler J. VanderWeele. Mediation analysis for a survival outcome with time-varying exposures, mediators, and confounders. Statistics in Medicine, 36(26):

4153–4166, November 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stephens-Shields:2017:BPR

[2867] Alisa J. Stephens-Shields, Andrew J. Spieker, Amanda Anderson, Paul Drawz, Michael Fischer, Stephen M. Sozio, Harold Feldman, Marshall Joffe, Wei Yang, Tom Greene, and Cric Study Investigators. Blood pressure and the risk of chronic kidney disease progression using multistate marginal structural models in the CRIC study. Statistics in Medicine, 36(26):4167–4181, November 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bind:2017:QCM

[2868] M.-A. Bind, T. J. VanderWeele, J. D. Schwartz, and B. A. Coull. Quantile causal mediation analysis allowing longitudinal data. Statistics in Medicine, 36(26):4182–4195, November 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hogg:2017:BAP

[2869] Tanja Högg, John Petkau, Yinshan Zhao, Paul Gustafson, José Ma Wijnands, and Helen Tremlett. Bayesian analysis of pair-matched case-control studies subject to outcome misclassification. Statistics in Medicine, 36(26):4196–4213, November 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dagne:2017:JTP

[2870] Getachew A. Dagne. Joint two-part tobit models for longitudinal and time-to-event data. Statistics in Medicine, 36(26):4214–4229, November 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Samawi:2017:NOM

[2871] Hani M. Samawi, Jingjing Yin, Haresh Rochani, and Viral Panchal. Notes on the overlap measure as an alternative to the Youden index: How are they related? Statistics in Medicine, 36(26):4230–4240, November 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIba

[2872] Anonymous. Issue information. Statistics in Medicine, 36(27):4241–4242, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Conlon:2017:LBC

[2873] Anna Conlon, Jeremy Taylor, Yun Li, Karla Diaz-Ordaz, and Michael Elliott. Links between causal effects and causal association for surrogacy evaluation in a Gaussian setting. Statistics in Medicine, 36(27):4243– 4265, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petropoulou:2017:CHV

[2874] Maria Petropoulou and Dimitris Mavridis. A comparison of 20 heterogeneity variance estimators in statistical synthesis of results from studies: a simulation study. Statistics in Medicine, 36(27):4266–4280, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Timkova:2017:SCC

[2875] Jana Timkova, Lukas Kotik, and Ladislav Tomasek. Study of coverage of confidence intervals for the standardized mortality ratio in studies with missing death certificates. Statistics in Medicine, 36(27):4281–4300, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barnett:2017:CSM

[2876] Helen Yvette Barnett, Helena Geys, Tom Jacobs, and Thomas Jaki. Comparing sampling methods for pharmacokinetic studies using model averaged derived parameters. Statistics in Medicine, 36(27):4301–4315, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2017:MCR

[2877] Qi Liu, Bryan E. Shepherd, Chun Li, and Frank E. Harrell, Jr. Modeling continuous response variables using ordinal regression. Statistics in Medicine, 36(27):4316–4335, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yoneoka:2017:MAS

[2878] Daisuke Yoneoka and Masayuki Henmi. Meta-analytical synthesis of regression coefficients under different categorization scheme of continuous covariates. Statistics in Medicine, 36(27):4336–4352, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ohneberg:2017:MTC

[2879] Kristin Ohneberg, Martin Schumacher, and Jan Beyersmann. Modelling two cause-specific hazards of competing risks in one cumulative proportional odds model? Statistics in Medicine, 36(27):4353–4363, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jonas:2017:TCC

[2880] Sarah Flora Jonas, Cyprien Mbogning, and Philippe Broët. A test for comparing current status survival data with crossing hazard functions and its application to immunogenicity of biotherapeutics. Statistics in Medicine, 36(27):4364–4377, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tothfalusi:2017:AER

[2881] Laszlo Tothfalusi and Laszlo Endrenyi. Algorithms for evaluating reference scaled average bioequivalence: power, bias, and consumer risk. Statistics in Medicine, 36(27):4378–4390, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2017:PRR

[2882] Peter C. Austin and Jason P. Fine. Practical recommendations for reporting Fine–Gray model analyses for competing risk data. Statistics in Medicine, 36(27):4391–4400, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomas:2017:UMM

[2883] Neal Thomas. Understanding MCP-MOD dose finding as a method based on linear regression. Statistics in Medicine, 36(27):4401–4413, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Doi:2017:CTE

[2884] Masaaki Doi, Daisuke Watanabe, Kazumasa Takenouchi, and Masaru Tsuchikawa. Comments on “A taxonomy of estimands for regulatory clinical trials with discontinuations”. Statistics in Medicine, 36(27):4414– 4416, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Permutt:2017:ARC

[2885] Thomas Permutt. Author’s reply to comments on “A taxonomy of estimands for regulatory clinical trials with discontinuations”. Statistics in Medicine, 36(27):4417, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:2017:C

[2886] John N. S. Matthews and N. H. Badi. Correction. Statistics in Medicine, 36(27):4418, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colantuoni:2017:C

[2887] Elizabeth Colantuoni and Michael Rosenblum. Correction. Statistics in Medicine, 34(19):4419, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Govindarajulu:2017:LCEb

[2888] Usha S. Govindarajulu, Marco Stillo, David Goldfarb, Michael E. Matheny, and Frederic S. Resnic. Learning curve estimation in medical devices and procedures: hierarchical modeling. Statistics in Medicine, 34(18):4420, November 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIbb

[2889] Anonymous. Issue information. Statistics in Medicine, 36(28):4421–4422, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dmitrienko:2017:EMI

[2890] Alex Dmitrienko and Ralph B. D’Agostino, Sr. Editorial: Multiplicity issues in clinical trials. Statistics in Medicine, 36(28):4423–4426, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chuang-Stein:2017:CSN

[2891] Christy Chuang-Stein and Jianjun (David) Li. Changes are still needed on multiple co-primary endpoints. Statistics in Medicine, 36(28):4427– 4436, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sankoh:2017:CME

[2892] Abdul J. Sankoh, Haihong Li, and Ralph B. D’Agostino, Sr. Composite and multicomponent end points in clinical trials. Statistics in Medicine, 36(28):4437–4440, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Snapinn:2017:SRC

[2893] Steven Snapinn. Some remaining challenges regarding multiple endpoints in clinical trials. Statistics in Medicine, 36(28):4441–4445, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dmitrienko:2017:MCS

[2894] Alex Dmitrienko, Brian Millen, and Ilya Lipkovich. Multiplicity considerations in subgroup analysis. Statistics in Medicine, 36(28):4446–4454, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2017:NRI

[2895] Michael J. Pencina, Ewout W. Steyerberg, and Ralph B. D’Agostino, Sr. Net reclassification index at event rate: properties and relationships. Statistics in Medicine, 36(28):4455–4467, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chipman:2017:SPI

[2896] J. Chipman and D. Braun. Simpson’s paradox in the integrated discrimination improvement. Statistics in Medicine, 36(28):4468–4481, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2017:DSI

[2897] Michael J. Pencina, Jason P. Fine, and Ralph B. D’Agostino, Sr. Discrimination slope and integrated discrimination improvement — properties, relationships and impact of calibration. Statistics in Medicine, 36(28): 4482–4490, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2017:STT

[2898] Stuart G. Baker. The summary test tradeoff: a new measure of the value of an additional risk prediction marker. Statistics in Medicine, 36(28): 4491–4494, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanSmeden:2017:ERN

[2899] Maarten van Smeden and Karel G. M. Moons. Event rate net reclassification index and the integrated discrimination improvement for studying incremental value of risk markers. Statistics in Medicine, 36(28):4495– 4497, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2017:CRR

[2900] Nancy R. Cook, Olga V. Demler, and Nina P. Paynter. Clinical risk reclassification at 10 years. Statistics in Medicine, 36(28):4498–4502, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kerr:2017:FTF

[2901] Kathleen F. Kerr and Holly Janes. First things first: risk model performance metrics should reflect the clinical application. Statistics in Medicine, 36(28):4503–4508, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilson:2017:RAN

[2902] Peter W. F. Wilson and Leslee J. Shaw. Risk assessment with newer statistical metrics. Statistics in Medicine, 36(28):4509–4510, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2017:ARC

[2903] Michael J. Pencina, Jonathan Chipman, Ewout W. Steyerberg, Danielle Braun, Jason P. Fine, and Ralph B. D’Agostino, Sr. Authors’ response to comments. Statistics in Medicine, 36(28):4511–4513, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sweeting:2017:URB

[2904] Michael J. Sweeting, Jessica K. Barrett, Simon G. Thompson, and Angela M. Wood. The use of repeated blood pressure measures for cardiovascular risk prediction: a comparison of statistical models in the ARIC study. Statistics in Medicine, 36(28):4514–4528, December 10, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vergouwe:2017:CTP

[2905] Yvonne Vergouwe, Daan Nieboer, Rianne Oostenbrink, Thomas P. A. Debray, Gordon D. Murray, Michael W. Kattan, Hendrik Koffijberg, Karel G. M. Moons, and Ewout W. Steyerberg. A closed testing procedure to select an appropriate method for updating prediction models. Statistics in Medicine, 36(28):4529–4539, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yue:2017:NVB

[2906] Jin Yue, Xin Lai, Liu Liu, and Paul B. S. Lai. A new VLAD-based control chart for detecting surgical outcomes. Statistics in Medicine, 36 (28):4540–4547, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McParland:2017:CHD

[2907] D. McParland, C. M. Phillips, L. Brennan, H. M. Roche, and I. C. Gormley. Clustering high-dimensional mixed data to uncover sub-phenotypes: joint analysis of phenotypic and genotypic data. Statistics in Medicine, 36(28):4548–4569, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubin:2017:JLR

[2908] Maria Laura Rubin, Wenyaw Chan, Jose-Miguel Yamal, and Claudia Sue Robertson. A joint logistic regression and covariate-adjusted continuoustime Markov chain model. Statistics in Medicine, 36(28):4570–4582, December 10, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIbc

[2909] Anonymous. Issue information. Statistics in Medicine, 36(29):4583–4584, December 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pallmann:2017:SCR

[2910] Philip Pallmann and Thomas Jaki. Simultaneous confidence regions for multivariate bioequivalence. Statistics in Medicine, 36(29):4585–4603, December 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Antonelli:2017:SEE

[2911] Joseph Antonelli, Bing Han, and Matthew Cefalu. A synthetic estimator for the efficacy of clinical trials with all-or-nothing compliance. Statistics in Medicine, 36(29):4604–4615, December 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2017:IPI

[2912] Chien-Ju Lin and James M. S. Wason. Improving phase II oncology trials using best observed RECIST response as an endpoint by modelling continuous tumour measurements. Statistics in Medicine, 36(29):4616– 4626, December 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:2017:MRI

[2913] John R. Thompson, Cosetta Minelli, Jack Bowden, Fabiola Del Greco M., Dipender Gill, Elinor M. Jones, Chin Yang Shapland, and Nuala A. Sheehan. Mendelian randomization incorporating uncertainty about pleiotropy. Statistics in Medicine, 36(29):4627–4645, December 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2017:SDS

[2914] Suhyun Kang, Wenbin Lu, and Rui Song. Subgroup detection and sample size calculation with proportional hazards regression for survival data. Statistics in Medicine, 36(29):4646–4659, December 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cruz:2017:RIT

[2915] Maricela Cruz, Miriam Bender, and Hernando Ombao. A robust interrupted time series model for analyzing complex health care intervention data. Statistics in Medicine, 36(29):4660–4676, December 20, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buhule:2017:BHJ

[2916] O. D. Buhule, A. S. Wahed, and A. O. Youk. Bayesian hierarchical joint modeling of repeatedly measured continuous and ordinal markers of disease severity: Application to Ugandan diabetes data. Statistics in Medicine, 36(29):4677–4691, December 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2017:SJM

[2917] Zhigang Li, H. R. Frost, Tor D. Tosteson, Lihui Zhao, Lei Liu, Kathleen Lyons, Huaihou Chen, Bernard Cole, David Currow, and Marie Bakitas. A semiparametric joint model for terminal trend of quality of life and survival in palliative care research. Statistics in Medicine, 36(29):4692– 4704, December 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rees:2017:EME

[2918] Jessica M. B. Rees, Angela M. Wood, and Stephen Burgess. Extending the MR-Egger method for multivariable Mendelian randomization to correct for both measured and unmeasured pleiotropy. Statistics in Medicine, 36(29):4705–4718, December 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crowther:2017:PMS

[2919] Michael J. Crowther and Paul C. Lambert. Parametric multistate survival models: Flexible modelling allowing transition-specific distributions with application to estimating clinically useful measures of effect differences. Statistics in Medicine, 36(29):4719–4742, December 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2017:GSM

[2920] Xing-Rong Liu, Yudi Pawitan, and Mark S. Clements. Generalized survival models for correlated time-to-event data. Statistics in Medicine, 36 (29):4743–4762, December 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2017:IIbd

[2921] Anonymous. Issue information. Statistics in Medicine, 36(30):4763–4764, December 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2017:MBC

[2922] Baiming Zou, Jianwen Cai, Gary G. Koch, Haibo Zhou, and Fei Zou. A model-based conditional power assessment for decision making in randomized controlled trial studies. Statistics in Medicine, 36(30):4765– 4776, December 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Molins:2017:TSD

[2923] Eduard Molins, Erik Cobo, and Jordi Ocaña. Two-stage designs versus European scaled average designs in bioequivalence studies for highly variable drugs: Which to choose? Statistics in Medicine, 36(30):4777– 4788, December 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Doi:2017:BNT

[2924] Masaaki Doi, Fumihiro Takahashi, and Yohei Kawasaki. Bayesian noninferiority test for 2 binomial probabilities as the extension of Fisher exact test. Statistics in Medicine, 36(30):4789–4803, December 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Atkinson:2017:ODS

[2925] Anthony Atkinson and David Pedrosa. Optimum design and sequential treatment allocation in an experiment in deep brain stimulation with sets of treatment combinations. Statistics in Medicine, 36(30):4804–4815, December 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2017:BHM

[2926] Liping Zhao and Zhiming Xia. Bayesian hierarchical model for analyzing multiresponse longitudinal pharmacokinetic data. Statistics in Medicine, 36(30):4816–4830, December 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fu:2017:STI

[2927] Wei Fu and Jeffrey S. Simonoff. Survival trees for interval-censored survival data. Statistics in Medicine, 36(30):4831–4842, December 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2017:MCD

[2928] Zhuoyu Wang, Nandini Dendukuri, Heather J. Zar, and Lawrence Joseph. Modeling conditional dependence among multiple diagnostic tests. Statistics in Medicine, 36(30):4843–4859, December 30, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Warasi:2017:GTR

[2929] Md S. Warasi, Christopher S. McMahan, Joshua M. Tebbs, and Christopher R. Bilder. Group testing regression models with dilution submodels. Statistics in Medicine, 36(30):4860–4872, December 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deOliveira:2017:RCP

[2930] Guilherme Lopes de Oliveira, Rosangela Helena Loschi, and Renato Martins Assunção. A random-censoring Poisson model for underreported data. Statistics in Medicine, 36(30):4873–4892, December 30, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2017:PEP

[2931] Minggen Lu and Chin-Shang Li. Penalized estimation for proportional hazards models with current status data. Statistics in Medicine, 36(30): 4893–4907, December 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ray:2017:IDP

[2932] Evan L. Ray, Krzysztof Sakrejda, Stephen A. Lauer, Michael A. Johansson, and Nicholas G. Reich. Infectious disease prediction with kernel conditional density estimation. Statistics in Medicine, 36(30):4908–4929, December 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2017:AEF

[2933] Feifei Wang, Jian Wang, Alan Gelfand, and Fan Li. Accommodating the ecological fallacy in disease mapping in the absence of individual exposures. Statistics in Medicine, 36(30):4930–4942, December 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2018:TED

[2934] Zhiwei Zhang, Ruizhe Chen, Guoxing Soon, and Hui Zhang. Treatment evaluation for a data-driven subgroup in adaptive enrichment designs of clinical trials. Statistics in Medicine, 37(1):1–11, January 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lemme:2018:ETA

[2935] Francesca Lemme, Gerard J. P. van Breukelen, and Math J. J. M. Candel. Efficient treatment allocation in 2 × 2 multicenter trials when costs and variances are heterogeneous. Statistics in Medicine, 37(1):12–27, January 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Konietschke:2018:SIF

[2936] Frank Konietschke, Randolph R. Aguayo, and Wieland Staab. Simultaneous inference for factorial multireader diagnostic trials. Statistics in Medicine, 37(1):28–47, January 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choi:2018:SAF

[2937] Sangbum Choi, Liang Zhu, and Xuelin Huang. Semiparametric accelerated failure time cure rate mixture models with competing risks. Statistics in Medicine, 37(1):48–59, January 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rhodes:2018:LIM

[2938] K. M. Rhodes, D. Mawdsley, R. M. Turner, H. E. Jones, J. Savović, and J. P. T. Higgins. Label-invariant models for the analysis of metaepidemiological data. Statistics in Medicine, 37(1):60–70, January 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jha:2018:CCR

[2939] Jayant Jha and Atanu Biswas. Circular-circular regression model with a spike at zero. Statistics in Medicine, 37(1):71–81, January 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mozumder:2018:DLI

[2940] Sarwar Islam Mozumder, Mark Rutherford, and Paul Lambert. Direct likelihood inference on the cause-specific cumulative incidence function: a flexible parametric regression modelling approach. Statistics in Medicine, 37(1):82–97, January 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andrews:2018:VES

[2941] Nichole Andrews and Hyunkeun Cho. Validating effectiveness of subgroup identification for longitudinal data. Statistics in Medicine, 37(1): 98–106, January 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Silva:2018:TEP

[2942] Ivair R. Silva. Type I error probability spending for post-market drug and vaccine safety surveillance with binomial data. Statistics in Medicine, 37 (1):107–118, January 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2018:GLM

[2943] Xianhong Xie, Xiaonan Xue, and Howard D. Strickler. Generalized linear mixed model for binary outcomes when covariates are subject to measurement errors and detection limits. Statistics in Medicine, 37(1): 119–136, January 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giordani:2018:RTB

[2944] Paolo Giordani and Henk A. L. Kiers. A review of tensor-based methods and their application to hospital care data. Statistics in Medicine, 37(1): 137–156, January 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guolo:2018:ILB

[2945] Annamaria Guolo. Improving likelihood-based inference in control rate regression. Statistics in Medicine, 37(1):157–166, January 15, 2018. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2018:MNM

[2946] Hongbin Zhang, Hubert Wong, and Lang Wu. A mechanistic nonlinear model for censored and mismeasured covariates in longitudinal models, with application in AIDS studies. Statistics in Medicine, 37(1):167–178, January 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gezmu:2018:P

[2947] Misrak Gezmu and Jing Qin. Preface. Statistics in Medicine, 37(2): 179–180, January 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Theiler:2018:GBO

[2948] James Theiler and Bette Korber. Graph-based optimization of epitope coverage for vaccine antigen design. Statistics in Medicine, 37(2):181– 194, January 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cybis:2018:BNC

[2949] Gabriela B. Cybis, Janet S. Sinsheimer, Trevor Bedford, Andrew Rambaut, Philippe Lemey, and Marc A. Suchard. Bayesian nonparametric clustering in phylogenetics: modeling antigenic evolution in influenza. Statistics in Medicine, 37(2):195–206, January 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2018:SPR

[2950] Zonghui Hu, Jing Qin, and Dean Follmann. Semiparametric pseudoscore for regression with multidimensional but incompletely observed regressor. Statistics in Medicine, 37(2):207–217, January 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goyal:2018:INS

[2951] Ravi Goyal and Victor De Gruttola. Inference on network statistics by restricting to the network space: applications to sexual history data. Statistics in Medicine, 37(2):218–235, January 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carnegie:2018:ECN

[2952] Nicole Bohme Carnegie. Effects of contact network structure on epidemic transmission trees: implications for data required to estimate network structure. Statistics in Medicine, 37(2):236–248, January 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Benkeser:2018:OCV

[2953] David Benkeser, Cheng Ju, Sam Lendle, and Mark van der Laan. Online cross-validation-based ensemble learning. Statistics in Medicine, 37(2): 249–260, January 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zheng:2018:CBC

[2954] Wenjing Zheng, Laura Balzer, Mark van der Laan, Maya Petersen, and the Search Collaboration. Constrained binary classification using ensemble learning: an application to cost-efficient targeted PrEP strategies. Statistics in Medicine, 37(2):261–279, January 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Benkeser:2018:IEC

[2955] David Benkeser, Marco Carone, and Peter B. Gilbert. Improved estimation of the cumulative incidence of rare outcomes. Statistics in Medicine, 37(2):280–293, January 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Halloran:2018:EPE

[2956] M Elizabeth Halloran and Michael G. Hudgens. Estimating population effects of vaccination using large, routinely collected data. Statistics in Medicine, 37(2):294–301, January 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2018:STF

[2957] Hana Lee, Joseph W. Hogan, Becky L. Genberg, Xiaotian K. Wu, Beverly S. Musick, Ann Mwangi, and Paula Braitstein. A state transition framework for patient-level modeling of engagement and retention in HIV care using longitudinal cohort data. Statistics in Medicine, 37(2):302– 319, January 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kitchenham:2018:CES

[2958] Barbara Kitchenham, Lech Madeyski, and François Curtin. Corrections to effect size variances for continuous outcomes of crossover clinical trials. Statistics in Medicine, 37(2):320–323, January 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Inan:2018:CMM

[2959] Gul Inan. Comments on “Marginalized multilevel hurdle and zeroinflated models for overdispersed and correlated count data with excess zeros”. Statistics in Medicine, 37(2):324–326, January 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hattori:2018:SAP

[2960] Satoshi Hattori and Xiao-Hua Zhou. Sensitivity analysis for publication bias in meta-analysis of diagnostic studies for a continuous biomarker. Statistics in Medicine, 37(3):327–342, February 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Breinegaard:2018:PRD

[2961] Nina Breinegaard, Sophia Rabe-Hesketh, and Anders Skrondal. Pairwise residuals and diagnostic tests for misspecified dependence structures in models for binary longitudinal data. Statistics in Medicine, 37(3):343– 356, February 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersson:2018:HME

[2962] Claes Andersson, Tuomas Rajala, and Aila Särkkä. Hierarchical models for epidermal nerve fiber data. Statistics in Medicine, 37(3):357–374, February 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2018:IPD

[2963] Xinrui Zhang, Keith E. Muller, Maureen M. Goodenow, and Yueh-Yun Chi. Internal pilot design for balanced repeated measures. Statistics in Medicine, 37(3):375–389, February 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cho:2018:CAU

[2964] Youngjoo Cho, Chen Hu, and Debashis Ghosh. Covariate adjustment using propensity scores for dependent censoring problems in the accelerated failure time model. Statistics in Medicine, 37(3):390–404, February 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chernyavskiy:2018:CPM

[2965] Pavel Chernyavskiy, Mark P. Little, and Philip S. Rosenberg. Correlated Poisson models for age-period-cohort analysis. Statistics in Medicine, 37 (3):405–424, February 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gravel:2018:WEC

[2966] Christopher A. Gravel and Robert W. Platt. Weighted estimation for confounded binary outcomes subject to misclassification. Statistics in Medicine, 37(3):425–436, February 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2018:DGE

[2967] Cen Wu, Yu Jiang, Jie Ren, Yuehua Cui, and Shuangge Ma. Dissecting gene-environment interactions: a penalized robust approach accounting for hierarchical structures. Statistics in Medicine, 37(3):437–456, February 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longford:2018:DTC

[2968] Nicholas T. Longford. Decision theory for comparing institutions. Statistics in Medicine, 37(3):457–472, February 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2018:ENF

[2969] Xiang Li, Shanghong Xie, Donglin Zeng, and Yuanjia Wang. Efficient ℓ_{0}-norm feature selection based on augmented and penalized minimization. Statistics in Medicine, 37(3):473–486, February 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DePalma:2018:BME

[2970] Glen DePalma and Bruce A. Craig. Bayesian monotonic errors-in-variables models with applications to pathogen susceptibility testing. Statistics in Medicine, 37(3):487–502, February 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Inan:2018:CLB

[2971] Gul Inan. Comments on “Longitudinal beta-binomial modeling using GEE for overdispersed binomial data”. Statistics in Medicine, 37(3): 503–505, February 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2018:DGE

[2972] Gwangsu Kim, Chao-Qiang Lai, Donna K. Arnett, Laurence D. Parnell, Jose M. Ordovas, Yongdai Kim, and Joungyoun Kim. Detection of gene-environment interactions in a family-based population using SCAD. Statistics in Medicine, 37(3):506, February 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2018:FCUa

[2973] Stuart G. Baker. Five criteria for using a surrogate endpoint to predict treatment effect based on data from multiple previous trials. Statistics in Medicine, 37(4):507–518, February 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fay:2018:MPB

[2974] Michael P. Fay, Michael C. Sachs, and Kazutoyo Miura. Measuring precision in bioassays: Rethinking assay validation. Statistics in Medicine, 37(8):519–529, February 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schnitzer:2018:CTL

[2975] Mireille E. Schnitzer and Matthew Cefalu. Collaborative targeted learning using regression shrinkage. Statistics in Medicine, 37(8):530–543, February 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burger:2018:RFB

[2976] Divan Aristo Burger and Robert Schall. Robust fit of Bayesian mixed effects regression models with application to colony forming unit count in tuberculosis research. Statistics in Medicine, 37(8):544–556, February 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2018:MRD

[2977] Xiaoyan Lin, Hua Chen, Don Edwards, and Kerrie P. Nelson. Modeling rater diagnostic skills in binary classification processes. Statistics in Medicine, 37(8):557–571, February 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2018:MCH

[2978] Peter C. Austin, Henrik Stryhn, George Leckie, and Juan Merlo. Measures of clustering and heterogeneity in multilevel Poisson regression analyses of rates/count data. Statistics in Medicine, 37(8):572–589, February 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Deng:2018:HAC

[2979] Yihao Deng and N. Rao Chaganty. Hierarchical Archimedean copula models for the analysis of binary familial data. Statistics in Medicine, 37(8):590–597, February 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2018:IGG

[2980] Mengyun Wu, Jian Huang, and Shuangge Ma. Identifying gene-gene interactions using penalized tensor regression. Statistics in Medicine, 37 (8):598–610, February 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2018:TPJ

[2981] Haocheng Li, John Staudenmayer, Tianying Wang, Sarah Kozey Keadle, and Raymond J. Carroll. Three-part joint modeling methods for complex functional data mixed with zero-and-one-inflated proportions and zeroinflated continuous outcomes with skewness. Statistics in Medicine, 37 (8):611–626, February 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2018:CMB

[2982] Qingxiang Yan, Leonidas E. Bantis, Janet L. Stanford, and Ziding Feng. Combining multiple biomarkers linearly to maximize the partial area under the ROC curve. Statistics in Medicine, 37(8):627–642, February 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitaker:2018:IAS

[2983] Heather J. Whitaker, Yonas Ghebremichael-Weldeselassie, Ian J. Douglas, Liam Smeeth, and C. Paddy Farrington. Investigating the assumptions of the self-controlled case series method. Statistics in Medicine, 37 (8):643–658, February 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2018:MAA

[2984] Wentao Lu, Xinlei Wang, Xiaowei Zhan, and Adi Gazdar. Meta-analysis approaches to combine multiple gene set enrichment studies. Statistics in Medicine, 37(8):659–672, February 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2018:MDC

[2985] Xueying Wang, Pengyue Zhang, Chien-Wei Chiang, Hengyi Wu, Li Shen, Xia Ning, Donglin Zeng, Lei Wang, Sara K. Quinney, Weixing Feng, and Lang Li. Mixture drug-count response model for the high-dimensional drug combinatory effect on myopathy. Statistics in Medicine, 37(8):673– 686, February 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:2018:REP

[2986] Weichung Joe Shih and Yong Lin. Relative efficiency of precision medicine designs for clinical trials with predictive biomarkers. Statistics in Medicine, 37(5):687–709, February 28, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hasler:2018:MAT

[2987] Mario Hasler and Ludwig A. Hothorn. Multi-arm trials with multiple primary endpoints and missing values. Statistics in Medicine, 37(5):710– 721, February 28, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bretz:2018:ASD

[2988] Frank Bretz, Kathrin Möllenhoff, Holger Dette, Wei Liu, and Matthias Trampisch. Assessing the similarity of dose response and target doses in two non-overlapping subgroups. Statistics in Medicine, 37(5):722–738, February 28, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoyer:2018:MAC

[2989] Annika Hoyer and Oliver Kuss. Meta-analysis for the comparison of two diagnostic tests — a new approach based on copulas. Statistics in Medicine, 37(5):739–748, February 28, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rauch:2018:WCE

[2990] Geraldine Rauch, Kevin Kunzmann, Meinhard Kieser, Karl Wegscheider, Jochem König, and Christine Eulenburg. A weighted combined effect measure for the analysis of a composite time-to-first-event endpoint with components of different clinical relevance. Statistics in Medicine, 37(5): 749–767, February 28, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2018:RAI

[2991] Han Zhang, Peijie Wang, and Jianguo Sun. Regression analysis of interval-censored failure time data with possibly crossing hazards. Statistics in Medicine, 37(5):768–775, February 28, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keown-Stoneman:2018:EDB

[2992] Charles Donald George Keown-Stoneman, Julie Horrocks, and Gerarda Darlington. Exponential decay for binary time-varying covariates in Cox models. Statistics in Medicine, 37(5):776–788, February 28, 2018. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yue:2018:SBH

[2993] Mu Yue, Jialiang Li, and Shuangge Ma. Sparse boosting for highdimensional survival data with varying coefficients. Statistics in Medicine, 37(5):789–800, February 28, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choo-Wosoba:2018:BAA

[2994] Hyoyoung Choo-Wosoba, Jeremy Gaskins, Steven Levy, and Somnath Datta. A Bayesian approach for analyzing zero-inflated clustered count data with dispersion. Statistics in Medicine, 37(5):801–812, February 28, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proudfoot:2018:JMC

[2995] James Proudfoot, Walter Faig, Loki Natarajan, and Ronghui Xu. A joint marginal-conditional model for multivariate longitudinal data. Statistics in Medicine, 37(5):813–828, February 28, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shardell:2018:JME

[2996] Michelle Shardell and Luigi Ferrucci. Joint mixed-effects models for causal inference with longitudinal data. Statistics in Medicine, 37(5): 829–846, February 28, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

An:2018:AUP

[2997] Chuankai An, A. James O’Malley, Daniel N. Rockmore, and Corey D. Stock. Analysis of the U.S. patient referral network. Statistics in Medicine, 37(5):847–866, February 28, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holzhauer:2018:ESA

[2998] Björn Holzhauer, Craig Wang, and Heinz Schmidli. Evidence synthesis from aggregate recurrent event data for clinical trial design and analysis. Statistics in Medicine, 37(6):867–882, March 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hemming:2018:MCT

[2999] Karla Hemming, Monica Taljaard, and Andrew Forbes. Modeling clustering and treatment effect heterogeneity in parallel and stepped-wedge cluster randomized trials. Statistics in Medicine, 37(6):883–898, March 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wright:2018:NCC

[3000] Stephen T. Wright, Louise M. Ryan, and Tung Pham. A novel casecontrol subsampling approach for rapid model exploration of large clustered binary data. Statistics in Medicine, 37(6):899–913, March 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2018:TED

[3001] Catherine Lee, Rebecca A. Betensky, and for the Alzheimer’s Disease Neuroimaging Initiative. Time-to-event data with time-varying biomarkers measured only at study entry, with applications to Alzheimer’s disease. Statistics in Medicine, 37(6):914–932, March 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xia:2018:BIU

[3002] Michelle Xia and Paul Gustafson. Bayesian inference for unidirectional misclassification of a binary response trait. Statistics in Medicine, 37(6): 933–947, March 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2018:IIB

[3003] Q. L. Yan, S. Y. Tang, and Y. N. Xiao. Impact of individual behaviour change on the spread of emerging infectious diseases. Statistics in Medicine, 37(6):948–969, March 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ainslie:2018:MLE

[3004] Kylie E. C. Ainslie, Michael J. Haber, Ryan E. Malosh, Joshua G. Petrie, and Arnold S. Monto. Maximum likelihood estimation of influenza vaccine effectiveness against transmission from the household and from the community. Statistics in Medicine, 37(6):970–982, March 15, 2018. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2018:JMM

[3005] Zhen Jiang, Yimeng Liu, Abdus S. Wahed, and Geert Molenberghs. Joint modeling of multiple ordinal adherence outcomes via generalized estimating equations with flexible correlation structure. Statistics in Medicine, 37(6):983–995, March 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2018:SRA

[3006] Chi Hyun Lee, Chiung-Yu Huang, Gongjun Xu, and Xianghua Luo. Semiparametric regression analysis for alternating recurrent event data. Statistics in Medicine, 37(6):996–1008, March 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2018:PLA

[3007] Weihua Zhao, Heng Lian, and Dipankar Bandyopadhyay. A partially linear additive model for clustered proportion data. Statistics in Medicine, 37(6):1009–1030, March 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barbieri:2018:EAE

[3008] Antoine Barbieri, Myriam Tami, Xavier Bry, David Azria, Sophie Gourgou, Caroline Bascoul-Mollevi, and Christian Lavergne. EM algorithm estimation of a structural equation model for the longitudinal study of the quality of life. Statistics in Medicine, 37(6):1031–1046, March 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gabriel:2018:BOR

[3009] Erin E. Gabriel, Martha Nason, Michael P. Fay, and Dean A. Follmann. A boundary-optimized rejection region test for the two-sample binomial problem. Statistics in Medicine, 37(7):1047–1058, March 30, 2018. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2018:CSR

[3010] Dan Jackson, Martin Law, Theo Stijnen, Wolfgang Viechtbauer, and Ian R. White. A comparison of seven random-effects models for metaanalyses that estimate the summary odds ratio. Statistics in Medicine, 37(7):1059–1085, March 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lyu:2018:ISR

[3011] Tianmeng Lyu, Xianghua Luo, Gongjun Xu, and Chiung-Yu Huang. Induced smoothing for rank-based regression with recurrent gap time data. Statistics in Medicine, 37(7):1086–1100, March 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomas:2018:PEW

[3012] Emma G. Thomas, Sarah B. Peskoe, and Donna Spiegelman. Prevalence estimation when disease status is verified only among test positives: Applications in HIV screening programs. Statistics in Medicine, 37(7):1101–1114, March 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pateras:2018:DGM

[3013] Konstantinos Pateras, Stavros Nikolakopoulos, and Kit Roes. Datagenerating models of dichotomous outcomes: Heterogeneity in simulation studies for a random-effects meta-analysis. Statistics in Medicine, 37(7):1115–1124, March 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fournier:2018:CEA

[3014] Marie-Cécile Fournier, Etienne Dantan, and Paul Blanche. An R^{2}- curve for evaluating the accuracy of dynamic predictions. Statistics in Medicine, 37(7):1125–1133, March 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2018:MST

[3015] Guowen Huang, Duncan Lee, and E. Marian Scott. Multivariate spacetime modelling of multiple air pollutants and their health effects accounting for exposure uncertainty. Statistics in Medicine, 37(7):1134–1148, March 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2018:BSL

[3016] Chanmin Kim, Michael Daniels, Yisheng Li, Kathrin Milbury, and Lorenzo Cohen. A Bayesian semiparametric latent variable approach to causal mediation. Statistics in Medicine, 37(7):1149–1161, March 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2018:UTR

[3017] Yan Li, Tao Xiao, Dandan Liao, and Mei-Ling Ting Lee. Using threshold regression to analyze survival data from complex surveys: With application to mortality linked NHANES III Phase II genetic data. Statistics in Medicine, 37(7):1162–1177, March 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Noma:2018:BTC

[3018] Hisashi Noma, Kengo Nagashima, Kazushi Maruo, Masahiko Gosho, and Toshi A. Furukawa. Bartlett-type corrections and bootstrap adjustments of likelihood-based inference methods for network meta-analysis. Statistics in Medicine, 37(7):1178–1190, March 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davies:2018:TKE

[3019] Tilman M. Davies, Jonathan C. Marshall, and Martin L. Hazelton. Tutorial on kernel estimation of continuous spatial and spatiotemporal relative risk. Statistics in Medicine, 37(7):1191–1221, March 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martinez-Camblor:2018:PNO

[3020] P. Martínez-Camblor. On the paper “Notes on the overlap measure as an alternative to the Youden index”. Statistics in Medicine, 37(7):1222– 1224, March 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yin:2018:RCP

[3021] Jingjing Yin, Hani M. Samawi, and Haresh Rochani. Response to comments by Pablo Martínez–Camblor on “Notes on Overlap Measure as an Alternative to Youden Index: How are they related?”. Statistics in Medicine, 37(7):1225–1226, March 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sudell:2018:BRI

[3022] Maria Sudell, Catrin Tudur Smith, François Gueyffier, and Ruwanthi Kolamunnage-Dona. Book review: Investigation of 2-stage meta-analysis methods for joint longitudinal and time-to-event data through simulation and real data application. Statistics in Medicine, 37(8):1227–1244, April 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gillaizeau:2018:IPW

[3023] Florence Gillaizeau, Thomas Sénage, Florent Le Borgne, Thierry Le Tourneau, Jean-Christian Roussel, Karen Leffondrè, Raphaël Porcher, Bruno Giraudeau, Etienne Dantan, and Yohann Foucher. Inverse probability weighting to control confounding in an illness-death model for interval-censored data. Statistics in Medicine, 37(8):1245–1258, April 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wong:2018:ERF

[3024] Benedict H. W. Wong, Sarah B. Peskoe, and Donna Spiegelman. The effect of risk factor misclassification on the partial population attributable risk. Statistics in Medicine, 37(8):1259–1275, April 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oh:2018:CAT

[3025] Eric J. Oh, Bryan E. Shepherd, Thomas Lumley, and Pamela A. Shaw. Considerations for analysis of time-to-event outcomes measured with error: Bias and correction with SIMEX. Statistics in Medicine, 37(8): 1276–1289, April 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hellton:2018:FFF

[3026] Kristoffer H. Hellton and Nils Lid Hjort. Fridge: Focused fine-tuning of ridge regression for personalized predictions. Statistics in Medicine, 37(8):1290–1303, April 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bein:2018:TSE

[3027] Edward Bein, Jonah Deutsch, Guanglei Hong, Kristin E. Porter, Xu Qin, and Cheng Yang. Two-step estimation in ratio-of-mediator-probability weighted causal mediation analysis. Statistics in Medicine, 37(8):1304– 1324, April 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernhardt:2018:MVI

[3028] Paul W. Bernhardt. Model validation and influence diagnostics for regression models with missing covariates. Statistics in Medicine, 37(8): 1325–1342, April 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martin:2018:MMG

[3029] Glen P. Martin, Mamas A. Mamas, Niels Peek, Iain Buchan, and Matthew Sperrin. A multiple-model generalisation of updating clinical prediction models. Statistics in Medicine, 37(8):1343–1358, April 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Papayiannis:2018:FSL

[3030] Georgios I. Papayiannis, Emmanuel A. Giakoumakis, Efstathios D. Manios, Spyros D. Moulopoulos, Kimon S. Stamatelopoulos, Savvas T. Toumanidis, Nikolaos A. Zakopoulos, and Athanasios N. Yannacopoulos. A functional supervised learning approach to the study of blood pressure data. Statistics in Medicine, 37(8):1359–1375, April 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leroux:2018:DPF

[3031] Andrew Leroux, Luo Xiao, Ciprian Crainiceanu, and William Checkley. Dynamic prediction in functional concurrent regression with an application to child growth. Statistics in Medicine, 37(8):1376–1388, April 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2018:SSD

[3032] Qing Yang, Wing K. Fung, and Gang Li. Sample size determination for jointly testing a cause-specific hazard and the all-cause hazard in the presence of competing risks. Statistics in Medicine, 37(8):1389–1401, April 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weber:2018:NRS

[3033] Kristina Weber, Florian Lasch, and Armin Koch. New research strategy with ambiguous implications: a comment on “Planning future studies based on the conditional power of a meta-analysis”. Statistics in Medicine, 37(8):1402–1404, April 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2018:DPB

[3034] Peter C. Austin, Douglas S. Lee, Ralph B. D’Agostino, and Jason P. Fine. Developing points-based risk-scoring systems in the presence of competing risks. Statistics in Medicine, 37(8):1405, April 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2018:FCUb

[3035] Stuart G. Baker. Five criteria for using a surrogate endpoint to predict treatment effect based on data from multiple previous trials. Statistics in Medicine, 35(22):1406, April 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laber:2018:IOD

[3036] Eric B. Laber, Fan Wu, Catherine Munera, Ilya Lipkovich, Salvatore Colucci, and Steve Ripa. Identifying optimal dosage regimes under safety constraints: an application to long term opioid treatment of chronic pain. Statistics in Medicine, 37(9):1407–1418, April 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morris:2018:MAG

[3037] Tim P. Morris, David J. Fisher, Michael G. Kenward, and James R. Carpenter. Meta-analysis of Gaussian individual patient data: Two-stage or not two-stage? Statistics in Medicine, 37(9):1419–1438, April 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2018:EBT

[3038] Chaeryon Kang, Holly Janes, Parvin Tajik, Henk Groen, Ben Mol, Corine Koopmans, Kim Broekhuijsen, Eva Zwertbroek, Maria van Pampus, and Maureen Franssen. Evaluation of biomarkers for treatment selection using individual participant data from multiple clinical trials. Statistics in Medicine, 37(9):1439–1453, April 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Silverman:2018:SPC

[3039] Rachel Kloss Silverman, Anastasia Ivanova, and Jason Fine. Sequential parallel comparison design with binary and time-to-event outcomes. Statistics in Medicine, 37(9):1454–1466, April 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2018:CPI

[3040] Yongqiang Tang. Controlled pattern imputation for sensitivity analysis of longitudinal binary and ordinal outcomes with nonignorable dropout. Statistics in Medicine, 37(9):1467–1481, April 30, 2018. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [3657].

Gleiss:2018:EVS

[3041] Andreas Gleiss, Michael Gnant, and Michael Schemper. Explained variation in shared frailty models. Statistics in Medicine, 37(9):1482–1490, April 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heimann:2018:NMD

[3042] Günter Heimann, Rossella Belleli, Jouni Kerman, Roland Fisch, Joseph Kahn, Sigrid Behr, and Conny Berlin. A nonparametric method to detect increased frequencies of adverse drug reactions over time. Statistics in Medicine, 37(9):1491–1514, April 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheng:2018:IEP

[3043] Wenting Cheng, Jeremy M. G. Taylor, Pantel S. Vokonas, Sung Kyun Park, and Bhramar Mukherjee. Improving estimation and prediction in linear regression incorporating external information from an established reduced model. Statistics in Medicine, 37(9):1515–1530, April 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lilleborge:2018:OPB

[3044] Marie Lilleborge, Solveig Hofvind, Sofie Sebuødegård, and Ragnar Hauge. Optimizing performance of BreastScreen Norway using value of information in graphical models. Statistics in Medicine, 37(9):1531–1549, April 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mistry:2018:RPA

[3045] Dipesh Mistry, Nigel Stallard, and Martin Underwood. A recursive partitioning approach for subgroup identification in individual patient data meta-analysis. Statistics in Medicine, 37(9):1550–1561, April 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pallmann:2018:SSS

[3046] Philip Pallmann, Christian Ritz, and Ludwig A. Hothorn. Simultaneous small-sample comparisons in longitudinal or multi-endpoint trials using multiple marginal models. Statistics in Medicine, 37(9):1562–1576, April 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ahn:2018:VSG

[3047] Kwang Woo Ahn and Soyoung Kim. Variable selection with group structure in competing risks quantile regression. Statistics in Medicine, 37(9): 1577–1586, April 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maurer:2018:CTE

[3048] Willi Maurer, Byron Jones, and Ying Chen. Controlling the type I error rate in two-stage sequential adaptive designs when testing for average bioequivalence. Statistics in Medicine, 37(10):1587–1607, May 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomas:2018:SID

[3049] Marius Thomas, Björn Bornkamp, and Heidi Seibold. Subgroup identification in dose-finding trials via model-based recursive partitioning. Statistics in Medicine, 37(10):1608–1624, May 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2018:PTC

[3050] Tianlei Chen and Pang Du. Promotion time cure rate model with nonparametric form of covariate effects. Statistics in Medicine, 37(10):1625– 1635, May 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Machado:2018:FMM

[3051] Robson J. M. Machado and Ardo van den Hout. Flexible multistate models for interval-censored data: Specification, estimation, and an application to ageing research. Statistics in Medicine, 37(10):1636–1649, May 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rusa:2018:MMM

[3052] Sárka Rusá, Arnost Komárek, Emmanuel Lesaffre, and Luk Bruyneel. Multilevel moderated mediation model with ordinal outcome. Statistics in Medicine, 37(10):1650–1670, May 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yuan:2018:TFS

[3053] Yan Yuan, Qian M. Zhou, Bingying Li, Hengrui Cai, Eric J. Chow, and Gregory T. Armstrong. A threshold-free summary index of prediction accuracy for censored time to event data. Statistics in Medicine, 37(10):

1671–1681, May 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Madden:2018:MSB

[3054] J. M. Madden, L. D. Browne, X. Li, P. M. Kearney, and A. P. Fitzgerald. Morning surge in blood pressure using a random-effects multiplecomponent cosinor model. Statistics in Medicine, 37(10):1682–1695, May 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2018:BCF

[3055] Qiaolin Chen, Catherine A. Sugar, and Robert E. Weiss. A Bayesian confirmatory factor model for multivariate observations in the form of two-way tables of data. Statistics in Medicine, 37(10):1696–1710, May 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Iorio:2018:BSM

[3056] Maria De Iorio, Natacha Gallot, Beatriz Valcarcel, and Lucy Wedderburn. A Bayesian semiparametric Markov regression model for juvenile dermatomyositis. Statistics in Medicine, 37(10):1711–1731, May 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lam:2018:EAE

[3057] K. F. Lam, Jiajun Xu, and Hongqi Xue. Estimation of age effect with change-points on survival of cancer patients. Statistics in Medicine, 37 (10):1732–1743, May 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindmark:2018:SAU

[3058] Anita Lindmark, Xavier de Luna, and Marie Eriksson. Sensitivity analysis for unobserved confounding of direct and indirect effects using uncertainty intervals. Statistics in Medicine, 37(10):1744–1762, May 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenblad:2018:BRS

[3059] Andreas Rosenblad. Book review: Statistical analysis of contingency tables . Fagerland, Morten W., Lydersen, Stian, Laake, Petter, Chapman and Hall/CRC, Boca Raton, FL, 2017. No. of pages: 634. Price: £77.00 (HB). ISBN 978-1-4665-8817-2. Statistics in Medicine, 37(10):1763–1764, May 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachin:2018:SSE

[3060] John M. Lachin. Sample size evaluation for a multiply matched casecontrol study using the score test from a conditional logistic (discrete Cox PH) regression model. Statistics in Medicine, 37(10):1765–1766, May 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Powers:2018:SMH

[3061] Scott Powers, Junyang Qian, Kenneth Jung, Alejandro Schuler, Nigam H. Shah, Trevor Hastie, and Robert Tibshirani. Some methods for heterogeneous treatment effect estimation in high dimensions. Statistics in Medicine, 37(11):1767–1787, May 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Islas:2018:AES

[3062] Clara Domínguez Islas and Kenneth M. Rice. Addressing the estimation of standard errors in fixed effects meta-analysis. Statistics in Medicine, 37(11):1788–1809, May 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:2018:CMA

[3063] Soojin Park and Esra Kürüm. Causal mediation analysis with multiple mediators in the presence of treatment noncompliance. Statistics in Medicine, 37(11):1810–1829, May 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yuan:2018:SAS

[3064] Ao Yuan, Xiaofei Chen, Yizhao Zhou, and Ming T. Tan. Subgroup analysis with semiparametric models toward precision medicine. Statistics in Medicine, 37(11):1830–1845, May 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2018:TCE

[3065] Bo Lu, Dingjiao Cai, and Xingwei Tong. Testing causal effects in observational survival data using propensity score matching design. Statistics in Medicine, 37(11):1846–1858, May 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martos:2018:DSA

[3066] Gabriel Martos and Miguel de Carvalho. Discrimination surfaces with application to region-specific brain asymmetry analysis. Statistics in Medicine, 37(11):1859–1873, May 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2018:APG

[3067] Peter C. Austin. Assessing the performance of the generalized propensity score for estimating the effect of quantitative or continuous exposures on binary outcomes. Statistics in Medicine, 37(11):1874–1894, May 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reiss:2018:CSV

[3068] Philip T. Reiss. Cross-sectional versus longitudinal designs for function estimation, with an application to cerebral cortex development. Statistics in Medicine, 37(11):1895–1909, May 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cekic:2018:TFT

[3069] Sezen Cekic, Didier Grandjean, and Olivier Renaud. Time, frequency, and time-varying Granger-causality measures in neuroscience. Statistics in Medicine, 37(11):1910–1931, May 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2018:SSR

[3070] Lu Chen, Yihan Sui, Chi Song, and Grzegorz A. Rempala. The sum of standardized residuals: Goodness-of-fit test for binary response models. Statistics in Medicine, 37(11):1932–1941, May 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Molenberghs:2018:RCM

[3071] Geert Molenberghs, Alvaro Florez Poveda, Wondwosen Kassahun, Thomas Neyens, Christel Faes, and Geert Verbeke. Response to comments on “Marginalized multilevel hurdle and zero-inflated models for overdispersed and correlated count data with excess zeros”. Statistics in Medicine, 37(11):1942–1946, May 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zeng:2018:DCT

[3072] Leilei Zeng, Richard J. Cook, and Ker-Ai Lee. Design of cancer trials based on progression-free survival with intermittent assessment. Statistics in Medicine, 37(12):1947–1959, May 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2018:PPI

[3073] Meng Liu and Emily V. Dressler. A predictive probability interim design for phase II clinical trials with continuous endpoints. Statistics in Medicine, 37(12):1960–1972, May 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:2018:ERI

[3074] Paul S. Albert. Estimating recurrence and incidence of preterm birth subject to measurement error in gestational age: a hidden Markov modeling approach. Statistics in Medicine, 37(12):1973–1985, May 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:2018:EER

[3075] Abigail R. Smith, Danting Zhu, Nathan P. Goodrich, Robert M. Merion, and Douglas E. Schaubel. Estimating the effect of a rare time-dependent treatment on the recurrent event rate. Statistics in Medicine, 37(12): 1986–1996, May 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wan:2018:BRG

[3076] Fei Wan, Dylan Small, and Nandita Mitra. Book review: A general approach to evaluating the bias of 2-stage instrumental variable estimators. Statistics in Medicine, 37(12):1997–2015, May 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mahar:2018:BML

[3077] Robert K. Mahar, John B. Carlin, Sarath Ranganathan, Anne-Louise Ponsonby, Peter Vuillermin, and Damjan Vukcevic. Bayesian modelling of lung function data from multiple-breath washout tests. Statistics in Medicine, 37(12):2016–2033, May 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wynants:2018:REM

[3078] L. Wynants, R. D. Riley, D. Timmerman, and B. Van Calster. Randomeffects meta-analysis of the clinical utility of tests and prediction models. Statistics in Medicine, 37(12):2034–2052, May 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hatfield:2018:SCM

[3079] Laura A. Hatfield and Alan M. Zaslavsky. Separable covariance models for health care quality measures across years and topics. Statistics in Medicine, 37(12):2053–2066, May 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hillis:2018:RBR

[3080] Stephen L. Hillis. Relationship between Roe and Metz simulation model for multireader diagnostic data and Obuchowski–Rockette model parameters. Statistics in Medicine, 37(13):2067–2093, June 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2018:SIR

[3081] Kai Yang and Peihua Qiu. Spatiotemporal incidence rate data analysis by nonparametric regression. Statistics in Medicine, 37(13):2094–2107, June 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2018:BRL

[3082] Xiaolei Lin, Robin J. Mermelstein, and Donald Hedeker. Book review: A 3-level Bayesian mixed effects location scale model with an application to ecological momentary assessment data. Statistics in Medicine, 37(13): 2108–2119, June 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zelnick:2018:LBA

[3083] Leila R. Zelnick, Jonathan S. Schildcrout, and Patrick J. Heagerty. Likelihood-based analysis of outcome-dependent sampling designs with longitudinal data. Statistics in Medicine, 37(13):2120–2133, June 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ren:2018:LNV

[3084] Xiaowei Ren, Shanshan Li, Changyu Shen, and Zhangsheng Yu. Linear and nonlinear variable selection in competing risks data. Statistics in Medicine, 37(13):2134–2147, June 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Krol:2018:MJF

[3085] Agnieszka Król, Christophe Tournigand, Stefan Michiels, and Virginie Rondeau. Multivariate joint frailty model for the analysis of nonlinear tumor kinetics and dynamic predictions of death. Statistics in Medicine, 37(13):2148–2161, June 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2018:JAL

[3086] Menghan Li and Lan Kong. Joint analysis of left-censored longitudinal biomarker and binary outcome via latent class modeling. Statistics in Medicine, 37(13):2162–2173, June 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hyun:2018:GME

[3087] Noorie Hyun, Joseph L. Gastwirth, and Barry I. Graubard. Grouping methods for estimating the prevalences of rare traits from complex survey data that preserve confidentiality of respondents. Statistics in Medicine, 37(13):2174–2186, June 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2018:PPM

[3088] Ming Zhou, Qi Tang, Lixin Lang, Jun Xing, and Kay Tatsuoka. Predictive probability methods for interim monitoring in clinical trials with longitudinal outcomes. Statistics in Medicine, 37(14):2187–2207, June 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2018:CRN

[3089] Heng Zhou, Thomas A. Murray, Haitao Pan, and Ying Yuan. Comparative review of novel model-assisted designs for phase I clinical trials. Statistics in Medicine, 37(14):2208–2222, June 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2018:QRV

[3090] Shiqi Zhang and Jørgen Holm Petersen. Quantifying rater variation for ordinal data using a rating scale model. Statistics in Medicine, 37(14): 2223–2237, June 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2018:PHM

[3091] Jing Xu, Jun Ma, Michael H. Connors, and Henry Brodaty. Proportional hazard model estimation under dependent censoring using copulas and penalized likelihood. Statistics in Medicine, 37(14):2238–2251, June 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schomaker:2018:BIW

[3092] Michael Schomaker and Christian Heumann. Bootstrap inference when using multiple imputation. Statistics in Medicine, 37(14):2252–2266, June 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jozani:2018:QRN

[3093] Mohammad Jafari Jozani, Olawale F. Ayilara, and William D. Leslie. Quantile regression with nominated samples: an application to a bone mineral density study. Statistics in Medicine, 37(14):2267–2283, June 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maringe:2018:EVE

[3094] Camille Maringe, Maja Pohar Perme, Janez Stare, and Bernard Rachet. Explained variation of excess hazard models. Statistics in Medicine, 37 (14):2284–2300, June 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andres:2018:CBO

[3095] Antonio Martín Andrés. Comments on “A boundary-optimized rejection region test for the two-sample binomial problem”. Statistics in Medicine, 37(14):2301–2302, June 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gabriel:2018:RLA

[3096] Erin E. Gabriel, Martha Nason, Michael P. Fay, and Dean Follmann. Response to letter by Antonio Martín Andrés on “A boundary-optimized rejection region test for the two-sample binomial problem”. Statistics in Medicine, 37(14):2303–2306, June 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Horiguchi:2018:FCT

[3097] Miki Horiguchi, Angel M. Cronin, Masahiro Takeuchi, and Hajime Uno. A flexible and coherent test/estimation procedure based on restricted mean survival times for censored time-to-event data in randomized clinical trials. Statistics in Medicine, 37(15):2307–2320, July 10, 2018. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:2018:SOA

[3098] Yinghao Pan, Jianwen Cai, Matthew P. Longnecker, and Haibo Zhou. Secondary outcome analysis for data from an outcome-dependent sampling design. Statistics in Medicine, 37(15):2321–2337, July 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tompsett:2018:URF

[3099] Daniel Mark Tompsett, Finbarr Leacy, Margarita Moreno-Betancur, Jon Heron, and Ian R. White. On the use of the not-at-random fully conditional specification (NARFCS) procedure in practice. Statistics in Medicine, 37(15):2338–2353, July 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moon:2018:TSC

[3100] Nathalie C. Moon, Leilei Zeng, and Richard J. Cook. Tracing studies in cohorts with attrition: Selection models for efficient sampling. Statistics in Medicine, 37(15):2354–2366, July 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newsome:2018:ELT

[3101] Simon J. Newsome, Ruth H. Keogh, and Rhian M. Daniel. Estimating long-term treatment effects in observational data: a comparison of the performance of different methods under real-world uncertainty. Statistics in Medicine, 37(15):2367–2390, July 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nguyen:2018:SPE

[3102] Ngoc T. Nguyen, Ebru K. Bish, and Hrayer Aprahamian. Sequential prevalence estimation with pooling and continuous test outcomes. Statistics in Medicine, 37(15):2391–2426, July 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Su:2018:RAT

[3103] Pei-Fang Su and Siu Hung Cheung. Response-adaptive treatment allocation for survival trials with clustered right-censored data. Statistics in Medicine, 37(16):2427–2439, July 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quintero:2018:CHM

[3104] Adrian Quintero and Emmanuel Lesaffre. Comparing hierarchical models via the marginalized deviance information criterion. Statistics in Medicine, 37(16):2440–2454, July 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittenberg:2018:SSR

[3105] Philipp Wittenberg, Fah Fatt Gan, and Sven Knoth. A simple signaling rule for variable life-adjusted display derived from an equivalent riskadjusted CUSUM chart. Statistics in Medicine, 37(16):2455–2473, July 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jensen:2018:CAS

[3106] Signe M. Jensen and Christian Ritz. A comparison of approaches for simultaneous inference of fixed effects for multiple outcomes using linear mixed models. Statistics in Medicine, 37(16):2474–2486, July 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:2018:RAS

[3107] J. A. Thompson, C. Davey, K. Fielding, J. R. Hargreaves, and R. J. Hayes. Robust analysis of stepped wedge trials using cluster-level summaries within periods. Statistics in Medicine, 37(16):2487–2500, July 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sinnott:2018:PAS

[3108] Jennifer A. Sinnott and Tianxi Cai. Pathway aggregation for survival prediction via multiple kernel learning. Statistics in Medicine, 37(16): 2501–2515, July 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ge:2018:ELT

[3109] Zhiyun Ge, Daniel F. Heitjan, David E. Gerber, Lei Xuan, and Sandi L. Pruitt. Estimating lead-time bias in lung cancer diagnosis of patients with previous cancers. Statistics in Medicine, 37(16):2516–2529, July 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luque-Fernandez:2018:TML

[3110] Miguel Angel Luque-Fernandez, Michael Schomaker, Bernard Rachet, and Mireille E. Schnitzer. Targeted maximum likelihood estimation for a binary treatment: a tutorial. Statistics in Medicine, 37(16):2530–2546, July 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Su:2018:RFI

[3111] Xiaogang Su, Annette T. Peña, Lei Liu, and Richard A. Levine. Random forests of interaction trees for estimating individualized treatment effects in randomized trials. Statistics in Medicine, 37(17):2547–2560, July 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2018:EHS

[3112] Jiacheng Wu, Forrest W. Crawford, David A. Kim, Derek Stafford, and Nicholas A. Christakis. Exposure, hazard, and survival analysis of diffusion on social networks. Statistics in Medicine, 37(17):2561–2585, July 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qiu:2018:AEB

[3113] Weiliang Qiu, Michael A. Sandberg, and Bernard Rosner. Application of empirical Bayes methods to predict the rate of decline in ERG at the individual level among patients with retinitis pigmentosa. Statistics in Medicine, 37(17):2586–2598, July 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Faya:2018:UAD

[3114] Paul Faya, John W. Seaman, Jr., and James D. Stamey. Using accelerated drug stability results to inform long-term studies in shelf life determination. Statistics in Medicine, 37(17):2599–2615, July 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanAert:2018:MEB

[3115] Robbie C. M. van Aert and Dan Jackson. Multistep estimators of the between-study variance: The relationship with the Paule–Mandel estimator. Statistics in Medicine, 37(17):2616–2629, July 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xue:2018:ISH

[3116] Hongqi Xue, Shuang Wu, Yichao Wu, Juan C. Ramirez Idarraga, and Hulin Wu. Independence screening for high dimensional nonlinear additive ODE models with applications to dynamic gene regulatory networks. Statistics in Medicine, 37(17):2630–2644, July 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2018:FQR

[3117] Xiaobing Zhao, Weiwei Wang, Lei Liu, and Ya-Chen T. Shih. A flexible quantile regression model for medical costs with application to medical expenditure panel survey study. Statistics in Medicine, 37(17):2645– 2666, July 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:2018:USP

[3118] A. Lawrence Gould. Unified screening for potential elevated adverse event risk and other associations. Statistics in Medicine, 37(18):2667– 2689, August 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2018:CPP

[3119] Libo Sun and Ying Wan. Conditional power and predictive power based on right censored data with supplementary auxiliary information. Statistics in Medicine, 37(18):2690–2699, August 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saha-Chaudhuri:2018:DTS

[3120] P. Saha-Chaudhuri and P. J. Heagerty. Dynamic thresholds and a summary ROC curve: Assessing prognostic accuracy of longitudinal markers. Statistics in Medicine, 37(18):2700–2714, August 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davenport:2018:FIB

[3121] Clemontina A. Davenport, Arnab Maity, and Veerabhadran Baladandayuthapani. Functional interaction-based nonlinear models with application to multiplatform genomics data. Statistics in Medicine, 37 (18):2715–2733, August 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehta:2018:PIC

[3122] Shraddha Mehta, Rowena F. Bastero-Caballero, Yijun Sun, Ray Zhu, Diane K. Murphy, Bhushan Hardas, and Gary Koch. Performance of intraclass correlation coefficient (ICC) as a reliability index under various distributions in scale reliability studies. Statistics in Medicine, 37 (18):2734–2752, August 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teng:2018:TSA

[3123] Ming Teng, Timothy D. Johnson, and Farouk S. Nathoo. Time series analysis of fMRI data: Spatial modelling and Bayesian computation. Statistics in Medicine, 37(18):2753–2770, August 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2018:JMR

[3124] Cong Xu, Vernon M. Chinchilli, and Ming Wang. Joint modeling of recurrent events and a terminal event adjusted for zero inflation and a matched design. Statistics in Medicine, 37(18):2771–2786, August 15, 2018. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troendle:2018:DCR

[3125] James F. Troendle, Eric S. Leifer, and Lauren Kunz. Dealing with competing risks in clinical trials: How to choose the primary efficacy analysis? Statistics in Medicine, 37(19):2787–2796, August 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Punzo:2018:MGH

[3126] Antonio Punzo, Salvatore Ingrassia, and Antonello Maruotti. Multivariate generalized hidden Markov regression models with random covariates: Physical exercise in an elderly population. Statistics in Medicine, 37(19):2797–2808, August 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2018:EAD

[3127] Di Lu, Chunxiao Zhou, Larry Tang, Ming Tan, Ao Yuan, and Leighton Chan. Evaluating accuracy of diagnostic tests without conditional independence assumption. Statistics in Medicine, 37(19):2809–2821, August 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2018:MLD

[3128] Tsung-I Lin, Victor H. Lachos, and Wan-Lun Wang. Multivariate longitudinal data analysis with censored and intermittent missing responses. Statistics in Medicine, 37(19):2822–2835, August 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simpkin:2018:DEL

[3129] Andrew J. Simpkin, Maria Durban, Debbie A. Lawlor, Corrie MacDonald-Wallis, Margaret T. May, Chris Metcalfe, and Kate Tilling. Derivative estimation for longitudinal data analysis: Examining features of blood pressure measured repeatedly during pregnancy. Statistics in Medicine, 37(19):2836–2854, August 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Euan:2018:SSB

[3130] Carolina Euán, Hernando Ombao, and Joaquín Ortega. Spectral synchronicity in brain signals. Statistics in Medicine, 37(19):2855–2873, August 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2018:GCI

[3131] Zonghui Hu and Jing Qin. Generalizability of causal inference in observational studies under retrospective convenience sampling. Statistics in Medicine, 37(19):2874–2883, August 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Storlie:2018:CVS

[3132] C. B. Storlie, S. M. Myers, S. K. Katusic, A. L. Weaver, R. G. Voigt, P. E. Croarkin, R. E. Stoeckel, and J. D. Port. Clustering and variable selection in the presence of mixed variable types and missing data. Statistics in Medicine, 37(19):2884–2899, August 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2018:CTS

[3133] Chuanwu Zhang, Matthew S. Mayo, and Byron J. Gajewski. Comments on “Tutorial on statistical considerations on subgroup analysis in confirmatory clinical trials”. Statistics in Medicine, 37(19):2900–2901, August 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2018:CIA

[3134] Lei Li, Matthew A. Rysavy, and Abhik Das. Comments on “Intermediate and advanced topics in multilevel logistic regression analysis”. Statistics in Medicine, 37(19):2902–2906, August 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Imai:2018:SAM

[3135] Kosuke Imai and Zhichao Jiang. A sensitivity analysis for missing outcomes due to truncation by death under the matched-pairs design. Statistics in Medicine, 37(20):2907–2922, September 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fay:2018:CEC

[3136] Michael P. Fay, Erica H. Brittain, Joanna H. Shih, Dean A. Follmann, and Erin E. Gabriel. Causal estimands and confidence intervals associated with Wilcoxon–Mann–Whitney tests in randomized experiments. Statistics in Medicine, 37(20):2923–2937, September 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [3658].

Weiss:2018:BMA

[3137] Robert E. Weiss, Xiaomao Xia, Nan Zhang, Hui Wang, and Eric Chi. Bayesian methods for analysis of biosimilar phase III trials. Statistics in Medicine, 37(20):2938–2953, September 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2018:UGA

[3138] Minsun Song. A unified genetic association test robust to latent population structure for a count phenotype. Statistics in Medicine, 37(20): 2954–2967, September 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moellenhoff:2018:RAD

[3139] Kathrin Moellenhoff, Holger Dette, Evangelos Kotzagiorgis, Stanislas Volgushev, and Olivier Collignon. Regulatory assessment of drug dissolution profiles comparability via maximum deviation. Statistics in Medicine, 37(20):2968–2981, September 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2018:MSB

[3140] Chun-Shu Chen and Chung-Wei Shen. Model selection based on resampling approaches for cluster longitudinal data with missingness in outcomes. Statistics in Medicine, 37(20):2982–2997, September 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spagnoli:2018:BFM

[3141] Alessandra Spagnoli, Maria Francesca Marino, and Marco Alfò. A bidimensional finite mixture model for longitudinal data subject to dropout. Statistics in Medicine, 37(20):2998–3011, September 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chatterjee:2018:GRZ

[3142] Saptarshi Chatterjee, Shrabanti Chowdhury, Himel Mallick, Prithish Banerjee, and Broti Garai. Group regularization for zero-inflated negative binomial regression models with an application to health care demand in Germany. Statistics in Medicine, 37(20):3012–3026, September 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanBreukelen:2018:EDC

[3143] Gerard J. P. van Breukelen and Math J. J. M. Candel. Efficient design of cluster randomized trials with treatment-dependent costs and treatmentdependent unknown variances. Statistics in Medicine, 37(21):3027–3046, September 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2018:NCD

[3144] Guoqiao Wang, Scott Berry, Chengjie Xiong, Jason Hassenstab, Melanie Quintana, Eric M. McDade, Paul Delmar, Matteo Vestrucci, Gopalan Sethuraman, Randall J. Bateman, and Dominantly Inherited Alzheimer Network Trials Unit. A novel cognitive disease progression model for clinical trials in autosomal-dominant Alzheimer’s disease. Statistics in Medicine, 37(21):3047–3055, September 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ryeznik:2018:CSR

[3145] Yevgen Ryeznik and Oleksandr Sverdlov. A comparative study of restricted randomization procedures for multiarm trials with equal or unequal treatment allocation ratios. Statistics in Medicine, 37(21):3056– 3077, September 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Winell:2018:GSI

[3146] Henric Winell and Johan Lindbäck. A general score-independent test for order-restricted inference. Statistics in Medicine, 37(21):3078–3090, September 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2018:NAD

[3147] Yayuan Zhu, Jerald F. Lawless, and Cecilia A. Cotton. Nonparametric analysis of dependently interval-censored failure time data. Statistics in Medicine, 37(21):3091–3105, September 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2018:QEP

[3148] Cai Wu and Liang Li. Quantifying and estimating the predictive accuracy for censored time-to-event data with competing risks. Statistics in Medicine, 37(21):3106–3124, September 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yen:2018:BME

[3149] Amy Ming-Fang Yen and Hsiu-Hsi Chen. Bayesian measurement-error-driven hidden Markov regression model for calibrating the effect of covariates on multistate outcomes: Application to androgenetic alopecia.

Statistics in Medicine, 37(21):3125–3146, September 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cameron:2018:ETS

[3150] Briana Cameron, Peter Peduzzi, and Denise Esserman. Extensions to the two-stage randomized trial design for testing treatment, selfselection, and treatment preference effects to binary outcomes. Statistics in Medicine, 37(22):3147–3178, September 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kimani:2018:PEF

[3151] Peter K. Kimani, Susan Todd, Lindsay A. Renfro, and Nigel Stallard. Point estimation following two-stage adaptive threshold enrichment clinical trials. Statistics in Medicine, 37(22):3179–3196, September 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2018:NSS

[3152] Yongqiang Tang. A noniterative sample size procedure for tests based on t distributions. Statistics in Medicine, 37(22):3197–3213, September 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jin:2018:DPC

[3153] Jin Jin, Lin Zhang, Ethan Leng, Gregory J. Metzger, and Joseph S. Koopmeiners. Detection of prostate cancer with multiparametric MRI utilizing the anatomic structure of the prostate. Statistics in Medicine, 37(22):3214–3229, September 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2018:MIC

[3154] Molei Liu, Ming Hu, and Xiao-Hua Zhou. Modeling individualized coefficient alpha to measure quality of test score data. Statistics in Medicine, 37(22):3230–3243, September 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Deng:2018:DDP

[3155] Qiqi Deng, Xiaofei Bai, and Naitee Ting. Dynamic development paths for expanding a proof-of-concept study to explore dose range. Statistics in Medicine, 37(22):3244–3253, September 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hwang:2018:MNM

[3156] Hyunsoo Hwang and Stacia M. DeSantis. Multivariate network metaanalysis to mitigate the effects of outcome reporting bias. Statistics in Medicine, 37(22):3254–3266, September 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2018:SIT

[3157] Ye He, Huazhen Lin, and Dongsheng Tu. A single-index threshold Cox proportional hazard model for identifying a treatment-sensitive subset based on multiple biomarkers. Statistics in Medicine, 37(23):3267–3279, October 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2018:IBM

[3158] Rengyi Xu, Devan V. Mehrotra, and Pamela A. Shaw. Incorporating baseline measurements into the analysis of crossover trials with timeto-event endpoints. Statistics in Medicine, 37(23):3280–3292, October 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ding:2018:SIA

[3159] Ying Ding, Shengchun Kong, Shan Kang, and Wei Chen. A semiparametric imputation approach for regression with censored covariate with application to an AMD progression study. Statistics in Medicine, 37 (23):3293–3308, October 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wendling:2018:CME

[3160] T. Wendling, K. Jung, A. Callahan, A. Schuler, N. H. Shah, and B. Gallego. Comparing methods for estimation of heterogeneous treatment effects using observational data from health care databases. Statistics in Medicine, 37(23):3309–3324, October 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Corpas-Burgos:2018:SFZ

[3161] Francisca Corpas-Burgos, Gonzalo García-Donato, and Miguel A. Martinez-Beneito. Some findings on zero-inflated and hurdle Poisson models for disease mapping. Statistics in Medicine, 37(23):3325–3337, October 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sutton:2018:SPL

[3162] Matthew Sutton, Rodolphe Thiébaut, and Benoît Liquet. Sparse partial least squares with group and subgroup structure. Statistics in Medicine, 37(23):3338–3356, October 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Satten:2018:MAS

[3163] Glen A. Satten, Maiying Kong, and Somnath Datta. Multisample adjusted U-statistics that account for confounding covariates. Statistics in Medicine, 37(23):3357–3372, October 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tomita:2018:BCE

[3164] Hiroaki Tomita, Hironori Fujisawa, and Masayuki Henmi. A biascorrected estimator in multiple imputation for missing data. Statistics in Medicine, 37(23):3373–3386, October 15, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sugitani:2018:FAA

[3165] Toshifumi Sugitani, Martin Posch, Frank Bretz, and Franz Koenig. Flexible alpha allocation strategies for confirmatory adaptive enrichment clinical trials with a prespecified subgroup. Statistics in Medicine, 37 (24):3387–3402, October 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miles:2018:CST

[3166] Caleb H. Miles, Joel Schwartz, and Eric J. Tchetgen Tchetgen. A class of semiparametric tests of treatment effect robust to confounder measurement error. Statistics in Medicine, 37(24):3403–3416, October 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qi:2018:SIM

[3167] Lihong Qi, Ying-Fang Wang, Rongqi Chen, Juned Siddique, John Robbins, and Yulei He. Strategies for imputing missing covariates in accelerated failure time models. Statistics in Medicine, 37(24):3417–3436, October 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheysson:2018:EEA

[3168] Felix Cheysson, Marie-Anne Vibet, Didier Guillemot, and Laurence Watier. Estimation of exposure-attributable fractions from time series: a simulation study. Statistics in Medicine, 37(24):3437–3454, October 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Enders:2018:RVC

[3169] Dirk Enders, Susanne Engel, Roland Linder, and Iris Pigeot. Robust versus consistent variance estimators in marginal structural Cox models. Statistics in Medicine, 37(24):3455–3470, October 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2018:ABE

[3170] Fang Liu. Assessment of Bayesian expected power via Bayesian bootstrap. Statistics in Medicine, 37(24):3471–3485, October 30, 2018. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hou:2018:HDV

[3171] Jiayi Hou, Anthony Paravati, Jue Hou, Ronghui Xu, and James Murphy. High-dimensional variable selection and prediction under competing risks with application to SEER-Medicare linked data. Statistics in Medicine, 37(24):3486–3502, October 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Molina:2018:MPS

[3172] J. Molina, M. Sued, and M. Valdora. Models for the propensity score that contemplate the positivity assumption and their application to missing data and causality. Statistics in Medicine, 37(24):3503–3518, October 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2018:TFG

[3173] Renjun Ma, Guohua Yan, and M. Tariqul Hasan. Tweedie family of generalized linear models with distribution-free random effects for skewed longitudinal data. Statistics in Medicine, 37(24):3519–3532, October 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wortman:2018:SRL

[3174] Joan Heck Wortman and Jerome P. Reiter. Simultaneous record linkage and causal inference with propensity score subclassification. Statistics in Medicine, 37(24):3533–3546, October 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehrotra:2018:HRE

[3175] Devan V. Mehrotra and Yiwei Zhang. Hazard ratio estimation and inference in clinical trials with many tied event times. Statistics in Medicine, 37(25):3547–3556, November 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hobbs:2018:BBT

[3176] Brian P. Hobbs and Rick Landin. Bayesian basket trial design with exchangeability monitoring. Statistics in Medicine, 37(25):3557–3572, November 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhan:2018:OUS

[3177] Zhuozhao Zhan, Geertruida H. de Bock, and Edwin R. van den Heuvel. Optimal unidirectional switch designs. Statistics in Medicine, 37(25): 3573–3588, November 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3651].

Claggett:2018:QTT

[3178] Brian Claggett, Lu Tian, Haoda Fu, Scott D. Solomon, and Lee-Jen Wei. Quantifying the totality of treatment effect with multiple event-time observations in the presence of a terminal event from a comparative clinical study. Statistics in Medicine, 37(25):3589–3598, November 10, 2018. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Somanchi:2018:DAP

[3179] Sriram Somanchi, Daniel B. Neill, and Anil V. Parwani. Discovering anomalous patterns in large digital pathology images. Statistics in Medicine, 37(25):3599–3615, November 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mahabadi:2018:SOL

[3180] Samaneh Eftekhari Mahabadi. Second-order local sensitivity to nonignorability in Bayesian inferences. Statistics in Medicine, 37(25):3616– 3636, November 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lefebvre:2018:BFM

[3181] Geneviève Lefebvre, Mariia Samoilenko, Isabelle Boucoiran, and Lucie Blais. A Bayesian finite mixture of bivariate regression model for causal mediation analyses. Statistics in Medicine, 37(25):3637–3660, November 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keogh:2018:MIC

[3182] Ruth H. Keogh and Tim P. Morris. Multiple imputation in Cox regression when there are time-varying effects of covariates. Statistics in Medicine, 37(25):3661–3678, November 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nghiem:2018:DEP

[3183] Linh Nghiem and Cornelis J. Potgieter. Density estimation in the presence of heteroscedastic measurement error of unknown type using phase function deconvolution. Statistics in Medicine, 37(25):3679–3692, November 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taylor:2018:TAD

[3184] Leslie Taylor, Xiao-Hua Zhou, and Peter Rise. A tutorial in assessing disclosure risk in microdata. Statistics in Medicine, 37(25):3693–3706, November 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2018:RLB

[3185] Anonymous. Retraction: Lu, T. Bayesian inference on mixed-effects location scale models with skew-t distribution and mismeasured covariates for longitudinal data. Statist. Med. 2017; 36:2614–2629. doi: 10.1002/sim.7315. Statistics in Medicine, 37(25):3707, November 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2750].

Kim:2018:BSR

[3186] Mi-Ok Kim, Nusrat Harun, Chunyan Liu, Jane C. Khoury, and Joseph P. Broderick. Bayesian selective response-adaptive design using the historical control. Statistics in Medicine, 37(26):3709–3722, November 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:2018:BAS

[3187] Boxian Wei, Thomas M. Braun, Roy N. Tamura, and Kelley M. Kidwell. A Bayesian analysis of small n sequential multiple assignment randomized trials (snSMARTs). Statistics in Medicine, 37(26):3723–3732, November 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2018:ECM

[3188] Lei Liu, Cheng Zheng, and Joseph Kang. Exploring causality mechanism in the joint analysis of longitudinal and survival data. Statistics in Medicine, 37(26):3733–3744, November 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mao:2018:PSW

[3189] Huzhang Mao, Liang Li, Wei Yang, and Yu Shen. On the propensity score weighting analysis with survival outcome: Estimands, estimation, and inference. Statistics in Medicine, 37(26):3745–3763, November 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wen:2018:GRP

[3190] Yalu Wen, Xiaoxi Shen, and Qing Lu. Genetic risk prediction using a spatial autoregressive model with adaptive lasso. Statistics in Medicine, 37(26):3764–3775, November 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2018:AOW

[3191] Ying Liu, Yuanjia Wang, Michael R. Kosorok, Yingqi Zhao, and Donglin Zeng. Augmented outcome-weighted learning for estimating optimal dynamic treatment regimens. Statistics in Medicine, 37(26):3776–3788, November 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2018:EZI

[3192] Rong Luo and Sudhir Paul. Estimation for zero-inflated beta-binomial regression model with missing response data. Statistics in Medicine, 37 (26):3789–3813, November 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Psioda:2018:BDS

[3193] Matthew A. Psioda and Joseph G. Ibrahim. Bayesian design of a survival trial with a cured fraction using historical data. Statistics in Medicine, 37(26):3814–3831, November 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Khan:2018:QRM

[3194] Manzoor Khan and Jake Olivier. Quantifying the regression to the mean effect in Poisson processes. Statistics in Medicine, 37(26):3832–3848, November 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Makela:2018:BIU

[3195] Susanna Makela, Yajuan Si, and Andrew Gelman. Bayesian inference under cluster sampling with probability proportional to size. Statistics in Medicine, 37(26):3849–3868, November 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liang:2018:EIO

[3196] Muxuan Liang, Ting Ye, and Haoda Fu. Estimating individualized optimal combination therapies through outcome weighted deep learning algorithms. Statistics in Medicine, 37(27):3869–3886, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gravel:2018:VSA

[3197] Christopher A. Gravel, Anup Dewanji, Patrick J. Farrell, and Daniel Krewski. A validation sampling approach for consistent estimation of adverse drug reaction risk with misclassified right-censored survival data. Statistics in Medicine, 37(27):3887–3903, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2018:SSC

[3198] Zhiguo Li, Xiaofei Wang, Yuan Wu, and Kouros Owzar. Sample size calculation for studies with grouped survival data. Statistics in Medicine, 37(27):3904–3917, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bebu:2018:PCT

[3199] Ionut Bebu and John M. Lachin. Properties of composite time to first event versus joint marginal analyses of multiple outcomes. Statistics in Medicine, 37(27):3918–3930, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kurum:2018:CMJ

[3200] Esra Kürüm, Daniel R. Jeske, Carolyn E. Behrendt, and Peter Lee. A copula model for joint modeling of longitudinal and time-invariant mixed outcomes. Statistics in Medicine, 37(27):3931–3943, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chaudhari:2018:SDS

[3201] Monica Chaudhari, Edwin H. Kim, Prabhashi W. Withana Gamage, Christopher S. McMahan, and Michael R. Kosorok. Study design with staggered sampling times for evaluating sustained unresponsiveness to peanut sublingual immunotherapy. Statistics in Medicine, 37(27):3944– 3958, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2018:SVC

[3202] Yang Li, Li Qi, and Yanqing Sun. Semiparametric varying-coefficient regression analysis of recurrent events with applications to treatment switching. Statistics in Medicine, 37(27):3959–3974, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marker:2018:SLE

[3203] David A. Marker, Russ Mardon, Frank Jenkins, Joanne Campione, Jennifer Nooney, Jane Li, Sharon Saydeh, Xuanping Zhang, Sundar Shrestha, and Deborah Rolka. State-level estimation of diabetes and prediabetes prevalence: Combining national and local survey data and clinical data. Statistics in Medicine, 37(27):3975–3990, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fay:2018:CIM

[3204] Michael P. Fay and Yaakov Malinovsky. Confidence intervals of the Mann–Whitney parameter that are compatible with the Wilcoxon– Mann–Whitney test. Statistics in Medicine, 37(27):3991–4006, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Domelen:2018:LRC

[3205] Dane R. Van Domelen, Emily M. Mitchell, Neil J. Perkins, Enrique F. Schisterman, Amita K. Manatunga, Yijian Huang, and Robert H. Lyles. Logistic regression with a continuous exposure measured in pools and subject to errors. Statistics in Medicine, 37(27):4007–4021, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bantis:2018:CTC

[3206] Leonidas E. Bantis and Ziding Feng. Comparison of two correlated ROC surfaces at a given pair of true classification rates. Statistics in Medicine, 37(27):4022–4035, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shestopaloff:2018:ADB

[3207] Konstantin Shestopaloff, Michael D. Escobar, and Wei Xu. Analyzing differences between microbiome communities using mixture distributions.

Statistics in Medicine, 37(27):4036–4053, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Psioda:2018:PBA

[3208] Matthew A. Psioda, Mat Soukup, and Joseph G. Ibrahim. A practical Bayesian adaptive design incorporating data from historical controls. Statistics in Medicine, 37(27):4054–4070, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gregg:2018:LRT

[3209] Mary E. Gregg, Somnath Datta, and Doug Lorenz. A log rank test for clustered data with informative within-cluster group size. Statistics in Medicine, 37(27):4071–4082, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2018:PSE

[3210] Wei Zhang, Aiyi Liu, Paul S. Albert, Robert D. Ashmead, Enrique F. Schisterman, and James L. Mills. A pooling strategy to effectively use genotype data in quantitative traits genome-wide association studies. Statistics in Medicine, 38(17):4083–4095, November 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teng:2018:OSP

[3211] Zhaoyang Teng, Liang Liang, Guohui Liu, and Yi Liu. Optimal seamless phase 2/3 oncology trial designs based on Probability of Success (PoS). Statistics in Medicine, 37(28):4097–4113, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2018:QPO

[3212] Lifeng Lin. Quantifying and presenting overall evidence in network metaanalysis. Statistics in Medicine, 37(28):4114–4125, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boatman:2018:ERC

[3213] Jeffrey A. Boatman, David M. Vock, and Joseph S. Koopmeiners. Efficiency and robustness of causal effect estimators when noncompliance is measured with error. Statistics in Medicine, 37(28):4126–4141, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sperrin:2018:UMS

[3214] Matthew Sperrin, Glen P. Martin, Alexander Pate, Tjeerd Van Staa, Niels Peek, and Iain Buchan. Using marginal structural models to adjust for treatment drop-in when developing clinical prediction models. Statistics in Medicine, 37(28):4142–4154, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klich:2018:UIG

[3215] Amna Klich, René Ecochard, and Fabien Subtil. Unequal intra-group variance in trajectory classification. Statistics in Medicine, 37(28):4155– 4166, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoque:2018:MRE

[3216] Md Erfanul Hoque and Mahmoud Torabi. Modeling the random effects covariance matrix for longitudinal data with covariates measurement error. Statistics in Medicine, 37(28):4167–4184, December 10, 2018. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Govindarajulu:2018:SAH

[3217] Usha Govindarajulu, Sandeep Bedi, Aaron Kluger, and Frederic Resnic. Survival analysis of hierarchical learning curves in assessment of cardiac device and procedural safety. Statistics in Medicine, 37(28):4185–4199, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jang:2018:OIA

[3218] Jeong Hoon Jang, Amita K. Manatunga, Andrew T. Taylor, and Qi Long. Overall indices for assessing agreement among multiple raters. Statistics in Medicine, 37(28):4200–4215, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2018:HHB

[3219] Jian Zou, Zhongqiang Zhang, and Hong Yan. A hybrid hierarchical Bayesian model for spatiotemporal surveillance data. Statistics in Medicine, 37(28):4216–4233, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bjornland:2018:PEP

[3220] Thea Bjørnland, Anja Bye, Einar Ryeng, Ulrik Wisløff, and Mette Langaas. Powerful extreme phenotype sampling designs and score tests for genetic association studies. Statistics in Medicine, 37(28):4234–4251, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2018:RCO

[3221] Debashis Ghosh. Relaxed covariate overlap and margin-based causal effect estimation. Statistics in Medicine, 37(28):4252–4265, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2018:BLV

[3222] Xue Li, Pankaj Kumar Choudhary, Swati Biswas, and Xinlei Wang. A Bayesian latent variable approach to aggregation of partial and topranked lists in genomic studies. Statistics in Medicine, 37(28):4266–4278, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ouyang:2018:BSF

[3223] Ming Ouyang, Xiaoqing Wang, Chunjie Wang, and Xinyuan Song. Bayesian semiparametric failure time models for multivariate censored data with latent variables. Statistics in Medicine, 37(28):4279–4297, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Happ:2018:IMA

[3224] Clara Happ, Sonja Greven, and Volker J. Schmid. The impact of model assumptions in scalar-on-image regression. Statistics in Medicine, 37(28): 4298–4317, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ford:2018:CBC

[3225] Whitney P. Ford and Philip M. Westgate. A comparison of bias-corrected empirical covariance estimators with generalized estimating equations in small-sample longitudinal study settings. Statistics in Medicine, 37(28): 4318–4329, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wan:2018:NNR

[3226] Fei Wan. A note on a naive regression-based test on the validity of an instrumental variable. Statistics in Medicine, 37(28):4330–4333, December 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chiu:2018:DEC

[3227] Yi-Da Chiu, Franz Koenig, Martin Posch, and Thomas Jaki. Design and estimation in clinical trials with subpopulation selection. Statistics in Medicine, 37(29):4335–4352, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Samuels:2018:BOO

[3228] Lauren R. Samuels and Robert A. Greevy, Jr. Bagged one-to-one matching for efficient and robust treatment effect estimation. Statistics in Medicine, 37(29):4353–4373, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Follmann:2018:RPB

[3229] Dean Follmann. Reliably picking the best endpoint. Statistics in Medicine, 37(29):4374–4385, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2018:AGE

[3230] Yang Li, Ruofan Bie, Sebastian J. Teran Hidalgo, Yichen Qin, Mengyun Wu, and Shuangge Ma. Assisted gene expression-based clustering with AWNCut. Statistics in Medicine, 37(29):4386–4403, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Legha:2018:IPD

[3231] Amardeep Legha, Richard D. Riley, Joie Ensor, Kym I. E. Snell, Tim P. Morris, and Danielle L. Burke. Individual participant data meta-analysis of continuous outcomes: a comparison of approaches for specifying and estimating one-stage models. Statistics in Medicine, 37(29):4404–4420, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leao:2018:IFC

[3232] Jeremias Leão, Víctor Leiva, Helton Saulo, and Vera Tomazella. Incorporation of frailties into a cure rate regression model and its diagnostics and application to melanoma data. Statistics in Medicine, 37(29):4421– 4440, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Henderson:2018:ALC

[3233] Nicholas C. Henderson and Paul J. Rathouz. AR(1) latent class models for longitudinal count data. Statistics in Medicine, 37(29):4441–4456, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neuhaus:2018:ALD

[3234] John M. Neuhaus, Charles E. McCulloch, and Ross D. Boylan. Analysis of longitudinal data from outcome-dependent visit processes: Failure of proposed methods in realistic settings and potential improvements. Statistics in Medicine, 37(29):4457–4471, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nautiyal:2018:SBC

[3235] Nisheet Nautiyal and Theodore R. Holford. A spatiotemporal backcalculation approach to estimate cancer incidence measures. Statistics in Medicine, 37(29):4472–4489, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ganyani:2018:AIB

[3236] Tapiwa Ganyani, Christel Faes, Gerardo Chowell, and Niel Hens. Assessing inference of the basic reproduction number in an SIR model incorporating a growth-scaling parameter. Statistics in Medicine, 37(29): 4490–4506, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bruckner:2018:NAE

[3237] Matthias Brückner, Hans U. Burger, and Werner Brannath. Nonparametric adaptive enrichment designs using categorical surrogate data. Statistics in Medicine, 37(29):4507–4524, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alonso:2018:MEA

[3238] Ariel Alonso, Wim Van der Elst, and Geert Molenberghs. A maximum entropy approach for the evaluation of surrogate endpoints based on causal inference. Statistics in Medicine, 37(29):4525–4538, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2018:EGL

[3239] Wan-Chen Lee, Sanjoy K. Sinha, Tye E. Arbuckle, and Mandy Fisher. Estimation in generalized linear models under censored covariates with an application to MIREC data. Statistics in Medicine, 37(29):4539–4556, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rhodes:2018:CEM

[3240] K. M. Rhodes, R. M. Turner, R. A. Payne, and I. R. White. Computationally efficient methods for fitting mixed models to electronic health records data. Statistics in Medicine, 37(29):4557–4570, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korre:2018:GFT

[3241] Antonia K. Korre and Vassilis G. S. Vasdekis. Goodness of fit tests for random effect models with binary responses. Statistics in Medicine, 37 (29):4571–4587, December 20, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2018:DCI

[3242] Zhenzhen Xu, Yongsoek Park, Boguang Zhen, and Bin Zhu. Designing cancer immunotherapy trials with random treatment time-lag effect. Statistics in Medicine, 37(30):4589–4609, December 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2018:AVE

[3243] Ting Wang, Xiaofei Wang, Haibo Zhou, Jianwen Cai, and Stephen L. George. Auxiliary variable-enriched biomarker-stratified design. Statistics in Medicine, 37(30):4610–4635, December 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kunz:2018:AMA

[3244] Cornelia Ursula Kunz, Thomas Jaki, and Nigel Stallard. An alternative method to analyse the biomarker-strategy design. Statistics in Medicine, 37(30):4636–4651, December 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Girling:2018:REU

[3245] Alan J. Girling. Relative efficiency of unequal cluster sizes in stepped wedge and other trial designs under longitudinal or cross-sectional sampling. Statistics in Medicine, 37(30):4652–4664, December 30, 2018. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:2018:URE

[3246] Hayley E. Jones, A. E. Ades, Alex J. Sutton, and Nicky J. Welton. Use of a random effects meta-analysis in the design and analysis of a new clinical trial. Statistics in Medicine, 37(30):4665–4679, December 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2018:BVC

[3247] Shelley H. Liu, Jennifer F. Bobb, Birgit Claus Henn, Chris Gennings, Lourdes Schnaas, Martha Tellez-Rojo, David Bellinger, Manish Arora, Robert O. Wright, and Brent A. Coull. Bayesian varying coefficient kernel machine regression to assess neurodevelopmental trajectories associated with exposure to complex mixtures. Statistics in Medicine, 37(30):4680– 4694, December 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2018:LBI

[3248] Woojoo Lee, Arvid Sjölander, Anton Larsson, and Yudi Pawitan. Likelihood-based inference for bounds of causal parameters. Statistics in Medicine, 37(30):4695–4706, December 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2018:MTV

[3249] Yihao Li, Danh V. Nguyen, Yanjun Chen, Connie M. Rhee, Kamyar Kalantar-Zadeh, and Damla Sentürk. Modeling time-varying effects of multilevel risk factors of hospitalizations in patients on dialysis. Statistics in Medicine, 37(30):4707–4720, December 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2018:NMR

[3250] Ming Sun and Yuanjia Wang. Nonlinear model with random inflection points for modeling neurodegenerative disease progression. Statistics in Medicine, 37(30):4721–4742, December 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ortega-Villa:2018:EOT

[3251] Ana M. Ortega-Villa and Paul S. Albert. Estimating onset time from longitudinal data in the presence of measurement error with application to estimating gestational age from maternal anthropometry during pregnancy. Statistics in Medicine, 37(30):4743–4757, December 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2018:MSL

[3252] Guohai Zhou and Lang Wu. Modeling semicontinuous longitudinal data with order constraints. Statistics in Medicine, 37(30):4758–4770, December 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kohler:2018:NAS

[3253] Meike Köhler, Nikolaus Umlauf, and Sonja Greven. Nonlinear association structures in flexible Bayesian additive joint models. Statistics in Medicine, 37(30):4771–4788, December 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maadooliat:2018:NCS

[3254] Mehdi Maadooliat, Ying Sun, and Tianbo Chen. Nonparametric collective spectral density estimation with an application to clustering the brain signals. Statistics in Medicine, 37(30):4789–4806, December 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dutta:2018:RBI

[3255] Sandipan Dutta and Somnath Datta. Rank-based inference for covariate and group effects in clustered data in presence of informative intra-cluster group size. Statistics in Medicine, 37(30):4807–4822, December 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Omidvar:2018:JPS

[3256] Sedigheh Omidvar, Mohammad Jafari Jozani, and Nader Nematollahi. Judgment post-stratification in finite mixture modeling: an example in estimating the prevalence of osteoporosis. Statistics in Medicine, 37(30): 4823–4836, December 30, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenberger:2019:RFC

[3257] William F. Rosenberger, Diane Uschner, and Yanying Wang. Randomization: The forgotten component of the randomized clinical trial. Statistics in Medicine, 38(1):1–12, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:2019:DRF

[3258] Peter Armitage. Discussion of “Randomization: The Forgotten Component of the Randomized Clinical Trial”. Statistics in Medicine, 38 (1):13, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittes:2019:CRF

[3259] Janet Turk Wittes. Commentary on randomization: The forgotten component of the randomized clinical trial. Statistics in Medicine, 38(1):

14–16, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Day:2019:RRF

[3260] Simon Day. Reflections on “Randomization: The Forgotten Component of the Randomized Clinical Trial”. Statistics in Medicine, 38(1):17–18, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Louis:2019:DRF

[3261] Thomas A. Louis. Discussion of “Randomization: The forgotten component of the randomized clinical trial”. Statistics in Medicine, 38(1): 19–22, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seidenfeld:2019:ACR

[3262] Teddy Seidenfeld. Ancillarity contra randomization as a basis for inference. Statistics in Medicine, 38(1):23–26, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenberger:2019:R

[3263] William F. Rosenberger, Diane Uschner, and Yanying Wang. Rejoinder. Statistics in Medicine, 38(1):27–30, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Follmann:2019:HBS

[3264] Dean Follmann, Erica Brittain, and Keith Lumbard. Half blind superiority tests for clinical trials of anti-infective drugs. Statistics in Medicine, 38(1):31–43, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hooper:2019:IEI

[3265] Richard Hooper and Charles Knowles. Improving the efficiency of individually randomized clinical trials by staggering the introduction of the intervention. Statistics in Medicine, 38(1):44–52, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finkelstein:2019:GWR

[3266] Dianne M. Finkelstein and David A. Schoenfeld. Graphing the win ratio and its components over time. Statistics in Medicine, 38(1):53–61, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2019:SIE

[3267] Mandi Yu, Benmei Liu, Yan Li, Zhaohui (Joe) Zou, and Nancy Breen. Statistical inferences of extended concentration indices for directly standardized rates. Statistics in Medicine, 38(1):62–73, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hubbard:2019:BLC

[3268] Rebecca A. Hubbard, Jing Huang, Joanna Harton, Arman Oganisian, Grace Choi, Levon Utidjian, Ihuoma Eneli, L. Charles Bailey, and Yong Chen. A Bayesian latent class approach for EHR-based phenotyping. Statistics in Medicine, 38(1):74–87, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scosyrev:2019:PAM

[3269] Emil Scosyrev and Ekkehard Glimm. Power analysis for multivariable Cox regression models. Statistics in Medicine, 39(2):88–99, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2019:ECP

[3270] Lu Wang and Ying Huang. Evaluating classification performance of biomarkers in two-phase case-control studies. Statistics in Medicine, 39 (2):100–114, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lian:2019:BAC

[3271] Qinshu Lian, James S. Hodges, Richard MacLehose, and Haitao Chu. A Bayesian approach for correcting exposure misclassification in metaanalysis. Statistics in Medicine, 39(2):115–130, January 15, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fellows:2019:RDS

[3272] Ian E. Fellows. Respondent-driven sampling and the homophily configuration graph. Statistics in Medicine, 39(2):131–150, January 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:2019:DAE

[3273] Paul S. Albert. Driving the analysis: an exciting opportunity for statistical innovation in driving research. Statistics in Medicine, 38(2):151, January 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dawson:2019:PSC

[3274] Jeffrey D. Dawson. Practical and statistical challenges in driving research. Statistics in Medicine, 38(2):152–159, January 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2019:SBM

[3275] Feng Guo, Inyoung Kim, and Sheila G. Klauer. Semiparametric Bayesian models for evaluating time-variant driving risk factors using naturalistic driving data and case-crossover approach. Statistics in Medicine, 38(2): 160–174, January 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:2019:IMN

[3276] Paul S. Albert. Innovative modeling of naturalistic driving data: Inference and prediction. Statistics in Medicine, 38(2):175–183, January 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bond:2019:TGB

[3277] Simon Bond. Theory of general balance applied to step wedge designs. Statistics in Medicine, 38(2):184–191, January 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2019:DMS

[3278] Xiaodong Luo, Xuezhou Mao, Xun Chen, Junshan Qiu, Steven Bai, and Hui Quan. Design and monitoring of survival trials in complex scenarios. Statistics in Medicine, 38(2):192–209, January 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brand:2019:CMI

[3279] Jaap Brand, Stef van Buuren, Saskia le Cessie, and Wilbert van den Hout. Combining multiple imputation and bootstrap in the analysis of cost-effectiveness trial data. Statistics in Medicine, 38(2):210–220, January 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Villain:2019:APB

[3280] Laura Villain, Daniel Commenges, Chloé Pasin, Mélanie Prague, and Rodolphe Thiébaut. Adaptive protocols based on predictions from a mechanistic model of the effect of IL7 on CD4 counts. Statistics in Medicine, 38(2):221–235, January 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiao:2019:AKM

[3281] Qian Xiao, Lin Wang, and Hongquan Xu. Application of kriging models for a drug combination experiment on lung cancer. Statistics in Medicine, 38(2):236–246, January 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sudell:2019:IOS

[3282] Maria Sudell, Ruwanthi Kolamunnage-Dona, François Gueyffier, and Catrin Tudur Smith. Investigation of one-stage meta-analysis methods for joint longitudinal and time-to-event data through simulation and real data application. Statistics in Medicine, 38(2):247–268, January 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rueten-Budde:2019:IHH

[3283] Anja J. Rueten-Budde, Hein Putter, and Marta Fiocco. Investigating hospital heterogeneity with a competing risks frailty model. Statistics in Medicine, 38(2):269–288, January 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wan:2019:MUA

[3284] Fei Wan. Matched or unmatched analyses with propensity-score-matched data? Statistics in Medicine, 38(2):289–300, January 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abberbock:2019:AES

[3285] Judah Abberbock, Stewart Anderson, Priya Rastogi, and Gong Tang. Assessment of effect size and power for survival analysis through a binary surrogate endpoint in clinical trials. Statistics in Medicine, 38(3):301– 314, February 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tao:2019:DRE

[3286] Yebin Tao and Haoda Fu. Doubly robust estimation of the weighted average treatment effect for a target population. Statistics in Medicine, 38(3):315–325, February 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2019:MAN

[3287] Ian R. White, Stephen Kaptoge, Patrick Royston, Willi Sauerbrei, and Emerging Risk Factors Collaboration. Meta-analysis of non-linear exposure-outcome relationships using individual participant data: a comparison of two methods. Statistics in Medicine, 38(3):326–338, February 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2019:EHT

[3288] Hang J. Kim, Bo Lu, Edward J. Nehus, and Mi-Ok Kim. Estimating heterogeneous treatment effects for latent subgroups in observational studies. Statistics in Medicine, 38(3):339–353, February 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qu:2019:CMA

[3289] Yongming Qu. Can a multiple ascending dose study serve as an informative proof-of-concept study? Statistics in Medicine, 38(3):354–362, February 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Happ:2019:OSS

[3290] Martin Happ, Arne C. Bathke, and Edgar Brunner. Optimal sample size planning for the Wilcoxon–Mann–Whitney test. Statistics in Medicine, 38(3):363–375, February 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2019:ADS

[3291] Peirong Xu, Youngjo Lee, Jian Qing Shi, and Janet Eyre. Automatic detection of significant areas for functional data with directional error control. Statistics in Medicine, 38(3):376–397, February 10, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2019:MMA

[3292] Qingzhao Yu, Xiaocheng Wu, Bin Li, and Richard A. Scribner. Multiple mediation analysis with survival outcomes: With an application to explore racial disparity in breast cancer survival. Statistics in Medicine, 38(3):398–412, February 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klein:2019:MBC

[3293] Nadja Klein, Thomas Kneib, Giampiero Marra, Rosalba Radice, Slawa Rokicki, and Mark E. McGovern. Mixed binary-continuous copula regression models with application to adverse birth outcomes. Statistics in Medicine, 38(3):413–436, February 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2019:ERO

[3294] Hui Xu, Jing Qian, Nina P. Paynter, Xuehong Zhang, Brian W. Whitcomb, Shelley S. Tworoger, Kathryn M. Rexrode, Susan E. Hankinson, and Raji Balasubramanian. Estimating the receiver operating characteristic curve in matched case control studies. Statistics in Medicine, 38 (3):437–451, February 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2019:CEL

[3295] Yanmei Xie and Biao Zhang. Constrained empirical-likelihood confidence regions in nonignorable covariate-missing data problems. Statistics in Medicine, 38(3):452–479, February 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gomes:2019:CSM

[3296] Manuel Gomes, Rosalba Radice, Jose Camarena Brenes, and Giampiero Marra. Copula selection models for non-Gaussian outcomes that are missing not at random. Statistics in Medicine, 38(3):480–496, February 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vickers:2019:CNR

[3297] Andrew J. Vickers. Comments on “Net reclassification index at event rate: Properties and relationships”. Statistics in Medicine, 38(3):497– 498, February 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2019:SNS

[3298] Michael J. Pencina, Ewout W. Steyerberg, and Ralph B. D’Agostino, Sr. Single-number summary and decision analytic measures can happily coexist. Statistics in Medicine, 38(3):499–500, February 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:2019:SRP

[3299] Paul S. Albert. Shared random parameter models: a legacy of the biostatistics program at the national heart, lung, and blood institute. Statistics in Medicine, 38(4):501–511, February 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gaynor:2019:MAC

[3300] Sheila M. Gaynor, Joel Schwartz, and Xihong Lin. Mediation analysis for common binary outcomes. Statistics in Medicine, 38(4):512–529, February 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2019:ITD

[3301] Song Yang. Improving testing and description of treatment effect in clinical trials with survival outcomes. Statistics in Medicine, 38(4):530– 544, February 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Loh:2019:SRT

[3302] Wei-Yin Loh, Michael Man, and Shuaicheng Wang. Subgroups from regression trees with adjustment for prognostic effects and postselection inference. Statistics in Medicine, 38(4):545–557, February 20, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ishwaran:2019:SEC

[3303] Hemant Ishwaran and Min Lu. Standard errors and confidence intervals for variable importance in random forest regression, classification, and survival. Statistics in Medicine, 38(4):558–582, February 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petersen:2019:DAA

[3304] Ashley Petersen and Daniela Witten. Data-adaptive additive modeling. Statistics in Medicine, 38(4):583–600, February 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2019:IMU

[3305] Song Yang. Interim monitoring using the adaptively weighted log-rank test in clinical trials for survival outcomes. Statistics in Medicine, 38 (4):601–612, February 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lai:2019:AED

[3306] Tze Leung Lai, Philip W. Lavori, and Ka Wai Tsang. Adaptive enrichment designs for confirmatory trials. Statistics in Medicine, 38(4): 613–624, February 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Caetano:2019:MIE

[3307] Samantha-Jo Caetano, David Dawe, Peter Ellis, Craig C. Earle, and Gregory R. Pond. Methods to improve the estimation of time-to-event outcomes when data is de-identified. Statistics in Medicine, 38(4):625– 635, February 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2019:PSS

[3308] Fan Li, Andrew B. Forbes, Elizabeth L. Turner, and John S. Preisser. Power and sample size requirements for GEE analyses of cluster randomized crossover trials. Statistics in Medicine, 38(4):636–649, February 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2019:AIT

[3309] Mimi Kim, Cuiling Wang, and Xiaonan Xue. Assessing the influence of treatment nonadherence on noninferiority trials using the tipping point approach. Statistics in Medicine, 38(4):650–659, February 20, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cursio:2019:LTS

[3310] John F. Cursio, Robin J. Mermelstein, and Donald Hedeker. Latent trait shared-parameter mixed models for missing ecological momentary assessment data. Statistics in Medicine, 38(4):660–673, February 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rover:2019:MAR

[3311] Christian Röver, Simon Wandel, and Tim Friede. Model averaging for robust extrapolation in evidence synthesis. Statistics in Medicine, 38 (4):674–694, February 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zapf:2019:WMB

[3312] Antonia Zapf, Marianne Huebner, Geraldine Rauch, and Meinhard Kieser. What makes a biostatistician? Statistics in Medicine, 38(4):695– 701, February 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weber:2019:QAB

[3313] Enya M. Weber and Andrew C. Titman. Quantifying the association between progression-free survival and overall survival in oncology trials using Kendall’s τ. Statistics in Medicine, 38(5):703–719, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mavridis:2019:AUD

[3314] Dimitris Mavridis, Georgia Salanti, Toshi A. Furukawa, Andrea Cipriani, Anna Chaimani, and Ian R. White. Allowing for uncertainty due to missing and LOCF imputed outcomes in meta-analysis. Statistics in Medicine, 38(5):720–737, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sheng:2019:ECE

[3315] Elisa Sheng, Wei Li, and Xiao-Hua Zhou. Estimating causal effects of treatment in RCTs with provider and subject noncompliance. Statistics in Medicine, 38(5):738–750, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2019:PSM

[3316] Peter C. Austin and Jason P. Fine. Propensity-score matching with competing risks in survival analysis. Statistics in Medicine, 38(5):751– 777, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seppa:2019:EMR

[3317] Karri Seppä, Håvard Rue, Timo Hakulinen, Esa Läärä, Mikko J. Sillanpää, and Janne Pitkäniemi. Estimating multilevel regional variation in excess mortality of cancer patients using integrated nested Laplace approximation. Statistics in Medicine, 38(5):778–791, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pham:2019:PCM

[3318] Tra My Pham, James R. Carpenter, Tim P. Morris, Angela M. Wood, and Irene Petersen. Population-calibrated multiple imputation for a binary/categorical covariate in categorical regression models. Statistics in Medicine, 38(5):792–808, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2019:WKM

[3319] Wei Zhang, Zhen Chen, Aiyi Liu, and Germaine M. Buck Louis. A weighted kernel machine regression approach to environmental pollutants and infertility. Statistics in Medicine, 38(5):809–827, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2019:BRM

[3320] Yu-Bo Wang, Zhen Chen, Jill M. Goldstein, Germaine M. Buck Louis, and Stephen E. Gilman. A Bayesian regularized mediation analysis with multiple exposures. Statistics in Medicine, 38(5):828–843, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2019:IUR

[3321] Kevin He, John D. Kalbfleisch, Yuan Yang, and Zhe Fei. Inter-unit reliability for nonlinear models. Statistics in Medicine, 38(5):844–854, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Araujo:2019:UWA

[3322] Joana Araújo, Elisabete Ramos, Gita D. Mishra, and Milton Severo. The use of weight adjusted for height rather than body mass index to assess growth trajectory: Results from a population-based cohort. Statistics in Medicine, 38(5):855–865, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Villanueva:2019:MDG

[3323] Nora M. Villanueva, Marta Sestelo, and Luís Meira-Machado. A method for determining groups in multiple survival curves. Statistics in Medicine, 38(5):866–877, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:2019:PVS

[3324] Eunyoung Park and Il Do Ha. Penalized variable selection for accelerated failure time models with random effects. Statistics in Medicine, 38(5): 878–892, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bracher:2019:CUR

[3325] Johannes Bracher. Comment on “Under-reported data analysis with INAR-hidden Markov chains”. Statistics in Medicine, 38(5):893–898, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fernandez-Fontelo:2019:RLU

[3326] Amanda Fernández-Fontelo, Alejandra Cabaña, Pedro Puig, and David Moriña. Response to the letter of “Under-reported data analysis with INAR-hidden Markov chains”. Statistics in Medicine, 38(5):899–900, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2019:CLC

[3327] Stuart G. Baker, Barnett S. Kramer, and Karen S. Lindeman. Correction to “Latent class instrumental variables: a clinical and biostatistical perspective”. Statistics in Medicine, 38(5):901, February 28, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2019:IIa

[3328] Anonymous. Issue information. Statistics in Medicine, 38(6):1–4, March 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bible:2019:PIU

[3329] Joe Bible, Paul S. Albert, Bruce G. Simons-Morton, and Danping Liu. Practical issues in using generalized estimating equations for inference on transitions in longitudinal data: What is being estimated? Statistics in Medicine, 38(6):903–916, March 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2019:MBC

[3330] L. Tian, F. Jiang, T. Hasegawa, H. Uno, M. Pfeffer, and Lj. Wei. Moving beyond the conventional stratified analysis to estimate an overall treatment efficacy with the data from a comparative randomized clinical study. Statistics in Medicine, 38(6):917–932, March 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cui:2019:EAS

[3331] Lu Cui and Lanju Zhang. On the efficiency of adaptive sample size design. Statistics in Medicine, 38(6):933–944, March 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lan:2019:SMP

[3332] Yu Lan, Gong Tang, and Daniel F. Heitjan. Statistical modeling and prediction of clinical trial recruitment. Statistics in Medicine, 38(6): 945–955, March 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dharmarajan:2019:SSE

[3333] Sai Dharmarajan, Joo-Yeon Lee, and Rima Izem. Sample size estimation for case-crossover studies. Statistics in Medicine, 38(6):956–968, March 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McGrath:2019:OSA

[3334] Sean McGrath, XiaoFei Zhao, Zhi Zhen Qin, Russell Steele, and Andrea Benedetti. One-sample aggregate data meta-analysis of medians. Statistics in Medicine, 38(6):969–984, March 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shapland:2019:BAM

[3335] Chin Yang Shapland, John R. Thompson, and Nuala A. Sheehan. A Bayesian approach to Mendelian randomisation with dependent instruments. Statistics in Medicine, 38(6):985–1001, March 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McLain:2019:PIP

[3336] Alexander C. McLain, Edward A. Frongillo, Juan Feng, and Elaine Borghi. Prediction intervals for penalized longitudinal models with multisource summary measures: an application to childhood malnutrition. Statistics in Medicine, 38(6):1002–1012, March 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OMalley:2019:MBR

[3337] A. James O’Malley, Peter James, Todd A. MacKenzie, Jinyoung Byun, S. V. Subramanian, and Jason P. Block. Modeling a bivariate residentialworkplace neighborhood effect when estimating the effect of proximity to fast-food establishments on body mass index. Statistics in Medicine, 38(6):1013–1035, March 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tawiah:2019:MMR

[3338] Richard Tawiah, Kelvin K. W. Yau, Geoffrey J. McLachlan, Suzanne K. Chambers, and Shu-Kay Ng. Multilevel model with random effects for clustered survival data with multiple failure outcomes. Statistics in Medicine, 38(6):1036–1055, March 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bartolucci:2019:SPC

[3339] Francesco Bartolucci and Alessio Farcomeni. A shared-parameter continuous-time hidden Markov and survival model for longitudinal data with informative dropout. Statistics in Medicine, 38(6):1056–1073, March 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachos:2019:FLL

[3340] Victor H. Lachos, Larissa A. Matos, Luis M. Castro, and Ming-Hui Chen. Flexible longitudinal linear mixed models for multiple censored responses data. Statistics in Medicine, 38(6):1074–1102, March 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2019:IIb

[3341] Anonymous. Issue information. Statistics in Medicine, 38(7):1–4, March 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grayling:2019:AMS

[3342] Michael J. Grayling, Adrian P. Mander, and James M. S. Wason. Admissible multiarm stepped-wedge cluster randomized trial designs. Statistics in Medicine, 38(7):1103–1119, March 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2019:CBD

[3343] Rui Tang, Jing Shen, and Ying Yuan. ComPAS: a Bayesian drug combination platform trial design with adaptive shrinkage. Statistics in Medicine, 38(7):1120–1134, March 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gabriel:2019:OEB

[3344] Erin E. Gabriel, Michael C. Sachs, Michael J. Daniels, and M. Elizabeth Halloran. Optimizing and evaluating biomarker combinations as trial-level general surrogates. Statistics in Medicine, 38(7):1135–1146, March 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Banbeta:2019:MPP

[3345] Akalu Banbeta, Joost van Rosmalen, David Dejardin, and Emmanuel Lesaffre. Modified power prior with multiple historical trials for binary endpoints. Statistics in Medicine, 38(7):1147–1169, March 30, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2019:BDP

[3346] Chenguang Wang, Gary L. Rosner, and Richard B. S. Roden. A Bayesian design for phase I cancer therapeutic vaccine trials. Statistics in Medicine, 38(7):1170–1189, March 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Daniels:2019:NCI

[3347] Michael J. Daniels and Xuan Luo. A note on compatibility for inference with missing data in the presence of auxiliary covariates. Statistics in Medicine, 38(7):1190–1199, March 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2019:ICO

[3348] Yifan Sun, Yu Jiang, Yang Li, and Shuangge Ma. Identification of cancer omics commonality and difference via community fusion. Statistics in Medicine, 38(7):1200–1212, March 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2019:MAC

[3349] Jian Wang, Jing Ning, and Sanjay Shete. Mediation analysis in a casecontrol study when the mediator is a censored variable. Statistics in Medicine, 38(7):1213–1229, March 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2019:PDA

[3350] Tianzhong Yang, Han Chen, Hongwei Tang, Donghui Li, and Peng Wei. A powerful and data-adaptive test for rare-variant-based geneenvironment interaction analysis. Statistics in Medicine, 38(7):1230– 1244, March 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peskoe:2019:TNI

[3351] S. B. Peskoe, D. Spiegelman, and M. Wang. There is no impact of exposure measurement error on latency estimation in linear models. Statistics in Medicine, 38(7):1245–1261, March 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2019:MSSa

[3352] Richard D. Riley, Kym I. E. Snell, Joie Ensor, Danielle L. Burke, Frank E. Harrell, Jr., Karel G. M. Moons, and Gary S. Collins. Minimum sample size for developing a multivariable prediction model: Part I — continuous outcomes. Statistics in Medicine, 38(7):1262–1275, March 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2019:MSSb

[3353] Richard D. Riley, Kym I. E. Snell, Joie Ensor, Danielle L. Burke, Frank E. Harrell, Jr., Karel G. M. Moons, and Gary S. Collins. Minimum sample size for developing a multivariable prediction model: Part II — binary and time-to-event outcomes. Statistics in Medicine, 38(7):1276–1296, March 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Castellares:2019:CSS

[3354] Fredy Castellares, Marcos O. Prates, and Ali Abolhassani. Comments on “A spatial scan statistic for compound Poisson data”. Statistics in Medicine, 38(30):1297–1299, March 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Raseta:2019:CGA

[3355] Marko Raseta and Alina Bazarova. Comments on “A general approach for sample size calculation for the three-arm ‘gold standard’ noninferiority design”. Statistics in Medicine, 38(30):1300–1302, March 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sloan:2019:DAC

[3356] Abigail Sloan, Yue Song, Mitchell H. Gail, Rebecca Betensky, Bernard Rosner, Regina G. Ziegler, Stephanie A. Smith-Warner, and Molin Wang. Design and analysis considerations for combining data from multiple biomarker studies. Statistics in Medicine, 38(8):1303–1320, April 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turner:2019:IEE

[3357] Rebecca M. Turner, Clara P. Domínguez-Islas, Dan Jackson, Kirsty M. Rhodes, and Ian R. White. Incorporating external evidence on betweentrial heterogeneity in network meta-analysis. Statistics in Medicine, 38 (8):1321–1335, April 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mathur:2019:NMM

[3358] Maya B. Mathur and Tyler J. VanderWeele. New metrics for metaanalyses of heterogeneous effects. Statistics in Medicine, 38(8):1336– 1342, April 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Su:2019:SAR

[3359] Pei-Fang Su, Junjiang Zhong, and Huang-Tz Ou. Semiparametric additive rates model for recurrent events data with intermittent gaps. Statistics in Medicine, 38(8):1343–1356, April 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qian:2019:LBA

[3360] Jing Qian, Evan Ray, Regina L. Brecha, Muredach P. Reilly, and Andrea S. Foulkes. A likelihood-based approach to transcriptome association analysis. Statistics in Medicine, 38(8):1357–1373, April 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2019:CBR

[3361] Zhen Chen, Beom Seuk Hwang, and Sungduk Kim. A correlated Bayesian rank likelihood approach to multiple ROC curves for endometriosis. Statistics in Medicine, 38(8):1374–1385, April 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lyu:2019:NST

[3362] Ting Lyu, Zhiliang Ying, and Hong Zhang. A new semiparametric transformation approach to disease diagnosis with multiple biomarkers. Statistics in Medicine, 38(8):1386–1398, April 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gabrio:2019:FBM

[3363] Andrea Gabrio, Alexina J. Mason, and Gianluca Baio. A full Bayesian model to handle structural ones and missingness in economic evaluations from individual-level data. Statistics in Medicine, 38(8):1399–1420, April 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones-Todd:2019:IPS

[3364] Charlotte M. Jones-Todd, Peter Caie, Janine B. Illian, Ben C. Stevenson, Anne Savage, David J. Harrison, and James L. Bown. Identifying prognostic structural features in tissue sections of colon cancer patients using point pattern analysis. Statistics in Medicine, 38(8):1421–1441, April 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qin:2019:UMS

[3365] Jing Qin, Tao Yu, Pengfei Li, Hao Liu, and Baojiang Chen. Using a monotone single-index model to stabilize the propensity score in missing data problems and causal inference. Statistics in Medicine, 38(8):1442– 1458, April 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quintana:2019:BMD

[3366] M. Quintana, J. Shrader, C. Slota, G. Joe, J. C. McKew, M. Fitzgerald, W. A. Gahl, S. Berry, and N. Carrillo. Bayesian model of disease progression in GNE myopathy. Statistics in Medicine, 38(8):1459–1474, April 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2019:CIU

[3367] Jihnhee Yu, Albert Vexler, and Kabir Jalal. A critical issue of using the variance of the total in the linearization method — in the context of unequal probability sampling. Statistics in Medicine, 38(8):1475–1483, April 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neve:2019:SLT

[3368] Jan De Neve, Olivier Thas, and Thomas A. Gerds. Semiparametric linear transformation models: Effect measures, estimators, and applications. Statistics in Medicine, 38(8):1484–1501, April 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qin:2019:SMB

[3369] Gengsheng Qin. Statistics in Medicine book review. Statistics in Medicine, 38(8):1502, April 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Asendorf:2019:SSR

[3370] Thomas Asendorf, Robin Henderson, Heinz Schmidli, and Tim Friede. Sample size re-estimation for clinical trials with longitudinal negative binomial counts including time trends. Statistics in Medicine, 38(9): 1503–1528, April 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bao:2019:ARS

[3371] Yanchun Bao, Paul S. Clarke, Melissa Smart, and Meena Kumari. Assessing the robustness of sisVIVE in a Mendelian randomization study to estimate the causal effect of body mass index on income using multiple SNPs from understanding society. Statistics in Medicine, 38(9): 1529–1542, April 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neelon:2019:MDF

[3372] Brian Neelon, Azza Shoaibi, and Sara E. Benjamin-Neelon. A multivariate discrete failure time model for the analysis of infant motor development. Statistics in Medicine, 38(9):1543–1557, April 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brookmeyer:2019:MML

[3373] Ron Brookmeyer and Nada Abdalla. Multistate models and lifetime risk estimation: Application to Alzheimer’s disease. Statistics in Medicine, 38(9):1558–1565, April 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cho:2019:MAC

[3374] Shu-Hsien Cho and Yen-Tsung Huang. Mediation analysis with causally ordered mediators using Cox proportional hazards model. Statistics in Medicine, 38(9):1566–1581, April 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Narisetty:2019:SNI

[3375] Naveen N. Narisetty, Bhramar Mukherjee, Yin-Hsiu Chen, Richard Gonzalez, and John D. Meeker. Selection of nonlinear interactions by a forward stepwise algorithm: Application to identifying environmental chemical mixtures affecting health outcomes. Statistics in Medicine, 38(9): 1582–1600, April 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deJong:2019:SSC

[3376] Valentijn M. T. de Jong, Marinus J. C. Eijkemans, Ben van Calster, Dirk Timmerman, Karel G. M. Moons, Ewout W. Steyerberg, and Maarten van Smeden. Sample size considerations and predictive performance of multinomial logistic prediction models. Statistics in Medicine, 38(9): 1601–1619, April 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2019:IGE

[3377] Xiaoyan Wang, Yonghong Xu, and Shuangge Ma. Identifying geneenvironment interactions incorporating prior information. Statistics in Medicine, 38(9):1620–1633, April 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2019:BAG

[3378] Kai Kang, Xinyuan Song, X. Joan Hu, and Hongtu Zhu. Bayesian adaptive group lasso with semiparametric hidden Markov models. Statistics in Medicine, 38(9):1634–1650, April 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grigg:2019:SCS

[3379] Olivia Aj Grigg. The STRAND chart: a survival time control chart. Statistics in Medicine, 38(9):1651–1661, April 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2019:BHM

[3380] Tianyi Cai and Alan M. Zaslavsky. Bayesian hierarchical modeling of substate area estimates from the medicare CAHPS survey. Statistics in Medicine, 38(9):1662–1677, April 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2019:ISD

[3381] Seonjin Kim, Hyunkeun Ryan Cho, and Xianyang Zhang. Initial severitydependent longitudinal model with application to a randomized controlled trial of women with depression. Statistics in Medicine, 38(9): 1678–1689, April 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alam:2019:SPS

[3382] Shomoita Alam, Erica E. M. Moodie, and David A. Stephens. Should a propensity score model be super? The utility of ensemble procedures for causal adjustment. Statistics in Medicine, 38(9):1690–1702, April 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2019:MLM

[3383] Zhiwei Zhang and Shujie Ma. Machine learning methods for leveraging baseline covariate information to improve the efficiency of clinical trials. Statistics in Medicine, 38(10):1703–1714, May 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2019:MDA

[3384] Yongqiang Tang. A monotone data augmentation algorithm for multivariate nonnormal data: With applications to controlled imputations for longitudinal trials. Statistics in Medicine, 38(10):1715–1733, May 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cruz:2019:AHC

[3385] Maricela Cruz, Daniel L. Gillen, Miriam Bender, and Hernando Ombao. Assessing health care interventions via an interrupted time series model: Study power and design considerations. Statistics in Medicine, 38(10): 1734–1752, May 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saint-Hilary:2019:PPS

[3386] Gaelle Saint-Hilary, Valentine Barboux, Matthieu Pannaux, Mauro Gasparini, Veronique Robert, and Gianluca Mastrantonio. Predictive probability of success using surrogate endpoints. Statistics in Medicine, 38(10): 1753–1774, May 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ying:2019:TSR

[3387] Andrew Ying, Ronghui Xu, and James Murphy. Two-stage residual inclusion for survival data and competing risks — an instrumental variable approach with application to SEER-Medicare linked data. Statistics in Medicine, 38(10):1775–1801, May 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grand:2019:JMD

[3388] Mia Klinten Grand, Koenraad Arndt Vermeer, Tom Missotten, and Hein Putter. A joint model for dynamic prediction in uveitis. Statistics in Medicine, 38(10):1802–1816, May 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Innocenti:2019:RET

[3389] Francesco Innocenti, Math J. J. M. Candel, Frans E. S. Tan, and Gerard J. P. van Breukelen. Relative efficiencies of two-stage sampling schemes for mean estimation in multilevel populations when cluster size is informative. Statistics in Medicine, 38(10):1817–1834, May 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shu:2019:WCI

[3390] Di Shu and Grace Y. Yi. Weighted causal inference methods with mismeasured covariates and misclassified outcomes. Statistics in Medicine, 38(10):1835–1854, May 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barrett:2019:EAB

[3391] Jessica K. Barrett, Raphael Huille, Richard Parker, Yuichiro Yano, and Michael Griswold. Estimating the association between blood pressure variability and cardiovascular disease: an application using the ARIC study. Statistics in Medicine, 38(10):1855–1868, May 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gayawan:2019:MSV

[3392] Ezra Gayawan, Samson B. Adebayo, and Elisabeth Waldmann. Modeling the spatial variability in the spread and correlation of childhood malnutrition in Nigeria. Statistics in Medicine, 38(10):1869–1890, May 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Santacatterina:2019:OPW

[3393] Michele Santacatterina, Celia García-Pareja, Rino Bellocco, Anders Sönnerborg, Anna Mia Ekström, and Matteo Bottai. Optimal probability weights for estimating causal effects of time-varying treatments with marginal structural Cox models. Statistics in Medicine, 38(10): 1891–1902, May 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2019:SMB

[3394] Michael J. Pencina, Chirag R. Parikh, Paul L. Kimmel, Nancy R. Cook, Josef Coresh, Harold I. Feldman, Andrea Foulkes, Phyllis A. Gimotty, Chi yuan Hsu, Kevin Lemley, Peter Song, Kenneth Wilkins, Daniel R. Gossett, Yining Xie, and Robert A. Star. Statistical methods for building better biomarkers of chronic kidney disease. Statistics in Medicine, 38 (11):1903–1917, May 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grantham:2019:ADC

[3395] Kelsey L. Grantham, Jessica Kasza, Stephane Heritier, Karla Hemming, and Andrew B. Forbes. Accounting for a decaying correlation structure in cluster randomized trials with continuous recruitment. Statistics in Medicine, 38(11):1918–1934, May 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roig:2019:NAS

[3396] Marta Bofill Roig and Guadalupe Gómez Melis. A new approach for sizing trials with composite binary endpoints using anticipated marginal values and accounting for the correlation between components. Statistics in Medicine, 38(11):1935–1956, May 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2019:ECS

[3397] Kevin He, Valarie B. Ashby, and Douglas E. Schaubel. Evaluating centerspecific long-term outcomes through differences in mean survival time: Analysis of national kidney transplant data. Statistics in Medicine, 38 (11):1957–1967, May 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2019:CBT

[3398] Hong Li and Constantine Gatsonis. Combining biomarker trajectories to improve diagnostic accuracy in prospective cohort studies with verification bias. Statistics in Medicine, 38(11):1968–1990, May 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pugh:2019:ESO

[3399] Sierra Pugh, Matthew J. Heaton, Brian Hartman, Candace Berrett, Chantel Sloan, Amber M. Evans, Tebeb Gebretsadik, Pingsheng Wu, Tina V. Hartert, and Rees L. Lee. Estimating seasonal onsets and peaks of bronchiolitis with spatially and temporally uncertain data. Statistics in Medicine, 38(11):1991–2001, May 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2019:CUE

[3400] Yizhen Xu, Tao Liu, Michael J. Daniels, Rami Kantor, Ann Mwangi, and Joseph W. Hogan. Classification using ensemble learning under weighted misclassification loss. Statistics in Medicine, 38(11):2002–2012, May 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nguyen:2019:UPS

[3401] Tri-Long Nguyen and Thomas P. A. Debray. The use of prognostic scores for causal inference with general treatment regimes. Statistics in Medicine, 38(11):2013–2029, May 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2019:EDL

[3402] Xuelin Huang, Lei Liu, Jing Ning, Liang Li, and Yu Shen. Estimation of the distribution of longitudinal biomarker trajectories prior to disease progression. Statistics in Medicine, 38(11):2030–2046, May 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hall:2019:MCT

[3403] Lauren M. Hall and Joshua P. French. A modified CUSUM test to control postoutbreak false alarms. Statistics in Medicine, 38(11):2047–2058, May 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiao:2019:RRO

[3404] W. Xiao, H. H. Zhang, and W. Lu. Robust regression for optimal individualized treatment rules. Statistics in Medicine, 38(11):2059–2073, May 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morris:2019:USS

[3405] Tim P. Morris, Ian R. White, and Michael J. Crowther. Using simulation studies to evaluate statistical methods. Statistics in Medicine, 38(11): 2074–2102, May 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2019:ACI

[3406] Chi Hyun Lee, Jing Ning, Richard J. Kryscio, and Yu Shen. Analysis of combined incident and prevalent cohort data under a proportional mean residual life model. Statistics in Medicine, 38(12):2103–2114, May 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2019:SDB

[3407] Li Zou, Albert Vexler, Jihnhee Yu, and Hongzhi Wan. A sequential density-based empirical likelihood ratio test for treatment effects. Statistics in Medicine, 38(12):2115–2125, May 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Silva:2019:ASH

[3408] Ivair R. Silva, Wilson M. Lopes, Philipe Dias, and W. Katherine Yih. Alpha spending for historical versus surveillance Poisson data with CMaxSPRT. Statistics in Medicine, 38(12):2126–2138, May 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2019:ELM

[3409] Hong Wang and Gang Li. Extreme learning machine Cox model for highdimensional survival analysis. Statistics in Medicine, 38(12):2139–2156, May 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kawamori:2019:SET

[3410] Ai Kawamori, Keiichi Fukaya, Masumi Kitazawa, and Makio Ishiguro. A self-excited threshold autoregressive state-space model for menstrual cycles: Forecasting menstruation and identifying within-cycle stages based on basal body temperature. Statistics in Medicine, 38(12):2157–2170, May 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Irimata:2019:PGL

[3411] Kyle M. Irimata, Jennifer Broatch, and Jeffrey R. Wilson. Partitioned GMM logistic regression models for longitudinal data. Statistics in Medicine, 38(12):2171–2183, May 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Adhikari:2019:HDL

[3412] Samrachana Adhikari, Fabrizio Lecci, James T. Becker, Brian W. Junker, Lewis H. Kuller, Oscar L. Lopez, and Ryan J. Tibshirani. High-dimensional longitudinal classification with the multinomial fused lasso. Statistics in Medicine, 38(12):2184–2205, May 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Knoth:2019:RAC

[3413] Sven Knoth, Philipp Wittenberg, and Fah Fatt Gan. Risk-adjusted CUSUM charts under model error. Statistics in Medicine, 38(12):2206– 2218, May 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Du:2019:SPM

[3414] Mingyue Du, Tao Hu, and Jianguo Sun. Semiparametric probit model for informative current status data. Statistics in Medicine, 38(12):2219– 2227, May 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boulet:2019:BVS

[3415] Sandrine Boulet, Moreno Ursino, Peter Thall, Anne-Sophie Jannot, and Sarah Zohar. Bayesian variable selection based on clinical relevance weights in small sample studies — application to colon cancer. Statistics in Medicine, 38(12):2228–2247, May 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martin:2019:MMA

[3416] Ivonne Martin, Hae-Won Uh, Taniawati Supali, Makedonka Mitreva, and Jeanine J. Houwing-Duistermaat. The mixed model for the analysis of a repeated-measurement multivariate count data. Statistics in Medicine, 38(12):2248–2268, May 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baart:2019:JML

[3417] Sara J. Baart, Eric Boersma, and Dimitris Rizopoulos. Joint models for longitudinal and time-to-event data in a case-cohort design. Statistics in Medicine, 38(12):2269–2281, May 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2019:SBH

[3418] Bei Wang, Yi Zheng, Di Fang, Yiannis Kamarianakis, and Jeffrey R. Wilson. Split bootstrap hierarchical modeling of antibiotics abuse in China. Statistics in Medicine, 38(12):2282–2291, May 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proschan:2019:RRT

[3419] Michael A. Proschan and Lori E. Dodd. Re-randomization tests in clinical trials. Statistics in Medicine, 38(12):2292–2302, May 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2019:SAA

[3420] Kwonsang Lee, Scott A. Lorch, and Dylan S. Small. Sensitivity analyses for average treatment effects when outcome is censored by death in instrumental variable models. Statistics in Medicine, 38(13):2303–2316, June 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2019:OTS

[3421] Stephen D. Walter, Robin M. Turner, and Petra Macaskill. Optimising the two-stage randomised trial design when some participants are indifferent in their treatment preferences. Statistics in Medicine, 38(13): 2317–2331, June 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ha:2019:AHI

[3422] Neung Soo Ha and Joseph Sedransk. Assessing health insurance coverage in Florida using the behavioral risk factor surveillance system. Statistics in Medicine, 38(13):2332–2352, June 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2019:GBS

[3423] Zhongxue Chen and Kai Wang. Gene-based sequential burden association test. Statistics in Medicine, 38(13):2353–2363, June 15, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fan:2019:AGM

[3424] Xinyan Fan, Kuangnan Fang, Shuangge Ma, Shuaichao Wang, and Qingzhao Zhang. Assisted graphical model for gene expression data analysis. Statistics in Medicine, 38(13):2364–2380, June 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2019:ESS

[3425] Chunling Wang and Timothy E. Hanson. Estimation of sensitivity and specificity of multiple repeated binary tests without a gold standard. Statistics in Medicine, 38(13):2381–2390, June 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Khan:2019:RMB

[3426] Manzoor Khan and Jake Olivier. Regression to the mean for the bivariate binomial distribution. Statistics in Medicine, 38(13):2391–2412, June 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Welchowski:2019:CAR

[3427] Thomas Welchowski, Verena Zuber, and Matthias Schmid. Correlationadjusted regression survival scores for high-dimensional variable selection. Statistics in Medicine, 38(13):2413–2427, June 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Caniglia:2019:ETJ

[3428] Ellen C. Caniglia, James M. Robins, Lauren E. Cain, Caroline Sabin, Roger Logan, Sophie Abgrall, Michael J. Mugavero, Sonia Hernández-Díaz, Laurence Meyer, Remonie Seng, Daniel R. Drozd, George R. Seage III, Fabrice Bonnet, Fabien Le Marec, Richard D. Moore, Peter Reiss, Ard van Sighem, William C. Mathews, Inma Jarrín, Belén Alejos, Steven G. Deeks, Roberto Muga, Stephen L. Boswell, Elena Ferrer, Joseph J. Eron, John Gill, Antonio Pacheco, Beatriz Grinsztejn, Sonia Napravnik, Sophie Jose, Andrew Phillips, Amy Justice, Janet Tate, Heiner C. Bucher, Matthias Egger, Hansjakob Furrer, Jose M.

Miro, Jordi Casabona, Kholoud Porter, Giota Touloumi, Heidi Crane, Dominique Costagliola, Michael Saag, and Miguel A. Hernán. Emulating a trial of joint dynamic strategies: an application to monitoring and treatment of HIV-positive individuals. Statistics in Medicine, 38(13): 2428–2446, June 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Capistrano:2019:BEA

[3429] Estelina S. M. Capistrano, Erica E. M. Moodie, and Alexandra M. Schmidt. Bayesian estimation of the average treatment effect on the treated using inverse weighting. Statistics in Medicine, 38(13):2447–2466, June 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2019:IAI

[3430] Sheng-Hsuan Lin, Yen-Tsung Huang, and Hwai-I Yang. On identification of agonistic interaction: Hepatitis B and C interaction on hepatocellular carcinoma. Statistics in Medicine, 38(13):2467–2476, June 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2019:ECA

[3431] Jialiang Li, Ming Gao, and Ralph D’Agostino. Evaluating classification accuracy for modern learning approaches. Statistics in Medicine, 38(13): 2477–2503, June 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bretz:2019:C

[3432] Frank Bretz. Correction. Statistics in Medicine, 38(13):2504, June 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leahy:2019:ISA

[3433] Joy Leahy, Howard Thom, Jeroen P. Jansen, Emma Gray, Aisling O’Leary, Arthur White, and Cathal Walsh. Incorporating single-arm evidence into a network meta-analysis using aggregate level matching: Assessing the impact. Statistics in Medicine, 38(14):2505–2523, June 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2019:RTP

[3434] S. D. Walter, G. H. Guyatt, D. Bassler, M. Briel, T. Ramsay, and H. D. Han. Randomised trials with provision for early stopping for benefit (or harm): The impact on the estimated treatment effect. Statistics in Medicine, 38(14):2524–2543, June 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mondol:2019:BRS

[3435] Momenul Haque Mondol and M. Shafiqur Rahman. Bias-reduced and separation-proof GEE with small or sparse longitudinal binary data. Statistics in Medicine, 38(14):2544–2560, June 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tanniou:2019:LEP

[3436] Julien Tanniou, Sanne C. Smid, Ingeborg van der Tweel, Steven Teerenstra, and Kit C. B. Roes. Level of evidence for promising subgroup findings: The case of trends and multiple subgroups. Statistics in Medicine, 38(14):2561–2572, June 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2019:BNC

[3437] Chenguang Wang and Gary L. Rosner. A Bayesian nonparametric causal inference model for synthesizing randomized clinical trial and real-world evidence. Statistics in Medicine, 38(14):2573–2588, June 30, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hayashi:2019:PID

[3438] Kenichi Hayashi and Shinto Eguchi. The power-integrated discriminant improvement: an accurate measure of the incremental predictive value of additional biomarkers. Statistics in Medicine, 38(14):2589–2604, June 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2019:MBM

[3439] Jingli Wang, Jialiang Li, Yaguang Li, and Weng Kee Wong. A modelbased multithreshold method for subgroup identification. Statistics in Medicine, 38(14):2605–2631, June 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qiu:2019:CIT

[3440] Xin Qiu and Yuanjia Wang. Composite interaction tree for simultaneous learning of optimal individualized treatment rules and subgroups. Statistics in Medicine, 38(14):2632–2651, June 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Menssen:2019:PIO

[3441] Max Menssen and Frank Schaarschmidt. Prediction intervals for overdispersed binomial data with application to historical controls. Statistics in Medicine, 38(14):2652–2663, June 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2019:OEC

[3442] Sungim Lee and Johan Lim. Online estimation of the case fatality rate using a run-off triangle data approach: an application to the Korean MERS outbreak in 2015. Statistics in Medicine, 38(14):2664–2679, June 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3656].

Ballout:2019:SEB

[3443] Nadim Ballout and Vivian Viallon. Structure estimation of binary graphical models on stratified data: Application to the description of injury tables for victims of road accidents. Statistics in Medicine, 38(14):2680– 2703, June 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barthel:2019:CDM

[3444] Friederike M-S Barthel, Abdel Babiker, Patrick Royston, and Mahesh Kb Parmar. Comments on design and monitoring of survival trials in complex scenarios. Statistics in Medicine, 38(14):2704, June 30, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:2019:MVS

[3445] Ben Brown, Timothy Weaver, and Julian Wolfson. MEBoost: Variable selection in the presence of measurement error. Statistics in Medicine, 38(15):2705–2718, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hjellvik:2019:AUC

[3446] Vidar Hjellvik, Marie L. De Bruin, Sven O. Samuelsen, Øystein Karlstad, Morten Andersen, Jari Haukka, Peter Vestergaard, Frank de Vries, and Kari Furu. Adjusting for unmeasured confounding using validation data: Simplified two-stage calibration for survival and dichotomous outcomes. Statistics in Medicine, 38(15):2719–2734, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diaz:2019:SID

[3447] Iván Díaz. Statistical inference for data-adaptive doubly robust estimators with survival outcomes. Statistics in Medicine, 38(15):2735–2748, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2019:ODM

[3448] Kim May Lee, Stefanie Biedermann, and Robin Mitra. D-optimal designs for multiarm trials with dropouts. Statistics in Medicine, 38(15):2749– 2766, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2019:TPM

[3449] Wenhui Liu, Gary K. Grunwald, and P. Michael Ho. Two-part models for cost with zeros to decompose effects of covariates on probability of cost, mean nonzero cost, and mean total cost. Statistics in Medicine, 38(15):2767–2782, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2019:BLI

[3450] Ching-Yun Wang and James Dai. Best linear inverse probability weighted estimation for two-phase designs and missing covariate regression. Statistics in Medicine, 38(15):2783–2796, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anoke:2019:ATE

[3451] Sarah C. Anoke, Sharon-Lise Normand, and Corwin M. Zigler. Approaches to treatment effect heterogeneity in the presence of confounding. Statistics in Medicine, 38(15):2797–2815, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:2019:ODH

[3452] Tao Feng, Pallavi Basu, Wenguang Sun, Hsun Teresa Ku, and Wendy J. Mack. Optimal design for high-throughput screening via false discovery rate control. Statistics in Medicine, 38(15):2816–2827, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2019:EAT

[3453] Xiaofang Yan, Younathan Abdia, Somnath Datta, K. B. Kulasekera, Beatrice Ugiliweneza, Maxwell Boakye, and Maiying Kong. Estimation of average treatment effects among multiple treatment groups by using an ensemble approach. Statistics in Medicine, 38(15):2828–2846, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Humphrey:2019:MLI

[3454] Colman H. Humphrey, Dylan S. Small, Shane T. Jensen, Kevin G. Volpp, David A. Asch, Jingsan Zhu, and Andrea B. Troxel. Modeling lottery incentives for daily adherence. Statistics in Medicine, 38(15):2847–2867, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheung:2019:PCI

[3455] Li C. Cheung, Qing Pan, Noorie Hyun, and Hormuzd A. Katki. Prioritized concordance index for hierarchical survival outcomes. Statistics in Medicine, 38(15):2868–2882, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zang:2019:BAM

[3456] Yong Zang, Beibei Guo, Yan Han, Sha Cao, and Chi Zhang. A Bayesian adaptive marker-stratified design for molecularly targeted agents with customized hierarchical modeling. Statistics in Medicine, 38(15):2883– 2896, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Deng:2019:TER

[3457] Ai Deng. On the test of exclusion restriction in linear instrumental variable regressions. Statistics in Medicine, 38(15):2897–2900, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cole:2019:RWG

[3458] T. J. Cole. Relating weight growth trajectory to height and age. Statistics in Medicine, 38(15):2901–2902, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Araujo:2019:RCU

[3459] Joana Araújo, Elisabete Ramos, Gita D. Mishra, and Milton Severo. Response to comments on “The use of weight adjusted for height rather than body mass index to assess growth trajectory: Results from a populationbased cohort”. Statistics in Medicine, 38(15):2903–2904, July 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sverdlov:2019:IUR

[3460] Oleksandr Sverdlov and Yevgen Ryeznik. Implementing unequal randomization in clinical trials with heterogeneous treatment costs. Statistics in Medicine, 38(16):2905–2927, July 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sofeu:2019:OSV

[3461] Casimir Ledoux Sofeu, Takeshi Emura, and Virginie Rondeau. Onestep validation method for surrogate endpoints using data from multiple randomized cancer clinical trials with failure-time endpoints. Statistics in Medicine, 38(16):2928–2942, July 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Katki:2019:QRS

[3462] Hormuzd A. Katki. Quantifying risk stratification provided by diagnostic tests and risk predictions: Comparison to AUC and decision curve analysis. Statistics in Medicine, 38(16):2943–2955, July 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kawabata:2019:TSA

[3463] Takanori Kawabata, Ryo Emoto, Jo Nishino, Kunihiko Takahashi, and Shigeyuki Matsui. Two-stage analysis for selecting fixed numbers of features in omics association studies. Statistics in Medicine, 38(16):2956– 2971, July 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lv:2019:QRE

[3464] Yang Lv, Guoyou Qin, Zhongyi Zhu, and Dongsheng Tu. Quantile regression and empirical likelihood for the analysis of longitudinal data with monotone missing responses due to dropout, with applications to quality of life measurements from clinical trials. Statistics in Medicine, 38(16):2972–2991, July 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Efthimiou:2019:NMA

[3465] Orestis Efthimiou, Gerta Rücker, Guido Schwarzer, Julian P. T. Higgins, Matthias Egger, and Georgia Salanti. Network meta-analysis of rare events using the Mantel–Haenszel method. Statistics in Medicine, 38 (16):2992–3012, July 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2019:STB

[3466] Shu Jiang and Richard J. Cook. Score tests based on a finite mixture model of Markov processes under intermittent observation. Statistics in Medicine, 38(16):3013–3025, July 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2019:VSS

[3467] Liuquan Sun, Shuwei Li, Lianming Wang, and Xinyuan Song. Variable selection in semiparametric nonmixture cure model with intervalcensored failure time data: an application to the prostate cancer screening study. Statistics in Medicine, 38(16):3026–3039, July 20, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2019:BAI

[3468] Kan Li, Sheng Luo, Sammy Yuan, and Shahrul Mt-Isa. A Bayesian approach for individual-level drug benefit-risk assessment. Statistics in Medicine, 38(16):3040–3052, July 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tallarita:2019:CRN

[3469] Marta Tallarita, Maria De Iorio, and Gianluca Baio. A comparative review of network meta-analysis models in longitudinal randomized controlled trial. Statistics in Medicine, 38(16):3053–3072, July 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sofrygin:2019:TLD

[3470] Oleg Sofrygin, Zheng Zhu, Julie A. Schmittdiel, Alyce S. Adams, Richard W. Grant, Mark J. van der Laan, and Romain Neugebauer. Targeted learning with daily EHR data. Statistics in Medicine, 38(16): 3073–3090, July 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mauro:2019:IMD

[3471] Christine Mauro, M. Katherine Shear, and Yuanjia Wang. Integrating multiple-domain rules for disease classification. Statistics in Medicine, 38(16):3091–3104, July 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Placzek:2019:CEF

[3472] Marius Placzek and Tim Friede. A conditional error function approach for adaptive enrichment designs with continuous endpoints. Statistics in Medicine, 38(17):3105–3122, July 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gajewski:2019:BHE

[3473] Byron J. Gajewski, Caitlyn Meinzer, Scott M. Berry, Gaylan L. Rockswold, William G. Barsan, Frederick K. Korley, and Renee’ H. Martin. Bayesian hierarchical EMAX model for dose-response in early phase efficacy clinical trials. Statistics in Medicine, 38(17):3123–3138, July 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scotina:2019:MAC

[3474] Anthony D. Scotina and Roee Gutman. Matching algorithms for causal inference with multiple treatments. Statistics in Medicine, 38(17):3139– 3167, July 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sall:2019:TCS

[3475] Alioune Sall, Karine Aubé, Xavier Trudel, Chantal Brisson, and Denis Talbot. A test for the correct specification of marginal structural models. Statistics in Medicine, 38(17):3168–3183, July 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Akushevich:2019:PTT

[3476] Igor Akushevich, Julia Kravchenko, Arseniy P. Yashkin, Fang Fang, and Anatoliy I. Yashin. Partitioning of time trends in prevalence and mortality of lung cancer. Statistics in Medicine, 38(17):3184–3203, July 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Desai:2019:MDS

[3477] Manisha Desai, Maria E. Montez-Rath, Kristopher Kapphahn, Vilija R. Joyce, Maya B. Mathur, Ariadna Garcia, Natasha Purington, and Douglas K. Owens. Missing data strategies for time-varying confounders in comparative effectiveness studies of non-missing time-varying exposures and right-censored outcomes. Statistics in Medicine, 38(17):3204–3220, July 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2019:PIS

[3478] Yang Li, Rong Li, Cunjie Lin, Yichen Qin, and Shuangge Ma. Penalized integrative semiparametric interaction analysis for multiple genetic datasets. Statistics in Medicine, 38(17):3221–3242, July 30, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Amro:2019:MCT

[3479] Lubna Amro, Frank Konietschke, and Markus Pauly. Multiplicationcombination tests for incomplete paired data. Statistics in Medicine, 38(17):3243–3255, July 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2019:SIH

[3480] Jialiang Li, Mu Yue, and Wenyang Zhang. Subgroup identification via homogeneity pursuit for dense longitudinal/spatial data. Statistics in Medicine, 38(17):3256–3271, July 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelson:2019:PKC

[3481] Kerrie P. Nelson and Don Edwards. A paired kappa to compare binary ratings across two medical tests. Statistics in Medicine, 38(17):3272– 3287, July 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seide:2019:SDG

[3482] Svenja E. Seide, Katrin Jensen, and Meinhard Kieser. Simulation and data-generation for random-effects network meta-analysis of binary outcome. Statistics in Medicine, 38(17):3288–3303, July 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lorenz:2019:CLR

[3483] Doug Lorenz. Correction for “A log rank test for clustered data with informative within-cluster group size” by Mary E. Gregg, Somnath Datta, and Doug Lorenz. Statistics in Medicine, 38(17):3304, July 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2019:AAN

[3484] Kim May Lee, James Wason, and Nigel Stallard. To add or not to add a new treatment arm to a multiarm study: a decision-theoretic framework. Statistics in Medicine, 38(18):3305–3321, August 15, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bujkiewicz:2019:BNM

[3485] Sylwia Bujkiewicz, Dan Jackson, John R. Thompson, Rebecca M. Turner, Nicolas Städler, Keith R. Abrams, and Ian R. White. Bivariate network meta-analysis for surrogate endpoint evaluation. Statistics in Medicine, 38(18):3322–3341, August 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kelly:2019:NSS

[3486] Thu-Lan Kelly and Nicole Pratt. A note on sample size calculations for cluster randomised crossover trials with a fixed number of clusters. Statistics in Medicine, 38(18):3342–3345, August 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Djordjilovic:2019:GTH

[3487] Vera Djordjilović, Christian M. Page, Jon Michael Gran, Therese H. Nøst, Torkjel M. Sandanger, Marit B. Veierød, and Magne Thoresen. Global test for high-dimensional mediation: Testing groups of potential mediators. Statistics in Medicine, 38(18):3346–3360, August 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2019:BNE

[3488] Rui Zhu and Subhashis Ghosal. Bayesian nonparametric estimation of ROC surface under verification bias. Statistics in Medicine, 38(18):3361– 3377, August 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2019:MBS

[3489] Zhongkai Wang, Babette A. Brumback, Adel A. Alrwisan, and Almut G. Winterstein. Model-based standardization using an outcome model with random effects. Statistics in Medicine, 38(18):3378–3394, August 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2019:SSC

[3490] Xiaohan Xu, Hong Zhu, and Chul Ahn. Sample size considerations for stratified cluster randomization design with binary outcomes and varying cluster size. Statistics in Medicine, 38(18):3395–3404, August 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Balan:2019:NHU

[3491] Theodor Adrian Balan and Hein Putter. Nonproportional hazards and unobserved heterogeneity in clustered survival data: When can we tell the difference? Statistics in Medicine, 38(18):3405–3420, August 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stander:2019:APV

[3492] Julian Stander, Luciana Dalla Valle, Charlotte Taglioni, Brunero Liseo, Angie Wade, and Mario Cortina-Borja. Analysis of paediatric visual acuity using Bayesian copula models with sinh-arcsinh marginal densities. Statistics in Medicine, 38(18):3421–3443, August 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luijken:2019:IPM

[3493] K. Luijken, R. H. H. Groenwold, B. Van Calster, E. W. Steyerberg, and M. van Smeden. Impact of predictor measurement heterogeneity across settings on the performance of prediction models: a measurement error perspective. Statistics in Medicine, 38(18):3444–3459, August 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:2019:ECI

[3494] Hayeon Park and Lawrence M. Leemis. Ensemble confidence intervals for binomial proportions. Statistics in Medicine, 38(18):3460–3475, August 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Charles-Nelson:2019:HAI

[3495] Anaïs Charles-Nelson, Sandrine Katsahian, and Catherine Schramm. How to analyze and interpret recurrent events data in the presence of a terminal event: an application on readmission after colorectal cancer surgery. Statistics in Medicine, 38(18):3476–3502, August 15, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vanbelle:2019:RMA

[3496] Sophie Vanbelle. Review of measuring agreement: Models, methods, and applications. Statistics in Medicine, 38(18):3503–3504, August 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shuster:2019:WHI

[3497] Jonathan J. Shuster and Mark Handler. Withdrawn: How to investigate an accused serial sexual harasser. Statistics in Medicine, 38(18):O3505, August 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ryan:2019:FPC

[3498] Louise Ryan. Four papers on child growth modelling. Statistics in Medicine, 38(19):3505–3506, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ohuma:2019:SMC

[3499] Eric O. Ohuma, Douglas G. Altman, and International Fetal and Newborn Growth Consortium for the 21st Century (INTERGROWTH-21st Project). Statistical methodology for constructing gestational age-related charts using cross-sectional and longitudinal data: The INTERGROWTH-21st Project as a case study. Statistics in Medicine, 38(19):3507–3526, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ohuma:2019:DOM

[3500] Eric O. Ohuma, Douglas G. Altman, and International Fetal and Newborn Growth Consortium for the 21st Century (INTERGROWTH-21st Project). Design and other methodological considerations for the construction of human fetal and neonatal size and growth charts. Statistics in Medicine, 38(19):3527–3539, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anderson:2019:UDM

[3501] Craig Anderson, Luo Xiao, and William Checkley. Using data from multiple studies to develop a child growth correlation matrix. Statistics in Medicine, 38(19):3540–3554, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anderson:2019:CPA

[3502] Craig Anderson, Ryan Hafen, Oleg Sofrygin, Louise Ryan, and members of the Hbgdki Community. Comparing predictive abilities of longitudinal child growth models. Statistics in Medicine, 38(19):3555–3570, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cole:2019:CMC

[3503] T. J. Cole. Commentary: Methods for calculating growth trajectories and constructing growth centiles. Statistics in Medicine, 38(19):3571– 3579, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ohuma:2019:RPC

[3504] Eric O. Ohuma. Response to Professor Cole’s commentary: Methods for calculating growth trajectories and constructing growth centiles. Statistics in Medicine, 38(19):3580–3583, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anderson:2019:RPT

[3505] Craig Anderson. Response to Professor Tim Cole’s commentary: Methods for calculating growth trajectories and constructing growth centiles. Statistics in Medicine, 38(19):3584–3585, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawrence:2019:FPF

[3506] John Lawrence. Familywise and per-family error rates of multiple comparison procedures. Statistics in Medicine, 38(19):3586–3598, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rennert:2019:BII

[3507] Lior Rennert and Sharon X. Xie. Bias induced by ignoring double truncation inherent in autopsy-confirmed survival studies of neurodegenerative diseases. Statistics in Medicine, 38(19):3599–3613, August 30, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morrison:2019:CSH

[3508] Doug Morrison, Oliver Laeyendecker, and Ron Brookmeyer. Crosssectional HIV incidence estimation in an evolving epidemic. Statistics in Medicine, 38(19):3614–3627, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mao:2019:PHR

[3509] Lu Mao. Proportional hazards regression of survival-sacrifice data with cause-of-death information in animal carcinogenicity studies. Statistics in Medicine, 38(19):3628–3641, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2019:EVJ

[3510] Yicheng Kang, Xiaodong Gong, Jiti Gao, and Peihua Qiu. Errors-in-variables jump regression using local clustering. Statistics in Medicine, 38(19):3642–3655, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2019:FRM

[3511] Xiangyu Liu, Jing Ning, Yu Cheng, Xuelin Huang, and Ruosha Li. A flexible and robust method for assessing conditional association and conditional concordance. Statistics in Medicine, 38(19):3656–3668, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hogg:2019:ADM

[3512] Tanja Högg, Yinshan Zhao, Paul Gustafson, John Petkau, John Fisk, Ruth Ann Marrie, and Helen Tremlett. Adjusting for differential misclassification in matched case-control studies utilizing health administrative data. Statistics in Medicine, 38(19):3669–3681, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baayen:2019:STC

[3513] C. Baayen, C. Volteau, C. Flamant, and P. Blanche. Sequential trials in the context of competing risks: Concepts and case study, with R and SAS code. Statistics in Medicine, 38(19):3682–3702, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2019:BAL

[3514] Chunjie Wang, Qun Li, Xinyuan Song, and Xiaogang Dong. Bayesian adaptive lasso for additive hazard regression with current status data. Statistics in Medicine, 38(20):3703–3718, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McCurdy:2019:FAS

[3515] Shannon McCurdy, Annette Molinaro, and Lior Pachter. Factor analysis for survival time prediction with informative censoring and diverse covariates. Statistics in Medicine, 38(20):3719–3732, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2019:ODT

[3516] Jingxia Liu, Lei Liu, and Graham A. Colditz. Optimal designs in threelevel cluster randomized trials with a binary outcome. Statistics in Medicine, 38(20):3733–3746, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bluhmki:2019:BCT

[3517] Tobias Bluhmki, Hein Putter, Arthur Allignol, Jan Beyersmann, and on behalf of the Combacte-Magnet consortium. Bootstrapping complex time-to-event data without individual patient data, with a view toward time-dependent exposures. Statistics in Medicine, 38(20):3747– 3763, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tekwe:2019:IVA

[3518] Carmen D. Tekwe, Roger S. Zoh, Miao Yang, Raymond J. Carroll, Gilson Honvoh, David B. Allison, Mark Benden, and Lan Xue. Instrumental variable approach to estimating the scalar-on-function regression model with measurement error with application to energy expenditure assessment in childhood obesity. Statistics in Medicine, 38(20):3764–3781, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Magirr:2019:MWL

[3519] Dominic Magirr and Carl-Fredrik Burman. Modestly weighted logrank tests. Statistics in Medicine, 38(20):3782–3790, September 10, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Segalas:2019:HTP

[3520] Corentin Segalas, Hélène Amieva, and Hélène Jacqmin-Gadda. A hypothesis testing procedure for random changepoint mixed models. Statistics in Medicine, 38(20):3791–3803, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sidik:2019:NEB

[3521] Kurex Sidik and Jeffrey N. Jonkman. A note on the empirical Bayes heterogeneity variance estimator in meta-analysis. Statistics in Medicine, 38(20):3804–3816, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Enserro:2019:MEP

[3522] Danielle M. Enserro, Olga V. Demler, Michael J. Pencina, and Ralph B. D’Agostino, Sr. Measures for evaluation of prognostic improvement under multivariate normality for nested and nonnested models. Statistics in Medicine, 38(20):3817–3831, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Conner:2019:ARM

[3523] Sarah C. Conner, Lisa M. Sullivan, Emelia J. Benjamin, Michael P. LaValley, Sandro Galea, and Ludovic Trinquart. Adjusted restricted mean survival times in observational studies. Statistics in Medicine, 38 (20):3832–3860, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [3652].

Spineli:2019:POG

[3524] Loukia M. Spineli, Chrysostomos Kalyvas, and Konstantinos Pateras. Participants’ outcomes gone missing within a network of interventions: Bayesian modeling strategies. Statistics in Medicine, 38(20):3861–3879, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vonCube:2019:PAF

[3525] Maja von Cube, Martin Schumacher, Sébastien Bailly, Jean-François Timsit, Alain Lepape, Anne Savey, Anais Machut, and Martin Wolkewitz. The population-attributable fraction for time-dependent exposures and competing risks — a discussion on estimands. Statistics in Medicine, 38(20):3880–3895, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kipourou:2019:EAC

[3526] Dimitra-Kleio Kipourou, Hadrien Charvat, Bernard Rachet, and Aurélien Belot. Estimation of the adjusted cause-specific cumulative probability using flexible regression models for the cause-specific hazards. Statistics in Medicine, 38(20):3896–3910, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cochran:2019:LVA

[3527] Amy L. Cochran, Paul J. Rathouz, Keith E. Kocher, and Gabriel Zayas-Cabán. A latent variable approach to potential outcomes for emergency department admission decisions. Statistics in Medicine, 38(20):3911– 3935, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2019:MDW

[3528] Zonglin He and Youyi Fong. Maximum diversity weighting for biomarkers with application in HIV-1 vaccine studies. Statistics in Medicine, 38(20): 3936–3946, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hirschfeld:2019:CBS

[3529] G. Hirschfeld and C. Thiele. Cloud-based simulation studies in R — a tutorial on using doRedis with Amazon spot fleets. Statistics in Medicine, 38(20):3947–3959, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baiocchi:2019:C

[3530] Michael Baiocchi, Jing Cheng, and Dylan S. Small. Correction. Statistics in Medicine, 38(20):3960, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Su:2019:ACF

[3531] Chien-Lin Su and Feng-Chang Lin. Analysis of clustered failure time data with cure fraction using copula. Statistics in Medicine, 38(21): 3961–3973, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steingrimsson:2019:SIU

[3532] Jon Arni Steingrimsson and Jiabei Yang. Subgroup identification using covariate-adjusted interaction trees. Statistics in Medicine, 38(21):3974– 3984, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sinha:2019:AGS

[3533] Arup K. Sinha, Lemuel Moye III, Linda B. Piller, Jose-Miguel Yamal, Carlos H. Barcenas, Jianchang Lin, and Barry R. Davis. Adaptive groupsequential design with population enrichment in phase 3 randomized controlled trials with two binary co-primary endpoints. Statistics in Medicine, 38(21):3985–3996, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Camilli:2019:SAE

[3534] Gregory Camilli and Eugene Geis. Stochastic approximation EM for large-scale exploratory IRT factor analysis. Statistics in Medicine, 38 (21):3997–4012, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2019:PMA

[3535] Chang Yu and Daniel Zelterman. A parametric meta-analysis. Statistics in Medicine, 38(21):4013–4025, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2019:FOC

[3536] Yifan Zhang, Lorenzo Trippa, and Giovanni Parmigiani. Frequentist operating characteristics of Bayesian optimal designs via simulation. Statistics in Medicine, 38(21):4026–4039, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schwall:2019:EPP

[3537] Philipp Schwall, Christian Meesters, and Jochen Hardt. Estimating person parameters via item response model and simple sum score in small samples with few polytomous items: a simulation study. Statistics in Medicine, 38(21):4040–4050, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2019:ICI

[3538] Peter C. Austin and Ewout W. Steyerberg. The integrated calibration index (ICI) and related metrics for quantifying the calibration of logistic regression models. Statistics in Medicine, 38(21):4051–4065, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2019:DMM

[3539] Jooyoung Lee and Richard J. Cook. Dependence modeling for multi-type recurrent events via copulas. Statistics in Medicine, 38(21):4066–4082, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2019:SIR

[3540] Mandi Yu, Yan Li, and Meng Qiu. Statistical inference of the relative concentration index for complex surveys. Statistics in Medicine, 38(21): 4083–4095, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2019:AIS

[3541] Yi Liu, Feng Guo, and Richard J. Hanowski. Assessing the impact of sleep time on truck driver performance using a recurrent event model. Statistics in Medicine, 38(21):4096–4111, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2019:CNM

[3542] Yuyin Liu, Denis Rybin, Timothy C. Heeren, and Gheorghe Doros. Comparison of novel methods in two-way enriched clinical trial design. Statistics in Medicine, 38(21):4112–4130, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yoneoka:2019:CHR

[3543] Daisuke Yoneoka and Masayuki Henmi. Clinical heterogeneity in random-effect meta-analysis: Between-study boundary estimate problem. Statistics in Medicine, 38(21):4131–4145, September 20, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eilertson:2019:EPM

[3544] Kirsten E. Eilertson, John Fricks, and Matthew J. Ferrari. Estimation and prediction for a mechanistic model of measles transmission using particle filtering and maximum likelihood estimation. Statistics in Medicine, 38(21):4146–4158, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pilz:2019:VAO

[3545] Maximilian Pilz, Kevin Kunzmann, Carolin Herrmann, Geraldine Rauch, and Meinhard Kieser. A variational approach to optimal two-stage designs. Statistics in Medicine, 38(21):4159–4171, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alam:2019:CCD

[3546] M. Iftakhar Alam, D. Stephen Coad, and Barbara Bogacka. Combined criteria for dose optimisation in early phase clinical trials. Statistics in Medicine, 38(21):4172–4188, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Infanger:2019:PVF

[3547] Denis Infanger and Arno Schmidt-Trucksäss. P value functions: an underused method to present research results and to promote quantitative reasoning. Statistics in Medicine, 38(21):4189–4197, September 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schuemie:2019:PSU

[3548] Martijn J. Schuemie, Patrick B. Ryan, Kenneth K. C. Man, Ian C. K. Wong, Marc A. Suchard, and George Hripcsak. A plea to stop using the case-control design in retrospective database studies. Statistics in Medicine, 38(22):4199–4208, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schneeweiss:2019:DSA

[3549] Sebastian Schneeweiss and Samy Suissa. Discussion of Schuemie et al: “A plea to stop using the case-control design in retrospective database studies”. Statistics in Medicine, 38(22):4209–4212, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rose:2019:COD

[3550] Sherri Rose. Considerations for outcome-dependent biased sampling in health databases. Statistics in Medicine, 38(22):4213–4215, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gault:2019:CPS

[3551] Nathalie Gault. A comment on “A plea to stop using the case-control design in retrospective database studies”. Statistics in Medicine, 38(22): 4216–4217, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vonesh:2019:MEM

[3552] Edward Vonesh, Hocine Tighiouart, Jian Ying, Hiddo L. Heerspink, Julia Lewis, Natalie Staplin, Lesley Inker, and Tom Greene. Mixed-effects models for slope-based endpoints in clinical trials of chronic kidney disease. Statistics in Medicine, 38(22):4218–4239, September 30, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2019:STC

[3553] Jung In Kim, Feng-Chang Lin, and Jason P. Fine. On the self-triggering Cox model for recurrent event data. Statistics in Medicine, 38(22):4240– 4252, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yarnell:2019:MVB

[3554] Christopher Yarnell, Ruxandra Pinto, and Rob Fowler. Measuring variability between clusters by subgroup: an extension of the median odds ratio. Statistics in Medicine, 38(22):4253–4263, September 30, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prescott:2019:TTS

[3555] Robin J. Prescott. Two-tailed significance tests for 2 × 2 contingency tables: What is the alternative? Statistics in Medicine, 38(22):4264– 4269, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [3648] and response [3650].

Meller:2019:JMP

[3556] Matthias Meller, Jan Beyersmann, and Kaspar Rufibach. Joint modeling of progression-free and overall survival and computation of correlation measures. Statistics in Medicine, 38(22):4270–4289, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steyerberg:2019:AHI

[3557] Ewout W. Steyerberg, Daan Nieboer, Thomas P. A. Debray, and Hans C. van Houwelingen. Assessment of heterogeneity in an individual participant data meta-analysis of prediction models: an overview and illustration. Statistics in Medicine, 38(22):4290–4309, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mandal:2019:STS

[3558] Saumen Mandal, Reza Arabi Belaghi, Akram Mahmoudi, and Minoo Aminnejad. Stein-type shrinkage estimators in gamma regression model with application to prostate cancer data. Statistics in Medicine, 38(22): 4310–4322, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gustafson:2019:RRT

[3559] Reka Gustafson, Paul Gustafson, and Patricia Daly. Reconciling randomized trial evidence on proximal versus distal outcomes, with application to trials of influenza vaccination for healthcare workers. Statistics in Medicine, 38(22):4323–4333, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:2019:CTC

[3560] Jeffrey M. Albert, Youjun Li, Jiayang Sun, Wojbor A. Woyczynski, and Suchitra Nelson. Continuous-time causal mediation analysis. Statistics in Medicine, 38(22):4334–4347, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jazic:2019:DAN

[3561] Ina Jazić, Sebastien Haneuse, Benjamin French, Gaëtan MacGrogan, and Virginie Rondeau. Design and analysis of nested case-control studies for recurrent events subject to a terminal event. Statistics in Medicine, 38 (22):4348–4362, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bastos:2019:MAC

[3562] Leonardo S. Bastos, Theodoros Economou, Marcelo F. C. Gomes, Daniel A. M. Villela, Flavio C. Coelho, Oswaldo G. Cruz, Oliver Stoner, Trevor Bailey, and Claudia T. Codeço. A modelling approach for correcting reporting delays in disease surveillance data. Statistics in Medicine, 38 (22):4363–4377, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:2019:CFD

[3563] Xianming Tan, Guanghan F. Liu, Donglin Zeng, William Wang, Guoqing Diao, Joseph F. Heyse, and Joseph G. Ibrahim. Controlling false discovery proportion in identification of drug-related adverse events from multiple system organ classes. Statistics in Medicine, 38(22):4378–4389, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ohneberg:2019:EDS

[3564] Kristin Ohneberg, Jan Beyersmann, and Martin Schumacher. Exposure density sampling: Dynamic matching with respect to a time-dependent exposure. Statistics in Medicine, 38(22):4390–4403, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fernandez-Fontelo:2019:USD

[3565] Amanda Fernández-Fontelo, Alejandra Cabaña, Harry Joe, Pedro Puig, and David Moriña. Untangling serially dependent underreported count data for gender-based violence. Statistics in Medicine, 38(22):4404–4422, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kurz:2019:IIS

[3566] Christoph F. Kurz and Laura A. Hatfield. Identifying and interpreting subgroups in health care utilization data with count mixture regression models. Statistics in Medicine, 38(22):4423–4435, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sysoev:2019:PDT

[3567] Oleg Sysoev, Krzysztof Bartoszek, Eva-Charlotte Ekström, and Katarina Ekholm Selling. PSICA: Decision trees for probabilistic subgroup identification with categorical treatments. Statistics in Medicine, 38(22): 4436–4452, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2019:MLE

[3568] Stuart G. Baker. Maximum likelihood estimation with missing outcomes: From simplicity to complexity. Statistics in Medicine, 38(22):4453–4474, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ono:2019:FPS

[3569] Yusuke Ono. Formula for Var(β̂_{1}) on page 5536 in Senn, Graf, and Caputo (2007). Statistics in Medicine, 38(22):4475, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gasparini:2019:IMM

[3570] Alessandro Gasparini, Mark S. Clements, Keith R. Abrams, and Michael J. Crowther. Impact of model misspecification in shared frailty survival models. Statistics in Medicine, 38(23):4477–4502, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilbert:2019:APM

[3571] Peter B. Gilbert, Yuanyuan Zhang, Erika Rudnicki, and Yunda Huang. Assessing pharmacokinetic marker correlates of outcome, with application to antibody prevention efficacy trials. Statistics in Medicine, 38 (23):4503–4518, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2019:RAV

[3572] Juexin Lin, Dewei Wang, and Qi Zheng. Regression analysis and variable selection for two-stage multiple-infection group testing data. Statistics in Medicine, 38(23):4519–4533, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chase:2019:AEP

[3573] Elizabeth C. Chase and Philip S. Boonstra. Accounting for established predictors with the multistep elastic net. Statistics in Medicine, 38(23): 4534–4544, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rivera:2019:MED

[3574] Roberto Rivera and Wolfgang Rolke. Modeling excess deaths after a natural disaster with application to Hurricane Maria. Statistics in Medicine, 38(23):4545–4554, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2019:NES

[3575] Kai Yang and Peihua Qiu. Nonparametric estimation of the spatiotemporal covariance structure. Statistics in Medicine, 38(23):4555–4565, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Joseph:2019:BCB

[3576] Lawrence Joseph and Patrick Bélisle. Bayesian consensus-based sample size criteria for binomial proportions. Statistics in Medicine, 38 (23):4566–4573, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2019:DFS

[3577] Hailin Huang, Yuanzhang Li, Hua Liang, and Colin O. Wu. Decomposition feature selection with applications in detecting correlated biomarkers of bipolar disorders. Statistics in Medicine, 38(23):4574–4582, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawless:2019:NPL

[3578] Jerald F. Lawless and Richard J. Cook. A new perspective on loss to follow-up in failure time and life history studies. Statistics in Medicine, 38(23):4583–4610, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rivera-Rodriguez:2019:ATP

[3579] C. Rivera-Rodriguez, D. Spiegelman, and S. Haneuse. On the analysis of two-phase designs in cluster-correlated data settings. Statistics in Medicine, 38(23):4611–4624, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2019:RSG

[3580] Mengyun Wu and Shuangge Ma. Robust semiparametric geneenvironment interaction analysis using sparse boosting. Statistics in Medicine, 38(23):4625–4641, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mandal:2019:ALT

[3581] Soutrik Mandal, Suojin Wang, and Samiran Sinha. Analysis of linear transformation models with covariate measurement error and interval censoring. Statistics in Medicine, 38(23):4642–4655, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xi:2019:ABG

[3582] Dong Xi and Paul Gallo. An additive boundary for group sequential designs with connection to conditional error. Statistics in Medicine, 38 (23):4656–4669, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2019:VSC

[3583] Erqian Li, Maozai Tian, and Man-Lai Tang. Variable selection in competing risks models based on quantile regression. Statistics in Medicine, 38(23):4670–4685, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kasza:2019:ICS

[3584] Jessica Kasza, Monica Taljaard, and Andrew B. Forbes. Information content of stepped-wedge designs when treatment effect heterogeneity and/or implementation periods are present. Statistics in Medicine, 38 (23):4686–4701, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proust-Lima:2019:JMM

[3585] Cécile Proust-Lima, Viviane Philipps, and Jean-François Dartigues. A joint model for multiple dynamic processes and clinical endpoints: Application to Alzheimer’s disease. Statistics in Medicine, 38(23):4702–4717, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wolter:2019:ETR

[3586] Kirk M. Wolter, N. Ganesh, Kennon R. Copeland, James A. Singleton, and Meena Khare. Estimation tools for reducing the impact of sampling and nonresponse errors in dual-frame RDD telephone surveys. Statistics in Medicine, 38(23):4718–4732, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gleiss:2019:QDN

[3587] Andreas Gleiss and Michael Schemper. Quantifying degrees of necessity and of sufficiency in cause-effect relationships with dichotomous and survival outcomes. Statistics in Medicine, 38(23):4733–4748, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zetterqvist:2019:DRC

[3588] Johan Zetterqvist, Karel Vermeulen, Stijn Vansteelandt, and Arvid Sjölander. Doubly robust conditional logistic regression. Statistics in Medicine, 38(23):4749–4760, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Magnusson:2019:BIP

[3589] Baldur P. Magnusson, Heinz Schmidli, Nicolas Rouyrre, and Daniel O. Scharfstein. Bayesian inference for a principal stratum estimand to assess the treatment effect in a subgroup characterized by postrandomization event occurrence. Statistics in Medicine, 38(23):4761–4771, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:2019:RAT

[3590] Changyong Feng, Bokai Wang, and Hongyue Wang. The relations among three popular indices of risks. Statistics in Medicine, 38(23):4772–4787, October 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:2019:QHD

[3591] Hayley E. Jones, Constantine A. Gatsonsis, Thomas A. Trikalinos, Nicky J. Welton, and A. E. Ades. Quantifying how diagnostic test accuracy depends on threshold in a meta-analysis. Statistics in Medicine, 38 (24):4789–4803, October 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [3655].

Li:2019:DPA

[3592] Kan Li and Sheng Luo. Dynamic prediction of Alzheimer’s disease progression using features of multiple longitudinal outcomes and time-to-event data. Statistics in Medicine, 38(24):4804–4818, October 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2019:SSC

[3593] Yongqiang Tang and Ronan Fitzpatrick. Sample size calculation for the Andersen–Gill model comparing rates of recurrent events. Statistics in Medicine, 38(24):4819–4827, October 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vansteelandt:2019:MAT

[3594] Stijn Vansteelandt, Martin Linder, Sjouke Vandenberghe, Johan Steen, and Jesper Madsen. Mediation analysis of time-to-event endpoints accounting for repeatedly measured mediators subject to time-varying confounding. Statistics in Medicine, 38(24):4828–4840, October 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tchetgen:2019:GAD

[3595] Eric J. Tchetgen Tchetgen, Xu Shi, Benedict H. W. Wong, and Tamar Sofer. A general approach to detect gene (G)–environment (E) additive interaction leveraging G–E independence in case-control studies. Statistics in Medicine, 38(24):4841–4853, October 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2019:FPM

[3596] Rui Huang, Liming Xiang, and Il Do Ha. Frailty proportional mean residual life regression for clustered survival data: a hierarchical quasilikelihood method. Statistics in Medicine, 38(24):4854–4870, October 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnson:2019:SDA

[3597] Olatunji Johnson, Peter Diggle, and Emanuele Giorgi. A spatially discrete approximation to log-Gaussian Cox processes for modelling aggregated disease count data. Statistics in Medicine, 38(24):4871–4887, October 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schomaker:2019:ULT

[3598] M. Schomaker, M. A. Luque-Fernandez, V. Leroy, and M. A. Davies. Using longitudinal targeted maximum likelihood estimation in complex settings with dynamic interventions. Statistics in Medicine, 38(24):4888– 4911, October 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hitt:2019:OFC

[3599] Brianna D. Hitt, Christopher R. Bilder, Joshua M. Tebbs, and Christopher S. McMahan. The objective function controversy for group testing: Much ado about nothing? Statistics in Medicine, 38(24):4912–4923, October 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sies:2019:EQO

[3600] Aniek Sies and Iven Van Mechelen. Estimating the quality of optimal treatment regimes. Statistics in Medicine, 38(25):4925–4938, November 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roy:2019:NBF

[3601] Akash Roy, Solomon W. Harrar, and Frank Konietschke. The nonparametric Behrens–Fisher problem with dependent replicates. Statistics in Medicine, 38(25):4939–4962, November 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Landsman:2019:OMC

[3602] V. Landsman, D. Landsman, C. S. Li, and H. Bang. Overdispersion models for correlated multinomial data: Applications to blinding assessment. Statistics in Medicine, 38(25):4963–4976, November 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2019:EAM

[3603] Jooyoung Lee and Richard J. Cook. On estimands arising from misspecified semiparametric rate-based analysis of recurrent episodic conditions. Statistics in Medicine, 38(25):4977–4998, November 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2019:TST

[3604] Yi Cai, Jing Huang, Jing Ning, Mei-Ling Ting Lee, Bernard Rosner, and Yong Chen. Two-sample test for correlated data under outcomedependent sampling with an application to self-reported weight loss data.

Statistics in Medicine, 38(25):4999–5009, November 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spooner:2019:LRE

[3605] Max Spooner and Anders Stockmarr. Likelihood ratios for evaluating DNA matches obtained from a database search when there is substructure in the population. Statistics in Medicine, 38(25):5010–5020, November 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grantham:2019:HMT

[3606] Kelsey L. Grantham, Jessica Kasza, Stephane Heritier, Karla Hemming, Edward Litton, and Andrew B. Forbes. How many times should a cluster randomized crossover trial cross over? Statistics in Medicine, 38(25): 5021–5033, November 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rao:2019:PBT

[3607] Kaidi Rao, Reza Drikvandi, and Benjamin Saville. Permutation and Bayesian tests for testing random effects in linear mixed-effects models. Statistics in Medicine, 38(25):5034–5047, November 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:2019:BAR

[3608] Yaoyuan Vincent Tan and Jason Roy. Bayesian additive regression trees and the general BART model. Statistics in Medicine, 38(25):5048–5069, November 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moerbeek:2019:WSI

[3609] Mirjam Moerbeek and Sander van Schie. What are the statistical implications of treatment non-compliance in cluster randomized trials: a simulation study. Statistics in Medicine, 38(26):5071–5084, November 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mittinty:2019:EDT

[3610] Murthy N. Mittinty, John W. Lynch, Andrew B. Forbes, and Lyle C. Gurrin. Effect decomposition through multiple causally nonordered mediators in the presence of exposure-induced mediator-outcome confounding. Statistics in Medicine, 38(26):5085–5102, November 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zeng:2019:IMS

[3611] X. Y. Zeng, P. H. Chau, and Paul S. F. Yip. Improving the monitoring of suicide incidence by estimating the probability of news reporting. Statistics in Medicine, 38(26):5103–5112, November 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hanley:2019:MIM

[3612] James A. Hanley. A more intuitive and modern way to compute a small-sample confidence interval for the mean of a Poisson distribution. Statistics in Medicine, 38(26):5113–5119, November 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Granger:2019:APW

[3613] Emily Granger, Jamie C. Sergeant, and Mark Lunt. Avoiding pitfalls when combining multiple imputation and propensity scores. Statistics in Medicine, 38(26):5120–5132, November 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2019:CEI

[3614] Xin Wang, Yingchao Zhong, Purna Mukhopadhyay, and Douglas E. Schaubel. Computationally efficient inference for center effects based on restricted mean survival time. Statistics in Medicine, 38(26):5133– 5145, November 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sugasawa:2019:EIT

[3615] Shonosuke Sugasawa and Hisashi Noma. Estimating individual treatment effects by gradient boosting trees. Statistics in Medicine, 38(26):5146– 5159, November 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Espasandin-Dominguez:2019:ARB

[3616] J. Espasandín-Domínguez, C. Cadarso-Suárez, T. Kneib, G. Marra, N. Klein, R. Radice, O. Lado-Baleato, A. González-Quintela, and F. Gude. Assessing the relationship between markers of glycemic control through flexible copula regression models. Statistics in Medicine, 38(27): 5161–5181, November 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nab:2019:MEC

[3617] L. Nab, R. H. H. Groenwold, P. M. J. Welsing, and M. van Smeden. Measurement error in continuous endpoints in randomised trials: Problems and solutions. Statistics in Medicine, 38(27):5182–5196, November 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2019:CAB

[3618] Ian R. White, Rebecca M. Turner, Amalia Karahalios, and Georgia Salanti. A comparison of arm-based and contrast-based models for network meta-analysis. Statistics in Medicine, 38(27):5197–5213, November 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lou:2019:ARM

[3619] Yiyue Lou, Michael P. Jones, and Wanjie Sun. Assessing the ratio of means as a causal estimand in clinical endpoint bioequivalence studies in the presence of intercurrent events. Statistics in Medicine, 38(27):5214– 5235, November 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cengiz:2019:NSE

[3620] Ünal Cengiz and Mehmet Karahasan. A new statistical early outbreak detection method for biosurveillance and performance comparisons. Statistics in Medicine, 38(27):5236–5258, November 30, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Watson:2019:ALA

[3621] David Watson and Amanda Nickel. Analysis of longitudinal adherence data: Hitting the target with consistency. Statistics in Medicine, 38(27): 5259–5267, November 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xi:2019:SGE

[3622] Dong Xi and Frank Bretz. Symmetric graphs for equally weighted tests, with application to the Hochberg procedure. Statistics in Medicine, 38 (27):5268–5282, November 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2019:REE

[3623] Abhik Ghosh and Ayanendranath Basu. Robust and efficient estimation in the parametric proportional hazards model under random censoring. Statistics in Medicine, 38(27):5283–5299, November 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2019:UBB

[3624] Yanhong Zhou, J. Jack Lee, and Ying Yuan. A utility-based Bayesian optimal interval (U-BOIN) phase I/II design to identify the optimal biological dose for targeted and immune therapies. Statistics in Medicine, 38 (28):S5299–S5316, December 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2019:QTE

[3625] Ying-Qi Zhao, Mary W. Redman, and Michael L. LeBlanc. Quantifying treatment effects using the personalized chance of longer survival. Statistics in Medicine, 38(28):5317–5331, December 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2019:NTM

[3626] Yuanyuan Han, Zhao-Hua Lu, and Wai-Yin Poon. Noninferiority testing for matched-pair ordinal data with misclassification. Statistics in Medicine, 38(28):5332–5349, December 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmidt:2019:SCI

[3627] Sylvia Schmidt and Werner Brannath. Simultaneous confidence intervals for ratios with application to the gold standard design with more than one experimental treatment. Statistics in Medicine, 38(28):5350–5360, December 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lancker:2019:EFB

[3628] Kelly Van Lancker, An Vandebosch, Stijn Vansteelandt, and Filip De Ridder. Evaluating futility of a binary clinical endpoint using early readouts. Statistics in Medicine, 38(28):5361–5375, December 10, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sewell:2019:ANI

[3629] Daniel K. Sewell and for the Cdc Mind-Healthcare Program. Analysis of network interventions with an application to hospital-acquired infections. Statistics in Medicine, 38(28):5376–5390, December 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siriwardhana:2019:PTS

[3630] Chathura Siriwardhana, K. B. Kulasekera, and Somnath Datta. Personalized treatment selection using data from crossover designs with carry-over effects. Statistics in Medicine, 38(28):5391–5412, December 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2019:SSC

[3631] Dateng Li, Song Zhang, and Jing Cao. Sample size calculation for clinical trials with correlated count measurements based on the negative binomial distribution. Statistics in Medicine, 38(28):5413–5427, December 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sera:2019:EME

[3632] Francesco Sera, Benedict Armstrong, Marta Blangiardo, and Antonio Gasparrini. An extended mixed-effects framework for meta-analysis. Statistics in Medicine, 38(29):5429–5444, December 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2019:TSE

[3633] Yong Lin, Weichung J. Shih, and Shou-En Lu. Two-stage enrichment clinical trial design with adjustment for misclassification in predictive biomarkers. Statistics in Medicine, 38(29):5445–5469, December 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2019:MSE

[3634] Wen Li, Jing Zhao, Xiaoyun Li, Cong Chen, and Robert A. Beckman. Multi-stage enrichment and basket trial designs with population selection. Statistics in Medicine, 38(29):5470–5485, December 20, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2019:DAB

[3635] Dake Yang, Elena Deych, Berkley Shands, Meghan C. Campbell, Joel S. Perlmutter, Steve Petersen, Bradley L. Schlaggar, and William Shannon. Detecting associations between intact connectomes and clinical covariates using recursive partitioning object-oriented data analysis. Statistics in Medicine, 38(29):5486–5496, December 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2019:CTT

[3636] Jingyuan Liu, Jason C. Legg, May Mo, and Xuwen Zhang. Considerations in testing treatment effects on transient event driven health status changes measured by patient reported outcomes. Statistics in Medicine, 38(29):5497–5511, December 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edwards:2019:NEC

[3637] Jessie K. Edwards, Giorgos Bakoyannis, Constantin T. Yiannoutsos, Margaret W. Mburu, and Stephen R. Cole. Nonparametric estimation of the cumulative incidence function under outcome misclassification using external validation data. Statistics in Medicine, 38(29):5512–5527, December 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2019:ADC

[3638] Minjung Lee, Eric J. Feuer, Zhuoqiao Wang, Hyunsoon Cho, Zhaohui Zou, Benjamin F. Hankey, Angela B. Mariotto, and Jason P. Fine. Analyzing discrete competing risks data with partially overlapping or independent data sources and nonstandard sampling schemes, with application to cancer registries. Statistics in Medicine, 38(29):5528–5546, December 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freeman:2019:IIN

[3639] Suzanne C. Freeman, David Fisher, Ian R. White, Anne Auperin, and James R. Carpenter. Identifying inconsistency in network meta-analysis: Is the net heat plot a reliable method? Statistics in Medicine, 38(29): 5547–5564, December 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [3653] and reply [3654].

Wu:2019:NBJ

[3640] Jing Wu, Ming-Hui Chen, Elizabeth D. Schifano, Joseph G. Ibrahim, and Jeffrey D. Fisher. A new Bayesian joint model for longitudinal count data with many zeros, intermittent missingness, and dropout with applications to HIV prevention trials. Statistics in Medicine, 38(30):5565–5586, December 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scheffler:2019:CAR

[3641] Aaron W. Scheffler, Donatello Telesca, Catherine A. Sugar, Shafali Jeste, Abigail Dickinson, Charlotte DiStefano, and Damla Sentürk. Covariateadjusted region-referenced generalized functional linear model for EEG data. Statistics in Medicine, 38(30):5587–5602, December 30, 2019. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Francq:2019:CPT

[3642] Bernard G. Francq, Dan Lin, and Walter Hoyer. Confidence, prediction, and tolerance in linear mixed models. Statistics in Medicine, 38(30):

5603–5622, December 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Papageorgiou:2019:IDP

[3643] Grigorios Papageorgiou, Mostafa M. Mokhles, Johanna J. M. Takkenberg, and Dimitris Rizopoulos. Individualized dynamic prediction of survival with the presence of intermediate events. Statistics in Medicine, 38(30):5623–5640, December 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Verbeeck:2019:GPC

[3644] J. Verbeeck, E. Spitzer, T. de Vries, G. A. van Es, W. N. Anderson, N. M. Van Mieghem, M. B. Leon, G. Molenberghs, and J. Tijssen. Generalized pairwise comparison methods to analyze (non)prioritized composite endpoints. Statistics in Medicine, 38(30):5641–5656, December 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xia:2019:NGS

[3645] Meng Xia, Susan Murray, and Nabihah Tayob. Nonparametric group sequential methods for recurrent and terminal events from multiple followup windows. Statistics in Medicine, 38(30):5657–5669, December 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jennison:2019:AR

[3646] Christopher Jennison and Bruce W. Turnbull. Authors’ reply. Statistics in Medicine, 38(30):5670–5671, December 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2019:CMS

[3647] Richard D. Riley. Correction to: Minimum sample size for developing a multivariable prediction model: Part II-binary and time-to-event outcomes by Riley RD, Snell KI, Ensor J, et al. Statistics in Medicine, 38(30):5672, December 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andres:2020:CTT

[3648] A. Martín Andrés, I. Herranz Tejedor, and F. Gayá Moreno. Comments on “Two-tailed significance tests for 2 × 2 contingency tables: What is the alternative?” by Robin J. Prescott, Statistics in Medicine 2019; 38:4264–4269. Statistics in Medicine, 39(4):510–513, February 20, 2020.

CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3555] and response [3650].

Walter:2020:CWS

[3649] Stephen D. Walter. Correction to Walter SD, Turner RM, Macaskill P. Optimising the two-stage randomised trial design when some participants are indifferent in their treatment preferences (2019). Statistics in Medicine 38, 2317–2331. Statistics in Medicine, 39(10):1591, May 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3421].

Prescott:2020:RCT

[3650] Robin J. Prescott. Response to comments on “Two-tailed significance tests for 2 × 2 contingency tables: What is the alternative?”. Statistics in Medicine, 39(1):99–101, January 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3555, 3648].

Zhan:2020:COU

[3651] Zhuozhao Zhan and Edwin R. van den Heuvel. Correction for “Optimal unidirectional switch designs” by Zhuozhao Zhan, Geertruida H. de Bock, and Edwin R. van den Heuvel. Statistics in Medicine, 39(19):2587, August 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3177].

Conner:2020:CAR

[3652] Sarah C. Conner and Ludovic Trinquart. Correction: Adjusted restricted mean survival times in observational studies. Statistics in Medicine, 39 (27):4105, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3523].

Krahn:2021:CII

[3653] Ulrike Krahn, Harald Binder, Gerta Rücker, and Jochem König. Comments on “Identifying inconsistency in network meta-analysis: Is the net heat plot a reliable method?”. Statistics in Medicine, 40(18):4161–4163, August 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3639, 3654].

Freeman:2021:ARC

[3654] Suzanne C. Freeman, David Fisher, Ian R. White, and James R. Carpenter. Authors’ reply to “Comments on Identifying inconsistency in network meta-analysis: Is the net heat plot a reliable method?”. Statistics in Medicine, 40(18):4164–4165, August 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3639, 3653].

Jones:2021:CQH

[3655] Hayley E. Jones. Correction: Quantifying how diagnostic test accuracy depends on threshold in a meta-analysis. Statistics in Medicine, 40(18): 4166, August 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3591].

Kim:2022:COE

[3656] Byungwon Kim, Seonghong Kim, Sungkyu Jung, Woncheol Jang, and Johan Lim. Comments on “Online estimation of the case fatality rate using a run-off triangle data approach: an application to the Korean MERS outbreak in 2015” by Sungim Lee and Johan Lim published in Statistics in Medicine (vol. 38, 2644–2679, 2019). Statistics in Medicine, 41(9):1728–1732, April 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3442].

Anonymous:2023:CCP

[3657] Anonymous. Correction to “Controlled pattern imputation for sensitivity analysis of longitudinal binary and ordinal outcomes with nonignorable dropout”. Statistics in Medicine, 42(12):2027–2028, May 30, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3040].

Anonymous:2023:CCE

[3658] Anonymous. Correction to “Causal estimands and confidence intervals associated with Wilcoxon–Mann–Whitney tests in randomized experiments”. Statistics in Medicine, 42(18):3316, August 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3136].